A Bibliography of Publications in Frontiers in Marine Science (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

10 April 2024 Version 1.10

Title word cross-reference

110m [2538]. ^{13} [866, 236]. ^{134} [2538]. ^{137} [2538]. ^{15} [1991, 591, 236, 2937]. ^{18} [351]. ^{2+} [1256, 1751]. ^{210} [2483, 2435, 2527, 2550, 2685, 2340, 2559, 2530, 2998, 2538, 2449, 2523]. ^{226} [2538]. ^{238} [2538]. ^{40} [2538]. ^{7} [2559]. ^{90} [2538]. ^{vp+} [392]. ^{vp−} [392]. _{2} [2431, 1737, 53, 2641, 1771, 2916, 1596, 814, 2378, 2983, 2456, 1828, 2700, 2288, 1894, 1111, 2617, 736, 1847, 425, 491, 2855, 86, 613, 409, 447, 781, 1725, 1917]. 2′ [2700]. _{xs} [2559]. α [222, 703]. β [2920, 2361, 2579, 300]. Δ [2700, 1991, 351, 591, 866, 236, 2937]. γ [2093, 703]. κ [330]. n [1297]. ′ [2920]. × [2595, 3146, 2124, 2870, 2388].

*Omics [2434].

-3 [1297, 1165]. -Associated [437]. -Butenolides [703]. -carotene-9 [2920]. -Catenin [2361]. -Conglycinin [2579, 300]. -Furanone [952]. -Hydroxy [703]. -Irradiated [2093]. -oxygenase [2920].

/Calmodulin [1256]. /Calmodulin-NOS/NO-TNFs [1256]. /N [2700].

0-Group [440].

1 [651, 131, 1381]. 1- [2948]. 1-D [651]. 1-M [131]. 100-Year [1397]. 1119/27 [1940]. 12^{o}N [723]. 14 [2704, 1711]. 150-m [1620]. 16S [1924, 2607, 2843]. 18S [2607, 2880, 2996]. 19 [1807, 2402, 2231, 2633, 1911, 793, 2240, 1789, 2229, 2242, 2697, 1730, 2426, 2152, 2834, 2367, 2030, 2364, 1983, 1912, 2731, 2156, 2701, 2851]. 1980s [345]. 1a [3183].

2 [522, 494, 952, 2153, 2508, 2156, 1404]. 2- [2948]. 2-Arachidonoylglycerol [2710]. 2-Ethylhexyl [2008]. 2.0 [2, 1330]. 2000 [2165]. 2000s [1029]. 2010/11 [1869]. 2010s [345, 1029]. 2017 [2724]. 2018/2019 [2316]. 2019 [1792, 2755]. 2020 [997, 1861]. 2020/21 [1869]. 2030 [270]. 20th [2059]. 21st [644, 1071, 2356, 2624, 2582, 217]. 25-Hydroxylase [65]. 26 [1626].

3 [2873, 2591]. 3-D [2873, 2591]. 3-Year [1887]. 30-Year [3101, 145]. 35-Year [1616]. 360-Degree [2593]. 3D [1417, 2948, 1795, 1876, 2041, 1602, 1090, 1216, 1888, 1504, 2914, 1797, 89, 1170, 81, 1732, 1814, 1870]. 3D-Structural [1417]. 3Rs [1096].

48^{o}S [1500]. 4DEMON [2313]. 4DVAR [1908].

5H [952].

65-Day [2731].

70^{o}S [723].

8Rs [280].

9^{o}N [97].

= [3029, 3188, 2331].

Abalone [1943, 2655, 2754, 2084]. Abandoned [3040]. ABCB1 [222]. ABCC [222]. Ability [3046, 1712, 475, 2016]. Abiotic [1706, 1420, 2121, 2496, 2185]. Abnormal [440, 1688, 2854, 1590]. Aboard [2370]. Aboriginal [2137]. Above [3064]. Absence [3049, 2989]. Absolute [1900, 3038]. Absorption [2643, 596]. Abu [643]. Abundance [907, 2179, 2736, 381, 1811, 3100, 2682, 1024, 610, 589, 531, 828, 1636, 2724, 2248, 919, 3089, 2469, 2836, 1828, 24, 1025, 3052, 1533, 2779, 2032, 1391, 921, 356, 2241, 1273, 2101, 2145, 1546, 2234, 140, 1178, 368, 167, 582, 74, 2860, 1577, 1696, 2438, 2551]. Abundances [920, 2175, 1860]. Abundant [1259, 2848, 1241]. Abyssal [2350, 2377, 2578, 1746, 2063, 784, 2183, 2200, 1075, 1747, 1553]. AC [445]. AC-S [445]. Academy [2790]. Acanthopagrus [1744, 3141]. Acartia [1940, 2564, 2781]. Accelerate [2045, 350]. Acceleration [2542, 1799]. Accelerometry [1799, 1637]. Acceptability [2626]. Access [367, 902, 2553]. Accessible [2552]. Accidents [159]. Acclimation [2800, 1149, 1963, 2406, 2038, 1947]. Acclimatization [1086, 2568]. Accommodation [3036]. Accompanied [1023]. Account [2739, 1166]. Accountability [2612]. Accounting [1374, 2735]. Accumulate [1516]. Accumulating [512]. Accumulation [329, 1761, 2074, 1039, 1491, 1547, 2082, 1631, 1159, 1720, 3086, 2894, 3176, 2811, 1468]. Accuracy [2010, 406, 2005, 1552, 701, 2503]. Accurate [203, 1093, 627, 1230]. Accurately [2421]. Acetylcholine [112]. Acetylene [513]. AChE [3046]. Achieve [2618]. Achievements [1055, 1774]. Achieving [109, 60, 210]. Acid [404, 1991, 76, 1215, 2639, 580, 744, 2342, 837, 512, 633, 1159, 2297, 1409, 941, 2380, 3027, 2671, 695, 2669, 2658, 3184, 184, 65]. Acid-Specific [404]. Acidic [1876, 425, 21]. Acidification [1233, 682, 2203, 1446, 1932, 1845, 230, 107, 3154, 967, 1213, 59, 1405, 1415, 2378, 3012, 2110, 1755, 2456, 650, 1058, 882, 1520, 2843, 974, 2443, 1893, 1398, 1442, 2296, 2615, 2833, 354, 1306, 1431, 2565, 2712, 707, 3211, 2788, 1218, 1495, 788, 806, 1825, 1247, 898, 3204, 1864]. Acidified [1086, 1454]. Acidifying [2965]. Acids [276, 939, 591, 1823, 1449, 2204, 1297, 2945, 2223, 943, 167, 695]. Acorn [1128]. acoroides [427]. Acoustic [1470, 2605, 203, 369, 2094, 3181, 654, 2821, 1695, 3070, 3000, 1116, 641, 3145, 2604, 921, 2081, 2555, 747, 2252, 2308, 2742, 2947, 1540, 1604, 1153, 2741, 1698, 3033, 1623, 2553, 1034, 459]. Acoustically [2173]. Acoustics [690]. Acquaintances [2273]. Acquisition [728, 2290, 2871, 687]. Acropora [2220, 2133, 2176, 1888, 1608, 2130, 195, 938, 2671, 1384, 2168, 2470]. Acroporid [2392]. Across [1268, 2698, 585, 230, 337, 2578, 1690, 807, 18, 453, 1094, 2272, 2462, 1692, 3171, 1763, 1746, 1568, 2342, 13, 769, 2333, 106, 1635, 478, 558, 1530, 2735, 1216, 2264, 2241, 2405, 3034, 2101, 1317, 2011, 450, 1717, 2892, 2296, 838, 595, 117, 103, 1854, 2924, 3194, 2824, 2664, 2798, 1495, 1641, 1098]. Act [2544, 893]. Actin [1713]. Acting [529]. Actinobacteria [2670]. Action [522, 1246, 193, 1570, 1722, 3032, 1121]. Actions [260, 166, 68, 1857]. Activated [295]. Activating [388]. Activation [2840, 330]. Activator [2361]. Active [2173, 231, 1769, 2486, 2488, 97, 812, 1971, 810, 90, 1976, 2128, 1588, 1567, 410, 1358]. Activities [3047, 2091, 2247, 1817, 1392, 2631, 2665, 899, 1321, 2364, 1144, 2007, 220, 80, 2138, 3068, 148]. Activity [1807, 987, 2325, 1250, 522, 258, 1083, 3099, 1863, 176, 2161, 868, 1117, 664, 2904, 3146, 2366, 2555, 1280, 1799, 2090, 1165, 2936, 2538, 955, 1751]. Acts [2033]. Actual [1144]. acuta [2644]. Acute [3122, 3045, 3061, 2761, 2038, 1056, 1384, 955, 2357, 2826, 2861]. Acutely [2627]. acutidens [1540]. acutorostrata [2034]. Acyl [220]. AD [1115]. Adam [1590]. Adaptability [2798]. Adaptation [656, 2784, 1638, 1033, 891, 2430, 2869, 2569, 1152, 2568, 2548, 2632, 200, 1949, 2694, 1674, 1450, 2826]. Adapted [1338]. Adapting [508, 957, 2058]. Adaptive [2203, 429, 1086, 2896, 2871, 583, 68, 1241]. ADCP [873]. Added [1787, 1444, 2745]. Added-Value [2745]. Adding [2390]. Addition [900, 521, 1989]. Additional [2038, 2925]. Additives [2481]. ADDOMEx [1827]. Address [1441, 1182, 1949]. Addressed [1350]. Addressing [888, 2534, 510, 1722, 401]. Aden [996]. Adhesive [2853, 766]. Adjacent [1372, 2986, 130, 1904, 106, 2538, 2769, 2287, 1989, 609, 1134, 1342]. Adjust [1752]. Adjustment [2448]. Administration [1760]. Admixture [2725]. ADR [1180]. Adriatic [1102, 1773, 1127, 1580, 2025, 1240, 1482, 2783, 1367, 819, 961, 1004, 2213, 2364, 187, 781]. adriatica [3037]. Adsorption [2208, 3102]. Adult [792, 2110, 1086, 241, 1286, 2945, 2681, 1384, 2565, 3160, 2540]. Adults [1017, 631]. aduncus [1464, 1550, 876]. ADV [2817]. ADV-Based [2817]. Advance [1236, 2741]. Advanced [2290]. Advancement [380, 2444]. Advances [250, 1795, 3125, 451, 1536, 679, 1288, 3059, 820, 1501]. Advancing [2640, 979, 1359, 144, 960, 2705, 630, 2533, 1874]. Advantage [2490, 578, 2267, 712]. Advantages [513]. Advection [2746, 593, 2117, 2766]. Adverse [408, 1213]. Advice [1445, 1279, 1913, 2524, 2600]. Aegean [2507]. Aerial [2948, 1256, 1804, 1852, 43]. Aeromonas [2899, 2982, 3157]. Aerosol [3060, 1209, 2713, 1912]. Aerosols [971, 3171]. Aetobatus [3092]. Affairs [2075]. Affect [887, 1337, 787, 381, 951, 2537, 1563, 2380, 2441, 1473, 1985]. Affected [2666, 1061, 666, 50, 1502, 2364, 1258, 2953, 2830]. Affecting [2688, 2513, 1662, 2536, 399]. Affects [2595, 1285, 1851, 808, 323, 1162, 2084, 2470, 263]. Affinity [602]. Afforestation [3057]. Africa [2551, 3066, 1027, 579, 1949, 876, 1829, 2989, 927]. African [1447, 293, 688]. After [2802, 2641, 591, 1481, 48, 1681, 121, 2697, 97, 1188, 1887, 159, 1406, 2668, 817, 3130, 1797, 3128, 3169, 95, 1952, 409, 955, 3088, 1805, 2388]. AG [880, 2538]. Against [2802, 2782, 2763, 162, 565, 846, 475, 448, 3, 1069, 2731]. agalactiae [2214]. Agarophyton [468]. Age [2334, 1102, 76, 1127, 1126, 1557, 1528, 1122, 290, 2653, 1295, 2559, 417]. Aged [3102]. Ageing [581]. Agelas [1251]. Agenda [141]. Agent [3061]. Agents [2695, 219, 871]. agglomerans [2214]. Aggregate [632, 1843, 878]. Aggregated [1109, 2518]. Aggregates [1675, 578, 2405]. Aggregating [543, 259]. Aggregation [2387, 1801, 1827, 680, 17, 1010]. Aggregations [2202, 1329, 45, 1274, 849]. Aggressive [181]. Aggressiveness [3099, 1286]. Agreement [3197, 2570, 2649]. Agricultural [1918]. Agriculture [418, 2352]. Agulhas [1224, 457]. Ahead [2546]. AHPND [392]. aibuhitensis [2486, 2033]. Aid [1807, 2112, 3070]. Aids [1195]. Aiming [336]. Aiptasia [1023]. Air [1311, 2916, 814, 2092, 2426, 2617, 1997, 1967, 1317, 2474, 1656]. Air-Breathing [2426]. Airborne [2491]. Aircraft [540]. Airgun [2006, 89]. AIS [1754, 1083, 212]. Aitutaki [248]. Akashiwo [2889]. al [997, 1261, 2266]. Alacranes [1106]. Alanine [695]. Alaska [805, 1975, 41, 1689, 2062, 1628, 1241, 1572, 1369]. Albacore [2805]. Albinism [825]. Alboran [2277, 456]. Alcyonacea [1227]. Alectryonella [2878]. Aleutian [1079]. Alexandrium [1940, 1281, 1137, 2782, 1505, 2961, 2835]. alfredi [660]. Alga [1110, 188, 3123, 2918, 2782, 135, 816, 2135, 2923, 2541, 496, 1428, 468]. Algae [1347, 1769, 2753, 2111, 2912, 2971, 1818, 1802, 3097, 3195, 2298, 2038, 2729, 454, 1041, 1709]. Algal [2787, 1700, 3012, 46, 209, 2842, 1105, 68, 138, 165, 2824, 98, 439, 715, 777]. Algicidal [2782]. Algicolous [2971]. Alginate [522]. alginolyticus [2214]. Algorithm [1838, 1232, 666, 1792, 1105]. Algorithms [701]. algosus [450]. Alicella [2896]. Alien [1021, 2256]. Aliger [748]. Alive [1343]. Alkalinity [321, 3021, 1003]. Allatostatin [2866]. Allee [1301]. Alleviated [1115]. Alleviation [3158]. Allocation [991, 2987, 2907]. Allocation-Mediated [2907]. Allocations [1416]. Allochthonous [784, 1503]. Allowing [1615]. Alone [2158, 1829]. Along [1892, 786, 656, 1142, 2269, 1433, 1395, 66, 1281, 473, 389, 370, 316, 611, 1047, 1512, 3063, 1309, 1394, 143, 2385, 1448, 1828, 2656, 2887, 1666, 717, 1535, 1088, 50, 3098, 2995, 1899, 1982, 1429, 2910, 2030, 2373, 2519, 2197, 1477, 2200, 477, 694, 671, 1756, 2416, 301, 788, 618, 2881, 2551, 1, 790]. Along-Path [477]. Alongside [402]. Alpha [191]. Alpheid [2267]. alrasheidi [1242]. Alter [1079]. Alteration [851, 1703]. Alterations [1517, 293, 3120, 2579]. Alternative [801, 1156, 140, 3141, 1180]. Alternatives [1096, 985]. Alters [163, 2328]. Altimeter [3073]. Altimetric [2294, 87]. Altimetry [1891, 2453, 1900, 906]. Aluminum [570]. Alveolata [1853, 1032]. Alvinellid [2166]. Amazon [1, 790, 292, 3098]. Amazonian [1026, 2198]. Ambassadors [2695]. Ambient [2983, 520, 2552, 1389, 2929]. amblycephala [388]. Amebocyte [1278, 869]. Ameliorates [806]. Amenable [475]. America [2174, 1592, 3153, 138]. American [2299, 881, 3059, 2526, 905, 2823]. americanus [2299]. Amery [1791]. Amid [549]. Amino [276, 404, 1991, 591, 2639, 2342, 2945, 2223, 943, 695]. Ammonia [2043, 520, 2366, 2676, 3187, 2124, 3193, 279]. Ammonia-Oxidizers [2366]. Ammonium [1280, 1399, 955, 1239]. AMOC [1830]. Amoebophrya [2889]. Among [2028, 2273, 1425, 2684, 1595, 196, 1359, 317, 2668, 1859, 1682, 1629, 2136, 220]. Amongst [1762]. amoyensis [3174]. Ampelisca [1841]. Amphibalanus [1128]. Amphidinium [2415]. Amphileptus [1837]. Amphioctopus [2802]. Amphipod [472, 3115, 2896]. Amphipoda [1841, 1818]. Amphipods [2578, 2895, 2949, 1154]. amphitrite [1128]. Amphora [293]. Amplicon [83]. Amplifies [1431]. ampullatus [1584]. Amundsen [62, 1442]. amyloliquefaciens [1204]. Anadromous [656, 1044]. Anaerobe [3121]. Anaerobic [3192, 635]. Analogs [1863]. Analogue [922]. Analyser [2492]. Analyses [275, 18, 83, 1655, 71, 2120, 1374, 2473, 1455, 358, 2749, 3111, 1382, 3166]. Analysis [2334, 500, 2499, 1526, 2596, 1093, 643, 2428, 324, 276, 3025, 588, 230, 1713, 2944, 1349, 1992, 2397, 2782, 591, 2691, 945, 531, 1998, 1213, 654, 541, 1759, 1507, 1624, 1405, 3106, 2920, 2375, 2853, 697, 1568, 212, 2256, 760, 75, 385, 181, 1373, 1659, 690, 2102, 3037, 1525, 3015, 825, 803, 321, 520, 1919, 3175, 2127, 1658, 3172, 2088, 2880, 145, 1145, 2725, 375, 2223, 1274, 1139, 2037, 3023, 2552, 2545, 1113, 3112, 1151, 1744, 384, 26, 3021, 1399, 938, 160, 3055, 2833, 259, 561, 1091, 1611, 1960, 197, 2759, 3184, 3160, 3020, 2875, 517, 2217, 2434, 405]. Analysis [2472, 1886, 1207, 136, 2758, 3026, 3158, 2489, 3038, 2372, 2996, 2762, 1941, 1482, 2258, 1009]. Analytical [1539, 902, 820]. Analyze [212, 738, 1744, 3129]. Analyzed [2503]. Analyzer [1003]. Analyzing [1758, 1460]. Anchovy [706]. Ancillary [2783]. Andaman [1620, 33]. Andres [1220, 3074, 936]. Androgenic [2114, 2271, 3127]. Androgens [2854]. Anemone [1225, 1650, 692, 98]. Anemones [2267, 692]. Angler [757]. anglica [163]. Anguilla [1362, 47, 905]. anguillarum [2802]. angulata [2907]. angulifera [1820]. AniBOS [2959]. Animal [531, 2347, 97, 2959, 1191, 1417, 2574, 136]. Animals [789, 1084, 519, 2426, 394]. Anisakids [1768]. Annelid [1198]. Annelida [1987, 545, 2166, 1842, 2343, 5, 2424]. Annotated [1676]. Annotating [2, 1330]. Annotation [2421, 1542, 3022]. Annual [2812, 3060, 1311, 554, 290, 2828, 973, 1278, 191, 417, 1192]. Anomalies [11, 3004]. Anomalous [1730, 593, 2117]. Anomaly [895, 2977, 195]. Anoplopoma [1241]. Anoxia [780, 426]. Antagonistic [2214]. Antarctic [1526, 3108, 1198, 1047, 1359, 1679, 1452, 37, 2794, 2369, 53, 1425, 1189, 2815, 2108, 3045, 1824, 1913, 1391, 1398, 603, 1452, 2615, 576, 518, 2191, 3149, 2177, 2625, 1034]. Antarctica [1030, 1246, 2714, 2527, 441, 1116, 2755, 582, 2547, 1486, 53, 62, 1806]. antarctica-Dominated [62]. antarcticus [3045]. Anthozoa [542]. Anthropause [2633, 1911]. Anthropized [1779]. Anthropocene [885, 466, 2174, 442, 2284]. Anthropogenic [1463, 682, 381, 2689, 3003, 1622, 957, 2403, 897, 1271, 2011, 732, 2118, 2441, 1473, 332, 1074, 1461, 2860]. Anthropogenically [514]. Anthropogenized [2932]. Anthropozoonotic [881]. Anti [1319, 415, 2634, 2707]. Anti-cancer [1319]. Anti-Diabetic [2707]. Anti-fouling [2634]. Anti-Hypertensive [2707]. Anti-phytopathogenic [415]. Antibacterial [1751]. Antibiotic [1972]. Antibiotics [515]. Anticipation [2294]. Anticyclones [1291, 1543]. Anticyclonic [1872, 1187, 595]. Antidepressant [1255]. Antifouling [219]. Antiglioma [1863]. Antillean [1885]. Antimicrobial [2310, 1115, 2417, 871]. Antioxidant [1250, 2595, 3205, 3157, 604, 3146, 2761, 3130, 2676, 3027, 2304]. Antipatharia [542, 2992]. Antipatharians [413]. Antiporter [1725]. Antiproliferative [2374, 709, 880]. Antiviral [709]. Aotearoa [2874]. APEI [2050]. Apex [2102]. APGWamide [2084]. Apherusa [3115]. Apical [1039]. Aplacophora [2158]. Apoptosis [2033, 880, 2124]. Apostichopus [2726, 2349]. Apparatuses [744]. Apparition [1797]. Appeal [1913]. Application [2820, 1071, 2435, 1846, 2641, 189, 2346, 1213, 2593, 710, 2700, 1305, 1426, 2120, 47, 1413, 1219, 2236, 2352, 2065, 1077, 1170, 2526, 2907, 2216, 2392, 3068, 1468]. Applications [3119, 71, 1594, 69, 2000, 1683, 1627, 1455, 1767, 2073, 1816, 407]. Applying [771, 2012, 1695, 3015, 1623]. Appraisal [2621]. Appraising [1889]. Approach [1892, 1472, 153, 1367, 2930, 644, 2991, 2061, 1350, 1539, 2133, 489, 45, 3147, 757, 903, 2699, 486, 2346, 1877, 2502, 2619, 722, 1929, 1441, 352, 1721, 257, 697, 3134, 1158, 837, 1659, 640, 690, 2599, 3116, 2315, 237, 925, 2957, 2119, 335, 2487, 1419, 1090, 1334, 566, 2776, 1494, 750, 908, 913, 1059, 467, 817, 2146, 2373, 559, 679, 841, 612, 1593, 506, 2672, 3128, 2447, 2987, 1172, 1227, 1623, 2529, 2822, 2446, 1298, 1884, 3068]. Approaches [643, 324, 1649, 2243, 2972, 2797, 2048, 2575, 557, 931, 579, 1441, 3016, 2047, 738, 2062, 1705, 140, 1514, 3020, 2434]. Approvals [3206]. April [2724]. Aqaba [2150]. Aqua [1179]. Aqua-Cultured [1179]. Aquaculture [2820, 1964, 821, 157, 842, 320, 2113, 156, 1532, 258, 1700, 1769, 2933, 1518, 2111, 502, 1353, 957, 2638, 71, 1294, 981, 1631, 2642, 1727, 2153, 747, 2943, 67, 2781, 2450, 38, 187, 1752, 1170, 165, 126, 2452, 319, 869, 985, 255, 1475, 2622, 1970, 164]. Aquaponics [3110]. Aquarium [368]. Aquatic [243, 789, 1633, 1084, 3084, 1695, 1393, 221, 981, 1015, 2093, 982, 1339, 2769, 2389, 136]. Aquiferous [216]. Aquifers [847]. Arab [643, 2670, 2171]. Arabia [2266]. Arabian [2479, 2499, 2324, 70, 426, 3150, 2623, 675, 909, 2373]. Aracati [1479]. Arachidonic [512, 65]. Arachidonic-Acid-Accumulating [512]. Arachidonoylglycerol [2710]. Arbacia [2859]. arborea [315]. Arc [397]. Archaea [2227]. Archaeal [54, 1747]. archetypal [438]. Archipelago [1647, 812, 1926, 3118, 665, 1208, 2383, 2109, 1108, 1117, 2081, 2556, 3074, 980, 1079, 2798, 1440, 1436]. Architecture [1873, 2686]. Archive [456]. Archivists [1933]. Arctic [2483, 2700, 2109, 304, 904, 2582, 86, 1931, 1484, 234, 346, 2044, 1123, 1282, 1986, 1527, 2435, 887, 1420, 708, 2520, 373, 234, 2099, 1813, 2575, 3062, 611, 639, 498, 31, 1388, 1779, 3115, 554, 545, 1830, 2884, 1755, 868, 345, 2895, 1483, 1625, 528, 1196, 1331, 2531, 1078, 423, 285, 874, 1411, 1442, 452, 597, 593, 2117, 1085, 592, 252, 1389, 530, 132, 55, 686, 306, 3083, 311, 2594, 2279, 326, 2525, 878, 64]. Arctica [2371, 2937]. Area [3108, 2202, 2510, 2041, 1779, 2375, 560, 2656, 2887, 349, 2756, 478, 934, 3067, 2648, 1403, 1273, 915, 1408, 559, 1509, 1691, 3194, 781, 1556, 1641, 1989, 3176, 2811, 146, 2224, 2182, 1927, 420, 436, 1290, 283, 1274, 2632, 894, 3128, 3113, 1607, 2147, 1743, 2442]. Area-Based [1509]. Areas [827, 224, 3042, 2641, 367, 1492, 1669, 1267, 1246, 1570, 1997, 1487, 78, 2398, 374, 2064, 1683, 2702, 2123, 15, 775, 2199, 1839, 3066, 601, 1741, 2327, 12, 2454, 2259, 1447, 2631, 1521, 2241, 100, 2536, 2906, 2144]. Arenas [2353]. argenteus [3077]. Argentina [1485, 801, 391, 2018]. Argentine [2253]. argentinus [2253]. Arginine [712, 3158]. Argo [1057, 1891, 223, 1512, 2262, 1758, 1055, 1139, 802]. argoFloats [1758]. Argopecten [1651, 2886]. argus [3160]. Argyrosomus [1068, 3175]. Arhynchobatidae [393]. ariakensis [1152]. Arise [2064]. Ark [2861]. ARMS [550, 1018]. ARMS-MBON [1018]. Aromatic [366, 158, 1925]. Array [2305, 802, 223]. Arrays [1139]. Arrest [388]. Arrival [753]. Arsenic [293]. Arsenic-Induced [293]. Art [308, 2262, 835, 442, 2335]. artemiae [2792]. Arthrinium [2971]. Arthropod [2534]. Artifact [2253]. Artificial [2882, 1363, 2048, 922, 1574, 480, 3081, 1257, 3016, 431, 2554, 384, 248, 822, 549, 1935]. Artisanal [557, 2137, 2021, 399, 2336]. Ascension [1743]. Ascidian [303, 2043, 3044, 1995]. Asconema [1504, 1506]. Ascothoracida [1537]. Ashmore [1516, 2017]. Asia [278, 3167, 1780]. asiaeorientalis [1370]. Asian [619, 2330, 2945, 1515, 3092]. asiaticus [2514]. Asking [1649]. Aspartic [76]. Aspects [1518, 1731, 27]. aspera [2232]. Aspergillus [415, 880]. Assay [513]. Assays [735, 1411, 290]. Assemblage [2039, 2249, 139, 1596, 2457, 2813, 283, 2635, 119, 1679, 103, 3076, 3207]. Assemblages [2812, 324, 18, 2730, 457, 1639, 3017, 1000, 2358, 2072, 24, 3009, 1809, 2443, 1289, 137, 1413, 2827, 2627, 1386, 681, 1796, 677, 1959, 425, 3094, 3118, 551, 1477, 2992, 1743, 254, 88, 148]. Assembling [326]. Assembly [146, 1936, 2355, 1366, 1567, 2810, 2849, 3149, 1542]. Assess [963, 1808, 1032, 1071, 2991, 1795, 771, 757, 48, 316, 2240, 2348, 934, 1796, 344, 2518, 841, 995, 1091, 2529, 840, 3048]. Assessed [1268, 998, 83, 2987]. Assessing [2511, 1964, 2224, 1410, 2428, 1845, 1560, 2235, 1496, 1712, 62, 370, 1622, 795, 1760, 1507, 1407, 2785, 2751, 3116, 1697, 3170, 2569, 236, 1139, 3151, 2819, 2236, 387, 1949, 894, 127, 532, 140, 168, 1011, 2412, 2275, 262, 3005, 612, 1298]. Assessment [2739, 1807, 1367, 23, 2203, 1539, 819, 1787, 1320, 336, 794, 2005, 3132, 206, 895, 2933, 2274, 1166, 1430, 1267, 2418, 277, 3081, 962, 3077, 1926, 3016, 1375, 2804, 2477, 1529, 1519, 1335, 2827, 912, 2661, 293, 356, 3023, 1701, 1775, 731, 1465, 1866, 2632, 1966, 915, 1432, 2834, 1504, 2030, 1408, 599, 3074, 1387, 2538, 2073, 1211, 1265, 2851, 114, 718, 197, 266, 2625, 3085, 844, 1303, 2696, 2757, 1941]. Assessments [406, 1351, 528, 1169, 416, 725, 78, 1544, 2740, 1774, 742, 2287, 831, 609]. Assignment [2744]. Assimilation [2428, 3140, 87, 2432, 1077, 1908, 955]. Assisted [2170, 2159]. Associated [163, 1433, 1456, 515, 2948, 1052, 188, 3124, 2955, 1817, 420, 3115, 513, 2390, 2378, 1766, 2457, 2498, 3114, 1244, 2072, 2895, 1272, 3192, 2928, 1818, 2288, 571, 1043, 209, 1119, 2550, 1274, 1159, 920, 437, 1409, 1871, 1796, 1959, 2059, 543, 1650, 3118, 1368, 2022, 195, 1165, 849, 249, 2276, 671, 2748, 311, 3129, 2953, 2935, 56, 1082, 2602, 2886, 2138]. Association [2467, 2863, 1392, 1419, 2115, 2886, 761]. Associations [2273, 986, 877, 1653, 2496, 2798]. Astaxanthin [1215, 3205]. Asteroids [1274]. astreoides [57, 2544, 1535]. Asymmetric [1616]. Asymmetry [2268]. Asymptotic [1038]. At-Risk [1541]. At-Sea [839, 103]. Atacama [2178]. Atlantic [424, 1994, 2060, 1956, 48, 1308, 949, 2378, 2469, 2513, 1830, 1862, 2165, 1195, 1126, 523, 966, 1485, 2947, 2560, 1548, 3118, 1080, 1356, 980, 1269, 1665, 2211, 91, 971, 1472, 1009, 2532, 329, 1262, 3060, 1486, 1209, 1311, 1360, 1612, 2482, 741, 337, 1161, 2039, 2728, 2510, 79, 1270, 2847, 2253, 2437, 1823, 1492, 2639, 1849, 3062, 1401, 1841, 1474, 1782, 2248, 2111, 1183, 128, 2019, 919, 2902, 2439, 277, 670, 1096, 1673, 1818, 3052, 106, 235, 3056, 460, 1169, 1557, 1196, 2713, 340, 2546, 2642, 1145, 375, 2631, 2735, 2036, 1314, 1729, 2425, 356]. Atlantic [3023, 2995, 208, 616, 1238, 1775, 9, 2105, 834, 2101, 2098, 2767, 2158, 1016, 1048, 2225, 341, 2906, 2934, 2227, 727, 2368, 1656, 753, 1177, 694, 3198, 1543, 1610, 113, 1738, 178, 199, 459, 1212, 1685, 2086, 378, 242, 935, 1277]. atlanticum [1462]. Atlantis [173]. Atlas [2751]. Atmosphere [2431, 16, 1210]. Atmosphere-Ocean [2431]. Atmospheric [1616, 1566, 1771, 1234, 1438, 1111, 899, 2998, 2594, 43, 1321]. Atmospheric-Driven [899, 1321]. Atoll [861, 956, 1197, 613, 1436]. Atolls [146]. ATP [1973]. atra [1503]. Atrina [2976]. Attached [1639]. Attachment [2710]. Attacks [2915]. Attempt [1083]. Attenuates [1977, 2927]. Attenuation [878]. Attitudes [746]. Atypical [2737, 1214]. audax [1886]. Aulacomya [1503]. Aulonocara [825]. Aurantiochytrium [1297, 3111]. aurata [408, 1988, 2024, 2090, 26, 2452]. Aurelia [2893, 2341]. Aureoumbra [2667, 3097]. Austral [2942, 37]. Australasia [1541]. Australia [672, 632, 1320, 389, 628, 278, 685, 1939, 1097, 555, 505, 1516, 624, 1666, 177, 1187, 2636, 200, 1059, 1883, 1191, 244, 953, 2376, 2468, 719, 2293, 2420]. Australian [1045, 192, 3143, 467, 2739, 947, 1166, 1001, 1671, 2154, 1869, 1546]. Australia’s [1698]. Authentication [2988]. Authorities [262]. Autochthonous [1081, 1503]. Autocompaction [2448]. Autocorrelation [71, 2545]. Autocrine [2939]. Autocrine/Paracrine [2939]. autoethanogenum [3065]. Automated [1470, 2605, 2421, 381, 2005, 1676, 1870]. Automatic [500, 1754, 2140, 2981]. Automating [531]. Automation [1896, 1511]. Autonomous [842, 550, 823, 499, 1003, 1457, 799, 2547]. Autophagy [465]. Autumn [2813, 534, 1317, 2040]. AUV [131]. Availability [868, 930, 2476, 1709]. Available [3109]. Aveiro [2513]. Aveiro-Figueira [2513]. Avenues [2464]. Average [2867]. Averse [1377]. Avicennia [1856, 1892, 2499, 2670, 2623]. Avilés [1274]. Avoid [1573, 1437]. Avoidance [1573, 27]. Awareness [2511, 258, 69, 1406, 1085]. Away [2441, 1473]. Axiidae [571]. Axiidean [571]. Axis [3127]. Azooxanthellate [50]. Azores [91, 1921, 1238, 2556, 980, 1440].

B [2082, 330]. B1 [1269]. BACI [3169]. Bacillus [2214, 2905, 1204, 2634]. Back [1617, 1916, 2613]. Backbarrier [2733]. Background [237, 1988, 1966]. Backgrounds [1762]. Backreef [2071]. Backscatter [873, 3181, 2821, 2390, 3000]. Bacteria [175, 188, 990, 1215, 1817, 2715, 1828, 209, 2497, 2227, 1627, 1011, 1283, 1620, 671, 372, 3141]. Bacterial [2377, 1462, 1924, 629, 815, 1639, 1225, 415, 2598, 952, 2978, 2016, 3094, 2090, 2940, 2022, 195, 1023, 2831, 368, 125, 2441, 1473, 1135, 703, 1747, 396, 2472, 1247, 761]. Bacterioplankton [478, 1162, 574]. Bacteriosyncytia [1788]. Bacteriovorous [475]. Bacterium [2782]. Bad [713, 2221, 2158]. baenschi [825]. Baffin [2099]. Bahamas [1290, 2317, 2229, 2168]. bahamense [3097]. Bahía [1387, 2318]. Bail [1481]. Bail-Out [1481]. Bait [852]. Baited [2005]. Baits [1156]. Baixas [2335]. Baja [1572]. Baker [1500]. Baker-Martínez [1500]. Balaena [2619]. Balaenoptera [438, 2034, 1607]. Balanced [173, 612, 1298]. Balancing [1474, 1035]. balanoides [3198]. Balanus [766]. Balearic [3031, 2261, 984, 1810]. Baleen [203, 2948, 2947]. Bali [1465]. Balikpapan [850]. Ball [776, 100]. Ballast [2427, 1794]. Baltic [2402, 2837, 2672, 1153, 2233, 2919, 807, 2001, 2110, 7, 1346, 2542, 2430, 2232, 1426, 2871, 1997, 912, 1900, 766, 1412, 2496, 1013, 665, 1427, 1283, 1019, 667, 1260, 3043]. Bamboo [3001]. Ban [28, 1487, 2731, 1524]. Banana [387]. Band [1792, 2963]. Bands [290]. Bangiales [32]. Bangladesh [3014, 2797, 1752, 2731]. Bank [1009, 748, 2773, 457]. Banks [632, 474]. Barcode [3098]. Barcodes [1742]. Barcoding [550, 2272, 962, 1336, 1907, 2493]. Bare [2913]. Barents [173, 1860, 2272, 828, 1518, 440, 833, 1396, 1299, 1802, 2809, 2922]. Baridi [3011]. Barnacle [2393, 1100, 698, 766, 1128, 3198, 2935]. Barnacles [2393, 1100, 655]. Barramundi [1920]. Barred [3104]. Barreirinhas [1543]. Barren [1875, 621]. barretti [1226, 1203]. Barrier [2286, 2991, 1115, 134, 1599, 3089, 2733, 2771, 997, 170, 98, 2969, 2620]. Barrier-Island [2991]. Barriers [157, 1749, 127, 870, 1752]. Base [1547]. Base-Extractable [1547]. Based [3014, 1892, 2802, 3028, 805, 2490, 2948, 607, 1560, 406, 3082, 1992, 2782, 2461, 2048, 2012, 757, 189, 1928, 1961, 895, 1213, 1759, 2753, 2390, 379, 931, 2795, 579, 1441, 2514, 492, 2539, 418, 94, 877, 2677, 1831, 2941, 12, 1583, 2804, 1279, 2877, 2537, 2638, 2129, 930, 734, 754, 1697, 1264, 2655, 1169, 1528, 335, 321, 1889, 2172, 2088, 2844, 2827, 2880, 2473, 2803, 2776, 2236, 1796, 2405, 2817, 913, 2398, 982, 2345, 1942, 1452, 140, 1515, 2878, 1455, 2058, 2744, 2998, 1509, 2276, 165, 2818, 2780, 2741, 2600, 1501, 1611, 3001, 197, 3184, 2875, 444, 1404, 1382]. Based [136, 2207, 1260, 609, 1134, 1303, 2696, 3038, 2752, 468, 1941, 935, 1890, 148, 115]. Baseline [771, 371, 1673, 2717, 2223, 1627, 1014, 2198]. Baselines [2883, 3199]. Basic [2820]. Basin [1492, 1517, 1775, 1572, 80, 675, 658, 256, 916, 1917]. Basin-Scale [1775]. Basin-Wide [1517]. Basins [588, 1906, 5, 2424, 326, 3171]. Basis [668, 2569]. Basket [901]. Basking [2095]. Bass [1539, 79, 299, 277, 3065, 977, 1067, 2669, 2239, 1199]. Batch [180]. Bath [3061, 3141]. Bather [876]. Bathybenthic [940]. Bathymetric [627, 1803, 1557]. Bathymetry [2872, 2390, 2453]. Bathypathes [581]. Bathypelagic [283, 836]. Bathytoshia [1822]. BATS [275]. Bay [1534, 1527, 1081, 1189, 2815, 628, 1160, 1061, 1939, 143, 2527, 2561, 3009, 1116, 850, 2779, 1157, 1651, 2827, 63, 2635, 2984, 2668, 2893, 195, 1983, 2468, 179, 1985, 1265, 2814, 197, 2442, 2675, 2951, 1825, 2860, 2886, 224, 2577, 1137, 2099, 2249, 189, 1070, 1817, 1948, 2242, 3170, 1744, 2560, 1548, 2580, 1726, 410, 1912, 2559, 245, 288, 2727, 1794, 978, 2769, 3155, 1732, 439, 2620]. Bayesian [2845, 2804, 3126, 734, 1169, 1909, 1593, 532, 609, 1303, 1941]. Bays [1395, 2855, 1881]. BBNJ [931, 2047]. BCC52274 [3111]. BCDO2 [2920]. Be [1350, 404, 3100, 107, 652, 108, 1742, 2546, 384, 1686, 1958, 2559, 2175, 2168, 2436]. Be/ [2559]. Be4 [1934]. Beach [316, 1022, 2344, 1158, 1826, 3015, 2819, 2817, 547, 2745, 2543, 1776, 1467, 2293, 435]. Beached [169]. Beaches [3031, 370, 2152, 1035, 2198]. Beagle [1750, 1592]. Beaked [2081, 1960]. Beaks [2334]. Bearing [31]. Beaufort [1976]. beauforti [851]. Because [108]. Become [436]. Becoming [3080]. Bed [1420, 1570, 2817, 239, 1989, 2475, 1217]. Bedload [838]. Bedrocks [1965]. Beds [1780, 2913]. been [1019]. Before [2402, 591, 3169, 95, 1012]. Before-After [3169]. Beginner [2434]. Begins [179]. Behavior [408, 1235, 472, 52, 370, 1208, 1444, 2034, 697, 1801, 2465, 181, 1099, 2102, 924, 2403, 3145, 2339, 1968, 1072, 2721, 433, 747, 526, 1380, 3162, 2921, 1361, 1607, 2368, 2040, 1079, 1206, 447, 1572, 1623, 459, 2597, 2470]. Behavioral [2235, 472, 36, 2006, 2460, 305, 2989, 27, 2741, 705]. Behaviors [1739, 1717, 2530, 1241]. Behaviour [1176, 1939, 1859]. Behavioural [1527, 3122, 2468, 2086]. Behind [1376]. Beibu [2804]. Being [1106, 1150]. Belgian [2313, 3137]. Beliefs [3191]. Belize [2633, 1911]. Below [44, 1483]. Belowground [3091]. Belt [1916, 374]. Benchmarks [2161]. Bend [436]. Beneath [1824, 2020, 626]. Beneficial [188]. Benefit [2903, 367, 1444, 1659, 2585, 2702, 2123]. Benefit-Sharing [367]. Benefits [2182, 2864, 2717, 1521, 2306, 2668, 3128, 2194, 2816, 1010]. Bengal [2249, 2561, 1912, 1985, 2814, 2620]. Benguela [2549, 2793, 357, 642, 165, 417]. Benthic [346, 1526, 1009, 2812, 2369, 1032, 1845, 577, 294, 887, 1420, 320, 1176, 2039, 373, 18, 2684, 62, 14, 1680, 31, 1779, 552, 1760, 1718, 3122, 1000, 919, 2789, 1458, 1746, 3114, 1824, 560, 837, 1851, 13, 284, 106, 1652, 528, 237, 2747, 1103, 784, 2348, 2827, 2995, 1899, 2536, 600, 483, 341, 988, 592, 1946, 506, 973, 2744, 1295, 2200, 937, 1154, 1985, 1524, 1201, 785, 2291, 3018, 1980, 2720, 372, 3088, 3149, 992, 118, 2814, 191, 2588, 2597, 1989, 3048, 1484, 2138, 2765]. Benthic-Pelagic [973, 2138]. Bentho [2060, 1375]. Bentho-Pelagic [2060, 1375]. Benthos [1189, 2815, 318, 326]. Benzo [222]. Berberine [2982, 3157]. Bering [1438, 1103, 227, 1613, 714]. Bermuda [537, 2544, 1371]. Beryllium [2403]. Best [643, 923, 312, 2964, 1578, 3013, 3005]. Bestiolina [3174]. Beta [3149]. Better [1106, 891, 1869, 1786, 2158, 758, 1133, 2157, 2608, 1422]. Between [1873, 827, 224, 188, 1372, 2845, 1750, 1332, 1255, 2614, 316, 420, 457, 1638, 1000, 1981, 710, 3063, 1766, 2837, 3081, 1523, 385, 2733, 886, 2768, 797, 798, 1965, 2082, 2839, 3153, 226, 340, 1392, 1150, 1274, 2923, 2554, 2302, 2149, 1897, 2497, 745, 2876, 2857, 2090, 598, 567, 2609, 1355, 858, 5, 2424, 1283, 29, 2445, 3127, 1784, 2740, 2769, 655, 2759, 2029, 2675, 3103, 783]. Beyond [3042, 1571, 367, 931, 2570, 144, 2702, 2123, 1944, 380, 774, 179]. BFT [3210]. Bias [1903]. Biased [1243, 2772, 35, 1339]. Bibliometric [1992, 145]. Biecheleriopsis [3037]. Bifunctional [522]. Bifurcate [3188]. Big [1787, 1488, 436, 101]. Bigeye [1528, 1101, 721]. Bigfin [351]. Bight [1045, 2774, 1458, 1673, 1803, 583, 1908, 126]. Bigmouth [2841]. BIIGLE [2, 1330, 3022]. Bilirubin [2280]. Billows [2301]. Binary [2944]. Bind [2399]. Binding [2550, 3127]. binucleata [1259, 1261]. Bio [1937, 2593, 1659, 1826, 1105]. Bio-labile [1826]. Bio-Logging [2593]. Bio-Optical [1937, 1105]. Bio-Oscillator [1659]. Bioaccumulation [963]. Bioactive [2864, 135, 2497, 712]. Bioactivity [1782, 368]. Bioavailability [2672, 1162]. Biochemical [2800, 1693, 948, 1297, 293, 382, 171, 869]. Biochemistry [2090, 2676]. Biocomposites [3039]. Bioconstructions [1087]. Biodegradable [2052]. Biodiscovery [2369]. Biodiversity [1526, 1808, 3066, 3209, 336, 1161, 1065, 1978, 1189, 2815, 936, 486, 48, 2244, 1351, 931, 1087, 1236, 2570, 1746, 3114, 1246, 560, 144, 2183, 2874, 3185, 2348, 2803, 318, 3023, 2585, 271, 3067, 1465, 1018, 2118, 2398, 341, 2064, 1508, 3194, 1513, 2913, 2572, 975, 1641, 1318]. Bioeconomic [30, 561]. Bioeconomy [1347]. Bioenergetic [1527, 1523, 1635, 1219]. Bioenergetics [2987]. Biofertilizers [2352]. Biofilm [2310, 364, 24, 952, 167]. Biofilms [352, 2405, 1881, 125]. Biofloc [3210]. Biofouling [3046, 3041, 492, 2035, 3016, 829, 539, 3206]. Biogenic [3060, 1308, 3186, 1785, 382, 171, 2422]. Biogeochemical [651, 2251, 946, 682, 2328, 1979, 223, 488, 1512, 2808, 2326, 2521, 1603, 599, 759, 477, 253, 1421, 521, 1348, 182]. Biogeochemistry [2492, 2692, 2879, 683, 1448, 2743, 874, 950, 1306, 820, 1201]. Biogeographic [1818, 677, 1661]. Biogeographical [1969]. Biogeographically [1926]. Biogeography [2578, 341, 374, 551, 1080, 1356]. Bioindicators [2360]. Bioinformatic [1113]. Bioinformatics [2434]. Bioinspired [1314]. Bioinvasion [2256]. Biological [2334, 1104, 3042, 2520, 79, 638, 3147, 890, 121, 1518, 3202, 3089, 2901, 1589, 647, 2617, 475, 487, 2154, 791, 1722, 2298, 1442, 1546, 623, 613, 2620, 1884]. Biologically [1769, 2128]. Biology [651, 789, 1670, 2993, 1693, 1196, 2068, 1312, 1288, 3092]. Bioluminescence [1731, 1050, 1842, 3051, 2343]. Bioluminescent [1462, 2051, 181]. Biomagnification [963]. Biomarker [837]. Biomarkers [1258, 2304, 2988, 2525]. Biomass [867, 522, 1024, 828, 1021, 2804, 49, 734, 1723, 161, 2082, 2232, 664, 948, 750, 2101, 626, 1452, 1177, 2734, 2669, 2422, 1287, 1134, 1303, 361, 210]. Biomaterial [1498]. Biomineralization [974]. Biomolecular [3013]. Biophysical [1741, 590, 2691, 2942, 1507, 1659, 1015, 729, 325, 123, 617, 1664, 3142]. Bioplastics [331]. Biopolymers [2550]. Bioprospecting [2374, 1895]. Biopsies [236]. Bioresource [260, 166]. Biosecurity [492, 2748]. Biosilica [1782]. Biosphere [748, 936, 2567, 82, 3113, 2055]. Biosynthesis [207, 2866, 941]. Biosynthetic [2658]. Biota [1532, 1999, 319]. Biotechnology [1312, 380, 1677, 2354, 1767]. Biotic [2827]. Biotopes [373]. Biotoxin [1700]. Biotoxins [1159]. Bioturbation [2597]. Bioturbator [2765]. Bird [1246, 2529]. Birth [1212]. Biscay [2560, 1548, 410, 978]. biscaya [3179, 3178]. Bivalve [2249, 1992, 2667, 797, 2318, 2937]. Bivalves [1480, 569, 2960, 1389, 1258, 2357, 154]. Bivalvia [450, 1080, 1356]. Black [542, 2589, 3099, 1073, 581, 2464, 274, 2913, 2992, 1229, 752, 2740, 197, 2859, 206, 2630, 977, 679, 3141]. Blade [1039]. Blades [1660]. blainvillei [801, 1284]. Blamed [652]. Blanks [46]. Bleached [1855]. Bleaching [134, 863, 922, 997, 1855, 1535, 448, 920, 616, 689, 2115, 822, 804, 347, 3, 1069, 1404]. Bleeding [1278]. Bloat [435]. Blood [1175, 2676, 1751]. Bloom [196, 7, 1346, 1792, 1426, 2842, 2077, 1105, 2167, 2307, 2922, 165, 2824, 3083]. Bloom-Forming [1426]. Blooms [1282, 2787, 3060, 1978, 2865, 1505, 1778, 1723, 1606, 2635, 68, 138, 2796, 1628, 518, 311, 439, 715]. Blowout [872]. Blubber [1449]. Blue [1347, 23, 314, 105, 3042, 807, 482, 371, 652, 1253, 492, 2309, 728, 505, 60, 2638, 2717, 735, 2406, 1957, 1634, 1916, 2545, 688, 1910, 2206, 147, 3121, 870, 2966, 27, 2664, 2712, 2191, 2381, 1034, 730, 2049]. Blue-Green [2406]. Bluefin [1145]. Bluenose [1486, 1806]. Blueprint [1120, 2278]. Blunt [388]. Boating [875]. Boats [1624, 455, 1563, 1231]. Bobtail [1351]. Bodied [3137]. Body [1534, 2948, 1783, 2952, 1951, 2676, 662, 1199, 2762]. BOEM [3033]. bogaraveo [2556]. Bohai [182, 2456, 3077, 2215, 354, 266]. Boldness [3099, 1286]. Bomb [76, 290, 1092, 245]. Bombay [2780]. Bonaire [448, 3, 1069]. Bone [2939]. Bones [1887]. Bony [2841]. Book [2511, 64, 2067]. Boosted [1019]. Borderland [1484]. Bore [453]. Borealization [593, 2525, 2117]. Boreo [1196]. Boreo-Arctic [1196]. Boring [1080, 1356, 154, 2002]. Born [1880]. Borne [2959]. Both [522, 1302, 1177]. Bottleneck [1658]. Bottlenecks [2690]. Bottlenose [1534, 225, 2974, 1464, 605, 855, 2079, 1584, 1181, 1550, 3170, 1490, 1551, 875, 2635, 1343, 2984, 1563, 876, 1688, 2653, 2854, 1621]. Bottom [854, 2794, 1753, 2328, 1249, 1388, 986, 1190, 823, 83, 1704, 1652, 335, 2978, 1990, 1018, 559, 1075, 1121, 1747, 1613, 1835]. Bottom-Moored [823]. Bottom-Mounted [1121]. Bottom-Set [1249]. Bottom-Towed [335]. Bottom-Trapped [1075]. Bottom-Trawling [559]. Bottom-Up [83]. Bottom-Water [2978]. Boulder [1950, 2926]. Boundaries [2536]. Boundary [1640, 1183, 1058, 43]. Bovichtidae [523]. Bowhead [1527, 2619]. Box [2268, 2958, 2464, 2949]. Box-Type [2268]. Brachionus [1515]. Brachyuran [3189]. Brackish [24, 2232, 1570]. Brain [2639, 3056, 1565, 2810, 707]. Brains [883]. Branch [1292, 2727]. Branched [860]. Branches [1902]. Branching [1888]. brasiliensis [1646, 935]. Brazil [2909, 583, 3025, 1476, 281, 2312, 1181, 1112, 514, 1284]. Brazilian [854, 1479, 601, 2021, 1863, 962, 2454]. Breadth [524]. Breakdown [1728, 2399]. Breaking [1824, 1671, 1882, 529, 251]. Breakwater [2268]. Bream [1720, 2024, 2497, 1744, 2090, 26, 2574, 388, 3141]. Breathing [2426]. Breeders [1988, 740]. Breeding [2094, 1359, 2170, 160, 167, 170, 2084, 1231, 2380]. Brest [2580]. Brevetoxin [1491]. brevicaudata [1822]. brevirostris [1556]. brevis [1491, 2635]. Bridge [306]. Bridging [987, 1416, 1133]. Brief [1324, 2712]. Brightness [323, 808]. Brine [675, 2835]. Brings [1875]. Bristles [1880]. British [2036, 21]. Brittle [2829]. Broad [2091, 877, 1413]. Broad-Scale [2091]. Broader [479]. Bronze [3114]. Brown [188, 2918, 1039, 2667, 2971, 1818, 999, 2011]. Browsing [2, 1330]. Brünnich [2556]. Bryde [438, 1934]. Bryopsidales [2912]. Bryozoans [1087]. BSM [532, 720, 725, 718, 742, 831, 844]. Budget [1007, 2384, 246, 1035, 2998, 326]. Budgets [2325]. Buffalo [2841]. Building [589, 144, 1373, 3201, 2587, 1182, 1654, 2225, 1947, 870]. Built [2264]. Bulk [1783, 2223, 2597]. Bull [2273, 1125, 2083]. Buoy [2140, 1139, 1967, 1648, 2701]. Buoyancy [377]. Buoyant [251, 2303]. Buoys [814]. Burbot [2840]. Burial [104]. Burrard [21]. Burrowing [2597]. Business [260, 166, 830]. Butenolides [703]. Butyrate [2579, 2870]. By-Products [2745]. Bycatch [2711, 3144, 801, 1846, 3132, 1249, 1559, 962, 673, 2717, 641, 2964, 387, 78, 876, 1691, 1513, 1514, 3005, 1049]. Bycatch/Interaction [3144]. Bycaught [2560, 1548].

C [815, 3110, 866, 42, 46, 236, 2866, 3027, 1165]. C-Terminal [1165]. C/N [3110]. C1Q [1805]. C2 [499]. Ca [1256]. Cabled [653]. Cabo [1921, 3023]. Cachucho [1274]. Cadiz [143, 229]. Cadmium [1539, 1979, 1112, 604, 208, 2962]. Cage [747, 1799]. Caicos [792]. Calanoid [376, 2781]. Calanoida [2564]. Calanus [868, 1248, 1858, 878]. Calappa [2958]. calcarifer [1920, 1515]. Calcification [230, 107, 1006, 3154, 1398, 916, 239]. Calcified [789]. Calcifiers [682, 1405, 2965]. Calcifying [2443, 1111]. Calcium [846, 2080, 333]. Calculating [2405]. Calculation [943]. Calculi [1401]. Caledonia [2487]. Caledonian [1089]. Calf [1551]. Calibration [500, 1675, 2069]. Calibration/Correction [500]. California [1458, 2149, 2668, 126, 911, 1572, 80, 1865, 2387, 1510, 1678, 2356, 418, 1244, 2186, 1768, 2318, 932, 101, 63, 691, 1611, 1934, 2613]. californianus [1244, 1017, 1768]. Call [2305, 1674]. Callinectes [735, 2712]. Calling [2034]. Callistoctopus [1195]. Callogorgia [2829]. Calls [1934]. Calmodulin-NOS [1256]. Calodium [437]. Calories [2333]. Camera [1867, 755, 2593, 1236, 2738, 1696]. Cameras [381, 2924]. Camouflage [1800]. Campaign [317]. Campaigns [128, 2019]. Can [1991, 3100, 107, 652, 108, 2246, 2238, 1204, 1441, 661, 3134, 1257, 2546, 1950, 1869, 384, 1686, 758, 1958, 822, 2741, 2553, 1886, 2436, 115]. Canada [2619, 485, 1281, 2243, 771, 1948, 2753, 1156, 753, 694, 3198, 360]. Canadian [3062, 2884, 597, 2427, 1448, 2700, 2109, 1976, 1442, 1509, 2525]. canadum [2214]. Canaria [1423]. Canary [1423, 2438, 94, 1117]. Cancer [1244, 1319, 101]. Candidate [2655, 562, 2886]. Canned [3180]. Canopies [676]. Canopy [842, 2975, 3029]. Canopy-Forming [3029]. Cantabrian [1506, 680]. Canyon [1009, 283, 693, 2425, 2222, 1477, 722, 1929]. CANYON-MED [722, 1929]. Canyon-Valley [1477]. Canyons [1009, 668, 2437, 710, 1274, 3059, 113]. Capabilities [979, 2143]. Capacities [1375]. Capacity [2203, 232, 2195, 2933, 3205, 811, 144, 886, 2627, 583, 1056, 2676, 3027]. Cape [1990]. Capelin [1396, 360]. capensis [927, 2540]. Capillaria [437]. Capital [2397, 1253, 160]. CAPSTONE [6, 92]. Capsular [1972]. capsulatus [3061, 3141]. Capsule [2048]. Captive [1461]. Capture [243, 2699, 1703, 1490, 374, 334, 539, 985]. Captured [2822]. Captures [153, 757]. Capturing [1642, 1696]. Carbohydrate [2342, 2537, 80]. Carbon [308, 163, 2044, 1053, 267, 2226, 329, 619, 104, 1456, 2882, 105, 279, 607, 44, 482, 3147, 866, 552, 1715, 1183, 3012, 2488, 666, 674, 2342, 730, 505, 1547, 1007, 3057, 2527, 3192, 1673, 690, 3000, 1819, 49, 798, 620, 2755, 1299, 810, 90, 246, 2172, 2871, 568, 2880, 2923, 2049, 910, 1871, 991, 452, 904, 1317, 2206, 1, 1013, 1003, 635, 306, 443, 2422, 326, 2476, 350, 777, 2449, 546, 2620, 1864, 1292, 1203, 878, 790]. Carbonate [632, 107, 2693, 1160, 722, 1929, 1192, 846, 2080, 130, 1904, 1589, 1141, 1371, 477, 1082]. Carbonate-Associated [1082]. Carbonates [2504]. Carboxylicivirga [2801]. Carcharhinus [228, 1125, 1092, 2083]. carcharias [2387, 370, 2902, 524, 1854]. Carcharodon [2387, 370, 2902, 524, 1854]. Carcinoscorpius [2967, 2093, 1076, 1498]. Cardiac [2540]. Cardinal [814]. Care [2330, 2854]. Career [794, 1385, 1749, 1719, 1182, 2407]. Caretta [1773, 808, 323]. Cariaco [916]. Caribbean [2137, 1820, 199, 2736, 585, 230, 2182, 2614, 778, 2122, 1290, 1909, 830, 1654, 2353, 2581, 1220, 448, 2829, 1644, 82, 2892, 2519, 3113, 2159, 1962, 3, 1069, 2168]. carinicauda [3021]. Carlo [532, 2822]. Carnivorous [1600, 1050]. Carnivory [1802]. caroli [2334]. carotene [2920]. Carotenoid [2920]. Carp [2603, 2489]. carpenteri [1506]. Carpet [2776]. Carrier [2639]. Carriers [2639]. Carrying [2933]. Carryover [1149]. cart [3056]. carterae [2415]. Cartilage [2562]. Cartography [1654]. Caryophyllia [50]. Cas12a-Mediated [1226]. Cascade [2332]. Case [3031, 1472, 2930, 3144, 842, 3191, 2202, 1360, 1052, 1161, 1065, 489, 278, 223, 2567, 1779, 1559, 1118, 1106, 228, 2383, 949, 928, 1721, 1843, 418, 2455, 505, 673, 3077, 30, 2265, 1742, 130, 1904, 1004, 1195, 1305, 1528, 2487, 2563, 1522, 2646, 1374, 2890, 3097, 2433, 318, 448, 2221, 1866, 2556, 2062, 1744, 442, 2345, 127, 1735, 2245, 1191, 1452, 2891, 2744, 927, 125, 1509, 2336, 561, 2649, 3, 1069, 27, 1180, 1611, 2168, 3036, 3078, 1917, 843, 614, 1404, 1205, 2620, 338, 2438].

Casey [582]. Cashing [762]. Cassiopea [2583]. Casual [2273]. Catalina [3074]. Catalog [2010]. Catastrophe [886]. Catch [1815, 3104, 2682, 371, 2356, 2137, 782, 3077, 503, 1122, 2065, 2058, 93, 2177, 2602, 532]. Catch-MSY [532]. Catchability [2086]. Catches [116, 3054, 2101, 3087]. Catching [442]. Catchment [1071, 2394]. Catchment-Estuary-Coastal [1071]. Catchment-Scale [2394]. catenella [1940, 1281, 1137, 2782, 1505]. Catenin [2361]. Catfish [3152, 293, 941, 2436]. Cauca [1387]. caudatus [1118]. Caudofoveate [2158]. Caulerpa [3123, 2302, 1240, 2541, 1364]. Causal [1919]. Causative [3061]. Cause [648]. Caused [3122, 1582, 1117, 226, 2261, 2767]. Causes [1885, 421, 1047, 650, 1689, 2561, 772, 2410, 3093]. Causing [1892, 392]. Cautionary [533]. Caveats [2412]. cavernosa [2509, 2666, 2687, 830, 2053, 2311, 2107]. Caves [1694]. Cavity [1791]. Cayman [2070]. CCAMLR [1421]. CCZ [1692]. Cecropin [1115]. Cell [2918, 1243, 3154, 1226, 2033, 3061, 356, 2271, 1288, 703, 388]. Cell-to-Cell [703]. Cells [1971, 3201, 2587]. Cellular [3188, 1728, 2460, 2306, 2987]. Celtic [1066]. Cement [1128]. Center [1329, 261, 273]. Centers [902]. Central [998, 2621, 2845, 1118, 112, 2771, 721, 2364, 2535, 2334, 2968, 471, 699, 2678, 1127, 2465, 1130, 335, 304, 1048, 38, 1506, 2609, 2780, 635, 911, 1985, 1239, 1034, 182]. Central-Southern [1239]. Centralized [2553]. Centrolophidae [1486]. Centrophoridae [1138]. Centropristis [977]. Centuries [134, 997]. Century [121, 637, 644, 1071, 2991, 2356, 2261, 2624, 2059, 2582, 217]. Cephalopod [3023]. Cephalopoda [1351, 284]. Cephalopods [1407, 2599, 534, 151, 366, 158]. Cephalorhynchus [921]. Cephalotoxins [1895]. Cerastoderma [2495]. Ceratonova [1713]. Cercopagis [1021]. cereus [2214]. Cerium [2131]. Certification [3180]. cervicornis [2220, 2133, 2176, 1888, 938]. Cesspit [3080]. Cetacea [1174]. Cetacean [963, 225, 1123, 2711, 3144, 1930, 1849, 2383, 1617, 212, 1915, 1831, 883, 1460, 435, 1832, 1273, 995, 1459, 490, 2275, 1461]. Cetacean-Vessel [212]. Cetaceans [786, 2235, 852, 469, 2165, 1460, 1175, 2036, 3067, 732, 2234]. Cetorhinus [2095]. cf [785]. ch25h [65]. Chaceon [178]. Chaetoceros [55]. Chaetognatha [498]. Chagos [1208]. Chain [1823, 882, 941, 2367, 2658, 2822, 1479]. Chains [428]. Chairs [1959]. Chalcone [3035]. Chalk [509]. Challenge [653, 2309, 918, 431, 2142, 2624]. Challenged [1262, 2037]. Challenges [885, 1350, 2784, 85, 819, 464, 2972, 2975, 2413, 1930, 260, 166, 508, 2346, 1559, 2122, 1441, 2035, 72, 1555, 1007, 957, 60, 1852, 1004, 2078, 297, 1701, 2097, 1588, 630, 1829, 2519, 1288, 1509, 1352, 3173, 820, 1611, 1619, 1774, 1475]. Challenging [2585]. chalumnae [2989]. Cham [2567]. Cham-Hoi [2567]. Chamelea [1102]. Chance [174]. Change [1640, 809, 3180, 2739, 1954, 1300, 2662, 2784, 485, 1921, 1753, 672, 1281, 2471, 1598, 1137, 2972, 2797, 2695, 2601, 1322, 2805, 713, 979, 1070, 2763, 457, 2840, 639, 2567, 1638, 2986, 2932, 1166, 2347, 196, 1441, 1830, 2498, 649, 2256, 1066, 41, 516, 342, 1489, 1582, 2386, 1331, 2531, 1528, 1315, 706, 1400, 568, 1434, 583, 2167, 1836, 684, 77, 2632, 200, 1271, 1949, 756, 2118, 1143, 3186, 2225, 1389, 227, 759, 1478, 1707, 132, 1295, 1752, 1538, 1041, 989, 179, 743, 1780, 334, 625, 262, 1738, 983, 2381, 1613, 2389, 737, 2946, 2830, 1062, 1618]. Change-Induced [759]. Changed [1665]. Changes [1378, 779, 1395, 53, 2243, 62, 107, 659, 2034, 552, 124, 1047, 2912, 728, 2066, 7, 1346, 1448, 345, 1471, 1803, 2363, 1666, 428, 696, 2127, 2323, 969, 573, 1139, 1502, 297, 950, 1650, 2817, 2090, 817, 3130, 626, 1023, 3055, 1912, 694, 493, 2818, 288, 2445, 667, 1810, 2529, 1212, 1369, 417, 2826, 1260]. Changing [2509, 3031, 3163, 1739, 2299, 643, 1425, 502, 976, 109, 1755, 683, 60, 2928, 1251, 2082, 2557, 503, 897, 483, 40, 2934, 2891, 64, 190, 2647, 191, 311]. Changjiang [1134, 3017, 2715, 2617, 1536, 361, 1342]. Channel [1526, 3108, 763, 2955, 2001, 3212, 2436, 1750, 767, 1198, 1592, 2355, 1366, 2602]. Channels [3007, 984, 1810]. Chara [2232]. Characiformes [110]. Character [2796]. Characteristic [2408]. Characteristics [1964, 3104, 755, 2137, 1512, 3069, 2771, 2164, 2779, 1392, 2172, 2890, 2956, 1072, 1314, 2154, 1092, 2341, 2058, 117, 1498, 2284, 1422, 3064, 3200]. Characterization [1462, 1955, 522, 2920, 2280, 3157, 441, 2550, 1015, 680, 2831, 397, 463, 2854, 2992, 392, 65, 2372]. Characterize [2736, 578]. Characterizing [3180, 1350, 485, 1052, 1458, 700, 1798, 933, 2963, 848, 3187, 3087, 2675]. Charged [308]. Charismatic [2510]. Charophyte [2232]. Charybdis [2982, 3157]. Check [1581, 2254]. Chelonia [3011, 1476, 75, 2009, 749, 775]. Chemical [682, 376, 2012, 3125, 565, 2409, 441, 620, 512, 2485, 600, 3109, 288, 2719, 171, 2969, 1498, 382]. Chemically [1232, 1355]. Chemicals [209]. Chemistry [1761, 2693, 519, 1109, 1471, 1141, 449, 1371, 95, 360]. Chemo [2302]. Chemo-Ecological [2302]. Chemosynthetic [2960]. Chemotaxis [629]. Chesapeake [224, 1070, 2769, 3155, 439]. CHHs [3166]. chierchiae [2973]. Children [2511]. Chile [1500, 2535, 1760, 1763, 635, 1621, 1239]. Chilean [1756, 786, 1753, 1505, 2385, 1503, 2191]. chilensis [14, 1995]. Chimeric [1660]. China [1566, 3124, 815, 2807, 2775, 2865, 2692, 1424, 2821, 1778, 2488, 3024, 1547, 2066, 2014, 1163, 1742, 3192, 2660, 3145, 321, 2978, 3117, 1586, 1733, 959, 645, 253, 2931, 2780, 2538, 613, 2657, 2007, 840, 788, 3133, 2976, 574, 2449, 2846, 546, 1303, 2015, 1657, 2316, 2696, 1342, 3105, 2804, 3009, 2779, 2646, 2958, 1324, 1261, 1374, 1400, 3072, 1327, 2725, 915, 2893, 493, 354, 2835, 197, 742, 2444, 843, 2603, 1672, 2951, 2860, 3026, 2898, 1028, 1205, 609, 3200, 2696, 2757, 2622, 2926, 1941, 2493]. chinensis [1392, 1969, 2205, 1380, 2897]. Chinese [3125, 1853, 2514, 1728, 2792, 416, 725, 636, 22, 65, 618, 831, 400, 2446, 263]. Chip [2492, 2485]. Chips [2918]. Chirikov [1572]. Chitosan [2595, 2093]. Chl [704]. Chl-a [704]. Chlamys [2877, 1951, 2725, 3184]. Chloroflexi [776]. Chlorophyll [2596, 773, 389, 1507, 1334, 566, 452, 377, 1358, 665, 2412, 2628, 911, 1457, 2370]. Chlorophyll-a [773, 1334, 566, 1358, 665, 2370]. Chlorophyta [156, 816, 2135]. Chloroplast [2889, 3172, 2135]. Choice [1316, 666, 1339]. Choices [3042, 947]. cholerae [1609]. Cholesterol [65]. Chondrichthyan [1703]. Chondrichthyans [3190]. Chondrichthyes [1646, 393]. Chondrocytes [2134]. Chondrosia [1280]. Chorus [2211]. Chromatic [2406]. Chromatographic [276]. Chromatography [333, 860]. Chromophoric [229, 2890]. Chromosome [2877]. Chromosomes [859]. Chronic [1056, 2861]. Chronicles [519]. Chronologies [417]. Chrysogorgia [1227]. Chub [1523, 2756]. chuii [180]. Chukchi [714, 1438, 1296, 1931, 1484]. Cichlid [825]. Cíes [3118]. Ciliate [990, 1259, 475, 1242, 1837]. Ciliated [1323]. Ciliates [1243, 1853, 1324, 3072, 1327, 1629, 1821]. Ciliature [1323]. Ciliophora [1259, 1261, 1341, 3072, 1327, 1629, 1590, 1243, 1853, 1324, 1242, 1837, 1304]. Circadian [2904, 2726]. Circular [1121]. Circulating [881, 95, 2854]. Circulation [3203, 2046, 451, 1830, 3095, 470, 2487, 1586, 487, 2201, 9, 2105, 2000, 1908, 395, 2594, 2212, 1672, 1436]. Circulations [1536, 2727]. Circumpolar [518]. Cirripedia [1100]. Cities [1906, 1707]. Citizen [2247, 2099, 1488, 1580, 2753, 1591, 1831, 2186, 3116, 2315, 1552, 1406, 1832, 1999]. Citizen-Sourced [2186]. Citizens [702, 2697]. Citreamicin [330]. Citrithrix [1261]. Citrithrixidae [1261]. City [2330]. Clade [2310, 1338]. Clades [1733]. Cladocopium [3159]. Cladorhizidae [1050]. Clam [1142, 1102, 3136, 2359, 2920, 2417, 1653, 1971, 996, 3159, 3187, 118, 1751, 2002, 2758, 618, 2757]. Clam-Dinoflagellate [1653]. Clams [3124, 2945, 1733, 567, 1178]. Clarias [293]. Clarifying [1342]. Clarion [2187, 1746, 2063, 2068, 1765, 2183, 1717, 3194, 2050, 1641, 1747, 1987, 560]. Clarion-Clipperton [1746, 2068, 2183, 1717, 3194, 2050, 1641, 1747]. Clarity [2633, 1911, 2723]. Class [196]. Classes [1821]. Classical [1195]. Classification [373, 2012, 608, 292, 934, 268, 467, 113, 1814]. Classifying [913, 2591]. Classroom [2511]. clavata [1127]. Claw [2958]. Claw-Only [2958]. Cleaner [829]. Cleaning [2267, 539, 3206]. Cleanup [2903, 2979]. Clearing [2155]. Cliff [2151, 131]. Climate [3031, 3163, 2739, 3119, 2096, 1954, 1300, 2662, 2784, 643, 1921, 1753, 672, 2778, 1281, 2471, 741, 1137, 3203, 185, 2601, 1322, 1595, 2805, 1070, 2763, 639, 2567, 895, 2274, 1638, 2986, 1719, 1166, 502, 1190, 109, 196, 1830, 2498, 649, 728, 1066, 7, 60, 41, 1375, 2928, 1251, 106, 1489, 2732, 1582, 2386, 2743, 1331, 2531, 1528, 2557, 1315, 1662, 830, 999, 226, 2630, 706, 1400, 2052, 896, 1434, 510, 1551, 583, 448, 2167, 1836, 77, 2632, 200, 1949, 950, 1143, 3186, 2225, 1389, 227, 759, 2754, 2237, 1230, 1707, 132, 2197, 2584, 1041, 3, 1069]. Climate [1674, 334, 625, 2160, 262, 1764, 1738, 983, 1736, 1613, 2389, 737, 1062, 1577, 1618, 1346]. Climate- [728]. Climate-Biogeochemistry [2743]. Climate-Change [1921]. Climate-Driven [1595, 999]. Climate-Induced [448, 950, 3, 1069]. Climate-Resilient [2732]. Climatic [1630, 133, 863, 1984, 2199]. Climatological [984]. Climatology [3028, 587]. Clinical [1409]. Clipperton [1987, 2187, 1746, 560, 2063, 2068, 1765, 2183, 1717, 3194, 2050, 1641, 1747].

Clocks [2653]. Clonality [2130]. Cloning [65]. Closed [2394]. Closer [2608]. Closing [1326]. Clostridium [3065]. Closure [1815, 2561, 3128]. Closures [1509, 1691]. Cloud [2596, 3060, 2453, 1202, 2508]. Cloud-Free [2596]. Cloud-Native [2508]. Clues [1437]. Clupea [1412, 1218]. Clustering [2930, 112, 3112]. Clymene [2018]. CMAR [2244]. CMEMS [1455, 1861]. CMEMS-Based [1455]. CMSY [720, 725, 532, 718, 742, 831, 844]. Cnidaria [1993, 2663, 1713, 2970, 63, 901, 2893, 1944]. Cnidarian [1713]. Cnidarians [3024]. cnidata [368]. Cnidome [2663]. Co [2387, 484, 152, 1215, 42, 2697, 317, 481, 2101, 2560, 1548, 2136, 2064, 2431, 1737, 1771, 2916, 814, 2378, 2983, 1894, 1111, 2617, 736, 1847, 2855, 86, 613, 409, 447, 781, 1725, 1917]. Co-construction [2697]. Co-culture [1215]. Co-governance [2136]. Co-limitation [42]. Co-movement [2101]. Co-Occurrence [2064, 2387, 484, 152, 2560, 1548]. Co-occurring [481]. Co-variability [317]. Coagulation [1175]. Coarsening [3040]. Coast [1235, 2930, 1433, 663, 1678, 2955, 685, 1415, 555, 2266, 212, 2018, 1792, 135, 1953, 2598, 2557, 2056, 1335, 959, 2141, 1131, 2030, 1820, 1604, 2197, 694, 125, 2855, 1091, 2926, 1941, 786, 854, 1281, 1309, 2385, 1995, 3126, 503, 50, 410, 845, 1776, 648]. Coastal [885, 2507, 1042, 1276, 2100, 34, 1739, 19, 926, 2930, 1071, 586, 888, 3191, 577, 279, 1464, 2641, 1978, 1146, 1228, 2633, 1911, 2684, 2872, 1332, 1595, 2397, 1397, 3136, 2359, 370, 1781, 1505, 2046, 2346, 2763, 685, 1377, 1749, 1249, 3181, 792, 1779, 1901, 986, 1948, 552, 541, 1344, 228, 1596, 2079, 519, 247, 2312, 1645, 434, 814, 2378, 2667, 2382, 1853, 1097, 1267, 1843, 3156, 2539, 277, 418, 1555, 2751, 1516, 650, 298, 1914, 1550, 186, 201, 1375, 2643, 2656, 1803, 929, 2318, 2140, 720, 2629, 2315, 2082, 1693, 850, 2957, 738, 2869, 2979]. Coastal [2487, 478, 461, 1324, 636, 1536, 1374, 1570, 2562, 2679, 1906, 959, 2261, 3072, 1327, 2031, 503, 1902, 772, 875, 1727, 1896, 3098, 1485, 437, 2241, 944, 2149, 8, 84, 2000, 1193, 900, 750, 2624, 2152, 425, 1171, 183, 2193, 1966, 2352, 2141, 2252, 894, 2398, 2508, 1540, 1380, 2580, 1883, 1307, 1905, 1683, 40, 637, 1958, 2448, 1821, 2296, 122, 906, 2694, 1455, 1707, 665, 1133, 753, 1752, 1288, 2690, 2584, 2618, 3086, 2835, 1283, 395, 1861, 1306, 3173, 74, 2731, 2538, 1153, 262, 1918, 2332, 1856, 1554, 360, 964, 1839, 1581, 1816, 3001, 3211, 2788]. Coastal [826, 1014, 1495, 2674, 3155, 174, 648, 2608, 847, 843, 614, 2181, 350, 777, 1287, 400, 2446, 361, 1618]. Coastal-Ocean [1902]. Coastal-Pelagic [1595]. Coastal-Plain [2679]. Coastalscapes [3209, 2174]. Coastline [1142, 1395, 2986, 3040]. Coastlines [644, 2554, 1707, 618]. CoastNet [2592, 2358]. Coasts [3163, 1071, 1538, 1352, 671, 301, 1894]. Cobalamin [991]. Cobia [2214, 1070]. Coccolithophore [2550]. Coccolithophore-Associated [2550]. Coccolithophores [2080]. Coccolithophorid [806]. Coccoliths [994]. Cockle [2495]. Cod [805, 1975, 3062, 1396, 2101, 667, 3083, 1685, 2503, 132]. Cod-Capelin [1396]. Coded [3162]. coding [3189, 2759, 2925]. Codon [1243]. Coelacanth [2989]. coerulea [2545, 2893]. Coexistence [3134]. coffeaeformis [293]. Cohesive [2994, 2463]. Cohort [2032]. COI [636].

Coilia [1044, 3078]. coioides [2362]. Cold [346, 763, 779, 2061, 390, 2437, 2240, 121, 456, 343, 770, 3192, 235, 3037, 548, 558, 2761, 2297, 1411, 1588, 901, 1858, 1368, 2057, 315, 3129, 2504, 3036, 1814, 733]. Cold-Seep [343]. Cold-Water [763, 2061, 390, 2240, 121, 456, 770, 235, 548, 2297, 1588, 901, 1368, 2057, 315, 3036, 1814, 733]. colemani [2565]. coli [1001]. Coliform [1817]. Collaboration [2155, 818, 144, 1182]. Collaborations [1385, 442]. Collaborative [1194, 1182, 1896, 1883, 2354]. Collagen [791]. Collapse [226, 2754]. Collapses [1118]. Collected [484, 152]. Collection [1123, 1488, 1757, 2743, 1999, 832, 2434, 405, 1325]. Collections [2, 1330, 184]. College [1305]. Collichthys [2446]. Collisions [212, 394]. Colombia [1220, 199, 1376, 748, 1604, 3074]. Colombian [82, 1820, 413]. Coloniality [563]. Colonies [881, 149, 596, 1384]. Colonisation [3108]. Colonization [545, 1204, 2264, 325, 751, 211]. Colonizers [565]. Colonizing [5, 2424]. Colony [2061, 1903, 2365, 1888]. Colony-Level [1888]. Color [971, 1316, 1843, 2952, 2512, 2842, 3162, 165, 2002, 2886, 2762]. Color-Coded [3162]. Colored [1789, 2643]. Columbia [21]. Column [2377, 1496, 337, 507, 722, 1929, 2677, 1881, 1454, 1952, 3093, 1342, 80]. Combat [1200]. Combination [1600, 3072, 1327, 2474, 898]. Combinations [2204]. Combine [1161]. Combined [713, 1349, 654, 513, 1596, 1391, 1847, 2615, 2980]. Combining [3025, 2453, 1279, 572, 2348, 578, 3023, 2298, 1679, 2600]. Come [2126]. Coming [536]. Command [499]. Commandments [1538]. Commentary [134, 997, 808, 2473, 448, 1069, 2600, 1298, 997]. Comments [2018]. Commerce [1693]. Commercial [821, 14, 2248, 2021, 213, 1533, 720, 1374, 969, 503, 3090, 913, 1868, 2616, 2468, 1229, 752, 2572, 2177]. Commercialization [2155, 3098]. Commercially [1027, 1219, 178, 2357]. Commerson [3153, 3104]. Commission [2132]. Commitments [100]. Common [1534, 1541, 2974, 2393, 1401, 640, 1302, 534, 3035, 875, 2635, 1859, 2857, 1546, 2560, 1548, 982, 2653, 2719, 2748, 1203]. Commonly [108]. Communication [1859, 405]. Communities [163, 3058, 3047, 1433, 1753, 1464, 336, 2173, 1598, 1977, 767, 2437, 48, 2612, 556, 1388, 3003, 1692, 1779, 1718, 855, 1624, 1047, 109, 2983, 3156, 3069, 2498, 2607, 2718, 1819, 2747, 478, 1088, 2978, 2703, 1020, 503, 1788, 2803, 583, 1485, 2425, 2995, 1836, 68, 1018, 1650, 2145, 2910, 626, 2206, 2022, 1307, 195, 592, 140, 1295, 1129, 1473, 582, 2731, 2849, 1201, 2291, 1784, 1764, 3088, 3149, 118, 1756, 191, 1597, 1747, 2434, 1247, 574, 2881, 3064, 3204, 2765, 2551, 2441]. Community [243, 779, 146, 329, 1410, 364, 1845, 275, 2974, 1612, 1030, 1420, 1425, 2255, 1936, 62, 107, 3147, 1397, 2575, 2691, 1680, 1022, 1225, 3135, 462, 1424, 1061, 2848, 1097, 2196, 1458, 97, 812, 3114, 1824, 1483, 1272, 1828, 2700, 2319, 2232, 317, 932, 3170, 1078, 1182, 1889, 256, 2355, 1366, 2803, 356, 495, 956, 908, 384, 2090, 597, 1942, 680, 357, 576, 3013, 2200, 368, 595, 549, 1275, 1336, 493, 302, 1306, 1620, 179, 1269, 86, 2966, 2849, 2824, 1497, 1925, 564, 191, 2588, 1369, 2029, 254, 2472, 715, 1124, 3048, 1553, 2438]. Community-Based [2803, 1942]. Community-Scale [3135]. Compact [2305]. Companions [2273]. Company [2158]. Comparative [129, 541, 744, 2066, 2971, 2599, 3037, 1520, 1396, 2068, 3172, 2135, 2961, 2956, 1731, 1274, 412, 3021, 358, 2573, 123, 517, 3166]. Compared [2722, 374, 2525]. Comparing [2948, 704, 1903, 1156, 533, 906, 1691, 2168, 788, 1696, 2333]. Comparison [1293, 1102, 2736, 2203, 406, 1789, 1523, 2463, 49, 330, 2056, 1727, 2923, 2554, 2661, 208, 2337, 662, 2340, 594, 2931, 392, 631, 1772, 1623, 2420, 3176, 2811, 3103, 1814]. Comparisons [1688]. Comparsion [220]. Compartmentalization [1112]. Compatible [147]. Compensation [1056]. Compensatory [667]. Competent [2670]. Competing [736]. Competition [42]. Competitive [2490, 2267, 2759]. Compilation [570]. Compiling [2748]. complanata [1373]. Complementarity [188]. Complementary [1032]. Complete [542, 1110]. Complex [3191, 2612, 2094, 2576, 1031, 241, 1888, 2954, 1829, 2908, 2282, 637, 838, 198, 671, 614, 2976, 1422, 3165, 2983]. Complexes [2062]. Complexity [1706, 2185, 1270, 1877, 1474, 1114, 2257, 1417, 2237]. Compliance [1072, 1927]. Component [2820, 1651, 2959, 908, 2744, 4, 2466]. Components [2918, 2195, 930]. Composition [1534, 2251, 1009, 1410, 2736, 2586, 364, 273, 2800, 1612, 337, 2039, 2520, 1250, 1716, 2682, 2840, 1718, 3017, 1596, 2853, 1068, 524, 2480, 877, 1036, 7, 1346, 3009, 620, 317, 1289, 2713, 633, 1335, 256, 1502, 3151, 1634, 356, 916, 597, 2676, 3027, 355, 680, 2892, 357, 530, 576, 2671, 187, 103, 2924, 2777, 1162, 86, 2720, 617, 191, 2007, 2929, 1482]. Composition-Based [877]. Compositional [1981]. Compositions [1819]. Compound [710, 2223, 943]. Compound-Specific [710, 2223, 943]. Compounds [1308, 1319, 512, 3140, 2708, 1312]. Comprehensive [3197, 282, 2274, 1373, 825, 2759]. compressa [816, 2135]. Compression [1002]. Compromises [1868]. Computational [820]. Computed [1876, 883]. Computers [1552]. Concealing [1735]. Concentrate [2595]. Concentration [1737, 1761, 2382, 1830, 1017, 2844, 2742, 665, 1587, 911, 3174]. Concentration-Dependent [1017]. Concentrations [722, 1929, 1254, 1973, 1280, 2742, 1688, 2198, 1493, 2935]. Concepción [635]. Concept [2699, 1652, 608, 510, 2500, 2997]. Concepts [2659, 2964, 1511, 2966, 1597]. Conceptual [1928, 1531, 2275, 1544]. Concern [1647]. Concerns [2993]. Conch [748]. Condition [408, 23, 2948, 2418, 1569, 2768, 2339, 2870]. Conditional [947]. Conditioned [2478]. Conditioning [1455]. Conditions [2233, 2919, 1425, 2046, 2623, 440, 2753, 863, 976, 36, 2475, 1217, 1149, 930, 2415, 1826, 106, 1520, 1073, 2546, 503, 681, 691, 2634, 1966, 1051, 483, 2496, 1389, 2934, 1984, 1364, 368, 2719, 631, 1028]. Conductivity [2099]. Conduits [847]. Cone [3079]. Conference [1719]. Conferences [1786]. Confidence [870]. Confirm [2295]. Confirms [2875]. Conflict [931]. Conflicts [888, 1757, 25]. Confluence [37]. Confronting [2612]. Confusing [1597]. Congeneric [2949]. Congeners [196]. Conglycinin [2579, 300].

Congruent [2607]. Conjugated [2853]. Conjugation [1323]. Conjunction [3012]. Connect [1690]. Connected [476]. Connecticut [2929]. Connecting [102, 2001, 503, 1611, 847]. Connections [3115, 5, 2424, 1269]. Connectivity [827, 3066, 1541, 2578, 3082, 2691, 498, 3161, 414, 3179, 1290, 710, 2900, 1806, 340, 3178, 487, 1026, 2709, 2910, 3194, 123, 199, 2929, 56]. Conopeptides [3079]. Consecutive [686]. Consensus [931, 2533]. Consequences [1433, 218, 741, 650, 882, 772, 40, 1230, 1427, 2391, 2389]. Conservation [1499, 91, 225, 466, 1185, 2640, 2202, 1320, 1647, 1549, 270, 2413, 370, 1492, 2246, 2612, 1252, 792, 2244, 124, 414, 2327, 2795, 2395, 1721, 72, 962, 1915, 2900, 1246, 3070, 2993, 2165, 1163, 1245, 2637, 2426, 960, 1008, 1354, 2259, 1460, 25, 2626, 1798, 2300, 731, 363, 1370, 2222, 688, 908, 2709, 344, 2210, 1829, 1942, 2064, 1797, 533, 2260, 1857, 2144, 1558, 711, 1041, 1513, 1138, 179, 1780, 1180, 1597, 289, 1835]. Conservation-Focused [370]. Conservative [393]. Conserved [2791, 938]. Considerations [2566, 2035, 2704, 1711, 2318, 2360, 2147, 3206, 1747, 1833]. Considering [2616]. Consistency [1007]. Consistent [1783]. Consortia [654]. Consortium [2670, 711, 54]. Conspecific [1523]. Conspicuous [519]. Constant [483]. Constituents [3125, 948]. Constrain [2449]. Constrained [497, 252]. Constrains [1554]. Constraint [2690]. Constraints [2520, 1557, 2472]. constricta [2904, 3187, 2758]. Construction [2091, 2514, 2877, 2655, 1380, 2354, 3050, 1450, 400, 2697]. Construction-Degraded [1450]. Consume [778]. Consumer [1793, 2892]. Consumers [1142]. Consumption [1539, 1633, 815, 1308, 958, 1766, 1248, 1222, 2892]. Contact [96]. Contact- [96]. Container [892]. Containing [2806]. Containment [793]. Contaminant [2785]. Contaminants [1244, 2409, 732]. Contaminated [2459, 2881]. Contamination [2330, 1476, 2321, 2313, 915, 3020]. Contemporary [336, 246, 1013, 2279]. Content [404, 807, 1250, 1830, 580, 49, 925, 375, 667, 695, 2759]. Contents [1894]. Contesting [2345]. Context [1123, 557, 2567, 2426, 1913, 1822, 2997, 1784]. Contextualizing [2577, 928]. Contiguous [960]. Continental [887, 611, 2821, 1448, 1589, 2477, 292, 1699, 1585, 2995, 780, 2200, 70, 1854, 113, 315, 1668, 777, 2211, 389, 1458, 559]. Continental-Scale [2477]. Continue [1450]. Continuous [1093, 2393, 1415, 385, 1859, 1212]. Continuum [1547, 2000]. Contraction [2794, 216]. Contraction-Expansion [216]. Contrasting [1643, 107, 2437, 2561, 810, 90, 2681, 2995, 2665, 1682, 1059, 1984, 2530]. Contribute [1915, 1886]. Contributes [1316, 1056]. Contribution [105, 867, 2085, 1109, 3151]. Contributions [2044, 1964, 2825, 2643, 2403, 1324, 159]. Contributors [271]. Control [390, 215, 2344, 499, 627, 235, 754, 1606, 2326, 1220, 1494, 1399, 3169, 1258, 617, 2175, 1287]. Control-Impact [3169]. Controlled [2401, 1134]. Controlling [137]. Controls [971, 1009, 2520, 590, 863, 3024, 2480, 83, 1589, 1918, 1825, 878]. Conus [3079]. Convention [2549]. Conventions [2136]. Convergence [57, 2239]. Converter [1648]. Conveyed [1023]. Convolutional [1470, 2605, 2686]. Cook [248]. Cookie [1646]. Cookie-Cutter [1646]. Cooler [822]. Cooling [2766, 753]. Cooperation [2244, 567]. Coordinated [2207]. Cope [127]. Copepod [1940, 2179, 376, 133, 2484, 2077, 2076, 1720, 433, 2781, 160, 582, 17, 3174]. Copepoda [2564]. Copepods [1527, 1978, 2014, 714, 1634, 878]. Copernicus [2569, 2548]. Copes [1493]. Coping [2043]. Copper [3049, 1017]. Coral [2509, 2799, 466, 763, 1009, 146, 1954, 1300, 2736, 2061, 390, 643, 2202, 1795, 230, 2673, 406, 1811, 2400, 584, 2182, 936, 107, 1481, 2041, 231, 2322, 1903, 2246, 2614, 2317, 1094, 2567, 856, 134, 2240, 778, 1006, 2666, 1249, 57, 1167, 951, 2365, 281, 2687, 513, 2344, 1718, 1599, 863, 922, 456, 2390, 96, 2576, 414, 1721, 72, 2453, 2544, 1237, 730, 1812, 1257, 2950, 1562, 1184, 770, 13, 886, 2768, 1272, 2771, 235, 3201, 2587, 1693, 193, 1602, 2319, 717, 1086, 548, 1043, 1535, 830, 474, 2323, 1654, 633, 562, 479, 1147, 2626, 50]. Coral [1950, 2568, 581, 968, 2554, 1798, 3138, 2681, 448, 1216, 2180, 920, 2432, 616, 1888, 2097, 2545, 689, 1002, 2115, 2522, 265, 2610, 3034, 1982, 2250, 412, 2709, 2940, 2591, 2130, 2810, 2337, 1368, 1762, 195, 2054, 1023, 551, 1947, 2257, 1797, 938, 1958, 2616, 838, 198, 2458, 2103, 2671, 2057, 822, 2491, 774, 804, 347, 463, 2159, 711, 1352, 1962, 2157, 2749, 2376, 2832, 3, 1069, 1384, 2125, 955, 2992, 1201, 2291, 2720, 2798, 98, 2311, 707, 1772, 2168, 2533, 975, 2107, 648, 3036, 254, 970, 289, 1702, 2139, 301, 2953, 1404]. Coral [1247, 233, 1303, 2015, 1864, 2926, 2762, 1814, 3150, 733, 248, 995]. Coral-Dissolved [1811]. Coral-Eroding [2115]. Coral-Symbiont [2798]. Corales [1376]. Coralligenous [88]. Coralline [2038, 454, 1041, 1154]. Corallistidae [680]. Corallium [584, 1721, 1797, 2125]. Coralporosis [770]. Corals [1233, 542, 1937, 563, 1708, 2682, 1923, 589, 2666, 121, 2983, 2768, 2170, 1855, 2779, 479, 2037, 2306, 1588, 745, 1996, 901, 2169, 597, 2225, 845, 3010, 2500, 2053, 315, 2664, 2740, 405, 2953, 2392]. Core [2802, 2435, 2974, 1022, 782]. Cores [2692]. coriacea [1104, 3132]. Coris [2455, 2908]. Coronavirus [1906]. Correction [500, 407]. Corrections [2370]. Correlated [563, 87]. Correlated-Error [87]. Correlates [2626]. Correlation [2839, 3145]. Corridor [1690, 2244]. Corridors [310]. Corrigendum [2185, 2226, 1954, 1337, 2633, 2815, 484, 2601, 3136, 260, 1929, 2019, 2475, 2704, 1346, 1904, 1340, 3201, 92, 1354, 2531, 810, 1330, 3072, 2355, 1334, 2343, 2648, 2105, 84, 1321, 2560, 2117, 2702, 1356, 366, 790, 804, 2424, 2441, 2188, 382, 2466, 3018, 1229, 3211, 378, 3148, 3176]. Corsica [763]. Corticosteroids [95]. Corticosterone [2436]. Cortisol [2854]. coruscus [222, 2843, 952, 2987, 458]. Corynactis [1650]. Coryphaena [1963, 1637]. Cosmetic [2330]. Cost [771, 2903, 2327, 1236, 1659, 1413, 995, 535]. Cost-Benefit [2903, 1659]. Cost-Effective [2327, 995]. Costa [809, 1104, 1361]. Costs [2406, 1588]. Cottonseed [2595, 3130]. Cottoperca [523]. Could [1742, 2871, 1935]. Count [2333]. Counter [1644]. Counteract [1802]. Counteracting [2534]. Counteracts [905].

Countering [1573]. Counting [2005, 1676]. Countries [3144]. Country [1787]. Counts [2628]. County [2668, 2749]. Coupled [1167, 1670, 41, 2873, 1697, 1684, 599, 623, 1390, 1210]. Coupling [622, 1312, 506, 1455, 372, 361, 2138]. Course [2782]. CoV [2153, 2156]. Covariance [221]. Cove [2755]. Cover [2736, 2884, 2928, 1662, 2536]. Coverage [1189, 2815, 13]. Covered [1198, 1860]. Covert [2181]. COVID [1807, 2231, 2633, 1911, 793, 2240, 1789, 2229, 2242, 1730, 2426, 2152, 2834, 2367, 2030, 2364, 1983, 1912, 2156, 2701, 2851, 2402, 2697, 2731]. COVID-19 [1807, 2231, 2633, 1911, 793, 2240, 1789, 2229, 2242, 1730, 2426, 2152, 2834, 2367, 2030, 2364, 1983, 1912, 2156, 2701, 2851, 2402, 2697, 2731]. Cozumel [1718]. CPIES [2290]. Crab [1414, 1372, 2003, 2772, 851, 3084, 2349, 101, 735, 1286, 2958, 2866, 145, 503, 1418, 205, 1131, 2921, 2750, 2658, 255, 2712, 705, 178, 2416, 1475, 136, 3193, 263]. Crabs [828, 1096, 3189, 101, 2967, 2880, 503, 3054, 1076, 1278, 869, 985, 1498]. Craniella [776]. crassirostris [2077, 2076]. Crassostrea [2907, 501, 967, 1204, 700, 1152, 67, 1258, 327, 2759, 2907, 2946, 2501]. Crassostreinae [2878]. Create [2041, 2697, 923]. Created [2769]. Credit [1513]. Creeks [798]. Cretaceous [1117]. Crete [1291]. Cribroelphidium [3140]. Crime [264]. Crimes [1150]. Crinoid [1537]. Crisis [1719]. CRISPR [1226]. CRISPR/Cas12a [1226]. CRISPR/Cas12a-Mediated [1226]. Criteria [1775]. Criterion [2308]. Critical [2089, 2651, 73, 1182, 2581, 2396, 2097, 2408, 1437]. Critically [1392, 3018, 1980]. Croaker [2761, 2181, 574, 20, 2898, 2446, 1941]. Croatia [2066]. crocea [1733, 2761, 2002, 574]. Crocodile [1126]. Crocosphaera [1661]. crosnieri [1136]. Cross [482, 2903, 818, 2477, 2411, 678]. Cross-Cultural [818]. Cross-Disciplinary [2411, 678]. Cross-Habitat [482]. Cross-Jurisdictional [2477]. Crossroads [2259]. Crown [2775]. Crown-of-Thorns [2775]. Crucial [2900]. Crude [1923]. Cruise [980]. Cruises [2036]. Cruising [1325]. Crumbling [770]. Crush [75]. Crustacea [3143, 1841, 851, 1818, 1537, 571]. Crustacean [1750, 2917, 3183, 2271]. Crustaceans [2292, 919, 202, 1487]. Cruz [2668]. Cryobanking [353]. Cryopreservation [2395, 1943]. Cryosphere [1674]. Cryptic [519, 3134, 2999, 2130, 1364, 1907]. Cryptophytes [2629]. Crystal [100]. Crystalline [974]. Cs [2538]. CSOMIO [1670]. CSSI [710]. Ctenochaetus [2323]. Ctenophore [3116]. Cu [2567, 862, 3102, 1751]. Cuba [13, 1098, 1885]. Cucumaria [1518, 2460]. Cucumber [2248, 2467, 2349, 1451, 917, 2936, 2860]. Cucumbers [3204]. Cue [2710]. Cues [2285, 27, 705]. Culling [621]. Cultivable [188]. Cultivar [859, 213, 2943]. Cultivated [1715, 930]. Cultivating [156]. Cultivation [213, 1590]. Culturable [2715]. Cultural [818, 772, 688, 817, 88]. Culture [2699, 3057, 1226, 2033, 475, 791, 2271, 2835, 2980, 3026, 1215]. Cultured [2688, 1179, 368, 3103, 291]. Cultures [1639, 46, 356, 180]. Culturing [3204]. Cumacean [2551]. cumingii [358]. Cumulative [129, 958, 988]. Cunha [1729]. Cupped [2878]. curema [268]. Current [1347, 1310, 2413, 2461, 1561, 2111, 1591, 928, 2884, 3016, 1279, 929, 2433, 312, 2548, 2097, 2549, 2722, 412, 724, 1908, 1075, 1352, 2600, 123, 443, 1348, 742, 3020, 439, 3133, 1205, 3022, 407, 1510, 185, 2356, 2290, 2318, 932, 101, 226, 691, 1644, 467, 518, 911, 1611, 2613]. Current- [2290]. Current-Mediated [123]. Currents [2294, 390, 1560, 1183, 1819, 2604, 1586, 503, 2219, 1665, 2606]. Curtain [1916]. Customary [2136]. Cut [1266]. Cutter [1646]. Cutting [629]. Cuttlebone [61]. Cuttlefish [215, 61, 1800]. Cyanide [348]. Cyanobacteria [238, 1426, 461, 2281, 2846]. Cyanobacteria-Dominated [2846]. Cyanobacterial [324, 3075, 1918]. Cyanobacterium [2131, 1661]. Cycle [2663, 946, 1311, 3115, 1945, 1533, 664, 46, 1103, 558, 304, 2904, 3054, 398, 626, 2357, 388, 1292]. Cycles [1838, 2952, 2824, 655]. Cyclic [2310]. Cyclical [2796]. Cycling [2483, 1979, 3063, 568, 1269, 388, 2946, 350, 2523]. Cyclone [861, 2288, 2766, 753]. Cyclone-Driven [753]. Cyclones [2786]. Cyclonic [1187]. Cyclopoid [2484]. Cycloprodigiosin [2940]. cygnus [1666]. Cylinder [1121]. cylindracea [3123, 2302, 1240, 2541, 1364]. cylindrica [2470]. cylindrus [1982, 1709]. Cymodocea [1240]. Cynoglossus [65]. Cyprids [2935]. cyprinellus [2841]. Cyrtophorian [1629]. Cyst [2513]. Cystoseira [3029, 2331]. Cytb [636]. cytherea [195]. Cytogenetic [2877]. Cytokines [300]. Cytometry [914, 1426, 2298]. Cytotoxic [3120]. Cytotoxicology [330].

D [651, 2873, 2591]. D40 [415]. Dade [2749]. Daily [2596, 1663, 587, 2844]. Dalatias [1731]. Dalatiidae [1646]. Dam [2823]. Damage [1115, 1031]. damicornis [2074, 633, 2568, 955]. Dana [2781]. Danafungia [231]. Danish [2722]. Dark [2230, 445, 2952, 2904]. Darkening [2512, 900]. Darkness [1168]. Darnley [1990]. Data [500, 1754, 854, 1123, 1367, 570, 1310, 3025, 1571, 2112, 2845, 1891, 873, 2955, 316, 1488, 1757, 1474, 881, 1344, 2821, 2390, 1491, 3106, 2358, 379, 1853, 257, 1843, 1407, 697, 833, 1695, 2900, 3077, 3008, 499, 587, 690, 3000, 1004, 2290, 1579, 1758, 1557, 2743, 2313, 101, 1182, 1152, 2604, 3107, 1105, 2326, 934, 2661, 87, 1494, 1836, 2545, 3067, 2522, 1866, 2069, 2062, 543, 1107, 1682, 2508, 2496, 1593, 1604, 1848, 2296, 2065, 1358, 2058, 1077, 1637, 902, 1177, 2524, 911, 362, 2125, 1774, 1623, 2278, 832, 2553, 405, 802, 2822, 3048, 2696, 2752, 1577, 1890]. Data [1325]. Data-Driven [362]. Data-Informed [257]. Data-Limited [3025, 3077, 3008, 1866, 1107, 2058]. Data-Moderate [2065]. Data-Poor [1344, 934, 2661, 1593, 2696]. Database [2401, 2751, 1775, 428]. Dataset [2592, 1584, 3101, 925, 365, 3050, 1470, 2605, 2254]. Datasets [1109, 1004, 1870]. Date [2435]. Davao [1434]. Davis [562]. Day [1986, 2731]. Daya [1744]. Days [3056]. Daytime [925]. Dead [46, 435, 2263]. Deadly [2614, 1167]. Deal [835]. Dealing [1004]. Deania [1138]. Death [2249, 1443]. Deaths [1922]. Debris [471, 364, 1773, 2482, 2689, 1636, 1587]. Decadal [2096, 2378, 516, 899, 1321, 3073]. Decade [2662, 1862, 918, 1120, 1578, 2075, 917, 3128, 3059, 174, 1010, 2055]. Decades [1771, 1617, 1536, 2241, 1273, 642, 1668, 3040]. Decapod [2292]. Decapoda [571]. Decapods [1932]. Decapterus [1163]. DeCarlo [997]. Decay [175]. Decelerate [822]. Decision [579, 492, 12, 2557, 1632, 1133, 2690, 165, 1605, 1144]. Decision-Making [579, 2557, 1632, 2690, 1144]. Decisions [2778, 2345]. Decline [206, 1982, 2429, 1656, 1223]. Declines [2754, 1041]. Declining [345, 2895, 1219]. Decommissioning [2636, 1833]. Decomposition [2329, 1651, 2049, 777]. Deconstructed [620]. Decoupling [361]. Decrease [713, 403, 983, 2826, 3040]. Decreasing [2620]. Deep [1328, 471, 1095, 240, 2151, 146, 2787, 23, 1045, 1456, 390, 3042, 1987, 1712, 872, 336, 3203, 776, 767, 2051, 1891, 2437, 1492, 903, 486, 497, 261, 1923, 531, 3115, 1782, 545, 2502, 1424, 2085, 202, 823, 710, 1236, 833, 2941, 2385, 1184, 1120, 2960, 2949, 6, 92, 528, 237, 571, 1196, 2068, 608, 803, 696, 693, 2546, 2348, 1179, 2844, 2020, 1366, 1788, 581, 423, 1050, 1314, 318, 934, 2425, 10, 2963, 3023, 1494, 2585, 1238, 1266, 1840, 2556, 2215, 2059, 412, 2158, 2118, 1717, 1016, 597, 1694, 341, 2706, 3162, 1080, 1356, 366, 158, 1785]. Deep [533, 530, 198, 3128, 1477, 822, 937, 845, 1223, 1138, 1952, 2125, 409, 2997, 1784, 178, 2740, 2814, 154, 3129, 2504, 2212, 184, 2029, 326, 970, 289, 405, 1493, 1440, 1136, 2216, 1484, 2138, 1203, 1696, 961, 2355]. Deep-Pelagic [903]. Deep-Sea [471, 240, 2151, 146, 1045, 1456, 1987, 336, 776, 2437, 1492, 486, 1923, 1782, 1424, 202, 710, 1236, 833, 2941, 1184, 1120, 2960, 6, 92, 528, 237, 571, 1196, 2068, 803, 693, 2546, 2348, 1366, 1788, 581, 1314, 934, 2425, 3023, 1238, 1266, 1840, 2556, 2059, 412, 2118, 1717, 1016, 341, 2706, 1080, 1356, 366, 158, 1785, 533, 530, 845, 2125, 178, 2740, 154, 2504, 184, 970, 289, 405, 1440, 1136, 1203, 1696, 2355]. Deep-Water [823, 10, 2963, 597, 1694, 3128, 409, 2814, 1493]. Deepest [495, 1266]. DeepWater [202, 903, 1681, 1564, 1193, 1785, 937, 1952]. Defense [565, 3130, 2719, 2907]. Defensin [2417]. Deficiencies [1579]. Deficient [944, 3120]. Deficit [2742]. Defining [139, 2327, 2638]. Deformities [2767]. Deformity [1345, 1988]. Degeneration [851]. Deglaciation [456]. Degradability [910]. Degradation [522, 1063, 972, 1827, 2858]. Degraded [1450]. Degrading [1627]. Degree [2593, 3081, 481]. Degrees [450]. DEHP [2008]. del [1387]. Delaware [2675]. Delayed [2326, 2891]. Delayed-Mode [2326]. Delays [1459]. Delivering [363]. Delphinids [379]. Delphinus [1541, 1401]. delphis [1541, 1401]. Delta [3040, 1902]. Demand [2585]. Demersal [457, 1692, 2375, 560, 1184, 1819, 2454, 561, 148]. Demersal-Inshore [2375]. Demo [1721]. Demo-Genetic [1721]. Demographic [2248, 1044, 876, 199]. Demonstrate [1111, 490]. Demonstrates [555, 710]. Demosponge [1313, 216]. Demosponges [1309, 2648, 1403, 1694]. Demospongiae [2848, 576]. Dendrogyra [1982]. Dendy [1508]. Denitrification [513]. Densely [2330, 2442]. Densities [1414, 2077, 2076]. Density [748, 704, 1804, 497, 728, 2429, 1408, 2373, 374, 1452, 198, 2597]. dentatus [1716]. Deoxygenation [104, 1161, 150, 1415, 2751, 1614, 1935, 788]. Depend [649, 753]. Dependence [2277]. Dependent [2299, 2371, 1783, 1316, 1033, 1237, 1017, 207, 1653, 1971, 2362, 68, 1756, 2008, 407]. Depends [314, 2052]. Depict [2971]. Depleted [674]. Depletion [354, 1287, 1342]. Deployed [381, 823, 43]. Deployment [312, 2054]. Depocenters [1681]. Deposition [2048, 1681, 2927, 2090, 751, 2533]. Depositions [1234]. Deposits [237, 2926]. Depredation [1559]. Deprivation [512, 1040]. Deprivation-Induced [512]. Depth [2812, 2371, 2099, 2277, 1763, 1302, 717, 2176, 968, 1002, 1967, 916, 2257, 2200, 117, 1620, 3018, 1980, 1756, 2086, 2419, 407]. Depth-Dependent [2371, 1756, 407]. Depth-Generalist [968]. Depth-Variable [1002]. Depths [44]. Depuration [2150, 1159]. Derivation [2461]. Derivatization [943]. Derived [2294, 2577, 1081, 2872, 415, 2384, 925, 209, 791, 2432, 1967, 1848, 1785, 122, 2559, 2931, 614]. Deriving [377]. Dermochelys [1104, 3132]. Deroux [1629]. Desaturases [220]. Description [767, 544, 1195, 803, 3072, 1327, 140, 1080, 1356, 1907, 1304]. Descriptions [1357]. Descriptors [1796]. Design [3042, 1928, 1531, 3008, 2176, 249, 892, 444, 1010, 3142]. Designation [1376, 668]. Designing [2327, 12, 2869, 262]. Despite [516, 1143, 368, 191, 2846]. Destination [1705]. Destroying [1751]. destructive [2041]. Desulfovibrio [509]. Detached [1388]. Detailed [2774]. Details [2434]. Detect [1991, 1792, 1698]. Detectable [1082]. Detected [1640]. Detecting [1158, 428, 2891, 2160, 1668]. Detection [1470, 2605, 1573, 203, 1783, 531, 666, 2718, 2792, 1489, 914, 1116, 2655, 2862, 3139, 2508, 2216, 1404]. Detections [823]. Detects [328]. Deter [994]. Deteriorates [691]. Deteriorating [1370]. Determinants [2249]. Determination [1093, 298, 333, 641, 1844]. Determine [2270, 916, 122, 289, 2446, 2765]. Determined [221, 317]. Determiner [1391]. Determines [576]. Determining [1779, 2463, 1827, 198, 1263]. Determinism [978]. Detoxification [222, 604, 2806, 3187]. Detrimental [368]. Detrital [939, 2276]. Detrital-Based [2276]. Detritivore [2323]. Detritus [939]. Develop [301]. Developed [274]. Developing [1961, 1474, 2619, 3161, 2569, 1866, 280, 2073]. Development [1954, 1300, 1420, 270, 1992, 453, 260, 166, 1213, 1670, 2772, 2292, 1232, 2395, 213, 746, 1158, 1285, 1483, 3152, 2792, 2659, 850, 2339, 145, 375, 2464, 1966, 698, 2615, 125, 2618, 2966, 3206, 871, 3033, 2529, 884, 504, 20, 300, 263, 733, 1578]. Developmental [1286, 21, 307]. Developments [1057, 1591]. Device [2268, 387]. Devices [3106, 99, 543, 2054]. Devoid [3077]. DFADs [543]. DHA [1720, 388]. Dhabi [643]. Di- [2008]. di-GMP [2310]. Diabetic [2707]. Diagenesis [1037, 2737, 1454]. Diagenetic [1631]. Diagnosing [607]. Diagnostic [406, 2439, 701]. Dialkyl [860]. Dialogue [1874]. Dialogues [2649]. Diameter [2856]. Diapause [1802, 3092]. Diapausing [160]. diaphana [1225, 98]. Diastereomers [330]. Diatom [294, 2329, 1639, 862, 2550, 596, 483, 305, 140, 1443, 55, 17, 191, 715, 1709].

Diatoms [1978, 2480, 1828, 2629, 2781, 1767, 2493]. Diazotroph [987, 238, 2983, 2432]. Diazotroph-Derived [2432]. Diazotrophs [1089]. Diazotrophy [737]. Dicentrarchus [1067, 2239]. Dictates [3084]. Dictyota [1818]. Dictyotales [1894]. Did [1906]. Didemnum [303, 2776]. Diego [2193]. Diel [2968, 2325, 1678, 2051, 2807, 107, 2693, 654, 2014, 1725, 2381, 2946, 535]. Diel-Cycling [2946]. Diet [1991, 939, 3075, 2484, 2927, 1600, 524, 2018, 465, 3061, 1635, 1101, 740, 355, 869, 1482]. Diet-Induced [2927]. Dietary [2586, 2253, 2589, 2595, 2361, 2537, 1299, 2967, 2880, 3065, 2098, 2090, 2676, 2777, 2574, 2304, 1199, 3160, 65, 905, 300, 263]. Diets [1372, 1716, 299, 3205, 1920, 837, 1819, 161, 2893, 2380, 1515]. Differ [1372, 3081, 2432]. Difference [2845, 2003, 1792, 2967]. Differences [472, 2614, 2195, 2116, 2365, 1793, 196, 1535, 2713, 2871, 1540, 1056, 3165]. Different [2663, 2802, 776, 1189, 2815, 1345, 1036, 2333, 2204, 1289, 1073, 2339, 3175, 943, 1503, 2011, 374, 1993, 594, 2759, 2177, 517, 220, 2888, 1028, 2258]. Differential [2074, 2786, 920, 2115, 1608, 327]. Differentially [951, 330, 2759]. Differentiation [3153, 1132, 2917, 996, 2114]. Differently [1258]. Differing [18, 877, 2340, 989]. Diffusive [1058, 2405]. Diflubenzuron [2450]. Digestibility [2586]. Digestion [735, 2956]. Digestive [2586, 2494, 1362, 2904, 2090]. Digital [2736, 406, 755, 627]. digitata [558]. digitifera [2671]. Dihydroeponemycin [1863]. Diking [2588]. Dilemma [538, 28]. Dimension [2649]. Dimensional [1291, 1654, 933, 1796, 1854]. Dimensionality [992]. Dimensions [2850, 1544, 1784]. Diminish [959]. Dimorphism [2416]. Dinitrogen [513]. Dinoflagellate [1940, 1281, 1790, 2684, 2889, 2716, 2683, 2513, 2080, 1653, 3037, 2857, 600, 2835, 671, 785]. Dinoflagellates [1828, 3201, 2587, 2806, 2563, 1805]. Dinophysistoxin [1159]. Dinophysistoxin-1 [1159]. Diodes [2339]. Dioxide [2172, 1317, 1, 790]. Dipeptide [2423]. Diplodus [489, 2302, 2540]. Diplomacy [193, 2075]. Direct [3161]. Directional [1968]. Directions [1588, 600, 1148, 3022]. Directive [2517, 1432, 1430]. Directly [2524]. Disappeared [1229, 752]. Disappearing [2919, 652]. Disaster [1839]. Discard [1766, 1487]. Discarded [1571]. Discerning [449, 1944]. Discharge [2286, 675, 2755, 2876, 2404, 2053, 1794]. Discharges [1201]. Disciplinary [2903, 2411, 678]. Disciplines [2278]. Disconnected [1454]. Discourses [278]. Discoveries [2078]. Discovery [2280, 47, 2904, 1312, 1357, 211]. Discrete [1329, 822, 2741]. Discrimination [2120, 1105, 2867]. discus [1943, 2956, 2084, 3053]. Discussions [2434]. Disease [406, 2589, 2322, 1769, 1167, 1875, 3091, 1237, 1031, 3061, 2319, 1906, 1216, 2180, 684, 265, 2250, 2169, 2157, 1148, 2107, 2509, 2673, 2322, 2614, 2317, 2666, 2687, 1718, 1043, 2616, 2749, 2311, 1702, 2139, 2953]. Diseased [195]. Diseases [2614]. Disentangling [1463, 2972, 2196, 765]. Disequilibria [2435, 2527, 2685, 2449]. Disequilibrium [2340, 2523]. Disjunct [2042].

Disko [1527]. Disparate [130, 2337, 1904]. Disparities [745]. Disparity [3035]. Dispersal [1104, 2482, 2775, 2691, 403, 670, 661, 241, 1088, 3178, 2545, 2331, 533, 1554, 2015, 1468]. Dispersant [1923]. Dispersants [1827]. Dispersed [1232, 705]. Dispersion [588, 2785, 119, 2832, 1277]. Displacement [510, 2295]. Displays [1947]. Dispute [1180]. Disputes [1180]. Disrupt [1000, 703, 898]. Disruption [1188, 1043]. Disruptors [573]. Disrupts [3181, 952]. Dissolution [107, 2693, 149, 304, 1442]. Dissolvable [2998]. Dissolved [308, 229, 329, 570, 1456, 273, 1093, 1811, 1250, 815, 1979, 2942, 1789, 462, 1183, 1254, 3084, 1830, 1054, 580, 94, 2342, 3192, 1673, 2643, 441, 1826, 2403, 3140, 2890, 568, 2708, 1502, 3151, 910, 2104, 1151, 1140, 2227, 530, 973, 443, 1467, 2476, 350, 1082, 182]. Dissostichus [2108]. Distal [2579]. Distance [1301, 1099, 2236]. Distance-Based [2236]. Distances [2071]. Distant [861, 648]. Distant-Source [648]. Distinct [1853, 1926, 2978, 1023, 3164, 3079]. Distinction [1087]. Distinctive [812]. Distributed [1100, 454]. Distribution [2507, 229, 1009, 146, 390, 2299, 672, 234, 2774, 1329, 3082, 2805, 2249, 389, 3104, 3105, 1024, 1804, 2437, 945, 3125, 1817, 497, 1646, 2462, 507, 124, 2724, 2248, 1424, 2356, 2390, 202, 1001, 2513, 3024, 833, 1746, 2186, 1550, 1995, 769, 2165, 1163, 604, 2960, 2949, 534, 151, 3056, 1965, 1723, 1108, 1164, 571, 717, 850, 2756, 548, 292, 2979, 1088, 969, 1714, 3107, 2366, 503, 1064, 875, 2425, 910, 2721, 208, 1899, 944, 543, 915, 1858, 2145, 2450, 2158, 1368, 2914, 2234, 2373, 1506, 1821, 1944, 2609, 37, 2200, 1128, 325, 103, 493, 5, 2424]. Distribution [1041, 1780, 315, 2741, 2184, 1556, 3071, 2023, 2293, 405, 1440, 3093, 2860, 1860, 400, 210]. Distributional [919]. Distributions [1712, 704, 1849, 251, 2383, 2095, 379, 2887, 1324, 691, 724, 2225, 58, 1034, 3165]. District [418]. Disturbance [218, 2328, 48, 1622, 2323, 455, 559, 2860, 2773]. Disturbance-Mediated [2323]. Disturbances [841, 1726]. Disturbing [455]. Diurnal [3181, 1151, 2846]. Dive [1663, 1552]. Diver [1006, 1801]. Diver-Portable [1006]. Divergence [656, 2277]. Divergent [1359]. Divers [1552, 374]. Diverse [1470, 2605, 1425, 1713, 2848, 29, 1227, 1338]. diversicolor [2204, 2127, 884]. Diversification [1955]. Diversity [3014, 243, 1009, 2377, 987, 3042, 1987, 2684, 1224, 1936, 3124, 2108, 14, 2322, 1319, 1022, 1646, 2272, 1225, 1001, 3089, 1853, 744, 812, 83, 2178, 3134, 2999, 2715, 1184, 186, 1851, 13, 345, 24, 2949, 932, 1635, 693, 2056, 256, 2554, 1634, 681, 1959, 1171, 1735, 2373, 1627, 1080, 1356, 2057, 187, 125, 1336, 5, 1154, 582, 382, 171, 2798, 1856, 1265, 3149, 199, 1756, 191, 2007, 2391, 1837, 1440, 22, 574, 233, 2015, 2757, 291, 1970, 1696, 1284, 2424]. Divide [1416, 858]. Diving [697, 2368, 1206, 2470]. djiboutiensis [1422]. DMS [1209]. DMSP [991]. DNA [1808, 1262, 2108, 48, 1783, 2272, 57, 967, 2789, 962, 760, 2417, 75, 1742, 1887, 1200, 2967, 2183, 2348, 2725, 3098, 2776, 3023, 920, 1113, 938, 2143, 1336, 3010, 328, 307, 1265, 2572, 327, 197, 618, 1907, 2493]. Do [2427, 3197, 1561, 1437, 730, 2333, 1888, 268, 2380, 1829, 1180, 1074, 2055, 3165]. DOC [568, 2020]. Documenting [2358, 923, 3050]. DOE [1673]. Does [765, 509, 2126, 1624, 580, 3081, 2045, 1302, 829, 954, 2304, 3155]. Dogfish [1138]. Dolabrifera [3075]. Doliolid [1628]. Dolphin [225, 1463, 2985, 1185, 2974, 369, 1332, 391, 855, 1915, 2018, 850, 1392, 875, 2635, 2555, 1343, 2984, 1370, 1563, 1380, 876, 1688, 2653, 1558, 762]. Dolphin-Swim [762]. Dolphin-Watching [1563]. Dolphins [1534, 1541, 801, 1464, 1320, 1633, 1401, 605, 1437, 2079, 1181, 385, 1550, 1112, 3153, 3170, 1969, 2205, 1490, 1551, 1090, 921, 455, 1114, 1546, 2560, 1548, 1563, 1829, 2854, 1621, 1619, 1556, 1450, 1461]. DOM [1811, 900, 443]. Domestic [371]. Domesticated [662]. Dominance [1081, 2329, 736, 409]. Dominant [778, 2629, 1389, 2744, 3129]. Dominate [776, 1035]. Dominated [3209, 62, 1875, 1547, 1589, 956, 1962, 1736, 2846]. Dominates [13]. Dominican [2133, 2159]. Domoic [580, 1409]. Domperidone [1008, 1354]. Donation [1049]. Done [108, 2055]. Dongsha [2693, 613]. Doppler [2604, 2308]. Dorian [753]. Doryteuthis [1329]. Dose [2362, 3120, 2538, 2008]. Dose-Dependent [2362, 2008]. Dose-Rate [3120]. Dosidicus [2724, 744]. Dover [1841]. Down [424, 83, 1287]. Downscaled [1510]. Downscaling [1582, 1902, 964, 1839]. Downstream [1861]. Downwelling [2046, 1935]. Downwelling/Upwelling [1935]. Draft [1451]. Dragonfishes [204]. Dragons [767, 2178]. Dramatic [206, 2229]. draw [649]. Dredge [2706]. Dredged [2990, 1603]. Dredging [1051, 118, 3068]. Dreissena [2409]. Dried [3191]. Drift [1560, 2112, 2026, 608, 2295, 407, 2482]. Drifters [1277]. Drifting [2081, 543]. Drilling [2345]. Drive [864, 155, 793, 782, 1809, 1435, 573, 2154, 874, 1275, 1214, 2913, 1661]. Driven [676, 2286, 19, 1595, 1488, 3024, 2863, 728, 7, 1346, 2430, 2598, 726, 999, 1141, 736, 1064, 2556, 899, 1321, 67, 2754, 838, 753, 2734, 362, 2723, 2814, 1736, 301, 2606, 546, 3073, 407]. Driver [2256, 2119, 119]. Drivers [2431, 3014, 809, 229, 172, 1706, 2185, 146, 485, 1598, 713, 669, 856, 124, 2095, 1290, 1237, 769, 1471, 528, 2031, 2735, 909, 1386, 495, 2721, 897, 2521, 2410, 1178, 729, 1656, 103, 2017, 743, 1144]. Drivers/Pressures [2031]. Drives [621, 1002, 553, 1206]. Driving [2096, 2091, 1224, 670]. Drone [385, 49]. Drones [1913]. Dronning [1317]. Drop [3122, 2194, 1696]. Drop-Induced [3122]. Droplet [1728, 396]. Drowned [1074]. Drowning [159]. Dual [2791]. Duck [2780]. Due [2269, 682, 2229, 2218, 2242, 1973, 2142, 1912, 2529, 1931, 2620]. Dump [2321]. Dumping [2990]. Dune [2930, 2819]. Dungeness [101, 503, 205]. Duplications [26]. Duration [2777]. During [1940, 2535, 971, 1807, 1573, 1323, 2402, 1209, 2150, 2091, 872, 663, 2592, 990, 1256, 451, 2942, 659, 1789, 569, 456, 1703, 1939, 2902, 3202, 2242, 2697, 746, 16, 36, 2656, 1730, 534, 714, 2755, 2288, 1299, 664, 2756, 46, 1855, 1535, 1522, 1733, 1882, 2642, 2261, 2036, 977, 1406, 904, 2828, 2876, 1858, 2809, 1898, 2030, 2766, 598, 357, 2936, 1983, 37, 780, 2040, 822, 1656, 2750, 3121, 1861, 2854, 1172, 426, 2084, 1153, 617, 1960, 2701, 2814, 1931, 1369, 2929, 2826, 1860, 3040, 2316, 3193, 1439, 300, 1590, 1436, 2438, 878]. Durusdinium [3159]. Dusky [391, 921]. Dust [149]. Dutch [448, 3, 1069, 2930, 2649].

Dwarf [1329]. Dynamic [1975, 1690, 1146, 1349, 3202, 932, 478, 2945, 2642, 677, 642, 1691, 743, 2657, 2579, 361, 3004, 1835]. Dynamical [1582, 964]. Dynamically [1510]. Dynamics [1640, 2509, 2783, 2044, 946, 1262, 2150, 383, 3209, 2974, 1030, 3025, 1469, 2060, 185, 1189, 2815, 45, 591, 903, 48, 1022, 1639, 1160, 456, 1715, 1793, 1181, 480, 2378, 2196, 343, 97, 2863, 1568, 2256, 3057, 1066, 2643, 2363, 235, 1625, 2873, 602, 3177, 1396, 850, 932, 226, 2958, 2032, 1586, 3097, 2355, 1366, 2581, 1220, 1216, 2167, 2495, 956, 1871, 2241, 2307, 3137, 398, 2722, 756, 1850, 2146, 3198, 1479, 2832, 1952, 2824, 2275, 1609, 17, 964, 1934, 1495, 3174, 254, 1702, 843, 1740, 764, 110, 1124, 3048, 1205, 182, 1134, 2838, 300, 3150, 1418]. Dysbiosis [1115].

E. [2595, 2124]. Eagle [3092]. Early [3014, 424, 794, 1037, 1385, 659, 1700, 1749, 1928, 1719, 2544, 101, 1182, 1168, 2842, 1733, 573, 2581, 2407, 2582, 1454, 21, 307, 392, 631, 1218, 301, 2888]. Early-Career [1719, 2407]. Earth [3119, 800, 2150, 1838, 33, 1723, 1594, 897, 526, 2508, 737, 1390]. Earthen [662]. Easier [902]. East [1281, 1566, 815, 2378, 1547, 375, 2828, 1075, 2780, 582, 742, 788, 574, 3064, 2532, 619, 2847, 2111, 97, 3095, 1234, 1953, 3177, 1525, 1296, 2598, 3145, 321, 1586, 834, 200, 3167, 467, 1080, 1356, 410, 2998, 1206, 3092, 1657, 1342]. East-Greenland [1206]. East-Siberian [2828]. East/Japan [3177, 1525]. Easter [2382, 2992]. Eastern [1640, 1526, 3108, 157, 867, 1841, 967, 2137, 1939, 505, 560, 700, 460, 2685, 413, 2245, 327, 3064, 2838, 1223, 2507, 1328, 1630, 1849, 3062, 1401, 659, 206, 3181, 1636, 2244, 2724, 2619, 2884, 2714, 2333, 2063, 235, 1126, 2890, 3006, 1626, 2201, 944, 78, 1679, 597, 1694, 2582, 1307, 1191, 227, 2040, 103, 667, 3088, 1572, 1613, 1917, 1482]. Eat [868, 375]. Eating [2321]. Ecdysone [2866]. Echinoderm [1386]. Echinodermata [2708, 901]. Echinoderms [1446]. Echinoidea [2708]. Echo [2290]. Echosounder [1960]. Ecklonia [2943]. Eco [3209, 736, 2106, 2237]. Eco-Engineering [2237]. Eco-Evolutionary [3209, 736, 2106]. Ecoengineering [2516]. Ecological [1123, 3119, 1032, 1433, 3191, 577, 1867, 1978, 771, 2689, 1270, 2847, 936, 979, 610, 1779, 1474, 2274, 1600, 2795, 928, 1087, 1746, 41, 2718, 2561, 2426, 1697, 572, 1909, 2563, 478, 1632, 2827, 896, 1788, 1667, 2236, 2302, 2025, 1588, 8, 84, 1866, 344, 817, 2558, 1942, 1905, 2282, 1230, 1833, 15, 2690, 1985, 2992, 1609, 1784, 3088, 687, 2740, 194, 2608, 3036, 2422, 2389, 220, 1124, 3068]. Ecologically [2131, 469, 55, 490, 2357, 476]. Ecologically-Relevant [55]. Ecology [346, 2479, 1630, 1486, 464, 1208, 1841, 531, 792, 2187, 537, 2902, 1555, 1568, 2186, 3070, 891, 1693, 1909, 1103, 204, 600, 533, 1944, 2989, 2143, 549, 1558, 2083, 190, 444, 2086, 2757, 2996]. Econometric [3026]. Economic [3180, 79, 2182, 363, 127, 727, 1857, 15, 334, 2147, 2444, 889, 3142]. Economically [2712]. Economy [371, 60, 2638, 688, 2722, 1910]. Ecopath [2404, 400]. Ecoregion [2992]. Ecoregions [268]. Ecosim [2404].

Ecosystem [809, 1472, 926, 805, 1350, 1209, 1445, 2202, 1337, 787, 464, 794, 185, 2601, 1322, 1270, 615, 903, 2346, 2933, 2932, 2132, 958, 480, 2795, 1441, 555, 257, 72, 1394, 1562, 683, 1066, 918, 2768, 1375, 1279, 2629, 754, 1697, 528, 1326, 335, 101, 430, 226, 1391, 562, 1521, 684, 1503, 271, 1238, 173, 688, 1171, 2809, 1059, 597, 2210, 817, 2206, 559, 679, 1705, 540, 2143, 1013, 454, 1288, 2447, 642, 1854, 2276, 3194, 1605, 1306, 635, 743, 1952, 1668, 1241, 1544, 1263, 1201, 81, 2600, 1611, 2335, 992, 190, 1774, 1014, 1495, 2674, 174, 444, 764]. Ecosystem [1062, 88]. Ecosystem-Based [805, 2795, 1279, 754, 335, 2600, 1611, 444]. Ecosystemic [16]. Ecosystems [885, 1499, 971, 34, 800, 577, 887, 320, 2972, 234, 936, 979, 1680, 2763, 1253, 1877, 1599, 2502, 1645, 109, 814, 2378, 555, 710, 2544, 186, 2901, 754, 1693, 1396, 1528, 896, 1727, 1798, 2585, 1775, 897, 2059, 425, 2352, 344, 341, 40, 198, 2199, 1576, 989, 2500, 826, 174, 1597, 2389, 350, 56, 400]. Ecotoxicological [225, 331]. EcoTroph [612, 1298]. Ecotypes [1044]. Ectocarpus [2918]. Ectopic [2271]. Ecuador [225, 855, 2152]. Eddies [556, 2724, 782, 769, 810, 90, 474, 546, 1543]. Eddy [1891, 2692, 1872, 221, 1187, 2630, 2004, 2977, 899, 1321, 2766, 595, 1791, 2822, 2216, 3004]. Eddy-Induced [3004]. Eddy-Mean [1791]. Eddy-Resolving [2630]. edeni [438]. Edge [2698, 813, 2604, 491, 1671]. Edible [2864, 2988, 504]. Editing [1226]. Editorial [3163, 34, 466, 3119, 789, 66, 3060, 264, 85, 250, 3209, 102, 3203, 2728, 2972, 150, 1532, 557, 2351, 739, 480, 3156, 1353, 2993, 647, 2078, 1315, 2965, 3185, 3131, 2652, 3195, 1238, 138, 2162, 2885, 1271, 2850, 3167, 2210, 3186, 679, 40, 2411, 1355, 3113, 1295, 217, 454, 3169, 1288, 1352, 3194, 1952, 1816, 306, 1348, 2279, 2389]. eDNA [48, 1783, 654, 548, 784]. Education [2157, 1173]. edule [2495]. edulis [939, 187, 3121, 1258]. Edward [1156]. Edwards [2658]. Eel [2280, 1362, 47, 905]. Eelgrass [3091, 2643, 1435, 2913]. Effect [1328, 1815, 314, 1761, 2882, 1911, 2944, 1771, 2692, 1397, 639, 2382, 1068, 2196, 1812, 1301, 2982, 3188, 797, 1248, 120, 930, 608, 948, 3146, 1988, 943, 1402, 3054, 2101, 1171, 305, 1797, 37, 249, 2987, 1278, 1878, 521, 98, 1239, 2357, 1197, 2898, 1657, 2393, 2633]. Effective [102, 2327, 2394, 875, 68, 1775, 1744, 384, 410, 995, 2584, 2628]. Effectively [685, 1742, 1866]. Effectiveness [2454, 2631, 1411, 2250, 1156, 894, 1427, 2505]. Effects [408, 2214, 1630, 1737, 2586, 2299, 1921, 1975, 1991, 1281, 1176, 1977, 156, 2255, 615, 2905, 2595, 2564, 3110, 2322, 1401, 129, 251, 1388, 1622, 828, 1933, 1718, 1838, 3205, 1596, 2079, 3106, 2121, 1793, 96, 1430, 958, 1345, 216, 1021, 1801, 3091, 1149, 1663, 1963, 201, 1099, 1943, 604, 465, 2415, 3061, 1625, 2843, 974, 1528, 1447, 1141, 475, 2172, 3185, 633, 2362, 1733, 3065, 2433, 896, 348, 2223, 2635, 1398, 2425, 2236, 422, 2024, 1266, 900, 1240, 2098, 387, 2011, 950, 2781, 1608, 954, 2090, 1432, 1380, 740, 988, 2676, 3027, 58, 841, 2921, 1604, 1067, 2615]. Effects [3128, 15, 307, 2053, 1384, 2565, 529, 671, 2669, 1925, 869, 1621, 1231, 1218, 2608, 3174, 3160, 3078, 2588, 2597, 806, 2008, 2476, 1247, 3148, 2680, 2870, 20, 2860, 905, 3053, 2501, 1864, 2773, 1618]. Efficacy [2089, 1255, 1781, 2634, 2947, 539]. Efficiency [1523, 1519, 461, 1067, 1162].

Efficient [2400, 1797, 637, 2628, 2207]. Effluent [2577]. Effort [1754, 371, 1466, 2184, 93, 2602, 783]. Efforts [2202, 2900]. Effrenium [2868]. Egg [2564, 2048, 133, 353, 28, 2497, 2496, 3174]. Eggs [1163]. Egypt [2688, 1626]. Eiconaxius [571]. Eight [266]. Eighty [2293]. Eiseman [2875]. Elasmobranch [610, 2395, 673, 1579, 2833, 3018, 1980]. Elasmobranchii [3098]. Elasmobranchs [2454, 1419, 2336]. Elderfield [217]. Electric [729]. Electrical [1176]. Electrode [298]. Electron [569]. Electronic [1571]. Electrospray [1684]. Element [2150, 1684, 526, 698, 2198]. Element-Selective [1684]. Elemental [61, 2562, 1335, 422, 55]. Elementary [2247]. Elements [33, 2737]. Elephant [839, 1122]. Eleutheronema [2925]. Elevated [1737, 1639, 1828, 2129, 51, 520, 1551, 1280, 425, 2038, 447, 1725, 1493, 2603]. Elevating [299]. Elevation [2453, 2448]. Elicit [705]. Elongases [220]. elongata [1542]. Elongate [1542]. Elopomorph [271]. Elucidate [261]. Elucidates [126]. Elucidation [2782]. Elusive [1909, 1698]. Elutriates [331]. Embayment [1081, 1372, 2046, 2014, 2957]. Emboli [1437]. Embolic [1437]. Embolism [544]. Embrace [2282]. Embryo [2339]. Embryonic [2048, 3152, 2615, 3092]. Embryos [353, 20]. Emergence [2673, 2244, 2385, 657, 437, 759]. Emerging [1347, 2570, 2973, 2993, 3016, 2429, 2210, 2278, 1747]. Emiliania [846, 806, 3165]. Emiliania/Gephyrocapsa [3165]. Emirates [643, 2670, 2171]. Emissions [2641, 2770, 3135, 2049, 1059, 1, 790, 1427]. Emitting [2339]. EMODnet [1109]. Emphasis [1894, 93]. Emphasizes [2853]. Emphasizing [1498]. Empirical [1954, 1300, 932, 701, 2397]. Empirically [614]. Empowering [2695]. Enable [857, 385, 1578]. Enables [698, 2940]. Enabling [2702, 2123, 410, 1884]. Enclosed [1157, 2827, 3043]. Encoding [1567]. Encounter [257, 2984]. Encourage [3013]. Encouraging [931, 1361]. Encrusting [1811]. End [1377, 560, 2318, 2584, 625, 1618]. End-to-End [1618]. End-User [2584]. Endangered [1885, 1104, 1690, 228, 1584, 1099, 850, 1392, 47, 3098, 1370, 1453, 1829, 1922, 1558, 1459, 1698, 3018, 1980, 1542, 1284, 1915]. Ende [2136]. Endemic [1863, 2331, 2107, 3092]. Endocrine [1920, 573]. Endogenous [185, 2710, 2759]. Endolithic [111, 1247]. Endolytic [522]. Endophytes [1769, 703]. Endophytic [2623, 135, 2888]. Endoplasmic [2927, 465]. Endosymbiont [1993, 2953]. Endosymbionts [2971]. Endosymbiotic [3124]. Endotoxin [1096, 255]. Endurance [43]. Enemy [1722]. Energetic [2333, 2380, 1543, 1231]. Energetics [1791]. Energy [1233, 2155, 611, 1022, 1838, 1669, 728, 1659, 2384, 693, 968, 2885, 1648, 2936, 2987, 3033, 2907, 1467, 517, 3053, 535]. Engagement [1488, 624, 908, 1605]. Engaging [685, 511]. Engine [1788, 2508]. Engineered [862]. Engineering [1928, 2539, 2626, 1314, 1883, 2237]. Engineers [454]. Engraulis [706]. Enhalus [427]. Enhance [509, 2566, 646, 730, 2733, 912, 1727, 69, 954, 1003, 2553, 1661]. Enhanced [19, 104, 1146, 1215, 2994, 2456, 2013, 1111, 1297, 596, 1222, 2237, 74].

Enhancement [2693, 1345, 384, 197, 1475, 2898]. Enhances [1420, 2329, 2950, 2199]. Enhancing [243, 2251, 925, 322, 2054]. Enhydra [515, 1689, 437, 684, 1409]. Enigmatic [10]. Enriched [1345, 2204, 1515]. Enrichment [1068, 797, 2045, 633, 2020, 626, 1307, 1227, 2980, 889]. Ensemble [651, 1510, 2805, 2735, 227, 1170, 2332, 1210]. Ensembles [1705]. Ensis [118]. ENSO [3054, 2686]. Ensuring [1901]. Entanglement [1360, 1584]. Entanglements [3126]. Entering [1719, 2303]. Enteritis [2388, 300]. Enterocytospora [2792]. Enteromorph [880]. Entrapment [605]. Environment [240, 2883, 364, 404, 813, 39, 818, 1319, 57, 1998, 2576, 2667, 2697, 1155, 1523, 2901, 2768, 3016, 2052, 2031, 2264, 2149, 2115, 2954, 1999, 2450, 1821, 2729, 3030, 1153, 142, 1581, 2951, 1970, 1835, 1680]. Environmental [3014, 1892, 702, 1808, 2511, 1104, 2330, 2784, 485, 864, 2299, 2074, 1975, 2919, 1787, 1994, 2113, 336, 381, 2990, 2972, 2695, 2775, 258, 48, 1817, 278, 828, 552, 2351, 1715, 1000, 196, 3202, 2513, 1267, 1033, 1021, 2455, 1158, 1394, 36, 2475, 1217, 143, 75, 2813, 2415, 2318, 2460, 1606, 528, 349, 1632, 1684, 3044, 2546, 2348, 909, 2776, 297, 1063, 3138, 1386, 2425, 3023, 1406, 2721, 3195, 684, 119, 691, 2634, 205, 972, 2536, 724, 2940, 2345, 127, 2210, 2496, 938, 1361, 2143, 1752, 103, 2690, 3010, 1214, 328, 743, 2849, 539]. Environmental [2073, 2572, 3078, 3207, 2050, 417, 3064, 1577, 761, 2654]. Environmentally [105, 573, 1064, 327, 291]. Environmentally-Friendly [105]. Environmentally-Induced [327]. Environments [653, 3084, 2983, 2455, 2993, 957, 2315, 1887, 3145, 1238, 897, 2580, 2690, 2452, 319, 2759]. Enzymatic [1646]. Enzyme [735, 2090, 2138]. Enzymes [3157, 604, 2904, 2761, 1567]. enzymogenes [1972]. Eolid [2439]. Epibenthic [3069, 2718, 2849, 671]. Epibiotic [2781]. Epidemio [1167]. Epidemio-Hydrodynamic [1167]. Epidemiology [265]. Epidermal [236]. Epifauna [256, 3118, 1484]. Epifaunal [986, 1696]. Epigenetic [700, 573, 938, 386]. Epigenetics [2351]. Epinephelus [2595, 2362, 3146, 2124, 2870, 2388]. Epipelagic [2968, 1410, 3076]. Epipelic [535]. Episodes [2712]. Episodic [3015]. Epistemology [1383]. Epithelial [1971]. EPS [396]. Equal [910]. Equalizing [1224]. Equations [108]. Equator [491]. Equatorial [2724, 1254, 1543, 2416]. Equatorially [2042]. Equilibria [675]. Equilibrium [421, 3096]. Equitable [1852, 1416, 2702, 2123]. Equitably [2144]. Equity [2640]. Equivocal [3128]. Era [1867, 451, 2932, 1900, 2911, 820]. ERA5 [3028]. Eradicate [2588]. Eretmochelys [909]. Ericaria [3029, 2331]. Eriocheir [3189, 2880, 3166, 263]. Erodes [3034]. Eroding [2115]. Erosion [2651, 2763, 1039, 2344, 1715, 2994, 2463, 3015, 2819, 2049, 2332]. Erratum [2739, 2605]. Error [2856, 87]. Erupted [131]. Eruption [273, 343, 97]. eruptive [2438]. Erythrocytes [3120]. erythrogramma [1149]. Escaping [2239]. Escherichia [1001]. Eschrichtius [1148]. ESM [606]. Esophagus [1362]. Espírito [1284]. Essential [939, 2421, 228, 2204, 2959, 2297, 2398, 2295, 2436, 1884]. Essentials [1677]. EST [1542]. EST-SSR [1542]. Established [1270]. Establishing [2852].

Establishment [3161, 1341, 1242, 98]. Estimate [1754, 391, 1491, 722, 1929, 1007, 2527, 3126, 256, 3159]. Estimated [2968, 591, 2545]. Estimates [203, 99, 2717, 921, 2049, 290, 374, 1452, 762, 210]. Estimating [2711, 44, 76, 961, 1070, 116, 418, 505, 2836, 49, 734, 2604, 2964, 2219, 2184]. Estimation [3052, 734, 1169, 1419, 2856, 2610, 2252, 2305, 2653, 906, 1501, 609, 1303]. Estimations [1489, 2542]. Estrogen [263]. Estrogens [2854]. Estuaries [1372, 2358, 2066, 130, 1904, 2045, 2887, 2394, 2679, 2521, 2296, 3078, 2592]. Estuarine [682, 1081, 1977, 2510, 1181, 1097, 2196, 2498, 1730, 798, 2403, 1141, 1152, 1536, 449, 2053, 1497, 1672, 1124]. Estuarine-Resident [1181]. Estuary [225, 163, 1071, 324, 1643, 2255, 1716, 1789, 2590, 2932, 3017, 272, 2795, 2837, 1547, 75, 527, 2415, 1392, 1570, 2679, 1490, 2000, 526, 1603, 2672, 865, 514, 2824, 2657, 2442, 1467, 2675, 1672, 2881, 2757, 1021, 1192, 2715, 2813, 2403, 1536, 2056, 2433, 2366, 3151, 3078, 3207, 2287, 2588, 609, 1134, 338, 361, 1342]. Estuary-Coastal-Open [2000]. Estuary-Wide [865]. Ethanol [1783]. Ethical [2816]. Ethics [2695, 2654, 2694]. Ethology [1450]. Ethylhexyl [2008]. Etmopterus [1840]. EU [561]. Euduboscquella [2716]. Euherdmania [1863]. Eukaryotes [2007]. Eukaryotic [2183, 1113]. Eukrohnia [498]. Eumetopias [1052]. Euphausia [1452, 2177, 2625]. Euphausiacea [202]. Euphausiid [2996]. Euphotic [866]. Euplotes [1323, 1655, 1324]. Euplotida [1259, 1324]. EuroGOOS [232]. Europe [1260, 1347, 1350, 286, 141, 1591, 1393, 981, 982, 380]. European [2506, 2688, 2678, 232, 1549, 2872, 3136, 2359, 282, 2462, 1430, 1853, 2469, 1109, 1529, 960, 1362, 428, 246, 2548, 1487, 1432, 1567, 2676, 1067, 2669]. Eurosion [428]. Euryhaline [1720]. Eutrophic [383, 1716, 3156, 445, 2066, 1570, 2672, 1825, 1777]. Eutrophication [34, 1161, 1430, 2607, 2283, 1589, 772, 954, 849, 1019, 372]. Evaluate [1975, 41, 825]. Evaluated [2990, 3100]. Evaluating [720, 2947, 3128, 1777, 768, 2505]. Evaluation [1954, 1300, 1761, 1787, 2805, 2589, 2461, 1083, 2656, 2700, 1615, 2384, 1887, 3107, 791, 1113, 2634, 200, 662, 2734, 1691, 606, 1861, 539, 1925, 985, 3141]. Event [1489, 1188, 1855, 1216, 68, 288, 409, 1287]. Events [1642, 484, 152, 389, 2763, 1522, 2261, 1551, 1836, 689, 2580, 2331, 2519, 122, 822, 2199, 2855, 1381, 964, 2926]. Ever [2647]. Everything [29]. Evidence [240, 314, 577, 2847, 591, 1208, 1757, 3135, 492, 419, 2999, 12, 770, 696, 2020, 1026, 2355, 1366, 1836, 2320, 2908, 780, 753, 3030, 1587, 2017, 2780, 2194, 154, 2422, 2752, 3068, 115]. Evidence-Based [492]. Evidences [2897]. Evolution [644, 1071, 2991, 563, 2230, 1872, 891, 2170, 1291, 428, 2896, 1175, 1990, 637, 2103, 477, 1776, 1207, 1890]. Evolutionary [3209, 1955, 2515, 2576, 1655, 736, 2106, 1629, 1174, 110, 2752]. Evolving [2209, 2132, 1578]. Ex [1709, 1008, 1354, 745, 2169, 2138]. Exacerbated [2723]. Exact [2764]. Exaiptasia [1225, 98]. Examination [73, 1969].

Examine [101]. Examining [2299, 503, 442, 307, 1241]. Example [1915, 2638, 2546, 2473, 1382, 136]. Examples [947, 3203]. Exceed [2540]. Excellent [475]. Exceptional [812, 2059]. Exceptionally [2213]. Excess [388]. Excessive [2927]. Exchange [2698, 2916, 2566, 1997, 2876, 2769]. Exchanges [2451, 1693, 1586]. Excitation [147]. Excreta [1020]. Excreted [95]. Excretion [2676]. Exercise [1883]. Exhibit [175, 261, 2754]. Exhibiting [2670]. Exhibits [996]. Existence [1131]. Existing [1669]. Exit [1643]. Exogenous [3158]. Exolytic [522]. Exopalaemon [3021]. Exopolymers [1827]. Exopolysaccharide [1972]. Exotic [3019]. Expand [2760]. Expanded [1063, 2065]. Expanding [16, 1428, 1325]. Expands [2015, 1879]. Expansion [157, 216, 2571, 658, 977, 3019, 1805]. Expectancy [3174]. Expected [79, 1265]. Expedite [550]. Expedition [2099, 16]. Expeditions [6, 92]. Expenses [535]. Experience [688, 1443]. Experienced [655]. Experiences [731]. Experiment [2150, 1845, 42, 1872, 572, 1072, 900, 1859, 2706, 372]. Experimental [258, 429, 2994, 2170, 1520, 2180, 1151, 740, 559, 2833, 3155, 3085]. Experimentally [1923, 2291]. Experiments [2662, 990, 545, 83, 3188, 2893]. Expert [1848]. Expert-Derived [1848]. Explain [2871, 1959, 325]. Explained [1422]. Explaining [2027]. Explains [868, 736]. Explanation [1936]. Explanations [1739]. Exploitation [2883, 821, 538]. Exploitation-Exploration [538]. Exploited [416, 725, 742, 831, 844, 2696]. Exploration [1433, 538]. Exploratory [1846, 1840]. Explore [928, 2455, 237, 1576, 743]. Exploring [1873, 3191, 2492, 1813, 48, 508, 1319, 2487, 1122, 173, 1129, 642, 2404, 3182, 1269, 785, 182, 3200]. Explosive [2961]. Export [329, 2483, 2882, 192, 607, 2527, 2020, 546, 2620, 2523]. Exports [2521, 3050, 1794, 2594]. Exposed [2074, 2235, 331, 1923, 3045, 957, 2718, 2806, 520, 1871, 3198, 3184, 2946]. Exposure [1940, 963, 1142, 485, 682, 2203, 1956, 1256, 1225, 1703, 604, 1887, 573, 2180, 2962, 205, 2668, 2767, 214, 2726, 1805, 1461, 20, 2489, 2503]. Exposure-Induced [1461]. Exposures [2006, 2615]. Expressed [2759]. Expression [53, 967, 2772, 1943, 2791, 2904, 941, 35, 2271, 2084, 2669, 2311, 3160, 1136, 1805, 1207, 2372, 2258]. Expressional [295]. Expressions [2134, 65]. Extended [940]. Extending [957]. Extension [782, 1888, 2004, 899, 1321, 3073]. Extensions [1579]. Extent [489, 2590, 1830, 2144]. Externalities [2608]. Extinction [2911, 29, 1466]. Extinctions [1229, 752]. Extirpation [2471, 855, 2017]. Extracellular [175, 1098, 396, 2138]. Extractable [1547]. Extracted [276, 2033]. Extraction [1659, 1894, 1113, 1866, 1131, 2337]. Extracts [2634, 2707, 1657]. Extraordinary [904]. Extreme [1642, 1928, 501, 2623, 1914, 3188, 2261, 1551, 1836, 365, 2580, 2766, 841, 1726, 122, 1984, 1295, 2199, 2441, 1473, 1306, 288, 2969, 964, 1890]. Extremely [1107, 2551]. Extremes [53, 1147]. Eye [2334, 76]. Eyes [2282]. Eyestalk [3127]. Eyestalk-Androgenic [3127].

F [884, 895]. Face [502, 1826, 2118, 1829, 2225]. Faced [794]. Faceted [3134]. Facilitate [2276]. Facilitates [2980]. Facilitating [758]. Facilitation [1879]. Facilities [2450]. Facing [150, 3186, 1538, 1618]. Factor [1224, 2940]. Factors [765, 3014, 1892, 1104, 864, 381, 3104, 828, 2467, 2121, 2513, 143, 2813, 2172, 137, 2300, 2536, 724, 399, 2022, 2496, 1275, 2445, 2748, 3207, 3064, 761]. Facultative [3121]. FADS2 [2669]. Failed [917]. Failing [1549]. Fails [994]. Failure [839, 603]. Failures [2985]. Falkland [379]. Fall [311, 1436]. Falls [5, 2424]. False [2854]. fam [1261]. Families [2639, 110]. Family [1955, 1600, 112, 3157, 295, 2970, 1261, 3035, 119, 2169, 3002, 1207]. fangsiao [2802]. fanii [1110]. Far [2079, 1824, 1850, 2441, 1473]. Far-Field [2079, 1850]. Farm [277, 674, 1704, 658, 626]. Farmed [1539, 2725]. Farmer [2397]. Farming [473, 3210, 2534, 1871, 2855, 1239]. Farmland [3200]. Farms [842, 3161, 3052, 1533, 2236, 2239, 1144, 2529, 164]. farreri [2877, 1951, 3184]. fascicularis [1864]. Fast [2528, 3103]. Fast-Growing [3103]. Faster [2354]. Fasting [3056]. Fat [2927, 465, 520, 2090]. Fatal [438]. Fatality [1922]. Fate [872, 1564, 1827, 2528, 970]. Fatty [939, 1823, 744, 837, 1449, 2204, 1297, 633, 2297, 941, 2380, 3027, 2671, 167, 2669, 2658, 220]. Fauna [1456, 1712, 2173, 1198, 554, 552, 739, 1851, 237, 2149, 2744, 2057]. Faunal [48, 343, 937, 1357, 1680]. faveolata [2687, 1654, 2311]. Favites [2565]. Favorable [2847]. Favoring [336]. Fe [42, 1119, 1709]. Feasibility [473, 965]. Feature [2012, 1477]. Feature-Based [2012]. Features [1032, 2663, 3047, 286, 1052, 857, 1259, 929, 9, 2105, 2380, 325, 2217, 3133]. February [2755]. Fecal [73]. Fecundity [1401]. Fed [2595, 1297, 941, 3027, 2574]. Feed [2820, 30, 2264, 248, 1067, 695, 2669, 2579]. Feedback [1363]. Feedbacks [1942]. Feedforward [1781]. Feeding [1630, 2273, 1750, 2667, 1763, 2494, 744, 2863, 2014, 2333, 1909, 2756, 1101, 2323, 2904, 969, 1714, 2880, 2956, 2163, 1503, 2098, 1744, 2893, 1735, 248, 1067, 2777, 667, 1450, 220, 3103, 2996, 878]. FeedKind [3141]. Feeds [1294, 326]. Female [1939, 1579, 3183, 2496, 662, 2936, 3174, 3160]. Females [662]. Fences [3001]. Fenneropenaeus [2897]. Feralization [2239]. Fermentation [663, 635]. Fernando [1436]. ferox [2349]. Ferries [2572]. FerryBox [3139]. Fertility [2489]. Fetal [1012]. Few [778]. Fibers [3039]. Fibrous [2785]. Fidelity [516, 1540, 1734]. Field [500, 2202, 767, 131, 2048, 48, 1507, 2079, 128, 2019, 1120, 317, 2303, 1850, 559, 81, 115]. Field-Based [115]. Fieldable [1808]. Fields [3058, 2068, 933, 302]. Fight [243]. Figueira [2513]. Fiji [2273, 2338, 648, 2420]. Fijian [1125]. Filament [2698, 1625]. Filamentous [1426, 1364, 2777]. Filet [161, 2669]. Fill [832]. Filling [1781]. Film [3200]. Films [1516]. Filtering [2849]. Filters [3135, 2679, 2097]. fimbria [1241]. Fin [2465, 1130, 1714, 1145, 1682, 1607]. Financial [853]. Finding [75, 1820]. Findings [2334, 786, 605, 732, 2320, 3113]. Fine [2294, 192, 1795, 2955, 2383, 414, 3069, 3057, 1101, 1858, 1368, 2257, 1230, 259, 27, 1856, 1702]. Fine-Scale [2294, 1795, 2383, 414, 3069, 1101, 1858, 1368, 2257, 259, 27]. Fine-Scaled [1856]. Fine-Structure [2955]. Fine-Tuning [1230]. Finfish [2113, 1700, 502, 38, 2835]. Fingerprinting [1924]. Finite [2487]. Finite-Volume [2487]. Finland [2233]. Finless [1025, 1370]. FIO [606]. FIO-ESM [606]. First [2334, 2324, 570, 2060, 3123, 489, 544, 945, 206, 2593, 277, 3114, 1824, 2178, 1158, 2999, 2018, 1426, 111, 523, 1951, 1731, 921, 1266, 2098, 3051, 2320, 2834, 2526, 2992, 115, 2991, 2190, 2847, 3114, 2261]. Fish [2820, 1378, 2214, 2071, 3180, 2028, 1472, 2379, 1027, 3191, 2202, 381, 2060, 185, 2845, 2595, 45, 52, 1024, 2005, 2730, 628, 457, 531, 1676, 353, 440, 1600, 2590, 3161, 2390, 2358, 2899, 2514, 2841, 2418, 1755, 674, 962, 1523, 1812, 3077, 2950, 2072, 181, 1184, 1163, 1742, 313, 1819, 829, 1809, 720, 161, 1525, 2957, 1565, 825, 1889, 416, 725, 1374, 1400, 969, 3146, 3107, 348, 573, 2803, 297, 2635, 1626, 681, 2297, 1487, 119, 1871, 2241, 271, 2221, 747, 2707, 2142, 2556, 543, 268, 1744, 2145, 384, 2398, 724, 1143, 248, 626, 3027, 374, 1946]. Fish [2103, 3128, 727, 386, 2058, 1477, 1336, 259, 2404, 74, 1524, 2741, 1012, 871, 2239, 707, 1265, 1229, 752, 2867, 718, 266, 3160, 3078, 2287, 459, 254, 2420, 1287, 831, 844, 476, 1303, 2211, 2696, 148, 2540]. Fish-Specific [1525]. Fisher [2145, 344, 2336]. Fisheries [3180, 1739, 2739, 748, 153, 805, 1541, 1350, 1445, 1310, 3025, 852, 2621, 716, 2682, 371, 3136, 2359, 2246, 557, 420, 1249, 1757, 2274, 1901, 1760, 1344, 1559, 1166, 2347, 116, 2021, 579, 1441, 1407, 2418, 419, 673, 36, 1393, 3008, 1279, 2228, 2318, 1579, 1809, 2732, 1697, 1594, 1416, 2259, 1029, 335, 2808, 1315, 1400, 1434, 2236, 2790, 200, 1949, 2722, 913, 3167, 2560, 1548, 1059, 127, 2834, 679, 1593, 1452, 1705, 2143, 858, 927, 561, 2188, 1333, 334, 2966, 401, 1431, 2600, 1611, 1738, 2647, 2444, 459, 417, 843, 2505, 2602, 2622, 1482]. Fisheries-Induced [417]. Fishermen [2136, 1241]. Fishers [1739, 1573, 381, 610, 579, 2665, 3087]. Fishery [1367, 1815, 805, 3191, 1360, 741, 1118, 728, 962, 2561, 2717, 3126, 101, 2958, 1699, 269, 2661, 1487, 68, 2149, 1418, 1060, 78, 543, 2722, 1546, 1156, 1453, 2367, 399, 917, 532, 727, 1361, 1509, 1691, 2445, 1229, 752, 742, 970, 3026, 609]. Fishery-Dependent [68]. Fishery-Driven [728]. Fishery-Independent [543]. Fishery-Seascape [970]. Fishes [2689, 1595, 2699, 1692, 519, 283, 1175, 1729, 3118, 836, 2541, 712, 328, 3076, 3212]. Fishing [1573, 264, 1360, 1846, 2972, 185, 79, 757, 121, 1624, 1703, 958, 538, 1066, 2338, 1447, 226, 1150, 1072, 2631, 583, 1859, 913, 399, 2146, 2206, 3162, 1868, 612, 2410, 410, 2364, 845, 2731, 1844, 983, 2184, 2381, 93, 832, 2177, 2644, 2602, 2696, 1298, 783]. Fishing-Capture [1703]. Fishmeal [3130]. Fit [2621]. Fitness [862, 1289, 2965, 1152, 1762, 1524, 1609, 2938]. Fits [384]. Fitted [2524]. Five [857, 1169, 1383, 2337]. Fixation [513, 2078, 491, 1661, 1864]. Fixed [1742, 374]. Fixing [2281]. Fjord [1753, 2575, 1662, 1502, 874, 2303, 1503, 2341, 1275, 686, 1500, 1756]. Fjords [2431, 1499, 2512, 495, 1497, 1740]. FL [2703]. Flashlight [181]. Flat [2887, 2568, 2495, 838, 521, 3096]. Flat/Lagoon [838]. Flatfish [449]. Flats [453, 494, 2747, 2172, 3040, 1439, 2773]. flavescens [881]. Flavin [2806]. Flavin-Containing [2806]. Flawed [108]. Fleets [2146]. Fleshy [2729]. flesus [1567]. Flight [2325]. Float [435, 378, 242, 1512]. Floating [2507, 1531, 1636, 1574, 1570, 1587, 892, 2073, 114, 1365]. Floats [1057, 2326, 802, 2262]. Flocculation [1664]. Flocs [2459]. Flood [1397, 2539, 3074, 1091, 2929]. Flooding [3031, 2100, 861, 201, 2261, 1958, 1352, 288, 301]. Florida [2319, 3097, 2796, 1726, 1702, 1534, 1808, 2220, 1433, 2202, 1167, 795, 414, 3089, 436, 640, 310, 1043, 1535, 2635, 1343, 677, 265, 1982, 1996, 2458, 2749, 2053, 2311, 2969, 2953]. Flounder [1716, 1567, 1207]. Flow [1276, 1093, 411, 2182, 813, 2001, 914, 2873, 1557, 1426, 2443, 2298, 412, 1399, 613, 211, 1791, 2675, 3053]. Flower [474]. Flows [676, 131]. Fluctuation [1977]. Fluctuations [3100, 2429]. Fluid [2770, 497, 211]. Fluorescence [704, 2877, 2890, 2610, 377, 1358, 2412]. Fluorescent [229, 1937, 2280, 1411, 1151]. Fluorometric [3139]. Fluted [3159]. Fluvial [3040]. Flux [2431, 2301, 2882, 44, 1838, 2488, 666, 99, 1704, 2617, 1317, 1501, 3040, 878]. Fluxes [2286, 1311, 552, 814, 2478, 460, 1296, 2172, 2995, 2405, 904, 2474, 252, 973, 1457, 2449]. Fly [2589, 2385]. Flying [2032]. FMAIa [1284]. Focal [385]. Focus [2580]. Focused [370]. Foissner [1590]. Following [1053, 903, 2840, 1703, 2994, 6, 92, 1951, 689, 2754, 1785, 937, 2441, 1473, 288, 751, 1524, 1764, 564]. Follows [385]. Food [907, 1964, 376, 607, 1189, 2815, 615, 1759, 1793, 837, 1663, 1066, 1375, 296, 882, 2013, 1953, 2957, 738, 2164, 1157, 2708, 2517, 1222, 1143, 1822, 1432, 227, 1339, 2276, 743, 3174, 1613, 889, 2408]. Food-Related [2708]. Food-Web [1066]. Food-Webs [2517]. Foodweb [963]. Foot [3058]. Footage [1676]. Footprint [1710, 1704, 1699]. Forage [1472, 2590, 2404]. Forager [190]. Foraging [2375, 1766, 2285, 661, 728, 2714, 432, 2009, 1551, 2914, 2040, 518, 1206, 1960, 199]. Foraminifera [31, 1746, 2256, 837, 1411, 1899, 1985, 1860, 3140, 687]. Foraminiferal [2044, 1845, 18, 1103, 2814]. Forces [1121]. Forcing [1616, 1771, 497, 3208, 2876]. Forecast [2805, 2735, 69, 1869]. Forecasted [3062, 1613]. Forecasting [2428, 2111, 205, 1861, 2147]. Forecasts [1064, 2686]. Forensics [2143]. Forest [3080, 1678, 853, 646, 124, 49, 177, 956, 2754, 1417, 307, 3155]. Forests [1921, 2611, 1252, 1875, 558]. Form [2868, 1364, 1590]. Formalin [1742]. Formalin-Fixed [1742]. Formation [2310, 509, 3060, 1713, 1675, 1437, 1202, 952, 1951, 1827, 358, 3036, 1545, 2762]. Formed [872]. Formigas [9, 2105]. Forming [842, 2975, 813, 3029, 196, 1721, 833, 1426, 816, 2135, 993]. Forms [1463, 1567, 2423]. Formulation [1675]. Forward [1617, 731]. Fossil [2504]. Fostering [1519]. Fouling [3161, 3044, 2634]. Found [1476, 1302, 1147, 3072, 1327]. Foundation [907, 565, 2386, 1585, 2852, 1962]. Four [134, 1163, 997, 1770, 1374, 3072, 1327, 1266, 2308, 894, 2305, 642, 2357, 3022]. Four-Receiver [2308]. Fourier [2418]. FOWTs [1349]. Foz [2513]. Fractal [2554]. Fractionated [867, 2366, 2422]. Fractionation [2116]. Fractions [18, 2672]. Fracture [1987, 2068]. Fragilariopsis [1709]. Fragmentation [3042, 2174, 2543]. Fragmented [2244, 3161]. Fragments [2176, 1856]. Fram [1293, 659, 2342, 1078, 452, 530, 1457]. Frameshifting [1243]. Framework [1975, 1560, 2903, 3147, 2274, 1430, 1695, 1393, 527, 2170, 981, 1889, 2517, 677, 982, 1432, 2065, 892, 2275, 1544, 2748, 1879]. Frameworks [1954, 1300, 2490, 1884]. Franciscana [801, 1558, 1284]. Francisco [189, 1794, 1732]. Fraser [1187]. Freak [1121]. Free [2596, 1639, 59, 1491, 1001, 657, 446, 2828, 2876, 2360, 2423]. Free-Living [1639, 2360]. Free-Ranging [1491, 1001, 657, 446]. French [120, 74, 1197, 1797, 902, 262]. Frenchman [628]. Frequency [2535, 2955, 403, 1316, 2161, 1426, 1882, 312, 2101, 1563, 1983, 1091, 2316]. Frequency-Dependent [1316]. Fresh [2286]. Freshened [2828]. Freshening [2794, 708]. Freshwater [1823, 1512, 1044, 2841, 2409, 1655, 1471, 1340, 879, 2755, 1662, 2642, 1603, 395, 1201, 2929]. Freshwater-Influenced [1340, 879]. Friendly [105]. Fringed [798, 547]. frondosa [2460]. Front [2955, 2272, 2478, 1423, 3177]. Frontal [478]. Frontier [2103]. Frontiers [508]. Fronts [1516, 1438, 2020]. Fucalean [2011]. Fucales [2331]. fucata [3103, 1970, 2791, 2573]. Fucoxanthin [1894]. Fucus [2110, 21]. fuegensis [424]. Fuegian [424]. Fuel [2957]. Fuels [518]. Fujian [3009]. Full [2514, 2841, 1264, 274]. Full-Length [2514, 2841, 1264, 274]. fulva [2848]. Function [408, 2812, 2595, 1388, 2920, 83, 2844, 2648, 1403, 1288, 3194, 2674]. Functional [243, 2421, 2133, 1224, 522, 3003, 112, 1981, 697, 1036, 298, 2747, 693, 3035, 1959, 1154, 2849, 2275, 574, 1484]. Functionality [2454]. Functioning [1337, 787, 2932, 528, 1171, 597, 559, 3088]. Functions [2365, 1695, 2791, 3148, 2680, 1124]. Funders [2566]. Fundy [1137, 1948]. Fungal [2377, 1424, 1135]. Fungi [135, 2777, 2858]. Fungus [415]. Fuqing [2538]. Furanone [952]. Further [1132]. furvus [1195]. fuscatus [2906]. fuscoguttatus [2595, 3146, 2124, 2870, 2388]. fusiforme [1657]. Future [1964, 66, 672, 53, 464, 1510, 2972, 1322, 2872, 1876, 339, 979, 1253, 3199, 1561, 3210, 59, 1190, 2312, 2347, 1415, 213, 2884, 2534, 1393, 770, 2165, 1448, 2426, 2013, 2082, 237, 981, 2078, 246, 2965, 1460, 1968, 835, 2652, 1588, 365, 100, 1271, 600, 724, 982, 1389, 227, 2360, 1627, 2694, 865, 1833, 1352, 1148, 1774, 443, 3020, 2279, 1747, 2613, 799, 301, 439, 233, 3022, 2601]. Futures [947, 928, 1059]. Fuvahmulah [856]. Fuzzy [1560]. Fuzzy-Based [1560].

G [880]. G. [1422]. gaditana [180]. Gadus [667, 2503]. gahi [1329]. Gain [3070, 2442]. Gained [1406]. Gaining [708]. Gains [505]. Galápagos [1245, 917]. Galaxea [1864]. Galicia [2335]. Galician [2469, 2647, 2851]. gallina [1102]. galloprovincialis [821, 1159, 2962, 1056, 2008]. Galveston [2984, 2559, 288]. Game [2385]. Gamete [2545]. Gametogenesis [3123]. Gametogenic [1389]. Gametophyte [859]. Gametophytes [35]. Gametophytic [496]. Gamifying [1870]. Gaming [1576]. Gamma [3120].

Gamma-Rays-Induced [3120]. gangetica [1370]. Gap [987, 3143, 1781, 1133, 401]. Gaps [2662, 1310, 1561, 1555, 1862, 2396, 127, 1417, 832, 402]. Garbage [1516]. Garcia [2193]. Garden [2443, 474, 562, 2719]. Gardening [774]. Gardens [597]. gardneri [21]. gariepinus [293]. Gas [2698, 273, 544, 1437, 2072, 2049, 2636, 2345, 2474, 1833, 1698]. Gastric [515]. Gastronautidae [1629]. Gastropod [1876, 3105, 1847, 898]. Gastropoda [3105, 2439, 325]. Gatekeepers [647]. Gates [2001]. Gateway [902]. Gauges [1091]. Gauging [1509]. Gauntlet [168]. Gaussian [2529]. Gazing [100]. Gazul [9, 2105]. GC [2979]. GC-MS [2979]. GC-MS/MS [2979]. GCOAST [1210]. GDGT [2677, 860]. GDGT-Based [2677]. GDGTs [2677]. Gear [1360, 608, 3162, 2410, 832]. Gear-Induced [608]. Gearing [51]. Gecarcoidea [2658]. Gel [507]. Gel-Like [507]. Gelidiales [1110]. Gelidiella [1110]. gelidus [55]. gen [1261, 1242]. Gender [2799, 399, 401]. Gene [2802, 53, 1955, 813, 967, 2772, 2920, 2134, 2607, 1226, 1943, 2791, 2843, 295, 2904, 3035, 35, 2314, 26, 1800, 2084, 3187, 2669, 2311, 1136, 65, 1207, 1304, 3079, 2372, 2258]. Genera [1537, 3034]. General [1592, 169, 327]. Generalist [2074, 2863, 2102, 968]. Generation [1086, 1848, 774]. Generational [2784, 386]. Generic [3010]. Generic-Level [3010]. Genes [294, 2791, 2655, 2806, 825, 636, 1175, 941, 3127, 1174, 2658, 3160, 1805, 2408, 2758, 2752, 2886, 2493, 935]. Genet [1608]. Genetic [1499, 3014, 998, 656, 1462, 1332, 2775, 2108, 1481, 1823, 14, 501, 1853, 1721, 2900, 1806, 2047, 2655, 3153, 735, 2874, 996, 47, 2056, 1988, 1026, 2725, 1996, 2943, 1018, 268, 1762, 2908, 2702, 2123, 187, 2125, 1856, 2572, 199, 197, 2002, 2875, 1136, 22, 56, 1422, 618, 476, 2757, 1284]. Genetically [225, 2626, 1993]. Genetics [891, 636, 796]. Genome [53, 32, 1955, 760, 572, 295, 47, 796, 2473, 1451, 26, 2810, 938, 2878, 1800, 2125, 392, 2533, 2897, 22, 1382, 2801, 2925, 2886]. Genome-Based [2473, 1382]. Genome-Wide [53, 32, 1955, 760, 295, 47, 796, 938, 2125, 2886]. Genomes [542, 1110, 816, 3172, 2135, 1660, 1338]. Genomic [1541, 1795, 1243, 2899, 101, 2399, 927, 2980, 3166]. Genomics [1795, 2859, 3179, 2971, 2135, 2352, 2911, 2331, 123, 2875, 476]. Genomics-Based [2875]. Genotoxic [3120]. Genotype [736, 2176]. Genotypes [2127]. Genotypic [98]. Genotyping [2655]. Genotyping-by-Sequencing [2655]. Gentian [2388]. Genus [162, 3189, 571, 1324, 2135, 3072, 1327, 393, 600, 1080, 1356, 1242, 2664, 1227, 2416, 1837, 2801]. Genypterus [14, 927]. Geochemical [33, 2403, 1976, 122]. Geochemistry [2737, 217, 1500]. Geodia [1226, 1196, 1203]. Geoengineered [922]. Geographic [2799, 2760, 1236, 1995, 1413, 977, 1191, 1934, 655, 3093, 1028]. Geographical [2578, 1853, 1893]. Geographically [2130]. Geological [1357]. Geology [131]. Geomagnetic [516]. Geometry [2344, 2270]. Geomicrobiological [2770]. Geomorphic [113]. Geomorphological [2425]. Geomorphologies [1074]. Geomorphology [2812, 453, 316, 1798].

Geomorphometric [292]. geonameCanary [2165]. Geopolitics [25]. Geoportal [2358]. George [441, 2755]. Georgia [2753, 1961]. Geoscience [1883]. Geospatial [706, 122]. Geostatistical [2350, 1507]. Geostrophic [1276]. GEOTRACES [1979, 670]. Gephyrocapsa [3165]. German [2402, 2774, 1803, 2320, 1908]. GFCM [1626]. GFCM-GSA [1626]. Ghost [1414]. giacconei [3029]. Giant [3124, 3045, 1653, 1971, 996, 1733, 1335, 3159, 567, 1178, 2723, 1698, 2002, 3148, 2680]. Gibraltar [2916, 1841]. Giga [270]. Giga-Project [270]. gigantea [2956, 2896]. Gigantism [2896]. Gigartinales [468]. gigas [748, 501, 2724, 1204, 744, 67, 2878, 1258, 2759, 2907, 2501]. Gill [502, 2982, 3193]. Gillnet [1249]. Gillnets [801, 641]. Gills [1631, 2831, 3055]. Gilthead [408, 1988, 2024, 2497, 2090, 26, 2574, 2452, 2258]. GIpr07 [1979]. Gironde [2532]. GIS [418]. Give [1364]. glaberrima [1451]. glacialis [3115]. Glacier [1502, 1740, 2242]. Gland [2114, 3127]. Gland-Testis [3127]. Glass [907, 2475, 1217, 956, 2528]. glauca [1957]. Glider [500, 2492, 2474, 1452, 2207, 2189]. Glider-Based [1452]. Gliders [1123, 203, 2092, 729, 377]. Global [1873, 1042, 2100, 1807, 267, 1964, 800, 570, 2428, 2778, 947, 1598, 3203, 2770, 824, 1741, 2126, 1891, 62, 979, 486, 3197, 3125, 2932, 116, 2453, 2751, 1007, 2385, 1120, 2637, 1594, 634, 738, 568, 1334, 566, 616, 897, 8, 84, 365, 3032, 2632, 1271, 1910, 2834, 2075, 2410, 553, 1478, 1848, 3013, 1295, 2852, 394, 858, 1538, 362, 1431, 2650, 1501, 1623, 2875, 802, 1325, 2226, 2959]. Global-Local-Comparison [1623]. Global-Scale [1594]. Globalization [772]. Globalized [557]. Globally [1190, 3186, 454, 768, 1661]. Gloria [1882, 1898, 1861, 1890]. Glutamic [695]. Glutamine [3187]. Glutaredoxin [2372]. Glycans [2853]. Glycerol [860, 2669]. Glycogen [2314, 2759]. Glycoglycerolipids [219]. Glyphosate [24]. Glyptonotus [3045]. GMP [2310]. GnRH [1008, 1354, 2084]. Go [2126, 1093]. GO-SHIP [1093]. Goa [1789]. Goal [2659]. Going [862]. Golden [1345, 2134, 3130, 3055]. Golf [776]. Gonad [2772, 3127, 2758]. Gonad-Inhibiting [3127]. Gonadal [3014, 851, 3189, 2084]. Gonadosomatic [3014]. Gonads [695]. gondii [438]. Goniopora [1422]. Good [713, 1799, 405]. Goods [1521]. Google [2508]. GOOS [4, 2466]. Gorda [131]. Gorgonians [803, 1962]. Gorgonocephalidae [901]. Goro [1902]. Gossypol [3130, 2870]. Gouic [1557]. Gould [1195]. Governance [1873, 1739, 888, 1745, 3197, 557, 1817, 1194, 639, 1877, 1638, 2244, 2122, 2132, 2570, 2309, 186, 349, 2353, 731, 2549, 1387, 1687, 1874, 678, 3142, 2136]. Governed [2144]. Governing [2618, 164]. Governs [910]. Gracilariopsis [3158]. gracilenta [2857]. Graded [941]. Gradient [473, 2462, 1512, 1763, 717, 2844, 50, 3034, 450, 1429, 2200, 117, 103, 788]. Gradients [66, 887, 1021, 1394, 143, 1435, 2909, 2257, 3109, 2798]. Graduate [630]. Gram [2016]. Gram-Negative [2016]. Grampus [385]. Gran [1423]. Grand [464, 1823]. granosa [1751]. Grant [1305]. granulifera [1590, 1590]. granulosus [1840]. Grass [2641, 2792, 2489]. gratilla [2339]. Gray [2333, 236, 1148, 1572]. Grayling [2676]. Grazer [1875]. Grazers [238, 2386, 994]. Grazing [990, 238, 2683, 2014, 2077, 2076, 2464, 2824, 17]. Great [1663, 2771, 997, 1191, 1045, 2231, 134, 1599, 170, 98]. Greater [1233, 24, 2442, 2596]. Greatly [163]. Green [3011, 1476, 792, 2280, 760, 2186, 75, 2009, 1894, 816, 2135, 2406, 749, 775, 1734, 199]. Green-Tide [816, 2135]. Greener [1767]. Greenhouse [2623, 1111, 2049]. Greenland [1293, 2151, 1527, 76, 2099, 541, 2619, 562, 1275, 1206]. Greenling [520]. gregaria [1750]. Grenadines [2137]. Grey [1212]. Grid [2487]. griseus [385]. Gross [622, 623]. Ground [627, 1163, 690, 1913, 706, 2032, 2163, 1266, 1231, 2696]. Ground-Truthing [690]. Groundfish [1509]. Grounding [2695]. Grounds [661, 621, 2475, 1217, 2333, 296, 1196, 518, 1844, 199, 2177]. Groundwater [2286, 19]. Group [440, 385, 2629, 1969, 2205, 2305]. Grouper [2595, 475, 2362, 3146, 259, 2124, 3148, 2680, 2579, 2870, 2388, 2762, 1010]. Groups [2421, 1750, 1959, 2744, 2275, 2741]. Growing [424, 111, 756, 3103, 783]. Grown [3037, 180]. Growth [2968, 424, 408, 2214, 19, 1102, 2586, 2641, 585, 2371, 1127, 2918, 990, 2905, 2589, 2595, 61, 2645, 590, 2670, 1039, 3205, 128, 2019, 2683, 580, 1068, 3063, 1569, 3091, 2309, 2080, 60, 2361, 1058, 207, 2537, 342, 930, 2415, 3061, 235, 161, 2655, 1126, 1602, 2232, 2756, 523, 1297, 1651, 1374, 2961, 2362, 1419, 2114, 3065, 1145, 2176, 1485, 791, 1038, 422, 1888, 2024, 2556, 290, 67, 2314, 1608, 1429, 2090, 1504, 1056, 3130, 180, 2676, 3027, 195, 355, 2921, 1515, 55, 1162, 1172, 2719, 2966, 667, 2160, 671, 2669, 3111, 1199, 2907, 2444, 1732, 417, 889, 2870, 20]. Growth [3026, 3103, 3053, 3038, 2516]. Growth-Defense [2907]. Growth-Promoting [2670]. Growth-Related [2655]. GSA [1626]. Guajira [1644]. Guangdong [2442]. guangzhouensis [1341]. Guarapiranga [3182]. Guatemala [2234]. Guayaquil [855, 225]. Guaymas [256, 80]. Guiana [1185, 1633, 1112]. guianensis [1185, 1633, 1112]. Guide [3008, 511, 2434]. Guidelines [1741, 1444, 2790, 2533]. Guinea [2551, 1276, 2730, 3050]. Guldberg [997]. Gulf [225, 2974, 615, 2760, 2730, 945, 139, 3179, 432, 2804, 2009, 1589, 675, 151, 111, 1855, 3178, 1434, 909, 1798, 749, 119, 1628, 1454, 1820, 1627, 366, 158, 2040, 549, 775, 2924, 2083, 1844, 3076, 1369, 254, 80, 2233, 2479, 2499, 809, 1276, 229, 805, 1433, 2299, 2150, 1975, 1846, 903, 451, 1580, 403, 654, 2623, 855, 196, 436, 1601, 72, 418, 1158, 1036, 2072, 1221, 310, 2808, 996, 1530, 204, 2652, 69, 681, 2149, 3112, 1442, 2062, 2876, 359, 2373, 594, 70, 1952, 123, 1934, 2528, 233, 2138, 3150, 2551]. Gulls [1991]. Gully [283]. Gustav [1526, 3108, 1198]. Gut [1773, 2595, 1600, 2843, 2642, 2024, 2497, 2098, 2090, 1076, 2777, 2574, 2304]. guttulatus [740]. Gwangyang [1081]. Gymnothorax [2280]. Gyre [1771, 2451, 1634, 1317, 3164]. Gyres [2422].

H. [2658]. Håkon [1606]. Habit [2408]. Habitat [1526, 1865, 2071, 1706, 2185, 2151, 286, 672, 1052, 2174, 1464, 1712, 2421, 2578, 1137, 482, 2089, 2847, 3100, 813, 133, 3062, 1070, 2691, 139, 986, 228, 379, 196, 1290, 1721, 2884, 833, 1695, 1301, 2950, 1550, 837, 770, 1221, 1086, 726, 2323, 966, 479, 2052, 909, 449, 934, 3054, 901, 2118, 2398, 1048, 1540, 1380, 2914, 359, 1947, 2341, 680, 2892, 1848, 2368, 1477, 325, 1854, 1610, 315, 2160, 1495, 3078, 1450, 2023, 574, 1870]. Habitat-Forming [1721, 833]. Habitat-Specific [837]. Habitat-Use [1380]. Habitats [1009, 1807, 656, 286, 1994, 1598, 1372, 1249, 1676, 2590, 3161, 2375, 1309, 3081, 1036, 1655, 1652, 1289, 1976, 2546, 3072, 1327, 1274, 2568, 2995, 1370, 2222, 1016, 3118, 2981, 988, 804, 347, 2924, 2526, 402]. Habits [2009, 2756, 1101, 1744, 220]. HABreports [1700]. HABs [138]. Hadal [2178, 2896, 2980, 1553]. Haddock [2767]. haeundaensis [1215]. Haima [3192, 2978]. Hainan [2926, 3124]. haitanensis [32, 517, 889, 2476]. Hake [2469, 1529, 1064]. Half [121, 2052]. Half-Life [2052]. Halichondria [216]. Haliclona [2848, 368]. Halifax [2243]. Halimeda [2923]. Haliotis [2956, 1943, 2127, 2956, 2084, 3053]. Hallmarks [1256]. Halocline [3043]. Halophila [602, 77, 2875, 3019, 402]. hamata [498]. Hammerhead [1630, 1663, 966, 1191]. Han [163]. Hands [258]. Hands-on [258]. hannai [1943, 2956, 2084, 3053]. Haplosporidium [484, 152]. Haplotypes [3165]. Haptophyte [991]. Haptoria [1837]. Harbor [421, 2091, 2243, 1561, 2129, 829, 446, 834, 1273, 2263, 89, 245, 1740]. Harbour [387, 2320]. Hard [2736, 1753, 48, 986, 2920, 1824, 886, 1652, 2945, 1418, 1018, 2671, 618, 458]. Hard-Bottom [1753, 1018]. Hard-Shelled [458]. harengus [1412]. Harmful [2787, 2782, 1700, 2753, 2111, 2534, 2842, 3097, 2077, 1105, 68, 138, 165, 439]. Harmonization [39]. Harmonized [819]. Harnessing [3067]. Harp [989]. Harpacticoids [674]. Harpadon [2780]. Harvest [3008, 754, 2367, 612, 1298]. Harvester [1142]. Harvester-Consumers [1142]. Harvesters [1859]. Harvesting [584, 173, 985]. Harvey [130, 1904, 288, 564]. harveyi [2214]. Hatch [603]. Hatchery [2264, 2757]. Hatchery-Released [2757]. Hatching [2802, 2564]. Hatchling [808, 323, 795]. Hatchlings [403]. Hato [338]. HAUSGARTEN [423, 530]. Hav [1606]. Hawai’i [3101, 2963]. Hawaiian [1706, 2185, 241, 412, 762]. Hawksbill [3101, 909]. Hazard [169, 3074]. Hazardous [1998]. Hazards [1267, 640, 1958]. Headland [3198]. Health [1539, 141, 406, 2589, 2864, 1006, 469, 502, 2751, 2362, 1335, 3065, 1701, 2668, 2754, 1748, 2441, 1473, 1148, 1461, 2870]. Healthy [2666, 109, 1855, 195, 1538]. Heart [995]. Heat [1053, 2074, 1708, 281, 1039, 1830, 2478, 3037, 1557, 2761, 2306, 3034, 1628, 2582, 2766, 2671, 822, 496, 955, 1501, 2357, 1931, 2726]. Heat-Stressed [2306]. Heat-Tolerant [1708, 3034]. Heatwave [805, 2462, 1046, 1869, 1369, 3168]. Heatwaves [250, 1161, 11, 3012, 565, 3188, 1443, 1656, 29, 848, 2723, 2613, 2540].

Heaviside [921]. heavisidii [921]. Heavy [2208, 2150, 819, 1004, 514, 1153, 2881]. Hediste [2204, 884]. Height [2215, 2289]. Heights [411]. heinsohni [1320]. Held [1762]. Helena [1729, 1866, 1966, 1048]. helgolandicus [1248]. helicina [53]. Helicobacter [515]. Heliconoides [916]. Helicostomella [2716]. Heliocidaris [1149]. Heliopora [2545, 2664]. Helmholtz [2301]. Help [2971, 1968]. Helps [3025, 959, 1698]. Hemato [293]. Hemato-Biochemical [293]. Hematological [95]. Hematology [2586]. Hematopoietic [2271]. Heme [1751]. Hemisphere [2837, 1682]. Hemispheres [1177]. Hemocyte [2946]. Hemoglobin [2033, 1751]. Hemolymph [1076, 985]. Henle [1125]. Henry [217]. Hepatic [1920, 2258]. hepatica [437]. hepaticum [437]. Hepatitis [437]. Hepatopancreas [851, 2952]. Hepatopancreatic [3061]. Herbivore [778, 307]. Herbivorous [2689, 1600, 2076]. Herbivory [2541, 898]. Herbst [3071]. Heritage [827]. Hermaphrodite [3183, 2114]. Hermatypic [1233, 1864]. Hermetia [2589, 2304]. Herpesvirus [1244]. Herring [1412, 2496, 727, 631, 1218]. Hervey [1939, 2468]. Hesionidae [5, 2424]. Heterarchy [1822]. Heterocystous [2281]. Heterogeneity [1526, 480, 1490, 285, 1368, 2050]. Heterogeneous [3084, 2059, 2521, 2686, 3026]. Heterogenic [656]. Heterokont [207]. Heteroscleromorpha [1440]. Heterotrophic [175, 2683, 2857]. Heuristic [639]. Hexacorallia [542, 2992]. Hexactinellid [1788, 1504, 1506, 1493]. Hexagrammos [520]. HF [1146, 594]. Hidden [1351, 2158, 975]. Hiding [271]. Hierarchal [2310]. Hierarchy [153]. Hierro [2438]. High [2535, 854, 1616, 224, 2379, 2151, 2203, 1262, 987, 2428, 337, 704, 1790, 1811, 1930, 2955, 2879, 2046, 1377, 1022, 2272, 1249, 501, 2927, 1084, 1945, 1507, 1160, 1267, 2983, 877, 333, 1272, 465, 235, 714, 221, 1426, 2756, 1535, 693, 3097, 1882, 2076, 312, 2681, 2213, 365, 1273, 2069, 205, 183, 2742, 2130, 2210, 1408, 2518, 2341, 1506, 2296, 2745, 368, 595, 2573, 2295, 992, 3083, 1467, 847, 2935, 715, 1338, 2870, 43, 905, 1134, 2316, 3158, 361, 2525, 2551, 1420, 2132, 735, 874, 525, 975]. High- [2525]. High-Density [1408]. High-End [1377]. High-Endurance [43]. High-Energy [1467]. High-Fat [2927, 465]. High-Frequency [2535, 2955, 1426, 2316]. High-Latitude [1272, 2341, 2295, 3083]. High-Level [2870]. High-Light-Adapted [1338]. High-Microbial [368]. High-Resolution [854, 1616, 2151, 987, 2428, 704, 2879, 2046, 1507, 1160, 877, 235, 3097, 365, 2069, 205, 2518, 2296, 1134]. High-Stock [2756]. High-Sustained [2742]. High-Temperature [714, 1535, 3158]. High-Throughput [2379, 2272, 1084]. High-Traffic [224]. High-Use [1273]. Higher [2385, 797]. Highlight [1237]. Highlighted [486]. Highlights [628, 2256, 1797]. Highly [2096, 2043, 2983, 2993, 3044, 969, 2776, 2169, 2118, 2458, 15, 1952, 1825, 361]. Hilgendorf [1012]. Hill [1615]. Hilsa [656]. Hindcast [594]. Hindcasting [2932]. Hippocampus [3125, 740, 2380]. hippurus [1963, 1637]. Histamine [2362, 905]. Histological [2380, 358]. Histology [3146]. Histomorphology [2304]. Histopathological [732]. Historical [2358, 1108]. Histories [519].

History [1378, 2985, 1486, 1329, 371, 2272, 1351, 2137, 1044, 2544, 2599, 2088, 1419, 1733, 2581, 977, 269, 1038, 834, 205, 2038, 1092, 1214, 1229, 752, 1685]. HMS [1925]. Hoegh [997]. Hoegh-Guldberg [997]. Hog [1371]. Hokkaido [1415]. Holistic [644]. Holobiont [156, 1373, 1693, 2500, 2533]. Holobionts [2901]. Holocene [2266]. Hologenome [2848]. Holograms [2459]. Holographic [755]. Holography [1015]. Holopelagic [2735]. Holothuria [2248, 1451, 2860]. Holothuroid [2460]. Homalogastra [1261]. Homarus [2299]. Home [296, 2267, 525, 179]. Homeostasis [2952, 2936, 3212]. Homodimer [1751]. Homology [3046]. Honeycomb [1959]. Hong [2442, 1025, 2967, 1380]. hongkongensis [2946]. Hooghly [1730]. Hook [2246]. Hook.f. [2875]. hookeri [2300]. Hope [1875]. Horizon [499, 903, 1681, 202, 1564, 1193, 1785, 937, 1952]. Horizontal [2324, 1424, 202, 1221, 119, 1610]. Horizontally [596]. Hormonal [2460]. Hormone [1179, 3183, 2114, 2271, 3127]. Hormone-Binding [3127]. horneri [999]. Horseshoe [1096, 2967, 145, 1131, 1076, 1278, 869, 985, 255, 1498, 1475]. Host [1480, 1043, 2497, 698, 2267, 576, 249, 98, 2888]. Host-Associated [249]. Hot [1487, 1590]. Hotspot [91, 1527, 2321, 176, 1895, 3023, 2913, 546]. Hotspots [949, 1714, 1667, 975]. Hotter [1517]. House [1399]. Houttuyn [1744]. HPMAs [15]. HRM [735]. hsp40 [1207]. Hukou [3017]. Human [1706, 2185, 2224, 3047, 1954, 1300, 1414, 203, 1539, 1921, 3209, 141, 2864, 1817, 1000, 109, 2517, 1701, 1273, 2850, 1558, 582, 1544, 1776, 2175, 1736]. Human-Dominated [3209]. Human-Driven [1736]. Human-Related [1558]. Human-Reviewed [203]. Humans [2421, 2989]. Humboldt [185, 2724, 928, 226, 1621]. Humic [1151]. Humic-Like [1151]. Humpback [1470, 1235, 2605, 224, 76, 3132, 1444, 2094, 1939, 516, 3126, 1392, 1969, 2205, 1714, 2163, 2721, 1380, 1829, 244, 1604, 2468, 1231]. Hunting [1206]. Hurricane [403, 3074, 130, 1904, 288, 564]. Hurricanes [448, 3, 1069, 350]. huxleyi [846, 806]. hyacinthus [1384, 2470]. hyblaea [3029]. Hybrid [2595, 1659, 2127, 3146, 2168, 2124, 2579, 2870, 2501]. Hybridization [2877, 1026]. Hybrids [2168]. Hydrate [31]. Hydraulic [118]. Hydro [463]. Hydro-Morphological [463]. Hydroacoustic [2547]. Hydrocarbon [1627, 1925]. Hydrocarbon-Degrading [1627]. Hydrocarbons [872, 1827, 366, 158, 1785]. Hydrochloride [2982, 3157]. Hydrochloride-Treated [3157]. Hydrodynamic [108, 856, 2268, 1167, 1063, 2518, 705]. Hydrodynamics [2061, 2344, 1803, 950, 396]. Hydrogen [175, 397]. Hydrogen-Oxidizing [397]. Hydrogenovibrio [397]. Hydrographic [2609, 984]. Hydrography [1093, 659, 1592, 2227]. Hydrolase [80]. Hydrologic [122]. Hydrological [2521]. Hydrolysate [2497]. Hydrolysates [2707]. Hydrolyzates [299]. Hydroperiod [1220]. hydrophila [2982, 3157]. Hydrophones [823, 2305]. Hydrothermal [3058, 273, 1994, 131, 176, 670, 2166, 1117, 237, 1088, 2737, 2910, 1947, 397, 302, 2650]. Hydroxamate [1119]. Hydroxy [703]. Hydroxylase [1285, 65]. Hydrozoa [63]. Hydrurga [3190].

Hypereutrophic [2394]. Hyperoglyphe [1486, 1806]. Hyperoodon [1584]. Hyperosmotic [2945]. Hypersaline [2674]. Hyperspectral [1032, 2262]. Hypertensive [2707]. Hypophthalmichthys [2603]. hypophthalmus [3152]. Hypotheses [404, 2744]. Hypothesis [1359, 1765, 1623]. Hypotrichia [1341, 1304]. Hypotrichous [1242]. Hypoxia [1262, 1845, 1037, 569, 2456, 2283, 1589, 2843, 959, 3121, 1258, 426, 2712, 1495, 1736, 2726, 2946, 777, 898, 2860, 361]. Hypoxia-Enhanced [2456]. Hypoxia-Reoxygenation [3121, 1258]. Hypoxia-Tolerant [3121]. Hypoxia/Anoxia [426]. Hypoxic [1192, 1565, 1103, 1454, 3055]. Hyriopsis [358].

i.e [83]. IAG [1132]. IAG-Switch [1132]. IAG2 [2114]. Iberia [3118]. Iberian [1643, 2513, 1529, 559, 849]. Icariin [263]. Ice [1282, 2435, 1425, 1189, 2815, 638, 1813, 1198, 3115, 554, 1830, 2453, 1359, 1824, 345, 2895, 1483, 2928, 2656, 441, 1625, 1802, 1919, 3185, 2844, 285, 1967, 904, 2828, 2876, 491, 1275, 1172, 1931, 1740, 1791, 1665, 715, 1709, 1860]. Ice-Associated [3115, 2895]. Ice-Covered [1198]. Ice-Edge [491]. Ice-Free [2828, 2876]. Ice-Growth [1172]. Ice-Rafted [2435]. Iceberg [1365]. Iceland [541, 2451, 1357, 1590]. Icelandic [2503]. Ichthyoplankton [3017, 2813, 3009, 1525, 932, 3207, 3064]. Iconic [1961, 2473, 1382]. Ictalurus [2436]. Ictiobus [2841]. Identical [631]. Identification [809, 1754, 827, 32, 1891, 1676, 228, 1584, 2321, 760, 2941, 1163, 2877, 313, 1234, 3116, 295, 1887, 932, 735, 2056, 562, 1667, 2101, 2977, 1138, 1207, 3079, 1542]. Identifications [1583]. Identified [225, 1255, 1742]. Identifies [3023, 2886]. Identify [3046, 1032, 1924, 1579, 1230]. Identifying [1492, 2502, 257, 2732, 1487, 68, 2222, 2062, 2560, 1548, 198, 2741, 2740, 402]. Idiosyncratic [572]. Ieodo [1566]. If [1722]. IFCB [1426]. IFK [2338]. II [3102, 623, 1338]. III [2656]. ilisha [3014, 656]. illecebrosus [1699]. Illegal [1573, 264, 1200, 1150]. Illex [2253, 1699]. illucens [2589, 2304]. Illuminate [1615, 401]. Illustrate [1004]. ILTER [8, 84]. Image [500, 1526, 406, 2941, 1583, 2, 1330, 3022]. Image-Based [406, 2941, 1583]. Imaged [2478]. Imagery [1903, 494, 2453, 2941, 627, 690, 49, 2512, 562, 126, 2740, 1404]. Images [666, 608, 2419]. Imaging [1795, 883, 1426, 933, 2298, 1402, 2610, 2491]. Imaging-in-Flow [1426]. IMBeR [2055]. imbricata [909]. Imitating [2016]. Immature [2847]. Immediate [2561]. Imminent [855]. Immune [408, 2802, 2905, 2595, 3154, 951, 3205, 2417, 3188, 2016, 1897, 2142, 3130, 358, 2938]. Immunity [2361, 2362, 3027, 2489]. Immunocytochemical [859]. Immunomodulation [2962]. Impact [2233, 2511, 2428, 279, 1773, 2231, 2673, 3203, 713, 2775, 62, 1444, 2566, 2567, 2240, 2131, 507, 3049, 746, 1803, 24, 3052, 2542, 2755, 3177, 706, 487, 1139, 1485, 2819, 1701, 1442, 1271, 1644, 2210, 1726, 2734, 3169, 187, 1509, 3030, 395, 334, 2156, 613, 2073, 1776, 118, 2444, 737, 2284, 3026, 3096, 1890]. Impacted [391, 2115, 514, 1952]. Impacts [1042, 1235, 1095, 2739, 800, 2224, 2800, 1446, 1932, 2577, 250, 1845, 320, 872, 1642, 1647, 79, 1532, 2805, 2397, 2322, 2462, 739, 1624, 1047, 1405, 1166, 1000, 949, 862, 3012, 343, 1755, 1158, 1149, 1066, 41, 1375, 2561, 2102, 882, 534, 2319, 2119, 658, 568, 1434, 503, 1551, 1836, 2264, 2097, 1193, 2152, 1949, 2850, 1546, 2834, 2367, 1850, 2518, 227, 1607, 21, 937, 1275, 2447, 1734, 2441, 1473, 2404, 2468, 582, 1576, 86, 1431, 2291, 319, 1764, 1738, 2851, 2938, 1613, 970, 1665, 2505, 2622, 2823]. Impair [1147]. Impaired [2489]. Impairment [59, 2516]. Impairs [991]. Impeded [751]. Impedes [1828]. Impediments [1478]. Imperfect [3052]. Implementable [260, 166]. Implementation [1441, 3139, 1133, 1003]. Implementing [2122, 444]. Implication [1400, 1487]. Implications [2233, 2071, 1739, 748, 3060, 1137, 1189, 2815, 1313, 2682, 472, 792, 2932, 1715, 414, 2570, 1394, 2677, 2417, 683, 2165, 797, 1058, 2477, 714, 237, 2443, 2563, 1460, 2088, 3035, 1788, 2803, 449, 1798, 2943, 1910, 1429, 1717, 592, 1726, 822, 1339, 1041, 2543, 1138, 1743, 3018, 1980, 2988, 2476, 1461, 1835]. Imply [1955]. Implying [2806]. Importance [1065, 1642, 2133, 1480, 2763, 247, 432, 1995, 558, 205, 1686, 2267, 2075, 1361, 775, 178, 3036, 2422, 1287, 1010]. Important [1009, 1027, 1595, 1094, 3115, 853, 1945, 1645, 1246, 479, 1219, 271, 2314, 1048, 2712, 2357, 1556, 768, 1661, 476, 1482]. Imposed [1789]. Imprint [2770, 462, 1331, 2531]. Imprinted [2772]. Improper [2470]. Improve [821, 580, 1562, 349, 758, 459]. Improved [215, 2005, 557, 685, 1488, 1375, 3000, 2069, 1515, 2690, 328, 806]. Improvement [232, 2390, 2629]. Improvements [1524]. Improves [2670, 161, 2300, 2093, 2831, 889]. Improving [2799, 2010, 1712, 1522, 1085]. Improvisus [766]. Impulsive [2079]. In-House [1399]. In-situ [1526, 2862, 406, 1006]. In-Water [1903, 539, 3206]. Inaccurate [990]. Inadequate [1391]. INBUS [2140]. Incentives [1857]. Inception [186]. Incident [1468]. incisa [512]. Including [2334, 2385, 365, 2200]. Inclusion [3065, 2188, 1333, 1879]. Inclusive [1488, 1120, 678]. Inconsistencies [2990]. Inconsistency [1332]. Incorporate [2147]. Incorporated [1919]. Incorporating [1544, 459, 2055]. Incorporation [1119, 422]. incrassata [2923]. Increase [2269, 238, 1184, 2016, 625, 695]. Increased [1596, 3102, 2950, 1828, 974, 1485, 1847, 953, 2237, 955, 1725]. Increases [682, 2070, 2386, 1165, 1764]. Increasing [1508, 845, 409]. Increasingly [1094]. Incubated [2307]. Incubation [1709]. Independent [1652, 645, 543]. Index [3014, 821, 3146, 2058, 1736, 2436, 148, 2503]. Index-Based [2058]. Indexes [1297]. India [2196, 1579, 135, 1131, 2030, 195, 2197, 1154, 2192]. Indian [3191, 324, 1186, 2786, 2166, 1234, 1088, 1101, 2890, 3006, 2703, 543, 2910, 780, 2200, 2734, 2156, 1201, 2191, 3064, 3014, 2794, 2377, 2393, 723, 1979, 2760, 1208, 3171, 812, 673, 1264, 3097, 1906, 1026, 1614, 2796, 3086, 617, 2184, 2701, 1886]. Indicate [61, 481]. Indicated [2742, 1500, 2926]. Indicates [813, 2129, 715]. Indications [188, 1467]. Indicative [2365]. Indicator [234, 257, 3008, 770, 2718, 1889, 1391, 1597]. Indicator-Testing [1889]. Indicators [577, 1228, 2510, 1430, 1830, 1489, 2517, 3159, 165, 1212, 417].

Indices [106, 2827, 2295]. indicus [1264]. Indigenous [2214, 771, 624, 2558, 2136, 1040, 280, 2526, 2338]. Indirect [2367]. Indirectly [952, 2545]. Individual [276, 828, 2232, 1072, 1056, 1067, 2565, 2741, 459]. Individual-Based [2741]. Individually [898]. Indo [1464, 589, 96, 1550, 1025, 1392, 1969, 2205, 1380, 876, 2664, 1010]. Indo-West [589, 2664]. Indonesia [850, 363, 1465, 2136]. Indoor [3188, 2665, 900, 985]. Induce [3155]. Induced [2301, 2100, 2865, 48, 1115, 134, 2927, 2986, 3122, 2912, 1423, 2660, 997, 1164, 512, 608, 448, 293, 2167, 950, 603, 2706, 2766, 759, 1912, 3120, 3, 1069, 2727, 327, 417, 1461, 2388, 300, 3004]. Induces [967, 388, 2124]. Inducible [2280]. Inducing [880]. Induction [1401, 2687]. Inductively [1684]. Indus [1370]. Industrial [2021, 2785, 1447, 2745, 1767, 1511, 2444, 215]. Industrial-Scale [1511]. Industrialized [2979]. Industries [1700]. Industry [1347, 2490, 2209, 2111, 1158, 322, 2485, 439]. Industry-Specific [2490]. Inequitable [1624]. Inert [751]. Inertial [2660]. Infauna [1458, 1989]. Infaunal [597, 1389]. Infected [851, 2982, 3157]. Infecting [2716]. Infection [2802, 2889, 846, 700, 2016, 2831, 2888, 2181]. Infections [2899, 446, 475]. Inference [1941]. Inferences [1571, 935]. Inferred [542, 45, 1909, 636, 2996]. Inferring [839, 1107, 1092]. Inflammation [1256, 1115]. Inflammatory [2124, 300]. inflatus [916]. Inflow [19, 395]. Inflows [183]. Influence [1940, 286, 2325, 2371, 3082, 389, 2195, 434, 1920, 2080, 130, 1904, 2952, 2164, 304, 474, 2032, 2909, 1906, 1020, 2176, 1796, 689, 1383, 750, 724, 766, 597, 1368, 341, 559, 122, 2729, 1605, 2832, 1878, 1497, 1201, 1014, 777]. Influenced [2778, 2643, 1340, 879, 209, 1985, 1856, 781, 2594]. Influences [1053, 907, 2698, 185, 2942, 2484, 1914, 3152, 1819, 1234, 1029, 246, 2849, 17, 2007]. Influencing [1104, 2689, 2467, 285, 836, 2448]. Influx [279]. Inform [661, 1632, 1219, 2548, 1775, 67, 1833]. Information [2820, 203, 1445, 2607, 925, 1060, 1886]. Informed [257]. Infragravity [453]. Infrared [925, 2418]. Infrastructure [1580, 499, 2636, 1701, 2584, 1698, 2816]. Infrastructures [1642, 282]. infrequens [776]. Ingestion [2689, 2253, 628, 795, 2683, 2080, 1729, 2893]. Inhabiting [1309, 1550, 2068, 2504]. Inherent [1984, 2289]. Inhibit [3046, 2076]. Inhibiting [2866, 3127]. Inhibition [1951]. Inhibits [1751]. iniae [2214]. Initial [1186, 2673, 1888, 539]. Initialization [488, 1381]. Initiative [451, 1182]. Initiatives [2312, 731]. Injuries [1464]. Injury [1463]. Injustice [2654]. Injustices [1786]. INLA [2529]. Inlet [644, 1071, 21]. Inlet-Interrupted [644, 1071]. Inlets [2986]. Inline [1121]. Innate [951, 3027]. Inner [225, 2771]. Innovation [2885]. Innovative [1084, 1950, 1752, 711]. Inoculated [331]. Inoculation [509]. inopinus [1720]. Inorganic [308, 1093, 723, 94, 1673, 1971, 3140, 1500]. inornata [50]. Input [434, 2998]. Inputs [1662, 1502, 874, 309, 506]. Inquiry [1846]. Insect [2304]. Inshore [457, 2375, 2319, 1997]. Insight [2060, 501, 11, 2178, 1547, 1179, 1934, 2801]. Insights [346, 2324, 1787, 2775, 544, 557, 685, 124, 2772, 456, 196, 1044, 1601, 1125, 2900, 2186, 3070, 1818, 6, 92, 571, 523, 2183, 1132, 2961, 1731, 2399, 2302, 2709, 491, 2989, 3055, 358, 2924, 1514, 719, 54, 1880, 2674, 737, 2888, 378, 242, 2925, 2230]. Instabilities [2532]. Instability [1872]. Institutional [3042, 2490, 853, 2649]. Institutionally [538]. Instrument [2795, 1236, 356, 2464]. Instrumented [2092]. insuetum [2835]. Insular [2812, 1125]. Insulin [2114, 3127]. Insulin-Like [2114, 3127]. Integral [908]. Integrated [2310, 1071, 2991, 2203, 1787, 156, 258, 282, 2933, 1948, 2502, 499, 527, 41, 2119, 2127, 1570, 2440, 1544, 4, 2466, 2907, 2625, 2446, 2055]. Integrating [2051, 391, 2313, 2522, 1724, 2296, 2434, 3048, 3142]. Integration [1146, 2142, 2885, 2831, 1133, 2649]. Integrative [1052, 2003, 1100, 2941, 3023, 1383, 1907, 1304, 2762, 1203]. Integrity [770]. Integument [1646]. Intelligence [1257]. Intelligent [2140]. Intense [1183]. Intensities [2978]. Intensity [314, 2882, 2459, 403, 305, 1748, 2501]. Intensive [3098, 2781]. Intentions [1647]. Inter [1293, 1789, 1192, 2828, 1056, 2988]. Inter-Annual [2828, 1192]. Inter-Comparison [1293]. Inter-Individual [1056]. Inter-Monsoon [1789]. Inter-Specific [2988]. Interacting [2172]. Interaction [3144, 2092, 226, 736, 2433, 2149, 2090, 3186, 1656, 518, 613, 2830, 338]. Interactions [376, 188, 1329, 420, 2094, 2576, 1801, 1359, 2455, 728, 2534, 797, 1031, 929, 2082, 2743, 1220, 2497, 1015, 2857, 1946, 758, 1355, 194, 1428, 1791, 2675, 2505]. Interactive [1737, 3202, 2962, 2565]. Interannual [1865, 1137, 3100, 895, 782, 1995, 899, 1321, 1917, 1825]. Interannual-to-Decadal [899, 1321]. Intercalary [1039]. Intercellular [3094]. Intercept [1772]. Interconnectivity [1784]. Interdecadal [2960]. Interdependencies [2733]. Interdisciplinarity [1749]. Interdisciplinary [857, 1182, 817]. Interest [1577]. Interface [888, 376, 16, 2618]. Interfaces [2405, 2411]. Interference [2392]. Interior [2218]. Intermediate [1328, 1270, 456, 477]. Intermediately [2687]. Internal [2583, 3041, 2821, 2439, 604, 1671, 3117, 3196, 433, 125, 70, 1587, 2818, 2832, 1019, 1918, 1210, 2752]. International [643, 2401, 8, 1786, 84, 2067]. Internesting [78]. Internodal [2499]. Interoceanic [1592]. Interoperability [1578]. Interplay [1897, 655]. Interpretation [2112, 3106, 666, 2705, 400]. Interrupted [644, 1071]. Intersecting [2260]. Intersection [1150]. Intersectional [1624]. Interspecies [1568]. Interspecific [2853, 2967, 2416, 1879]. Intertidal [163, 2028, 314, 946, 1337, 787, 2421, 2393, 494, 1388, 3003, 1981, 352, 3081, 143, 1723, 2119, 2164, 1893, 1413, 1335, 622, 1836, 140, 623, 167, 694, 1275, 1154, 521, 2198, 1369, 1989, 1124]. Intervals [1145]. Intervention [2107]. Interventions [2246, 2394, 1343]. Interview [757]. Interview-Based [757]. Intestinal [2589, 1115, 2952, 2362, 2761, 2579, 2870, 300]. Intestine [2905, 2388, 1970]. Intra [2812, 2205, 2305, 973, 2416, 2988, 1879]. Intra- [2416, 2988, 1879]. Intra-Annual [2812, 973]. Intra-Group [2305]. Intra-Population [2205]. Intracapsular [2048]. Intracellular [930, 880]. Intraseasonal [1075, 3133]. Intraskeletal [490]. Intraspecific [1716, 2949, 816]. Intrinsic [899, 1321, 3121, 3073]. Introduced [325]. Introductions [2427, 1480]. Intrusion [527, 959, 2007, 1545]. Intrusions [2727]. Inuit [1416]. Inundation [648]. Invaded [3136, 2359, 2175]. Invariant [1389]. Invasion [2471, 2256, 1722, 1240]. Invasions [890, 3202, 1722]. Invasive [163, 2471, 3123, 303, 2043, 1480, 537, 2285, 2409, 3044, 2776, 2302, 77, 2541, 2588, 402, 2516, 881]. Inventories [336]. Inventory [49]. Inversion [2350]. Invertebrate [2071, 2413, 2691, 3114, 416, 481, 1881, 1431, 447, 844]. Invertebrates [2369, 1176, 59, 3161, 3122]. Inverted [2290, 1453, 2822]. Investigate [3144, 903, 48, 1426]. Investigating [2061, 669, 2131, 1965, 2839, 2555, 564]. Investigation [2334, 3014, 2247, 331, 149, 2627, 2817, 1417, 1984, 2304]. Investigations [2487]. Investing [1253]. Investment [2484]. Invisible [1282, 1686]. Involved [2639, 1505, 222, 1971, 825, 517, 2939]. Involvements [2806]. Involving [2345]. Iodine [723, 3171]. Ion [333]. Ionian [819, 1004, 2507]. Ionic [862]. Ionization [1684]. Ionocytes [811]. Ions [2945]. IoT [1896]. IPCC [1059, 1674]. iPSC [791]. IQuOD [2401]. Ira [2874]. Ireland [1792, 1490]. Iridocytes [567]. Irish [2144]. Iron [1293, 946, 580, 670, 1871, 553, 377, 37, 1661, 2838]. Iron-Limited [1661]. Iron-Stressed [580]. Irradiance [445]. irradians [1651, 2886, 1651, 2886]. Irradiated [2093]. Irradiation [3120]. Irrawaddy [850, 1556]. Irrigation [418]. Isfjorden [874, 686]. Isistius [1646]. Island [2286, 1616, 2991, 105, 3191, 936, 2730, 316, 1531, 1901, 3089, 2382, 1875, 1818, 120, 441, 2755, 2487, 1187, 2554, 1220, 1156, 82, 397, 3050, 1878, 2073, 2992, 1201, 1784, 1738, 2969, 1197, 2139, 768, 2926, 1807, 2693, 1778, 33, 1620, 1743]. islandica [2371, 2937]. Islands [3031, 146, 1956, 3124, 856, 2383, 1574, 2733, 1423, 3118, 1303, 2438, 1706, 2185, 2070, 2673, 2322, 1961, 792, 379, 176, 94, 2165, 2265, 1926, 3153, 241, 2261, 248, 37, 1336, 93, 762]. Iso [2514, 274]. Iso-Seq [2514, 274]. Isobaric [3038]. Isoform [2514, 1264, 274]. Isolate [1040]. Isolated [2214, 1464, 2623, 1106, 1972, 1863, 2109, 2263, 880, 2835, 2832, 2801, 1284]. Isolates [392, 1338]. Isolation [276, 1125]. Isomers [860]. Isopod [3045]. Isopora [1247]. Isoprenoid [2677, 860]. Isoscapes [698]. Isotocin [712]. Isotope [2150, 2116, 537, 2375, 710, 744, 2480, 1909, 2957, 1145, 2223, 836, 780, 1339, 2780, 1989, 777, 2996, 2503]. Isotopes [1268, 1750, 1254, 2403, 2880, 943, 309, 1191]. Isotopic [2520, 674, 1819, 1957, 916, 490, 154, 1482]. Israeli [1309]. issue [890]. Issues [702, 102, 1574, 2345, 1577]. Italian [50, 88]. Italy [1037, 2891, 1539, 484, 152, 2025]. Iterative [2822]. iTRAQ [3184]. iTRAQ-Based [3184]. Ivermectin [2300].

J [1501, 205]. J-OFURO3 [1501]. J-SCOPE [205]. Jack [226]. Jackson [342]. Jankowski [1629]. January [2755, 1861]. Japan [3095, 2998, 1075, 742, 1415, 2643, 3177, 1525, 2756, 1029, 2554, 964, 1825, 844]. Japanese [2349, 725, 47, 706, 2032, 1207]. japonica [2982, 3157, 47, 1028]. japonicas [2726]. japonicum [2125]. japonicus [1523, 2349, 2756, 3175, 706, 3038]. Jardines [13, 1098]. Jaw [1345]. Jellies [63]. Jellyfish [1808, 2583, 2641, 196, 2970, 2414, 791, 2988]. Jersey [328, 2211]. Jet [3024, 899, 1321]. Jets [2606]. Jewel [1650]. Jiaozhou [1817, 2893, 2951]. Jiulong [2646]. Job [2407]. johnsonii [2875]. Johnston [956]. Johor [3071]. Jointly [2270, 2696]. Jones [1951]. Journey [1705]. jubatus [1052]. julis [2455, 2908]. Junction [300, 2212]. June [2724, 1792]. Jurisdiction [3042, 367, 931, 2570, 144, 2702, 2123]. Jurisdictional [2477]. Jurisdictions [888]. Just [1106, 2144]. Justice [1760]. Juvenile [2387, 1052, 1975, 1115, 2515, 828, 3205, 2375, 524, 2134, 1149, 2537, 2349, 1579, 2967, 1286, 2362, 966, 3146, 2880, 449, 749, 901, 359, 1610, 1241, 2083, 707, 3141, 2392, 2388, 2489]. Juveniles [2494, 3045, 2297, 1762].

Kafrelsheikh [2688]. Kairei [3058]. Kajikia [1886]. kamoharai [1126]. Kara [2876]. Karenia [1491, 1792, 2635]. Karstic [1220]. Karyorelictea [3072, 1327]. Katsuwonus [998, 2845]. Kattegat [1674]. Kazakhstan [2298]. KCNK3 [3212]. Kelp [1420, 1678, 669, 2462, 853, 646, 1715, 2884, 1394, 2349, 177, 1168, 558, 796, 35, 2943, 2754, 2276, 126, 307, 2723, 703, 2888, 210]. Kelp-Associated [2276]. Kelps [842, 2975, 213]. Kelvin [2301]. Kemp [316, 403, 661, 1887, 355]. kempii [316, 1887]. Kenya [1832]. kenyoni [1689]. Kerguelen [3153, 848, 743]. Kermadec [397]. Key [2179, 1978, 708, 3123, 2365, 2993, 797, 1163, 1274, 285, 1701, 944, 200, 398, 2064, 2966, 2254, 1212, 2765]. Keys [1808, 1535, 677, 2319, 1702]. Keystone [1875]. kHz [1960]. Kidney [1971]. Killer [2089, 1099, 1909, 95]. Kilometer [3178]. Kilometer-Scale [3178]. Kimberley [1320]. Kinase [295, 2314]. Kinases [112]. Kinematic [705]. Kinetochore [859]. King [828, 2755, 2914, 441]. Kingdom [1831]. Kingfish [2264]. Kingklip [927]. Kinorhyncha [767, 2178]. Kiribati [861]. Kit [1999]. Kitefin [1731]. KJ [522]. KJ-2 [522]. Kleptoplastidic [2857]. Kleptoplasts [687]. km [50]. Knockdown [1285]. Knoll [10, 211]. Know [1722, 1107, 692, 402]. Knowing [2282]. Knowledge [765, 2662, 485, 2190, 3143, 3025, 1994, 2695, 1647, 610, 2346, 2566, 1561, 2697, 1862, 2228, 1195, 1909, 2338, 127, 344, 2558, 1508, 2336, 1352, 402]. Known [1324, 2192]. Kogia [2081]. Kölliker [1880]. Kong [2442, 1025, 1606, 2967, 1380]. Kongsfjorden [591]. Konkan [135]. Konosirus [636]. Korea [1081, 1953, 2979, 2141, 2855, 1907]. Korean [725]. Krill [2010, 1116, 603, 1452, 2615, 2160, 2625, 2177]. Krka [2066]. Kuakata [2797]. Kuroshio [2301, 2269, 782, 2598, 1586, 959, 2004, 899, 1321, 2727, 2007, 3093, 3073]. Kuruma [3038]. Kuwait [125]. KwaZulu [876]. KwaZulu-Natal [876].

L [3158]. L-Arginine [3158]. L. [215, 584, 1823, 3056, 1988, 2098, 300]. Lab [2492, 2485]. Lab-on-Chip [2492]. Labeling [1222, 1870]. Labelling [2918]. Labile [3192, 777, 1826]. Laboratory [1102, 258, 2973, 1872, 1105, 356, 1402, 3139, 2672, 715]. Labrador [1396, 10, 1416, 1418]. labrax [1067, 2239]. Labridae [2908]. Lack [765]. Lactic [1215]. lactis [2024]. Lactococcus [2024]. Lacuna [898]. Lag [2101]. Lagenorhynchus [921]. Lagoon [1892, 19, 383, 279, 320, 1037, 2693, 519, 3089, 2283, 2318, 3097, 2703, 2193, 2796, 248, 2891, 838, 74, 1918, 2969, 2674, 3047, 2633, 1911, 1902, 1089, 2481]. Lagrangian [575, 2112, 2842, 1683, 2331, 1390]. lagunensis [2667, 3097]. Lahille [1181]. Lake [1185, 2001, 1633, 1823, 2928]. Lakes [2231]. lalandi [2264]. lalandii [2658]. Laminaria [558]. Laminariales [2888, 1135]. Lamna [1610]. Lamnid [1571]. lanceolatu [2124]. lanceolatus [2595, 3146, 2870, 2388]. Land [1961, 2453, 1831, 930, 1662, 2618, 1918, 2658, 2951, 2830, 1317]. Land-Based [1961, 1831, 930]. Landfills [2624]. Landform [1554]. Landings [2021, 1579]. Landlocked [1823]. Landsat [2577]. Landsat-Derived [2577]. Landscape [1395, 1310, 2244, 1851, 3189]. Landscapes [2891]. Landscaping [774]. Lane [2528]. Lanka [2283, 1878]. Lao [2567]. Lapse [1986]. Laptev [3181, 1340, 879, 2828]. Large [1414, 2269, 638, 1927, 420, 1877, 206, 1901, 853, 1793, 1290, 72, 2863, 621, 2161, 2717, 1579, 283, 2, 1330, 2679, 2761, 2631, 2627, 934, 920, 2722, 1686, 1143, 1540, 2245, 858, 529, 2674, 1736, 2603, 574, 1062, 20, 2898, 3176, 2811, 3022, 1618]. Large-Scale [638, 1927, 206, 853, 621, 2631, 2627, 858, 1062, 1618]. Largemouth [299, 3065, 1199]. Larger [1750, 756]. Largest [3080, 1731, 2906, 3113]. Larimichthys [2761, 1941]. Larvae [2071, 2589, 3154, 353, 2486, 1345, 2469, 1068, 1017, 1742, 237, 1362, 2756, 1086, 475, 633, 1720, 2297, 941, 271, 2940, 1515, 117, 2832, 2712, 631, 705, 3111, 2528, 20, 3053, 2501]. Larval [2802, 2775, 1345, 1285, 2950, 3152, 474, 2270, 3178, 2581, 2681, 681, 2545, 2767, 2709, 359, 533, 2609, 2295, 360, 3076, 2935, 458]. Last [456, 1996, 1668, 2814, 2951]. Lasting [261]. Late [1703, 1299, 359, 2266]. Late-Larval [359]. Latent [2529]. Lateolabrax [2927]. Lateral [1141, 2303, 1371]. Lates [1920, 1515]. Latimeria [2989]. Latin [138]. latipes [3120]. latissima [2462, 1715, 930, 1168, 35, 1135]. Latitude [155, 1272, 271, 2341, 2295, 3083, 788, 2525]. Latitudes [2203]. Latitudinal [2462, 3063, 450, 694, 3165]. Latreille [2093, 1498]. latus [1744]. Lau [273]. Lava [131]. Law [2067]. Lawrence [485, 2808]. Layer [1116, 1530, 2828, 1717, 2727, 3133, 43, 2620]. Layered [847]. Layers [155, 2051, 1058, 1724, 3043]. LBB [734, 1169, 416, 266, 2287, 840, 609]. LC [2537]. Leach [56]. Leachates [1011]. Lead [765, 1539, 2285, 565, 298, 1112, 2101]. Lead-Lag [2101]. Leading [880]. Leads [215, 990, 1859, 917, 806]. Learn [1869]. Learned [2985, 2346, 2025, 1370, 894, 1459]. Learning [2787, 369, 1891, 1349, 2048, 2012, 1194, 531, 1695, 3008, 1723, 608, 1919, 2844, 1494, 2219, 701, 1784, 3129, 2216, 136]. Learning-Based [2048]. Learnt [634, 1519, 262]. Lease [1669]. Leatherback [1104, 3132, 78, 908, 2040]. Leaves [296]. Leaving [2230]. Lee [1291]. Legacy [2743]. Legal [1574]. Legged [1991]. Legislation [3144]. leithii [3105]. lemaneiformis [3158]. Length [2514, 2841, 2804, 734, 1264, 1169, 1626, 1840, 1051, 274, 2065, 1177, 2524, 3071, 609, 1303, 2696]. Length-Based [2804, 734, 1169, 609, 1303, 2696]. Length-Structured [2524]. Lengths [2281]. Lens [76, 1749, 2206, 2282]. Lenses [2334]. Lentiviral [2271]. Lentiviral-Transduced [2271]. leonina [697]. Leopard [2762, 3190]. leopardus [2762]. Lepeophtheirus [3052]. Lepidochelys [316, 1887]. Lepidopus [1118]. Leptobotia [1542]. Leptocephalus [1362]. leptolepis [2999]. leptonyx [3190]. Lesion [2687, 1216]. Less [1243, 616, 764]. Lessening [640]. Lesser [2973]. Lesson [2025, 1370]. lessoniana [351]. Lessons [800, 2985, 794, 2346, 1617, 213, 2637, 634, 1519, 894, 1459, 719, 262, 1774, 2442, 1173, 1180]. Leste [2228]. Lethal [2148]. leucas [1125, 2083]. Leucine [2639, 2423]. leucospilota [2860]. Levantine [2507, 1917, 658]. Level [2100, 1110, 1377, 1928, 895, 2933, 1622, 59, 976, 1830, 2266, 661, 2733, 2928, 441, 2363, 2542, 1438, 1582, 46, 2869, 2569, 2646, 2550, 2261, 583, 2548, 2171, 1888, 1701, 2306, 2977, 2817, 1900, 2090, 1898, 2448, 906, 865, 1908, 3010, 1674, 2870, 905]. Leveling [1786]. Levels [155, 215, 2229, 2242, 1550, 797, 2537, 1179, 573, 941, 365, 2624, 2193, 1966, 3159, 2676, 1306, 2538, 409, 517]. Leverage [1182]. Leveraging [1795, 2733]. lewini [1630]. LHCX [294]. Library [328]. Lice [1285, 3052, 829, 1533]. License [1005]. licha [1731]. licheniformis [2634]. Lidao [197]. Lidar [1032, 316, 933, 147, 595]. Lidar-Compatible [147]. Liepaja [2001]. Life [1378, 2663, 1486, 585, 776, 1329, 2272, 3115, 519, 2868, 2544, 877, 2599, 101, 558, 2088, 1419, 1733, 2052, 573, 2581, 977, 1038, 834, 398, 1092, 2093, 1443, 2852, 1214, 631, 1218, 940, 3174, 2528, 1685]. Life-History [1329, 2272, 2581, 977, 1038, 1685]. Lift [1624]. Lifting [2140]. Light [243, 1282, 1737, 314, 2800, 643, 294, 155, 2459, 1344, 2912, 445, 1653, 1031, 930, 2952, 2771, 1770, 1615, 1435, 2339, 2904, 2862, 431, 596, 1002, 1051, 2118, 2940, 483, 305, 2331, 2500, 1172, 81, 1732, 614, 2389, 1338, 2501]. Light-Dependent [1653]. Light/Dark [2952, 2904]. lignano [1955]. Like [507, 3192, 2791, 2114, 1151, 525, 3127, 2953]. Likely [243, 939, 736]. lima [671]. Limacina [53, 1398]. limacinum [161, 3111]. Limb [2212]. Limifossorid [2158]. Limit [2590, 728, 2004]. Limitation [1276, 699, 1420, 1002, 377, 2015, 2838, 42]. Limitations [2346, 513, 2534, 2406]. Limited [2799, 3025, 813, 379, 760, 3077, 3008, 846, 1026, 1957, 1866, 1107, 2058, 1661]. Limiting [2690]. Limits [2847, 488, 196, 1455, 2540]. Limnological [2012]. Limpet [1893]. Limulus [1278, 869]. Line [921, 1772]. Lineages [83, 1995, 2130]. Linear [2532, 1888, 1334, 566, 2675]. Lined [1352, 301]. Lingling [1522]. Link [2087, 3059, 1577]. Linkage [2655, 1062]. Linkages [3094, 2029]. Linked [605, 2227, 2897]. Linking [716, 552, 186, 891, 2232, 634, 478, 911, 1685]. Links [681, 1629]. Linnaeus [1102, 1127, 2859, 2048, 1345, 816, 2302, 2908, 3055]. Lion [2300]. Lionfish [537, 2175]. Lions [1052, 881, 1244, 1768]. Lipid [2840, 2912, 1728, 941, 2497, 2098, 2090, 2676]. Lipidomes [1425]. Lipidomic [3111]. Lipids [1449, 1720]. Lipolytic [941]. Liquid [276, 914]. List [2650]. Listening [3033]. Literacy [2247, 102, 261, 1910]. Literature [2202, 3197, 881, 2221, 1905, 1318]. Lith [2266]. Lithistid [1694]. Lithistids [1440]. Lithogenic [309]. Litopenaeus [2905, 2589, 3205, 2952, 2314, 662, 2372]. Litter [1095, 793, 2126, 2903, 1379, 835, 423, 410, 2543, 2194]. Little [101, 63]. Littoraria [1820]. Littorinidae [3105, 1820]. Live [1491, 1602, 2750]. Lived [2724, 2841, 309, 2058]. Livelihood [1739, 2397, 688]. Liver [1728, 465, 520, 3065, 3160, 2423]. Lives [519]. lividus [2708, 1897]. Living [66, 1639, 1655, 479, 1193, 2360, 536, 1985, 190, 2769, 1618]. lixula [2859]. Loach [1542]. Loading [1918]. lobata [1247, 1422]. Lobosphaera [512]. Lobster [2299, 1360, 1750, 2321, 2661, 2241, 1156, 2367, 3071, 1136]. Lobsters [56]. Local [2100, 1761, 1598, 813, 610, 589, 3161, 1000, 1523, 2228, 1582, 572, 2430, 1152, 2031, 689, 2059, 3186, 179, 2913, 1623]. Localization [859]. Localized [634]. Location [1761, 753]. Locations [1866, 2458]. Lockdown [1789, 2697, 1730, 1906, 2030, 2364, 1912, 2734, 2156]. Locomotion [1206]. Locomotor [2936]. Locus [2772]. Loggerhead [1846, 1690, 432, 310, 808, 323, 768]. Logging [2593]. Loliginid [1329]. Long [1042, 1470, 2605, 1845, 1823, 1817, 457, 1998, 1584, 3106, 2717, 24, 161, 1666, 850, 696, 2032, 1400, 1551, 1343, 2963, 2495, 1836, 2025, 941, 8, 84, 2250, 2101, 309, 1540, 876, 2496, 368, 1509, 1278, 3182, 1138, 2156, 2658, 190, 1810, 93, 1197, 2980, 768, 764, 2898, 2547, 3200, 3193, 2388, 1884]. Long-Chain [941, 2658]. Long-Lived [309]. Long-Period [1998]. Long-Read [3182, 2658, 2980, 2388]. Long-Snouted [1138]. Long-Standing [1509]. Long-Term [1042, 1470, 2605, 1845, 1817, 457, 1584, 3106, 2717, 24, 161, 850, 696, 2032, 1400, 1551, 1343, 2963, 2495, 1836, 2025, 8, 84, 2250, 1540, 876, 2496, 368, 1278, 2156, 190, 1810, 93, 1197, 768, 764, 2898, 2547, 3200, 3193, 1884]. longa [2716]. Longest [2841]. Longest-Lived [2841]. longimanus [1092]. longirostris [2555]. Longitudinal [103]. Longline [1559, 2717, 2184]. Longlines [801]. Longqi [3058]. Longshore [2606]. Look [1649]. Looking [1617]. Loop [1456, 1349]. Lophelia [315]. Loser [3029]. Losin [233]. Losing [1174]. Loss [2509, 2673, 1161, 2322, 2614, 2317, 2666, 2687, 1718, 770, 2319, 1043, 3185, 2844, 1216, 2180, 265, 2616, 880, 2833, 780, 2749, 426, 2723, 2311, 2107, 2442, 1702, 2139, 2953]. Lost [2145, 2410, 1229, 752, 2823]. Lota [2840]. Louisiana [111]. Low [172, 1414, 1081, 155, 771, 1811, 1903, 1236, 3045, 1149, 2161, 929, 1438, 2957, 1413, 271, 904, 3130, 2921, 1983, 3120, 2573, 2175, 2357, 3155, 2870, 1709, 1577, 1284, 2551, 2742]. Low- [2870]. Low-Cost [1236, 1413]. Low-Fe [1709]. Low-Frequency [2161, 1983].

Low-Gossypol [3130]. Low-Latitude [155, 271]. Low-Level [1438]. Low-Lying [3155]. Low-Tech [1413]. Low-Turbidity [1081, 2957]. Lowe [2908]. Lower [866, 976, 2212, 806, 1547, 1, 790, 1702, 1134]. Lowered [1485]. LPS [2016]. LSTM [1781]. LTER [2025]. LTER-Italy [2025]. lucidus [2446]. Luciferases [2230]. Lucinid [115]. lucioperca [2304]. Lucky [48]. Luminous [1731]. Lunar [2325, 2915]. Lungfish [2752]. luporinii [1304]. lusoria [2945]. lutea [991, 2644]. Lutein [3205]. lutris [1689, 437, 515, 684, 1409]. Luzon [1778]. Lyase [522]. Lying [3155]. Lymnaea [1255]. Lysate [1278, 869]. Lysine [1199, 2423]. Lysis [2121].

m [717, 1620, 131]. M. [2311]. M2 [3196]. M32A2M [2782]. Macao [2442]. Macaronesia [1956]. Machine [369, 2048, 2012, 3135, 1695, 1723, 1919, 2219, 136]. Mackerel [1815, 741, 3104, 1523, 2756, 226, 727]. Macondo [872]. Macro [1680, 2901, 1314, 2817, 1035, 2894, 118]. Macro- [1680]. Macro-Benthic [118]. Macro-Plastics [1035, 2894]. Macro-Scale [2901]. Macroalga [813, 1039]. Macroalgae [663, 778, 96, 880]. Macroalgal [939, 156, 1811, 2611, 590, 124, 1875, 2443, 785, 1918]. Macroalgal- [1811]. Macrobenthic [3108, 1926, 1756]. Macrobenthos [1783, 3003, 2192]. Macrobrachium [3127]. macrocephalus [1108]. Macrocystis [2723]. Macroevolutionary [2599]. Macrofauna [946, 1309, 1158, 1036, 506, 521, 2597]. Macrofaunal [550, 2827, 2995, 751, 1641]. Macroinvertebrate [1433, 2498]. Macrophyte [565]. Macrophytes [320, 891, 2013, 2923, 2738, 1355]. Macroplastic [114, 893]. Macrostomum [1955]. Macrotidal [2140, 804, 347]. maculatus [2927]. Madagascar [419]. MaDCrow [1580]. Madeira [2350, 2383]. Maërl [239]. Magallana [2878]. Magdalena [2318]. magister [101]. Magnetic [883, 1500]. Magnetometer [854]. Magnitude [7, 1346, 2307]. Mahi [1963, 1637]. Mahi-Mahi [1963, 1637]. Main [1367, 1706, 2185, 241, 2056, 2574]. Maine [2299, 69]. Mainstream [2443]. Mainstreaming [1222]. Maintain [2846]. Maintained [1406]. Maintenance [1316, 3161, 687]. Mainwaringia [3105]. Maize [3200]. Major [364, 281, 795, 1029, 2323, 1906, 616, 1734, 1019, 3149, 1720]. Makawa [2338]. Make [2246]. Makes [1945]. Making [579, 492, 12, 2557, 1632, 2552, 2690, 1605, 1144]. Mako [1601, 2368]. Malacca [137, 1899]. Málaga [1387]. Malaysia [28, 1335, 2022]. MALDI [1924, 1907]. MALDI-TOF [1924, 1907]. Maldives [2100, 856, 1878]. Maldivian [863, 689]. Male [945, 385, 2084, 2595, 3146, 2124, 2870, 2501, 2388]. Malformation [603]. Mallorca [2923]. Malta [1787]. maltophilia [522]. Mammal [1622, 958, 2836, 2067, 1623, 1556]. Mammals [218, 1437, 1491, 657, 2964, 2373, 1466, 3005]. mammata [2248]. MANACO [711]. Manage [1379, 1437, 1418, 1866]. Managed [2144, 2854, 1918]. Management [1499, 225, 34, 2739, 1472, 1807, 748, 805, 1350, 643, 2190, 1360, 1994, 741, 2089, 2621, 584, 2427, 2975, 2682, 2359, 1849, 2699, 557, 2346, 685, 129, 1769, 792, 1474, 1948, 429, 124, 2619, 2137, 1715, 1166, 1430, 1617, 931, 2795, 579, 1441, 2035, 2418, 527, 1279, 2477, 1004, 2318, 1615, 1697, 2259, 335, 2808, 2394, 1400, 2088, 1974, 966, 2803, 1798, 1219, 2300, 1487, 1701, 1775, 2632, 1060, 2152, 2850, 894, 1453, 1432, 127, 2210, 917, 592, 1593, 1705, 953, 2440, 1944, 995, 927, 711, 1459, 1691, 561, 1387, 1576, 1079, 1743, 3018]. Management [1980, 2600, 1784, 1611, 2553, 444, 768, 405, 678, 1886, 2505, 3142, 3136]. Management-Relevant [1944]. Managers [1257]. Managing [153, 2972, 1559, 3016, 335, 448, 721, 3, 1069, 3005]. Manatee [1026]. Manatees [1885, 2760, 1026]. Mandatory [1222]. Mandovi [1789]. Manganese [670, 2306]. Mangkhut [2433]. mangle [3002]. Mangrove [3080, 1892, 699, 2173, 2730, 1252, 3199, 2623, 429, 1036, 3057, 1852, 49, 798, 479, 1220, 1820, 1856, 2442, 1498]. Mangrove-Fringed [798]. Mangroves [2174, 3105, 1945, 2539, 479, 1554, 2608]. Manihiki [390]. Manila [3136, 2359, 2417, 893]. Manipulating [884]. Manipulation [46]. Manner [207]. Manta [827, 1801, 660, 2552]. Mantle [967, 1971, 2002]. Manual [1093]. Many [2901]. Map [649, 886, 2877, 2655, 548, 2914]. MAPKKK [295]. Mapping [2151, 373, 2728, 2400, 131, 486, 818, 3069, 1723, 2655, 69, 622, 169, 1910, 988, 1848, 623, 2690, 701, 1172, 211, 1870]. Maps [2795, 627]. Mar [849, 2481]. Maracaibo [1633, 1185]. March [1673]. Margin [2044, 2249, 2513, 1547, 1448, 3177, 2995, 113, 178, 2027, 2211, 192, 292, 559, 2200]. Marginal [1625, 1582, 2115, 1614, 2281, 3048, 1205]. Marginalised [579]. Margins [1673]. Marian [2755]. Mariana [2212, 2081]. Mariculture [824, 2643, 915]. marina [1892, 2499, 2670, 2623, 1031, 602, 3035, 865]. Marine [676, 1940, 885, 1499, 91, 308, 971, 702, 1095, 1123, 1053, 3046, 926, 2379, 2883, 267, 2226, 668, 2369, 805, 364, 2764, 2662, 586, 286, 1209, 888, 85, 2640, 250, 1921, 3042, 324, 1052, 3209, 857, 218, 464, 1787, 947, 2482, 320, 1176, 2113, 794, 1161, 1649, 704, 629, 1642, 2972, 234, 331, 1372, 793, 150, 1956, 1532, 1250, 2182, 3082, 270, 1936, 2155, 2126, 2249, 2413, 2782, 2903, 3110, 979, 166, 2012, 14, 231, 1849, 1480, 1379, 508, 2346, 3125, 2566, 2195, 299, 1194, 39, 282, 1308, 818, 1319, 2131, 1877, 1961, 1259, 1998]. Marine [1757, 3171, 1638, 2515, 1622, 1948, 2351, 2623, 11, 1759, 507, 2244, 739, 1316, 1507, 1437, 569, 1972, 3161, 1405, 1491, 1793, 502, 949, 1645, 116, 2021, 1430, 109, 958, 1981, 480, 1863, 931, 2593, 1853, 3012, 2374, 1591, 555, 257, 492, 2035, 1068, 2789, 666, 72, 1766, 2697, 2570, 1895, 2455, 3081, 3114, 877, 1575, 649, 565, 3134, 1155, 2677, 1516, 2836, 1393, 1246, 3070, 75, 499, 2993, 957, 624, 2971, 2901, 3016, 891, 1655, 657, 2047, 690, 3000, 1245, 2700, 2102, 2717, 2637, 296, 882, 2426, 71, 2013, 798, 1294, 135, 1953, 960, 1652, 283]. Marine [149, 3168, 2839, 2659, 1326, 1331, 2531, 634, 1887, 46, 73, 1519, 1565, 1528, 2078, 2, 1330, 608, 2313, 430, 2965, 1447, 1182, 2713, 1297, 558, 2874, 475, 1119, 3185, 2120, 706, 1974, 1391, 562, 2052, 2406, 3107, 2042, 3072, 1327, 2128, 2517, 912, 25, 2031, 835, 510, 1274, 578, 2803, 423, 1159, 2923, 1485, 965, 2154, 2705, 2399, 2297, 437, 2025, 2264, 1847, 1046, 2241, 1869, 677, 1383, 2964, 2405, 2149, 1238, 1775, 2648, 1403, 731, 1113, 1193, 1465, 1866, 1018, 200, 972, 2407, 398, 450, 268, 2193, 1966, 1999, 630, 2450, 3094, 3167, 709]. Marine [1724, 1628, 3118, 1453, 1432, 1694, 127, 344, 2834, 38, 817, 2518, 1827, 2373, 1191, 860, 1868, 1905, 1312, 2360, 540, 2411, 357, 2702, 2123, 2738, 1417, 380, 1677, 2354, 198, 2745, 1983, 1355, 995, 880, 1230, 3113, 1011, 2852, 2143, 249, 217, 1656, 902, 1129, 394, 15, 1133, 518, 3030, 2690, 2199, 1214, 170, 3121, 2649, 1154, 1620, 2067, 582, 820, 1576, 1258, 2500, 179, 328, 2194, 29, 848, 334, 2192, 2538, 751, 1241, 1431, 2723, 1180, 447, 319, 1466, 871, 332, 1764, 142, 2335, 2254, 1265, 2748, 1005, 992, 190, 2572, 1774, 826]. Marine [1623, 1556, 2278, 3005, 1348, 2444, 1498, 1597, 2613, 1874, 1369, 2434, 1173, 737, 1318, 65, 56, 1062, 3026, 3085, 3176, 2811, 2858, 1304, 2801, 3079, 2622, 1884, 3142, 2055, 1870, 2540, 3066, 2224, 601, 1741, 260, 1927, 420, 1430, 1290, 2327, 12, 2454, 2259, 2631, 1521, 100, 2632, 2536, 894, 2906, 2702, 3128, 1607, 2144, 2147, 1743]. Marine-Based [3082]. Marine-Related [1194]. marinus [700, 111, 2801]. Maritime [2309, 2136, 1180]. Marked [560]. Marker [506]. Markers [2108, 469, 1044, 1017, 2655, 1299, 2348, 796, 268, 1785, 3010, 618, 1542]. Market [2246, 2407]. Markets [3180]. Markings [1584]. Markov [428, 2822]. Marlin [1886]. Marmara [1229, 752, 1248]. Marsh [505, 2571, 2873, 2887, 2049, 598, 2891, 3155, 2588]. Marshall [93]. Marshes [942, 2164, 2516]. Marsupenaeus [3038]. martensii [3103, 1970]. Martínez [1500]. Marvelous [507]. Masirah [768]. Mask [1384]. Masking [1604]. Mass [1924, 484, 152, 1930, 456, 2085, 1254, 120, 1188, 714, 1684, 304, 1585, 1858, 467, 2754, 1035, 804, 347, 1878, 984, 1810, 1197]. Masses [2598, 341]. Massive [1309, 849, 780, 426]. Mat [461]. Match [3083]. Match-Mismatch [3083]. Matching [3107]. Mate [1316]. Material [676, 135, 2052, 1950, 766, 1399]. Maternal [2825, 1703, 3002, 307]. Matla [2196]. Mats [2729, 2674]. Matter [2596, 229, 2224, 1834, 1541, 1081, 1710, 2328, 1811, 867, 1250, 2942, 1252, 1789, 3063, 1547, 2066, 2045, 1302, 2643, 441, 1730, 2415, 1826, 3140, 2890, 1502, 3151, 874, 2995, 1503, 2104, 1151, 183, 1140, 1454, 506, 2227, 530, 122, 973, 1162, 1467, 777, 1082]. Matters [18, 2646, 1060, 289]. Maturation [3014, 2003, 2084, 978, 2939]. Maturity [2678, 2845, 1126, 3071]. Maud [1317]. Maui [2555]. Mauritania [2661]. mawsoni [2108]. Maxima [773, 996, 2293]. Maximal [411]. Maximization [500]. Maximize [1523]. Maximizing [2816]. maximus [1115, 2095, 1179, 2423, 300]. May [864, 1690, 2847, 403, 2590, 2082, 2016, 2175]. Mayo [418]. MBON [1018]. Me [261, 1072]. Meadow [602, 1240, 1082]. Meadows [172, 2693, 2467, 3091, 730, 2643, 221, 3211, 2788]. Meager [1068]. Meal [2589, 2595, 1115, 3130, 2304, 3141, 2388]. Meal-Induced [1115, 2388]. Mean [2100, 1146, 1791]. Meander [2269]. Meandrinidae [2169]. Measure [1602, 3001]. Measured [2138]. Measurement [312, 1552, 2405, 1399, 712]. Measurement-Based [2405]. Measurements [987, 2948, 704, 2400, 2592, 1006, 513, 445, 1990, 1967, 3137, 594, 595, 2370, 378, 242, 1260, 1890].

Measures [2089, 793, 2165, 236, 2410, 1427, 1079]. Measuring [2231, 2240, 2687, 3106, 2308]. Mechanism [2782, 1638, 2244, 59, 2292, 3049, 110, 2138]. Mechanisms [2802, 3060, 607, 1363, 2786, 1401, 2571, 1728, 1653, 51, 2896, 2961, 2303, 2458, 3121, 2175, 2907, 3184]. Mechanistic [1936, 447]. Med [2254, 722, 1929]. Medaka [1565, 3120]. Media [3067, 3087]. Mediate [3134, 567]. Mediated [2094, 96, 1226, 460, 2323, 1355, 123, 2907, 110]. Mediates [2726]. Medicine [3125]. Mediterranean [2507, 3031, 2334, 1378, 763, 2678, 223, 1118, 2085, 1309, 335, 2923, 1626, 2201, 724, 1694, 38, 2364, 2609, 477, 1581, 3036, 1917, 1482, 1328, 19, 1095, 3028, 577, 320, 588, 3203, 1127, 484, 152, 2859, 489, 2611, 1024, 1492, 2879, 206, 124, 3029, 2502, 1083, 2238, 722, 1929, 1517, 2439, 1087, 2285, 2455, 621, 1251, 1489, 221, 1519, 696, 1882, 3131, 297, 1796, 9, 2105, 1897, 2962, 2011, 950, 1898, 2429, 1307, 599, 679, 2331, 280, 1984, 2057, 2729, 1748, 1223, 561, 2526, 319, 2894, 3088, 2254, 2175, 781, 2822, 1287, 2516]. Medium [2033, 2215]. Medium-Term [2215]. Meet [1741]. Meeting [2413]. Meetings [2621]. Meets [2143]. Mega [1680, 3069, 1851]. Mega-Epibenthic [3069]. Mega-faunal [1680]. Mega-Nourishment [1851]. Megabenthic [1753, 2437, 256, 495]. Megafauna [91, 1009, 632, 942, 649, 2941, 1465, 1016, 1214, 332]. Megafaunal [2187, 1553]. Megahabitat [292]. Megalobrama [388]. Megalopae [205]. Megaptera [1235, 224, 3132, 1939, 2163, 2468]. Meiobenthic [2196, 582]. Meiofauna [2789, 302]. Meiofaunal [2200]. Meiotic [1951]. Melanoma [162]. melanura [2908]. melastigma [1565]. Melatonin [2820]. Melt [2435, 735, 1502, 715]. Melting [639]. Member [2310, 3212]. Members [111]. Membrane [880]. Memory [2101, 2977]. Menhaden [1472, 1270, 1474]. Menor [849, 2481]. Mercury [225, 1539, 1112, 236, 2811, 3176]. Meretrix [2920, 604, 2945, 618, 2757]. Merging [623]. Meridional [1830, 3095, 2212]. Meristems [1039]. Merluccius [2469, 1529, 417]. Meroplankton [2272]. Meso [2060, 283, 2817, 1452, 836]. Meso- [2060, 283, 836]. Meso-Macro [2817]. Meso-Zooplankton [1452]. Mesocentrotus [695]. Mesocosm [3188, 900, 1306, 2291, 372]. Mesopelagic [2968, 2329, 52, 873, 1024, 556, 1729, 933, 119, 2707, 2722, 1717, 3076]. Mesophotic [2812, 1937, 936, 1599, 1087, 2544, 1309, 1302, 1798, 1386, 413, 2610, 1962, 2139]. Mesopredators [652]. Mesoscale [575, 556, 1507, 782, 769, 1291, 2004, 2977, 546, 3004]. Mesozooplankton [2488, 90, 2824, 564, 810]. Messengers [512]. Meta [1539, 2074, 230, 1213, 1405, 1568, 384, 2833]. Meta-Analysis [230, 1213, 1405, 1568, 384, 2833]. Meta-Analytical [1539]. Meta-Organism [2074]. Metabarcoding [275, 336, 1783, 541, 2789, 2607, 1525, 2967, 2183, 2348, 1679, 3010, 1265, 2996]. Metabolic [172, 188, 2693, 1006, 3084, 1523, 1973, 318, 1165, 2750, 667, 409, 1212, 2826, 2258]. Metabolism [3099, 2920, 1920, 1097, 143, 2715, 2952, 2497, 991, 3130, 3187, 1495, 517, 3103, 3053, 1203]. Metabolite [1312]. Metabolites [2864, 1769, 3122, 415, 1684, 880, 95, 785, 1212, 1098]. Metabolome [2423, 2579]. Metabolomic [352, 1384]. Metabolomics [2666, 2537, 1693, 1312, 2907, 2762]. Metacommunity [2015]. Metacoupling [716]. Metagenomic [1744, 2674]. Metal [2208, 1761, 273, 2150, 1476, 2313, 1631, 2585]. Metalloid [1476]. Metallothionein [604]. Metals [819, 1054, 1004, 915, 514, 2881]. Metamorphosis [2408, 458]. Metapopulation [3179, 3178, 1667]. Metapopulations [3161]. Metatranscriptome [83, 1373]. Metatranscriptomes [1425]. Metatranscriptomic [541]. Metaurostylopsis [1242]. Metazoan [2018]. Methane [308, 663, 2774, 31, 3192, 2978, 1997, 645, 426, 178, 182]. Methane-Charged [308]. Methanotroph [3141]. Methionine [991]. Method [1052, 1891, 1349, 2459, 1804, 2041, 418, 333, 2804, 2792, 734, 1965, 1169, 1894, 416, 2862, 1413, 375, 2252, 2373, 1399, 2818, 2628, 1772, 266, 2287, 289, 136, 609]. Methodological [1903, 738, 1969]. Methodology [2941, 587, 1585, 169, 2518, 988]. Methods [1293, 1093, 2428, 218, 3025, 76, 1928, 2789, 1096, 2455, 2534, 2463, 923, 720, 634, 1913, 101, 2979, 725, 912, 1727, 578, 2661, 2236, 2298, 1113, 2062, 2337, 532, 2340, 1417, 3109, 2147, 401, 1772, 718, 742, 831, 844]. Methylation [1262, 53, 57, 967, 3202, 2653, 938, 307, 327, 2726]. Methylmercury [20]. Methylococcus [3061, 3141]. Métier [913]. Métier-Based [913]. Metric [1683]. Metrics [1435, 2947]. Mexican [268, 2188, 1333, 671]. Mexico [436, 1158, 2318, 418, 1846, 2974, 2684, 615, 2760, 903, 945, 451, 139, 403, 654, 2344, 1106, 196, 3179, 1601, 72, 1158, 432, 1221, 2009, 1589, 310, 151, 111, 3178, 1530, 204, 2652, 1798, 269, 681, 749, 119, 3090, 2149, 709, 359, 1454, 1627, 366, 158, 2040, 549, 775, 2924, 1952, 123, 2083, 1844, 1619, 3076, 254, 2138, 3142]. Miami [2458, 2749]. Miami-Dade [2749]. Michigan [2928]. Micro [1042, 1876, 298, 2901, 2839, 2894, 1468, 2481]. Micro- [2894]. Micro-Computed [1876]. Micro-Needle [298]. Micro-Phytoplankton [1042]. Micro-Plastics [1468]. Micro-Raman [2839]. Micro-Scale [2901]. Microalga [512]. Microalgae [1297, 991, 2291, 504, 1657, 291]. Microalgal [2484, 161]. Microbe [2576]. Microbenthic [143]. Microbes [946, 1936, 2063, 1973, 2399, 944, 1355, 2276, 2980, 2504]. Microbial [2479, 163, 3058, 3047, 1612, 2329, 1977, 1811, 867, 62, 1115, 1022, 1061, 42, 3156, 3063, 343, 1575, 565, 24, 1973, 1976, 1088, 1157, 461, 1788, 1727, 3151, 910, 2142, 2104, 1650, 626, 1827, 357, 368, 2777, 2574, 288, 1269, 1497, 2669, 1925, 1764, 211, 2674, 2434, 1124, 2881, 3204]. Microbiological [2153]. Microbiome [156, 303, 1600, 96, 1981, 1033, 2642, 2264, 1023, 1947, 368, 2015]. Microbiome-Dependent [1033]. Microbiomes [1855, 2337, 249, 2500]. Microbiota [1773, 2905, 2595, 2374, 2952, 2843, 2761, 2024, 2497, 2098, 1076, 3182, 2574, 1135, 2304, 2579, 1970]. Microchemistry [228]. Microcosm [83, 2463, 396]. Microcosms [24]. Microenvironment [2583]. Microenvironments [629, 149]. Microeukaryotes [541]. Microeukaryotic [83]. Microfiber [3135]. Microflora [1924]. Microfluidic [2918, 396]. Microfragments [1602, 1608]. Microgels [507]. Microglia [791]. Microhabitat [655]. Microlayer [2698, 1202, 2472]. Micronekton [810, 90, 551, 836]. Microorganism [1570, 972]. Microorganisms [2479]. Microornamentation [1313]. Micropaleontologic [2256]. Microphytobenthic [352]. Microphytobenthos [314, 494, 272, 797, 664, 431, 622, 598, 623, 1288, 521, 372, 535]. Microplankton [1224, 409]. Microplastic [1142, 1045, 2321, 2683, 3102, 1248, 1164, 2839, 73, 2646, 1729, 2221, 3030, 2420]. Microplastics [2507, 963, 1027, 2330, 231, 3049, 1155, 2409, 2109, 914, 2779, 2414, 2862, 2679, 375, 3131, 1999, 2320, 1427, 2987, 1748, 3020, 3200, 3043]. Micropterus [299, 3065, 1199]. Microrefugia [2054]. MicroRNAs [32]. Microsatellite [2108, 2725, 618]. Microsatellites [268]. Microscale [287, 664, 756, 1686]. Microscopic [1462]. Microscopy [140]. Microscopy-Based [140]. Microsporidian [2792]. Microstructural [1314]. Microstructure [2085, 631]. Microzooplankton [994]. Mid [1994, 48, 868, 2646, 2910, 592, 980, 686, 254, 2525, 91, 1102, 2266, 670, 1673]. Mid-Atlantic [1994, 48, 980, 91, 670, 1673]. Mid-Late [2266]. Mid-Latitude [2525]. Mid-Level [2646]. Mid-Ocean [2910, 592]. Mid-Shelf [254]. Mid-Summers [686]. Mid-Western [1102]. Mid-Winter [868]. Middle [2042]. Midwestern [3024]. MIG [2545]. MIG-seq [2545]. Migration [2325, 1469, 873, 2760, 654, 1939, 2902, 36, 310, 461, 2040, 167, 123]. Migrations [2051, 3115, 1540, 535]. Migrators [168]. Migratory [656, 314, 516, 1020, 1540, 1698, 1572]. Mile [2293]. miles [537]. Milk [1212]. Millennial [2814]. Millennial-Scale [2814]. Millepora [1373, 1384]. Million [802]. Milne [2658]. Mimic [1950]. Min [609]. Mind [1549]. Minds [2347]. Mine [751]. Mine-Tailings [751]. Mineral [1704, 1664]. Mineralization [1710, 1871]. Mini [771]. Minima [866]. Minimal [996]. Minimize [1913]. Minimum [104, 867, 462, 2715, 635, 3033, 211]. Mining [560, 237, 349, 934, 2585, 1717, 592, 3194, 2997, 1784, 1641]. MinION [550]. MinION-in-ARMS [550]. Minireview [2369]. Minke [2034]. minor [1370, 118]. Mira [2592]. miRNA [1073, 2127]. Mirounga [697]. Mischief [1336]. Misidentified [2878]. Misinterpreted [108]. Mismatch [1766, 3083]. MISNAC [184]. Misorientation [808, 323]. Misperceptions [765]. Missing [2768, 2042, 1724, 1577]. Mission [2294, 2099]. Mississippi [757, 2026]. Miter [840]. Mitigate [801, 3211, 2788, 1935, 115]. Mitigates [974]. Mitigating [1514, 1049]. Mitigation [153, 370, 1638, 922, 2899, 673, 2569, 2819, 965, 972, 387, 280, 1691, 3033]. Mitochondria [2927]. Mitochondrial [1233, 542, 1110, 2108, 569, 1044, 1887, 1200, 816, 880, 2878, 3010, 3121, 1258, 197, 2925, 935]. Mitogen [295]. Mitogen-Activated [295]. Mitogenome [542, 2878]. Mitogenomic [1655, 571]. Mitten [263]. Mix [55]. Mixed [986, 654, 1550, 3057, 1652, 634, 1530, 1603, 1762, 2275]. Mixed-Bottom [1652]. Mixed-Species [2275]. Mixing [2532, 961, 2645, 2218, 1254, 2478, 1589, 1909, 2288, 2598, 3117, 304, 1530, 756, 1339]. Mixoplankton [2106, 1129]. Mixotrophic [968]. Mixotrophy [247, 1821, 409, 154]. Mixtures [2651]. Moana [2874]. Mobile [2096, 823, 1447, 3118]. Mobility [3030]. Mobilizes [1040]. Mobula [660]. Modal [1870]. Mode [2882, 522, 2326]. Model [308, 651, 2787, 1071, 2764, 682, 607, 331, 2845, 1270, 2126, 1804, 2691, 1675, 895, 1474, 1670, 1507, 1838, 862, 1430, 2899, 2973, 1237, 1301, 1375, 2873, 3015, 1635, 1522, 1419, 1902, 573, 3138, 1219, 1451, 2405, 2000, 173, 205, 1051, 1059, 1648, 1023, 532, 227, 1705, 280, 1077, 1908, 2734, 642, 2524, 81, 1381, 98, 1390, 3129, 1732, 1210, 2207, 1203, 2808]. Modeled [1890]. Modeling [963, 3163, 644, 2991, 1846, 232, 1956, 615, 979, 2046, 1167, 124, 922, 2383, 2312, 1564, 2001, 1407, 746, 833, 1066, 918, 527, 41, 1279, 1803, 534, 2873, 1697, 1396, 932, 2270, 3107, 875, 1667, 622, 3112, 2954, 67, 2146, 2331, 1797, 1922, 637, 122, 623, 2447, 1479, 1041, 1172, 743, 362, 2600, 123, 1839, 2894, 1556, 2529, 2528, 405, 843, 1618, 2608, 2823]. Modelling [2100, 2248, 2121, 839, 2808, 2487]. Models [2251, 3119, 800, 672, 1712, 716, 3082, 3136, 2359, 260, 166, 552, 2390, 1793, 746, 2656, 3126, 1594, 1909, 1632, 1122, 1314, 2591, 599, 1604, 533, 2065, 906, 1339, 1611, 737, 400]. Moderate [1262, 2065, 518]. Modern [957, 161, 711]. Modes [2880, 1735]. Modification [1414, 2726]. Modifications [2085]. Modified [527, 751]. Modifies [991]. Modify [1275]. Modular [1531, 892]. Modulate [2448]. Modulated [1788]. Modulates [2800, 2494, 35, 2777, 1495]. Modulating [765]. Modulation [1972, 3117, 2761, 2497, 1563, 2305, 2750]. Modulations [782]. Mohn [592]. mokarran [1663]. Molecular [2334, 324, 1250, 2575, 2003, 1351, 3049, 3157, 1163, 1728, 1655, 1653, 441, 1195, 1693, 620, 237, 1152, 803, 2550, 1654, 1261, 3072, 1327, 1175, 1159, 2298, 3159, 1785, 530, 2458, 1138, 1258, 1467, 1837, 1207, 3193, 2372, 935]. Molecular-Level [441]. Molecule [2388]. Molecules [3046, 2791, 712]. molitrix [2603]. Mollusca [3108, 2439, 2791, 450, 2158, 1820, 1080, 1356]. Molluscs [3075, 2650]. Mollusk [2937]. Mollusks [1992, 3020]. Molting [1285, 2314]. Monanchocidin [162]. Monanchora [162]. Mondego [2592]. Monitor [1580, 1792, 1413, 747]. Monitored [3067]. Monitoring [2736, 842, 85, 2070, 218, 1186, 1065, 2112, 369, 550, 2051, 1930, 2048, 189, 1961, 1998, 755, 2753, 1491, 3106, 1415, 1267, 94, 12, 2315, 349, 912, 2153, 1916, 1585, 1722, 2069, 1018, 2398, 1540, 2738, 2296, 259, 2543, 2526, 1743, 1844, 1581, 3033, 1034, 1747, 2547, 3068, 1010]. Monitors [989]. Mono [356]. Mono-Specific [356]. monoceros [1206]. Monocular [2419]. Monoculture [3057, 658]. Monodon [1206, 1073, 274, 3111]. Monooxygenase [2806]. Monopile [529]. Monoraphidium [1040]. Monourostylopsi [1242]. Monsoon [1789, 33, 2598, 1878, 617, 2814, 2830]. Monsoon-Driven [2598, 2814]. Monsoonal [750]. Montastraea [2509, 2666, 2687, 830, 2053, 2107]. Monte [2822, 532]. Monterey [63, 2668, 1983, 245]. Monthly [637]. Monument [1868, 668]. Moon [2915]. Mo’orea [120]. Moored [823, 1139, 3139, 2701]. Mooriganga [324]. Mooring [1990]. Mopping [2654]. Moray [2280]. morhua [667, 2503]. Morpho [656]. Morpho-Genetic [656]. Morphodynamic [2930, 1803]. Morphofunctional [1735]. Morphogenesis [1323]. Morphogenetic [2939]. Morphological [427, 2663, 1876, 1481, 2575, 856, 1351, 1163, 2363, 1888, 2908, 463, 1422, 3096]. Morphologically [1227]. Morphology [2586, 669, 744, 1537, 717, 803, 1327, 1090, 2740, 3072]. Morphometric [3071]. Morphotype [1329]. Mortality [2688, 1885, 1463, 421, 1571, 484, 152, 2564, 133, 839, 3188, 1689, 2129, 1188, 1551, 1122, 2581, 1585, 2300, 684, 956, 2964, 2754, 2749, 392, 3155, 2181]. Mosaic [479]. Most [1143]. Mother [3050]. Mother-of-Pearl [3050]. Motile [461]. Motility [1713, 2840, 59, 1248]. Motion [3123, 1903, 1435, 2591, 3162]. Motions [2316]. Motivations [586, 624, 2665]. Motor [1851]. Mound [456, 3036, 733, 592]. Mounds [10]. Mounted [2268, 2738, 1121]. Mouth [224, 2860]. Mouths [3176, 2811, 2606, 893]. Move [1617, 257]. Move-on [257]. Movement [351, 1208, 2034, 2902, 1601, 2721, 698, 598, 525, 1610, 2381, 2101]. Movements [2324, 864, 1221, 2477, 966, 1957, 749, 660, 2906, 2368, 1854]. Moving [141, 510, 1786]. Mozambique [767, 2355, 1366, 2602]. MPA [1901, 233]. MPAs [1797]. mRNA [2127, 2759]. MS [2537, 2979, 1907]. MS-Based [2537]. MS/MS [2979]. MSC [1468]. MSP [2764]. MSY [532]. MtDNA [468, 636]. MtDNA-Based [468]. Much [2055]. Mucormycosis [657]. Mucosal [2264]. Mucus [2414, 871]. Mud [946, 2651, 2003, 2772, 851, 343, 2178, 1976, 2866, 9, 2105, 2416, 3071]. Mudbanks [2197]. Mudflat [946, 1337, 787]. Mudflats [352, 598]. Mudskipper [1335]. Mugil [268]. Mulching [3200]. Müller [1125, 884]. Multi [1892, 748, 156, 2182, 2955, 258, 1680, 1531, 1430, 2863, 3081, 1755, 1237, 3134, 3008, 516, 1519, 1289, 1419, 479, 2827, 1147, 356, 1775, 698, 2508, 1059, 2591, 2373, 2653, 506, 2447, 2749, 1172, 2445, 1431, 2332, 847, 1404, 2686, 1304, 1870, 164]. Multi-Approach [1892]. Multi-Biotic [2827]. Multi-Channel [2955]. Multi-Criteria [1775]. Multi-Decadal [516]. Multi-Ecosystem [1059]. Multi-Element [698]. Multi-Ensemble [2332]. Multi-environment [1680]. Multi-Faceted [3134]. Multi-Factors [2445]. Multi-Gene [1304]. Multi-Indicator [3008]. Multi-Instrument [356]. Multi-Layered [847]. Multi-Marker [506]. Multi-Method [2373]. Multi-Modal [1870]. Multi-Model [1430, 1419]. Multi-Purpose [1531, 2447]. Multi-Scale [3081]. Multi-Sensor [1172]. Multi-Species [2863, 1237, 1289]. Multi-Stressor [1755, 1431, 1147]. Multi-Temporal [2508, 1404]. Multi-Tissue [2653]. Multi-Trophic [156, 258]. Multi-Use [2182, 1519, 164]. Multi-Variate [2749]. Multi-View [2591]. Multi-Year [748, 479, 2686]. Multiannual [561]. Multibeam [2390, 1960]. Multichannel [147]. Multidecadal [2542, 2630, 2937]. Multidimensional [2217]. Multidisciplinary [903, 740, 980, 1514, 3068]. Multifaceted [908, 3128]. Multifunctionality [926]. Multigene [110]. Multigenerational [573]. Multilevel [2711, 2845, 2731]. Multiparametric [2592]. Multiphase [1545]. Multiple [1954, 1300, 515, 888, 1845, 1975, 615, 391, 1757, 2670, 2576, 1695, 1394, 587, 675, 2718, 1965, 1299, 1435, 2119, 1632, 1460, 816, 2348, 2635, 1271, 2062, 398, 2282, 518, 170, 1985, 1012, 184, 2493]. Multiple-Use [1757, 170]. Multipotent [2820]. Multiscalar [716]. multiseries [2082]. Multispecies [2732, 227, 1705, 2524, 843]. Multispectral [2453, 2262, 622, 2738]. Multivariable [235]. Multivariate [1489, 750, 2062]. Municipal [1638]. Munida [1750, 56]. Muscle [2586, 1819]. Muscular [1362]. Mushroom [231]. Musical [1959]. Mussel [19, 2710, 222, 3188, 2843, 952, 1159, 2962, 450, 1051, 2831, 2987, 3121, 2008, 458]. Mussels [2150, 807, 652, 2994, 2409, 2430, 1056]. mutabilis [2048, 2623]. Mutant [392]. Mutation [522]. Mutualism [115]. My [2443, 1072]. Mycobacterium [484, 152]. Mycotic [446]. Myctophidae [119]. mydas [3011, 1476, 75, 2009, 749, 775]. mykiss [3027, 2777]. mysticetus [2619]. mystus [3078]. Mytilid [2430]. Mytilisepta [3188]. Mytilus [821, 939, 807, 473, 222, 1017, 2843, 974, 952, 1159, 2962, 1056, 2987, 3121, 1258, 2008, 458]. Myxocyprinus [2514]. Myxozoa [1713].

N [2700, 1242, 1304, 1991, 2641, 591, 3110, 2456, 1770, 236, 491, 2937]. N2 [2078, 2281]. N2-Fixing [2281]. NAAMES [971, 1209, 128, 2019, 317]. Naive [1621]. Namibia [921]. nana [2484]. Nannochloropsis [207, 180]. Nano [862, 1113, 2481, 2007]. Nano- [1113, 2007]. Nanoflagellates [990, 247]. Nanofragmentation [1063]. Nanoparticle [2586]. Nanoparticles [2131, 862, 3049]. Nanoplastic [2615]. Nanoplastics [1897]. Nanopore [550]. Nanotubes [3094]. Nansha [1336, 1303]. NAOS [1057, 223, 1055]. Naphthalenes [2979]. Naples [594]. Narragansett [2577]. Narrative [2282]. Narrow [3104]. Narrow-Barred [3104]. narutobiei [3092]. Narwhals [2006, 1206]. Nassau [259]. nasus [1044, 1610]. Natal [760, 360, 876]. Nation [3050]. National [1376, 668, 264, 3042, 947, 367, 931, 2570, 144, 2265, 677, 3118, 2702, 2123, 953, 1983, 1387, 1154, 1621, 1236, 1305]. Nations [861, 2190]. Native [2471, 77, 2508, 514, 2175, 1428, 402, 468, 2823, 2923]. Natura [2335, 2165]. Natural [1376, 3039, 1463, 1571, 1253, 3091, 3081, 1520, 46, 1119, 1974, 1122, 2554, 2634, 2307, 2011, 2429, 147, 1387, 1265, 1074, 1493, 2181, 551]. Naturally [1811, 1086, 425]. Nature [2273, 2539, 2638, 2259, 3001]. Nature-Based [2539, 2638, 3001]. Naturelle [1721]. Nauplii [603]. Naval [2235, 3114]. Navigate [2282]. Navigating [1877]. Navigation [2189]. Navy [2034]. NE-Atlantic [2248]. Neap [664]. Near [500, 224, 203, 2328, 2437, 795, 59, 2321, 2475, 1217, 2660, 2129, 120, 1447, 2303, 1389, 1543, 2524, 27, 2418]. Near-Automatic [500]. Near-bed [2475, 1217]. Near-Equatorial [1543]. Near-Field [2303]. Near-Future [59, 1389]. Near-Inertial [2660]. Near-Infrared [2418]. Near-Real [2524].

Near-Shore [2129]. Nearby [131, 2321]. Nearshore [1808, 1865, 411, 369, 1843, 749, 684, 1193, 2924, 426, 2720]. Necessary [2175]. Necropsied [515]. Necropsy [605]. Necrosis [3061]. Need [1561, 2704, 1711, 2119, 965, 2585, 3032, 1230, 2694, 975]. Needed [1755]. Needle [298]. Needs [2872, 2137, 1555, 2636, 127, 1417, 1707]. Negaprion [1540]. Negative [713, 974, 2016, 1428, 905]. Negatively [2710, 1285, 458]. Neglect [990]. Negombo [2283]. Negotiations [1719, 1318]. nehereus [2780]. Nekton [647]. Nematode [3134, 582]. Nematodes [674, 693, 2355, 1366, 1679, 2360]. NeMO [1870]. NeMO-Net [1870]. NEMOVAR [1077]. Neodymium [1254]. Neonate [1104, 1703]. Neophocaena [1025, 1370]. Nephrocytes [1971]. Nephrops [2321]. Nephtheidae [901]. nereis [515, 684, 1409, 437]. Neritic [1289]. Nest [316, 808, 323]. Nested [2795]. Nesting [1104, 2070, 3101, 2906, 775, 1734, 768, 3011, 661]. Net [329, 107, 143, 2700, 221, 3023, 1501, 2644, 1325, 1870]. Netarts [1160]. Nets [829, 876]. Network [1470, 2605, 2310, 3066, 2182, 1992, 1492, 2691, 1759, 722, 1929, 3179, 746, 2165, 2477, 2102, 321, 934, 8, 2215, 1018, 380, 2335, 2701, 2759, 2050, 2391, 650, 84]. Network-Based [321]. Networks [2379, 1741, 1781, 1194, 1638, 2709, 3162, 2354, 832, 2686]. Neural [1470, 2605, 1781, 722, 1929, 321, 3162, 2686]. Neuroendocrine [2408]. Neuroimaging [883]. Neurons [1255]. Neuston [1268, 1636]. Neutral [105]. Neva [1021, 2837]. Newborn [740, 2380]. Newfoundland [1396, 1418, 360]. Newly [3105]. News [713]. Next [436, 886, 2103, 774, 2553]. Next-Generation [774]. Nexus [1873]. NF [330]. NF- [330]. Nha [1265]. ni [2338]. nicaraguana [3075]. Niche [2028, 1938, 744, 524, 2227]. Niche-Relationships [2028]. Niches [2623, 481, 318, 3118]. Nickel [208]. Nicotinic [112]. Nicoya [809]. Niger [1276]. Night [1986, 591, 238, 1168, 808, 431, 323, 1860]. Nighttime [52]. Nile [2586]. niloticus [2586, 465]. Nine [1482]. Ninety [3064]. Ningaloo [685]. Ningde [2538]. Niño [1933, 3054, 40]. nipponica [602]. NIS [1722]. Nisin [2024]. Nisin-Producing [2024]. Nitrate [2301, 2715, 2045, 1826, 633, 2326, 2069, 3159, 180, 1307, 252]. Nitrate-Enrichment [1307]. Nitrate-Utilizing [2715]. Nitric [458]. Nitrification [1280, 491]. Nitrifying [2063]. Nitrogen [699, 1456, 807, 1039, 418, 2715, 1819, 930, 1234, 2082, 512, 1103, 948, 2880, 1788, 1727, 2432, 2087, 3002, 780, 3187, 667, 2969, 1661, 1864, 1203]. Nitrous [1496]. Nitzschia [580, 2082, 3097]. No [998, 1741, 713, 1901, 917, 525]. No-Take [1741, 1901]. NO-TNFs [1256]. NOAA [6, 92, 3033]. nobilis [484, 152, 1188, 2725]. Noble [2725]. Noctiluca [2563, 3137, 993, 2663]. Nocturnal [2712]. Nodavirus [2181]. nodosa [1240]. Nodule [560, 349, 2068, 3194]. Nodules [1747]. Noise [2243, 629, 2089, 2240, 2079, 2552, 1604, 1983, 245]. Nomenclature [1583]. Non [2387, 1009, 2711, 2532, 1927, 2041, 890, 251, 438, 881, 1920, 2983, 1915, 3189, 1334, 566, 2923, 543, 244, 280, 2835, 2526, 2370, 2759, 3078, 2675, 2925]. Non-archetypal [438]. Non-breaking [251]. Non-coding [3189, 2759, 2925]. Non-complex [2983]. Non-compliance [1927]. Non-destructive [2041]. Non-endangered [1915]. Non-indigenous [280, 2526]. Non-invasive [881]. Non-issue [890]. Non-Linear [2532, 1334, 566, 2675]. Non-native [2923]. Non-paralytic [2835]. Non-photochemical [2370]. Non-protein [1920]. Non-random [2387]. Non-reef [1009]. Non-representative [2711]. Non-song [244]. Non-spirolide [2835]. Non-stationary [3078]. Non-tuna [543]. Nonapeptides [712]. Nonlinearity [1672]. Nordic [659, 234, 173]. Normal [1688, 2854]. Norms [3165]. Noronha [1436]. North [2596, 2044, 329, 1849, 1401, 1841, 1580, 1636, 128, 2019, 2902, 2465, 1130, 2333, 1819, 106, 1671, 966, 896, 2495, 526, 2947, 3118, 1129, 1479, 1610, 2649, 2281, 1572, 1935, 788, 764, 2523, 971, 2532, 3060, 1209, 1360, 1612, 2482, 2847, 1771, 131, 1492, 1308, 895, 121, 1782, 2344, 2111, 1183, 1254, 2378, 1568, 760, 2063, 2363, 460, 1196, 2313, 2713, 2546, 2909, 2020, 2685, 1957, 375, 1314, 356, 1238, 1775, 9, 2105, 834, 1273, 3137, 915, 1694, 341, 2766, 1946, 1604, 2227, 1080, 1356, 89, 2734, 2922, 606, 1777, 1269, 2156, 2526, 2506, 2422, 1917]. North [843, 614, 1665, 378, 242, 1062, 2996, 3004, 1468]. North-Atlantic [1196]. North-East [2532, 2847, 2111, 834, 1080, 1356]. North-Eastern [1841, 2063, 1694]. North-to-South [788]. Northeast [976, 2700, 1293, 2179, 668, 273, 1360, 741, 1669, 2165, 3168, 1699, 1131, 727, 132, 3083, 459]. Northeastern [2837, 1557, 963, 1410, 1198, 432, 2643, 796, 2706, 514, 2924, 199, 840, 2449]. Northern [2203, 2150, 713, 1584, 1689, 1416, 1152, 2020, 1699, 2025, 911, 1239, 2431, 2324, 390, 2174, 1994, 1566, 615, 2692, 2786, 2760, 3132, 139, 986, 2821, 1047, 3179, 1763, 2837, 962, 2166, 2014, 1163, 1272, 2660, 1589, 2561, 1529, 106, 151, 2756, 3196, 3178, 959, 2703, 3072, 1327, 119, 2241, 1240, 2943, 1682, 1454, 1452, 366, 158, 549, 253, 980, 613, 2720, 781, 1756, 3076, 2007, 1369, 417, 788, 3133, 1028, 1860, 546, 3200, 2696]. Northernmost [1284]. Northwest [1311, 1778, 657, 2192, 2437, 3062, 541, 1582, 246, 999, 1522, 2425, 2995, 205, 532, 1656, 2083, 718, 468]. Northwestern [2288, 111, 1773, 2009, 1662, 896, 749, 724, 70, 775, 1776, 254, 476]. norvegicus [2321]. Norway [495, 240, 1749, 986, 30, 1533, 1944]. Norwegian [873, 2431, 1395, 373, 2321, 434, 2451, 754, 1714, 2922, 315]. NOS/NO [1256]. Notable [2125]. Note [1324]. Notes [1537, 1837, 1590]. Nothoholosticha [1304]. Nothoholostichinae [1304]. Nourishment [1851]. nov [1972, 1261, 1341, 2801]. Nova [2243, 1116, 694]. novaeangliae [1235, 224, 3132, 1939, 2163, 2468]. Novel [821, 2435, 2919, 45, 2716, 2864, 522, 1804, 544, 438, 57, 605, 2772, 710, 1695, 222, 2900, 2186, 1264, 1525, 2970, 1324, 921, 2405, 2522, 691, 2707, 2263, 2062, 2252, 274, 2474, 2518, 759, 1364, 1003, 3010, 536, 2835, 2157, 1514, 703, 3079, 2372, 2925]. novo [1567, 2810, 1542]. NTOU1 [475]. Nuclear [1955, 2538, 197]. Nucleic [184]. Nucleocytoplasmic [920]. Nucleotide [998, 47]. Nudibranchs [2439]. nudus [2537, 695]. Nuf2p [859]. Nui [2382, 2555]. Number [2543]. Numerical [2787, 2061, 2944, 1675, 3112, 2706, 1408, 906, 1121]. Nurseries [2168]. Nursery [2375, 2176, 449]. Nutraceuticals [2864]. Nutrient [2286, 427, 1093, 383, 279, 39, 951, 552, 1061, 722, 1929, 1430, 797, 930, 2415, 1770, 1684, 736, 874, 2995, 2521, 1432, 3130, 2793, 1040, 126, 2657, 1918, 1457, 521, 1199, 889, 2476]. Nutrient-Sensing [1199]. Nutrients [1093, 1781, 94, 2066, 2643, 1485, 764]. Nutritional [713, 524, 1012]. Nutritive [299, 2484]. NW [3031, 1643, 2046, 2335, 2851, 599].

O [2641, 351, 2456, 2700]. O&M [1144]. O. [2311, 884]. O2 [2474]. Oak [1925]. Ob [2876]. obesus [1101]. Object [531]. Objectives [1474, 3107]. obscurus [921]. Observation [1236, 2140, 1723, 159, 2213, 8, 84, 1018, 2141, 1501, 832, 2207]. Observational [570, 3147, 11, 1587]. Observations [3163, 2294, 2428, 1566, 1146, 453, 1903, 282, 1507, 2085, 128, 2019, 1755, 1671, 2569, 696, 3023, 412, 559, 1604, 1908, 2734, 1003, 1479, 2818, 984, 799, 802, 1134, 3073, 1884, 3022]. Observatories [653, 1182]. Observatory [2874, 423, 530]. Observed [971, 1560, 2821, 1673, 2512, 729, 1089, 2835, 1457, 3133]. Observers [584]. Observing [3039, 451, 2312, 1896, 2959, 2485, 2558, 3013, 2852]. Obstacles [794]. Obtain [3067]. Obtained [3039, 1782, 2924]. Occurrence [2179, 1281, 2786, 1669, 2469, 1831, 2409, 1221, 714, 2555, 1899, 205, 2064, 2341, 995, 2197, 2191, 3200, 2387, 484, 152, 2560, 1548]. Occurrences [311]. occurring [481]. Ocean [2431, 1873, 1233, 2535, 971, 1616, 2044, 651, 2251, 779, 2787, 800, 1834, 2294, 104, 287, 682, 2203, 1446, 1932, 1845, 2492, 607, 1745, 2247, 102, 141, 1186, 230, 2421, 3203, 1566, 2695, 2944, 2126, 1876, 713, 1771, 107, 3147, 866, 1492, 1781, 2401, 1379, 451, 1194, 282, 1877, 3154, 967, 3171, 1213, 1670, 2229, 2244, 59, 1507, 1838, 3029, 2122, 1405, 949, 488, 2312, 1415, 2378, 3012, 1830, 1843, 1236, 99, 2570, 16, 2110, 2092, 1755, 2751, 2456, 650, 1862, 1007, 587, 770, 918, 2265, 923, 1058, 1120, 1673, 2047, 1819, 2637, 882, 1489, 2384, 1520]. Ocean [2843, 620, 974, 2512, 2743, 2443, 73, 1893, 784, 2808, 3145, 2965, 2842, 568, 348, 503, 2353, 1902, 772, 312, 1139, 1896, 69, 159, 2959, 1398, 2464, 2004, 1494, 2552, 1238, 9, 2485, 8, 84, 2000, 3032, 900, 2549, 1442, 2059, 1910, 1578, 2910, 1143, 2210, 2075, 2558, 1850, 592, 553, 2440, 2615, 410, 386, 3013, 1637, 1003, 3169, 253, 5, 2424, 1479, 3059, 2219, 165, 1861, 1306, 1687, 701, 2194, 1674, 625, 1431, 2565, 382, 171, 1263, 1784, 1839, 707, 3211, 2788, 1218, 2023, 1874, 444, 2029, 799, 2816, 678, 1665, 806, 1825, 1247, 2830, 43, 1864, 3004]. Ocean [1577, 1436, 2189, 1042, 424, 471, 2794, 773, 2483, 1410, 2377, 2435, 1186, 337, 794, 1598, 1690, 708, 2520, 2393, 2728, 2060, 2621, 2845, 2601, 1322, 2805, 3100, 2786, 1979, 1208, 1308, 498, 1692, 1636, 2724, 1405, 42, 919, 949, 1254, 2378, 3063, 812, 1568, 1755, 2166, 1862, 683, 2165, 345, 1483, 2700, 1195, 1234, 1606, 1126, 2288, 1528, 1973, 523, 1182, 1522, 304, 1101, 2890, 3006, 966, 2631, 1386, 2829, 3023, 208, 956, 944, 897, 2105, 1614, 452, 1317, 78, 543, 2947, 309, 2910, 2582, 593, 2117, 721, 2706, 2766, 491, 1085, 252, 1080, 1356, 1443, 2744]. Ocean [2368, 1089, 549, 2734, 103, 1605, 1162, 1269, 86, 2156, 2160, 1201, 1421, 1764, 617, 2191, 2184, 1774, 1931, 2279, 326, 289, 1665, 1886, 56, 1338, 476, 3064, 2523, 2996]. Ocean-Oil-Sediment-Biology [1670]. OceanGliders [4, 2466]. OceanGNS [2189]. Oceanian [2997]. Oceanic [1268, 1123, 570, 155, 3203, 1891, 3100, 611, 507, 2383, 1097, 2478, 1818, 2562, 1957, 1220, 681, 910, 271, 1092, 359, 2381, 2198, 443, 1931, 1935, 2216, 1738]. oceanica [207, 221, 1667, 954, 88]. Oceanids [499]. Oceanographic [3039, 146, 1527, 155, 2249, 3104, 2753, 863, 877, 516, 534, 1625, 69, 681, 3048]. Oceanography [2096, 1448, 2818, 1422]. Oceans [102, 638, 1385, 16, 2013, 2082, 2627, 322, 1271, 2295, 1661, 723, 1169]. Octocoral [1721, 2628, 1227]. Octocorallia [803, 1227]. Octocorals [585, 1272, 2581, 413, 692, 2740]. Octopodoidea [284]. Octopus [2973, 1195, 940, 1880, 2973]. Octopuses [284]. Ocypode [1414]. ODATIS [902]. Odontocetes [641]. OECMs [1509]. Off [2535, 1865, 1678, 389, 2955, 1512, 2513, 212, 2018, 1792, 2598, 1187, 921, 2963, 2654, 2742, 1380, 525, 2989, 927, 845, 2749, 635, 1342, 1941, 2246, 2813, 1953, 1906, 876, 125, 2293, 2139]. Offer [1257]. Offers [2143]. officinalis [51, 215]. Offs [1978, 2599, 1762, 2740, 2907]. Offshore [2151, 842, 2091, 1710, 632, 1741, 1681, 1704, 720, 111, 658, 645, 481, 2636, 2236, 2885, 1317, 67, 2345, 126, 529, 332, 1144, 3033, 2529, 847, 233, 164]. Offspring [1149]. Often [2413]. OFURO3 [1501]. Ogives [2845]. ohtsukai [2564]. Oil [1675, 1923, 1681, 1670, 1232, 202, 1564, 1575, 1158, 2072, 161, 3146, 2636, 169, 2767, 3139, 2345, 1408, 1827, 3027, 1785, 1515, 1833, 937, 1698, 1390, 3160, 396, 504, 1664]. Oil-Based [1515]. Oil-Derived [1785]. Oil-Mineral [1664]. Oil-Particle-Aggregates [1675]. Oil-Residues [1681]. Oils [965]. Oithona [133]. Okadaic [1159, 3184]. Okhotsk [2656, 715]. Okinawa [2737, 2554]. Old [1265]. Oleaginous [2669]. Oligoalginate [522]. Oligohymenophorea [1261]. Oligomeric [905]. Oligosaccharide [2595]. Oligotrich [990]. Oligotrophic [699, 779, 389, 1938, 2066, 658, 1634, 1821, 2530, 319, 1014, 2422, 2846]. Oligotrophy [1224]. olivacea [851]. Olive [36]. Oman [768]. Omega [2087]. Omega-3 [2087]. Omics [2142, 820]. Ommastrephes [2334]. Omnivorous [2541]. OMP [1673]. OMZ [2200]. Onboard [584, 965]. Once [1397]. Oncorhynchus [3027, 2777]. One [1803, 1615, 495, 1869, 384, 2906, 2922]. One-Size-Fits-All [384]. Ongoing [484, 152, 28]. Online [1700, 1786]. Only [2958]. Onset [1656, 2239, 2525]. Onshore [1276, 411, 847, 2935]. Ontario [3135]. Onto [3102]. Ontogenetic [2562, 969, 2937, 3018, 1980, 2083]. Ontogeny [811, 1635, 2264, 2767, 1242]. Ontogeny-Specific [2767]. Onychoprion [2906]. Oocyte [884, 2939, 263]. Open [2507, 2974, 866, 451, 2751, 1583, 1182, 1238, 2000, 1637, 1587, 701]. Open-Ocean [866]. Opening [2464, 2025, 1868]. Operated [771, 385, 322, 1477, 1696]. Operating [653]. Operation [499, 312]. Operational [232, 1139, 69, 1408, 1861, 2376, 832]. Operationalized [1393, 981]. Operations [159, 1170]. opercularis [687]. Ophiuroids [2068, 1274]. Opisthobranch [3075]. Opportunities [1921, 404, 1867, 2975, 1930, 508, 1948, 72, 60, 297, 630, 2206, 2064, 2519, 2412, 845, 3173, 820, 692, 2572, 832, 678]. Opportunity [814, 1326, 2549, 995, 2237, 1003, 2370]. Optical [2233, 1937, 704, 2942, 99, 578, 2069, 1140, 1417, 1105]. Optically [44]. Optimal [2400, 1843, 1301, 3037, 641]. Optimisation [1908]. Optimization [2764, 2903, 1928]. Optimized [2979, 2610, 2215, 1958, 2207]. Optimizes [461]. Optimizing [531, 2204, 2107]. Options [102, 2534, 349, 634, 200, 1752]. OPUS [2069]. Orange [851]. Orbicella [2687, 1654]. Orca [2305, 95]. Orcaella [1320, 1556]. Orchard [2352]. Orchestrating [2265]. Orchestration [1648]. Orcinus [95]. Order [542, 202, 662]. Order-Wide [542]. Ordinal [2845]. Oreochromis [2586, 465]. Organellar [22]. Organelle [3165]. Organic [225, 229, 163, 329, 1834, 619, 104, 1456, 279, 1081, 1710, 1811, 867, 1250, 866, 2942, 1308, 1789, 1183, 3089, 3063, 99, 2342, 1547, 2066, 2527, 3192, 3000, 2045, 2960, 2643, 1704, 441, 1826, 620, 512, 2713, 3140, 2890, 568, 1502, 3151, 874, 2399, 910, 2995, 1871, 1503, 452, 2104, 1151, 1140, 626, 1454, 506, 2227, 530, 122, 1013, 635, 820, 1162, 382, 171, 443, 1348, 1467, 2476, 350, 777, 2449, 1082, 546]. Organic-Rich [3089]. Organically [751]. Organisation [542]. Organism [2074, 1312]. Organismal [1272]. Organisms [1294, 1965, 2443, 2120, 2128, 933, 2162, 2742, 709, 2538, 3085]. Organization [1822, 733]. Organizational [857]. Organohalides [2104]. Organoleptic [713]. Organs [1880]. orientalis [1261]. Orientation [287, 1065, 705]. Orientation-Specific [705]. Orientations [2665]. Oriented [3057, 596]. Origin [563, 1584, 2785, 760, 1694, 3002, 2988]. Origins [259]. Orkney [37]. Ormonde [9, 2105]. Ornamental [348]. oroides [1251]. Orphan [1508, 10]. Orthoreovirus [2263]. Oryzias [1565, 3120]. Oscillating [1648]. Oscillation [3054, 1933, 40]. Oscillations [1725, 2289]. Oscillator [1659]. Osculum [2648, 1403]. Osedax [545]. Osmolality [2840]. Osmolyte [2399]. Osmolytes [2399]. Osmoregulatory [811]. Osmotic [32]. Osteoblasts [2134]. Ostrea [187]. Ostreinae [2878]. Ostreopsis [785, 600]. otakii [520]. Otaria [881]. Other [484, 152, 1243, 121, 1237, 735, 2183, 2665, 2162, 2538]. OtHV1 [1244]. Otolith [519, 449, 422, 667, 631, 360, 417, 2503]. Otters [515, 864, 1689, 437, 684, 1409]. Our [3143, 1712, 322, 2500]. Outbreak [1789, 1718, 1875, 2148, 2749, 1702]. Outbreaks [2317, 2426]. Outcome [1213, 2690]. Outcome-Limiting [2690]. Outcomes [3180, 270, 685, 363, 758]. Outdoor [2665, 985]. Outer [1971, 254]. Outer-Shelf [254]. Outflow [1517, 2646, 9, 2105]. Outplanting [2133]. Outplants [2220]. Outreach [774]. ovalis [2875]. Ovaprim [1008, 1354]. Ovary [2003]. ovata [785]. ovatus [1345, 2134, 1297, 3130, 3055]. Over-Summering [1858]. Over-the-Horizon [499]. Overcoming [794, 870]. Overexploitation [2238, 2809, 917]. Overfishing [419, 625]. Overlap [224, 2375]. Overlapping [2177]. Overlooked [1991, 2413, 890, 328]. Overrides [119]. Overseas [2036]. Overstated [2788, 3211]. Overturning [1830, 3095, 2212]. Overview [1185, 1186, 2770, 824, 353, 1830, 835, 2189]. Overwintering [1527, 574]. OWC [2268]. Own [1173]. Oxidation [163, 635, 3148, 2680]. Oxidative [2971, 2124]. Oxide [1496, 2131, 458]. Oxidization [3192]. Oxidized [3146, 3027]. Oxidizers [2366]. Oxidizing [397]. Oxygen [104, 867, 1781, 462, 3084, 1830, 2751, 2715, 1826, 528, 2326, 944, 2742, 2793, 2833, 1223, 354, 635, 211, 1342, 3150, 2503]. Oxygen-Deficient [944]. oxygenase [2920]. Oxygenation [2814]. Oxytricha [1590]. Oxytrichidae [1590]. Oxytrichids [1341]. Oyashio [1046]. Oyster [509, 157, 1256, 890, 3154, 501, 967, 1213, 1569, 700, 2270, 2056, 1727, 67, 758, 2878, 327, 2759, 2293, 2946, 3103, 1970]. Oysters [1204, 1152, 514].

P [1770]. P. [537, 393]. p53 [3120, 388]. p53-Deficient [3120]. P53/Cycling [388]. Paced [1214]. Pacific [471, 1410, 2621, 2845, 2805, 1771, 3171, 1901, 1636, 2244, 2724, 1254, 97, 16, 1568, 1995, 1234, 3168, 1973, 1522, 1386, 1634, 413, 956, 944, 205, 78, 721, 2706, 2766, 491, 1604, 1089, 103, 289, 56, 1338, 476, 2523, 2996, 1042, 963, 998, 1630, 2179, 509, 805, 1975, 1464, 2578, 1690, 663, 867, 2589, 3100, 1813, 14, 589, 3154, 501, 895, 1692, 121, 1213, 881, 2095, 96, 1100, 3063, 2973, 2465, 1523, 36, 2900, 2999, 760, 2166, 1550, 1130, 2333, 1448, 1943, 657, 1819, 2700, 2717, 1449, 1025, 3061, 6, 92, 2655, 714, 1582, 1169, 2338, 2288, 1528]. Pacific [999, 803, 1392, 1969, 2205, 2020, 896, 2685, 1957, 1064, 2829, 904, 908, 67, 2314, 1682, 1380, 876, 2234, 2245, 532, 2878, 1607, 1177, 3050, 606, 2084, 1091, 2664, 2992, 2997, 1784, 2281, 1572, 1218, 311, 718, 2759, 2212, 1931, 2938, 2422, 1935, 1062, 2838, 2372, 468, 3004, 1010]. pacificum [2961]. pacificus [2032]. Pack [554]. Package [1758]. Padre [316]. Pagellus [2556]. Pagrus [1720]. Pair [1187]. Pairing [780]. Palaemonetes [2792]. Palawan [2031]. Paleo [2926]. Paleo-Typhoon [2926]. Paleoenvironmental [2266]. Paleontological [1313]. Paleothermometry [2677]. palifera [1247]. Palk [195]. pallasi [1218]. Palm [1515]. palmata [1608]. Pam [861]. PAML [136]. Pampus [3077]. Pan [887, 708, 1430, 2895, 1331, 2531, 1432, 2279, 2872, 1261, 252]. Pan-Arctic [887, 708, 2895, 1331, 2531, 2279, 252]. Pan-European [1430, 1432, 2872]. Panama [1235, 1647, 36, 2900]. Pandemic [1807, 2402, 2242, 2426, 3032, 630, 2834, 2364, 1983, 2734, 2731, 2701, 2152]. Pandora [2949]. Pangasianodon [3152]. panicea [216]. Pantoea [2214]. Panulirus [1666, 3071]. Papa [1162]. Papua [3050]. Paracalanid [3174]. Paracentrotus [2708, 1897]. Paracoccus [1215]. Paracrine [2939]. Paracyclopina [2484]. Paradigm [1194, 2132, 12, 280, 1005]. paradoxus [417]. Paragorgia [315]. parahaemolyticus [1881, 2589, 3061, 392, 1609]. Paralichthys [1716]. Parallel [2686]. paralytic [2835]. paramamosain [2003, 2772, 2866, 2939, 3166, 2826]. Parameterization [2656, 614]. Parameterizations [81]. Parameters [990, 2088, 2856, 1906, 731, 834, 1908, 2855, 95, 1985]. Paramuricea [3179, 3178]. Parasite [1480, 2792]. Parasites [881, 2018]. Parasitic [1713, 2716, 1533, 1537]. Parasitizing [1768]. Park [1376, 1516, 3118, 551, 995, 170, 1387, 1154]. Parks [2154]. Parrotfishes [1735]. Parry [3135]. Part [1282, 2428, 3115, 2313, 622, 3137, 623]. Partial [2760, 2581, 2855]. Partially [1469, 1090, 1603]. Participants [2665]. Participation [2397, 746, 1406, 401]. Participatory [1649, 3132, 928, 746, 2315, 200]. Particle [2483, 2882, 44, 1675, 251, 666, 1704, 1296, 2713, 304, 2842, 2020, 2077, 933, 17, 3093, 400, 2523, 407]. Particles [963, 2208, 192, 575, 507, 99, 3102, 1119, 2366, 578, 2162, 1015, 686, 2953]. Particular [3140]. Particulate [2596, 1456, 1081, 2328, 815, 1979, 866, 2342, 2527, 2415, 620, 2403, 1973, 452, 973, 1035, 2476, 777, 2449]. Parties [1719]. Partitioning [1938, 133, 379, 726, 359, 2227, 1283]. Partly [806]. Partner [209]. Partnership [2485]. Partnerships [2558]. Parts [2917]. Parturition [1012]. parvacauda [393]. Parvocalanus [2077, 2076]. Pasir [1516]. Passage [1592]. Passive [1470, 2605, 203, 575, 369, 231, 2486, 641, 3033, 1623, 2553, 1034]. Past [66, 464, 1771, 1992, 3210, 213, 2652, 642, 3050, 3040, 2622]. Patagonia [1763, 1485, 1239, 1505, 391, 1512, 1662, 1503, 1756]. Patagonian [1499, 1065, 2018, 1502, 1497]. Patch [3069, 2319, 1457]. Patches [1516, 412, 1835]. Patchiness [494, 665]. Patchy [2194]. patens [2516]. Paternal [2825]. Path [1472, 2155, 9, 2105, 477, 753, 1005]. Pathogen [437]. Pathogenic [2214]. Pathogenicity [2899]. Pathogens [484, 152, 2035]. Pathological [786]. Pathology [1437, 1768, 1409, 2953]. Pathway [1256, 1213, 2361, 1088, 330, 388, 2939]. Pathways [163, 3180, 498, 951, 414, 2732, 3178, 731, 982, 2087, 1827, 1605, 2894, 1199, 2594, 2212]. Pattern [2249, 2693, 2772, 313, 1302, 2200, 493, 2657, 2177, 2002, 3160, 1484]. Patterns [2507, 2179, 2096, 267, 2226, 1433, 1795, 2325, 1110, 155, 2774, 351, 2048, 2026, 2691, 2046, 1680, 2317, 228, 2248, 3161, 919, 1853, 1601, 2837, 812, 2342, 1831, 2148, 143, 2385, 1130, 1221, 2949, 151, 2232, 2747, 101, 2183, 2562, 896, 2635, 3054, 2610, 2162, 1429, 2158, 1682, 2796, 2914, 374, 3162, 3164, 2892, 973, 2609, 995, 1089, 1477, 1734, 1610, 126, 1154, 1687, 911, 686, 2741, 1012, 2083, 1466, 564, 2894, 1660, 2381, 327, 1756, 1428, 2416, 1641, 1747, 1207]. patula [581]. Pave [3197]. Pb [2559, 2483, 2435, 2150, 2527, 2550, 2685, 2340, 2530, 2998, 2538, 2449, 2523]. PCB [2313]. pCO [53, 1771, 1596, 1828, 2288, 425]. PCR [2792]. PDS [2290]. Peak [167]. Peanut [2537]. Pearl [1992, 1569, 1192, 358, 3050, 2293, 3103, 2388, 1970, 2813, 2403, 2433, 2366, 2657, 1672, 338]. Peat [860]. pectinata [2976]. Pelagia [2663]. Pelagic [1378, 2783, 1463, 1027, 672, 887, 632, 2060, 185, 1750, 1595, 873, 903, 420, 498, 1559, 3024, 744, 1963, 1375, 924, 1169, 1157, 2223, 1634, 1916, 3090, 944, 2149, 2722, 724, 1143, 2981, 359, 506, 973, 727, 1952, 1743, 372, 190, 940, 459, 1287, 2138]. Pelagicoccus [1972]. Pelagification [1143]. Pelagophytes [1770]. Pelagos [1211]. pelamis [998, 2845]. Pellets [73]. Pelliciera [2174]. Pelodiscus [1728]. Pen [829, 2976]. Penaeus [2361, 3061, 1264, 1073, 274, 3111]. Penetrating [1402]. pengoi [1021]. Penguin [2914, 1621, 1569]. Penguins [1359, 1246]. Peninsula [1643, 709, 37, 1526, 3108, 1198, 1047, 1359, 1529, 1391, 1679, 1452, 849]. Peninsular [1433, 1335, 2022]. Pennatulacea [1944]. Pens [1944]. Pensacola [1881]. People [102, 649, 1131]. Peptide [1115, 1165]. Peptides [2128, 2707]. Peptidomics [993]. Per-Recruit [1374, 1941]. Peracarid [919]. Peracarida [1818]. Peracute [605]. Perceived [1434]. Perception [261, 186, 2708, 1131, 262]. Perceptions [765, 702, 240, 1252, 624, 2637, 2031, 248]. Perceptual [27]. Perciformes [2804, 523]. Perdido [1881]. Performance [821, 2586, 1924, 2621, 215, 2905, 2589, 2268, 2623, 3205, 1345, 1569, 2110, 1523, 1285, 1963, 2537, 3061, 1297, 2270, 1733, 1882, 3065, 1147, 3130, 662, 2669, 1199, 806, 2870, 1696]. Perinereis [2486, 2033]. Period [1789, 1998, 2756, 904, 1810, 2603, 1260, 878]. Periodic [3007, 2289]. Periodicity [2101]. Periods [1571, 2828, 1439]. Periophthalmodon [1335]. Peripheral [3120]. Peritrichs [1243]. Periwinkle [1820]. Perkinsus [700]. Perlas [1647]. Permeability [847]. Permits [2936]. Perna [2710]. Peroxidase [1751]. Peroxide [175]. Perpetuated [1019]. Perseverance [2701]. Persian [1036, 675]. Persian/Arabian [675]. Persistence [1189, 2815, 1584, 2450, 2458, 2773]. Persistent [225, 45, 1927, 1244, 2960, 1974, 1487, 705]. Personal [2330]. Personality [136]. Perspective [1964, 287, 601, 708, 232, 2397, 1749, 1033, 1575, 1655, 2585, 630, 953, 1707, 1857, 2584, 701, 443, 2444, 2279, 2816, 3038, 733]. Perspectives [308, 2784, 2209, 367, 1719, 2312, 2899, 1096, 1862, 1562, 60, 923, 835, 2106, 2548, 689, 727, 1558, 2278, 3020, 1173, 1475, 2757]. Perturbed [651]. pertusa [315]. Peru [852, 3132, 2742]. Peruvian [1933, 2480, 783]. Pervasive [1836]. Pest [280]. Pesticides [1239]. Petroleum [872]. pH [3122, 1830, 1149, 1485, 2326, 2474, 2938]. Phaeocystis [1978, 62]. Phaeodactylum [862]. Phaeophyceae [2110, 558, 1135]. Phaeophyta [859, 930]. Pharaoh [1800]. pharaonis [61, 1800]. Pharmacology [3125]. Pharynx [1362]. Pharynx/Esophagus [1362]. Phase [1057, 2915, 207, 1055, 726, 1906, 860]. Phase-Dependent [207]. Phases [2902]. Phenolic [1250, 1894]. Phenology [1070, 1606, 2922, 378, 242]. Phenomena [1653]. Phenomenon [489, 2385, 1686]. Phenotype [3121, 2867]. Phenotypes [920, 2867]. Phenotypic [501, 1893, 2239]. Phenylalanine [1285]. Pheronema [1506]. philargius [2958]. philippinarum [3136, 2359, 2417]. Philippine [2488, 2709, 2798]. Philippines [1815, 1434, 2031, 2803, 2536, 334, 1466]. Phoca [421]. phocaenoides [1025]. Phocarctos [2300]. Phocoena [1561, 446, 387, 2320, 89]. Phoenix [146, 1108]. Phosphate [2492, 1971, 208]. Phospholipids [1068]. Phosphorous [807]. Phosphorus [383, 1871, 2672, 1019, 2969]. Phosphorylation [569]. Photo [1584]. Photo-Identification [1584]. Photoacclimation [1429, 483]. Photobleaching [2868]. Photoblepharon [181]. Photochemical [1140, 2370]. Photogenic [1646]. Photogrammetric [1504]. Photogrammetry [2736, 2948, 2041, 1965, 1216, 1888, 1796, 1797, 1772]. Photogrammetry-Based [2948, 1796]. Photographic [827]. Photography [1986, 2186]. Photoinactivation [2846]. Photomosaics [2319]. Photonic [567]. Photooxidation [1390]. Photoperiod [314, 884, 3174]. Photophore [1646]. Photophysiological [2110, 1993]. Photophysiology [1358]. Photoprotection [305]. Photosensitivity [2940]. Photosynthesis [1937, 633, 1485, 2412, 239, 1709]. Photosynthetic [2889, 272, 3201, 2587, 664, 461, 2627, 567, 1011, 2846]. Phototaxis [2501]. Phototrophic [3037]. Phthalate [2008]. Phycotoxins [1790]. Phyllopharyngea [1629]. Phylogenetic [1313, 284, 2599, 1818, 803, 3035, 1227, 2875, 1440, 1207, 2752]. Phylogenetics [542, 2925]. Phylogenomics [2752]. Phylogeny [1768, 295, 571, 1261, 3072, 1327, 5, 2424, 1242, 1837, 1304]. Phylogeographic [1658, 1682]. Phylogeography [2684, 2664, 2976, 468]. Phylotypes [1733]. physalus [1607]. Physeter [1108]. Physical [2061, 887, 2328, 2520, 638, 976, 3081, 877, 2941, 340, 622, 897, 1648, 599, 623, 288, 1421]. Physical-Biological [623]. Physico [1471, 1498]. Physico-Chemical [1498]. Physico-Chemistry [1471]. Physicochemical [2255]. Physics [651]. Physiological [427, 2365, 862, 2418, 1535, 1297, 2223, 1056, 305, 2671, 117, 806, 2644, 2470, 3193, 3158, 2861]. Physiology [408, 96, 1149, 1043, 2956, 910, 2921]. phytopathogenic [415]. Phytoplankton [1042, 2251, 1282, 2269, 287, 383, 376, 1081, 1030, 1425, 337, 2427, 2255, 2865, 1397, 2645, 1596, 1047, 2121, 128, 272, 1778, 2837, 7, 1346, 1828, 2629, 1606, 2755, 3177, 317, 2957, 1111, 2598, 736, 2871, 137, 2627, 1485, 356, 2167, 677, 750, 2307, 425, 1171, 183, 756, 2796, 1850, 1307, 1358, 2412, 595, 2734, 2922, 1283, 701, 86, 2291, 617, 311, 1732, 764, 715, 242, 2846, 1134, 2838, 361, 2019, 378]. Phytoplankton-Derived [1081]. Phytoplanktonic [2329]. Pi [1971]. Pico [2871, 1113, 2007, 2422]. Pico-Eukaryotes [2007]. Pico-Eukaryotic [1113]. Pico-Phytoplankton [2871]. Pico-Sized [2422]. Picochlorum [2077]. Picophytoplankton [1737, 1938]. Picoplankton [2968, 2807, 1423]. Picture [584]. Piers [757]. Pigment [2912, 2970, 2940]. Pigmentation [825]. Pigmentation-Related [825]. Pigmented [247]. Pigments [317, 701]. Pikeperch [2304]. pilchardus [2678, 978]. Pile [2091]. Pile-Driving [2091]. Pillar [745, 1982, 1996]. Pilot [3135]. Pinctada [2791, 2573, 2293, 3103, 1970]. Pingers [387]. Pink [2661]. Pinna [484, 152, 1188]. Pinnacle [233]. Pinniped [1001]. Pinpointing [2017]. Pioneering [217]. Pipeline [1946]. Pipelines [1113]. pirAB [392]. piRNA [2791]. piRNA-Like [2791]. Pisces [1118]. Piscivorous [1101, 1400]. pistillata [968, 117]. Pit [302, 961]. Pitcairn [1807, 1901]. PIWI [2791].

PIWI-Like [2791]. Place [674, 886, 2560, 1548, 525, 2086]. Placing [2426, 1129]. Plague [2614]. Plain [2679, 271, 2200, 2350]. Plan [3032, 561]. Planet [40]. Planktic [2044, 1860]. Plankton [779, 3060, 1977, 1308, 755, 1061, 958, 1097, 769, 1078, 356, 1113, 900, 2162, 1015, 3137, 1306, 686, 2422, 1325]. Planktonic [324, 659, 1853, 2901, 24, 829, 1533, 1821]. Planning [2764, 601, 270, 1492, 1757, 2327, 2795, 2659, 430, 875, 2709, 1724, 2649, 2147, 1144]. Planoglabratella [687]. Plans [2739, 1166, 1722, 2632]. Plant [243, 3110, 2670]. Plantigrades [952]. Planting [1301]. Plants [1570, 2819, 953, 2538, 2588]. Plasma [1684, 2090]. Plastic [3080, 963, 364, 1395, 1773, 1956, 2510, 3197, 1379, 628, 2515, 795, 1636, 2321, 1516, 972, 2654, 26, 1587, 3086, 2543, 2452, 2481, 2644, 2858, 3200, 1292]. Plasticity [2825, 3045, 2544, 1302, 1893, 1286, 558, 481, 2936, 1993, 386, 2780, 307, 1136, 2258]. Plasticizers [1155]. Plastics [3197, 2382, 2052, 2396, 1035, 1011, 2894, 1468, 2481]. Plastid [1725]. Plastidial [207]. Plata [2510]. Platanista [1370]. Plate [1402]. Plateau [390, 848, 743]. Plates [3204]. Platform [2994, 8, 84, 1839, 832]. Platforms [3039, 842, 369, 2072, 1182, 995, 2447]. Platichthys [1567]. Platygyra [2810]. Play [1094, 112]. Players [1945, 1089]. Playing [646]. Plectropomus [2762]. Pleurostomatid [1837]. plicatilis [1515]. plicatula [2878]. plumbea [1829]. Plume [1276, 2532, 872, 670, 2617, 2303, 3112, 3208, 1051, 2929, 1134]. Plumes [273, 1088, 645, 1545, 2289]. Plunging [529]. Pluralism [2190]. Po/ [2527, 2685, 2340]. POC [2523]. Pocillopora [2074, 633, 2568, 955, 2644]. Pocket [547, 1751]. Pockmark [767]. Point [3053]. Points [2802, 1270, 1474, 983]. Poisoning [348]. Pol [2772]. Polar [1986, 638, 591, 2272, 1926, 1168, 1951, 1720, 55, 1860, 3164]. Polarized [2862]. Poleward [2660]. Policies [105, 1857, 1748]. Policy [79, 2397, 2566, 511, 2265, 634, 1913, 2353, 965, 2171, 1910, 1133, 1041, 3206]. Policy-Relevant [2171]. Political [981, 982, 2345, 2147, 1393]. Politics [1745]. Pollutant [225]. Pollutants [225, 1862, 2960, 1966, 953]. Pollution [1045, 85, 1991, 793, 1956, 2510, 1379, 2240, 951, 2515, 2839, 965, 2396, 3034, 972, 2152, 2654, 2193, 280, 1035, 1748, 2657, 1925, 3026, 761, 893]. polyactis [1941]. Polyamine [2623]. Polyamine-Producing [2623]. Polycentric [2549]. Polycentricity [2570]. Polychaete [1981, 884]. Polychaetes [812, 2457, 2204]. Polychlorinated [2979]. Polycomb [291]. Polycyclic [366, 158, 1925]. Polyethylene [3102]. Polymeric [396]. Polymetallic [560, 349, 2068, 1747]. Polymetallic-Nodule [2068]. Polymorphism [1316, 393]. Polymorphisms [998, 47]. Polynesia [120, 74, 262, 1197]. Polynoid [1842, 2343]. Polynoidae [1987, 1842, 2343]. Polynya [62, 3181, 1116]. polyphagus [3071]. Polypropylene [3102]. Polyps [1481, 2893, 2341]. Polystyrene [2208, 2414, 1063]. Polyunsaturated [1823, 1297, 2658]. POM [2329]. Pompano [1345, 2134, 3130, 3055]. Pond [2435, 2933, 2571]. Ponds [1631, 662]. Pontevedra [2046]. Pontoporia [801, 1284]. Pool [1859]. Pools [2393]. Poor [1367, 1344, 1407, 934, 2661, 1593, 1752, 2723, 2696]. Poorly [1463, 1324, 3067]. Pop [2290]. Pop-Up [2290]. Populated [1142, 2330, 2442]. Population [3014, 998, 421, 1104, 1541, 1185, 2070, 2974, 1110, 218, 185, 2859, 1332, 2108, 14, 391, 1622, 2619, 3161, 1181, 1044, 3101, 1875, 1359, 661, 1394, 2900, 2999, 760, 868, 717, 850, 523, 996, 1658, 2958, 2270, 47, 636, 2205, 479, 796, 1219, 2495, 956, 1982, 2429, 1822, 841, 2331, 3002, 1067, 259, 2924, 1524, 1609, 123, 1621, 2867, 2529, 1934, 3174, 2875, 768, 1685, 1136, 56, 618, 476, 3103, 2757, 1284]. Population-Level [1110, 1622, 661]. Populations [2379, 2684, 813, 610, 2836, 1806, 1188, 572, 2725, 77, 1682, 1191, 3128, 187, 1041, 178, 844, 761, 2823]. Porbeagle [375]. Porbeagles [1610]. Porcupine [1009]. Porewater [530]. Porifera [2848, 1196, 680, 576, 1440]. Porites [57, 951, 2544, 1535, 1247, 2644, 2470]. porosus [228]. Porphyra [1660]. Porpoise [1561, 2129, 834, 1370, 1273, 89, 245]. Porpoises [2091, 1549, 1025, 446, 387, 2320]. Port [2001, 2526, 342, 2458]. Portable [1006, 1325]. Portraying [2257]. Portrays [717]. Ports [1722]. Portugal [2255, 1249, 277, 2222, 2345, 239]. Portuguese [2358, 2336, 1668]. Portunus [2933, 3189, 1286, 2016, 2750, 136, 3166, 3193]. Posidonia [221, 1667, 954, 88]. Position [1991, 591, 385, 943]. Positions [1759]. Positive [314, 1363, 3101, 2323, 1942, 758, 194, 1428]. Positively [2595, 1244]. Positives [646]. Possessing [522]. Possession [994]. Possibilities [2534]. Possible [2479, 1401, 1746, 2866, 2097, 409, 719, 3093]. Post [3011, 2711, 864, 1571, 795, 661, 1943, 2642, 2152, 698, 3198, 1223, 3204, 2438]. Post-eastern [1223]. Post-eruptive [2438]. Post-hatchling [795]. Post-nesting [3011, 661]. Post-pandemic [2152]. Post-recruitment [3198]. Post-release [864, 1571]. Post-seawater [2642]. Post-settlement [3204]. Post-stratification [2711]. Post-telemetry [698]. Postlarvae [2699]. Postmortem [883]. Postsynaptic [112]. Pot [3126]. Potassium [3212]. Potent [1863]. Potential [2820, 963, 2214, 3066, 2698, 2330, 286, 1462, 1975, 2471, 741, 663, 473, 2864, 1823, 299, 1769, 1638, 2515, 1518, 922, 1405, 2321, 949, 730, 505, 650, 3000, 1234, 882, 2426, 1294, 1188, 2659, 2403, 51, 1086, 1029, 292, 1447, 487, 912, 1139, 1727, 159, 3090, 1949, 2722, 1546, 442, 709, 219, 355, 1627, 2616, 880, 132, 1734, 245, 3086, 2824, 178, 3211, 2788, 2050, 459, 715, 2758, 2846, 2858, 2752, 1468]. Potentially [1945, 2298]. Potentials [3159]. pourtalesii [2475, 1217, 1493]. Poverty [783]. Power [2439, 3067, 2538]. Practical [2459, 1219, 1884]. Practice [2566, 1638, 2557, 1182, 2790, 3142]. Practices [643, 757, 624, 923, 312, 2964, 1370, 913, 1578, 2136, 3013, 3005, 405]. Practitioner [3008]. Practitioners [1687]. Pragmatic [2744]. PRDX [3157]. Pre [215, 2380, 1339, 1777]. Pre-breeding [2380]. Pre-eutrophic [1777]. Pre-industrial [215]. Pre-treatment [1339]. Precautionary [2619]. Precious [584, 2682, 2125]. Precipitation [2340]. Precise [1093]. Precision [333, 1552]. Preconditioning [2213]. Predating [3190]. Predation [1527, 1052, 951, 2285, 2349, 2175]. Predator [2096, 1936, 2593, 1766, 2863, 2080, 2102, 1913]. Predators [1961, 1528, 1765, 1447, 684, 518, 2404]. Predatory [1763, 1524]. Predict [1712, 1695, 3178, 1794]. Predictability [1209, 488, 2004, 691, 2506]. Predictable [525]. Predicted [351, 2546, 2225, 2914]. Predicting [2930, 1410, 1281, 2235, 234, 1781, 1849, 379, 2751, 2129, 658, 100, 2307, 1191, 1361, 1390]. Prediction [1560, 895, 1919, 2844, 2101, 2977, 2215, 463, 606, 2376, 3173, 2419]. Predictions [488, 2884, 3000, 2735, 3138, 1637]. Predictive [1491]. Predictors [1000, 2984]. Predicts [3099, 1443, 1041]. Preference [196]. Preferences [3075, 1550, 1245]. Preferential [192, 1939]. Pregnancies [1688]. Pregnancy [1703, 2854]. Preliminary [2621, 1579, 914, 2193, 1357]. Preparation [1008, 1354, 504]. Preparing [1057, 1055]. Preselection [1567]. Presence [203, 990, 45, 3049, 1184, 2298, 2947, 1698, 1197]. Present [66, 3210, 124, 2652, 3050]. Preservation [2337, 1339]. Preservative [1783]. Press [2345]. Pressure [3003, 2290, 2855]. Pressure-Recording [2290]. Pressured [2452]. Pressures [1921, 2993, 2718, 2517, 2031, 689, 897]. Preston [1951]. Prevailing [2061]. Prevalence [2322, 2139]. Prevalent [1243]. Prevent [1204]. Prevention [2410]. Preventive [1722]. Previously [1198, 2720, 2504]. Prey [2096, 1936, 728, 2333, 868, 2967, 2635, 1219, 2101, 2914, 553, 3083]. Primary [1282, 44, 815, 221, 3006, 622, 119, 252, 3164, 2271, 2744, 623, 2412, 2931, 2156, 142, 1260]. Primer [1525]. Priming [2013]. Primnoa [315]. Prince [1526, 3108, 1198, 1156]. Principal [283]. Príncipe [2730]. Principles [2534, 430, 444, 3142]. Print [492]. Prionace [1957]. Priorities [586, 584, 2502, 1047, 109, 1883]. Prioritization [953]. Priority [2222, 1883]. Private [2155]. Pro [1647]. Pro-Conservation [1647]. Proactive [1708]. Proanthocyanidins [905]. Probabilistic [1071, 2991, 2332]. Probability [1184]. Probing [291]. Probiotic [2214, 2024]. Problem [2985, 628, 1974, 1722]. Procedural [1760]. Procedures [2069]. Process [1323, 2003, 2494, 746, 670, 1858, 410, 2438]. Processed [2574]. Processes [2532, 887, 1710, 2328, 3203, 1642, 1936, 670, 1631, 3178, 2735, 756, 1898, 1827, 2448, 1180, 3149, 1874, 43]. Processing [1811, 674, 1826, 3151, 2069, 1358, 517]. Prochlorococcus [2131, 1011, 1338]. Prochlorophyte [2076]. Produce [2204, 2669]. Produces [2229, 1863]. Producing [1940, 2623, 2806, 2024]. Production [1233, 1347, 1282, 2310, 329, 3191, 175, 815, 1215, 2564, 133, 1308, 2590, 2456, 3192, 2700, 1533, 221, 2232, 512, 3006, 2642, 622, 991, 626, 252, 3164, 2745, 880, 623, 2931, 1040, 1511, 613, 617, 142, 3174, 777, 1260, 2620, 361]. Productive [2530, 777, 878]. Productivity [243, 2800, 1945, 128, 2019, 272, 1669, 958, 2837, 1435, 2781, 2412, 103, 74, 2156, 2381, 2814]. Products [2330, 1924, 258, 1345, 1200, 2093, 2745]. Profile [2840, 1402, 2669, 3087]. Profiles [2930, 2099, 1891, 57, 2912, 295, 1073, 2252, 2742, 2380, 2257, 1384, 802].

Profiling [1057, 2802, 1200, 2961, 2326, 941, 2141, 1312, 729, 802, 1805, 378, 242]. Profundidad [1376]. Progeny [1733]. Progestagen [1688]. Progesterone [1688]. Program [1915, 1831, 1120, 1673, 1916, 1585, 197, 1049, 444, 1305]. Programmed [1243]. Programmes [1757]. Programs [2051, 12, 3033]. Progress [2799, 2662, 1586, 600, 1752, 3020]. Progression [2687, 285, 1216]. Project [644, 1186, 270, 1406, 2055, 1057, 1055, 2874]. Projected [230, 2356, 2363, 950, 1041, 334]. Projection [1415]. Projections [1510, 1393, 981, 1968, 365, 227, 1674, 2332, 1764]. Projects [1305]. Prokaryoplankton [462]. Prokaryotic [1020, 2029]. Prolonged [1940, 1703]. Promising [162]. Promote [2511, 258, 261, 579, 804, 347]. Promoted [149, 1686]. Promotes [2623, 1039]. Promoting [2247, 2695, 1363, 2670, 2464, 3195, 2702, 2123, 1040]. Pronounced [1266]. Proof [2699, 1652, 510]. Proof-of-Concept [2699, 1652, 510]. Propaedeutic [3039]. Propagation [3122, 2660, 3168, 2954, 1604]. Propagules [3002, 826]. Propel [3019]. Properties [2233, 2535, 229, 3060, 1937, 638, 1876, 2942, 1202, 1950, 2201, 1140, 477, 869, 521, 984, 1810]. Proportions [1579]. Proposal [1261]. Proposals [2152]. Proposed [1669, 2047]. Prorocentrum [1790, 2684, 671]. Prospect [2453, 504]. Prospective [963, 2284]. Prospects [213]. Protandric [1132, 2917, 3183]. Protect [1561]. Protected [3066, 146, 2224, 601, 2182, 1741, 1927, 1492, 420, 1106, 1290, 2327, 2514, 538, 12, 2454, 283, 2259, 1447, 2631, 1274, 1521, 934, 2241, 100, 2632, 2536, 894, 2906, 3128, 3113, 1607, 2144, 15, 775, 2199, 2147, 1743, 1098, 970]. Protecting [1459]. Protection [2397, 2763, 2502, 1550, 846, 875, 2097, 1775, 876, 536, 179, 3001, 975, 1010]. Protections [1868, 845]. Protective [475, 1946]. Protein [2586, 2595, 2840, 299, 859, 569, 2280, 3192, 2537, 1971, 3061, 295, 3065, 2761, 2223, 2707, 2676, 2574, 3127, 517, 3141, 2939, 1920]. Protein-Like [3192]. Proteins [2853, 1128, 2574, 517, 291, 300]. Proteome [2423, 3148, 2680, 2258]. Proteomic [1713, 1565, 3044, 330, 2037, 2681, 3021, 707, 3184]. Proteomics [1159, 1076, 2831, 3038]. Protist [1323, 345, 1078, 1129]. Protista [1243, 1259, 1629, 1242]. Protistal [238]. Protists [659, 2857, 761]. Protocol [2494, 1145, 2526, 1581]. Protocols [2736, 1496, 336, 1065, 2395, 16, 943, 3013]. Protozoa [1261, 1341, 3072, 1327, 1590]. Protracted [68]. Provenance [1500]. Provide [203, 188, 584, 1044, 2900, 2607, 1279, 2717, 290, 358, 2600, 2888]. Provided [2763, 88]. Providencia [3074]. Provides [1233, 2802, 501, 2772, 456, 846, 2267, 3055, 2572, 3038, 2752]. Providing [1123]. Province [2688, 763, 3009]. Provinces [1268, 759, 2057]. Provisioned [1822]. Provisioning [1663, 2102, 1551]. Provoke [2323]. Proximity [1856]. Proxy [62, 2321, 3114, 1799, 377]. pruinosa [2130]. Prydz [2527]. Psammobatis [393]. Pseudo [580, 2082, 3097]. Pseudo-Nitzschia [580, 2082, 3097]. Pseudocalanus [878].

Pseudocarcharias [1126]. Pseudodiaptomus [1720]. Pseudodiploria [2148]. pseudonana [1725]. Pseudoruegeria [2782]. Pseudosciaena [574]. Pseudotanais [3143]. PSII [2846]. PST [2806]. PST-Producing [2806]. Psychological [2626]. Psychosocial [1739]. Pteria [1569]. Pterois [537]. Pteropod [2044, 53, 1398, 1442, 916]. Pteropods [2203, 1860]. PtP38 [2016]. Public [765, 240, 2155, 1252, 261, 2637, 2626, 3013, 262, 1577]. Published [2878]. Puerto [2812, 2720, 2139, 2173]. Pufferfish [2175]. Puget [1761, 2263]. Pugettia [2349]. Pulau [1516]. Pulling [1916]. Pulse [1176, 24]. Pulses [383, 822]. Pump [308, 647]. Pumping [2195, 2648, 1403]. Pumps [3147]. punctatus [636, 2436]. Punctual [2816]. Punic [3114]. Pup [2300]. Pups [1001, 1212]. Pure [2651]. Purification [1570]. Purple [803, 2658]. Purplish [3188]. Purpose [1531, 2447, 1772]. Purposes [2120]. Putative [2853, 1881, 3166]. Puzzle [2917]. Pycnocline [44]. Pygmy [2191, 1034]. Pygoscelis [1359]. pyrene [222]. pyrifera [2723]. Pyrodinium [3097]. Pyropia [32, 295, 496, 517, 22, 889, 2476]. Pyrosoma [1462]. Python [1358]. pyura [1995].

Qatar [909, 2479]. Qoli [2338]. QTL [2655]. quadrata [1414]. Qualitative [401]. Quality [2233, 376, 2231, 713, 2401, 2484, 1580, 277, 2498, 1943, 882, 2629, 161, 1988, 1906, 2827, 2326, 1494, 750, 2193, 1966, 2030, 305, 2496, 1399, 662, 2227, 1258, 1211, 2720, 889]. Quality-Controlled [2401]. Quantification [1928, 548, 2628, 2481]. Quantifications [2748]. Quantifies [2421, 1051]. Quantifying [1300, 1771, 1801, 690, 309, 2340, 506, 1477, 1013, 170, 737, 2547, 1954]. Quantitation [3038]. Quantitation-Based [3038]. Quantitative [1083, 2792, 401, 3184, 2217]. Quantitatively [3129]. Quantities [2308]. Quantity [376]. Queen [748]. Quenching [2370]. Quercetin [2927]. quercinus [3079]. Questions [1182, 1383, 1587, 678]. Quilt [1687]. quinquedens [178]. Quintana [2344]. Quorum [2310]. Quota [538]. Quotas [2966].

R [1758]. R.Br. [2875]. Ra [2538]. Rabigh [1892]. Racemization [76]. Rachycentron [2214]. RAD [259]. RAD-Seq [259]. Radar [1146, 1882, 312, 594]. radiata [2943]. Radiation [408, 2082, 2538]. Radiocarbon [276, 404, 76, 290, 1092]. Radiocarbonscapes [619]. Radioisotope [508]. Radiolaria [2116]. Radiometry [337, 2262]. Radionuclides [2538]. Rafted [2435]. raikovi [1323]. Rainbow [2455, 3027, 2777]. Raja [1127]. Ram [3114]. Raman [2839, 147]. Ranch [1953]. Ranching [1475]. Random [2457, 2355, 1366, 2146, 985, 2387]. Range [175, 1329, 2847, 1595, 2760, 769, 1579, 999, 1026, 977, 2545, 1428, 2027, 2823]. Range-Limits [2847]. Ranges [196, 1191, 2175]. Ranging [1491, 1001, 657, 446]. Rangiroa [1197]. Ranking [2220]. Rapa [2408, 2382]. Rapana [2408]. Rapid [1807, 1395, 1595, 2418, 418, 735, 977, 1465, 1982, 78, 804, 347, 2295, 2851, 114]. Rapidly [2571, 214]. Rapido [1367]. Rare [2150, 2393, 2514, 33, 1658, 526]. RAS [2642, 3011]. Rate [3104, 2840, 1596, 3084, 1248, 3061, 2755, 422, 2648, 1403, 2305, 1504, 1056, 1067, 3120]. Rates [2968, 163, 19, 172, 175, 1496, 2564, 61, 1006, 2683, 580, 2129, 528, 1602, 1490, 2984, 2893, 1056, 1223]. Ratio [2559]. Ratio-Derived [2559]. Ration [1663, 422]. Rationalization [2444]. Ratios [3110, 61, 450]. Ray [1127, 1801, 363, 2911, 3092]. Rays [827, 2395, 2088, 3098, 660, 3120]. Razor [1142, 3187, 118, 2758]. RCP [950]. RCP8.5 [1059]. RCS1 [2214]. RCS2 [2214]. RCS3 [2214]. rDNA [1924, 2880]. Re [649]. Re-draw [649]. Reach [1603]. Reaction [89, 3165]. Reactivation [176]. Reactive [1643]. Reactors [180]. Read [3182, 2658, 2980, 2388]. Real [203, 2524]. Real-Time [203]. Realignment [1918]. Realistic [584, 1803]. Realm [1784]. Realms [1741]. Reanalysis [3028, 2879, 2630]. Reared [864, 2880, 662, 631]. Recapture [698, 1278]. Receiver [2308]. Receptor [1955, 2084]. Receptors [112, 3166]. Recipes [961]. Recirculating [2642, 662, 3053]. Reclamation [2951, 3096]. Recognising [931]. Recognize [3129]. Recolonization [48, 602]. Recombination [2002]. Recommendations [232, 1583, 1852, 960, 634, 1683, 2412, 3033]. Reconciliations [2579]. Reconciling [1527, 270]. Reconstruct [610]. Reconstructing [134, 2256, 997, 566, 297, 1777, 783, 1334]. Reconstruction [2021, 2266, 2455, 1504]. Reconstructions [1814]. Record [440, 2018, 1882, 3051, 1590]. Record-Breaking [1882]. Record-Warm [440]. Recorded [3105, 472, 1824, 2989]. Recorders [2081]. Recording [44, 2290, 2120, 2829]. Recordings [256]. Records [2223, 332, 2503]. Recovery [1053, 1571, 1481, 1875, 2868, 1188, 448, 689, 2809, 917, 804, 347, 3, 1069, 751, 300]. Recreational [1142, 79, 116, 1072, 2665, 127, 344, 2834, 3087]. Recreationally [1595]. Recruit [1374, 1941]. Recruited [51]. Recruitment [2950, 1188, 101, 2610, 3198, 2295, 1428, 2027, 2860]. Recruits [1529]. Recurrence [2213]. Recurrent [2215]. Recurring [3202]. Recycling [965, 972, 1454, 553]. Red [3011, 2968, 699, 2150, 1110, 584, 828, 2920, 1721, 2266, 640, 1792, 63, 1797, 2729, 496, 2650, 178, 468, 1892, 172, 2583, 270, 1253, 1491, 181, 193, 996, 1658, 1720, 2635, 269, 3112, 1178, 2855, 1734, 993, 1657]. Redefine [339]. Redescription [1324, 2908]. Redistribution [1681]. Redondo [211]. Redox [2737]. Redox-Sensitive [2737]. Reduce [1360, 2089, 2427, 3135, 2238, 2033, 2082, 641, 2206, 1922, 1958, 2712, 898]. Reduced [230, 2683, 2242, 2285, 1412]. Reduces [1255, 1225, 952, 2627, 2038]. Reducing [2539, 1579, 2631]. Reduction [2800, 663, 2229, 1096, 621, 1730, 2443, 87, 1871, 1983, 2194, 1839, 1931]. Reductive [149]. REE [217]. Reef [2799, 2028, 1095, 827, 1706, 2185, 1954, 1300, 2736, 2202, 585, 2689, 1708, 2684, 615, 550, 453, 351, 2699, 2246, 890, 2240, 589, 778, 57, 513, 2344, 1106, 2390, 1000, 414, 436, 2453, 1812, 1516, 1257, 1562, 1373, 13, 886, 2768, 997, 3201, 2587, 193, 1602, 2319, 2779, 1535, 2323, 2270, 1654, 633, 479, 1147, 348, 2568, 2803, 2037, 1798, 2097, 689, 1002, 2115, 1982, 214, 2536, 2709, 660, 1371, 1368, 2225, 248, 195, 758, 1947, 547, 1958, 838, 2103, 3128, 386, 2491, 774, 804, 347, 1336, 2441, 1473, 1352, 1201, 2291, 707, 1772, 1098]. Reef [2139, 301, 1404, 233, 3048, 1303, 2926, 3150, 1835, 1009, 466, 134, 1599, 2771, 1336, 170, 2017, 98, 1581, 2254]. Reef-Building [589, 1373, 3201, 2587, 1654, 2225, 1947]. Reef-Fringed [547]. Reef-Lined [1352, 301]. Reefs [466, 907, 146, 1032, 2220, 643, 1795, 1065, 2400, 2182, 590, 1094, 2567, 281, 1718, 2576, 1843, 72, 730, 770, 2771, 1693, 548, 1950, 2652, 3138, 448, 413, 616, 2545, 1796, 1002, 1959, 2610, 384, 2591, 3186, 2257, 549, 463, 711, 1962, 2376, 3, 1069, 2720, 975, 254, 970, 289, 1814]. Reefscape [1795]. Reference [1270, 1474, 1399, 328, 983, 1870]. Refinement [1096]. Refining [1445, 2049]. Reflect [268, 1725, 3165]. Reflectin [1800]. Reflecting [466]. Reflections [1649, 1361]. Reflectors [641]. Reflects [654]. Refractory [1834]. Reframing [2883]. Refrigerated [2093]. Refuge [1765]. Refugia [1190, 2544, 2443, 616]. Regeneration [1451]. Regime [186, 1029, 2796, 2582, 1668, 2830]. Regimes [2594]. Region [1320, 2621, 863, 2122, 2643, 1340, 879, 2487, 2779, 2353, 2555, 2004, 1046, 2742, 848, 2997, 360, 2416, 742, 3093, 2602, 546, 2925]. Regional [1873, 3119, 570, 2203, 1320, 708, 2621, 2903, 2691, 589, 1877, 1692, 2244, 2122, 722, 1929, 128, 2019, 2570, 2714, 1004, 1489, 2384, 1594, 1522, 1957, 2353, 1959, 2549, 950, 660, 355, 2245, 1077, 3194, 1091, 1794, 2444, 1210]. Regionalization [1421]. Regions [2377, 685, 2183, 2897, 1874]. Registry [2522]. regius [1068]. Regression [2711, 2845, 2207]. Regularized [2711]. Regulate [2971, 2617]. Regulated [3160]. Regulates [1469, 2710, 2134, 1199, 458]. Regulating [2310, 492, 2084]. Regulation [294, 1846, 569, 2791, 1111, 2945, 1487, 1732, 3212, 2408]. Regulations [3202, 965]. Regulator [2451]. Regulatory [2927, 2292, 1748, 3127]. Rehabilitated [3105]. Rehabilitation [429, 505]. reidi [2380]. Reina [13, 1098]. Reinforcement [1349]. Rejuvenated [1117]. Relate [699]. Related [243, 1414, 1030, 108, 659, 1194, 2915, 36, 1653, 2655, 1537, 825, 1101, 2961, 969, 2708, 2665, 626, 1558, 2855, 2758]. Relation [1539, 704, 1250, 3104, 589, 385, 2302, 3162, 662, 2519, 167]. Relational [2224]. Relationship [279, 1037, 1936, 316, 2840, 2813, 1626, 208, 355, 2609, 2445, 3071, 3064]. Relationships [2028, 2224, 381, 1000, 2837, 284, 1965, 2829, 3137, 3127, 1227, 1440, 1207, 2752]. Relative [2972, 615, 1636, 1695, 1913, 753, 3038, 1995]. Release [2208, 2699, 94, 2476, 864, 1571]. Released [1140, 2757]. Relevance [3057]. Relevant [469, 573, 2171, 1944, 1230, 55, 490]. Reliable [2418, 210]. Reliance [1802]. Relocatable [1839]. Remains [970]. Remarkably [904]. Remineralization [2474]. Remodeling [1174]. Remote [2233, 842, 2231, 1566, 2060, 2400, 2975, 1930, 1927, 1507, 1353, 1236, 418, 2290, 2109, 2487, 2036, 622, 67, 147, 623, 1637, 1263, 1816, 439]. Remotely [771, 322, 1477, 2720, 1696]. Removal [231, 2994, 1727, 1546]. Remove [539]. Removing [73]. Renewable [2155, 2885]. reniformis [1280]. Reoxygenation [569, 3055, 3121, 1258]. Repeat [1093]. Repeated [728, 689, 2169, 984]. Repertoire [2094, 2658]. Replaced [3130]. Replacement [2595, 3205, 1096, 161, 1959]. Replicate [2159]. Repopulate [1370]. Report [3123, 3114, 2290, 111, 1674]. Reporting [3144]. Reports [2202]. Repository [3013]. Representation [2799, 1793]. representative [2711]. Representativeness [1230]. Representatives [262]. Representativity [934, 2050]. Reproducible [1399]. Reproduction [1940, 2678, 215, 2563, 50, 2754, 2936, 2159, 2357, 1302]. Reproductive [376, 2484, 1196, 2068, 2958, 204, 1048, 1540, 1412, 662, 496, 2357, 3092, 1028]. Republic [2133, 371, 2159, 93]. Requirements [1474, 2297]. Requires [1488, 2394]. Rescue [159, 1996, 1683]. Research [91, 34, 586, 857, 1186, 1649, 1642, 1566, 1992, 3147, 486, 2566, 451, 282, 1749, 1561, 1047, 2137, 109, 1617, 2418, 1555, 2092, 1915, 1852, 2637, 1305, 1460, 145, 431, 2036, 297, 791, 2636, 2171, 2396, 2464, 1383, 8, 84, 1832, 1910, 1883, 679, 2411, 2702, 2123, 2440, 3013, 1707, 2143, 217, 1833, 1288, 3030, 2157, 1466, 1348, 3020, 678, 2055]. Researchers [794, 1385, 2697, 630]. resedaeformis [315]. Reserve [748, 82, 1797, 1621, 3142, 936, 2567, 1245, 487, 3113, 1721]. Reserves [2803]. Reservoir [2057, 2504, 3182]. Reshaping [2574]. Residence [676, 2559]. Residency [2902, 1957, 525]. Resident [2089, 1181, 1621]. Residents [225]. Residues [1681]. Resilience [1954, 1300, 133, 1094, 2733, 675, 51, 1435, 3007, 2270, 2635, 448, 689, 1726, 2237, 1230, 937, 2199, 3, 1069, 2731, 625, 1776, 1725]. Resilient [1022, 2732, 2584]. Resistance [1233, 2329, 2589, 196, 3061, 2635, 29, 3148, 2680]. Resolution [854, 1616, 2151, 987, 2428, 337, 704, 131, 1930, 2879, 2046, 1507, 1160, 1267, 877, 235, 735, 3097, 365, 2069, 205, 2518, 1506, 2296, 595, 1180, 1134]. Resolve [1757]. Resolved [120]. Resolving [1195, 2630, 3109]. Resonance [883]. Resorts [1406]. Resource [2802, 1669, 2514, 1695, 3134, 1264, 1974, 481, 2236, 1859, 274, 953, 1283, 1784, 405, 2898]. Resources [1499, 2238, 1393, 2047, 470, 981, 1029, 510, 1193, 1866, 982, 2702, 2123, 2445, 1482, 1618]. Respiration [1248, 633, 239, 1014, 1342]. Respirometry [1006, 1084]. Respond [2304]. Responding [200, 1162]. Responds [1213]. Response [1616, 2071, 2930, 2802, 2074, 1845, 53, 1256, 2889, 2595, 1349, 2786, 501, 2034, 440, 2724, 3205, 1517, 1812, 2982, 3157, 2006, 2361, 1826, 3056, 3015, 51, 850, 1565, 1535, 830, 3044, 2323, 2517, 1552, 68, 2024, 2142, 3208, 425, 183, 1412, 593, 2117, 3130, 3021, 1389, 3002, 1478, 938, 122, 2671, 865, 125, 358, 253, 1040, 95, 1069, 1241, 2311, 372, 2814, 2423, 2726, 65, 2951, 715, 2124, 3040, 3204]. Responses [427, 2091, 376, 2235, 1425, 1598, 2905, 856, 1923, 3003, 976, 3202, 2417, 2607, 3188, 130, 1904, 2460, 700, 2363, 1770, 1438, 572, 2806, 1913, 1086, 2598, 1168, 736, 2871, 1551, 1159, 2681, 2180, 3195, 433, 77, 2145, 248, 841, 1230, 2833, 29, 496, 2452, 409, 705, 707, 2946, 2888, 2644, 2470, 3193, 2861].

Responsibility [2695]. Responsive [32, 2258]. Restoration [885, 2071, 586, 286, 1921, 1228, 1363, 1708, 2611, 339, 3199, 890, 853, 646, 429, 1645, 414, 1721, 1301, 2950, 3057, 1562, 2013, 1602, 634, 177, 430, 2270, 2241, 1588, 2522, 817, 758, 540, 2491, 774, 536, 692, 870, 719, 194, 174, 1879, 1618]. Restorations [1958]. Restorative [824]. Restore [339, 826]. Restored [1918]. Restoring [845, 174]. Restricted [2796]. Restriction [735]. Restructure [1306]. Result [1628]. Resultant [411]. Results [1186, 670, 914, 2340, 594, 1479]. Resurrection [2908]. Resuspension [1681, 2138, 2765]. Retelling [269]. Retention [1914, 474, 1427]. Reticulum [2927, 465]. Retracing [698]. Retrieving [445]. Retrospective [2320]. retroversa [1398]. Return [1561]. Réunion [1201]. Reuse [1884]. Reveal [827, 2974, 1464, 1110, 1425, 1329, 1022, 2116, 1351, 1584, 2186, 2791, 1693, 1152, 1101, 2298, 26, 2496, 259, 1764, 2525]. Revealed [381, 1750, 615, 2575, 2272, 1831, 1819, 2967, 2866, 2725, 1159, 3098, 1990, 3112, 595, 2125, 1227, 1736, 3160, 2758, 618]. Revealing [3162]. Reveals [224, 230, 1713, 1243, 522, 2691, 2003, 206, 1249, 1624, 1512, 3179, 1100, 3101, 2848, 2375, 2868, 1568, 877, 1017, 1926, 2477, 2102, 51, 558, 330, 2961, 2114, 2223, 1216, 1888, 384, 1540, 1822, 938, 3128, 1734, 1800, 123, 2658, 360, 3149, 2907, 3184, 1034, 1989, 1287, 476, 2489, 1618]. Revenue [762]. Reverse [860]. Reversible [569]. Review [1350, 218, 601, 1994, 1228, 2975, 2872, 3197, 108, 3041, 881, 1405, 2467, 1100, 555, 2789, 2539, 1555, 3070, 2385, 891, 1653, 2637, 2659, 738, 2042, 2721, 2097, 2221, 1015, 600, 2910, 2558, 1905, 2921, 394, 3109, 3030, 536, 64, 2067, 692, 1148, 2275, 319, 2335, 1318, 2858]. Reviewed [203]. Reviewing [1521, 2192, 402, 28]. Revision [2976]. Revisit [1900]. Revisited [589]. Revisiting [1948]. Rewilding [2199]. Reykjanes [1357]. rhadinum [2925]. Rhincodon [2324, 2900, 1221, 1048]. Rhine [114]. Rhizaria [687]. Rhizocephala [851]. Rhizophora [3002]. Rhizosphere [2670]. Rhizosphere-Competent [2670]. Rhizostome [2970]. Rhizostomins [2970]. Rhodobacterales [2310]. Rhodoliths [1190, 111, 1686]. Rhodomonas [948]. Rhodophyta [1110, 32, 111, 468]. Rhodosorus [111]. Rhombic [313]. Rhopalonematidae [63]. Rhythm [314, 2014, 2901, 2904]. Rhythms [2901, 2726]. Ría [1643, 2046]. Rías [2335]. Ribosomal [1243]. Rica [809]. Rican [1104, 1361]. Rich [550, 1249, 3089, 135, 1297, 2128]. Richness [589, 681, 2551]. rickettsi [3075]. Rico [2812, 2173, 2720, 2139]. Ridge [48, 2451, 1016, 1357, 3064, 1994, 131, 670, 1088, 956, 592, 980]. Ridges [2910]. Ridley [316, 403, 661, 36, 1887, 355]. Right [1360, 2086]. rigida [156]. Ringed [2525]. Rings [1224]. Río [418, 2510]. Rise [2100, 1377, 2733, 2363, 1582, 2869, 2569, 2261, 2548, 2171, 1701, 2448, 1364, 97]. Rising [1624, 2627, 2624, 396]. Risk [963, 1541, 1539, 1360, 2778, 2919, 2990, 2601, 1322, 3132, 1700, 1377, 855, 2539, 2785, 1237, 212, 2842, 1335, 2153, 2300, 1866, 3186, 1408, 988, 58, 1922, 170, 1091, 1466, 1839, 2748, 3068]. Risk-Averse [1377]. Risks [682, 2203, 1991, 2253, 2426, 1869, 2668, 164]. Risky [983]. Risso [385]. River [163, 2511, 1192, 1547, 2715, 527, 2813, 1589, 1471, 2403, 2646, 2056, 3097, 1906, 2703, 2433, 2366, 1902, 2303, 2298, 3208, 2876, 2796, 2404, 1985, 2657, 2720, 1736, 2594, 1074, 2603, 2675, 1672, 3176, 2811, 2606, 2289, 609, 1134, 338, 361, 1276, 2532, 3017, 2196, 1370, 3112, 1, 790, 395, 2929, 3040, 1542, 893]. River-Dominated [1547, 1589, 1736]. River-Estuary [527]. River-Influenced [1985]. Riverine [192, 434, 2646, 2679, 781]. rivulatus [2285]. RNA [825, 2183, 3175, 2759, 2392]. RNA-seq [825, 3175]. RNAi [1285, 2114]. RNAs [3189]. RNAseq [1017]. Road [3197]. Roadmap [2611, 2171]. Robotic [45]. Robots [2421]. Robust [913, 289]. robusta [294]. robustus [1148]. Rocas [1436]. Rock [2393, 680]. Rockfish [3099, 2062, 197]. Rocks [2393]. Rocky [1065, 2421, 2119, 1413, 413, 1585, 1836, 1796, 1275, 1756, 655, 1369]. Role [2096, 364, 1921, 2482, 2235, 2173, 1978, 3123, 299, 1094, 569, 1061, 112, 502, 811, 2035, 1087, 2256, 650, 1202, 1188, 512, 73, 1111, 1103, 2866, 2114, 479, 2077, 2581, 2721, 363, 1114, 2831, 880, 2541, 1656, 1587, 307, 3187, 785, 3149, 884, 1665, 1664, 2858, 1439]. Roles [2479, 1713, 2972, 1746, 2417, 3183, 2876, 2892, 2188, 1333, 687, 2838]. Romanian [2913]. Roo [2344]. Rookeries [1734]. Room [1122]. Root [3057, 1220]. Roots [479, 3109]. ROS [512, 880]. rosenbergii [3127]. Roseobacter [2310]. Ross [1189, 2815, 1030, 1680, 2942, 1054, 2714, 2023, 2547]. Rossby [1075]. Rossellidae [1493]. rostrata [905]. Rotation [575]. Rotifer [1068, 1515, 2835]. Rotifers [1345]. rotundicauda [2967, 2093, 1076, 1498]. Rough [653]. Roughness [766]. Round [2906]. Route [516]. Routine [500, 1722]. Routing [611]. ROV [771, 256]. Roving [374]. ROVs [322]. Royal [1925]. rRNA [2607, 2843]. Rubble [1009, 2554]. Rubrioxytricha [1341]. rubrum [584, 1721, 1797]. Ruditapes [3136, 2359, 2417]. Rules [1824, 754, 2058]. Ruminant [663]. Running [168]. Runoff [418, 2969, 2594]. Runup [463]. Rural [1701]. Russian [1518].

S [445, 500]. Sabellaria [1032, 1835]. Sablefish [3126, 1241]. Sac [358]. Saccharina [2462, 859, 1715, 930, 1168, 35, 1135]. Sacculina [851]. Sacculinidae [851]. Safe [2443, 1538]. Safety [539]. Saga [2701]. Sagar [3191]. sagax [628]. Saildrone [1813]. Sailing [2099]. Saint [2322, 2137]. SalaciaML [1494]. salar [1262, 1823, 2639, 3056, 2098]. Salinities [2208, 3175, 2213, 2357]. Salinity [500, 2350, 2393, 473, 2564, 527, 3152, 675, 2204, 929, 2415, 572, 2430, 932, 1073, 2945, 2909, 1220, 2984, 2101, 2580, 1023, 3002, 86, 2573, 517, 802, 2822, 2946, 806, 2826, 2489]. Salinity-Conveyed [1023]. Salish [682, 2089, 1561, 3168]. Salmo [1262, 1823, 2639, 3056, 2098]. salmoides [299, 3065, 1199]. Salmon [1262, 2190, 1823, 2639, 1568, 1285, 2534, 30, 3052, 829, 1533, 3056, 161, 1008, 1354, 2642, 2098, 1239, 444]. salmonicida [2899, 2899]. Salmonid [2450]. salmonis [3052]. Salt [942, 2839, 2887, 2164, 2049, 598, 2891, 2588, 2516]. Salt-Marsh [2891]. Saltmarsh [3007, 2604]. Saltmarshes [2733, 2769]. Sample [2522, 1999, 1339, 2434]. Sampled [2554, 291]. Sampler [1652, 1265]. Samples [2711, 1924, 550, 760, 914, 1684, 1265, 2875]. Sampling [2235, 2400, 3106, 16, 877, 2941, 1145, 3023, 2236, 2977, 2572, 459, 2472]. San [1220, 3074, 936, 189, 1794, 1732]. Sanctuary [2717, 677, 1211, 1290, 1983]. Sand [2651, 1851, 2856, 1866]. Sand-Mud [2651]. Sander [2304]. Sandwich [1961, 1926]. Sandy [3031, 1414, 1022, 1158, 1826, 526, 637, 1538, 1835]. Sandy-Bottom [1835]. sanguinea [2889]. Sansha [3009]. Santa [2668, 3074]. Santo [1284]. Sanya [2779]. São [2222, 1016]. sapidus [735, 2712]. SAR [83]. Sarasota [1915, 1534, 3170, 2635]. Sardina [2678, 978]. Sardine [1815, 2678, 2356, 978]. Sardines [628, 2228]. Sardinian [2334]. Sardinops [628]. Sargasso [2132]. Sargassum [2735, 1140, 2519, 999, 3090, 219, 1657]. sargus [489, 2302]. SARS [2153, 2156, 3143]. SARS-CoV-2 [2153, 2156]. SARTRAC [2735]. Satellite [3011, 827, 224, 607, 2872, 2461, 494, 1843, 2453, 661, 1755, 212, 925, 2512, 2842, 1105, 87, 677, 1900, 906, 2931, 1734, 165, 701, 911, 1263, 1501]. Satellite-2 [2453]. Satellite-Based [607, 2461]. Satellite-Derived [2872, 925, 2931]. Satellites [842, 1930, 2384]. Satiation [1469]. Satsunan [2756]. Saudi [2266, 702, 270, 2373]. Save [2511, 1326, 1857]. Saving [1996, 1385]. Scaffolding [1314]. Scale [153, 1414, 2294, 1541, 2299, 2091, 3191, 1795, 1987, 575, 3025, 2173, 852, 638, 2903, 813, 1927, 1492, 557, 451, 206, 1901, 853, 3135, 2277, 1760, 124, 2383, 414, 1441, 3069, 419, 3081, 621, 13, 2901, 2768, 2477, 1809, 1594, 2232, 2394, 1101, 3178, 1413, 1434, 2631, 2627, 2495, 1888, 1842, 2343, 1775, 2162, 1015, 1060, 78, 1858, 2722, 756, 1686, 1368, 2146, 1593, 2257, 1848, 1361, 858, 1275, 2447, 259, 1511, 2966, 27, 1609, 1844, 2814, 93, 2647, 1702, 1062, 2602, 3022, 1618, 783, 2790, 2188, 1333]. Scale-Specific [2232]. Scale-Worms [1987]. Scaled [475, 1856]. Scaled-Up [475]. Scales [2578, 1843, 2011, 268, 637, 594, 1685]. Scaling [800, 1648]. Scallop [2877, 1951, 1651, 2725, 1805, 2886]. Scalloped [1630]. Scallops [2806]. Scaly [3058]. Scaly-Foot [3058]. Scan [572]. Scandola [1721]. Scanning [1602]. Scape [1975]. Scapharca [2861]. Scapulocoracoids [393]. Scarcity [991]. Scat [3160]. Scatophagus [3160]. Scatterers [873]. Scattering [155, 2051, 1116, 2862, 412, 1717, 595]. Scavenger [560]. Scavengers [1986, 1692, 512, 1765]. Scavenging [2578, 2949]. Scenario [1441, 2895, 830, 910, 950]. Scenario-Based [1441]. Scenarios [682, 947, 741, 2763, 1377, 1430, 928, 1393, 3015, 981, 1893, 982, 1432, 865, 301]. Scenes [1376]. Schaefer [532]. Scheduled [2977]. Schemes [634, 2216]. Schizochytrium [161]. schlegelii [3099, 1012, 197, 3141]. schlosseri [1335]. Schmidt [1493]. School [2247]. Schools [2845]. Sciaenidae [2446]. Science [2799, 225, 2247, 794, 2099, 486, 2566, 1488, 261, 1580, 2753, 2347, 931, 1591, 1831, 144, 1862, 624, 3116, 2315, 193, 1182, 2353, 2464, 1406, 2025, 1832, 2407, 2850, 1999, 1910, 442, 1578, 894, 1829, 2075, 2260, 1133, 1459, 1816, 2278, 2553, 1318]. Science-Based [931]. Science-Policy [2353, 1133]. Sciences [1015, 2210, 2055]. Scientific [668, 584, 746, 1055, 322, 2702, 2123, 980]. Scientist [511, 1182]. Scientists [1552]. scintillans [2563, 3137, 993].

Scleractinian [563, 2432, 3010, 955]. Sclerochronology [297]. Sclerotization [1783]. Scomber [1523, 2756]. Scomberomorus [3104, 935]. Scope [2517, 205]. Scophthalmus [1115, 1179, 2423, 300]. Scoping [2353, 536]. Score [1722, 1683]. Scotia [2243, 37, 694, 37]. Scotian [2475, 1217, 729]. Scotland [240, 1831, 2451, 690]. Scottish [1700]. Scour [1275]. Screen [2459]. Screening [3046, 2481]. scruposa [231]. SCS01 [2868]. SCTLD [2319, 1216, 2180, 2107, 2673]. Scuticociliates [1261]. Scylla [2003, 2772, 851, 2866, 2416, 2939, 3166, 2826]. Scyphozoa [2663, 196, 2970, 2893]. SDG [2704, 1711]. SE [3143]. Sea [2334, 3011, 471, 2100, 763, 240, 153, 2151, 146, 1815, 1045, 1456, 2698, 515, 864, 1539, 1921, 1311, 1052, 2070, 1846, 1030, 1987, 1186, 1712, 336, 1924, 1690, 2990, 1476, 776, 1189, 2815, 2859, 638, 767, 270, 2847, 62, 131, 2437, 1938, 1492, 757, 2026, 486, 3197, 544, 945, 497, 2916, 1094, 261, 1377, 1923, 1928, 1580, 3181, 895, 403, 1782, 1225, 795, 881, 554, 545, 2248, 1424, 653, 2467, 202, 2132, 2001, 814, 2382, 1830, 1763, 2853, 2266, 710, 1236, 2837, 277, 833, 1359, 2092, 661, 621, 1244, 3095, 2941, 2186, 1149, 75, 432]. Sea [2451, 2733, 1184, 839, 1202, 527, 345, 2895, 1483, 1058, 1120, 2009, 1768, 1689, 2960, 2349, 310, 28, 2949, 829, 1533, 1305, 2363, 2542, 1438, 6, 92, 528, 1582, 237, 974, 925, 571, 1196, 1887, 1029, 2068, 1289, 784, 2869, 2569, 803, 2617, 2339, 1919, 693, 2546, 2348, 2844, 808, 2708, 1997, 2261, 2355, 1366, 1150, 1788, 581, 423, 1050, 1314, 2776, 285, 318, 934, 1220, 1451, 1626, 2548, 2425, 2171, 3023, 2300, 437, 684, 1409, 1701, 1046, 1238, 2024, 1266, 365, 747, 2000, 2497, 1840, 2556, 2101, 2977, 2624, 1967, 2215, 2059, 904, 1317, 214, 1799, 1744]. Sea [412, 2118, 2090, 1717, 1898, 26, 1016, 2474, 341, 2706, 2766, 82, 917, 355, 3162, 1067, 2448, 1080, 1356, 366, 158, 1785, 533, 1944, 530, 198, 2936, 2364, 2609, 37, 1912, 594, 865, 1656, 694, 2998, 103, 1748, 845, 493, 1075, 2924, 2618, 126, 2574, 2017, 354, 1620, 323, 307, 1172, 1674, 1153, 695, 2125, 2669, 81, 2997, 1784, 2239, 1764, 178, 2740, 199, 154, 2504, 1931, 184, 970, 289, 405, 843, 2472, 1440, 1136, 715, 1709, 2860, 2547, 1860, 2289, 1260, 1205, 3204, 1482, 1890, 1203, 1696, 1325, 2233, 2507, 3011, 2968, 1378, 1892, 2596, 163]. Sea [2783, 172, 699, 1367, 2583, 2324, 3028, 1102, 2402, 2678, 682, 2150, 577, 1773, 2919, 230, 3203, 807, 2089, 1566, 1127, 484, 152, 3123, 2859, 270, 3124, 815, 2807, 2775, 2865, 2692, 873, 1024, 2879, 1680, 1253, 2942, 206, 2272, 828, 1518, 1561, 440, 2277, 2502, 1424, 2821, 456, 1083, 2085, 722, 1929, 919, 1778, 2488, 3024, 1054, 833, 2456, 621, 1547, 2714, 837, 2014, 181, 1066, 7, 1346, 1163, 1742, 3192, 2660, 1251, 1248, 2656, 1340, 879, 534, 2363, 1489, 2542, 3168, 714, 3177, 1525, 1296, 193, 2430, 1396, 1299, 1802, 1537, 2232, 1426, 1103, 335, 2313, 1976]. Sea [3145, 321, 696, 1889, 2978, 3117, 996, 1658, 2630, 3196, 2871, 1586, 1733, 1714, 959, 1997, 912, 3131, 1720, 645, 977, 269, 2829, 2201, 10, 2495, 2213, 2025, 1897, 1273, 1240, 2215, 3137, 2011, 2876, 526, 950, 1858, 1900, 2809, 724, 766, 1694, 38, 551, 1946, 227, 680, 2519, 1506, 2672, 1178, 280, 2609, 1984, 89, 1013, 665, 2144, 1427, 2729, 1129, 187, 1748, 253, 2931, 1734, 1223, 2780, 1283, 2649, 1777, 1019, 426, 2913, 1153, 613, 2526, 319, 781, 118, 199, 2007, 3036, 840, 1613, 1917, 614, 788, 764, 3141, 3133, 2976, 715, 574, 2449, 2846]. Sea [1860, 844, 182, 546, 1303, 2015, 1657, 2316, 2696, 1342, 3043, 1484, 878, 1468]. Sea-Air [1997, 1317]. Sea-Cage [747, 1799]. Sea-Ice [1189, 2815, 638, 554, 345, 1709, 1860]. Sea-Level [2100, 1377, 1928, 2733, 2363, 1582, 2261, 2548, 2171, 2448]. Sea-Surface [2382, 2766]. Sea-Urchins [1921]. Seabass [2688, 2927, 1515]. Seabed [1807, 292, 2585, 592, 1784, 1641, 1814]. Seabird [2325, 103]. Seabream [408, 489, 1988, 1799, 2452, 2258]. SEADRAGON [1365]. Seafloor [1712, 2328, 1598, 2728, 2390, 480, 690, 784, 835, 506, 845, 113, 1074]. SeaFlower [936, 748, 82, 3113, 487]. Seafood [2209, 1924, 278, 1624, 1222, 2851, 1049]. Seagrass [172, 1053, 1433, 2800, 105, 1363, 2693, 2691, 1676, 2467, 3091, 730, 1031, 296, 1809, 798, 602, 3035, 1667, 77, 1429, 954, 2022, 1726, 865, 3109, 1780, 719, 2291, 194, 3211, 2788, 3019, 402, 1989, 1028, 1082, 1879, 115, 221]. Seagrass-Associated [1433, 2022]. Seagrass-Lucinid [115]. Seagrass-Sea [2467]. Seagrasse [1478]. Seagrasses [2923, 115]. Seahorse [2380, 2429]. Seahorses [740]. Seal [421, 839, 1219, 245, 1212]. Seals [1933, 2714, 2263, 989, 1740, 2525, 3190, 2023]. Seamount [2301, 121, 1806, 1245, 1117, 1557, 1765, 2425, 956, 1508, 1479, 2832, 176]. Seamounts [146, 668, 2039, 2060, 1100, 3069, 803, 340, 518, 1597, 970, 289, 9, 2105]. Search [159, 1683]. Seas [552, 106, 960, 1582, 1614, 593, 2117, 3164, 906, 3086, 2506, 2281, 2507, 619, 819, 234, 232, 659, 2132, 1109, 3077, 1004, 1438, 1374, 1400, 173, 2828, 679, 2922, 1674, 1229, 752, 975, 266, 88]. Seascape [771, 3082, 3179, 2892, 970]. Seascapes [1410, 677]. Season [2535, 2288, 357, 1172, 2084]. Seasonal [2387, 1815, 946, 2800, 279, 1978, 708, 1678, 2371, 1189, 2815, 2807, 351, 1813, 2575, 2026, 494, 1308, 1022, 895, 1061, 2196, 1601, 2110, 1217, 1221, 1163, 1589, 1471, 2561, 101, 2713, 321, 1157, 1374, 969, 966, 1064, 2735, 1502, 1990, 944, 543, 1650, 1603, 2090, 2893, 3164, 506, 938, 525, 780, 1089, 694, 493, 2855, 606, 239, 2376, 426, 1980, 2506, 564, 2191, 142, 1556, 191, 2212, 2029, 1917, 788, 1825, 1260, 2815, 2475, 3018]. Seasonal-to-Interannual [895]. Seasonality [275, 1612, 2272, 3151, 874, 1967, 214, 2672]. Seasonally [2590]. Seasons [558, 2876]. SEATS [2931]. Seawalls [2237]. Seawater [1093, 1209, 2567, 3041, 1225, 3102, 33, 333, 2839, 2288, 2779, 1119, 2550, 2069, 2668, 2307, 1151, 1999, 1371, 2340, 21, 2998, 1170, 2481, 1664, 3204, 2642]. Seawater-Dissolvable [2998]. Seaway [1198]. Seaweed [522, 415, 572, 2386, 999, 1511, 2719, 889, 164]. Seaweed-Derived [415]. Seaweeds [2864, 1388, 1033, 2011, 3118, 1177, 29]. Sebastes [3099, 1012, 197]. Seco [2139]. Secondary [880, 1545]. Section [3017]. Sector [2794, 1347, 1979, 919, 1949, 2851]. Sectoral [947]. Sectors [888, 93]. Secure [1253]. Security [1964, 2013]. Sediment [1053, 267, 2226, 3047, 2377, 314, 390, 1710, 2990, 1670, 2486, 1972, 1981, 2607, 2873, 2979, 2183, 2172, 2703, 1871, 1402, 1999, 1051, 2267, 2706, 598, 1726, 2891, 530, 1925, 1500, 2769, 2980, 3036, 1732, 2597, 1493, 3040, 2801, 1439, 2773, 2138, 2765].

Sedimentary [619, 279, 690, 2995, 1342]. Sedimentation [1002]. Sediments [308, 104, 1045, 2435, 276, 1037, 331, 31, 1424, 3089, 2513, 2785, 674, 2463, 1547, 3000, 528, 2313, 2737, 2355, 1366, 2405, 1679, 1881, 626, 1454, 860, 1785, 2559, 302, 2913, 521, 2198, 1747, 2284, 1124, 2881, 3176, 2811]. Seed [2773]. Seeding [715]. Seedlings [427, 2623]. Seeking [2566]. Seems [2221]. Seen [1749]. Seep [343, 2355, 1366, 1411, 211, 3129, 1136]. Seepage [1826, 2978]. Seeps [346, 2437, 3192, 2978, 645, 178, 2504]. Segment [131]. Segmentation [2591]. Segmented [1399]. Segregation [2981]. Seismic [2350, 2955, 2006, 1557, 89, 2818, 2822]. Seismically [2478]. Seismicity [176]. Selaroides [2999]. Select [1128, 761]. Selected [2152]. Selecting [1708]. Selection [316, 1316, 1695, 2718, 2430, 808, 2473, 67, 2491, 1854, 323, 1382]. Selective [1763, 2993, 1684, 3103]. Selectively [2170]. Selectivity [2486, 162, 2014, 2967]. Self [187, 733]. Self-Organization [733]. Self-Sustainable [187]. selseyense [3140]. Semantic [2591]. Semi [1157, 2827, 1729, 3043]. Semi-Enclosed [1157, 2827, 3043]. Semi-Synthetic [1729]. Semibalanus [3198]. Semienclosed [2693, 1061]. semilaevis [65]. Semimytilus [450]. Seminavis [294]. senegalensis [2494]. Senegalese [2494]. Senescent [97]. Sense [640]. Sensed [2720]. Sensing [2233, 2310, 842, 2231, 2400, 2975, 2639, 1507, 1353, 418, 622, 67, 147, 623, 1263, 1816, 1199, 439]. Sensing-Based [418]. Sensing-Driven [67]. Sensitive [515, 107, 1017, 342, 2737, 1143]. Sensitivities [1520, 1214, 705]. Sensitivity [2203, 3084, 1635, 1105, 1847, 1683, 81]. Sensor [2485, 2069, 1172]. Sensors [1867, 39, 2959, 3139, 2474]. Sent [179]. Sentinel [437, 494, 2508, 1381, 1404]. Sentinel-1 [1381]. Sentinel-2 [494, 2508, 1404]. Sentinels [1753, 1740]. Separate [1750]. Separated [2130]. Separating [2825]. Separation [860]. Sepia [215, 61, 1800]. Sepiolidae [1351]. Sepioteuthis [351]. Septifer [3188]. Seq [259, 825, 3175, 2545, 2514, 274]. Sequencing [1808, 1924, 550, 1243, 2003, 2514, 2841, 1264, 2843, 2655, 26, 274, 3182, 2658, 2980, 2388, 1542]. Sequential [1145, 180]. Sequestration [3075, 1715, 3012, 3057, 2414]. Series [2041, 1237, 1666, 1089, 3048]. Seriola [2264]. Serious [1463]. Serotonin [2084]. Serpulidae [849]. Serrana [748, 82]. Serranilla [82]. Sertraline [1255]. Serum [1688, 95]. Service [2763, 902, 1998]. Services [2778, 2601, 1322, 2872, 2346, 555, 1394, 1562, 186, 1375, 2557, 1521, 2548, 2585, 817, 2206, 902, 2584, 1605, 1861, 262, 2335, 88]. Sessile [1824, 1965]. Set [2802, 1249, 379, 1525, 1800]. Setipinna [3077]. Setting [1883, 869]. Settings [2437, 1520, 2487]. Settlement [1058, 1666, 952, 766, 2940, 785, 3204]. Settlers [3153]. Settling [1664]. setubalense [1504, 1506]. Setup [1084]. Seven [94, 2296, 686, 2287]. Seven-Year [94]. Several [2546, 2857]. Severe [861, 1442, 1381]. Sewage [1161]. Sex [472, 139, 2772, 1099, 2917, 3183, 1038, 2302, 35, 450, 1540, 2653, 2758]. Sex-Biased [2772, 35]. Sex-Differentiation [2917]. Sex-Related [2758]. Sex-Specific [139, 2772, 1038, 2302]. Sexual [1323, 2678, 1316, 2292, 1766, 1126, 2563, 1132, 2114, 2416]. Seychelles [371, 660]. Shad [3014, 2823, 656]. Shading [521]. Shall [3197]. Shallow [308, 1095, 1986, 2770, 2667, 445, 94, 1309, 1302, 1195, 1997, 2705, 1796, 2610, 904, 2954, 1307, 1947, 2671, 3085, 476, 1439]. Shallow-Water [1986, 2770, 1195, 2671, 3085]. Shallows [1074]. Shannon [1490]. Shape [2010, 666, 2120, 2554, 268, 2022, 1428]. Shaped [1600]. Shapes [583, 2995]. Shaping [364, 2173, 3210]. Share [3087]. Shared [932, 982, 2694]. Shared-Socioeconomic-Pathways [982]. Shares [538]. Sharing [485, 2695, 367, 1087, 2585, 1060, 2702, 2123]. Shark [1630, 2273, 370, 1646, 139, 228, 2095, 2902, 2915, 1290, 1125, 1695, 342, 734, 1126, 726, 2223, 1731, 2473, 363, 1840, 1092, 2911, 1540, 1382]. Sharks [765, 1865, 2387, 2324, 672, 2273, 1571, 1208, 2375, 2395, 1601, 524, 2900, 1663, 1221, 2477, 1008, 1354, 2562, 966, 1957, 375, 3098, 1038, 290, 1048, 1191, 525, 1854, 2083, 2381]. Sharp [887]. shasta [1713]. Shear [2651]. Shed [2331]. Shedding [1282, 2118]. Shelf [2821, 653, 1669, 2469, 1824, 1589, 441, 106, 528, 738, 1671, 246, 321, 1586, 1699, 1679, 2093, 906, 70, 2530, 1854, 81, 1839, 1736, 254, 1791, 2284, 777, 1134, 2551, 878, 1293, 389, 3181, 976, 2475, 1217, 2746, 137, 904, 729, 780, 315, 1668, 2506, 2211]. Shelf-Coastal [1839]. Shelf-Sea [653]. Shell [1316, 2920, 222, 846, 2080, 2843, 1418, 1442, 698, 1498, 2976, 2886, 2861]. Shelled [1728, 1860, 458]. Shellfish [1964, 1761, 824, 1700, 353, 2111, 957, 2153, 399, 2855, 2835, 439]. Shellfish-Farming [2855]. Shells [1876, 916, 2937]. Sheltered [1372]. Shelters [829]. Shelves [611, 1396]. Shetland [37]. Shift [1130, 1101, 736, 3034, 2796]. Shifted [3199]. Shifting [2883, 1194, 3177]. Shifts [793, 1595, 2253, 3062, 1070, 967, 2356, 2498, 2629, 3007, 999, 2059, 2145, 2225, 2227, 1089, 368, 595, 667, 1668, 1241]. Shimiella [2857]. Shining [1344]. Shinkaia [1136]. Ship [337, 3041, 2229, 58, 1003, 170, 539, 3206, 2284, 43, 1093]. Ship-Deployed [43]. Ship-Towed [337]. Shipboard [1399]. Shipping [739, 64, 1153]. Ships [965, 58, 2370, 1794]. Shipwreck [1925]. Shoaling [2818]. Shock [2761]. Shore [189, 795, 2994, 2129, 1413]. Shore-Based [189]. Shorebirds [787, 167, 1337]. Shoreline [2797, 3015, 428, 2396, 547, 2276, 536, 3173]. Shorelines [536, 2769]. Shores [1414, 1585, 655]. Shorewatch [1831]. Short [2955, 1006, 1225, 2724, 1068, 1423, 1919, 2844, 2250, 2101, 2215, 2058, 1607, 2445, 613, 20]. Short- [2250, 2215]. Short-Lived [2724, 2058]. Short-Streamer [2955]. Short-Term [1006, 1225, 1068, 1423, 1919, 2844, 2101, 1607, 2445, 613]. Shorten [2603]. Shortfin [2253, 1601, 1699, 2368]. Shotgun [1159, 2037]. Should [2168]. Show [572]. Showing [1631]. Shown [3032]. Shows [1761, 1903, 108, 1212, 289]. Shrimp [713, 2589, 3110, 3210, 962, 2792, 3061, 1264, 3145, 1631, 1132, 1073, 3183, 2114, 2314, 274, 2921, 2835, 123, 2938, 3038, 2372]. Shrimps [571, 2267]. Shutdown [2231]. Shuttle [2290]. Siberian [1296, 2828]. Sibling [2908, 631]. Siboglinid [1976]. Siboglinidae [545]. SIBS [1652]. Sicily [1118, 610, 1407]. Side [2230, 2915]. Siderophore [149, 1119]. Siderophore-Promoted [149]. Sides [2859]. Sieves [2856]. Siganus [2285]. Sight [1549, 271]. Sightings [2186, 2129, 2036]. sign [2595, 3146, 2124, 2870, 2501, 2388]. Signal [2371, 3181]. Signaling [3154, 703, 1199]. Signals [2630, 2160]. Signature [674, 386, 777]. Signatures [651, 2116, 1976, 2568, 836, 2891]. Significance [106, 966, 293, 2275]. Significant [2131, 94, 2725, 1041]. Significantly [2134]. Signs [1409, 1947]. sikamea [2501]. Silencing [2791]. Silica [2586, 2422]. Silicon [2520, 2116, 1945, 2480, 304]. Silky [1208]. Silled [495]. Silled-Fjords [495]. Silver [2603]. Similar [374, 2177]. Similarities [2614]. similis [133]. Simple [1301, 1413, 2215, 2058]. SIMS [2503]. Simulate [979, 2487]. Simulated [2800, 565, 1063, 77, 865, 409, 1144]. Simulating [2787]. Simulation [2711, 2706, 841, 1170, 1121, 1404, 2207]. Simulations [2112, 2679, 1683, 1381, 1210]. Simultaneous [2786, 513, 3183]. Simultaneously [823]. Since [1406]. sinensis [1728, 2792, 2880, 2810, 3166, 263]. Sines [277]. Singapore [2022]. Singing [2465]. Single [998, 1243, 2365, 1279, 3061, 47, 2776, 2412, 2600, 2505, 2388]. Single-Cell [1243]. Single-Molecule [2388]. Single-Nucleotide [47]. Single-Species [1279, 2776, 2600, 2505]. Single-Turnover [2412]. Singular [1272]. sinicus [1858]. Sink [1007, 435, 2657]. Sinking [1596, 99, 924]. Sinks [2923, 1467, 893]. Sinonovacula [2904, 3187, 2758]. SINTEX [895]. SINTEX-F [895]. Sipunculan [3051]. Sipunculus [2537]. Sirsoe [5, 2424]. Sister [1893]. Site [2273, 522, 316, 277, 1663, 2164, 808, 2025, 67, 1540, 38, 2491, 1734, 323, 211]. Site-Specific [522]. Sites [2387, 2133, 1643, 1801, 235, 1953, 2152, 132, 1357]. Siting [71, 126]. situ [1526, 2294, 406, 2112, 2400, 453, 39, 1006, 99, 2877, 1008, 1354, 2862, 3023, 2648, 1403, 2610, 1015, 745, 2069, 2169, 2893, 1504, 259, 2931, 2924, 1479, 184, 1709, 1260, 2138, 1890]. Situational [1085]. Six [1894, 1180, 220]. Size [1761, 867, 18, 391, 1783, 666, 1237, 877, 717, 1289, 1078, 1101, 1969, 2205, 2856, 2366, 2077, 2554, 3137, 384, 662, 227, 1821, 325, 3030, 1660, 978, 311, 3071, 2625, 2422, 289, 3093, 400]. Size- [877]. Size-Dependent [1237]. Size-Fractionated [867, 2366, 2422]. Size-Related [1101]. Size-Structure [717]. Size-Structured [2625]. Sized [1750, 2014, 2422]. Sizes [356]. Skagerrak [1674]. Skates [2088, 393]. Skeletal [2767, 195]. Skeletonema [1639]. Skeletons [490]. Skill [895, 1507, 599, 1683]. Skills [1522, 606]. Skin [408, 23, 1174, 871, 1880]. Skipjack [998, 2845]. Sky [808, 323]. SLC38 [2639]. SLC7 [2639]. Sleep [868]. Slicks [2382]. Slocum [203]. Slope [2812, 887, 2568, 546, 2551, 3143, 1458, 559]. Slopes [1296, 1016]. Slow [1420]. Slower [2229]. Small [1378, 3046, 153, 1027, 3191, 3025, 1787, 852, 185, 369, 3124, 1804, 2730, 557, 451, 856, 1084, 1901, 2515, 2986, 1760, 1181, 3101, 419, 2014, 13, 3189, 1471, 1809, 1169, 996, 2856, 1434, 2495, 2149, 1273, 2162, 1015, 1060, 3137, 78, 3208, 724, 2146, 2234, 1593, 1361, 858, 1275, 2447, 1459, 179, 2966, 1609, 1844, 93, 2647, 289, 2181, 1287, 2602, 1941, 783, 2790, 2188, 1333]. Small- [858]. Small-Scale [153, 3191, 3025, 852, 557, 451, 1901, 1760, 419, 13, 1809, 1434, 2495, 1060, 78, 2146, 1593, 1361, 2966, 1609, 1844, 93, 2647, 2602, 783, 2790, 2188, 1333]. Small-Sized [2014]. smalli [1261]. SmartBuoy [3139]. Smearing [1402]. Smith [2540]. Smooth [966, 1822]. SMS [237]. Snail [1316, 3079]. Snails [3058, 924]. Snake [302]. Snakes [2017]. Snapping [3145]. Snapshot [2407, 993, 2697]. Snout [388]. Snouted [1138]. Snow [2435, 704, 1575, 73, 578, 1418, 1967, 1827]. Snowball [639]. SNP [2772, 2655, 2125]. Snubfin [1320]. SOCCOM [2326]. Social [3180, 2379, 2640, 3191, 1464, 1867, 2797, 2274, 1939, 2347, 928, 746, 2455, 2704, 1711, 41, 2102, 1697, 1150, 583, 3067, 1114, 1822, 127, 1942, 2206, 2282, 2260, 2147, 1005, 3087, 2391, 2055]. Social-Ecological [3191, 2274, 928, 41, 1697, 1942, 2282]. Sociality [385]. Socializing [1450]. Socially [2094, 1519]. Societal [1488, 1521, 2411, 2194]. Societies [2103]. Society [176, 1236]. Socio [1332, 2795, 1393, 981, 982, 1905, 1230, 15, 1784, 3142]. Socio-Ecological [2795, 1905, 1230, 1784]. Socio-Economic [15, 3142]. Socio-Political [981, 982, 1393]. Socio-Spatial [1332]. Sociocultural [1724]. Socioeconomic [2561, 982]. Sodium [1971, 2579, 2870]. Sodium-Dependent [1971]. Soft [1388, 1728, 1519, 562, 2703, 1418, 901, 3137, 302]. Soft-Bodied [3137]. Soft-Bottom [1388]. Soft-Sediment [2703]. Soft-Shell [1418]. Soft-Shelled [1728]. Software [1349, 2552]. Software-in-the-Loop [1349]. Sognefjord [495]. Soil [2641]. Soils [860, 3200]. Soldier [2589]. Sole [2494, 1391, 65]. Solea [2494]. Solitary [2821, 2818]. Solomon [2265]. Soluble [3171]. Solute [1643, 2639]. Solution [2856]. Solutions [2576, 2327, 2638, 630, 2516]. Somatic [3189, 355]. Sonar [2235]. Sonder [2302]. Song [1470, 2605, 1130, 2163, 1604, 244]. Soot [2449]. Soothsaying [443]. Sooty [2906]. Sorting [736]. Sotalia [1185, 1633, 1112]. Sound [155, 2240, 2229, 3135, 2242, 2161, 2154, 2552, 2263, 2954, 1461, 1761]. Sounders [2290]. Sounding [3000]. Sounds [2071]. Soundscape [2154, 2963, 2217]. Soundscapes [2402, 1678, 2705, 2547]. Source [1328, 2330, 2864, 1823, 1769, 3161, 527, 296, 1234, 135, 2082, 2128, 2707, 280, 2574, 1019, 2657, 255, 871, 648, 3141, 1082]. Source-to-Sea [527]. Sourced [1811, 611, 2186]. Sources [1862, 587, 1299, 2403, 968, 2154, 1503, 2062, 2352, 836, 1748, 1467, 182]. Sousa [1392, 1969, 2205, 1380, 1829]. South [424, 2174, 3124, 2865, 2692, 1901, 2821, 949, 3024, 2014, 1423, 1163, 3192, 2660, 2617, 2978, 2056, 959, 1729, 944, 200, 1048, 876, 103, 253, 613, 2657, 617, 2007, 840, 788, 3133, 2449, 546, 1303, 2316, 2696, 935, 1027, 2974, 1464, 2039, 2060, 867, 1127, 961, 2805, 389, 2807, 2775, 14, 1961, 881, 3029, 1424, 579, 1778, 2488, 1592, 505, 1862, 1926, 1742, 1792, 3153, 1973, 3117, 725, 1733, 583, 1485, 645, 1634, 616, 1465, 1949, 1546, 1829, 491, 410, 2989, 37, 1089, 2931, 3059, 1543, 2780, 2086, 2976, 2846, 2015]. South-Central [2780]. South-East [410]. South-Mediterranean [3029]. South-Western [617, 2974, 389]. Southeast [2955, 2030, 849, 1223, 192, 2202, 650, 1995, 1620, 2053, 1780, 2191]. Southeastern [594, 2992, 854, 1476, 1112, 2454, 1103, 3151, 1386, 732, 1628, 2934, 1607, 2855, 426, 1844, 2316]. Southern [515, 2089, 1933, 1939, 1592, 839, 2318, 2756, 2747, 1152, 1026, 437, 684, 1409, 3054, 1682, 40, 1865, 2387, 2794, 2224, 773, 1486, 794, 1598, 1678, 185, 2601, 1322, 723, 815, 873, 1979, 945, 1249, 1405, 42, 919, 1181, 1458, 2186, 683, 1768, 1606, 714, 2737, 2958, 1324, 2725, 2495, 1487, 897, 1317, 526, 2141, 1682, 309, 2793, 357, 1627, 2672, 1443, 2744, 1129, 927, 642, 165, 126, 1605, 239, 2160, 1421, 1764, 1239, 1774, 764, 574, 3040, 609, 2926, 3150]. Southwest [742, 424, 2812, 1186, 2253, 2095, 949, 1088, 1343, 2197, 289, 56]. Southwestern [1818, 2222, 523, 1798, 1075]. Sovereignty [2997]. Soy [2388]. Soybean [1115]. sp [2214, 2583, 509, 990, 2889, 2782, 1639, 1972, 1863, 2409, 415, 1818, 948, 1297, 475, 1261, 2077, 2341, 880, 1040, 1242, 1304, 2801, 2772]. Sp-Pol [2772]. Space [1930, 1995, 1066, 1519, 1635, 726, 2569, 1230, 3036]. Space-Time [1230]. Spacer [2439]. Spain [2046, 2335, 2851, 484, 152, 143, 127]. Spanish [3104, 561]. Sparid [2540]. Sparidae [489]. Sparsely [1142]. Spartina [163, 2471, 2516]. Sparus [408, 1988, 2024, 2090, 26, 2452]. Spatial [424, 224, 2764, 2299, 1360, 1030, 601, 3025, 1332, 270, 1024, 945, 1817, 659, 494, 2317, 1308, 792, 1757, 755, 2095, 1290, 2795, 2469, 2513, 1568, 2148, 2063, 690, 3000, 929, 1234, 534, 71, 1723, 1296, 2659, 717, 932, 1519, 335, 2183, 478, 430, 1157, 3006, 285, 2545, 265, 2162, 2011, 915, 2709, 1724, 2146, 2234, 2022, 1593, 1506, 249, 514, 2276, 2649, 582, 1079, 686, 2125, 1743, 3018, 1980, 2083, 1621, 781, 1660, 1623, 1960, 832, 655, 2608, 1702, 1260, 1134, 210, 733]. Spatial-Related [1030]. Spatial-Temporal [915]. Spatially [2059, 2822]. Spatio [2096, 2248, 272, 2902, 1778, 1569, 143, 1529, 1078, 69, 834, 876, 1263, 1844, 2184, 1034]. Spatio-Temporal [2096, 2248, 272, 2902, 1778, 1569, 143, 1529, 1078, 69, 834, 876, 1263, 1844, 2184, 1034]. Spatiotemporal [987, 2682, 1843, 1423, 2409, 1528, 3097, 944, 2560, 1548, 3162, 126, 1776]. Spawning [2202, 3123, 1329, 45, 59, 1163, 296, 2032, 1988, 745, 2169, 662, 132, 1637, 3078, 2603, 1010]. Spawns [662]. Speak [1060]. spec [1341]. Special [1674]. Specialization [1978, 1600, 1551, 1661]. Specialize [1523]. Speciation [1293, 723, 3171, 571]. Specie [2027]. Species [2028, 2783, 907, 1367, 146, 1027, 1410, 2736, 1761, 2471, 585, 331, 2510, 369, 3082, 767, 1224, 3105, 1804, 14, 3062, 544, 1841, 589, 778, 1783, 2687, 545, 3017, 1596, 1703, 96, 949, 1001, 1100, 2863, 1237, 1695, 3134, 1915, 1036, 2999, 3077, 2385, 2971, 1163, 1279, 1655, 313, 1302, 1195, 296, 3009, 720, 734, 1615, 283, 1169, 1537, 571, 1196, 1887, 1289, 1893, 726, 803, 693, 2958, 736, 1324, 1261, 3172, 2056, 3107, 1341, 3072, 2076, 1327, 875, 2473, 1050, 2776, 481, 1626, 2180, 1585, 2432, 691, 1959, 2062, 543, 398, 915, 1429, 2158, 26, 1453, 2908, 1868, 758]. Species [140, 1508, 2892, 280, 1993, 576, 2744, 249, 5, 2424, 1041, 1962, 1384, 2275, 2526, 671, 2600, 2969, 2254, 1229, 752, 1428, 718, 266, 742, 3129, 1597, 405, 1837, 2505, 2976, 1382, 1422, 1907, 1304, 993, 1542, 2996, 1618, 2765, 2551, 3165]. Species-Specific [26, 1384, 1618]. Specific [1445, 2490, 404, 704, 472, 522, 2116, 139, 1600, 2772, 710, 837, 2792, 1525, 2232, 2223, 943, 1038, 356, 2302, 2767, 26, 1540, 841, 1384, 705, 1121, 2988, 1618]. Specification [2401]. Specificity [2798, 2888]. Specimen [550, 2533]. Specimen-Rich [550]. Specimens [1742]. Spectra [971, 1968, 412, 400]. Spectral [838, 595]. Spectrometer [147]. Spectrometry [1924, 1684]. Spectroscopy [2418, 2839, 2491]. Spectrum [314, 877, 837, 227, 1129]. Speed [2229, 1381, 1325]. Speeds [212]. Sperm [59, 2395, 1943, 1108, 3148, 2680]. Spermatogenesis [1179]. Spermatozoon [2840]. Sphyrna [1630, 1663]. Spicules [1313]. Spill [1681, 202, 1564, 169, 3139, 1785, 937]. Spilling [529]. Spillover [2608]. Spills [1232]. Spindle [1951]. Spine [1145, 1475]. Spinner [2555]. Spinners [762]. spinulosa [1835]. Spiny [2661, 3071]. Spinyhead [2446]. Spiral [375]. spirolide [2835]. Spit [856]. Spitsbergen [686]. Spleen [3146]. Split [2908]. Sponge [907, 1456, 497, 2576, 2457, 2475, 1217, 1184, 1226, 1302, 571, 1196, 2546, 1274, 1050, 956, 1266, 2115, 1280, 597, 1504, 680, 2057, 368, 2441, 1473, 1962, 1265, 2528, 1493, 1203]. Sponge-Associated [2457, 571]. Sponge-Dominated [1962]. Sponges [776, 1811, 2195, 1782, 162, 833, 1251, 1788, 1314, 2104, 1506, 576, 2500, 54, 405, 1456]. Spongia [51]. Sporadically [786]. Spore [1713]. Spot [1487]. Spotted [2927, 2958, 3160]. spp [515, 279, 2174, 807, 473, 1792, 735, 2081, 2829, 3090, 1178, 2168, 1660]. Sprat [424]. Sprattus [424]. Sprawl [549]. Spread [1167]. Spreading [273, 2303, 2516]. Spring [1789, 7, 1346, 868, 1483, 2813, 664, 285, 2167, 598, 74, 3083, 3207, 715, 2316, 1590, 1436]. Spring-Bloom [3083]. Spring-Neap [664]. Squamation [313]. squamosa [1971, 1733, 3159]. squamosina [1658]. Squat [1750, 1136, 56]. Squid [1329, 2253, 351, 2724, 744, 2032, 1699, 840]. Squids [3100, 1351]. Squirt [2776]. Sr [2538]. Src [112]. Sri [2283, 1878]. SSF [2188, 1333, 2790]. SSPs [982]. SSR [1542]. St [2808, 1729, 1866, 1966]. St. [485, 1048]. Stability [992, 190]. Stable [1268, 1750, 537, 2375, 710, 744, 1909, 2957, 2880, 2223, 943, 916, 450, 1191, 836, 1339, 1543, 490, 2780, 191, 154, 1989, 3040, 2996]. Stage [1750, 101, 307]. Stages [2663, 585, 776, 2544, 829, 558, 1733, 21, 1218]. Staghorn [938]. Staircases [2218, 2085]. Stakeholder [260, 166, 367, 931, 60, 248, 1605, 1874]. Stakeholders [1377]. Stalked [1100, 1537]. Standard [2736, 2308]. Standardisation [1583]. Standardization [1358]. Standardized [2398, 113, 2526]. Standing [1394, 452, 1509, 2422]. Star [1763]. Starfish [2775]. Stars [2829, 901]. Starvation [948]. State [308, 1469, 2421, 2012, 1591, 212, 1158, 2262, 1697, 835, 2263, 2911, 1352, 1777, 2335, 2613, 1825, 1890]. State-of-the-Art [2262, 835, 2335]. States [2673, 36, 3126, 845, 1808, 316, 2161, 511, 1697, 2703, 796, 677, 2263, 2668, 732, 2796, 2296, 1707, 113, 328, 2675]. Static [3016, 1691]. Station [1042, 582, 1162, 1566, 2931]. stationary [3078]. Stations [423]. Statistical [570, 1122, 1334, 566, 750]. Statistics [3077, 706, 1623]. Statolith [351]. Status [1347, 1032, 577, 1787, 1994, 1320, 2346, 1779, 124, 2111, 72, 962, 1195, 720, 734, 1169, 1889, 2827, 912, 2548, 1107, 3167, 344, 2380, 2360, 1780, 1524, 1211, 742, 3020, 1989, 2898]. STAVIRO [2398]. Staying [1343]. Steady [2582]. Steatosis [465]. Steinahóll [1357]. steinitzi [181]. Steller [1052]. Stenella [2018, 2555]. Stenotrophomonas [522]. Step [928, 1803, 2262]. Stepping [1080, 1356]. Steps [301]. Stereo [2924]. Stereo-Video [2924]. Sterile [1225]. Sterraster [1313]. Stewardship [2209]. Stickiness [1664]. Stillness [1797]. Stimulated [2016]. Stimulation [1176]. Stingray [1822]. stipulacea [77, 3019, 402]. Stirring [756]. Stochastic [3149]. Stochasticity [2027]. Stock [1445, 2005, 206, 3077, 2804, 720, 734, 1169, 2756, 226, 2120, 2661, 2065, 2058, 197, 266, 742, 2625, 840, 1475, 2898, 609, 2446, 2757, 1941]. Stock-Specific [1445]. Stocks [1053, 267, 2226, 482, 1755, 1394, 1523, 798, 416, 725, 452, 1107, 1059, 532, 2065, 2058, 1013, 2287, 2422, 831, 1303, 2696]. Stoichiometry [553, 55, 1263]. Stokes [2482]. Stolon [3044]. Stomiidae [204]. Stomiiformes [204]. Stones [1080, 1356]. Stony [265, 2509, 2673, 2322, 2614, 2317, 2666, 2687, 1718, 2319, 1043, 1216, 2180, 2616, 2749, 2311, 2107, 1702, 2139, 2953]. Stony-Coral-Tissue-Loss [265]. Stop [1243, 1450]. Stopover [1939]. Storage [3148, 2680]. Storm [2286, 411, 389, 108, 2763, 201, 3015, 2433, 2819, 365, 362, 964, 301, 2675, 1186, 1882, 1898, 1861, 1890]. Storm-Driven [2286]. Storms [2930, 253, 2441, 1473]. Story [2511, 2159]. Strain [2310, 1823, 1204, 2899, 1872, 2024, 3103, 2214]. Strains [397, 2573, 2002, 889]. Strait [2958, 562, 1293, 610, 2916, 659, 1841, 2753, 1118, 1407, 2342, 470, 1078, 137, 1899, 3054, 452, 530, 1457, 142, 3071, 2830]. Straits [310]. Stranded [786, 1476, 1768, 2263, 732, 1035]. Stranding [2026, 1408]. Strandings [1930, 2129, 1460]. Strange [489]. Stranger [1272]. Strategic [1379]. Strategies [1573, 2325, 3106, 2714, 3008, 1305, 1973, 2869, 2569, 478, 1038, 2432, 1370, 1429, 1214, 536, 1241, 1028]. Strategy [1978, 2043, 282, 1430, 1562, 1615, 1168, 969, 2517, 510, 1453, 1432, 2367, 902, 1691, 1674, 3092]. Stratification [1340, 879, 1603, 2935, 2711]. Stratified [2677]. Stream [470, 2528]. Streamer [2955]. Streamers [396]. Streamflow [2720]. Streamlined [2628]. Streamlining [902]. Streams [1183]. Strength [2010, 1830, 412, 766]. Streptococcus [2214]. Streptomyces [2623, 1863]. Stress [1892, 805, 2074, 1975, 32, 2043, 2651, 1039, 569, 3122, 1703, 2868, 3045, 2971, 465, 3037, 1770, 572, 1565, 1535, 830, 2761, 1216, 3195, 1847, 2962, 2038, 3021, 2671, 3055, 822, 1040, 3121, 2376, 1258, 95, 2053, 496, 2084, 1384, 955, 2357, 655, 1665, 2436, 2124, 777, 1461, 2644, 3193, 3158, 2861, 2258, 115]. Stress-Promoting [3195, 1040]. Stress-Responsive [32]. Stress-Scape [1975]. Stressed [580, 2306]. Stresses [3202, 1033, 3044, 761]. Stressful [1947, 1364]. Stressor [1755, 1431, 1147]. Stressors [1845, 1975, 615, 2460, 1460, 1271, 2011, 398, 2258]. striata [977]. striatus [2323]. strigosa [2148]. Strike [48, 170, 27]. Strikes [159, 2245, 1922]. Striking [58]. Striped [1102, 79, 2973, 3152, 1114, 1886]. Stripping [298]. Stroke [1206]. Strombidium [990, 475]. Strong [1014]. Stronger [730]. Strongly [666]. Structural [2195, 1809, 2022, 2257, 1417]. Structurally [198]. Structure [1268, 998, 3047, 1541, 1030, 1464, 1420, 155, 2173, 1332, 2775, 2108, 2575, 2955, 14, 1903, 556, 1388, 462, 3012, 2196, 1458, 746, 3081, 877, 1394, 1066, 2813, 929, 1953, 283, 2319, 717, 2164, 1078, 1528, 523, 1157, 1889, 996, 495, 119, 1796, 2943, 2828, 268, 1679, 2910, 2591, 3118, 836, 1857, 249, 595, 1275, 259, 493, 2818, 2125, 1497, 1925, 564, 1856, 311, 3076, 1751, 2444, 3207, 2391, 1369, 1685, 1136, 56, 574, 1422, 618, 1124, 3048, 476, 2446, 1553, 1436, 2438]. Structure-From-Motion [1903, 2591]. Structured [579, 2524, 2625]. Structures [789, 550, 121, 480, 3016, 2120, 393, 2885, 1307, 1946, 332, 2472]. Structuring [2999]. Student [1173]. Students [2247]. Studied [1463, 1732]. Studies [3144, 2026, 1559, 1437, 2899, 1843, 3077, 47, 1586, 1411, 2236, 3169, 1339, 3059, 1180]. Study [1892, 971, 1616, 1472, 3039, 1293, 2930, 2711, 1986, 842, 1209, 2202, 1360, 1052, 1161, 2797, 2046, 278, 2567, 129, 1106, 228, 2383, 949, 814, 928, 2789, 744, 418, 505, 673, 30, 2066, 2006, 1742, 130, 1904, 2537, 1673, 1004, 1305, 235, 1652, 470, 1528, 2487, 1187, 1522, 1374, 2890, 3097, 2433, 2956, 1451, 448, 1402, 1866, 1840, 2062, 2193, 442, 2345, 740, 2245, 1191, 1452, 249, 927, 125, 1509, 2336, 561, 1306, 3, 1069, 2573, 27, 1609, 1611, 2168, 2198, 3036, 3078, 2434, 1917, 843, 614, 1404, 715, 2881, 43, 1205, 2620, 338, 2886, 670]. Studying [789]. Stunted [699]. Stylocheilus [3075]. Stylonematales [111]. Stylophora [968, 117]. Sub [854, 588, 373, 234, 2575, 792, 1779, 2109, 1489, 1398, 214, 2796, 593, 2117, 1457, 3149, 832, 3040, 3164]. Sub- [214]. Sub-Adult [792]. Sub-Antarctic [1398, 3149]. Sub-arctic [234, 373, 2575, 1779, 593, 2117]. Sub-Basins [588]. Sub-Bottom [854]. Sub-Delta [3040]. Sub-Polar [3164]. Sub-Regional [1489]. Sub-Surface [1457, 832, 2109]. Sub-Tropical [2796]. Subantarctic [1505, 3063, 1592]. Subarctic [1415, 3063, 162, 2643, 2020, 2163, 1935, 2838, 2700]. Subcellular [1112]. subcrenata [2861]. subcurvata [580]. Subeuphotic [1973]. Subfamily [3212, 1304]. Subject [1589, 1546]. Submarine [2286, 1009, 273, 2770, 2374, 710, 94, 283, 693, 2222, 1477, 3059, 74, 2438]. Submerged [632, 3016, 1887]. Submergence [2393]. Submerging [2571]. Submersed [2769]. Submersible [755]. Submesoscale [2277, 769, 1423, 929, 1625, 1530]. Submicron [2713]. Subpolar [42, 3177]. Subregions [2890]. Subsequent [815]. Subsidies [939, 482, 858]. Subsidy [2276]. Subsistence [2137, 2136]. subsp [2899]. Subspecies [3153, 2191]. Substance [396]. Substantial [1310, 481, 1457, 80]. Substrate [1824, 1652, 2405, 1608]. Substrate-Independent [1652]. Substrate-Water [2405]. Substrates [2048, 986, 513, 671, 3129]. Substratum [48]. Subsurface [773, 11, 949, 587, 802, 788]. Subterranean [1643, 2045, 526, 2521, 1467]. Subtidal [621, 1851, 1965, 1756, 1672]. Subtle [967, 1793, 2102]. Subtropical [1268, 779, 1771, 1397, 2955, 590, 2014, 130, 1904, 2288, 2020, 1634, 2934, 2824, 409, 2603, 2860, 1082]. Success [1228, 2013, 1343, 1412, 2295, 194, 174]. Successes [2985]. Successful [857, 1305, 2558, 1003, 2159]. Succession [1828, 1088, 1918, 564]. Such [149]. Sucker [2514]. Sugar [3091, 796]. Suggest [196, 1957, 667, 409]. Suggests [1853, 2544, 1658, 479, 1038, 2175]. Suitability [1712, 139, 2884, 38, 2360, 2023]. Suitable [672, 2590, 315]. Suite [800, 2062]. Sulfate [1871]. Sulfide [115]. Sulfur [382, 171]. Sultanate [768]. Summarized [1637]. Summer [2688, 1716, 2865, 2942, 659, 2462, 714, 2755, 1299, 2617, 285, 1046, 37, 1736, 182, 3150]. Summering [1858]. Summers [686]. Summertime [2342, 3196, 452, 354]. Sunda [137]. Sundarbans [2196, 3191, 324]. Sunscreens [2097]. Super [1850, 1522, 2433]. superba [1452, 2177, 2625]. Supercritical [1894]. Superfamilies [3079]. Supergiant [2896]. Supervised [369, 701]. Supplement [1115, 2033]. Supplementation [2905, 3027, 2870, 263]. Supplemented [3111]. Supplementing [2595]. Supplements [939]. Supplied [1931]. Supply [2950, 1826, 2270, 2581, 2367, 2087, 2891, 3036]. Support [225, 1846, 107, 1849, 2699, 1757, 853, 1430, 1441, 492, 1562, 1697, 1432, 1133, 165, 2966, 1180, 1611, 1839, 439, 2207, 1870]. Supported [1600]. Supporting [2659, 1494, 1593, 927]. Supports [1223]. Suppress [3154]. Surf [1164]. Surface [2535, 3047, 2698, 2577, 2482, 588, 1065, 1313, 2975, 723, 2944, 1771, 2461, 2041, 1505, 541, 1838, 2382, 1097, 2513, 1830, 33, 1202, 2384, 1438, 925, 1029, 73, 784, 321, 137, 1046, 2101, 2215, 452, 2059, 1151, 2828, 2766, 2448, 2609, 1035, 1912, 594, 694, 2219, 126, 785, 1457, 1501, 1764, 2727, 832, 2029, 2472, 1665, 574, 43, 2289, 1436, 1277, 1325, 2109]. Surface-Canopy [2975]. Surface-Layer [2727]. Surfers [2668]. Surfing [2064]. Surge [2433, 964, 2675, 338]. Surges [201, 365, 362, 1439]. Surrogate [864]. Surrogate-Reared [864]. Surrounding [364, 2944, 1981, 3164, 37, 2913]. Suruga [2727]. Surveillance [2776]. Survey [542, 1804, 1680, 1236, 1913, 784, 921, 2081, 89, 1010]. Surveying [374, 1365]. Surveys [1903, 1579, 995, 1772, 2572, 1960]. Survival [1940, 2220, 1176, 2393, 713, 1068, 1251, 3061, 1168, 1490, 2300, 2054, 2750, 1278, 2712, 3092, 3053]. Survive [2847]. Surviving [1443]. Survivors [2754]. Survivorship [898]. Susceptibility [2386, 1500]. Susceptible [2687, 616, 2169]. Suspended [2596, 867, 1730, 2415, 1726, 2559, 1493]. Suspends [2328]. Sustainability [2113, 258, 278, 1901, 30, 2659, 127, 2075, 1005, 2647, 459]. Sustainable [3110, 2246, 639, 2244, 1096, 60, 2659, 1519, 875, 1859, 1578, 1222, 187, 1538, 2618, 1387, 1687, 255, 678, 64]. Sustained [2490, 2742, 368, 1457]. Sustaining [1578]. Sustenance [2701]. Svalbard [2044, 591, 2512, 874]. SW [2510, 143, 1016, 1161]. Swath [87]. Sweden [2990]. Swelling [2840]. Swells [648]. Swim [1619, 762]. Swimming [52, 2465, 1285, 1963, 924, 1286, 747, 3054, 2750, 2468, 136, 3193]. Switch [1132]. Switching [801]. Swordfish [782, 1691]. SWOT [2294, 2428].

Symbiodiniaceae [3124, 209, 2115, 1993, 2628, 2868, 1272, 1733]. Symbiodiniaceae-Associated [209]. Symbiodinium [3159]. Symbiont [955, 2798]. Symbionts [1462, 3034, 2104, 567]. Symbiosis [563, 3201, 2587, 2157, 98]. Symbiotic [2583, 576]. Sympagic [554]. Sympatric [1841, 379, 2375, 2477, 726, 1090, 2191]. syn [437]. Synanthropic [881]. Synaptic [1255]. Synaptogenesis [1255]. Syndinean [2716]. Syndrome [2688, 1237, 2148, 392]. Synechococcus [1737, 2406]. Synergies [646, 25]. Synergistic [1570, 348, 2053]. Syngnathid [3118]. Synoptic [272]. Synthase [2314, 3035]. Synthesis [1446, 1932, 2202, 1823, 1746, 912, 2399, 2236, 1193, 1827, 2065, 937, 1558, 3030, 3194, 2017, 2419, 263]. synthetase [3187]. Synthetic [1008, 1354, 1729]. System [1754, 3119, 800, 1071, 1185, 485, 2428, 3191, 1753, 155, 1978, 1633, 1510, 2645, 189, 2730, 1006, 554, 1838, 1160, 569, 247, 722, 1929, 2378, 579, 2899, 492, 2196, 1236, 216, 1192, 2463, 527, 130, 1904, 1673, 1589, 1730, 2140, 1594, 932, 2642, 1139, 2735, 69, 874, 2495, 944, 897, 691, 1897, 2610, 2885, 1171, 2141, 2252, 2398, 1408, 2793, 2852, 477, 911, 613, 1844, 1500, 781, 1390, 3083, 2674, 2938, 2139, 1917, 737, 2408, 1205, 3053, 361, 2189, 2001, 928, 782, 13, 1274, 2959, 1644, 1098]. System-Based [579]. Systematic [601, 1994, 1228, 1492, 3197, 108, 2327, 3070, 738, 1905, 2572, 1318]. Systematics [284, 1138, 1304]. Systemic [438]. Systems [1640, 1095, 2991, 3025, 2427, 2684, 3136, 2359, 2012, 3041, 1928, 2274, 2312, 434, 3156, 3081, 1695, 499, 41, 1279, 441, 2119, 1896, 2819, 2781, 2087, 1905, 2282, 540, 1230, 1513, 2600, 985, 3033, 301, 2547].

T. [1733]. Tachai [1620]. Tachypleus [2967, 1076]. Tackle [2869]. Tag [839, 670, 302]. Tagged [2235, 1278]. Tagging [827, 1571]. Tagoro [2374, 94, 2438]. Tags [3038]. Tagus [2255]. Tahiti [74]. Tailings [751]. Taiwan [2958, 2830, 2944, 3104, 2682, 1094, 1998, 1272, 720, 734, 470, 2598, 645, 3054, 142]. Taiwanese [1460]. Take [1741, 1901, 2104]. Taken [404]. Taking [136, 2546]. Tale [533]. Tales [2228]. Tanaidacea [3143, 1818]. Tank [1709]. Tanks [662]. Tanzania [2224, 525]. Tap [2654]. Tara} [16, 16]. Target [1367, 2010, 1379, 1227]. Targeted [2776, 934, 3194, 1641]. Targeting [1684]. Targets [1741, 1246]. Tasman [129, 1537]. taty [3077]. Taxa [2179, 234, 257, 1184, 2718, 1391, 2042, 1368, 1162]. Taxon [704, 2116]. Taxon-Specific [704, 2116]. Taxonomic [1583, 1195, 2747, 2120, 681, 2976]. Taxonomically [2337]. Taxonomy [1841, 1100, 962, 1195, 1324, 1261, 1341, 1679, 249, 1138, 2740, 1837]. TCAP [1165]. Teahitia [176]. Tech [1413]. Technical [2797, 965, 3206]. Technique [710, 780]. Techniques [1102, 369, 2012, 391, 881, 49, 1588, 774, 1814]. Technological [1057, 653, 1591]. Technologies [2155, 1896, 2210, 679, 1417]. Technology [3210, 2593, 144, 2262, 2485, 3182, 1576, 504, 2392]. Tectonically [812]. Tedania [1508]. Tedaniopsis [1508]. Teflubenzuron [2450]. Tegillarca [1751]. Tejo [2592]. Teleconnection [2837]. Telemetry [224, 1695, 3070, 747, 2741, 1698, 698]. Teleost [1796, 65]. Teleostei [489, 2908]. Teleosts [811, 220]. Tell [2511]. Telomere [1840]. Temperate [2800, 1081, 3003, 552, 2590, 555, 602, 2957, 2747, 1157, 1959, 2130, 55, 15, 1854, 239, 1306, 1756, 1495]. Temperature [427, 2179, 2350, 1262, 2577, 2393, 1977, 2371, 2099, 3123, 3100, 2564, 1781, 669, 2401, 2840, 2567, 134, 501, 1830, 2256, 2134, 587, 1963, 342, 1031, 2204, 997, 3037, 1438, 714, 925, 1435, 932, 1029, 1535, 2909, 2627, 50, 1552, 422, 2984, 1847, 1046, 35, 2556, 2101, 1967, 2038, 2252, 2090, 603, 2921, 1067, 1912, 865, 694, 1275, 1170, 126, 86, 1384, 1764, 2381, 884, 802, 2822, 3158, 2503]. Temperature-Driven [2556]. Temperature-Induced [134, 997]. Temperature-Salinity [802]. Temperatures [1639, 974, 51, 2681, 655, 2603]. Temporal [424, 2096, 2955, 1817, 2317, 755, 2248, 272, 2902, 1778, 2469, 1569, 143, 201, 2129, 1529, 1296, 1078, 3145, 478, 423, 69, 834, 2162, 915, 2508, 1371, 876, 626, 1821, 158, 642, 1780, 1263, 1621, 1844, 3088, 781, 2184, 978, 832, 1034, 1702, 1404, 3048, 366]. Temporally [2642]. Temporarily [1493]. Temporary [893]. Ten [2566, 1538]. Teneurin [1165]. Tenualosa [3014, 656]. Terebellida [2166]. Terengganu [28]. Terephthalate [3102]. Term [1042, 1470, 2605, 1845, 1817, 457, 1006, 1225, 1584, 3106, 1068, 1423, 2717, 24, 161, 850, 696, 1919, 2032, 1400, 2844, 1551, 1343, 2963, 2495, 1836, 2025, 8, 84, 2250, 2101, 2215, 1540, 876, 2496, 1607, 368, 1278, 2445, 2156, 613, 190, 1810, 93, 1197, 768, 764, 20, 2898, 2547, 3200, 3193, 1884]. Terminal [1165, 892]. Terms [536]. Terns [2906]. Terra [1116]. Terrain [1965]. Terrestrial [710, 1887, 874, 183, 2752]. Terrigenous [900]. Territorial [181]. Territories [2036]. Test [805, 223, 2455, 318, 1648]. Tested [404]. Testing [3082, 2390, 1096, 1765, 1889, 1585, 2518, 255, 1623]. Testis [2003, 3127]. Tethys [1189, 2815]. Tetractinellida [1196]. Tetraether [860]. Tetrameristaceae [2174]. Tetraselmis [180]. TEX86 [2677]. Texas [2984]. Thailand [2072, 1855, 1620, 392, 233]. Thalassiosira [3172, 1725, 715]. Thaliacea [1628]. Thalli [496]. Thaw [1943]. Thecostraca [1537]. Their [2214, 1739, 656, 3060, 286, 250, 102, 2770, 1250, 2847, 3124, 3100, 1804, 2760, 1252, 1561, 2382, 162, 1801, 1746, 3069, 674, 1523, 2417, 3070, 1819, 929, 3056, 237, 2403, 2806, 512, 209, 1460, 2127, 47, 2904, 2056, 896, 1727, 1219, 2829, 2267, 3186, 2672, 567, 2729, 3059, 692, 2275, 671, 178, 2175, 2168, 3087, 2759, 1173, 1805, 2408, 3166, 1062, 3064, 2823]. Them [649]. Theoretical [404, 1879]. Theory [2182, 3142]. There [2057]. Thermal [2577, 1690, 1137, 2825, 813, 3062, 501, 3084, 2868, 3045, 1855, 558, 1216, 1847, 2962, 2038, 1056, 1993, 1443, 822, 2376, 2053, 2084, 2798, 655, 2935, 2861, 2540, 3165]. Thermally [2037, 347, 804]. Thermodynamic [3004]. Thermohaline [2233, 2955, 2218, 2085, 2201]. Thermooxidative [1063]. Thermotolerance [3034, 1023]. Theta [500]. Theta/S [500]. Thick [222, 2843]. Thickness [1830]. Things [1272]. Think [1385]. Thinking [2867]. Thomas [2322]. Thoondu [856]. Thoracica [1100]. Thorium [309]. Thornback [1127]. Thorns [2775]. Those [2145]. Thousands [1742]. Threat [281, 795, 1584, 1722, 245]. Threatened [2174, 1464, 2130]. Threatens [1190, 2969, 3019]. Threats [1185, 310, 2429, 3186, 168, 1558, 692]. Three [1102, 331, 767, 3124, 1243, 544, 3003, 96, 1001, 1617, 928, 1995, 1291, 1533, 2887, 803, 1654, 1261, 2890, 2880, 1341, 2036, 423, 1485, 933, 2661, 1796, 731, 3054, 1429, 1854, 1668, 1357, 1239, 1837, 2216, 2996]. Three-Dimensional [1291, 1654, 933, 1796, 1854]. Threshold [2463, 3138]. Thresholds [2883, 1446, 1932, 257, 2718, 198, 2765]. Throughout [1706, 2185, 2663, 776, 2889, 1359, 1533, 796, 2050]. Throughput [2379, 2272, 1084]. Thunnus [1101]. Thy [692]. Thymallus [2676]. Tidal [2255, 2986, 272, 505, 1471, 2363, 235, 2887, 470, 664, 3007, 2747, 726, 2172, 2495, 2817, 412, 1587, 655, 3040, 1205, 1439, 3096, 2773]. Tide [1624, 2667, 726, 816, 2135, 1091, 2675, 338, 1491, 2635, 2855, 993, 1657]. Tide-Surge [338]. Tides [1671, 3117, 3196, 3151, 365, 598, 125, 2929, 640]. Tidewater [1740]. Ties [2399]. Tiger [1073, 274]. Tight [300]. Tigriopus [1907]. Tilapia [2586, 465]. Time [676, 2509, 1986, 2802, 1410, 203, 575, 2782, 2041, 1237, 3095, 1995, 1130, 1066, 193, 1666, 1635, 2904, 1552, 2101, 1859, 2560, 1548, 2892, 637, 1230, 1089, 2559, 642, 1691, 2524, 743, 2086]. Time-Area [1691]. Time-Course [2782]. Time-Dynamic [642, 743]. Time-Lapse [1986]. Time-Varying [1410]. Timeframe [2204]. Times [51, 745, 2647, 2177]. Timeseries [1781]. Timing [7, 1346, 2307, 1540]. Timor [2228]. Timor-Leste [2228]. Tintinnid [2716]. Tip [1592]. Tisochrysis [991]. Tissue [2509, 2673, 2322, 2614, 2317, 2666, 967, 2687, 1718, 2982, 1031, 2319, 1043, 520, 375, 1314, 1216, 2180, 265, 2653, 2616, 2749, 2311, 2107, 1702, 2139, 2953]. Tissues [3189]. TNFs [1256]. Tocopherols [1040]. Todarodes [2032]. TOF [1924, 1907]. Together [730, 2941]. Tokyo [1825]. Tolerance [501, 196, 2038, 1993, 2573]. Tolerant [1708, 3034, 3121]. Tomography [1876, 883, 2252]. Tomorrow [1767]. Tongue [65]. Tonkin [3112]. tonsa [1940, 2781]. Tool [963, 2511, 2699, 1491, 3070, 75, 1602, 2808, 256, 791, 747, 1999, 988, 2354, 2412, 2524, 2376, 1144, 3068]. Tool-Kit [1999]. Toolbox [1358]. Tools [1032, 579, 2395, 1267, 1852, 2142, 384, 2411, 927, 1133, 711, 840]. Toothed [1559, 2391]. Toothfish [2108]. Top [2096, 1766, 83, 1528, 1287]. Top-Down [83, 1287]. Topics [1348]. Topographic [2237, 1075, 1479]. Topography [1146, 613]. Topography-Flow [613]. Topology [907]. Tortoiseshell [1200]. Total [44, 1730, 321, 236, 1888, 1003]. Tough [489]. Tourism [270, 1617, 363, 1621, 762]. Tourism-Naive [1621]. Tourist [1647, 1245, 2102]. Touristic [105, 1406]. Tours [2468]. Towed [337, 335]. Toxic [1281, 2111, 2082, 2298, 1805]. Toxicity [862, 1011, 2835, 3184, 2124]. Toxicology [472]. Toxicosis [1409]. Toxigenic [2684]. Toxigenicity [671]. Toxin [1940, 2835]. Toxin-Producing [1940]. Toxoplasma [438]. Toxoplasmosis [438]. Trace [1761, 462, 1054, 1299, 915, 612, 2198, 1298]. Traceability [2047, 2988]. Tracer [2700, 696, 1989]. Tracers [2483, 1155, 33, 2562, 1957]. Trachelocercid [3072, 1327]. Trachinotus [1345, 2134, 1297, 3130, 3055]. Trachymedusae [63]. Tracing [508, 278, 1254, 2066, 2998]. Track [1180]. Tracking [3011, 2379, 2126, 710, 2455, 661, 75, 467, 1734, 407]. Tracks [578, 2004]. Tract [1362, 414, 436, 1535, 1982]. Trade [1978, 1624, 2599, 28, 1200, 1762, 2740, 1794, 2907]. Trade-Offs [1978, 2599, 1762, 2740, 2907]. Traditional [3125, 1195, 1679, 1156, 2558, 1784]. Traffic [224, 2243, 189, 2242, 1408, 58, 1153, 2284]. Trafficking [1857]. Tragedy [159]. Training [2034, 1105, 1848]. Trait [1978, 3003, 1759, 1419, 3002]. Trait-Based [1759]. Traits [1378, 2689, 2272, 2670, 1600, 1703, 2121, 1812, 1036, 1272, 2599, 1809, 2655, 1651, 977, 2496, 249, 1467, 148]. Traits-Based [148]. Trajectories [339, 979, 2842]. Trajectory [1564, 2518]. Trang [1265]. Trans [2784, 264, 129, 1086, 1957]. Trans-Generation [1086]. Trans-Generational [2784]. Trans-National [264]. Trans-Oceanic [1957]. Trans-Tasman [129]. Transboundary [1974, 908, 1180]. Transcribed [2439]. Transcript [3159]. Transcriptional [3202, 1017, 3189, 2806, 2024, 1725, 2861]. Transcriptome [2802, 2782, 2003, 2514, 2841, 2361, 3037, 1264, 1073, 2961, 3146, 26, 1567, 274, 3055, 358, 3148, 2680, 2758, 2579, 3158, 2489, 1542, 2752]. Transcriptomes [2917]. Transcriptomic [1923, 2365, 3045, 2982, 700, 1770, 51, 1132, 520, 1168, 3021, 307, 2053, 3160, 3166]. Transcriptomics [2971, 160, 2831, 2907, 2762]. Transcripts [3056, 1567]. Transdisciplinary [1846, 1649, 1173]. Transduced [2271]. Transect [1979, 3063, 921, 1772]. Transects [374]. Transfer [1456, 144, 2642, 1720]. Transform [2418, 2104]. Transformation [1496, 453, 2868, 186, 1140, 382, 171]. Transformations [1745, 2647]. Transformative [925]. Transformed [2557]. Transgenerational [2825, 573]. Transgression [3155]. Transient [2900, 696, 1223]. Transit [2001]. Transition [30, 1212, 2408]. Transitional [2012]. Translating [2790]. Translation [3055]. Translocation [922, 2035, 1971, 2943, 187]. Transmission [1255, 285, 2311]. Transnational [1638, 1556]. Transparency [349, 953]. Transport [872, 1643, 569, 810, 90, 2646, 3159, 1051, 2267, 2706, 838, 3030, 1587, 2750, 114, 1732, 2212, 2935, 1439]. Transporter [1971, 3159]. Transporters [222]. Transports [2582, 984]. Transposon [2791, 110]. Transregional [2136]. Trap [2847, 2515, 1061]. Trapped [1075]. Trapping [2793, 3043]. Traps [2417]. Trash [1833]. Travelling [535]. Trawl [1367, 962, 1266, 1696]. Trawlers [673]. Trawling [2328, 1083, 1190, 559, 1524, 148]. Trawls [2177]. Treasure [592, 1833]. Treated [3157]. Treatment [427, 570, 793, 2982, 3052, 953, 1339]. Treatments [2250, 2107]. Treaty [144, 2047]. Tree [49, 3155]. Trees [699]. Trench [2178, 1553]. Trenches [1173]. Trend [3106, 3101, 768]. Trends [2662, 1992, 610, 486, 1359, 1831, 2768, 2599, 1025, 1108, 428, 1662, 423, 2167, 2552, 1273, 1900, 876, 3113, 2937, 2156, 93, 3022]. Tri [1475]. Tri-Spine [1475]. Triacylglycerol [207]. Triangle [688, 995]. triangula [2716]. Tribute [217]. Trichiuridae [1118]. Trichodesmium [149]. Trichome [2281]. tricornutum [862]. Tridacna [1971, 996, 1658, 1733, 3159, 1178, 2002]. Tridacninae [567]. tridentatus [2967, 1076]. Trieste [1580]. Trigger [1161]. Triggered [861]. Triggering [2786]. Triggers [3154]. trigloides [523]. Trimodal [614]. Trindade [1818]. Trioecy [450]. Tripneustes [2339]. Triportheidae [110]. Triportheus [110]. Tristan [1729]. Trithigmostoma [1629]. Trithorax [291]. Tritia [2048]. Triton} [211]. trituberculatus [2933, 1286, 2016, 2750, 136, 3166, 3193]. Trochus [325]. Trophic [1268, 1991, 156, 591, 258, 976, 2375, 744, 1359, 1568, 797, 1953, 1299, 2223, 943, 481, 2302, 3118, 836, 1129, 2780, 1524, 2083, 190, 2505, 1989]. Tropical [1268, 998, 1630, 1311, 632, 867, 2807, 2786, 1945, 2244, 2467, 3084, 1843, 812, 2999, 2018, 2283, 2228, 798, 460, 2288, 2743, 803, 1147, 2685, 348, 2627, 2735, 413, 2297, 1899, 1847, 750, 2796, 2367, 2766, 2245, 2934, 1089, 1780, 617, 2281, 54, 2422, 402, 3085, 1277, 214, 861]. Trough [2737]. Trout [3027, 2777]. True [918, 289]. truncatus [225, 2974, 605, 1490, 2635, 2984, 1688, 2854, 1621]. Truthing [690]. Tsunami [108]. Tsunami-Related [108]. Tsushima [742]. Tubes [1981]. Tubeworm [1976]. Tubular [1971]. Tuna [998, 2985, 2845, 420, 2717, 241, 1528, 1101, 1145, 543, 721, 2609, 1738, 2086]. Tuna-Dolphin [2985]. Tunas [359]. Tunicata [1628]. Tunicate [1863]. Tuning [1230]. Tunisian [1155, 1894]. Turbid [281, 2771, 1002, 2508]. Turbidity [1081, 797, 2957, 3043]. Turbine [481]. Turbines [529]. Turbot [1115, 1179, 2423, 300]. Turbulence [575, 470, 2443, 1063, 2308]. Turbulent [287, 3117, 252]. Turkey [1229, 752]. Turkish [1229, 752]. Turks [792]. Turn [726]. Turning [2654, 280]. Turnover [3095, 2045, 1959, 1603, 2412]. Tursiops [225, 2974, 1464, 605, 1550, 1490, 2635, 2984, 876, 1688, 2854, 1621]. Turtle [153, 2070, 757, 2026, 3132, 403, 36, 2186, 3070, 75, 1728, 28, 1887, 1289, 808, 2518, 355, 3162, 775, 1734, 2924, 323]. Turtles [3011, 1104, 1846, 1186, 1690, 1476, 2847, 544, 945, 792, 2515, 795, 661, 760, 432, 2009, 310, 909, 749, 82, 2040, 199]. Tuvalu [861]. Twelve [960]. Twenty [2922, 2991, 2261]. Twenty-First [2991, 2261]. Twenty-One [2922]. Two [827, 663, 1643, 2859, 3100, 1771, 2614, 2687, 1559, 855, 1100, 1044, 1843, 3069, 1523, 2333, 1272, 130, 1904, 1302, 2949, 1289, 726, 2127, 1324, 330, 636, 1733, 3072, 1327, 2154, 2241, 1273, 2169, 2908, 532, 2340, 397, 1089, 2966, 2573, 964, 1960, 2002, 802, 889, 3176, 2811, 935, 1180]. Two-Year [1089]. TWS119 [2361]. TX [316]. Type [3124, 2268, 2164, 1608, 529]. Types [2578, 3107, 2748, 1495]. Typhoon [2660, 1522, 2433, 1850, 2926, 338]. Typhoon-Induced [2660]. Typhoons [2944, 253]. Typical [915, 3200]. Typically [289]. Typological [2930]. Typology [2778]. typus [2324, 2900, 1221, 1048]. Tyrrhenian [594, 3123, 2218, 118].

U [2538]. U.S [2322, 3151]. U.S. [1360, 1559, 212, 650, 503]. UAE1 [2623]. UAS [540]. UAV [1961, 2738]. UAV-Mounted [2738]. Ubiquitous [628, 1259, 665]. Ubiquity [1867]. UK [466, 179]. Ulcers [515]. Ulloa [1846]. Ultra [658]. Ultra-Oligotrophic [658]. Ultrastructural [1259].

Ultrastructure [2889, 974]. Ultraviolet [408, 2082]. Ulva [156, 1723, 816, 2135]. Ulvophyceae [816, 2135]. Unaffected [1188]. Unbaited [2398]. Unbiasing [1382, 2473]. Uncertainties [2167, 2694, 1777]. Uncertainty [2948, 1560, 2235, 2401, 3199, 1928, 1122, 2735, 2332]. Unconstrained [1676]. Unconventional [1029]. Uncover [2917]. Uncovering [59, 1634]. Under-Ice [1282]. Under-Reporting [3144]. Undercurrent [2909]. Underexplored [1599]. Undergoing [3108, 2993]. Underlying [3060, 62, 1693, 1973, 3121]. Underpin [2500]. Underpins [512]. Underreporting [2245]. Underscores [775]. Undersized [1487]. Understand [3025, 2026, 1379, 1167, 1755, 3134, 891, 1869, 1383, 2157, 2217]. Understanding [2783, 2802, 250, 279, 708, 2413, 1927, 451, 552, 1803, 2717, 2031, 677, 2142, 2947, 1898, 399, 1604, 1478, 2103, 2690, 2500, 2741, 447, 1623, 2980]. Understudied [1599, 2940]. Underutilized [1294]. Underwater [2736, 2492, 2243, 2005, 605, 653, 2242, 445, 2186, 627, 2161, 2771, 9, 2105, 2954, 2398, 1417, 1153, 81, 1457, 332, 1581, 1772, 614, 2419]. Underway [16, 2572]. Undescribed [1050, 2720]. Undulating [337]. UNESCO [827]. Unexpected [2768]. Unexplored [6, 92]. Unfolding [1719]. unguis [880]. Unicellular [238]. Unicorns [1996]. Unifying [2279]. Unique [62, 2693, 456, 1818, 1326, 193, 698, 1040, 1800]. Unit [2215, 2602]. United [1808, 643, 316, 1831, 2703, 796, 677, 2263, 2668, 732, 2796, 328, 2675, 2673, 2670, 2161, 511, 3126, 1697, 2171, 2296, 1707, 845, 113]. Units [936, 335]. Unity [3042]. Universal [1525]. Unknown [6, 92, 2829, 2192, 1292]. Unmanned [43]. Unobserved [1083]. Unoccupied [1804, 540]. Unpaid [2407]. Unpaired [2419]. Unprecedented [667]. Unravel [1066, 1968]. Unraveling [3060, 1175, 214, 2580]. Unravels [2158]. Unresolved [2076]. Unsettled [1195]. Unstructured [2487, 1902]. Unstructured-Grid [2487]. Unsynchronized [3038]. Untangling [2228]. Untargeted [352]. Unusually [1214]. Unvegetated [2172]. Unveiling [2609, 1333, 2188]. Upcoming [2294]. Update [2373]. Upflow [3053]. Upgrading [2444]. Upon [104, 2131, 2945]. Upper [976, 2092, 1850, 253, 1263, 3133, 3093, 1436, 1192, 2319, 1855, 2984]. Upper-Layer [3133]. Upper-Ocean [1263]. Upregulated [565]. Upscaling [894]. Uptake [383, 1971, 528, 2755, 2617, 2221, 894, 180, 1239]. Upwelled [2746]. Upwelling [1640, 2535, 1276, 2882, 2865, 2645, 2046, 922, 247, 1061, 928, 1914, 1141, 2617, 944, 1644, 1171, 2087, 2793, 357, 822, 753, 1479, 635, 409, 1935, 3176, 2811, 2480]. Urabá [1820]. Ural [2298]. Uramba [1387]. Urban [2174, 1555, 2352, 1427, 2442, 2881]. Urbanized [926, 2458]. Urbanizing [1820]. Urbanness [2173]. Urchin [2859, 2853, 621, 1149, 1058, 2339, 2708, 269, 1897, 695]. Urchins [1921, 1094, 307]. Urgent [1707]. Uronema [1261]. Uronychia [1259]. Usage [1243]. Use [1865, 1464, 852, 2182, 2760, 2026, 1070, 1757, 755, 2244, 228, 1290, 1663, 1562, 1221, 1852, 1008, 1519, 726, 2487, 966, 909, 1105, 449, 1796, 1273, 1540, 344, 1380, 122, 2745, 2368, 1477, 1610, 170, 2524, 3018, 1980, 3033, 2529, 164, 1354]. Used [1573, 108, 161, 1602, 2097, 68, 2024, 1239, 2168, 417]. User [2202, 2872, 186, 2584]. User-Perception [186]. Using [2334, 500, 1754, 998, 1123, 3046, 2930, 2511, 2787, 2010, 2061, 3144, 1093, 2577, 324, 2492, 1560, 2231, 704, 2112, 2400, 716, 1363, 369, 2182, 2975, 1270, 2805, 1876, 351, 2108, 2459, 522, 1804, 2048, 610, 2041, 316, 2094, 39, 261, 531, 1757, 646, 1507, 1083, 2327, 579, 1407, 2418, 744, 212, 2463, 333, 587, 2265, 2804, 690, 313, 3126, 2290, 1195, 1792, 720, 734, 914, 2873, 1723, 1594, 1693, 2655, 2839, 1582, 1525, 1169, 1666, 2319, 925, 883, 634, 1894, 1289, 428, 101, 735, 1632, 416, 725, 475, 2842, 706, 2604, 562, 2856, 2827, 796, 1552, 2473, 921, 1334, 943]. Using [566, 2081, 934, 2545, 2610, 173, 2215, 205, 916, 2062, 543, 268, 3159, 1999, 2141, 2252, 2308, 3139, 2709, 309, 2508, 2591, 274, 2305, 1504, 2914, 1191, 860, 3162, 1399, 612, 2282, 532, 227, 2257, 147, 506, 2738, 729, 995, 1477, 2491, 1637, 2734, 711, 1509, 642, 2931, 2404, 1576, 911, 2147, 401, 2669, 1144, 1844, 1227, 826, 718, 266, 742, 1732, 2287, 840, 396, 2822, 1382, 2686, 1907, 1709, 2547, 43, 831, 844, 609, 1303, 1298, 1941]. Usual [830]. Usually [1437]. USV [1365]. USVs [1813]. Utilities [1873, 1358]. Utility [842, 1491]. Utilization [383, 2449]. Utilizing [2715]. UV [3102, 2097, 214]. UV-Aged [3102]. UVR [285].

V. [392]. v0.1 [2401]. v1.0 [606]. V1.1 [1501]. V9 [2996]. Vaginal [1401]. Valenciennes [2408]. Validate [1683]. Validated [1109, 290]. Validation [2202, 1127, 1564, 3139, 1848, 2370, 405, 1542]. Validity [393]. Valle [1387]. Valley [1477, 1074]. Valorization [972, 2745]. Valuable [178]. Value [1787, 1790, 713, 258, 833, 3116, 322, 2665, 363, 2745, 1886]. Values [2224, 3191, 351, 866, 299, 2333, 1724, 1361, 2931, 490, 1890]. Valuing [817]. Valve [375]. vannamei [2905, 2589, 3205, 2361, 2952, 3061, 2314, 662, 2372]. Vanuatu [2487]. Vanuatu/New [2487]. Vaquita [1857]. Variabilities [3100]. Variability [2233, 2535, 2968, 424, 1630, 2698, 708, 1137, 3203, 1678, 2371, 638, 2255, 1716, 2575, 2026, 2955, 133, 494, 1308, 2365, 1507, 1061, 128, 2019, 272, 2902, 2469, 2853, 524, 1192, 2475, 1217, 13, 1423, 201, 2063, 1471, 1340, 879, 1449, 1234, 2542, 528, 664, 1973, 3145, 321, 2205, 2909, 1400, 896, 3151, 3138, 2495, 119, 834, 750, 3112, 899, 1321, 2059, 1317, 2828, 1371, 467, 1762, 58, 547, 506, 1821, 366, 158, 729, 1984, 2197, 2937, 514, 642, 253, 1075, 1214, 2719, 613, 1263, 617, 781, 978, 1736, 2212, 1210, 1917, 2389, 2472, 3133, 1825]. Variability [1062, 3073, 1436, 317]. Variable [2251, 585, 1595, 2983, 1826, 285, 2180, 1002, 1056, 2496, 804, 347, 2412, 2588]. Variables [2421, 722, 1929, 1965, 2398, 1884]. Variance [1651]. Variances [1969]. Variate [2749]. Variation [1640, 2812, 748, 1433, 1795, 2807, 2682, 1692, 1981, 2439, 1569, 1394, 2677, 2714, 1995, 839, 2409, 798, 1289, 1893, 1157, 3006, 256, 977, 1959, 3034, 2556, 2101, 2977, 355, 836, 3002, 938, 490, 582, 2053, 1743, 3018, 1980, 3088, 1934, 2416, 655, 2988, 1685, 22, 3093]. Variations [1978, 107, 1789, 1778, 1296, 2288, 1078, 816, 1334, 566, 1151, 2855, 1384, 142, 2814, 311, 1062, 1134]. Varies [1761]. Various [2208, 2978, 2672, 1767, 3212]. Vary [1660, 2381]. Varying [2071, 427, 1410, 453]. Vasotocin [712]. Vast [518]. Vazella [2475, 1217, 1493]. Vector [2207]. Vectors [2616]. Vegetation [2173, 1331, 2531, 2164, 2769]. Vehicle [771, 1477, 1457, 1696]. Vehicles [1804, 322, 43]. Veined [2408]. velezensis [2905]. Velocimeter [2604]. Velocimetry [2308]. Velocities [411]. Velocity [802, 3053]. Venezuela [1185, 1633]. Venice [3047, 1037, 2891]. venosa [2408]. Vent [3058, 131, 48, 237, 2910, 1947, 302, 1357, 2650, 1136]. Ventilation [696, 1223]. Vents [97, 2166, 397]. Venus [1102, 2595, 3146, 2124, 2870, 2501, 2388]. Verde [1921, 3023]. Verification [417]. vermiculophyllum [468]. Versus [1924, 2421, 862, 2530, 2606]. Vertebrae [228, 2134, 2562, 2223, 290]. Vertebral [1090, 1092]. Vertebrate [1731]. Vertical [2233, 2301, 2151, 329, 314, 1469, 337, 704, 961, 2051, 873, 133, 3115, 654, 922, 1424, 202, 2478, 151, 2288, 461, 1002, 1858, 2141, 1429, 2252, 1371, 168, 2227, 1821, 1610, 1457, 2381, 535, 1436]. Very [1930, 1267, 2742, 1439]. vesiculosus [2110]. Vessel [224, 2091, 2243, 2089, 189, 2242, 492, 2035, 1766, 212, 728, 2161, 2006, 1099, 1408, 2245, 1604, 1922, 1983, 394, 1079, 27, 1844, 2748]. Vessel-Associated [1766]. Vessels [1444, 538, 2616, 2184, 2572]. vexillum [303, 2776]. via [615, 1039, 1316, 710, 2020, 1667, 900, 2101, 2977, 2671, 1199, 2726]. Viability [1181, 3170, 479, 1411, 372]. Viable [1492, 1881, 2524]. Vibrio [2214, 2802, 2589, 1204, 3061, 475, 1881, 392, 1609]. Vicente [2222]. Vicinity [202]. Vicious [46]. Victorgorgiidae [803]. Video [1676, 1236, 562, 256, 2398, 2989, 2924, 3022]. Videos [2005]. Vieques [2139]. Vietnam [1265, 3001, 2696]. View [2591, 1848]. Views [1527]. Vigo [1643]. VII [1606]. Vincent [2137]. vincta [898]. Viral [2121, 2848, 846, 2953]. Viral-Like [2953]. virgata [3188]. virgatus [3188]. Virgin [2673, 2322]. Virginia [157]. virginica [967, 700, 327]. viridis [1650, 2710]. Virtual [3046, 2621]. Viruses [1828, 920]. Visayan [1815]. Vision [270, 799]. Visitor [2018]. Visitors [685]. Visual [921, 332]. Visualising [2332]. Visualize [1576]. Vital [3101, 296]. Vitamin [3027, 1269]. Vitellogenesis [263]. vitro [663, 162, 3201, 2587, 3148, 2680]. vitulina [421]. Viviparous [1012]. VME [562]. Vocal [1114]. Vocalizations [1123, 823, 244]. VOCs [1308]. VOI [1886]. Voice [1318]. Volatile [1308, 512]. Volatilomes [209]. Volcanic [97, 397]. Volcanism [1117]. Volcano [2374, 343, 94, 1976, 2438, 9, 2105]. Volcanogenic [192]. volitans [537]. Voltammetric [1293, 298]. Volturno [395]. Volume [44, 1830, 2487, 2582, 2212]. Voluntary [1757, 2244, 100]. voratum [2868]. Vorticity [2277]. Voucher [2533]. vs [2867, 1577]. Vuku [2338]. vulgare [219]. Vulnerability [813, 2274, 3015, 1315, 1391, 1434, 583, 2206, 2381, 3085, 148]. Vulnerable [234, 121, 257, 1391, 1775, 2118, 198, 1597, 56, 562]. vulnificus [1204, 1881, 1609].

W [1538, 2897]. W-Linked [2897]. W/Healthy [1538]. Wadden [837, 2363]. Waginella [1537]. Wake [1616, 484, 152]. Walk [2146]. Wall [2918]. Walrus [1449]. Wander [2145]. War [3114]. Warehou [1486, 1806]. Warm [779, 813, 440, 782, 1149, 924, 602, 904, 742]. Warm-Core [782]. Warm-Low [1149]. Warmed [2291]. Warmer [1876, 51]. Warmest [1938]. Warming [346, 1042, 2794, 800, 2800, 1845, 230, 1977, 2859, 3029, 1596, 949, 1415, 3012, 2110, 1251, 2082, 2843, 348, 616, 214, 954, 2809, 1412, 386, 2833, 1734, 686, 2565, 1218, 2023, 3019, 1661, 3085, 3204]. Warning [1700, 1928, 2842, 301]. Was [3130]. Washing [3135]. Washington [2263, 212]. Waste [965, 2745, 1049]. Wastewater [953]. Wasting [3091, 1031]. Watching [1754, 1235, 1647, 1444, 2850, 1563, 1604, 1607, 1387, 1211, 1621, 1619, 1231]. Water [2233, 1328, 763, 2794, 1986, 2377, 2061, 390, 1761, 2231, 1496, 337, 1924, 2393, 2633, 1911, 2770, 3123, 2427, 1903, 1505, 2240, 1580, 121, 507, 1838, 922, 456, 2085, 96, 722, 1929, 823, 1254, 277, 2498, 2677, 1547, 2385, 770, 2928, 1471, 2717, 1195, 235, 2629, 924, 714, 2403, 1435, 2746, 46, 548, 2598, 2978, 304, 1570, 2862, 1906, 1552, 449, 2705, 1634, 2297, 10, 2963, 2264, 2984, 1990, 1588, 2405, 9, 2105, 750, 901, 2954, 2152, 2876, 2193, 1966, 1858, 467, 597, 1694, 1881, 2030, 1368, 341, 1454, 906, 3128, 2671, 2057, 822, 477, 1908]. Water [2530, 3182, 2835, 1138, 315, 1952, 2723, 1497, 539, 3206, 409, 2720, 2727, 984, 1810, 2381, 2814, 1794, 3036, 1931, 2029, 843, 1493, 2181, 2951, 3093, 3085, 1439, 1342, 80, 1814, 733]. Water- [1547]. Water-Column [507, 722, 1929]. Water-Mediated [96]. Waterborne [604, 2831]. Waters [2507, 2334, 2968, 1808, 1095, 2974, 373, 234, 2427, 2944, 3104, 1505, 792, 986, 1518, 795, 541, 2382, 1853, 1097, 1843, 3063, 1125, 445, 298, 1550, 2161, 2018, 1302, 1025, 2140, 2109, 720, 734, 2454, 2512, 2979, 1460, 416, 725, 1324, 636, 658, 749, 1487, 1899, 750, 452, 214, 2193, 1546, 660, 2320, 2234, 2373, 1307, 147, 2227, 122, 37, 15, 1223, 426, 2538, 3076, 2287, 1747, 184, 614, 1287, 831, 609, 2446, 3190]. Wave [676, 1053, 453, 251, 2092, 1659, 2928, 2656, 2873, 1164, 3015, 2433, 442, 1628, 1648, 838, 2197, 1077, 463, 3198, 2832, 2553, 1121, 1210, 301, 2606]. Wave-Atmosphere [1210]. Wave-Driven [676, 838, 301, 2606]. Wave-Exposed [3198]. Wave-Flow-Sediment [2873]. Wave-Ice [2656]. Wave-Induced [1164]. Wavelength [2501]. Wavelengths [314]. Waves [2100, 3028, 411, 2944, 669, 108, 1998, 281, 2821, 2660, 2604, 1968, 433, 594, 1275, 70, 1587, 1075, 2818, 179, 529, 2357, 2773]. WAVEWATCH [2656]. Way [1519, 1744]. Ways [1820, 2282]. Weak [30]. Weakening [1223]. Weaker [1517]. Wealth [555]. Weaning [1212]. Weather [122, 1381]. Weathered [1011]. Weathering [2994, 1063]. Web [907, 607, 1189, 2815, 1793, 1066, 1953, 2164, 1157, 227, 2276, 743, 1613]. Webbnesia [1921]. Webs [615, 1759, 1375, 2957, 738, 2517, 1143, 1432, 1339]. Weddell [919, 2714, 1317, 37]. Weight [828, 1626, 2543, 3071, 3068]. Weird [956]. Well [1150]. Well-Being [1150]. Wellbeing [2224]. Wenchang [2066]. West [2619, 3126, 1671, 562, 503, 1485, 3118, 845, 686, 2506, 1931, 2551, 76, 2760, 589, 541, 1234, 1447, 1335, 1026, 208, 2910, 559, 1985, 2664].

West-Atlantic [208]. West-Central [1985]. Western [2334, 471, 1320, 2249, 628, 1183, 3177, 495, 2766, 1089, 477, 617, 2422, 2293, 1917, 1378, 2499, 1328, 998, 3066, 329, 2483, 1102, 2974, 2621, 2845, 2099, 389, 2865, 2187, 2085, 128, 2019, 2902, 1512, 1100, 1097, 2488, 1359, 2110, 2999, 1819, 1195, 1666, 2746, 1765, 803, 1101, 966, 1882, 50, 2923, 1869, 543, 908, 2947, 1898, 1412, 721, 244, 1726, 780, 2200, 1610, 561, 86, 1698, 2281, 2814, 3036, 1935, 1338, 2830, 1287, 2838, 2523, 935]. Wet [1234]. Wetland [163, 2643, 2448, 122, 3001]. Wetland-Derived [122]. Wetland-Influenced [2643]. Wetlands [1228]. Whale [1470, 1754, 1235, 2605, 2324, 1527, 23, 203, 1647, 3132, 1444, 438, 823, 2465, 2900, 1130, 2333, 1221, 516, 3126, 1108, 2163, 1038, 2721, 290, 2850, 2947, 1682, 1048, 58, 1604, 1922, 525, 1607, 5, 2424, 1387, 2468, 95, 1148, 27, 1211, 1621, 2191, 1619, 1572, 1960, 1231, 1934]. Whale-Watching [1647, 2850, 1604, 1607, 1387, 1231]. Whaler [1909]. Whales [1235, 224, 2948, 1360, 76, 2089, 2094, 2034, 1559, 2619, 1584, 1939, 728, 2333, 1099, 1909, 236, 1714, 2081, 2245, 244, 58, 1922, 1607, 170, 2468, 2191, 1231, 1034, 2391]. Whaling [2137]. Whelk [2408]. Where [2126, 1759, 1645, 2126, 1577]. Which [1772, 806]. While [2695, 845]. Whip [2992]. Whistle [1563]. White [489, 2589, 370, 2614, 2902, 2375, 524, 1237, 2148, 3061, 1264, 2497, 2314, 1854, 2938, 2372]. Whitespace [500]. Whitetip [1092]. Who [1649, 1759, 649, 1060, 2145, 3087]. Whole [1425, 356]. Whom [1786]. Wide [542, 53, 175, 32, 1955, 1208, 1517, 760, 295, 47, 796, 2963, 87, 938, 865, 2125, 2886]. Wide-Band [2963]. Wide-Swath [87]. Wider [2122, 2353]. Widespread [1110, 2791, 1442, 3186]. Wiegmann [3188]. Wild [1437, 1244, 1362, 1551, 1762, 187, 2452, 985, 762, 2622, 2503]. Wildlife [1867, 1804, 2665, 1857, 1743]. Wildlife-Related [2665]. Wildmeat [1633]. Will [793, 436, 649]. Win [845]. Wind [2100, 3028, 2692, 611, 1669, 3024, 1704, 1141, 1968, 481, 2236, 3208, 2876, 911, 529, 1381, 1144, 3033, 2529, 3073, 164, 407]. Wind-Based [3028]. Wind-Driven [3024, 1141, 3073, 407]. Wind-Sourced [611]. Wind-Waves [2100]. Windfarm [2091, 1710]. Window [1326, 1238]. Windrows [1379, 410]. Winds [1560, 2617]. Winged [1569]. Winning [1233]. Winter [1979, 1778, 1192, 2714, 868, 1483, 2656, 1802, 2032, 2819, 2580, 2040, 2506, 1931, 2316]. Winter-Spring [1483]. Wintering [706]. Winters [779, 1412]. Wintertime [1153]. Within [2028, 146, 2784, 1372, 1224, 189, 1646, 2365, 1901, 2687, 541, 2244, 1600, 1290, 12, 2715, 1279, 2638, 2904, 2545, 2241, 1959, 913, 2906, 2458, 386, 113, 1668, 2600, 1227, 1660, 211, 1702, 2765, 3165]. Within- [2784]. Within-Generational [386]. Without [627, 612, 1298]. Wnt [2361]. Wnt/ [2361]. Women [2188, 1333, 401]. Won’t [1829]. Wood [1080, 1356, 154]. Wood-Boring [154]. Wooden [1080, 1356]. Work [108, 1915, 2941, 2407]. Workflow [2398, 1848]. Working [1182]. World [1640, 3080, 1739, 827, 2841, 495, 1735]. Worldwide [303, 1100]. Worm [2166, 2537, 3051]. Worms [1987, 1842, 2343, 1959]. Wounding [496]. Wrack [2745]. Wrasse [2455]. WRF [1381]. Wrong [2560, 1548].

Xisha [3124]. Xylophagaidae [1080, 1356]. Xylophagaids [154].

Y4I [2310]. Yangtze [1547, 2715, 2056, 361, 1370, 3078, 3207, 2287, 2588, 1542]. Yap [2212, 1553]. Year [1616, 1865, 748, 1397, 3101, 94, 1887, 479, 145, 448, 2906, 1089, 3, 1069, 2686]. Year-Round [2906]. Years [2596, 466, 2070, 1992, 2313, 1460, 2546, 1406, 1266, 1797, 2922, 3088, 2951, 2622, 3022]. Yellow [163, 1991, 3077, 3196, 2761, 266, 843, 2181, 574, 20, 2898, 3040, 1941, 534, 1889, 1858, 493, 788, 1205]. Yellow-Legged [1991]. Yellowfin [241, 1744, 2086]. Yellowtail [2264]. yezoensis [295, 496, 22]. Yield [3000, 948]. Young [1865, 2695, 3105, 1854]. Yourself [1722]. Yucatan [2746, 1776, 709, 1158].

Z [2897]. Z- [2897]. Zalophus [1244, 1768]. Zealand [23, 2955, 1617, 492, 3069, 957, 284, 2300, 2943, 2723, 719, 1034, 1074, 3190]. Zebrafish [3049, 1165]. zetlandica [776]. Zhejiang [2056]. Zhikong [2877, 1951, 1805]. Zinc [208]. Zoe [1468]. Zonation [1429]. Zone [867, 2397, 1813, 866, 462, 1746, 1914, 2715, 1625, 1164, 2068, 2183, 1335, 2052, 933, 1717, 2508, 1454, 3194, 635, 1421, 2198, 211, 2107, 2050, 1641, 1747, 1553, 1987, 2187, 560, 2063, 1765]. Zones [104, 418, 3086, 1154]. Zoning [2764, 731]. zonipectis [2280]. Zoonotic [2426, 437]. Zooplankton [376, 275, 704, 1750, 591, 2575, 2382, 1483, 1625, 1520, 647, 460, 1299, 1802, 73, 1020, 1452, 553, 729, 493, 1287, 2438]. Zooplankton-Mediated [460]. Zootechnology [215]. Zooxanthellae [1043, 2798]. Zostera [1031, 602, 3035, 1429, 865, 1028]. zosteroides [2331].

References

Sawakuchi:2017:CDE

[1] Henrique O. Sawakuchi, Vania Neu, Nicholas D. Ward, Maria de Lourdes C. Barros, Aline M. Valerio, William Gagne-Maynard, Alan C. Cunha, Diani F. S. Less, Joel E. M. Diniz, Daimio C. Brito, Alex V. Krusche, and Jeffrey E. Richey. Carbon dioxide emissions along the Lower Amazon River. Frontiers in Marine Science, 4, March 2017. ISSN 2296-7745. See corrigendum [790].

Langenkamper:2017:PBB

[2] Daniel Langenkämper, Martin Zurowietz, Timm Schoening, and Tim W. Nattkemper. BIIGLE 2.0 — browsing and annotating large marine image collections. Frontiers in Marine Science, 4, March 2017. ISSN 2296-7745. See corrigendum [1330].

Steneck:2019:MRR

[3] Robert S. Steneck, Suzanne N. Arnold, Robert Boenish, Ramón de León, Peter J. Mumby, Douglas B. Rasher, and Margaret W. Wilson. Managing recovery resilience in coral reefs against climate-induced bleaching and hurricanes: a 15 year case study from Bonaire, Dutch Caribbean. Frontiers in Marine Science, 6, June 2019. ISSN 2296-7745. See commentary [448].

Testor:2019:OCI

[4] Pierre Testor, Brad de Young, Daniel L. Rudnick, Scott Glenn, Daniel Hayes, Craig M. Lee, Charitha Pattiaratchi, Katherine Hill, Emma Heslop, Victor Turpin, Pekka Alenius, Carlos Barrera, John A. Barth, Nicholas Beaird, Guislain Bécu, Anthony Bosse, François Bourrin, J. Alexander Brearley, Yi Chao, Sue Chen, Jacopo Chiggiato, Laurent Coppola, Richard Crout, James Cummings, Beth Curry, Ruth Curry, Richard Davis, Kruti Desai, Steve DiMarco, Catherine Edwards, Sophie Fielding, Ilker Fer, Eleanor Frajka-Williams, Hezi Gildor, Gustavo Goni, Dimitri Gutierrez, Peter Haugan, David Hebert, Joleen Heiderich, Stephanie Henson, Karen Heywood, Patrick Hogan, Loïc Houpert, Sik Huh, Mark E. Inall, Masso Ishii, Shin ichi Ito, Sachihiko Itoh, Sen Jan, Jan Kaiser, Johannes Karstensen, Barbara Kirkpatrick, Jody Klymak, Josh Kohut, Gerd Krahmann, Marjolaine Krug, Sam McClatchie, Frédéric Marin, Elena Mauri, Avichal Mehra, Michael P. Meredith, Thomas Meunier, Travis Miles, Julio M. Morell, Laurent Mortier, Sarah Nicholson, Joanne O’Callaghan, Diarmuid O’Conchubhair, Peter Oke, Enric Pallàs-Sanz, Matthew Palmer, JongJin Park, Leonidas Perivoliotis, Pierre-Marie Poulain, Ruth Perry, Bastien Queste, Luc Rainville, Eric Rehm, Moninya Roughan, Nicholas Rome, Tetjana Ross, Simon Ruiz, Grace Saba, Amandine Schaeffer, Martha Schönau, Katrin Schroeder, Yugo Shimizu, Bernadette M. Sloyan, David Smeed, Derrick Snowden, Yumi Song, Sebastian Swart, Miguel Tenreiro, Andrew Thompson, Joaquin Tintore, Robert E. Todd, Cesar Toro, Hugh Venables, Taku Wagawa, Stephanie Waterman, Roy A. Watlington, and Doug Wilson. OceanGliders: a component of the integrated GOOS. Frontiers in Marine Science, 6, October 2019. ISSN 2296-7745. See corrigendum [2466].

Shimabukuro:2019:DDP

[5] Mauricio Shimabukuro, Orlemir Carrerette, Joan Manel Alfaro-Lucas, Alexandra Elaine Rizzo, Kenneth M. Halanych, and Paulo Yukio Gomes Sumida. Diversity, distribution and phylogeny of Hesionidae (Annelida) colonizing whale falls: New species of Sirsoe and connections between ocean basins. Frontiers in Marine Science, 6, August 2019. ISSN 2296-7745. See corrigendum [2424].

Kennedy:2019:UUI

[6] Brian R. C. Kennedy, Kasey Cantwell, Mashkoor Malik, Christopher Kelley, Jeremy Potter, Kelley Elliott, Elizabeth Lobecker, Lindsay McKenna Gray, Derek Sowers, Michael P. White, Scott C. France, Steven Auscavitch, Christopher Mah, Virginia Moriwake, Sarah R. D. Bingo, Meagan Putts, and Randi D. Rotjan. The unknown and the unexplored: Insights into the Pacific deep-sea following NOAA CAPSTONE Expeditions. Frontiers in Marine Science, 6, August 2019. ISSN 2296-7745. See corrigendum [92].

Hjerne:2019:CDC

[7] Olle Hjerne, Susanna Hajdu, Ulf Larsson, Andrea S. Downing, and Monika Winder. Climate driven changes in timing, composition and magnitude of the Baltic Sea phytoplankton spring bloom. Frontiers in Marine Science, 6, August 2019. ISSN 2296-7745. See corrigendum [1346].

Muelbert:2019:IIL

[8] José H. Muelbert, Nicholas J. Nidzieko, Alicia T. R. Acosta, Stace E. Beaulieu, Angelo F. Bernardino, Elmira Boikova, Thomas G. Bornman, Bruno Cataletto, Klaas Deneudt, Erika Eliason, Alexandra Kraberg, Masahiro Nakaoka, Alessandra Pugnetti, Olivier Ragueneau, Mirco Scharfe, Thomas Soltwedel, Heidi M. Sosik, Angela Stanisci, Kremena Stefanova, Pierre Stéphan, Adrian Stier, Johan Wikner, and Adriana Zingone. ILTER — the international long-term ecological research network as a platform for global coastal and ocean observation. Frontiers in Marine Science, 6, August 2019. ISSN 2296-7745. See corrigendum [84].

MosqueraGimenez:2019:OCN

[9] Ángela Mosquera Giménez, Pedro Vélez-Belchí, Jesús Rivera, Safo Piñeiro, Noelia Fajar, Verónica Caínzos, Rosa Balbín, Juan Antonio Jiménez Aparicio, Carlos Dominguez-Carrió, Jordi Blasco-Ferre, Marina Carreiro-Silva, Telmo Morato, Patricia Puerta, and Covadonga Orejas. Ocean circulation over North Atlantic underwater features in the path of the Mediterranean outflow water: The Ormonde and Formigas Seamounts, and the Gazul Mud Volcano. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745. See corrigendum [2105].

Meredyk:2020:EDW

[10] Shawn P. Meredyk, Evan Edinger, David J. W. Piper, Veerle A. I. Huvenne, Shannon Hoy, and Alan Ruffman. Enigmatic deep-water mounds on the Orphan Knoll, Labrador Sea. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Elzahaby:2019:OIS

[11] Youstina Elzahaby and Amandine Schaeffer. Observational insight into the subsurface anomalies of marine heatwaves. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Hayes:2019:DMP

[12] Keith R. Hayes, Geoffrey R. Hosack, Emma Lawrence, Paul Hedge, Neville S. Barrett, Rachel Przeslawski, M. Julian Caley, and Scott D. Foster. Designing monitoring programs for Marine Protected Areas within an evidence based decision making paradigm. Frontiers in Marine Science, 6, November 2019. ISSN 2296-7745.

Hernandez-Fernandez:2019:SSV

[13] Leslie Hernández-Fernández, Roberto González de Zayas, Laura Weber, Amy Apprill, and Maickel Armenteros. Small-scale variability dominates benthic coverage and diversity across the Jardines de La Reina, Cuba Coral Reef System. Frontiers in Marine Science, 6, November 2019. ISSN 2296-7745.

Cordova-Alarcon:2019:GDP

[14] Valentina R. Córdova-Alarcón, Cristián Araneda, Felipe Jilberto, Piero Magnolfi, María Isabel Toledo, and Natalia Lam. Genetic diversity and population structure of Genypterus chilensis, a commercial benthic marine species of the South Pacific. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Schratzberger:2019:ESE

[15] Michaela Schratzberger, Suzanna Neville, Suzanne Painting, Keith Weston, and Lucille Paltriguera. Ecological and socio-economic effects of highly protected marine areas (HPMAs) in temperate waters. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Gorsky:2019:ETO

[16] Gabriel Gorsky, Guillaume Bourdin, Fabien Lombard, Maria Luiza Pedrotti, Samuel Audrain, Nicolas Bin, Emmanuel Boss, Chris Bowler, Nicolas Cassar, Loic Caudan, Genevieve Chabot, Natalie R. Cohen, Daniel Cron, Colomban De Vargas, John R. Dolan, Eric Douville, Amanda Elineau, J. Michel Flores, Jean Francois Ghiglione, Nils Haëntjens, Martin Hertau, Seth G. John, Rachel L. Kelly, Ilan Koren, Yajuan Lin, Dominique Marie, Clémentine Moulin, Yohann Moucherie, Stéphane Pesant, Marc Picheral, Julie Poulain, Mireille Pujo-Pay, Gilles Reverdin, Sarah Romac, Mathew B. Sullivan, Miri Trainic, Marc Tressol, Romain Troublé, Assaf Vardi, Christian R. Voolstra, Patrick Wincker, Sylvain Agostini, Bernard Banaigs, Emilie Boissin, Didier Forcioli, Paola Furla, Pierre E. Galand, Eric Gilson, Stéphanie Reynaud, Shinichi Sunagawa, Olivier P. Thomas, Rebecca Lisette Vega Thurber, Didier Zoccola, Serge Planes, Denis Allemand, and Eric Karsenti. Expanding Tara oceans protocols for underway, ecosystemic sampling of the ocean– atmosphere interface during Tara Pacific Expedition (2016–2018). Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Toullec:2019:CGI

[17] Jordan Toullec, Dorothée Vincent, Laura Frohn, Philippe Miner, Manon Le Goff, Jérémy Devesa, and Brivaëla Moriceau. Copepod grazing influences diatom aggregation and particle dynamics. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Cappelli:2019:SMA

[18] Elena Lo Giudice Cappelli and William E. N. Austin. Size matters: Analyses of benthic foraminiferal assemblages across differing size fractions. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Andrisoa:2019:EGR

[19] Aladin Andrisoa, Franck Lartaud, Valentí Rodellas, Ingrid Neveu, and Thomas C. Stieglitz. Enhanced growth rates of the Mediterranean mussel in a coastal lagoon driven by groundwater inflow. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Yu:2019:EST

[20] Xiang Yu, Fangzhu Wu, Xiaoqun Xu, Quanzhen Chen, Lin Huang, Berhane Teklehaimanot Tesfai, Liang Cao, Xudan Xu, Shuozeng Dou, and Wei Huang. Effects of short term methylmercury exposure on growth and development of the large yellow croaker embryos and larvae. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Schiltroth:2019:IAS

[21] Braeden W. J. Schiltroth, Kyle T. Ohori, and Sherryl R. Bisgrove. Impacts of acidic seawater on early developmental stages of Fucus gardneri at Burrard Inlet, British Columbia. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Xu:2019:OGV

[22] Kuipeng Xu, Xinzi Yu, Xianghai Tang, Fanna Kong, and Yunxiang Mao. Organellar genome variation and genetic diversity of Chinese Pyropia yezoensis. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Barlow:2019:SDA

[23] Dawn R. Barlow, Acacia L. Pepper, and Leigh G. Torres. Skin deep: an assessment of New Zealand blue whale skin condition. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Janssen:2019:GPB

[24] René Janßen, Wael Skeff, Johannes Werner, Marisa A. Wirth, Bernd Kreikemeyer, Detlef Schulz-Bull, and Matthias Labrenz. A glyphosate pulse to brackish long-term microcosms has a greater impact on the microbial diversity and abundance of planktonic than of biofilm assemblages. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Mackelworth:2019:GMC

[25] Peter Charles Mackelworth, Yael Teff Seker, Tomás Vega Fernández, Márcia Marques, Fátima Lopes Alves, Giovanni D’Anna, Darren A. Fa, David Goldborough, Zacharoula Kyriazi, Cristina Pita, Michelle E. Portman, Bob Rumes, Stephen J. Warr, and Draško Holcer. Geopolitics and marine conservation: Synergies and conflicts. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Perez-Sanchez:2019:GST

[26] Jaume Pérez-Sánchez, Fernando Naya-Català, Beatriz Soriano, M. Carla Piazzon, Ahmed Hafez, Toni Gabaldón, Carlos Llorens, Ariadna Sitjà-Bobadilla, and Josep A. Calduch-Giner. Genome sequencing and transcriptome analysis reveal recent species-specific gene duplications in the plastic gilthead sea bream (Sparus aurata). Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Szesciorka:2019:CSN

[27] Angela R. Szesciorka, Ann N. Allen, John Calambokidis, James Fahlbusch, Megan F. McKenna, and Brandon Southall. A case study of a near vessel strike of a blue whale: Perceptual cues and fine-scale aspects of behavioral avoidance. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Jani:2020:BBR

[28] Jarina Mohd Jani, Muhammad Allim Jamalludin, and Seh Ling Long. To ban or not to ban? Reviewing an ongoing dilemma on sea turtle egg trade in Terengganu, Malaysia. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Straub:2019:REE

[29] Sandra C. Straub, Thomas Wernberg, Mads S. Thomsen, Pippa J. Moore, Michael T. Burrows, Ben P. Harvey, and Dan A. Smale. Resistance, extinction, and everything in between — the diverse responses of seaweeds to marine heatwaves. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Hansen:2019:WSS

[30] Lillian Hansen. The weak sustainability of the salmon feed transition in Norway — a bioeconomic case study. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Dessandier:2019:BFA

[31] Pierre-Antoine Dessandier, Chiara Borrelli, Dimitri Kalenitchenko, and Giuliana Panieri. Benthic foraminifera in Arctic methane hydrate bearing sediments. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Cao:2019:GWI

[32] Min Cao, Dongmei Wang, Fanna Kong, Junhao Wang, Kuipeng Xu, and Yunxiang Mao. A genome-wide identification of osmotic stress-responsive MicroRNAs in Pyropia haitanensis (Bangiales, Rhodophyta). Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Hathorne:2020:REE

[33] Ed C. Hathorne, Martin Frank, and P. M. Mohan. Rare earth elements in Andaman Island surface seawater: Geochemical tracers for the monsoon? Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Andersen:2019:ERM

[34] Jesper H. Andersen, Jacob Carstensen, Marianne Holmer, Dorte Krause-Jensen, and Katherine Richardson. Editorial: Research and management of eutrophication in coastal ecosystems. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Monteiro:2019:TMS

[35] Cátia Monteiro, Sandra Heinrich, Inka Bartsch, Klaus Valentin, Erwan Corre, Jonas Collén, Lars Harms, Gernot Glöckner, and Kai Bischof. Temperature modulates sex-biased gene expression in the gametophytes of the kelp Saccharina latissima. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Guzman:2019:BSR

[36] Hector M. Guzman, Grant Rogers, and Catalina G. Gomez. Behavioral states related to environmental conditions and fisheries during olive Ridley turtle migration from Pacific Panama. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Sanchez:2019:WSC

[37] Nicolas Sanchez, Christian S. Reiss, Osmund Holm-Hansen, Christopher D. Hewes, Kemal C. Bizsel, and Murat V. Ardelan. Weddell–Scotia confluence effect on the iron distribution in waters surrounding the South Shetland (Antarctic Peninsula) and South Orkney (Scotia Sea) Islands during the austral summer in 2007 and 2008. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Porporato:2020:SSF

[38] Erika M. D. Porporato, Roberto Pastres, and Daniele Brigolin. Site suitability for finfish marine aquaculture in the Central Mediterranean Sea. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Daniel:2020:THU

[39] Anne Daniel, Agathe Laës-Huon, Carole Barus, Alexander D. Beaton, Daniel Blandfort, Nathalie Guigues, Marc Knockaert, Dominique Munaron, Ian Salter, E. Malcolm S. Woodward, Naomi Greenwood, and Eric P. Achterberg. Toward a harmonization for using in situ nutrient sensors in the marine environment. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Riascos:2019:ENN

[40] José M. Riascos, Dimitri Gutiérrez, Rubén Escribano, and Sven Thatje. Editorial: El Niño–Southern Oscillation on a changing planet: Consequences for coastal ecosystems. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Hollowed:2020:IME

[41] Anne Babcock Hollowed, Kirstin Kari Holsman, Alan C. Haynie, Albert J. Hermann, Andre E. Punt, Kerim Aydin, James N. Ianelli, Stephen Kasperski, Wei Cheng, Amanda Faig, Kelly A. Kearney, Jonathan C. P. Reum, Paul Spencer, Ingrid Spies, William Stockhausen, Cody S. Szuwalski, George A. Whitehouse, and Thomas K. Wilderbuer. Integrated modeling to evaluate climate change impacts on coupled social-ecological systems in Alaska. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Fourquez:2020:MCS

[42] Marion Fourquez, Matthieu Bressac, Stacy L. Deppeler, Michael Ellwood, Ingrid Obernosterer, Thomas W. Trull, and Philip W. Boyd. Microbial competition in the subpolar Southern Ocean: an Fe–C colimitation experiment. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Zappa:2020:USD

[43] Christopher J. Zappa, Scott M. Brown, Nathan J. M. Laxague, Tejendra Dhakal, Ryan A. Harris, Aaron M. Farber, and Ajit Subramaniam. Using ship-deployed high-endurance unmanned aerial vehicles for the study of ocean surface and atmospheric boundary layer processes. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Bochdansky:2019:ECF

[44] Alexander B. Bochdansky, Robert B. Dunbar, Dennis A. Hansell, and Gerhard J. Herndl. Estimating carbon flux from optically recording total particle volume at depths below the primary pycnocline. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Cherubin:2020:FSA

[45] Laurent M. Chérubin, Fraser Dalgleish, Ali Khaleel Ibrahim, Michelle Schärer-Umpierre, Richard S. Nemeth, Anthony Matthews, and Richard Appeldoorn. Fish spawning aggregations dynamics as inferred from a novel, persistent presence robotic approach. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Kusch:2019:DWV

[46] Stephanie Kusch, Albert Benthien, Klaus-Uwe Richter, Björn Rost, and Gesine Mollenhauer. Dead in the water: The vicious cycle of blanks during natural level 14 C manipulation of marine algal cultures. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Liu:2020:GWD

[47] Bing-Jian Liu, Yu-Long Li, Bai-Dong Zhang, and Jin-Xian Liu. Genomewide discovery of single-nucleotide polymorphisms and their application in population genetic studies in the endangered Japanese eel (Anguilla japonica). Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Cowart:2020:EED

[48] Dominique A. Cowart, Marjolaine Matabos, Miriam I. Brandt, Julien Marticorena, and Jozée Sarrazin. Exploring environmental DNA (eDNA) to assess biodiversity of hard substratum faunal communities on the Lucky Strike Vent Field (Mid-Atlantic Ridge) and investigate recolonization dynamics after an induced disturbance. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Jones:2020:EMT

[49] Alice R. Jones, Ramesh Raja Segaran, Kenneth D. Clarke, Michelle Waycott, William S. H. Goh, and Bronwyn M. Gillanders. Estimating mangrove tree biomass and carbon content: a comparison of forest inventory techniques and drone imagery. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Marchini:2020:RAC

[50] Chiara Marchini, Giada Tortorelli, Elena Guidi, Valentina Airi, Giuseppe Falini, Zvy Dubinsky, and Stefano Goffredo. Reproduction of the azooxanthellate coral Caryophyllia inornata is not affected by temperature along an 850 km gradient on the Western Italian Coast. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Koutsouveli:2020:GWT

[51] Vasiliki Koutsouveli, Tereza Manousaki, Ana Riesgo, Jacques Lagnel, Spyros Kollias, Costas S. Tsigenopoulos, Christos Arvanitidis, Costas Dounas, Antonios Magoulas, and Thanos Dailianis. Gearing up for warmer times: Transcriptomic response of spongia officinalis to elevated temperatures reveals recruited mechanisms and potential for resilience. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Christiansen:2019:NSB

[52] Svenja Christiansen, Josefin Titelman, and Stein Kaartvedt. Nighttime swimming behavior of a mesopelagic fish. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Bogan:2020:CGW

[53] Samuel N. Bogan, Kevin M. Johnson, and Gretchen E. Hofmann. Changes in genome-wide methylation and gene expression in response to future pCO_{2} extremes in the Antarctic pteropod Limacina helicina antarctica. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Turon:2020:NIA

[54] Marta Turon and Maria J. Uriz. New insights into the archaeal consortium of tropical sponges. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Schiffrine:2020:GES

[55] Nicolas Schiffrine, Jean-Éric Tremblay, and Marcel Babin. Growth and elemental stoichiometry of the ecologically-relevant Arctic diatom Chaetoceros gelidus: a mix of polar and temperate. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Yan:2020:PSG

[56] Ruo-Jin Yan, Kareen E. Schnabel, Ashley A. Rowden, Xiang-Zhao Guo, and Jonathan P. A. Gardner. Population structure and genetic connectivity of squat lobsters (Munida Leach, 1820) associated with vulnerable marine ecosystems in the Southwest Pacific Ocean. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Dimond:2020:CDM

[57] James L. Dimond and Steven B. Roberts. Convergence of DNA methylation profiles of the reef coral Porites astreoides in a novel environment. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Redfern:2020:EVS

[58] Jessica V. Redfern, Elizabeth A. Becker, and Thomas J. Moore. Effects of variability in ship traffic and whale distributions on the risk of ships striking whales. Frontiers in Marine Science, 6, February 2020. ISSN 2296-7745.

Esposito:2020:SMI

[59] Maria Consiglia Esposito, Raffaele Boni, Alessia Cuccaro, Elisabetta Tosti, and Alessandra Gallo. Sperm motility impairment in free spawning invertebrates under near-future level of ocean acidification: Uncovering the mechanism. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Hoerterer:2020:SPO

[60] Christina Hoerterer, Maximilian F. Schupp, Andreas Benkens, Dustin Nickiewicz, Gesche Krause, and Bela H. Buck. Stakeholder perspectives on opportunities and challenges in achieving sustainable growth of the blue economy in a changing climate. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Chung:2020:ERC

[61] Ming-Tsung Chung, Kuo-Fang Huang, Chen-Feng You, Chuan-Chin Chiao, and Chia-Hui Wang. Elemental ratios in cuttlebone indicate growth rates in the cuttlefish Sepia pharaonis. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Cho:2020:UBM

[62] Hyeyoun Cho, Chung Yeon Hwang, Jong-Geol Kim, Sanghoon Kang, Katrin Knittel, Ayeon Choi, Sung-Han Kim, Sung-Keun Rhee, Eun Jin Yang, SangHoon Lee, and Jung-Ho Hyun. A unique benthic microbial community underlying the Phaeocystis antarctica-dominated Amundsen Sea Polynya, Antarctica: a proxy for assessing the impact of global changes. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Matsumoto:2020:LRJ

[63] George I. Matsumoto, Bastian Bentlage, Rob Sherlock, Kristine Walz, and Bruce H. Robison. “Little Red Jellies” in Monterey Bay, California (Cnidaria: Hydrozoa: Trachymedusae: Rhopalonematidae). Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Stadtlander:2020:BRS

[64] Christian T. K.-H. Stadtländer. Book review: Sustainable Shipping in a Changing Arctic. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Xu:2020:CCC

[65] Houguo Xu, Bo Sun, Linlin Jia, Yuliang Wei, Zhangbin Liao, and Mengqing Liang. Cloning and characterization of cholesterol 25hydroxylase (ch25h) from a marine teleost, Chinese tongue sole (Cynoglossus semilaevis), and its gene expressions in response to dietary arachidonic acid. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Bathmann:2020:ELA

[66] Ulrich Bathmann, Hendrik Schubert, Elinor Andrén, Laura Tuomi, Teresa Radziejewska, Karol Kulinski, and Irina Chubarenko. Editorial: Living along gradients: Past, present, future. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Palmer:2020:RSD

[67] Stephanie C. J. Palmer, Pierre M. Gernez, Yoann Thomas, Stefan Simis, Peter I. Miller, Philippe Glize, and Laurent Barillé. Remote sensingdriven Pacific oyster (Crassostrea gigas) growth modeling to inform offshore aquaculture site selection. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Moore:2020:HAB

[68] Kathleen M. Moore, Edward H. Allison, Stacia J. Dreyer, Julia A. Ekstrom, Sunny L. Jardine, Terrie Klinger, Stephanie K. Moore, and Karma C. Norman. Harmful algal blooms: Identifying effective adaptive actions used in fishery-dependent communities in response to a protracted event. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Masetti:2020:AGM

[69] Giuseppe Masetti, Michael J. Smith, Larry A. Mayer, and John G. W. Kelley. Applications of the Gulf of Maine operational forecast system to enhance spatio-temporal oceanographic awareness for ocean mapping. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Senafi:2020:IWC

[70] Fahad Al Senafi and Ayal Anis. Internal waves on the continental shelf of the Northwestern Arabian Gulf. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Jossart:2020:ASA

[71] Jonathan Jossart, Seth J. Theuerkauf, Lisa C. Wickliffe, and James A. Morris Jr. Applications of spatial autocorrelation analyses for marine aquaculture siting. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Gil-Agudelo:2020:CRG

[72] Diego L. Gil-Agudelo, Carlos E. Cintra-Buenrostro, Jorge Brenner, Patricia González-Díaz, William Kiene, Caitlin Lustic, and Horacio Pérez-España. Coral reefs in the Gulf of Mexico large marine ecosystem: Conservation status, challenges, and opportunities. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Kvale:2020:CER

[73] Karin F. Kvale, A. E. Friederike Prowe, and Andreas Oschlies. A critical examination of the role of marine snow and zooplankton fecal pellets in removing ocean surface microplastic. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Starke:2020:EPF

[74] Claudia Starke, Werner Ekau, and Nils Moosdorf. Enhanced productivity and fish abundance at a submarine spring in a coastal lagoon on Tahiti, French Polynesia. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Harper:2020:FCE

[75] Kirsten J. Harper, Kelly D. Goodwin, Lynsey R. Harper, Erin L. La-Casella, Amy Frey, and Peter H. Dutton. Finding crush: Environmental DNA analysis as a tool for tracking the green sea turtle Chelonia mydas in a marine estuary. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Boye:2020:EAW

[76] Tenna K. Boye, Eva Garde, Julius Nielsen, Rasmus Hedeholm, Jesper Olsen, and Malene Simon. Estimating the age of West Greenland humpback whales through aspartic acid racemization and eye lens bomb radiocarbon methods. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Nguyen:2020:RIN

[77] Hung Manh Nguyen, Narendra Singh Yadav, Simon Barak, Fernando P. Lima, Yuval Sapir, and Gidon Winters. Responses of invasive and native populations of the seagrass Halophila stipulacea to simulated climate change. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Ortiz-Alvarez:2020:RAL

[78] Clara Ortiz-Alvarez, Mariela Pajuelo, Daniel Grados, Marino E. Abrego, Ana Rebeca Barragán-Rocha, Mariano Barrantes, Alejandro Cotto Sánchez, Luis G. Fonseca, Velkiss Gadea Espinal, Jeffrey C. Mangel, Juan M. Rguez-Baron, Pilar Santidrian-Tomillo, Laura Sarti, Heriberto Santana-Hernández, George Shillinger, Myriam Prado, Bryan Wallace, Amanda S. Williard, Alan A. Zavala-Norzagaray, and Joanna Alfaro-Shigueto. Rapid assessments of leatherback small-scale fishery bycatch in internesting areas in the Eastern Pacific Ocean. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Carr-Harris:2020:EEB

[79] Andrew Carr-Harris and Scott Steinback. Expected economic and biological impacts of recreational Atlantic striped bass fishing policy. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Ziervogel:2020:SCH

[80] Kai Ziervogel and Carol Arnosti. Substantial carbohydrate hydrolase activities in the water column of the Guaymas Basin (Gulf of California). Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Thewes:2020:SSS

[81] Daniel Thewes, Emil V. Stanev, and Oliver Zielinski. Sensitivity of a 3D shelf sea ecosystem model to parameterizations of the underwater light field. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Ramirez-Gallego:2020:STS

[82] Cristian Ramirez-Gallego and Karla G. Barrientos-Muñoz. Sea turtles at Serrana Island and Serranilla Island, Seaflower Biosphere Reserve, Colombian Caribbean. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Grattepanche:2020:TBC

[83] Jean-David Grattepanche and Laura A. Katz. Top-down and bottomup controls on microeukaryotic diversity (i.e., amplicon analyses of SAR lineages) and function (i.e., metatranscriptome analyses) assessed in microcosm experiments. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Muelbert:2020:CII

[84] José H. Muelbert, Nicholas J. Nidzieko, Alicia T. R. Acosta, Stace E. Beaulieu, Angelo F. Bernardino, Elmira Boikova, Thomas G. Bornman, Bruno Cataletto, Klaas Deneudt, Erika Eliason, Alexandra Kraberg, Masahiro Nakaoka, Alessandra Pugnetti, Olivier Ragueneau, Mirco Scharfe, Thomas Soltwedel, Heidi M. Sosik, Angela Stanisci, Kremena Stefanova, Pierre Stéphan, Adrian Stier, Johan Wikner, and Adriana Zingone. Corrigendum: ILTER — the International Long-Term Ecological Research Network as a platform for global coastal and ocean observation. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745. See [8].

Bellas:2020:ENC

[85] Juan Bellas, Ketil Hylland, and Thierry Burgeot. Editorial: New challenges in marine pollution monitoring. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Sugie:2020:ITC

[86] Koji Sugie, Amane Fujiwara, Shigeto Nishino, Sohiko Kameyama, and Naomi Harada. Impacts of temperature, CO_{2}, and salinity on phytoplankton community composition in the Western Arctic Ocean. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Metref:2020:WSA

[87] Sammy Metref, Emmanuel Cosme, Florian Le Guillou, Julien Le Sommer, Jean-Michel Brankart, and Jacques Verron. Wide-swath altimetric satellite data assimilation with correlated-error reduction. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Zunino:2020:CES

[88] Serena Zunino, Donata Melaku Canu, Francesco Marangon, and Stefania Troiano. Cultural ecosystem services provided by coralligenous assemblages and Posidonia oceanica in the Italian Seas. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Sarnocinska:2020:HPP

[89] Joanna Sarnocińska, Jonas Teilmann, Jeppe Dalgaard Balle, Floris M. van Beest, Matthieu Delefosse, and Jakob Tougaard. Harbor porpoise (Phocoena phocoena) reaction to a 3D seismic airgun survey in the North Sea. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Kwong:2020:MMA

[90] Lian E. Kwong, Natasha Henschke, Evgeny A. Pakhomov, Jason D. Everett, and Iain M. Suthers. Mesozooplankton and micronekton active carbon transport in contrasting eddies. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745. See corrigendum [810].

Afonso:2020:AMA

[91] Pedro Afonso, Jorge Fontes, Eva Giacomello, Maria C. Magalhães, Helen R. Martins, Telmo Morato, Verónica Neves, Rui Prieto, Ricardo S. Santos, Mónica A. Silva, and Frédèric Vandeperre. The Azores: a Mid-Atlantic hotspot for marine megafauna research and conservation. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Kennedy:2020:CUU

[92] Brian R. C. Kennedy, Kasey Cantwell, Mashkoor Malik, Christopher Kelley, Jeremy Potter, Kelley Elliott, Elizabeth Lobecker, Lindsay McKenna Gray, Derek Sowers, Michael P. White, Scott C. France, Steven Auscavitch, Christopher Mah, Virginia Moriwake, Sarah R. D. Bingo, Meagan Putts, and Randi D. Rotjan. Corrigendum: The unknown and the unexplored: Insights into the Pacific deep-sea following NOAA CAPSTONE Expeditions. Frontiers in Marine Science, 6, February 2020. ISSN 2296-7745. See [6].

Vianna:2020:LTF

[93] Gabriel M. S. Vianna, E. James Hehre, Rachel White, Lincoln Hood, Brittany Derrick, and Dirk Zeller. Long-term fishing catch and effort trends in the Republic of the Marshall Islands, with emphasis on the small-scale sectors. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Gonzalez-Vega:2020:SRD

[94] Alba González-Vega, Eugenio Fraile-Nuez, J. Magdalena Santana-Casiano, Melchor González-Dávila, José Escánez-Pérez, María Gómez-Ballesteros, Olvido Tello, and Jesús M. Arrieta. Significant release of dissolved inorganic nutrients from the shallow submarine volcano Tagoro (Canary Islands) based on seven-year monitoring. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Steinman:2020:CEC

[95] K. J. Steinman, T. R. Robeck, G. A. Fetter, T. L. Schmitt, S. Osborn, S. DiRocco, H. H. Nollens, and J. K. O’Brien. Circulating and excreted corticosteroids and metabolites, hematological, and serum chemistry parameters in the killer whale (Orcinus orca) before and after a stress response. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Fong:2020:CWM

[96] Jenny Fong, Lindsey K. Deignan, Andrew G. Bauman, Peter D. Steinberg, Diane McDougald, and Peter A. Todd. Contact- and watermediated effects of macroalgae on the physiology and microbiome of three Indo–Pacific coral species. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Gollner:2020:ACD

[97] Sabine Gollner, Breea Govenar, Pedro Martinez Arbizu, Lauren S. Mullineaux, Susan Mills, Nadine Le Bris, Markus Weinbauer, Timothy M. Shank, and Monika Bright. Animal community dynamics at senescent and active vents at the 9^{o}N East Pacific Rise after a volcanic eruption. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Tortorelli:2020:HGE

[98] Giada Tortorelli, Roy Belderok, Simon K. Davy, Geoffrey I. McFadden, and Madeleine J. H. van Oppen. Host genotypic effect on algal symbiosis establishment in the coral model, the anemone Exaiptasia diaphana, from the Great Barrier Reef. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Giering:2020:SOP

[99] Sarah Lou Carolin Giering, Emma Louise Cavan, Sünnje Linnéa Basedow, Nathan Briggs, Adrian B. Burd, Louise J. Darroch, Lionel Guidi, Jean-Olivier Irisson, Morten H. Iversen, Rainer Kiko, Dhugal Lindsay, Catarina R. Marcolin, Andrew M. P. McDonnell, Klas Ove Möller, Uta Passow, Sandy Thomalla, Thomas William Trull, and Anya M. Waite. Sinking organic particles in the ocean — flux estimates from in situ optical devices. Frontiers in Marine Science, 6, February 2020. ISSN 2296-7745.

Nocito:2020:GCB

[100] Emily S. Nocito, Cassandra M. Brooks, and Aaron L. Strong. Gazing at the crystal ball: Predicting the future of Marine Protected Areas through voluntary commitments. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Lee:2020:BFB

[101] Elizabeth M. J. Lee and Kathleen G. O’Malley. Big fishery, big data, and little crabs: Using genomic methods to examine the seasonal recruitment patterns of early life stage Dungeness crab (Cancer magister) in the California Current ecosystem. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Borja:2020:ECP

[102] Angel Borja, Francesca Santoro, Gail Scowcroft, Stephen Fletcher, and Pierre Strosser. Editorial: Connecting people to their oceans: Issues and options for effective ocean literacy. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Serratosa:2020:EDS

[103] Juan Serratosa, K. David Hyrenbach, Diego Miranda-Urbina, Matías Portflitt-Toro, Nicolás Luna, and Guillermo Luna-Jorquera. Environmental drivers of seabird at-sea distribution in the Eastern South Pacific Ocean: Assemblage composition across a longitudinal productivity gradient. Frontiers in Marine Science, 6, February 2020. ISSN 2296-7745.

Baroni:2020:EOC

[104] Itzel Ruvalcaba Baroni, Virginia Palastanga, and Caroline P. Slomp. Enhanced organic carbon burial in sediments of oxygen minimum zones upon ocean deoxygenation. Frontiers in Marine Science, 6, January 2020. ISSN 2296-7745.

Bedulli:2020:CSB

[105] Camila Bedulli, Paul S. Lavery, Matt Harvey, Carlos M. Duarte, and Oscar Serrano. Contribution of seagrass blue carbon toward carbon neutral policies in a touristic and environmentally-friendly island. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Johnson:2020:SCI

[106] Clare Johnson, Mark Inall, Stefan Gary, and Stuart Cunningham. Significance of climate indices to benthic conditions across the Northern North Atlantic and adjacent shelf seas. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Chou:2020:CCD

[107] Wen-Chen Chou, Pi-Jen Liu, Ying-Hsuan Chen, and Wei-Jen Huang. Contrasting changes in diel variations of net community calcification support that carbonate dissolution can be more sensitive to ocean acidification than coral calcification. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Cox:2020:SRS

[108] Rónadh Cox, Fabrice Ardhuin, Frédéric Dias, Ronan Autret, Nicole Beisiegel, Claire S. Earlie, James G. Herterich, Andrew Kennedy, Raphaël Paris, Alison Raby, Pál Schmitt, and Robert Weiss. Systematic review shows that work done by storm waves can be misinterpreted as tsunami-related because commonly used hydrodynamic equations are flawed. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Friedman:2020:RPA

[109] Whitney R. Friedman, Benjamin S. Halpern, Elizabeth McLeod, Michael W. Beck, Carlos M. Duarte, Carrie V. Kappel, Arielle Levine, Robert D. Sluka, Steven Adler, Casey C. O’Hara, Eleanor J. Sterling, Sebastian Tapia-Lewin, Iñigo J. Losada, Tim R. McClanahan, Linwood Pendleton, Margaret Spring, James P. Toomey, Kenneth R. Weiss, Hugh P. Possingham, and Jensen R. Montambault. Research priorities for achieving healthy marine ecosystems and human communities in a changing climate. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Yano:2020:EDM

[110] Cassia F. Yano, Manuel A. Merlo, Silvia Portela-Bens, Marcelo de B. Cioffi, Luiz A. C. Bertollo, Célio D. Santos-Júnior, and Laureana Rebordinos. Evolutionary dynamics of multigene families in Triportheus (Characiformes, Triportheidae): a transposon mediated mechanism? Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Krayesky-Self:2020:FRE

[111] Sherry Krayesky-Self, Delena Phung, William Schmidt, Thomas Sauvage, Luke Butler, and Suzanne Fredericq. First report of endolithic members of Rhodosorus marinus (Stylonematales, Rhodophyta) growing inside rhodoliths offshore Louisiana, Northwestern Gulf of Mexico. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Flynn:2020:SFK

[112] Nichole Flynn and Naweed I. Syed. Src family kinases play a role in the functional clustering of central postsynaptic nicotinic acetylcholine receptors. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Sowers:2020:SGC

[113] Derek C. Sowers, Giuseppe Masetti, Larry A. Mayer, Paul Johnson, James V. Gardner, and Andrew A. Armstrong. Standardized geomorphic classification of seafloor within the United States Atlantic canyons and continental margin. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Vriend:2020:RAF

[114] Paul Vriend, Caroline van Calcar, Merel Kooi, Harm Landman, Remco Pikaar, and Tim van Emmerik. Rapid assessment of floating macroplastic transport in the Rhine. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

vanderGeest:2020:FFB

[115] Matthijs van der Geest, Tjisse van der Heide, Marianne Holmer, and Rutger de Wit. First field-based evidence that the seagrass-lucinid mutualism can mitigate sulfide stress in seagrasses. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Freire:2020:EGC

[116] Kátia Meirelles Felizola Freire, Dyhia Belhabib, Jeniffer C. Espedido, Lincoln Hood, Kristin M. Kleisner, Vicky W. L. Lam, Michel L. Machado, Jocemar Tomasino Mendonça, Jessica J. Meeuwig, Pietro S. Moro, Fábio S. Motta, Maria-Lourdes D. Palomares, Nicola Smith, Lydia Teh, Dirk Zeller, Kyrstn Zylich, and Daniel Pauly. Estimating global catches of marine recreational fisheries. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Scucchia:2020:PCS

[117] Federica Scucchia, Hagai Nativ, Maayan Neder, Gretchen Goodbody-Gringley, and Tali Mass. Physiological characteristics of Stylophora pistillata larvae across a depth gradient. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Vasapollo:2020:IMB

[118] Claudio Vasapollo, Massimo Virgili, Giada Bargione, Andrea Petetta, Rocco De Marco, Elisa Punzo, and Alessandro Lucchetti. Impact on macro-benthic communities of hydraulic dredging for razor clam Ensis minor in the Tyrrhenian Sea. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Milligan:2020:DOE

[119] Rosanna J. Milligan and Tracey T. Sutton. Dispersion overrides environmental variability as a primary driver of the horizontal assemblage structure of the mesopelagic fish family Myctophidae in the Northern Gulf of Mexico. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

James:2020:IME

[120] Anna K. James, Libe Washburn, Chris Gotschalk, Stéphane Maritorena, Alice Alldredge, Craig E. Nelson, James L. Hench, James J. Leichter, Alex S. J. Wyatt, and Craig A. Carlson. An island mass effect resolved near Mo’orea, French Polynesia. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

DuPreez:2020:CWC

[121] Cherisse Du Preez, Kelly D. Swan, and Janelle M. R. Curtis. Coldwater corals and other vulnerable biological structures on a North Pacific seamount after half a century of fishing. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Rudolph:2020:UGH

[122] Jacob C. Rudolph, Carli A. Arendt, Alexandria G. Hounshell, Hans W. Paerl, and Christopher L. Osburn. Use of geospatial, hydrologic, and geochemical modeling to determine the influence of wetland-derived organic matter in coastal waters in response to extreme weather events. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Timm:2020:CPG

[123] Laura E. Timm, Lys M. Isma, Matthew W. Johnston, and Heather D. Bracken-Grissom. Comparative population genomics and biophysical modeling of shrimp migration in the Gulf of Mexico reveals currentmediated connectivity. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Fabbrizzi:2020:MMF

[124] Erika Fabbrizzi, Michele Scardi, Enric Ballesteros, Lisandro Benedetti-Cecchi, Emma Cebrian, Giulia Ceccherelli, Francesco De Leo, Alan Deidun, Giuseppe Guarnieri, Annalisa Falace, Silvia Fraissinet, Chiara Giommi, Vesna Mačić, Luisa Mangialajo, Anna Maria Mannino, Luigi Piazzi, Mohamed Ramdani, Gil Rilov, Luca Rindi, Lucia Rizzo, Gianluca Sarà, Jamila Ben Souissi, Ergun Taskin, and Simonetta Fraschetti. Modeling macroalgal forest distribution at Mediterranean scale: Present status, drivers of changes and insights for conservation and management. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Senafi:2020:DDB

[125] Fahad Al Senafi, Raeid M. M. Abed, Thirumahal Muthukrishnan, Ayal Anis, Ahmed N. Al Rawahi, and Huda Mahmoud. Development and diversity of bacterial biofilms in response to internal tides, a case study off the coast of Kuwait. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Snyder:2020:SST

[126] Jordan N. Snyder, Tom W. Bell, David A. Siegel, Nicholas J. Nidzieko, and Kyle C. Cavanaugh. Sea surface temperature imagery elucidates spatiotemporal nutrient patterns for offshore kelp aquaculture siting in the Southern California Bight. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Pita:2020:AKG

[127] Pablo Pita, Josep Alós, Manel Antelo, Iñaki Artetxe, Sebastián Biton-Porsmoguer, Arnau Carreño, Amalia Cuadros, Toni Font, José Beiro, Jose A. García-Charton, Ana Gordoa, Kieran Hyder, Josep Lloret, Beatriz Morales-Nin, Estanis Mugerza, Oscar Sagué, José J. Pascual-Fernández, Jon Ruiz, Virginia Sandoval, Elena Santolini, Lucía Zarauz, and Sebastián Villasante. Assessing knowledge gaps and management needs to cope with barriers for environmental, economic, and social sustainability of marine recreational fisheries: The case of Spain. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745.

Fox:2020:PGP

[128] James Fox, Michael J. Behrenfeld, Nils Haëntjens, Alison Chase, Sasha J. Kramer, Emmanuel Boss, Lee Karp-Boss, Nerissa L. Fisher, W. Bryce Penta, Toby K. Westberry, and Kimberly H. Halsey. Phytoplankton growth and productivity in the Western North Atlantic: Observations of regional variability from the NAAMES field campaigns. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745. See corrigendum [2019].

Davies:2020:TTC

[129] Kathryn K. Davies, Karen T. Fisher, Gemma Couzens, Andrew Allison, Elizabeth Ingrid van Putten, Jeffrey M. Dambacher, Melissa Foley, and Carolyn J. Lundquist. Trans-Tasman cumulative effects management: a comparative study. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Hu:2020:DRC

[130] Xinping Hu, Hongming Yao, Cory J. Staryk, Melissa R. McCutcheon, Michael S. Wetz, and Lily Walker. Disparate responses of carbonate system in two adjacent subtropical estuaries to the influence of Hurricane Harvey — a case study. Frontiers in Marine Science, 7, January 2020. ISSN 2296-7745. See corrigendum [1904].

Clague:2020:LFE

[131] David A. Clague, Jennifer B. Paduan, David W. Caress, James McClain, and Robert A. Zierenberg. Lava flows erupted in 1996 on North Gorda Ridge segment and the geology of the nearby sea cliff hydrothermal vent field from 1-M resolution AUV mapping. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Sando:2020:CCN

[132] Anne Britt Sandø, Geir Odd Johansen, Asgeir Aglen, Jan Erik Stiansen, and Angelika H. H. Renner. Climate change and new potential spawning sites for Northeast Arctic cod. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Cornwell:2020:RCO

[133] Louise Elisabeth Cornwell, Elaine S. Fileman, John T. Bruun, Andrew Garwood Hirst, Glen Adam Tarran, Helen S. Findlay, Ceri Lewis, Timothy James Smyth, A. J. McEvoy, and A. Atkinson. Resilience of the copepod Oithona similis to climatic variability: Egg production, mortality, and vertical habitat partitioning. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

DeCarlo:2020:CRF

[134] Thomas M. DeCarlo. Commentary: Reconstructing four centuries of temperature-induced coral bleaching on the Great Barrier Reef. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745. See [997].

Kamat:2020:EFM

[135] Siya Kamat, Madhuree Kumari, Sidhartha Taritla, and C. Jayabaskaran. Endophytic fungi of marine alga from Konkan Coast, India — a rich source of bioactive material. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Yang:2020:NMA

[136] Chao Yang, Xianpeng Su, Dapeng Liu, Zhongwen Guo, Fang Wang, and Yunliang Lu. A new method of aquatic animal personality analysis based on machine learning (PAML): Taking swimming crab Portunus trituberculatus as an example. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Liu:2020:SPA

[137] Haijiao Liu, Chao Wu, Wenzhe Xu, Xingzhou Wang, Satheeswaran Thangaraj, Guicheng Zhang, Xiaodong Zhang, Yuying Zhao, and Jun Sun. Surface phytoplankton assemblages and controlling factors in the Strait of Malacca and Sunda Shelf. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Muller:2020:EHA

[138] Marius N. Müller, Jorge I. Mardones, and Juan J. Dorantes-Aranda. Editorial: Harmful algal blooms (HABs) in Latin America. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Drymon:2020:DSS

[139] J. M. Drymon, S. Dedman, J. T. Froeschke, E. A. Seubert, A. E. Jefferson, A. M. Kroetz, J. F. Mareska, and S. P. Powers. Defining sex-specific habitat suitability for a Northern Gulf of Mexico shark assemblage. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Ribeiro:2020:AAM

[140] Lourenço Ribeiro, Vanda Brotas, Tania Hernández-Fariñas, Bruno Jesus, and Laurent Barillé. Assessing alternative microscopy-based approaches to species abundance description of intertidal diatom communities. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Borja:2020:MTA

[141] Angel Borja, Mathew P. White, Elisa Berdalet, Nikolaj Bock, Claire Eatock, Peter Kristensen, Anne Leonard, Josep Lloret, Sabine Pahl, Mariluz Parga, Julia Vera Prieto, Susanne Wuijts, and Lora E. Fleming. Moving toward an agenda on ocean health and human health in Europe. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Tseng:2020:SVM

[142] Hsiao-Chun Tseng, Wan-Lynn You, Wei Huang, Chih-Ching Chung, An-Yi Tsai, Tzong-Yueh Chen, Kuo-Wei Lan, and Gwo-Ching Gong. Seasonal variations of marine environment and primary production in the Taiwan Strait. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Haro:2020:MNM

[143] Sara Haro, Miguel Lara, Irene Laiz, Carlos José González, Julio Bohórquez, Emilio Garcia-Robledo, Alfonso Corzo, and Sokratis Papaspyrou. Microbenthic net metabolism along intertidal gradients (Cadiz Bay, SW Spain): Spatio-temporal patterns and environmental factors. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Harden-Davies:2020:SCC

[144] Harriet Harden-Davies and Paul Snelgrove. Science collaboration for capacity building: Advancing technology transfer through a treaty for biodiversity beyond national jurisdiction. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Luo:2020:RDH

[145] Zhen Luo, Fengze Miao, Menghong Hu, and Youji Wang. Research development on horseshoe crab: a 30-year bibliometric analysis. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Auscavitch:2020:ODD

[146] Steven R. Auscavitch, Mary C. Deere, Abigail G. Keller, Randi D. Rotjan, Timothy M. Shank, and Erik E. Cordes. Oceanographic drivers of deep-sea coral species distribution and community assembly on seamounts, islands, atolls, and reefs within the Phoenix Islands Protected Area. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Ribeiro:2020:RSN

[147] Andréa de Lima Ribeiro and Helen Pask. Remote sensing of natural waters using a multichannel, lidar-compatible Raman spectrometer and blue excitation. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

deJuan:2020:VDF

[148] Silvia de Juan, Hilmar Hinz, Paolo Sartor, Sergio Vitale, Luis Bentes, Jose M. Bellido, Claudia Musumeci, Daniela Massi, Vita Gancitano, and Montserrat Demestre. Vulnerability of demersal fish assemblages to trawling activities: a traits-based index. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Kessler:2020:ISP

[149] Nivi Kessler, Stephan M. Kraemer, Yeala Shaked, and Walter D. C. Schenkeveld. Investigation of siderophore-promoted and reductive dissolution of dust in marine microenvironments such as Trichodesmium colonies. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Capet:2020:EFM

[150] Arthur Capet, Perran Cook, Emilio Garcia-Robledo, Babette Hoogakker, Aurélien Paulmier, Christophe Rabouille, and Raquel Vaquer-Sunyer. Editorial: Facing marine deoxygenation. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Judkins:2020:VDP

[151] Heather Judkins and Michael Vecchione. Vertical distribution patterns of cephalopods in the Northern Gulf of Mexico. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Carella:2020:WOM

[152] Francesca Carella, Elisabetta Antuofermo, Simone Farina, Fulvio Salati, Daniela Mandas, Patricia Prado, Rossella Panarese, Fabio Marino, Eleonora Fiocchi, Tobia Pretto, and Gionata De Vico. In the wake of the ongoing mass mortality events: Co-occurrence of Mycobacterium, Haplosporidium and other pathogens in Pinna nobilis collected in Italy and Spain (Mediterranean Sea). Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745. See corrigendum [484].

Arlidge:2020:MHA

[153] William N. S. Arlidge, Dale Squires, Joanna Alfaro-Shigueto, Hollie Booth, Jeffrey C. Mangel, and E. J. Milner-Gulland. A mitigation hierarchy approach for managing sea turtle captures in small-scale fisheries. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Voight:2020:SIE

[154] Janet R. Voight, Jacob C. Cooper, and Raymond W. Lee. Stable isotopic evidence of mixotrophy in xylophagaids, deep-sea wood-boring bivalves. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Boswell:2020:OSL

[155] Kevin M. Boswell, Marta D’Elia, Matthew W. Johnston, John A. Mohan, Joseph D. Warren, R. J. David Wells, and Tracey T. Sutton. Oceanographic structure and light levels drive patterns of sound scattering layers in a low-latitude oceanic system. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Califano:2020:CMH

[156] Gianmaria Califano, Michiel Kwantes, Maria Helena Abreu, Rodrigo Costa, and Thomas Wichard. Cultivating the macroalgal Holobiont: Effects of integrated multi-trophic aquaculture on the microbiome of Ulva rigida (Chlorophyta). Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Beckensteiner:2020:BEO

[157] Jennifer Beckensteiner, David M. Kaplan, and Andrew M. Scheld. Barriers to eastern oyster aquaculture expansion in Virginia. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Romero:2020:TVP

[158] Isabel C. Romero, Heather Judkins, and Michael Vecchione. Temporal variability of polycyclic aromatic hydrocarbons in deep-sea cephalopods of the Northern Gulf of Mexico. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745. See corrigendum [366].

Mateus:2020:WTS

[159] Marcos Mateus, Ricardo Canelas, Ligia Pinto, and Nuno Vaz. When tragedy strikes: Potential contributions from ocean observation to search and rescue operations after drowning accidents. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Roncalli:2020:CBD

[160] Vittoria Roncalli, Matthew C. Cieslak, Russell R. Hopcroft, and Petra H. Lenz. Capital breeding in a diapausing copepod: a transcriptomics analysis. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Katerina:2020:MSL

[161] Kousoulaki Katerina, Gerd Marit Berge, Mørkøre Turid, Krasnov Aleksei, Baeverfjord Grete, Ytrestøyl Trine, Carlehög Mats, Sweetman John, and Ruyter Bente. Microalgal Schizochytrium limacinum biomass improves growth and filet quality when used long-term as a replacement for fish oil, in modern salmon diets. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Gogineni:2020:MSS

[162] Vedanjali Gogineni, Joonseok Oh, Amanda L. Waters, Michelle Kelly, Robert Stone, and Mark T. Hamann. Monanchocidin A from subarctic sponges of the genus Monanchora and their promising selectivity against melanoma in vitro. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

An:2020:ISA

[163] Sung-Uk An, Hyeyoun Cho, Ui-Jung Jung, Bomina Kim, Hyeonji Lee, and Jung-Ho Hyun. Invasive Spartina anglica greatly alters the rates and pathways of organic carbon oxidation and associated microbial communities in an intertidal wetland of the Han River Estuary, Yellow Sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

vandenBurg:2020:GRM

[164] Sander W. K. van den Burg, Christine Röckmann, Jennifer L. Banach, and Luc van Hoof. Governing risks of multi-use: Seaweed aquaculture at offshore wind farms. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Smith:2020:SOC

[165] Marié E. Smith and Stewart Bernard. Satellite ocean color based harmful algal bloom indicators for aquaculture decision support in the Southern Benguela. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Collins:2020:MBD

[166] Jane Eva Collins, Thomas Vanagt, Isabelle Huys, and Helena Vieira. Marine bioresource development — stakeholder’s challenges, implementable actions, and business models. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745. See corrigendum [260].

Schnurr:2020:PAF

[167] Peter J. Schnurr, Mark C. Drever, Robert W. Elner, John Harper, and Michael T. Arts. Peak abundance of fatty acids from intertidal biofilm in relation to the breeding migration of shorebirds. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Robison:2020:RGA

[168] Bruce H. Robison, Rob E. Sherlock, Kim R. Reisenbichler, and Paul R. McGill. Running the gauntlet: Assessing the threats to vertical migrators. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Neves:2020:GMB

[169] Antonio A. Sepp Neves, Nadia Pinardi, Antonio Navarra, and F. Trotta. A general methodology for beached oil spill hazard mapping. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Smith:2020:QSS

[170] Joshua N. Smith, Natalie Kelly, Simon Childerhouse, Jessica V. Redfern, Thomas J. Moore, and David Peel. Quantifying ship strike risk to breeding whales in a multiple-use marine park: The Great Barrier Reef. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Tang:2020:CDB

[171] Kai Tang. Chemical diversity and biochemical transformation of biogenic organic sulfur in the ocean. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745. See corrigendum [382].

Anton:2020:DLM

[172] Andrea Anton, Kimberlee Baldry, Darren J. Coker, and Carlos M. Duarte. Drivers of the low metabolic rates of seagrass meadows in the Red Sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Nilsen:2020:EBH

[173] Ina Nilsen, Jeppe Kolding, Cecilie Hansen, and Daniel Howell. Exploring balanced harvesting by using an Atlantis ecosystem model for the Nordic and Barents Seas. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Waltham:2020:DER

[174] Nathan J. Waltham, Michael Elliott, Shing Yip Lee, Catherine Lovelock, Carlos M. Duarte, Christina Buelow, Charles Simenstad, Ivan Nagelkerken, Louw Claassens, Colin K-C Wen, Mario Barletta, Rod M. Connolly, Chris Gillies, William J. Mitsch, Matthew B. Ogburn, Jemma Purandare, Hugh Possingham, and Marcus Sheaves. UN decade on ecosystem restoration 2021–2030 — what chance for success in restoring coastal ecosystems? Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Bond:2020:HBE

[175] Robin J. Bond, Colleen M. Hansel, and Bettina M. Voelker. Heterotrophic bacteria exhibit a wide range of rates of extracellular production and decay of hydrogen peroxide. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

German:2020:HAS

[176] Christopher R. German, Joseph A. Resing, Guangyu Xu, Isobel A. Yeo, Sharon L. Walker, Colin W. Devey, James W. Moffett, Gregory A. Cutter, Olivier Hyvernaud, and Dominique Reymond. Hydrothermal activity and seismicity at Teahitia Seamount: Reactivation of the Society Islands hotspot? Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Layton:2020:KFR

[177] Cayne Layton, Melinda A. Coleman, Ezequiel M. Marzinelli, Peter D. Steinberg, Stephen E. Swearer, Adriana Vergés, Thomas Wernberg, and Craig R. Johnson. Kelp forest restoration in Australia. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Turner:2020:MSU

[178] Phillip J. Turner, Bernard Ball, Zoie Diana, Andrea Fariñas-Bermejo, Ian Grace, Doreen McVeigh, Megan M. Powers, Loïc Van Audenhaege, Svetlana Maslakova, Craig M. Young, and Cindy L. Van Dover. Methane seeps on the US Atlantic margin and their potential importance to populations of the commercially valuable deep-sea red crab, Chaceon quinquedens. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Stewart:2020:MCB

[179] Bryce D. Stewart, Leigh M. Howarth, Howard Wood, Kerri Whiteside, William Carney, Éilís Crimmins, Bethan C. O’Leary, Julie P. Hawkins, and Callum M. Roberts. Marine conservation begins at home: How a local community and protection of a small bay sent waves of change around the UK and beyond. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Rafay:2020:GNU

[180] Ramis Rafay, Joao M. Uratani, Hector H. Hernandez, and Jorge Rodríguez. Growth and nitrate uptake in Nannochloropsis gaditana and Tetraselmis chuii cultures grown in sequential batch reactors. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Hellinger:2020:ATA

[181] Jens Hellinger, Peter Jägers, Katharina Spoida, Linda C. Weiss, Melanie D. Mark, and Stefan Herlitze. Analysis of the territorial aggressive behavior of the bioluminescent flashlight fish Photoblepharon steinitzi in the Red Sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Zhang:2020:ESB

[182] Yong Zhang, Bing Chen, and Wei dong Zhai. Exploring sources and biogeochemical dynamics of dissolved methane in the Central Bohai Sea in summer. Frontiers in Marine Science, 7, February 2020. ISSN 2296-

Paczkowska:2020:RCP

[183] Joanna Paczkowska, Sonia Brugel, Owen Rowe, Robert Lefébure, Andreas Brutemark, and Agneta Andersson. Response of coastal phytoplankton to high inflows of terrestrial matter. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Wei:2020:MSN

[184] Zhan-Fei Wei, Wen-Li Li, Jun Li, Jun Chen, Yong-Zhi Xin, Li-Sheng He, and Yong Wang. Multiple in situ nucleic acid collections (MISNAC) from deep-sea waters. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Canales:2020:ECF

[185] T. Mariella Canales, Mauricio Lima, Rodrigo Wiff, Javier E. Contreras-Reyes, Ursula Cifuentes, and José Montero. Endogenous, climate, and fishing influences on the population dynamics of small pelagic fish in the Southern Humboldt Current ecosystem. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Herbst:2020:LUP

[186] Dannieli F. Herbst, Leopoldo Cavaleri Gerhardinger, and Natalia Hanazaki. Linking user-perception diversity on ecosystems services to the inception of coastal governance regime transformation. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Segvic-Bubic:2020:TAI

[187] Tanja Šegvić-Bubić, Iva Žužul, Igor Talijančić, Nika Ugrin, Ivana Lepen Pleić, Luka Žuvić, Nika Stagličić, and Leon Grubišić. Translocation and aquaculture impact on genetic diversity and composition of wild selfsustainable Ostrea edulis populations in the Adriatic Sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Burgunter-Delamare:2020:MCB

[188] Bertille Burgunter-Delamare, Hetty KleinJan, Clémence Frioux, Enora Fremy, Margot Wagner, Erwan Corre, Alicia Le Salver, Cédric Leroux, Catherine Leblanc, Catherine Boyen, Anne Siegel, and Simon M. Dittami. Metabolic complementarity between a brown alga and associated cultivable bacteria provide indications of beneficial interactions. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Cope:2020:ANS

[189] Samantha Cope, Ellen Hines, Roger Bland, Jerry D. Davis, Brendan Tougher, and Virgil Zetterlind. Application of a new shore-based vessel traffic monitoring system within San Francisco Bay. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Vales:2020:LTS

[190] Damián G. Vales, Luis Cardona, Rocío Loizaga, Néstor A. García, and Enrique A. Crespo. Long-term stability in the trophic ecology of a pelagic forager living in a changing marine ecosystem. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Virta:2020:SSA

[191] Leena Virta, Janne Soininen, and Alf Norkko. Stable seasonal and annual alpha diversity of benthic diatom communities despite changing community composition. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Bayon:2020:PRE

[192] Germain Bayon, Grant B. Douglas, Geoff J. Denton, Laurence Monin, and Patrick De Deckker. Preferential riverine export of fine volcanogenic particles to the Southeast Australian Margin. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Kleinhaus:2020:SDR

[193] Karine Kleinhaus, Ali Al-Sawalmih, Daniel J. Barshis, Amatzia Genin, Lola N. Grace, Ove Hoegh-Guldberg, Yossi Loya, Anders Meibom, Eslam O. Osman, Jean-Daniel Ruch, Yonathan Shaked, Christian R. Voolstra, Assaf Zvuloni, and Maoz Fine. Science, diplomacy, and the Red Sea’s unique coral reef: It’s time for action. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Valdez:2020:PEI

[194] Stephanie R. Valdez, Y. Stacy Zhang, Tjisse van der Heide, Mathew A. Vanderklift, Flavia Tarquinio, Robert J. Orth, and Brian R. Silliman. Positive ecological interactions and the success of seagrass restoration. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Rajasabapathy:2020:BCA

[195] Raju Rajasabapathy, Kesava Priyan Ramasamy, Balakrishnan Manikandan, Chellandi Mohandass, and Rathinam Arthur James. Bacterial communities associated with healthy and diseased (skeletal growth anomaly) reef coral Acropora cytherea from Palk Bay, India. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Frolova:2020:IBF

[196] Alexandra Frolova and Maria Pia Miglietta. Insights on bloom forming jellyfish (class: Scyphozoa) in the Gulf of Mexico: Environmental tolerance ranges and limits suggest differences in habitat preference and resistance to climate change among congeners. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Wang:2020:GAB

[197] Lijuan Wang, Zhihao Wu, Yingjun Wang, Mengxia Liu, Aihuan Song, Hongjun Liu, and Feng You. Genetic assessment of a black rockfish, Sebastes schlegelii, stock enhancement program in Lidao Bay, China based on mitochondrial and nuclear DNA analysis. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Rowden:2020:DCD

[198] Ashley A. Rowden, Tabitha R. R. Pearman, David Anthony Bowden, Owen F. Anderson, and Malcolm Ross Clark. Determining coral density thresholds for identifying structurally complex vulnerable marine ecosystems in the deep sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Vasquez-Carrillo:2020:GDD

[199] Catalina Vásquez-Carrillo, Carmen L. Noriega-Hoyos, Leidy Hernandez-Rivera, Guiomar Aminta Jáuregui-Romero, and Kathleen Sullivan Sealey. Genetic diversity and demographic connectivity of Atlantic green sea turtles at foraging grounds in Northeastern Colombia, Caribbean Sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Ogier:2020:RCC

[200] Emily Ogier, Sarah Jennings, Anthony Fowler, Stewart Frusher, Caleb Gardner, Paul Hamer, Alistair J. Hobday, Adrian Linanne, Stephan Mayfield, Craig Mundy, Andrew Sullivan, Geoff Tuck, Tim Ward, and Gretta Pecl. Responding to climate change: Participatory evaluation of adaptation options for key marine fisheries in Australia’s South East. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Hoffken:2020:ETV

[201] Jorid Höffken, Athanasios T. Vafeidis, Leigh R. MacPherson, and Sönke Dangendorf. Effects of the temporal variability of storm surges on coastal flooding. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Frank:2020:VHD

[202] Tamara M. Frank, Charles D. Fine, Eric A. Burdett, April B. Cook, and Tracey T. Sutton. The vertical and horizontal distribution of deep-sea crustaceans in the order Euphausiacea in the vicinity of the DeepWater Horizon oil spill. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Baumgartner:2020:SGP

[203] Mark F. Baumgartner, Julianne Bonnell, Peter J. Corkeron, Sofie M. Van Parijs, Cara Hotchkin, Ben A. Hodges, Jacqueline Bort Thornton, Bryan L. Mensi, and Scott M. Bruner. Slocum gliders provide accurate near real-time estimates of baleen whale presence from human-reviewed passive acoustic detection information. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Marks:2020:RED

[204] Alex D. Marks, David W. Kerstetter, David M. Wyanski, and Tracey T. Sutton. Reproductive ecology of dragonfishes (stomiiformes: Stomiidae) in the Gulf of Mexico. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Norton:2020:IEE

[205] Emily L. Norton, Samantha Siedlecki, Isaac C. Kaplan, Albert J. Hermann, Jennifer L. Fisher, Cheryl A. Morgan, Suzanna Officer, Casey Saenger, Simone R. Alin, Jan Newton, Nina Bednaršek, and Richard A. Feely. The importance of environmental exposure history in forecasting Dungeness crab megalopae occurrence using J-SCOPE, a high-resolution model for the US Pacific Northwest. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Demirel:2020:FLS

[206] Nazli Demirel, Mustafa Zengin, and Aylin Ulman. First large-scale Eastern Mediterranean and Black Sea stock assessment reveals a dramatic decline. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Huang:2020:PTB

[207] Dan Huang, Da-Wei Li, Srinivasan Balamurugan, Jian-Wei Zheng, Wan-Jun Liu, Li-Gong Zou, Wei-Dong Yang, Jie-Sheng Liu, and Hong-Ye Li. Plastidial and ER triacylglycerol biosynthesis in a growth phasedependent manner in the heterokont Nannochloropsis oceanica. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Middag:2020:DNW

[208] Rob Middag, Hein J. W. de Baar, Kenneth W. Bruland, and Steven M. A. C. van Heuven. The distribution of nickel in the West-Atlantic Ocean, its relationship with phosphate and a comparison to cadmium and zinc. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Lawson:2020:VSA

[209] Caitlin A. Lawson, Justin R. Seymour, Malcolm Possell, David J. Suggett, and Jean-Baptiste Raina. The volatilomes of symbiodiniaceaeassociated bacteria are influenced by chemicals derived from their algal partner. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

vanSon:2020:ARE

[210] Thijs Christiaan van Son, Nikolaos Nikolioudakis, Henning Steen, Jon Albretsen, Birgitte Rugaard Furevik, Sigrid Elvenes, Frithjof Moy, and Kjell Magnus Norderhaug. Achieving reliable estimates of the spatial distribution of kelp biomass. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Wagner:2020:DMT

[211] Jamie K. S. Wagner, Clara Smart, and Christopher R. German. Discovery and mapping of the Triton Seep Site, Redondo Knoll: Fluid flow and microbial colonization within an oxygen minimum zone. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Greig:2020:USA

[212] Nathan C. Greig, Ellen M. Hines, Samantha Cope, and XiaoHang Liu. Using satellite AIS to analyze vessel speeds off the coast of Washington State, U.S., as a risk analysis for cetacean-vessel collisions. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Goecke:2020:CDK

[213] Franz Goecke, Gunnar Klemetsdal, and Åshild Ergon. Cultivar development of kelps for commercial cultivation — past lessons and future prospects. Frontiers in Marine Science, 8, February 2020. ISSN 2296-7745.

Overmans:2020:USU

[214] Sebastian Overmans and Susana Agustí. Unraveling the seasonality of UV exposure in reef waters of a rapidly warming (sub-)tropical sea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Capaz:2020:CZL

[215] Juan C. Capaz, Daniel Hernández-Brooke, Safia Balvet, Ana T. Couto, Alexandra C. Alves, Rui A. Gonçalves, Paulo A. Frias, José P. Andrade, and António V. Sykes. Control of zootechnology leads to improved cuttlefish (Sepia officinalis, L.) reproduction performance up to pre-industrial levels. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Goldstein:2020:CEE

[216] Josephine Goldstein, Nicklas Bisbo, Peter Funch, and Hans Ulrik Riisgård. Contraction-expansion and the effects on the aquiferous system in the demosponge Halichondria panicea. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Schijf:2020:ERM

[217] Johan Schijf, Catherine Jeandel, Karen H. Johannesson, and Anne H. Osborne. Editorial: REE marine geochemistry in the 21st Century: a tribute to the pioneering research of Henry Elderfield (1943–2016). Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Booth:2020:MMP

[218] Cormac G. Booth, Rachael R. Sinclair, and John Harwood. Methods for monitoring for the population consequences of disturbance in marine mammals: a review. Frontiers in Marine Science, 7, February 2020. ISSN 2296-7745.

Plouguerne:2020:GSV

[219] Erwan Plouguerné, Lauro M. de Souza, Guilherme L. Sassaki, Claire Hellio, Rozenn Trepos, Bernardo A. P. da Gama, Renato Crespo Pereira, and Eliana Barreto-Bergter. Glycoglycerolipids from Sargassum vulgare as potential antifouling agents. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Xie:2020:CAF

[220] Dizhi Xie, Jialing Ye, Meisong Lu, Shuqi Wang, Cuihong You, and Yuanyou Li. Comparsion of activities of fatty acyl desaturases and elongases among six teleosts with different feeding and ecological habits. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Koopmans:2020:HNP

[221] Dirk Koopmans, Moritz Holtappels, Arjun Chennu, Miriam Weber, and Dirk de Beer. High net primary production of Mediterranean Seagrass (Posidonia oceanica) meadows determined with aquatic eddy covariance. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Guo:2020:NAA

[222] Baoying Guo, Zhongtian Xu, Xiaojun Yan, Isabella Buttino, Jiji Li, Chao Zhou, and Pengzhi Qi. Novel ABCB1 and ABCC transporters are involved in the detoxification of benzo(α)pyrene in thick shell mussel, Mytilus coruscus. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

DOrtenzio:2020:BAT

[223] Fabrizio D’Ortenzio, Vincent Taillandier, Hervé Claustre, Louis Marie Prieur, Edouard Leymarie, Alexandre Mignot, Antoine Poteau, Christophe Penkerc’h, and Catherine Marie Schmechtig. Biogeochemical Argo: The test case of the NAOS Mediterranean Array. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Aschettino:2020:STR

[224] Jessica M. Aschettino, Daniel T. Engelhaupt, Amy G. Engelhaupt, Andrew DiMatteo, Todd Pusser, Michael F. Richlen, and Joel T. Bell. Satellite telemetry reveals spatial overlap between vessel high-traffic areas and humpback whales (Megaptera novaeangliae) near the mouth of the Chesapeake Bay. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Alava:2020:POP

[225] Juan José Alava, Paola Calle, Ana Tirapé, George Biedenbach, Omar Alvarado Cadena, Keith Maruya, Wenjian Lao, Windsor Aguirre, Pedro J. Jiménez, Gustavo A. Domínguez, Gregory D. Bossart, and Patricia A. Fair. Persistent organic pollutants and mercury in genetically identified inner estuary bottlenose dolphin (Tursiops truncatus) residents of the Guayaquil Gulf, Ecuador: Ecotoxicological science in support of pollutant management and cetacean conservation. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Lima:2020:IBS

[226] Mauricio Lima, T. Mariella Canales, Rodrigo Wiff, and José Montero. The interaction between stock dynamics, fishing and climate caused the collapse of the jack mackerel stock at Humboldt Current ecosystem. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Reum:2020:EPF

[227] Jonathan C. P. Reum, Julia L. Blanchard, Kirstin K. Holsman, Kerim Aydin, Anne B. Hollowed, Albert J. Hermann, Wei Cheng, Amanda Faig, Alan C. Haynie, and André E. Punt. Ensemble projections of future climate change impacts on the Eastern Bering Sea food web using a multispecies size spectrum model. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Feitosa:2020:HUP

[228] Leonardo Manir Feitosa, Valderi Dressler, and Rosangela Paula Lessa. Habitat use patterns and identification of essential habitat for an endangered coastal shark with vertebrae microchemistry: The case study of Carcharhinus porosus. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Amaral:2020:DOM

[229] Valentina Amaral, Cristina Romera-Castillo, and Jesús Forja. Dissolved organic matter in the Gulf of Cádiz: Distribution and drivers of chromophoric and fluorescent properties. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Bove:2020:MAR

[230] Colleen B. Bove, James Umbanhowar, and Karl D. Castillo. Metaanalysis reveals reduced coral calcification under projected ocean warming but not under acidification across the Caribbean Sea. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Corona:2020:PAR

[231] Elena Corona, Cecilia Martin, Ramona Marasco, and Carlos M. Duarte. Passive and active removal of marine microplastics by a mushroom coral (Danafungia scruposa). Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Capet:2020:OMC

[232] Arthur Capet, Vicente Fernández, Jun She, Tomasz Dabrowski, Georg Umgiesser, Joanna Staneva, Lőrinc Mészáros, Francisco Campuzano, Laura Ursella, Glenn Nolan, and Ghada El Serafy. Operational modeling capacity in European Seas — an EuroGOOS perspective and recommendations for improvement. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Yucharoen:2020:CDL

[233] Mathinee Yucharoen, Anchalee Chankong, Niphon Phongsuwan, Santi Ninwat, Udomsak Darumas, Suthira Thongkao, and Supaporn Phasombun. Coral diversity at Losin Pinnacle, an offshore reef in the Gulf of Thailand: Toward a future MPA. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Burgos:2020:PDI

[234] Julian M. Burgos, Lene Buhl-Mortensen, Pål Buhl-Mortensen, Steinunn H. Ólafsdóttir, Petur Steingrund, Stefán Á. Ragnarsson, and Øystein Skagseth. Predicting the distribution of indicator taxa of vulnerable marine ecosystems in the Arctic and sub-arctic waters of the Nordic Seas. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Juva:2020:TDC

[235] Katriina Juva, Sascha Flögel, Johannes Karstensen, Peter Linke, and Wolf-Christian Dullo. Tidal dynamics control on cold-water coral growth: a high-resolution multivariable study on Eastern Atlantic cold-water coral sites. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Lian:2020:ATM

[236] Marianne Lian, J. Margaret Castellini, Debra Miller, Brooke Griff, Vladimir V. Vertyankin, Jennifer Dupont, Koen Broker, Céline A. J. Godard-Codding, and Todd M. O’Hara. Assessing δ^{13}C, δ^{15}N and total mercury measures in epidermal biopsies from gray whales. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Klunder:2020:MAE

[237] Lise Klunder, Henko de Stigter, Marc S. S. Lavaleye, Judith D. L. van Bleijswijk, Henk W. van der Veer, Gert-Jan Reichart, and Gerard C. A. Duineveld. A molecular approach to explore the background benthic fauna around a hydrothermal vent and their larvae: Implications for future mining of deep-sea SMS deposits. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Deng:2020:PGI

[238] Lixia Deng, Shunyan Cheung, and Hongbin Liu. Protistal grazers increase grazing on unicellular cyanobacteria diazotroph at night. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Sordo:2020:SPR

[239] Laura Sordo, Rui Santos, Isabel Barrote, Cátia Freitas, and João Silva. Seasonal photosynthesis, respiration, and calcification of a temperate maërl bed in Southern Portugal. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Ankamah-Yeboah:2020:PPD

[240] Isaac Ankamah-Yeboah, Bui Bich Xuan, Stephen Hynes, and Claire W. Armstrong. Public perceptions of deep-sea environment: Evidence from Scotland and Norway. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Lam:2020:CDA

[241] Chi Hin Lam, Clayward Tam, Donald R. Kobayashi, and Molly E. Lutcavage. Complex dispersal of adult yellowfin tuna from the main Hawaiian Islands. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Yang:2020:PPN

[242] Bo Yang, Emmanuel S. Boss, Nils Haëntjens, Matthew C. Long, Michael J. Behrenfeld, Rachel Eveleth, and Scott C. Doney. Phytoplankton phenology in the North Atlantic: Insights from profiling float measurements. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745. See corrigendum [378].

Angove:2020:FCL

[243] Charlotte Angove, Alf Norkko, and Camilla Gustafsson. The fight to capture light: Functional diversity is related to aquatic plant community productivity likely by enhancing light capture. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Recalde-Salas:2020:NSV

[244] Angela Recalde-Salas, Christine Erbe, Chandra Salgado Kent, and Miles Parsons. Non-song vocalizations of humpback whales in Western Australia. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Simonis:2020:SBN

[245] Anne E. Simonis, Karin A. Forney, Shannon Rankin, John Ryan, Yanwu Zhang, Andrew DeVogelaere, John Joseph, Tetyana Margolina, Anna Krumpel, and Simone Baumann-Pickering. Seal bomb noise as a potential threat to Monterey Bay harbor porpoise. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Legge:2020:CNE

[246] Oliver Legge, Martin Johnson, Natalie Hicks, Tim Jickells, Markus Diesing, John Aldridge, Julian Andrews, Yuri Artioli, Dorothee C. E. Bakker, Michael T. Burrows, Nealy Carr, Gemma Cripps, Stacey L. Felgate, Liam Fernand, Naomi Greenwood, Susan Hartman, Silke Kröger, Gennadi Lessin, Claire Mahaffey, Daniel J. Mayor, Ruth Parker, Ana M. Queirós, Jamie D. Shutler, Tiago Silva, Henrik Stahl, Jonathan Tinker, Graham J. C. Underwood, Johan Van Der Molen, Sarah Wakelin, Keith Weston, and Phillip Williamson. Carbon on the Northwest European shelf: Contemporary budget and future influences. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Figueiras:2020:APN

[247] Francisco G. Figueiras, Belén Arbones, Carmen G. Castro, María Froján, and Isabel G. Teixeira. About pigmented nanoflagellates and the importance of mixotrophy in a coastal upwelling system. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Prinz:2020:FFC

[248] Natalie Prinz, Richard Story, Stephen Lyon, Sebastian C. A. Ferse, and Sonia Bejarano. To feed or not to feed? Coral reef fish responses to artificial feeding and stakeholder perceptions in the Aitutaki Lagoon, Cook Islands. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Schellenberg:2020:SSM

[249] Lisa Schellenberg and Laurence J. Clarke. Spatial structure of marine host-associated microbiomes: Effect of taxonomy, species traits, and study design. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Benthuysen:2020:EAU

[250] Jessica A. Benthuysen, Eric C. J. Oliver, Ke Chen, and Thomas Wernberg. Editorial: Advances in understanding marine heatwaves and their impacts. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

DiBenedetto:2020:NBW

[251] Michelle H. DiBenedetto. Non-breaking wave effects on buoyant particle distributions. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Randelhoff:2020:PAO

[252] Achim Randelhoff, Johnna Holding, Markus Janout, Mikael Kristian Sejr, Marcel Babin, Jean-Éric Tremblay, and Matthew B. Alkire. Pan-Arctic Ocean primary production constrained by turbulent nitrate fluxes. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Shih:2020:BVU

[253] Yung-Yen Shih, Chin-Chang Hung, Szu-Yu Huang, François L. L. Muller, and Yu-Hsuan Chen. Biogeochemical variability of the upper ocean response to typhoons and storms in the Northern South China Sea. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Wetmore:2020:CDF

[254] Lynne S. Wetmore, Michael A. Dance, Ronald L. Hill, and Jay R. Rooker. Community dynamics of fish assemblages on mid-shelf and outer-shelf coral reefs in the Northwestern Gulf of Mexico. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Tinker-Kulberg:2020:HCA

[255] Rachel Tinker-Kulberg, Kristen Dellinger, Terry E. Brady, Lee Robertson, Jerrold H. Levy, Sarah K. Abood, Frank M. LaDuca, Christopher L. Kepley, and Anthony L. Dellinger. Horseshoe crab aquaculture as a sustainable endotoxin testing source. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Lopez-Garrido:2020:RVR

[256] Pedro H. López-Garrido, James P. Barry, Juan Ignacio González-Gordillo, and Elva Escobar-Briones. ROV’s video recordings as a tool to estimate variation in megabenthic epifauna diversity and community composition in the Guaymas Basin. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Geange:2020:DIA

[257] Shane W. Geange, Ashley A. Rowden, Simon Nicol, Tiffany Bock, and Martin Cryer. A data-informed approach for identifying move-on encounter thresholds for vulnerable marine ecosystem indicator taxa. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Correia:2020:IMT

[258] Marta Correia, Isabel Costa Azevedo, Helena Peres, Rui Magalhães, Aires Oliva-Teles, Cristina Marisa Ribeiro Almeida, and Laura Guimarães. Integrated multi-trophic aquaculture: a laboratory and hands-on experimental activity to promote environmental sustainability awareness and value of aquaculture products. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Sherman:2020:RSA

[259] Krista D. Sherman, Josephine R. Paris, Robert Andrew King, Karen A. Moore, Craig P. Dahlgren, Lindy C. Knowles, Kristine Stump, Charles R. Tyler, and Jamie R. Stevens. RAD-seq analysis and in situ monitoring of Nassau grouper reveal fine-scale population structure and origins of aggregating fish. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Collins:2020:CMB

[260] Jane Eva Collins, Thomas Vanagt, Isabelle Huys, and Helena Vieira. Corrigendum: Marine bioresource development — stakeholder’s challenges, implementable actions, and business models. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745. See [166].

Darr:2020:DSM

[261] Katherine D. Darr, Jennifer L. East, Sarah Seabrook, Steven J. Dundas, and Andrew R. Thurber. The deep sea and me: Using a science center exhibit to promote lasting public literacy and elucidate public perception of the deep sea. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Terorotua:2020:APC

[262] Heitea Terorotua, Virginie K. E. Duvat, Aurélie Maspataud, and Jehane Ouriqua. Assessing perception of climate change by representatives of public authorities and designing coastal climate services: Lessons learnt from French Polynesia. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Zheng:2020:DSI

[263] Xiaochuan Zheng, Wenbin Liu, Jiadai Liu, Caiyan Zhang, Ling Zhang, Fan Gao, Dingdong Zhang, and Cheng Chi. Dietary supplementation with icariin affects estrogen synthesis, vitellogenesis, and oocyte development in the Chinese mitten crab, Eriocheir sinensis. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Belhabib:2020:EIF

[264] Dyhia Belhabib and Philippe Le Billon. Editorial: Illegal fishing as a trans-national crime. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Muller:2020:SES

[265] Erinn M. Muller, Constance Sartor, Nicholas I. Alcaraz, and Robert van Woesik. Spatial epidemiology of the stony-coral-tissue-loss disease in Florida. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Wang:2020:SAU

[266] Yibang Wang, Yuanchao Wang, Shude Liu, Cui Liang, Hui Zhang, and Weiwei Xian. Stock assessment using LBB method for eight fish species from the Bohai and Yellow Seas. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Atwood:2020:GPM

[267] Trisha B. Atwood, Andrew Witt, Juan Mayorga, Edd Hammill, and Enric Sala. Global patterns in marine sediment carbon stocks. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745. See corrigendum [2226].

Pacheco-Almanzar:2020:DFS

[268] Eloísa Pacheco-Almanzar, Nadia Loza-Estrada, and Ana L. Ibáñez. Do the fish scales shape of Mugil curema reflect the genetic structure using microsatellites markers and the Mexican marine ecoregions classification? Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Medellin-Ortiz:2020:RHR

[269] Alfonso Medellín-Ortiz, Gabriela Montaño-Moctezuma, Carlos Alvarez-Flores, and Eduardo Santamaria del Angel. Retelling the history of the Red Sea urchin fishery in Mexico. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Chalastani:2020:RTD

[270] Vasiliki I. Chalastani, Panos Manetos, Abdulaziz M. Al-Suwailem, Jason A. Hale, Abhishekh P. Vijayan, John Pagano, Ian Williamson, Scott D. Henshaw, Raed Albaseet, Faisal Butt, Russell E. Brainard, Harry Coccossis, Vasiliki K. Tsoukala, and Carlos M. Duarte. Reconciling tourism development and conservation outcomes through marine spatial planning for a Saudi giga-project in the Red Sea (the Red Sea Project, Vision 2030). Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Moore:2020:HPS

[271] Jon A. Moore, Dante B. Fenolio, April B. Cook, and Tracey T. Sutton. Hiding in plain sight: Elopomorph larvae are important contributors to fish biodiversity in a low-latitude oceanic ecosystem. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Frankenbach:2020:SST

[272] Silja Frankenbach, João Ezequiel, Sandra Plecha, Johannes W. Goessling, Leandro Vaz, Michael Kühl, João Miguel Dias, Nuno Vaz, and João Serôdio. Synoptic spatio-temporal variability of the photosynthetic productivity of microphytobenthos and phytoplankton in a tidal estuary. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Baumberger:2020:DGM

[273] Tamara Baumberger, Marvin D. Lilley, John E. Lupton, Edward T. Baker, Joseph A. Resing, Nathaniel J. Buck, Sharon L. Walker, and Gretchen L. Früh-Green. Dissolved gas and metal composition of hydrothermal plumes from a 2008 submarine eruption on the Northeast Lau Spreading Center. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Pootakham:2020:NFL

[274] Wirulda Pootakham, Tanaporn Uengwetwanit, Chutima Sonthirod, Kanchana Sittikankaew, and Nitsara Karoonuthaisiri. A novel full-length transcriptome resource for black tiger shrimp (Penaeus monodon) developed using isoform sequencing (Iso-Seq). Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Blanco-Bercial:2020:MAS

[275] Leocadio Blanco-Bercial. Metabarcoding analyses and seasonality of the zooplankton community at BATS. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Blattmann:2020:LCI

[276] Thomas M. Blattmann, Daniel B. Montluçon, Negar Haghipour, Naoto F. Ishikawa, and Timothy I. Eglinton. Liquid chromatographic isolation of individual amino acids extracted from sediments for radiocarbon analysis. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Gomes:2020:CWQ

[277] Mara Gomes, Alexandre Correia, Lígia Pinto, Carolina Sá, Vanda Brotas, and Marcos Mateus. Coastal water quality in an Atlantic sea bass farm site (Sines, Portugal): a first assessment. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Cullen-knox:2020:TES

[278] Coco Cullen-knox, Aysha Fleming, Libby Lester, and Emily Ogier. Tracing environmental sustainability discourses: an Australia–Asia seafood case study. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Bhadury:2020:UIS

[279] Punyasloke Bhadury and Areen Sen. Understanding impact of seasonal nutrient influx on sedimentary organic carbon and its relationship with Ammonia spp. in a coastal lagoon. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Rotter:2020:NIS

[280] Ana Rotter, Katja Klun, Janja Francé, Patricija Mozetič, and Martina Orlando-Bonaca. Non-indigenous species in the Mediterranean Sea: Turning from pest to source by developing the 8Rs model, a new paradigm in pollution mitigation. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Duarte:2020:HWM

[281] Gustavo A. S. Duarte, Helena D. M. Villela, Matheus Deocleciano, Denise Silva, Adam Barno, Pedro M. Cardoso, Caren L. S. Vilela, Phillipe Rosado, Camila S. M. A. Messias, Maria Alejandra Chacon, Erika P. Santoro, Daniele B. Olmedo, Marcelo Szpilman, Luiz A. Rocha, Michael Sweet, and Raquel S. Peixoto. Heat waves are a major threat to turbid coral reefs in Brazil. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Danobeitia:2020:TCI

[282] Juan Jose Dañobeitia, Sylvie Pouliquen, Truls Johannessen, Alberto Basset, Mathilde Cannat, Benjamin Gerrit Pfeil, Maria Incoronata Fredella, Paola Materia, Claire Gourcuff, Giuseppe Magnifico, Eric Delory, Joaquin del Rio Fernandez, Ivan Rodero, Laura Beranzoli, Ilaria Nardello, Daniele Iudicone, Thierry Carval, Juan M. Gonzalez Aranda, George Petihakis, Jerome Blandin, Werner Leo Kutsch, Janne-Markus Rintala, Andrew R. Gates, and Paolo Favali. Toward a comprehensive and integrated strategy of the European marine research infrastructures for ocean observations. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Kenchington:2020:MBF

[283] Trevor John Kenchington, Daphne E. Themelis, Shannon Colleen De-Vaney, and Ellen Lorraine Kenchington. The meso- and bathypelagic fishes in a large submarine canyon: Assemblage structure of the principal species in the Gully Marine Protected Area. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Ibanez:2020:SPR

[284] Christian M. Ibáñez, Mark Fenwick, Peter A. Ritchie, Sergio A. Carrasco, and M. Cecilia Pardo-Gandarillas. Systematics and phylogenetic relationships of New Zealand benthic octopuses (Cephalopoda: Octopodoidea). Frontiers in Marine Science, 7, March 2020. ISSN 2296-

Matthes:2020:SHK

[285] Lisa C. Matthes, C. J. Mundy, S. L.-Girard, M. Babin, G. Verin, and J. K. Ehn. Spatial heterogeneity as a key variable influencing spring– summer progression in UVR and PAR transmission through Arctic sea ice. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Bekkby:2020:HFT

[286] Trine Bekkby, Nadia Papadopoulou, Dario Fiorentino, Chris J. McOwen, Eli Rinde, Christoffer Boström, Marina Carreiro-Silva, Cristina Linares, Guri Sogn Andersen, Elizabeth Grace Tunka Bengil, Meri Bilan, Emma Cebrian, Carlo Cerrano, Roberto Danovaro, Camilla With Fagerli, Simonetta Fraschetti, Karine Gagnon, Cristina Gambi, Hege Gundersen, Silvija Kipson, Jonne Kotta, Telmo Morato, Henn Ojaveer, Eva Ramirez-Llodra, and Christopher J. Smith. Habitat features and their influence on the restoration potential of marine habitats in Europe. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Basterretxea:2020:POT

[287] G. Basterretxea, J. S. Font-Muñoz, and I. Tuval. Phytoplankton orientation in a turbulent ocean: a microscale perspective. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Steichen:2020:MPC

[288] Jamie L. Steichen, Jessica M. Labonté, Rachel Windham, David Hala, Karl Kaiser, Samantha Setta, Patricia C. Faulkner, Hernando Bacosa, Ge Yan, Manoj Kamalanathan, and Antonietta Quigg. Microbial, physical, and chemical changes in Galveston Bay following an extreme flooding event, Hurricane Harvey. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Williams:2020:TSM

[289] Alan Williams, Franziska Althaus, Mark Green, Kylie Maguire, Candice Untiedt, Nick Mortimer, Chris J. Jackett, Malcolm Clark, Nicholas Bax, Roland Pitcher, and Thomas Schlacher. True size matters for conservation: a robust method to determine the size of deep-sea coral reefs shows they are typically small on seamounts in the Southwest Pacific Ocean. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Ong:2020:ABV

[290] Joyce J. L. Ong, Mark G. Meekan, Hua Hsun Hsu, L. Paul Fanning, and Steven E. Campana. Annual bands in vertebrae validated by bomb radiocarbon assays provide estimates of age and growth of whale sharks. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Zhao:2020:PDP

[291] Xue Zhao, Anne Flore Deton Cabanillas, Alaguraj Veluchamy, Chris Bowler, Fabio Rocha Jimenez Vieira, and Leila Tirichine. Probing the diversity of polycomb and trithorax proteins in cultured and environmentally sampled microalgae. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Lavagnino:2020:GSC

[292] Ana Carolina Lavagnino, Alex Cardoso Bastos, Gilberto Menezes Amado Filho, Fernando Coreixas de Moraes, Lais Silva Araujo, and Rodrigo Leão de Moura. Geomorphometric seabed classification and potential megahabitat distribution in the Amazon Continental Margin. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Mekkawy:2020:SAA

[293] Imam A. Mekkawy, Usama M. Mahmoud, Rehab H. Moneeb, and Alaa El-Din H. Sayed. Significance assessment of Amphora coffeaeformis in arsenic-induced hemato-biochemical alterations of African catfish (Clarias gariepinus). Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Blommaert:2020:LRL

[294] Lander Blommaert, Emmelien Vancaester, Marie J. J. Huysman, Cristina M. Osuna-Cruz, Sofie D’hondt, Johann Lavaud, Bernard Lepetit, Per Winge, Atle M. Bones, Klaas Vandepoele, Wim Vyverman, and Koen Sabbe. Light regulation of LHCX genes in the benthic diatom Seminavis robusta. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Kong:2020:GWI

[295] Fanna Kong, Min Cao, Na Li, Bin Sun, Meijuan Sun, and Yunxiang Mao. Genome-wide identification, phylogeny, and expressional profiles of the mitogen-activated protein kinase kinase kinase (MAPKKK) gene family in Pyropia yezoensis. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Jiang:2020:HMS

[296] Zhijian Jiang, Delian Huang, Yang Fang, Lijun Cui, Chunyu Zhao, Songlin Liu, Yunchao Wu, Qiming Chen, Chanaka Isuranga Premarathne Maha Ranvilage, Jialu He, and Xiaoping Huang. Home for marine species: Seagrass leaves as vital spawning grounds and food source. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Matic-Skoko:2020:FSR

[297] Sanja Matić-Skoko, Melita Peharda, Dario Vrdoljak, Hana Uvanović, and Krešimir Markulin. Fish and sclerochronology research in the Mediterranean: Challenges and opportunities for reconstructing environmental changes. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Han:2020:SVD

[298] Haitao Han, Dawei Pan, Ying Li, Jie Wang, and Chenchen Wang. Stripping voltammetric determination of lead in coastal waters with a functional micro-needle electrode. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Dai:2020:PRM

[299] Min Dai, Songlin Li, Cefeng Fu, Hongjie Qiu, and Naisong Chen. The potential role of marine protein hydrolyzates in elevating nutritive values of diets for largemouth bass, Micropterus salmoides. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Zheng:2020:DII

[300] Jing Zheng, Pei Yang, Jihong Dai, Guijuan Yu, Weihao Ou, Weiqi Xu, Kangsen Mai, and Yanjiao Zhang. Dynamics of intestinal inflammatory cytokines and tight junction proteins of turbot (Scophthalmus maximus L.) during the development and recovery of enteritis induced by dietary β-conglycinin. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Winter:2020:SDE

[301] Gundula Winter, Curt Storlazzi, Sean Vitousek, Ap van Dongeren, Robert McCall, Ron Hoeke, William Skirving, John Marra, Johan Reyns, Jerome Aucan, Matthew Widlansky, Janet Becker, Chris Perry, Gerd Masselink, Ryan Lowe, Murray Ford, Andrew Pomeroy, Fernando Mendez, Ana Rueda, and Moritz Wandres. Steps to develop early warning systems and future scenarios of storm wave-driven flooding along coral reef-lined coasts. Frontiers in Marine Science, 7, March 2020. ISSN 2296-7745.

Spedicato:2020:MCS

[302] Adriana Spedicato, Nuria Sánchez, Lucie Pastor, Lenaick Menot, and Daniela Zeppilli. Meiofauna community in soft sediments at TAG and snake pit hydrothermal vent fields. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Casso:2020:MWI

[303] Maria Casso, Marta Turon, Núria Marco, Marta Pascual, and Xavier Turon. The microbiome of the worldwide invasive ascidian Didemnum vexillum. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Liguori:2020:IWM

[304] Bianca T. P. Liguori, Claudia Ehlert, and Katharina Pahnke. The influence of water mass mixing and particle dissolution on the silicon cycle in the Central Arctic Ocean. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Prins:2020:ELI

[305] Antoine Prins, Paul Deleris, Cédric Hubas, and Bruno Jesus. Effect of light intensity and light quality on diatom behavioral and physiological photoprotection. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Vernet:2020:ECB

[306] Maria Vernet, Jacob Carstensen, Marit Reigstad, and Camilla Svensen. Editorial: Carbon bridge to the Arctic. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Strader:2020:ERD

[307] Marie E. Strader, Logan C. Kozal, Terence S. Leach, Juliet M. Wong, Jannine D. Chamorro, Madeline J. Housh, and Gretchen E. Hofmann. Examining the role of DNA methylation in transcriptomic plasticity of early stage sea urchins: Developmental and maternal effects in a kelp forest herbivore. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Akam:2020:DIC

[308] Sajjad A. Akam, Richard B. Coffin, Hussain A. N. Abdulla, and Timothy W. Lyons. Dissolved inorganic carbon pump in methane-charged shallow marine sediments: State of the art and new model perspectives. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Perez-Tribouillier:2020:QLI

[309] Habacuc Pérez-Tribouillier, Taryn L. Noble, Ashley T. Townsend, Andrew R. Bowie, and Zanna Chase. Quantifying lithogenic inputs to the Southern Ocean using long-lived thorium isotopes. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Iverson:2020:MCT

[310] Autumn R. Iverson, Allison M. Benscoter, Ikuko Fujisaki, Margaret M. Lamont, and Kristen M. Hart. Migration corridors and threats in the Gulf of Mexico and Florida Straits for loggerhead sea turtles. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Waga:2020:COF

[311] Hisatomo Waga and Toru Hirawake. Changing occurrences of fall blooms associated with variations in phytoplankton size structure in the Pacific Arctic. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Mantovani:2020:BPH

[312] Carlo Mantovani, Lorenzo Corgnati, Jochen Horstmann, Anna Rubio, Emma Reyes, Céline Quentin, Simone Cosoli, Jose Luis Asensio, Julien Mader, and Annalisa Griffa. Best practices on high frequency radar deployment and operation for ocean current measurement. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Ibanez:2020:FSI

[313] Ana L. Ibáñez, Ebenezer Guerra, and Eloísa Pacheco-Almanzar. Fish species identification using the rhombic squamation pattern. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Barnett:2020:VMR

[314] Alexandre Barnett, Vona Méléder, Christine Dupuy, and Johann Lavaud. The vertical migratory rhythm of intertidal microphytobenthos in sediment depends on the light photoperiod, intensity, and spectrum: Evidence for a positive effect of blue wavelengths. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Sundahl:2020:DSH

[315] Hanna Sundahl, Pål Buhl-Mortensen, and Lene Buhl-Mortensen. Distribution and suitable habitat of the cold-water corals Lophelia pertusa, Paragorgia arborea, and Primnoa resedaeformis on the Norwegian Continental Shelf. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Culver:2020:ULD

[316] Michelle Culver, James C. Gibeaut, Donna J. Shaver, Philippe Tissot, and Michael Starek. Using lidar data to assess the relationship between beach geomorphology and Kemp’s Ridley (Lepidochelys kempii) nest site selection along Padre Island, TX, United States. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Kramer:2020:PCC

[317] Sasha J. Kramer, David A. Siegel, and Jason R. Graff. Phytoplankton community composition determined from co-variability among phytoplankton pigments from the NAAMES field campaign. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

McClain:2020:MNB

[318] Craig R. McClain, Thomas J. Webb, Clifton C. Nunnally, S. River Dixon, Seth Finnegan, and James A. Nelson. Metabolic niches and biodiversity: a test case in the deep sea benthos. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Ticina:2020:MAI

[319] Vjekoslav Tičina, Ivan Katavić, and Leon Grubišić. Marine aquaculture impacts on marine biota in oligotrophic environments of the Mediterranean Sea — a review. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Boudouresque:2020:IML

[320] Charles-François Boudouresque, Aurélie Blanfuné, Gérard Pergent, Christine Pergent-Martini, Michèle Perret-Boudouresque, and Thierry Thibaut. Impacts of marine and lagoon aquaculture on macrophytes in Mediterranean benthic ecosystems. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Li:2020:NNB

[321] Xiaoshuang Li, Richard G. J. Bellerby, Philip Wallhead, Jianzhong Ge, Jie Liu, Jing Liu, and Anqiang Yang. A neural network-based analysis of the seasonal variability of surface total alkalinity on the East China Sea shelf. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

McLean:2020:ESV

[322] Dianne L. McLean, Miles J. G. Parsons, Andrew R. Gates, Mark C. Benfield, Todd Bond, David J. Booth, Michael Bunce, Ashley M. Fowler, Euan S. Harvey, Peter I. Macreadie, Charitha B. Pattiaratchi, Sally Rouse, Julian C. Partridge, Paul G. Thomson, Victoria L. G. Todd, and Daniel O. B. Jones. Enhancing the scientific value of industry remotely operated vehicles (ROVs) in our oceans. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Stanley:2020:BNS

[323] Thomas R. Stanley, Jeremy M. White, Susan Teel, and Mark Nicholas. Brightness of the night sky affects loggerhead (Caretta caretta) sea turtle hatchling misorientation but not nest site selection. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745. See commentary [808].

Bhadury:2020:AMP

[324] Punyasloke Bhadury and Tarkeshwar Singh. Analysis of marine planktonic cyanobacterial assemblages from Mooriganga Estuary, Indian Sundarbans using molecular approaches. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Seinor:2020:BHF

[325] Kate M. Seinor, Stephen D. A. Smith, Murray Logan, and Steven W. Purcell. Biophysical habitat features explain colonization and size distribution of introduced Trochus (Gastropoda). Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Wiedmann:2020:WFB

[326] Ingrid Wiedmann, Elizaveta Ershova, Bodil Annikki Bluhm, Eva-Maria Nöthig, Rolf R. Gradinger, Ksenia Kosobokova, and Antje Boetius. What feeds the benthos in the Arctic basins? Assembling a carbon budget for the deep Arctic Ocean. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Venkataraman:2020:GDM

[327] Yaamini R. Venkataraman, Alan M. Downey-Wall, Justin Ries, Isaac Westfield, Samuel J. White, Steven B. Roberts, and Kathleen E. Lotterhos. General DNA methylation patterns and environmentally-induced differential methylation in the eastern oyster (Crassostrea virginica). Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Stoeckle:2020:IED

[328] Mark Y. Stoeckle, Mithun Das Mishu, and Zachary Charlop-Powers. Improved environmental DNA reference library detects overlooked marine fishes in New Jersey, United States. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Baetge:2020:NCP

[329] Nicholas Baetge, Jason R. Graff, Michael J. Behrenfeld, and Craig A. Carlson. Net community production, dissolved organic carbon accumulation, and vertical export in the Western North Atlantic. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Liu:2020:PCC

[330] Ling-Li Liu, Jin Sun, Ying Xu, Li-Sheng He, Chandramouli Kondethimmanahalli, and Pei-Yuan Qian. Proteomic comparison of the cytotoxicology of two diastereomers of citreamicin reveals differentially activation of NF-κB pathway. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Campani:2020:EIT

[331] Tommaso Campani, Silvia Casini, Ilaria Caliani, Carlo Pretti, and Maria Cristina Fossi. Ecotoxicological investigation in three model species exposed to elutriates of marine sediments inoculated with bioplastics. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Todd:2020:UVR

[332] Victoria L. G. Todd, Laura Lazar, Laura D. Williamson, Ingrid T. Peters, Aimee L. Hoover, Sophie E. Cox, Ian. B. Todd, Peter I. Macreadie, and Dianne L. McLean. Underwater visual records of marine megafauna around offshore anthropogenic structures. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

He:2020:HPM

[333] Hangqi He, Yifan Li, Shuyuan Wang, Qianwei Ma, and Yiwen Pan. A high precision method for calcium determination in seawater using ion chromatography. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Suh:2020:PEI

[334] David Suh and Robert Pomeroy. Projected economic impact of climate change on marine capture fisheries in the Philippines. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Lauria:2020:SMU

[335] Valentina Lauria, Michele Gristina, Fabio Fiorentino, Martin J. Attrill, and Germana Garofalo. Spatial management units as an ecosystem-based approach for managing bottom-towed fisheries in the Central Mediterranean Sea. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Brandt:2020:AEM

[336] Miriam I. Brandt, Blandine Trouche, Nicolas Henry, Cathy Liautard-Haag, Lois Maignien, Colomban de Vargas, Patrick Wincker, Julie Poulain, Daniela Zeppilli, and Sophie Arnaud-Haond. An assessment of environmental metabarcoding protocols aiming at favoring contemporary biodiversity in inventories of deep-sea communities. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Bracher:2020:HRW

[337] Astrid Bracher, Hongyan Xi, Tilman Dinter, Antoine Mangin, Volker Strass, Wilken-Jon von Appen, and Sonja Wiegmann. High resolution water column phytoplankton composition across the Atlantic Ocean from ship-towed vertical undulating radiometry. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Zheng:2020:TSI

[338] Peng Zheng, Ming Li, Caixia Wang, Judith Wolf, Xueen Chen, Michela De Dominicis, Peng Yao, and Zhan Hu. Tide-surge interaction in the Pearl River Estuary: a case study of Typhoon Hato. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Coleman:2020:RRF

[339] Melinda Ann Coleman, Georgina Wood, Karen Filbee-Dexter, Antoine J. P. Minne, Hugh Douglas Goold, Adriana Vergés, Ezequiel Miguel Marzinelli, Peter David Steinberg, and Thomas Wernberg. Restore or redefine: Future trajectories for restoration. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Lima:2020:PCB

[340] Maria João Lima, Iria Sala, and Rui M. A. Caldeira. Physical connectivity between the NE Atlantic seamounts. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Puerta:2020:IWM

[341] Patricia Puerta, Clare Johnson, Marina Carreiro-Silva, Lea-Anne Henry, Ellen Kenchington, Telmo Morato, Georgios Kazanidis, José Luis Rueda, Javier Urra, Steve Ross, Chih-Lin Wei, José Manuel González-Irusta, Sophie Arnaud-Haond, and Covadonga Orejas. Influence of water masses on the biodiversity and biogeography of deep-sea benthic ecosystems in the North Atlantic. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Izzo:2020:PJS

[342] Christopher Izzo and Bronwyn May Gillanders. Port Jackson shark growth is sensitive to temperature change. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Girard:2020:IEC

[343] Fanny Girard, Jozée Sarrazin, and Karine Olu. Impacts of an eruption on cold-seep microbial and faunal dynamics at a mud volcano. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Pita:2020:URF

[344] Pablo Pita, Manel Antelo, Kieran Hyder, José Vingada, and Sebastián Villasante. The use of recreational fishers’ ecological knowledge to assess the conservation status of marine ecosystems. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Hop:2020:CSI

[345] Haakon Hop, Mikko Vihtakari, Bodil A. Bluhm, Philipp Assmy, Michel Poulin, Rolf Gradinger, Ilka Peeken, Cecilie von Quillfeldt, Lasse Mork Olsen, Ludmila Zhitina, and Igor A. Melnikov. Changes in sea-ice protist diversity with declining sea ice in the Arctic Ocean from the 1980s to 2010s. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Aastrom:2020:CSW

[346] Emmelie K. L. Åström, Arunima Sen, Michael L. Carroll, and JoLynn Carroll. Cold seeps in a warming Arctic: Insights for benthic ecology. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Schoepf:2020:TVM

[347] Verena Schoepf, Maria U. Jung, Malcolm T. McCulloch, Nicole E. White, Michael Stat, and Luke Thomas. Thermally variable, macrotidal reef habitats promote rapid recovery from mass coral bleaching. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745. See corrigendum [804].

Madeira:2020:SEO

[348] Diana Madeira, Joana Andrade, Miguel C. Leal, Violeta Ferreira, Rui J. M. Rocha, Rui Rosa, and Ricardo Calado. Synergistic effects of ocean warming and cyanide poisoning in an ornamental tropical reef fish. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Komaki:2020:OIT

[349] Kanae Komaki and David Fluharty. Options to improve transparency of environmental monitoring governance for polymetallic nodule mining in the area. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Yan:2020:HAD

[350] Ge Yan, Jessica M. Labonté, Antonietta Quigg, and Karl Kaiser. Hurricanes accelerate dissolved organic carbon cycling in coastal ecosystems. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Chiang:2020:SMP

[351] Chun-I Chiang, Ming-Tsung Chung, Jen-Chieh Shiao, Pei-Ling Wang, Tin-Yam Chan, Atsuko Yamaguchi, and Chia-Hui Wang. Seasonal movement patterns of the bigfin reef squid Sepioteuthis lessoniana predicted using statolith δ^{18}O values. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Gaubert-Boussarie:2020:UMA

[352] Julie Gaubert-Boussarie, Soizic Prado, and Cédric Hubas. An untargeted metabolomic approach for microphytobenthic biofilms in intertidal mudflats. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Diwan:2020:CFS

[353] Arvind D. Diwan, Sanjay N. Harke, Gopalkrishna, and Archana N. Panche. Cryobanking of fish and shellfish egg, embryos and larvae: an overview. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Song:2020:SOD

[354] Guisheng Song, Liang Zhao, Fei Chai, Fangfang Liu, Mengting Li, and Huixiang Xie. Summertime oxygen depletion and acidification in Bohai Sea, China. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Ramirez:2020:RVK

[355] Matthew D. Ramirez, Larisa Avens, Lisa R. Goshe, Melissa L. Snover, Melissa Cook, and Selina S. Heppell. Regional variation in Kemp’s Ridley sea turtle diet composition and its potential relationship with somatic growth. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Menden-Deuer:2020:MIA

[356] Susanne Menden-Deuer, Françoise Morison, Amanda L. Montalbano, Gayantonia Franzè, Jacob Strock, Ewelina Rubin, Heather McNair, Colleen Mouw, and Pierre Marrec. Multi-instrument assessment of phytoplankton abundance and cell sizes in mono-specific laboratory cultures and whole plankton community composition in the North Atlantic. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Rocke:2020:MMC

[357] Emma Rocke, Shunyan Cheung, Zimkhita Gebe, Nicole R. Dames, Hongbin Liu, and Coleen L. Moloney. Marine microbial community composition during the upwelling season in the Southern Benguela. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Shen:2020:HCT

[358] Wenying Shen, Yiwei Hu, Zhengkai He, Shuting Xu, Xinyue Xu, Genfang Zhang, and Gang Ren. Histological and comparative transcriptome analyses provide insights into the immune response in pearl sac formation of Hyriopsis cumingii. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Pruzinsky:2020:PHP

[359] Nina M. Pruzinsky, Rosanna J. Milligan, and Tracey T. Sutton. Pelagic habitat partitioning of late-larval and juvenile tunas in the oceanic Gulf of Mexico. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Tripp:2020:OCR

[360] Ashley Tripp, Hannah M. Murphy, and Gail K. Davoren. Otolith chemistry reveals natal region of larval capelin in coastal Newfoundland, Canada. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Zhou:2020:CDH

[361] Feng Zhou, Fei Chai, Daji Huang, Mark Wells, Xiao Ma, Qicheng Meng, Huijie Xue, Jiliang Xuan, Pengbin Wang, Xiaobo Ni, Qiang Zhao, Chenggang Liu, Jilan Su, and Hongliang Li. Coupling and decoupling of high biomass phytoplankton production and hypoxia in a highly dynamic coastal system: The Changjiang (Yangtze River) Estuary. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Tadesse:2020:DDM

[362] M. Tadesse, T. Wahl, and A. Cid. Data-driven modeling of global storm surges. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Mustika:2020:EVS

[363] Putu Liza Kusuma Mustika, Muhammad Ichsan, and Hollie Booth. The economic value of shark and ray tourism in Indonesia and its role in delivering conservation outcomes. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Basili:2020:MRS

[364] Marco Basili, Grazia Marina Quero, Donato Giovannelli, Elena Manini, Carla Vignaroli, Carlo Giacomo Avio, Rocco De Marco, and Gian Marco Luna. Major role of surrounding environment in shaping biofilm community composition on marine plastic debris. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Muis:2020:HRG

[365] Sanne Muis, Maialen Irazoqui Apecechea, Job Dullaart, Joao de Lima Rego, Kristine Skovgaard Madsen, Jian Su, Kun Yan, and Martin Verlaan. A high-resolution global dataset of extreme sea levels, tides, and storm surges, including future projections. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Romero:2020:CTV

[366] Isabel C. Romero, Heather Judkins, and Michael Vecchione. Corrigendum: Temporal variability of polycyclic aromatic hydrocarbons in deepsea cephalopods of the Northern Gulf of Mexico. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745. See [158].

Collins:2020:SPA

[367] Jane Eva Collins, Thomas Vanagt, and Isabelle Huys. Stakeholder perspectives on access and benefit-sharing for areas beyond national jurisdiction. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Schellenberg:2020:BML

[368] Johannes Schellenberg, Jessica Reichert, Martin Hardt, Ines Klingelhöfer, Gertrud Morlock, Patrick Schubert, Mina Bižić, Hans-Peter Grossart, Peter Kämpfer, Thomas Wilke, and Stefanie P. Glaeser. The bacterial microbiome of the long-term aquarium cultured high-microbial abundance sponge Haliclona cnidata — sustained bioactivity despite community shifts under detrimental conditions. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Caruso:2020:MNS

[369] Francesco Caruso, Lijun Dong, Mingli Lin, Mingming Liu, Zining Gong, Wanxue Xu, Giuseppe Alonge, and Songhai Li. Monitoring of a nearshore small dolphin species using passive acoustic platforms and supervised machine learning techniques. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Colefax:2020:AWS

[370] Andrew P. Colefax, Brendan P. Kelaher, Daniel E. Pagendam, and Paul A. Butcher. Assessing white shark (Carcharodon carcharias) behavior along coastal beaches for conservation-focused shark mitigation. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Christ:2020:BBE

[371] Hanna J. Christ, Rachel White, Lincoln Hood, Gabriel M. S. Vianna, and Dirk Zeller. A baseline for the blue economy: Catch and effort history in the Republic of Seychelles’ domestic fisheries. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Tsikopoulou:2020:RMB

[372] Irini Tsikopoulou, Ioulia Santi, Panagiotis D. Dimitriou, Nafsika Papageorgiou, Paraskevi Pitta, and Ioannis Karakassis. Response of microphytobenthos and benthic bacteria viability to eutrophication in a benthic– pelagic coupling mesocosm experiment. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Buhl-Mortensen:2020:CMB

[373] Pål Buhl-Mortensen, Margaret F. J. Dolan, Rebecca E. Ross, Genoveva Gonzalez-Mirelis, Lene Buhl-Mortensen, Lilja Run Bjarnadóttir, and Jon Albretsen. Classification and mapping of benthic biotopes in Arctic and sub-Arctic Norwegian waters. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Rassweiler:2020:RDS

[374] A. Rassweiler, A. K. Dubel, G. Hernan, D. J. Kushner, J. E. Caselle, J. L. Sprague, L. Kui, T. Lamy, S. E. Lester, and R. J. Miller. Roving divers surveying fish in fixed areas capture similar patterns in biogeography but different estimates of density when compared with belt transects. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Maes:2020:YWY

[375] Thomas Maes, Jael van Diemen de Jel, A. Dick Vethaak, Marieke Desender, Victoria A. Bendall, Martin van Velzen, and Heather A. Leslie. You are what you eat, microplastics in porbeagle sharks from the North East Atlantic: Method development and analysis in spiral valve content and tissue. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Bi:2020:FQQ

[376] Rong Bi and Ulrich Sommer. Food quantity and quality interactions at phytoplankton–zooplankton interface: Chemical and reproductive responses in a calanoid copepod. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Ryan-Keogh:2020:DPI

[377] Thomas J. Ryan-Keogh and Sandy J. Thomalla. Deriving a proxy for iron limitation from chlorophyll fluorescence on buoyancy gliders. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Yang:2020:CPP

[378] Bo Yang, Emmanuel S. Boss, Nils Haëntjens, Matthew C. Long, Michael J. Behrenfeld, Rachel Eveleth, and Scott C. Doney. Corrigendum: Phytoplankton phenology in the North Atlantic: Insights from profiling float measurements. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745. See [242].

Franchini:2020:HPS

[379] Filippo Franchini, Sophie Smout, Clint Blight, Lars Boehme, Grant Munro, Marina Costa, and Sonja Heinrich. Habitat partitioning in sympatric delphinids around the Falkland Islands: Predicting distributions based on a limited data set. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Rotter:2020:NNA

[380] Ana Rotter, Ariola Bacu, Michèle Barbier, Francesco Bertoni, Atle M. Bones, M. Leonor Cancela, Jens Carlsson, Maria F. Carvalho, Marta Cegłowska, Meltem Conk Dalay, Thanos Dailianis, Irem Deniz, Dragana Drakulovic, Arita Dubnika, Hjörleifur Einarsson, Ayşegül Erdoğan, Orhan Tufan Eroldoğan, David Ezra, Stefano Fazi, Richard J. FitzGerald, Laura M. Gargan, Susana P. Gaudêncio, Nadica Ivošević DeNardis, Danijela Joksimovic, Marija Kataržytė, Jonne Kotta, Manolis Mandalakis, Inga Matijošytė, Hanna Mazur-Marzec, Alexia Massa-Gallucci, Mohamed Mehiri, Søren Laurentius Nielsen, Lucie Novoveská, Donata Overlingė, Michelle E. Portman, Krzysztof Pyrc, Céline Rebours, Thorsten Reinsch, Fernando Reyes, Baruch Rinkevich, Johan Robbens, Vita Rudovica, Jerica Sabotič, Ivo Safarik, Siret Talve, Deniz Tasdemir, Xenia Theodotou Schneider, Olivier P. Thomas, Anna Toruńska-Sitarz, Giovanna Cristina Varese, and Marlen I. Vasquez. A new network for the advancement of marine biotechnology in Europe and beyond. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Brooker:2020:EAF

[381] Michael A. Brooker, Simon de Lestang, David V. Fairclough, Dianne McLean, Dirk Slawinski, Matthew B. Pember, and Tim J. Langlois. Environmental and anthropogenic factors affect fish abundance: Relationships revealed by automated cameras deployed by fishers. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Tang:2020:CCD

[382] Kai Tang. Corrigendum: Chemical diversity and biochemical transformation of biogenic organic sulfur in the ocean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745. See [171].

Berthold:2020:PDE

[383] Maximilian Berthold and Rhena Schumann. Phosphorus dynamics in a eutrophic lagoon: Uptake and utilization of nutrient pulses by phytoplankton. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Paxton:2020:MAR

[384] Avery B. Paxton, Kyle W. Shertzer, Nathan M. Bacheler, G. Todd Kellison, Kenneth L. Riley, and J. Christopher Taylor. Meta-analysis reveals artificial reefs can be effective tools for fish community enhancement but are not one-size-fits-all. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Hartman:2020:CFG

[385] Karin Hartman, Pieter van der Harst, and Raul Vilela. Continuous focal group follows operated by a drone enable analysis of the relation between sociality and position in a group of male Risso’s dolphins (Grampus griseus). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Ryu:2020:ESW

[386] Taewoo Ryu, Heather D. Veilleux, Philip L. Munday, Imgook Jung, Jennifer M. Donelson, and Timothy Ravasi. An epigenetic signature for within-generational plasticity of a reef fish to ocean warming. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Omeyer:2020:AEB

[387] Lucy C. M. Omeyer, Philip D. Doherty, Sarah Dolman, Robert Enever, Allan Reese, Nicholas Tregenza, Ruth Williams, and Brendan J. Godley. Assessing the effects of banana pingers as a bycatch mitigation device for harbour porpoises (Phocoena phocoena). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Wang:2020:EDI

[388] Cong cong Wang, Wen bin Liu, Xiu fei Cao, Yang yang Huang, Xi Wang, Kang Xiao, Xiang fei Li, Ding dong Zhang, and Guang zhen Jiang. Excess DHA induces cell cycle arrest by activating the P53/cycling pathway in blunt snout bream (Megalobrama amblycephala). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Chen:2020:ISE

[389] Miaoju Chen, Charitha B. Pattiaratchi, Anas Ghadouani, and Christine Hanson. Influence of storm events on chlorophyll distribution along the oligotrophic Continental Shelf off South-Western Australia. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Bashah:2020:CDC

[390] Sara Bashah, Kimberly C. Galvez, Gregor P. Eberli, and Kasey Cantwell. Control of deep currents on sediment and cold-water coral distribution on the Northern Manihiki Plateau. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Degrati:2020:IMT

[391] Mariana Degrati, Rocio Loizaga, Mariano A. Coscarella, Nicolás Sueyro, Enrique A. Crespo, and Silvana L. Dans. Integrating multiple techniques to estimate population size of an impacted dusky dolphin’s population in Patagonia, Argentina. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Thadtapong:2020:GCC

[392] Nalumon Thadtapong, Marvin Bryan Segundo Salinas, Varodom Charoensawan, Vanvimon Saksmerprome, and Soraya Chaturongakul. Genome characterization and comparison of early mortality syndrome causing Vibrio parahaemolyticus pirAB^{vp−} mutant from Thailand with V. parahaemolyticus pirAB^{vp+} AHPND isolates. Frontiers in Marine Science, 7, April 2020. ISSN 2296-7745.

Mabragana:2020:PCS

[393] E. Mabragaña, M. González-Castro, V. Gabbanelli, D. M. Vazquez, and J. M. Díaz de Astarloa. Polymorphism in conservative structures? The scapulocoracoids in skates genus Psammobatis (Chondrichthyes, Arhynchobatidae) and the validity of P. parvacauda. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Schoeman:2020:GRV

[394] Renée P. Schoeman, Claire Patterson-Abrolat, and Stephanie Plön. A global review of vessel collisions with marine animals. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Sorgente:2020:IFI

[395] Roberto Sorgente, Antonia Di Maio, Federica Pessini, Alberto Ribotti, Sergio Bonomo, Angelo Perilli, Ines Alberico, Fabrizio Lirer, Antonio Cascella, and Luciana Ferraro. Impact of freshwater inflow from the Volturno River on coastal circulation. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

White:2020:HRO

[396] Andrew R. White, Maryam Jalali, and Jian Sheng. Hydrodynamics of a rising oil droplet with bacterial extracellular polymeric substance (EPS) streamers using a microfluidic microcosm. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Sass:2020:CTH

[397] Katharina Sass and Mirjam Perner. Characterization of two hydrogenoxidizing hydrogenovibrio strains from Kermadec Volcanic Island Arc hydrothermal vents. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Otto:2020:LCD

[398] Saskia A. Otto, Susa Niiranen, Thorsten Blenckner, Maciej T. Tomczak, Bärbel Müller-Karulis, Gunta Rubene, and Christian Möllmann. Life cycle dynamics of a key marine species under multiple stressors. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Purcell:2020:UGF

[399] Steven W. Purcell, Alejandro Tagliafico, Brian R. Cullis, and Beverley J. Gogel. Understanding gender and factors affecting fishing in an artisanal shellfish fishery. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Zhai:2020:CIP

[400] Lu Zhai and Daniel Pauly. Construction and interpretation of particle size distribution spectra from 19 Ecopath models of Chinese coastal ecosystems. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Szymkowiak:2020:AGG

[401] Marysia Szymkowiak and Melissa Rhodes-Reese. Addressing the gender gap: Using quantitative and qualitative methods to illuminate women’s fisheries participation. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Winters:2020:TSH

[402] Gidon Winters, Sven Beer, Demian A. Willette, Inés G. Viana, Kelcie L. Chiquillo, Pedro Beca-Carretero, Betty Villamayor, Tomás Azcárate-García, Rachamim Shem-Tov, Bridget Mwabvu, Luciana Migliore, Alice Rotini, Michelle A. Oscar, Jonathan Belmaker, Inbal Gamliel, Ana Alexandre, Aschwin H. Engelen, Gabriele Procaccini, and Gil Rilov. The tropical seagrass Halophila stipulacea: Reviewing what we know from its native and invasive habitats, alongside identifying knowledge gaps. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

DuBois:2020:HFI

[403] Morgan J. DuBois, Nathan F. Putman, and Susan E. Piacenza. Hurricane frequency and intensity may decrease dispersal of Kemp’s Ridley sea turtle hatchlings in the Gulf of Mexico. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Blattmann:2020:TAA

[404] Thomas M. Blattmann and Naoto F. Ishikawa. Theoretical amino acidspecific radiocarbon content in the environment: Hypotheses to be tested and opportunities to be taken. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Winship:2020:GPS

[405] Arliss J. Winship, James T. Thorson, M. Elizabeth Clarke, Heather M. Coleman, Bryan Costa, Samuel E. Georgian, David Gillett, Arnaud Grüss, Mark J. Henderson, Thomas F. Hourigan, David D. Huff, Nissa Kreidler, Jodi L. Pirtle, John V. Olson, Matthew Poti, Christopher N. Rooper, Michael F. Sigler, Shay Viehman, and Curt E. Whitmire. Good practices for species distribution modeling of deep-sea corals and sponges for resource management: Data collection, analysis, validation, and communication. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Burns:2020:CDA

[406] John H. R. Burns, Grady Weyenberg, Travis Mandel, Sofia B. Ferreira, Drew Gotshalk, Chad K. Kinoshita, Micah J. Marshall, Nicholas A. V. Del Moral, Shane J. Murphy, Kailey H. Pascoe, Alexandra Runyan, Alexander J. Spengler, Brittany D. Wells, Danielle K. Wilde, and Roberto Pelayo. A comparison of the diagnostic accuracy of in-situ and digital image-based assessments of coral health and disease. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

vanderMheen:2020:DDC

[407] Mirjam van der Mheen, Charitha Pattiaratchi, Simone Cosoli, and Moritz Wandres. Depth-dependent correction for wind-driven drift current in particle tracking applications. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Alves:2020:AEU

[408] Ricardo N. Alves, Asaad H. Mahamed, Jorge F. Alarcon, Abdulaziz Al Suwailem, and Susana Agustí. Adverse effects of ultraviolet radiation on growth, behavior, skin condition, physiology, and immune function in gilthead seabream (Sparus aurata). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Tames-Espinosa:2020:MRS

[409] Mayte Tames-Espinosa, Ico Martínez, Vanesa Romero-Kutzner, Josep Coca, María Algueró-Muñiz, Henriette G. Horn, Andrea Ludwig, Jan Taucher, Lennart Bach, Ulf Riebesell, Theodore T. Packard, and May Gómez. Metabolic responses of subtropical microplankton after a simulated deep-water upwelling event suggest a possible dominance of mixotrophy under increasing CO_{2} levels. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Ruiz:2020:LWS

[410] Irene Ruiz, Oihane C. Basurko, Anna Rubio, Matthias Delpey, Igor Granado, Amandine Declerck, Julien Mader, and Andrés Cózar. Litter windrows in the South-East Coast of the Bay of Biscay: an ocean process enabling effective active fishing for litter. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Bujan:2020:MHN

[411] Nans Bujan and Rónadh Cox. Maximal heights of nearshore storm waves and resultant onshore flow velocities. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Parrish:2020:COC

[412] Frank A. Parrish and Thomas A. Oliver. Comparative observations of current flow, tidal spectra, and scattering strength in and around Hawaiian deep-sea coral patches. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Mejia-Quintero:2020:OAM

[413] Katherine Mejía-Quintero and Luis Chasqui. Octocorals and antipatharians in the mesophotic rocky reefs of Colombian Pacific (Eastern Tropical Pacific). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Frys:2020:FSC

[414] Charles Frys, Antoine Saint-Amand, Matthieu Le Hénaff, Joana Figueiredo, Alyson Kuba, Brian Walker, Jonathan Lambrechts, Valentin Vallaeys, David Vincent, and Emmanuel Hanert. Fine-scale coral connectivity pathways in the Florida Reef Tract: Implications for conservation and restoration. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Huang:2020:APB

[415] Rui-Huan Huang, Wei Lin, Peng Zhang, Jian-Yang Liu, Dan Wang, Yi-Qiang Li, Xiao-Qiang Wang, Cheng-Sheng Zhang, Wei Li, and Dong-Lin Zhao. Anti-phytopathogenic bacterial metabolites from the seaweedderived fungus Aspergillus sp. D40. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Liang:2020:AEFa

[416] Cui Liang, Weiwei Xian, Shude Liu, and Daniel Pauly. Assessments of 14 exploited fish and invertebrate stocks in Chinese waters using the LBB method. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Wilhelm:2020:NBM

[417] Margit R. Wilhelm, Bryan A. Black, Tarron Lamont, Sarah C. Paulus, Chris Bartholomae, and Deon C. Louw. Northern Benguela Merluccius paradoxus annual growth from otolith chronologies used for age verification and as indicators of fisheries-induced and environmental changes. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Gonzalez-Rivas:2020:ENR

[418] David A. González-Rivas, Felipe O. Tapia-Silva, José Jesús Bustillos-Guzmán, Daniel A. Revollo-Fernández, Luis F. Beltrán-Morales, Daniel B. Lluch-Cota, and Alfredo Ortega-Rubio. Estimating nitrogen runoff from agriculture to coastal zones by a rapid GIS and remote sensing-based method for a case study from the irrigation district río mayo, Gulf of California, méxico. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Gough:2020:EOS

[419] Charlotte L. A. Gough, Katrina M. Dewar, Brendan J. Godley, Erude Zafindranosy, and Annette C. Broderick. Evidence of overfishing in smallscale fisheries in Madagascar. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Curnick:2020:IBL

[420] David J. Curnick, Ben Collen, Heather J. Koldewey, Kate E. Jones, Kirsty M. Kemp, and Francesco Ferretti. Interactions between a large Marine Protected Area, pelagic tuna and associated fisheries. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Ashley:2020:CMH

[421] Elizabeth A. Ashley, Jennifer K. Olson, Tessa E. Adler, Stephen Raverty, Eric M. Anderson, Steven Jeffries, and Joseph K. Gaydos. Causes of mortality in a harbor seal (Phoca vitulina) population at equilibrium. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Miller:2020:GRR

[422] Jessica A. Miller and Thomas P. Hurst. Growth rate, ration, and temperature effects on otolith elemental incorporation. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Martinez:2020:TTM

[423] Karla B. Parga Martínez, Mine B. Tekman, and Melanie Bergmann. Temporal trends in marine litter at three stations of the HAUSGARTEN observatory in the Arctic deep sea. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Alonso:2020:GSS

[424] Virginia A. García Alonso, Daniel R. Brown, Marcelo Pájaro, and Fabiana L. Capitanio. Growing up down south: Spatial and temporal variability in early growth of Fuegian sprat Sprattus fuegensis from the Southwest Atlantic Ocean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Osma:2020:RPA

[425] Natalia Osma, Laura Latorre-Melín, Bárbara Jacob, Paulina Y. Contreras, Peter von Dassow, and Cristian A. Vargas. Response of phytoplankton assemblages from naturally acidic coastal ecosystems to elevated pCO_{2}. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Sudheesh:2020:MML

[426] V. Sudheesh, G. V. M. Gupta, and S. W. A. Naqvi. Massive methane loss during seasonal hypoxia/anoxia in the nearshore waters of Southeastern Arabian Sea. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Artika:2020:MPR

[427] Suci Rahmadani Artika, Rohani Ambo-Rappe, Mirta Teichberg, Agustín Moreira-Saporiti, and Inés G. Viana. Morphological and physiological responses of Enhalus acoroides seedlings under varying temperature and nutrient treatment. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

LeCozannet:2020:DCE

[428] Gonéri Le Cozannet, Carlos Oliveros, Olivier Brivois, Audrey Giremus, Manuel Garcin, and Franck Lavigne. Detecting changes in European shoreline evolution trends using Markov chains and the Eurosion Database. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ellison:2020:MRR

[429] Aaron M. Ellison, Alexander J. Felson, and Daniel A. Friess. Mangrove rehabilitation and restoration as experimental adaptive management. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Lester:2020:SPP

[430] Sarah E. Lester, Alexandra K. Dubel, Gema Hernán, Jennifer McHenry, and Andrew Rassweiler. Spatial planning principles for marine ecosystem restoration. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Maggi:2020:ALN

[431] Elena Maggi and João Serôdio. Artificial light at night: a new challenge in microphytobenthos research. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Hart:2020:ING

[432] Kristen M. Hart, Margaret M. Lamont, Autumn R. Iverson, and Brian J. Smith. The importance of the Northeastern Gulf of Mexico to foraging loggerhead sea turtles. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Mohaghar:2020:CBR

[433] Mohammad Mohaghar, Seongyu Jung, Kevin A. Haas, and Donald R. Webster. Copepod behavior responses around internal waves. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Frigstad:2020:IRI

[434] Helene Frigstad, Øyvind Kaste, Anne Deininger, Karsten Kvalsund, Guttorm Christensen, Richard G. J. Bellerby, Kai Sørensen, Marit Norli, and Andrew L. King. Influence of riverine input on Norwegian coastal systems. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Moore:2020:DCB

[435] Michael J. Moore, Glenn H. Mitchell, Teri K. Rowles, and Greg Early. Dead cetacean? Beach, bloat, float, sink. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Furman:2020:WFB

[436] Bradley T. Furman, Bradley J. Peterson, and Kenneth L. Heck. Will the Florida Big Bend Area become the next Gulf of Mexico Reef Tract? Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Miller:2020:EZP

[437] Melissa A. Miller, Pádraig J. Duignan, Erin Dodd, Francesca Batac, Michelle Staedler, Joseph A. Tomoleoni, Michael Murray, Heather Harris, and Chris Gardiner. Emergence of a zoonotic pathogen in a coastal marine sentinel: Capillaria hepatica (syn. Calodium hepaticum)-associated hepatitis in southern sea otters (Enhydra lutris nereis). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Diaz-Delgado:2020:FST

[438] Josué Díaz-Delgado, Kátia R. Groch, Hernani G. C. Ramos, Adriana Castaldo Colosio, Bruna Farias Alves, Hilda F. J. Pena, and José L. Catão-Dias. Fatal systemic toxoplasmosis by a novel non-archetypal Toxoplasma gondii in a Bryde’s whale (Balaenoptera edeni). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Wolny:2020:CFR

[439] Jennifer L. Wolny, Michelle C. Tomlinson, Stephanie Schollaert Uz, Todd A. Egerton, John R. McKay, Andrew Meredith, Kimberly S. Reece, Gail P. Scott, and Richard P. Stumpf. Current and future remote sensing of harmful algal blooms in the Chesapeake Bay to support the shellfish industry. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Eriksen:2020:RWB

[440] Elena Eriksen, Espen Bagøien, Espen Strand, Raul Primicerio, Tatiana Prokhorova, Alexander Trofimov, and Irina Prokopchuk. The recordwarm Barents Sea and 0-group fish response to abnormal conditions. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Jang:2020:MLC

[441] Jiyi Jang, Jiyeon Park, Seohee Ahn, Ki-Tae Park, Sun-Yong Ha, Jongkwan Park, and Kyung Hwa Cho. Molecular-level chemical characterization of dissolved organic matter in the ice shelf systems of King George Island, Antarctica. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Paterson:2020:EPA

[442] Shona K. Paterson, Martin Le Tissier, Hester Whyte, Lisa B. Robinson, Kristin Thielking, Mrill Ingram, and John McCord. Examining the potential of art–science collaborations in the Anthropocene: a case study of catching a wave. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Wagner:2020:SDC

[443] Sasha Wagner, Florence Schubotz, Karl Kaiser, Christian Hallmann, Hannelore Waska, Pamela E. Rossel, Roberta Hansman, Marcus Elvert, Jack J. Middelburg, Anja Engel, Thomas M. Blattmann, Teresa S. Catalá, Sinikka T. Lennartz, Gonzalo V. Gomez-Saez, Silvio Pantoja-Gutiérrez, Rui Bao, and Valier Galy. Soothsaying DOM: a current perspective on the future of oceanic dissolved organic carbon. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Wells:2020:IEB

[444] Brian K. Wells, David D. Huff, Brian J. Burke, Richard D. Brodeur, Jarrod A. Santora, John C. Field, Kate Richerson, Nathan J. Mantua, Kurt L. Fresh, Michelle M. McClure, William H. Satterthwaite, Forbes Darby, Su J. Kim, Richard W. Zabel, and Steven T. Lindley. Implementing ecosystem-based management principles in the design of a salmon ocean ecology program. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Goncalves-Araujo:2020:LDR

[445] Rafael Gonçalves-Araujo and Stiig Markager. Light in the dark: Retrieving underwater irradiance in shallow eutrophic waters from AC-s measurements. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Kapetanou:2020:MIF

[446] Athanasia Kapetanou, Lonneke L. IJsseldijk, Dorien S. Willems, Els M. Broens, Eligius Everaarts, Jochem B. Buil, Paul E. Verweij, Marja J. L. Kik, and Andrea Gröne. Mycotic infections in free-ranging harbor porpoises (Phocoena phocoena). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Thomas:2020:TMU

[447] Jodi T. Thomas, Philip L. Munday, and Sue-Ann Watson. Toward a mechanistic understanding of marine invertebrate behavior at elevated CO_{2}. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Meesters:2020:CMR

[448] Erik H. Meesters, Rolf P. M. Bak, and Fleur C. van Duyl. Commentary: Managing recovery resilience in coral reefs against climate-induced bleaching and hurricanes: a 15 year case study from Bonaire, Dutch Caribbean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745. See [3].

Martinho:2020:WOC

[449] Filipe Martinho, Beatriz Pina, Margarida Nunes, Rita P. Vasconcelos, Vanessa F. Fonseca, Daniel Crespo, Ana Lígia Primo, Ana Vaz, Miguel A. Pardal, Bronwyn M. Gillanders, Susanne E. Tanner, and Patrick Reis-Santos. Water and otolith chemistry: Implications for discerning estuarine nursery habitat use of a juvenile flatfish. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Oyarzun:2020:TMM

[450] Pablo A. Oyarzún, José J. Nuñez, Jorge E. Toro, and Jonathan P. A. Gardner. Trioecy in the marine mussel Semimytilus algosus (Mollusca, Bivalvia): Stable sex ratios across 22 degrees of a latitudinal gradient. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

DAsaro:2020:AOU

[451] Eric A. D’Asaro, Daniel F. Carlson, Marcelo Chamecki, Ramsey R. Harcourt, Brian K. Haus, Baylor Fox-Kemper, M. Jeroen Molemaker, Andrew C. Poje, and Di Yang. Advances in observing and understanding small-scale open ocean circulation during the Gulf of Mexico research initiative era. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Nothig:2020:SCP

[452] Eva-Maria Nöthig, Simon Ramondenc, Antonie Haas, Laura Hehemann, Andreas Walter, Astrid Bracher, Catherine Lalande, Katja Metfies, Ilka Peeken, Eduard Bauerfeind, and Antje Boetius. Summertime chlorophyll a and particulate organic carbon standing stocks in surface waters of the Fram Strait and the Arctic Ocean (1991–2015). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Cheriton:2020:SOW

[453] Olivia M. Cheriton, Curt D. Storlazzi, and Kurt J. Rosenberger. In situ observations of wave transformation and infragravity bore development across reef flats of varying geomorphology. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Schubert:2020:ECA

[454] Nadine Schubert, Kathryn M. Schoenrock, Julio Aguirre, Nicholas A. Kamenos, João Silva, Paulo A. Horta, and Laurie C. Hofmann. Editorial: Coralline algae: Globally distributed ecosystem engineers. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

New:2020:DBW

[455] Leslie New, David Lusseau, and Robert Harcourt. Dolphins and boats: When is a disturbance, disturbing? Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Fentimen:2020:CWC

[456] Robin Fentimen, Eline Feenstra, Andres Rüggeberg, Torsten Vennemann, Irka Hajdas, Thierry Adatte, David Van Rooij, and Anneleen Foubert. Cold-water coral mound archive provides unique insights into intermediate water mass dynamics in the Alboran Sea during the last deglaciation. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Currie:2020:LTC

[457] Jock C. Currie, Lara J. Atkinson, Kerry J. Sink, and Colin G. Attwood. Long-term change of demersal fish assemblages on the inshore Agulhas Bank between 1904 and 2015. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Zhu:2020:NON

[458] You-Ting Zhu, Ya Zhang, Yu-Zhu Liu, Yi-Feng Li, Asami Yoshida, Kiyoshi Osatomi, Jin-Long Yang, and Xiao Liang. Nitric oxide negatively regulates larval metamorphosis in hard-shelled mussel (Mytilus coruscus). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Wassermann:2020:PIS

[459] Sophia N. Wassermann and Mark P. Johnson. The potential to improve the sustainability of pelagic fisheries in the Northeast Atlantic by incorporating individual fish behavior into acoustic sampling. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Kiko:2020:ZMF

[460] Rainer Kiko, Peter Brandt, Svenja Christiansen, Jannik Faustmann, Iris Kriest, Elizandro Rodrigues, Florian Schütte, and Helena Hauss. Zooplankton-mediated fluxes in the eastern tropical North Atlantic. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Lichtenberg:2020:VMO

[461] Mads Lichtenberg, Paulo Cartaxana, and Michael Kühl. Vertical migration optimizes photosynthetic efficiency of motile cyanobacteria in a coastal microbial mat. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Faull:2020:ITD

[462] Luis Medina Faull, Paraskevi Mara, Gordon T. Taylor, and Virginia P. Edgcomb. Imprint of trace dissolved oxygen on prokaryoplankton community structure in an oxygen minimum zone. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Scott:2020:HMC

[463] Fred Scott, Jose A. A. Antolinez, Robert McCall, Curt Storlazzi, Ad Reniers, and Stuart Pearson. Hydro-morphological characterization of coral reefs for wave runup prediction. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Borja:2020:PFG

[464] Angel Borja, Jesper H. Andersen, Christos D. Arvanitidis, Alberto Basset, Lene Buhl-Mortensen, Susana Carvalho, Katherine A. Dafforn, Michelle J. Devlin, Elva G. Escobar-Briones, Christian Grenz, Tilmann Harder, Stelios Katsanevakis, Dongyan Liu, Anna Metaxas, Xosé Anxelu G. Morán, Alice Newton, Chiara Piroddi, Xavier Pochon, Ana M. Queirós, Paul V. R. Snelgrove, Cosimo Solidoro, Michael A. St. John, and Heliana Teixeira. Past and future grand challenges in marine ecosystem ecology. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Jia:2020:EHF

[465] Rui Jia, Li-Ping Cao, Jin-Liang Du, Qin He, Zheng-Yan Gu, Galina Jeney, Pao Xu, and Guo-Jun Yin. Effects of high-fat diet on steatosis, endoplasmic reticulum stress and autophagy in liver of tilapia (Oreochromis niloticus). Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Andradi-Brown:2020:ECR

[466] Dominic A. Andradi-Brown, Anastazia T. Banaszak, Thomas K. Frazer, Hannah Gilchrist, Alastair R. Harborne, Catherine E. I. Head, Heather J. Koldewey, Emma Levy, Kirsty Richards, Rebecca Short, Michael Sweet, Kristian Teleki, Christian R. Voolstra, Bryan Wilson, Elizabeth Wood, Robert T. Yarlett, and David J. Curnick. Editorial: Coral reefs in the Anthropocene — reflecting on 20 years of Reef Conservation UK. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Phillips:2020:WMC

[467] Lachlan R. Phillips, Gemma Carroll, Ian Jonsen, Robert Harcourt, and Moninya Roughan. A water mass classification approach to tracking variability in the East Australian Current. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Zhong:2020:MBP

[468] Kai-Le Zhong, Xiao-Han Song, Han-Gil Choi, Shimada Satoshi, Florian Weinberger, Stefano G. A. Draisma, De-Lin Duan, and Zi-Min Hu. MtDNA-based phylogeography of the red alga Agarophyton vermiculophyllum (Gigartinales, Rhodophyta) in the native Northwest Pacific. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Derous:2020:TNE

[469] Davina Derous, Mariel ten Doeschate, Andrew C. Brownlow, Nicholas J. Davison, and David Lusseau. Toward new ecologically relevant markers of health for cetaceans. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Korotenko:2020:SCT

[470] Konstantin A. Korotenko, Peter O. Zavialov, Yang-Yih Chen, and Hsien Hua Lee. A study of circulation, turbulence, and tidal stream resources in the Taiwan Strait. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Amon:2020:DSD

[471] Diva J. Amon, Brian R. C. Kennedy, Kasey Cantwell, Kelley Suhre, Deborah Glickson, Timothy M. Shank, and Randi D. Rotjan. Deep-sea debris in the Central and Western Pacific Ocean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Cherry:2020:SSD

[472] Taela-Rae Cherry, Shanelle A. Kohler, and Alex T. Ford. Sex specific differences recorded in the behavior of an amphipod: Implications for behavioral toxicology. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Buer:2020:PFM

[473] Anna-Lucia Buer, Marie Maar, Mark Nepf, Lukas Ritzenhofen, Sven Dahlke, René Friedland, Peter Krost, Florian Peine, and Gerald Schernewski. Potential and feasibility of Mytilus spp. farming along a salinity gradient. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Limer:2020:IEC

[474] Benjamin D. Limer, Jeanne Bloomberg, and Daniel M. Holstein. The influence of eddies on coral larval retention in the flower garden banks. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Lin:2020:BPE

[475] Hung-Yun Lin, Wei-Yu Yeh, Sheng-Fang Tsai, Kuo-Ping Chiang, John Han-You Lin, Che-Chia Tsao, and Han-Jia Lin. Biological protective effects against Vibrio infections in grouper larvae using the Strombidium sp. NTOU1, a marine ciliate amenable for scaled-up culture and with an excellent bacteriovorous ability. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Zhang:2020:PGR

[476] Bai-Dong Zhang, Yu-Long Li, Dong-Xiu Xue, and Jin-Xian Liu. Population genomics reveals shallow genetic structure in a connected and ecologically important fish from the Northwestern Pacific Ocean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Schroeder:2020:APE

[477] Katrin Schroeder, Stefano Cozzi, Malek Belgacem, Mireno Borghini, Carolina Cantoni, Sara Durante, Antonio Petrizzo, Annalisa Poiana, and Jacopo Chiggiato. Along-path evolution of biogeochemical and carbonate system properties in the intermediate water of the western mediterranean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Lemonnier:2020:LST

[478] Clarisse Lemonnier, Morgan Perennou, Damien Eveillard, Antonio Fernandez-Guerra, Aude Leynaert, Louis Marié, Hilary G. Morrison, Laurent Memery, Christine Paillard, and Lois Maignien. Linking spatial and temporal dynamic of bacterioplankton communities with ecological strategies across a coastal frontal area. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Lord:2020:MYV

[479] Karina Scavo Lord, Kathryn C. Lesneski, Zachary A. Bengtsson, Kirsten M. Kuhn, Joshua Madin, Beatrice Cheung, Roseline Ewa, Jacob F. Taylor, Elizabeth M. Burmester, Joshua Morey, Les Kaufman, and John R. Finnerty. Multi-year viability of a reef coral population living on mangrove roots suggests an important role for mangroves in the broader habitat mosaic of corals. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Fujii:2020:ESH

[480] Toyonobu Fujii, Daniel J. Pondella, Victoria L. G. Todd, and Andrew Guerin. Editorial: Seafloor heterogeneity: Artificial structures and marine ecosystem dynamics. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Mavraki:2020:RNC

[481] Ninon Mavraki, Ilse De Mesel, Steven Degraer, Tom Moens, and Jan Vanaverbeke. Resource niches of co-occurring invertebrate species at an offshore wind turbine indicate a substantial degree of trophic plasticity. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Bulmer:2020:BCS

[482] Richard Hugh Bulmer, Fabrice Stephenson, Hannah F. E. Jones, Michael Townsend, Jenny R. Hillman, Luitgard Schwendenmann, and Carolyn J. Lundquist. Blue carbon stocks and cross-habitat subsidies. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Pniewski:2020:PCC

[483] Filip Pniewski and Iwona Piasecka-Jedrzejak. Photoacclimation to constant and changing light conditions in a benthic diatom. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Carella:2020:CWO

[484] Francesca Carella, Elisabetta Antuofermo, Simone Farina, Fulvio Salati, Daniela Mandas, Patricia Prado, Rossella Panarese, Fabio Marino, Eleonora Fiocchi, Tobia Pretto, and Gionata De Vico. Corrigendum: In the wake of the ongoing mass mortality events: Co-occurrence of mycobacterium, haplosporidium and other pathogens in Pinna nobilis collected in Italy and Spain (Mediterranean Sea). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745. See [152].

Beauchesne:2020:CES

[485] David Beauchesne, Rémi M. Daigle, Steve Vissault, Dominique Gravel, Andréane Bastien, Simon Bélanger, Pascal Bernatchez, Marjolaine Blais, Hugo Bourdages, Clément Chion, Peter S. Galbraith, Benjamin S. Halpern, Camille Lavoie, Christopher W. McKindsey, Alfonso Mucci, Simon Pineault, Michel Starr, Anne-Sophie Ste-Marie, and Philippe Archambault. Characterizing exposure to and sharing knowledge of drivers of environmental change in the St. Lawrence System in Canada. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Costa:2020:GDS

[486] Corrado Costa, Emanuela Fanelli, Simone Marini, Roberto Danovaro, and Jacopo Aguzzi. Global deep-sea biodiversity research trends highlighted by science mapping approach. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Lopera:2020:CSR

[487] Luisa Lopera, Yuley Cardona, and Paula A. Zapata-Ramírez. Circulation in the Seaflower Reserve and its potential impact on biological connectivity. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Fransner:2020:OBP

[488] Filippa Fransner, François Counillon, Ingo Bethke, Jerry Tjiputra, Annette Samuelsen, Aleksi Nummelin, and Are Olsen. Ocean biogeochemical predictions — initialization and limits of predictability. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Casadevall:2020:SCT

[489] Margarida Casadevall, Conxi Rodríguez-Prieto, Josep Pueyo, Carolina Martí, Roberto Merciai, Marc Verlaque, Enric Real, Jordi Torres, and Jonathan Richir. The strange case of tough white seabream (Diplodus sargus, teleostei: Sparidae): a first approach to the extent of the phenomenon in the Mediterranean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Smith:2020:CSD

[490] Kerri J. Smith, Jed P. Sparks, Zena L. Timmons, and Markus J. Peterson. Cetacean skeletons demonstrate ecologically relevant variation in intraskeletal stable isotopic values. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Raes:2020:FNI

[491] Eric J. Raes, Jodie van de Kamp, Levente Bodrossy, Allison A. Fong, Jessica Riekenberg, Bronwyn H. Holmes, Dirk V. Erler, Bradley D. Eyre, Sarah-Sophie Weil, and A. M. Waite. N_{2} fixation and new insights into nitrification from the ice-edge to the equator in the South Pacific Ocean. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Georgiades:2020:RVB

[492] Eugene Georgiades, Daniel Kluza, Tracey Bates, Katie Lubarsky, Jennie Brunton, Abraham Growcott, Trecia Smith, Simon McDonald, Brendan Gould, Naomi Parker, and Andrew Bell. Regulating vessel biofouling to support New Zealand’s marine biosecurity system — a blue print for evidence-based decision making. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Shi:2020:SCZ

[493] Yongqiang Shi, Jun Wang, Tao Zuo, Xiujuan Shan, Xianshi Jin, Jianqiang Sun, Wei Yuan, and Evgeny A. Pakhomov. Seasonal changes in zooplankton community structure and distribution pattern in the Yellow Sea, China. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Daggers:2020:SSV

[494] Tisja D. Daggers, Peter M. J. Herman, and Daphne van der Wal. Seasonal and spatial variability in patchiness of microphytobenthos on intertidal flats from Sentinel-2 satellite imagery. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Meyer:2020:DMC

[495] Heidi K. Meyer, Emyr M. Roberts, Furu Mienis, and Hans T. Rapp. Drivers of megabenthic community structure in one of the world’s deepest silled-fjords, Sognefjord (Western Norway). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Suda:2020:RRW

[496] Masahiro Suda and Koji Mikami. Reproductive responses to wounding and heat stress in gametophytic thalli of the red alga Pyropia yezoensis. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Culwick:2020:SDD

[497] Timothy Culwick, Jeremy Phillips, Claire Goodwin, Emily J. Rayfield, and Katharine R. Hendry. Sponge density and distribution constrained by fluid forcing in the deep sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

DeHart:2020:PPC

[498] Hayley M. DeHart, Leocadio Blanco-Bercial, Mollie Passacantando, Jennifer M. Questel, and Ann Bucklin. Pathways of pelagic connectivity: Eukrohnia hamata (Chaetognatha) in the Arctic Ocean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Harris:2020:OCI

[499] Catherine A. Harris, Alvaro Lorenzo-Lopez, Owain Jones, Justin J. H. Buck, Alexandra Kokkinaki, Stephen Loch, Thomas Gardner, and Alexander B. Phillips. Oceanids C2: an integrated command, control, and data infrastructure for the over-the-horizon operation of marine autonomous systems. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Allen:2020:NAR

[500] John T. Allen, Cristian Munoz, Jim Gardiner, Krissy A. Reeve, Eva Alou-Font, and Nikolaos Zarokanellos. Near-automatic routine field calibration/correction of glider salinity data using whitespace maximization image analysis of Theta/S data. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ding:2020:PGR

[501] Fangfang Ding, Ao Li, Rihao Cong, Xinxing Wang, Wei Wang, Huayong Que, Guofan Zhang, and Li Li. The phenotypic and the genetic response to the extreme high temperature provides new insight into thermal tolerance for the Pacific oyster Crassostrea gigas. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Foyle:2020:WGH

[502] Kevin L. Foyle, Sybille Hess, Mark D. Powell, and Neill A. Herbert. What is gill health and what is its role in marine finfish aquaculture in the face of a changing climate? Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Magel:2020:CCC

[503] Caitlin L. Magel, Elizabeth M. J. Lee, Astrea M. Strawn, Kelsey Swieca, and Andrew D. Jensen. Connecting crabs, currents, and coastal communities: Examining the impacts of changing ocean conditions on the distribution of U.S. West Coast Dungeness crab commercial catch. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Xue:2020:DPP

[504] Zhaohui Xue, Yue Yu, Wancong Yu, Xin Gao, Yixia Zhang, and Xiaohong Kou. Development prospect and preparation technology of edible oil from microalgae. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Gulliver:2020:EPB

[505] Anne Gulliver, Paul E. Carnell, Stacey M. Trevathan-Tackett, Micheli Duarte de Paula Costa, Pere Masqué, and Peter I. Macreadie. Estimating the potential blue carbon gains from tidal marsh rehabilitation: a case study from South Eastern Australia. Frontiers in Marine Science, 7, May 2020. ISSN 2296-7745.

Rodil:2020:SVB

[506] Iván F. Rodil, Paloma Lucena-Moya, Tobias Tamelander, Joanna Norkko, and Alf Norkko. Seasonal variability in benthic–pelagic coupling: Quantifying organic matter inputs to the seafloor and benthic macrofauna using a multi-marker approach. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Engel:2020:MMM

[507] Anja Engel, Sonja Endres, Luisa Galgani, and Markus Schartau. Marvelous marine microgels: On the distribution and impact of gel-like particles in the oceanic water-column. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Cresswell:2020:ENF

[508] Tom Cresswell, Marc Metian, Nicholas S. Fisher, Sabine Charmasson, Roberta L. Hansman, Wokil Bam, Christian Bock, and Peter Wolfgang Swarzenski. Exploring new frontiers in marine radioisotope tracing — adapting to new opportunities and challenges. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Banker:2020:IDS

[509] Roxanne M. W. Banker and David Coil. Inoculation with Desulfovibrio sp. does not enhance chalk formation in the Pacific oyster. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Mangano:2020:MTS

[510] M. Cristina Mangano, Nova Mieszkowska, Brian Helmuth, Tiago Domingos, Tania Sousa, Giuseppe Baiamonte, Giuseppe Bazan, Angela Cuttitta, Fabio Fiorentino, Antonio Giacoletti, Magnus Johnson, Giuseppe D. Lucido, Marco Marcelli, Riccardo Martellucci, Simone Mirto, Bernardo Patti, Fabio Pranovi, Gray A. Williams, and Gianluca Sarà. Moving toward a strategy for addressing climate displacement of marine resources: a proof-of-concept. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Hetherington:2020:SGE

[511] Elizabeth D. Hetherington and Alexandra A. Phillips. A scientist’s guide for engaging in policy in the United States. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Kumari:2020:NDI

[512] Puja Kumari, Alon Cna’ani, Shoshana Didi-Cohen, Vered Tzin, and Inna Khozin-Goldberg. Nitrogen deprivation-induced production of volatile organic compounds in the arachidonic-acid-accumulating microalga Lobosphaera incisa underpins their role as ROS scavengers and chemical messengers. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

El-Khaled:2020:SMD

[513] Yusuf C. El-Khaled, Florian Roth, Nils Rädecker, Najeh Kharbatia, Burton H. Jones, Christian R. Voolstra, and Christian Wild. Simultaneous measurements of dinitrogen fixation and denitrification associated with coral reef substrates: Advantages and limitations of a combined acetylene assay. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Senez-Mello:2020:SVH

[514] Thaise Machado Senez-Mello, Mírian Araújo Carlos Crapez, Carlos Augusto Ramos e Silva, Eduardo Teixeira da Silva, and Estefan Monteiro da Fonseca. Spatial variability of heavy metals in native oysters from an anthropogenically impacted estuary in Northeastern Brazil. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Batac:2020:HSN

[515] Francesca I. Batac, Melissa A. Miller, Megan E. Moriarty, Zeli Shen, James G. Fox, and Karen M. Ottemann. Helicobacter spp. in necropsied southern sea otters (Enhydra lutris nereis) is associated with gastric ulcers and sensitive to multiple antibiotics. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Horton:2020:MDH

[516] Travis W. Horton, Alexandre N. Zerbini, Artur Andriolo, Daniel Danilewicz, and Federico Sucunza. Multi-decadal humpback whale migratory route fidelity despite oceanographic and geomagnetic change. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Wen:2020:CAP

[517] Jian Wen, Wenlei Wang, Kai Xu, Dehua Ji, Yan Xu, Changsheng Chen, and Chaotian Xie. Comparative analysis of proteins involved in energy metabolism and protein processing in Pyropia haitanensis at different salinity levels. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Sergi:2020:IAC

[518] Sara Sergi, Alberto Baudena, Cédric Cotté, Mathieu Ardyna, Stéphane Blain, and Francesco d’Ovidio. Interaction of the Antarctic Circumpolar Current with seamounts fuels moderate blooms but vast foraging grounds for multiple marine predators. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Feyrer:2020:CLC

[519] Frederick Feyrer, Matthew Young, Darren Fong, Karin Limburg, and Rachel Johnson. Cryptic lives of conspicuous animals: Otolith chemistry chronicles life histories of coastal lagoon fishes. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Li:2020:TAL

[520] Li Li, Fengxiang Gao, Yuxia Jian, Xue Wang, Xiaolong Wang, Lei Pan, Wen Guo, Dong Liu, and Fawen Hu. Transcriptomic analysis of liver tissue in fat greenling (Hexagrammos otakii) exposed to elevated ambient ammonia. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Tolhurst:2020:ESN

[521] Trevor J. Tolhurst, M. G. Chapman, and R. J. Murphy. The effect of shading and nutrient addition on the microphytobenthos, macrofauna, and biogeochemical properties of intertidal flat sediments. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Chu:2020:FCN

[522] Yoo Jung Chu, Hae Sol Kim, Myung Sun Kim, Eun Yeol Lee, and Hee Sook Kim. Functional characterization of a novel oligoalginate lyase of Stenotrophomonas maltophilia KJ-2 using site-specific mutation reveals bifunctional mode of action, possessing both endolytic and exolytic degradation activity toward alginate in seaweed biomass. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Lattuca:2020:FIG

[523] María Eugenia Lattuca, Facundo Llompart, Esteban Avigliano, Marta Renzi, Ileana De Leva, Claudia Clementina Boy, Fabián Alberto Vanella, María Eugenia Barrantes, Daniel Alfredo Fernández, and Cristiano Queiroz de Albuquerque. First insights into the growth and population structure of Cottoperca trigloides (Perciformes, Bovichtidae) from the Southwestern Atlantic Ocean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Grainger:2020:DCN

[524] Richard Grainger, Victor M. Peddemors, David Raubenheimer, and Gabriel E. Machovsky-Capuska. Diet composition and nutritional niche breadth variability in juvenile white sharks (Carcharodon carcharias). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Rohner:2020:NPL

[525] Christoph A. Rohner, Jesse E. M. Cochran, E. Fernando Cagua, Clare E. M. Prebble, Stephanie K. Venables, Michael L. Berumen, Baraka L. Kuguru, Jason Rubens, Juerg M. Brunnschweiler, and Simon J. Pierce. No place like home? High residency and predictable seasonal movement of whale sharks off Tanzania. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Paffrath:2020:REE

[526] Ronja Paffrath, Katharina Pahnke, Melanie K. Behrens, Anja Reckhardt, Claudia Ehlert, Bernhard Schnetger, and Hans-Jürgen Brumsack. Rare earth element behavior in a sandy subterranean estuary of the Southern North Sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Hoagland:2020:SIM

[527] P. Hoagland, A. Beet, D. Ralston, G. Parsons, Y. Shirazi, and E. Carr. Salinity intrusion in a modified river-estuary system: an integrated modeling framework for source-to-sea management. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Kiesel:2020:VBE

[528] Joshua Kiesel, Christina Bienhold, Frank Wenzhöfer, and Heike Link. Variability in benthic ecosystem functioning in Arctic shelf and deep-sea sediments: Assessments by benthic oxygen uptake rates and environmental drivers. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Tang:2020:ESP

[529] Ye Tang, Wei Shi, Dezhi Ning, Jikun You, and Constantine Michailides. Effects of spilling and plunging type breaking waves acting on large monopile offshore wind turbines. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Rossel:2020:MCD

[530] Pamela E. Rossel, Christina Bienhold, Laura Hehemann, Thorsten Dittmar, and Antje Boetius. Molecular composition of dissolved organic matter in sediment porewater of the Arctic deep-sea observatory HAUS-GARTEN (Fram Strait). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ditria:2020:AAF

[531] Ellen M. Ditria, Sebastian Lopez-Marcano, Michael Sievers, Eric L. Jinks, Christopher J. Brown, and Rod M. Connolly. Automating the analysis of fish abundance using object detection: Optimizing animal ecology with deep learning. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ren:2020:ANP

[532] Qing qiang Ren and Min Liu. Assessing Northwest Pacific fishery stocks using two new methods: The Monte Carlo Catch-MSY (CMSY) and the Bayesian Schaefer Model (BSM). Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ross:2020:CDS

[533] Rebecca E. Ross, W. Alex M. Nimmo-Smith, Ricardo Torres, and Kerry L. Howell. Comparing deep-sea larval dispersal models: a cautionary tale for ecology and conservation. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Jin:2020:MOI

[534] Yue Jin, Xianshi Jin, Harry Gorfine, Qiang Wu, and Xiujuan Shan. Modeling the oceanographic impacts on the spatial distribution of common cephalopods during autumn in the Yellow Sea. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

daSilva:2020:TEE

[535] Jorge Marques da Silva, Bernardo Duarte, and Andrei Borissovitch Utkin. Travelling expenses: The energy cost of diel vertical migrations of epipelic microphytobenthos. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Smith:2020:CTL

[536] Carter S. Smith, Morgan E. Rudd, Rachel K. Gittman, Emily C. Melvin, Virginia S. Patterson, Julianna J. Renzi, Emory H. Wellman, and Brian R. Silliman. Coming to terms with living shorelines: a scoping review of novel restoration strategies for shoreline protection. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Eddy:2020:SIE

[537] Corey Eddy, Joanna M. Pitt, Jennifer Larkum, Mark A. Altabet, and Diego Bernal. Stable isotope ecology of invasive lionfish (Pterois volitans and P. miles) in Bermuda. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Hannevig:2020:EED

[538] Hilde Hannevig and Bernt Arne Bertheussen. The exploitationexploration dilemma of fishing vessels with institutionally protected quota shares. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Tamburri:2020:WCC

[539] Mario N. Tamburri, Ian C. Davidson, Matthew R. First, Christopher Scianni, Katherine Newcomer, Graeme J. Inglis, Eugene T. Georgiades, Janet M. Barnes, and Gregory M. Ruiz. In-water cleaning and capture to remove ship biofouling: an initial evaluation of efficacy and environmental safety. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ridge:2020:UAS

[540] Justin T. Ridge and David W. Johnston. Unoccupied aircraft systems (UAS) for marine ecosystem restoration. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Elferink:2020:CMM

[541] Stephanie Elferink, Sylke Wohlrab, Stefan Neuhaus, Allan Cembella, Lars Harms, and Uwe John. Comparative metabarcoding and metatranscriptomic analysis of microeukaryotes within coastal surface waters of West Greenland and Northwest Iceland. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Barrett:2020:PMO

[542] Nick J. Barrett, Raissa I. Hogan, A. Louise Allcock, Tina Molodtsova, Kevin Hopkins, Andrew J. Wheeler, and Chris Yesson. Phylogenetics and mitogenome organisation in black corals (Anthozoa: Hexacorallia: Antipatharia): an order-wide survey inferred from complete mitochondrial genomes. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Orue:2020:SDT

[543] Blanca Orúe, Maria Grazia Pennino, Jon Lopez, Gala Moreno, Josu Santiago, Lourdes Ramos, and Hilario Murua. Seasonal distribution of tuna and non-tuna species associated with drifting fish aggregating devices (DFADs) in the Western Indian Ocean using fishery-independent data. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Crespo-Picazo:2020:NIG

[544] Jose L. Crespo-Picazo, Mariluz Parga, Yara Bernaldo de Quirós, Danielle Monteiro, Vicente Marco-Cabedo, Cristina Llopis-Belenguer, and Daniel García-Párraga. Novel insights into gas embolism in sea turtles: First description in three new species. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Eilertsen:2020:NSO

[545] Mari Heggernes Eilertsen, Thomas G. Dahlgren, and Hans Tore Rapp. A new species of Osedax (Siboglinidae: Annelida) from colonization experiments in the Arctic deep sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Zhang:2020:HOC

[546] Miao Zhang, Ying Wu, Fuqiang Wang, Dongfeng Xu, Sumei Liu, and Meng Zhou. Hotspot of organic carbon export driven by mesoscale eddies in the slope region of the Northern South China Sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Risandi:2020:SVR

[547] Johan Risandi, Jeff E. Hansen, Ryan J. Lowe, and Dirk P. Rijnsdorp. Shoreline variability at a reef-fringed pocket beach. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Kutti:2020:QEM

[548] Tina Kutti, Ingrid Askeland Johnsen, Katrine Sandnes Skaar, Jessica Louise Ray, Vivian Husa, and Thomas G. Dahlgren. Quantification of eDNA to map the distribution of cold-water coral reefs. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Schulze:2020:ARN

[549] Anja Schulze, Deana L. Erdner, Candace J. Grimes, Daniel M. Holstein, and Maria Pia Miglietta. Artificial reefs in the Northern Gulf of Mexico: Community ecology amid the “Ocean sprawl”. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Chang:2020:MAN

[550] Jia Jin Marc Chang, Yin Cheong Aden Ip, Andrew G. Bauman, and Danwei Huang. MinION-in-ARMS: Nanopore sequencing to expedite barcoding of specimen-rich macrofaunal samples from autonomous reef monitoring structures. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Receveur:2020:BMA

[551] Aurore Receveur, Elodie Vourey, Anne Lebourges-Dhaussy, Christophe Menkes, Frédéric Ménard, and Valerie Allain. Biogeography of micronekton assemblages in the Natural Park of the Coral Sea. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ehrnsten:2020:UEC

[552] Eva Ehrnsten, Xiaole Sun, Christoph Humborg, Alf Norkko, Oleg P. Savchuk, Caroline P. Slomp, Karen Timmermann, and Bo G. Gustafsson. Understanding environmental changes in temperate coastal seas: Linking models of benthic fauna to carbon and nutrient fluxes. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Richon:2020:PSD

[553] Camille Richon, Olivier Aumont, and Alessandro Tagliabue. Prey stoichiometry drives iron recycling by zooplankton in the global ocean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ehrlich:2020:SFU

[554] Julia Ehrlich, Fokje L. Schaafsma, Bodil A. Bluhm, Ilka Peeken, Giulia Castellani, Angelika Brandt, and Hauke Flores. Sympagic fauna in and under Arctic pack ice in the annual sea-ice system of the new Arctic. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Gaylard:2020:RCM

[555] Sam Gaylard, Michelle Waycott, and Paul Lavery. Review of coast and marine ecosystems in temperate Australia demonstrates a wealth of ecosystem services. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

DellaPenna:2020:MES

[556] Alice Della Penna and Peter Gaube. Mesoscale eddies structure mesopelagic communities. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Crona:2020:ESS

[557] Beatrice I. Crona, Robert S. Pomeroy, and Steven W. Purcell. Editorial: Small-scale and artisanal fisheries: Insights and approaches for improved governance and management in a globalized context. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Liesner:2020:TPK

[558] Daniel Liesner, Lisa N. S. Shama, Nora Diehl, Klaus Valentin, and Inka Bartsch. Thermal plasticity of the kelp Laminaria digitata (Phaeophyceae) across life cycle stages reveals the importance of cold seasons for marine forests. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Ramalho:2020:EFU

[559] Sofia P. Ramalho, Lidia Lins, Karline Soetaert, Nikolaos Lampadariou, Marina R. Cunha, Ann Vanreusel, and Ellen Pape. Ecosystem functioning under the influence of bottom-trawling disturbance: an experimental approach and field observations from a Continental Slope area in the West Iberian Margin. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Harbour:2020:BDS

[560] Rob P. Harbour, Astrid B. Leitner, Carsten Ruehlemann, Annemiek Vink, and Andrew K. Sweetman. Benthic and demersal scavenger biodiversity in the eastern end of the Clarion–Clipperton Zone — an area marked for polymetallic nodule mining. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Sola:2020:BAE

[561] Iván Sola, Francesc Maynou, and José Luis Sánchez-Lizaso. Bioeconomic analysis of the EU Multiannual Management Plan for demersal fisheries in the Western Mediterranean. Spanish fisheries as a case study. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Long:2020:ISC

[562] Stephen Long, Bridget Sparrow-Scinocca, Martin E. Blicher, Nanette Hammeken Arboe, Mona Fuhrmann, Kirsty M. Kemp, Rasmus Nygaard, Karl Zinglersen, and Chris Yesson. Identification of a soft coral garden candidate Vulnerable Marine Ecosystem (VME) using video imagery, Davis Strait, West Greenland. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Campoy:2020:OCE

[563] Ana Navarro Campoy, Anna Maria Addamo, Annie Machordom, Andrew Meade, Marcelo M. Rivadeneira, Cristián E. Hernández, and Chris Venditti. The origin and correlated evolution of symbiosis and coloniality in scleractinian corals. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Topor:2020:ISS

[564] Zachary M. Topor, Kelly L. Robinson, and Andrea Turcu. Investigating seasonal succession patterns in mesozooplankton community structure following Hurricane Harvey. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Guan:2020:SHL

[565] Chi Guan, Mahasweta Saha, and Florian Weinberger. Simulated heatwaves lead to upregulated chemical defense of a marine foundation macrophyte against microbial colonizers. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Martinez:2020:RGC

[566] Elodie Martinez, Thomas Gorgues, Matthieu Lengaigne, Clement Fontana, Raphaëlle Sauzède, Christophe Menkes, Julia Uitz, Emanuele Di Lorenzo, and Ronan Fablet. Reconstructing global chlorophyll-a variations using a non-linear statistical approach. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745. See corrigendum [1334].

Rossbach:2020:IMP

[567] Susann Rossbach, Ram Chandra Subedi, Tien Khee Ng, Boon S. Ooi, and Carlos M. Duarte. Iridocytes mediate photonic cooperation between giant clams (Tridacninae) and their photosynthetic symbionts. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Lonborg:2020:IGC

[568] Christian Lønborg, Cátia Carreira, Tim Jickells, and Xosé Antón Álvarez-Salgado. Impacts of global change on ocean dissolved organic carbon (DOC) cycling. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Falfushynska:2020:RRP

[569] Halina I. Falfushynska, Eugene Sokolov, Helen Piontkivska, and Inna M. Sokolova. The role of reversible protein phosphorylation in regulation of the mitochondrial electron transport system during hypoxia and reoxygenation stress in marine bivalves. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Barraqueta:2020:FGO

[570] Jan-Lukas Menzel Barraqueta, Saumik Samanta, Eric P. Achterberg, Andrew R. Bowie, Peter Croot, Ryan Cloete, Tara De Jongh, Maria D. Gelado-Caballero, Jessica K. Klar, Rob Middag, Jean C. Loock, Tomas A. Remenyi, Bernhard Wenzel, and Alakendra N. Roychoudhury. A first global oceanic compilation of observational dissolved aluminum data with regional statistical data treatment. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kou:2020:NSD

[571] Qi Kou, Peng Xu, Gary C. B. Poore, Xinzheng Li, and Chunsheng Wang. A new species of the deep-sea sponge-associated genus Eiconaxius (Crustacea: Decapoda: Axiidae), with new insights into the distribution, speciation, and mitogenomic phylogeny of Axiidean shrimps. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kinnby:2020:CEE

[572] Alexandra Kinnby, Per R. Jonsson, Olga Ortega-Martinez, Mats Töpel, Henrik Pavia, Ricardo T. Pereyra, and Kerstin Johannesson. Combining an ecological experiment and a genome scan show idiosyncratic responses to salinity stress in local populations of a seaweed. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Major:2020:ELE

[573] Kaley M. Major, Bethany M. DeCourten, Jie Li, Monica Britton, Matthew L. Settles, Alvine C. Mehinto, Richard E. Connon, and Susanne M. Brander. Early life exposure to environmentally relevant levels of endocrine disruptors drive multigenerational and transgenerational epigenetic changes in a fish model. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Yang:2020:SFD

[574] Wen Yang, Shi-Zhan Zheng, Shou-Heng Zhou, Li Zhao, Jin-Yong Zhu, Betina Lukwambe, Regan Nicholaus, Cheng-Hua Li, and Zhong-Ming Zheng. Structure and functional diversity of surface bacterioplankton communities in an overwintering habitat for large yellow croaker, Pseudosciaena crocea, of the Southern East China Sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Bordoloi:2020:LTS

[575] Ankur D. Bordoloi, Evan Variano, and Gautier Verhille. Lagrangian time scale of passive rotation for mesoscale particles in turbulence. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Sacristan-Soriano:2020:HSD

[576] Oriol Sacristán-Soriano, Natalia Pérez Criado, and Conxita Avila. Host species determines symbiotic community composition in Antarctic sponges (Porifera: Demospongiae). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Bevilacqua:2020:SCB

[577] Stanislao Bevilacqua, Stelios Katsanevakis, Fiorenza Micheli, Enric Sala, Gil Rilov, Gianluca Sarà, Dania Abdul Malak, Ameer Abdulla, Vasilis Gerovasileiou, Elena Gissi, Antonios D. Mazaris, Carlo Pipitone, Maria Sini, Vanessa Stelzenmüller, Antonio Terlizzi, Valentina Todorova, and Simonetta Fraschetti. The status of coastal benthic ecosystems in the Mediterranean Sea: Evidence from ecological indicators. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Markussen:2020:TSA

[578] Thor N. Markussen, Christian Konrad, Christoph Waldmann, Marius Becker, Gerhard Fischer, and Morten H. Iversen. Tracks in the snow — advantage of combining optical methods to characterize marine particles and aggregates. Frontiers in Marine Science, 7, June 2020. ISSN 2296-

Gammage:2020:USD

[579] Louise Carin Gammage and Astrid Jarre. Using structured decisionmaking tools with marginalised fishers to promote system-based fisheries management approaches in South Africa. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Geuer:2020:DDA

[580] Jana K. Geuer, Scarlett Trimborn, Florian Koch, Tina Brenneis, Bernd Krock, and Boris P. Koch. Dissolved domoic acid does not improve growth rates and iron content in iron-stressed Pseudo-Nitzschia subcurvata. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Marriott:2020:ADS

[581] Peter Marriott, Dianne Margaret Tracey, Helen Bostock, Nicholas Hitt, and Stewart J. Fallon. Ageing deep-sea black coral bathypathes patula. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Stark:2020:DAS

[582] Jonathan S. Stark, Mahadi Mohammad, Andrew McMinn, and Jeroen Ingels. Diversity, abundance, spatial variation, and human impacts in marine meiobenthic nematode and copepod communities at Casey Station, East Antarctica. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Martins:2020:ACL

[583] Ivan Machado Martins and Maria A. Gasalla. Adaptive capacity level shapes social vulnerability to climate change of fishing communities in the South Brazil Bight. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Carugati:2020:OSO

[584] Laura Carugati, Alessandro Cau, Maria Cristina Follesa, Riccardo Melis, Davide Moccia, Cristina Porcu, Susanna Salvadori, and Rita Cannas. Onboard scientific observers provide a realistic picture of harvesting and management priorities for the precious red coral (Corallium rubrum L.). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Borgstein:2020:VGA

[585] Natascha Borgstein, Diana M. Beltrán, and Carlos Prada. Variable growth across species and life stages in Caribbean reef octocorals. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Bayraktarov:2020:PMM

[586] Elisa Bayraktarov, Shantala Brisbane, Valerie Hagger, Carter S. Smith, Kerrie A. Wilson, Catherine E. Lovelock, Chris Gillies, Andrew D. L. Steven, and Megan I. Saunders. Priorities and motivations of marine coastal restoration research. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Hemming:2020:DSO

[587] Michael P. Hemming, Moninya Roughan, and Amandine Schaeffer. Daily subsurface ocean temperature climatology using multiple data sources: New methodology. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Bouzaiene:2020:ASD

[588] Maher Bouzaiene, Milena Menna, Pierre-Marie Poulain, Antonio Bussani, and Dalila Elhmaidi. Analysis of the surface dispersion in the Mediterranean sub-basins. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

DeVantier:2020:SRA

[589] Lyndon DeVantier, Emre Turak, and Robert Szava-Kovats. Species richness and abundance of reef-building corals in the Indo-West Pacific: The local–regional relation revisited. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Cordeiro:2020:BCM

[590] Cesar Augusto Marcelino Mendes Cordeiro, Alastair R. Harborne, and Carlos Eduardo Leite Ferreira. The biophysical controls of macroalgal growth on subtropical reefs. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Choi:2020:TDZ

[591] Hyuntae Choi, Sun-Yong Ha, Seunghan Lee, Jee-Hoon Kim, and Kyung-Hoon Shin. Trophic dynamics of zooplankton before and after polar night in the Kongsfjorden (Svalbard): Evidence of trophic position estimated by δ^{15}N analysis of amino acids. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ramirez-Llodra:2020:BCM

[592] Eva Ramirez-Llodra, Ana Hilario, Emil Paulsen, Carolina Ventura Costa, Torkild Bakken, Geir Johnsen, and Hans Tore Rapp. Benthic communities on the Mohn’s Treasure Mound: Implications for management of seabed mining in the Arctic Mid-Ocean Ridge. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Polyakov:2020:BAO

[593] Igor V. Polyakov, Matthew B. Alkire, Bodil A. Bluhm, Kristina A. Brown, Eddy C. Carmack, Melissa Chierici, Seth L. Danielson, Ingrid Ellingsen, Elizaveta A. Ershova, Katarina Gårdfeldt, Randi B. Ingvaldsen, Andrey V. Pnyushkov, Dag Slagstad, and Paul Wassmann. Borealization of the Arctic Ocean in response to anomalous advection from sub-Arctic seas. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745. See corrigendum [1382].

Saviano:2020:HRM

[594] Simona Saviano, Francesco De Leo, Giovanni Besio, Enrico Zambianchi, and Marco Uttieri. HF radar measurements of surface waves in the Gulf of Naples (Southeastern Tyrrhenian Sea): Comparison with hindcast results at different scales. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Schulien:2020:SPC

[595] Jennifer A. Schulien, Alice Della Penna, Peter Gaube, Alison P. Chase, Nils Haëntjens, Jason R. Graff, Johnathan W. Hair, Chris A. Hostetler, Amy Jo Scarino, Emmanuel S. Boss, Lee Karp-Boss, and Michael J. Behrenfeld. Shifts in phytoplankton community structure across an anticyclonic eddy revealed from high spectral resolution lidar scattering measurements. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

McFarland:2020:ELA

[596] Malcolm McFarland, Aditya R. Nayak, Nicole Stockley, Michael Twardowski, and James Sullivan. Enhanced light absorption by horizontally oriented diatom colonies. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Pierrejean:2020:IDW

[597] Marie Pierrejean, Cindy Grant, Bárbara de Moura Neves, Gwénaëlle Chaillou, Evan Edinger, F. Guillaume Blanchet, Frédéric Maps, Christian Nozais, and Philippe Archambault. Influence of deep-water corals and sponge gardens on infaunal community composition and ecosystem functioning in the Eastern Canadian Arctic. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Redzuan:2020:MMS

[598] Nurul S. Redzuan and Graham J. C. Underwood. Movement of microphytobenthos and sediment between mudflats and salt marsh during spring tides. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ramirez-Romero:2020:ASC

[599] Eduardo Ramirez-Romero, Gabriel Jordà, Angel Amores, Susan Kay, Mariona Segura-Noguera, Diego M. Macias, Francesc Maynou, Ana Sabatés, and Ignacio A. Catalán. Assessment of the skill of coupled physical–biogeochemical models in the NW Mediterranean. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Pavaux:2020:CEB

[600] Anne-Sophie Pavaux, Elisa Berdalet, and Rodolphe Lemée. Chemical ecology of the benthic dinoflagellate genus Ostreopsis: Review of progress and future directions. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Borges:2020:SRS

[601] Rebecca Borges, Indira Eyzaguirre, Roberta Sá Leitão Barboza, and Marion Glaser. Systematic review of spatial planning and Marine Protected Areas: a Brazilian perspective. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Kim:2020:RDW

[602] Seung Hyeon Kim, Young Kyun Kim, Hyegwang Kim, and Kun-Seop Lee. Recolonization dynamics of warm affinity Halophila nipponica in a temperate seagrass meadow with Zostera marina. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Perry:2020:TIH

[603] Frances A. Perry, So Kawaguchi, Angus Atkinson, Sévrine F. Sailley, Geraint A. Tarling, Daniel J. Mayor, Cathy H. Lucas, R. King, and A. Cooper. Temperature–induced hatch failure and nauplii malformation in Antarctic krill. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Huang:2020:EWC

[604] Yao Huang, Hongchao Tang, Jianyu Jin, Meng bi Fan, Alan K. Chang, and Xueping Ying. Effects of waterborne cadmium exposure on its internal distribution in Meretrix meretrix and detoxification by metallothionein and antioxidant enzymes. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Epple:2020:NNF

[605] Alexandra L. Epple, Joanna T. Daniel, Susan G. Barco, David S. Rotstein, and Alexander M. Costidis. Novel necropsy findings linked to peracute underwater entrapment in bottlenose dolphins (Tursiops truncatus). Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Song:2020:EFE

[606] Yajuan Song, Yiding Zhao, Xunqiang Yin, Ying Bao, and Fangli Qiao. Evaluation of FIO-ESM v1.0 seasonal prediction skills over the North Pacific. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Bisson:2020:DMO

[607] Kelsey Bisson, David A. Siegel, and Timothy DeVries. Diagnosing mechanisms of ocean carbon export in a satellite-based food web model. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Langenkamper:2020:GIC

[608] Daniel Langenkämper, Robin van Kevelaer, Autun Purser, and Tim W. Nattkemper. Gear-induced concept drift in marine images and its effect on deep learning classification. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Zhang:2020:FSA

[609] Lan lan Zhang, Qing qiang Ren, Min Liu, Qing Xu, Bin Kang, and Xiao bin Jiang. Fishery stock assessments in the Min River Estuary and its adjacent waters in Southern China using the length-based Bayesian estimation (LBB) method. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Colloca:2020:ULE

[610] Francesco Colloca, Virginia Carrozzi, Andrea Simonetti, and Manfredi Di Lorenzo. Using local ecological knowledge of fishers to reconstruct abundance trends of elasmobranch populations in the Strait of Sicily. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Danielson:2020:ORW

[611] Seth L. Danielson, Tyler D. Hennon, Katherine S. Hedstrom, Andrey V. Pnyushkov, Igor V. Polyakov, Eddy Carmack, Kirill Filchuk, Markus Janout, Mikhail Makhotin, William J. Williams, and Laurie Padman. Oceanic routing of wind-sourced energy along the Arctic continental shelves. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Rehren:2020:FTA

[612] Jennifer Rehren and Didier Gascuel. Fishing without a trace? Assessing the balanced harvest approach using EcoTroph. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745. See commentary [1298].

Tai:2020:STV

[613] Jen-Hua Tai, Wen-Chen Chou, Chin-Chang Hung, Kuan-Chieh Wu, Ying-Hsuan Chen, Tzong-Yueh Chen, Gwo-Ching Gong, Fuh-Kwo Shiah, and Chun Hoe Chow. Short-term variability of biological production and CO_{2} system around Dongsha Atoll of the Northern South China Sea: Impact of topography-flow interaction. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Wollschlager:2020:EDT

[614] Jochen Wollschläger, Beke Tietjen, Daniela Voß, and Oliver Zielinski. An empirically derived trimodal parameterization of underwater light in complex coastal waters — a case study in the North Sea. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Chagaris:2020:REM

[615] David D. Chagaris, William F. Patterson, and Micheal S. Allen. Relative effects of multiple stressors on reef food webs in the Northern Gulf of Mexico revealed via ecosystem modeling. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Mies:2020:SAC

[616] Miguel Mies, Ronaldo B. Francini-Filho, Carla Zilberberg, Amana G. Garrido, Guilherme O. Longo, Eduarda Laurentino, Arthur Z. Güth, Paulo Y. G. Sumida, and Thomás N. S. Banha. South Atlantic coral reefs are major global warming refugia and less susceptible to bleaching. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Tripathy:2020:BCV

[617] Sarat C. Tripathy, P. Sabu, Sivaji Patra, Ravidas K. Naik, Amit Sarkar, Vankara Venkataramana, Anvita U. Kerkar, and Pandi Sudarsanarao. Biophysical control on variability in phytoplankton production and composition in the south-western tropical Indian Ocean during monsoon 2014. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ye:2020:GPS

[618] Yingying Ye, Chengrui Yan, Wansuk Senanan, Baoying Guo, Kaida Xu, and Zhenming Lü. Genetic population structure of the hard clam Meretrix meretrix along the Chinese coastlines revealed by microsatellite DNA markers. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Bao:2020:RSO

[619] Rui Bao and Thomas M. Blattmann. Radiocarbonscapes of sedimentary organic carbon in the East Asian Seas. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kharbush:2020:POC

[620] Jenan J. Kharbush, Hilary G. Close, Benjamin A. S. Van Mooy, Carol Arnosti, Rienk H. Smittenberg, Frédéric A. C. Le Moigne, Gesine Mollenhauer, Barbara Scholz-Böttcher, Igor Obreht, Boris P. Koch, Kevin W. Becker, Morten H. Iversen, and Wiebke Mohr. Particulate organic carbon deconstructed: Molecular and chemical composition of particulate organic carbon in the ocean. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Guarnieri:2020:LSS

[621] Giuseppe Guarnieri, Stanislao Bevilacqua, Neus Figueras, Laura Tamburello, and Simonetta Fraschetti. Large-scale sea urchin culling drives the reduction of subtidal barren grounds in the Mediterranean Sea. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Meleder:2020:MIM

[622] Vona Méléder, Raphael Savelli, Alexandre Barnett, Pierre Polsenaere, Pierre Gernez, Philippe Cugier, Astrid Lerouxel, Anthony Le Bris, Christine Dupuy, Vincent Le Fouest, and Johann Lavaud. Mapping the intertidal microphytobenthos gross primary production part i: Coupling multispectral remote sensing and physical modeling. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Savelli:2020:MIM

[623] Raphaël Savelli, Vona Méléder, Philippe Cugier, Pierre Polsenaere, Christine Dupuy, Johann Lavaud, Alexandre Barnett, and Vincent Le Fouest. Mapping the intertidal microphytobenthos gross primary production, part II: Merging remote sensing and physical-biological coupled modeling. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Hedge:2020:PMP

[624] Paul Hedge, Elizabeth Ingrid van Putten, Cass Hunter, and Mibu Fischer. Perceptions, motivations and practices for indigenous engagement in marine science in Australia. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Sumaila:2020:EOI

[625] U. Rashid Sumaila and Travis C. Tai. End overfishing and increase the resilience of the ocean to climate change. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Quero:2020:TCM

[626] Grazia Marina Quero, Francesca Ape, Elena Manini, Simone Mirto, and Gian Marco Luna. Temporal changes in microbial communities beneath fish farm sediments are related to organic enrichment and fish biomass over a production cycle. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Hatcher:2020:ABM

[627] Gerald A. Hatcher, Jonathan A. Warrick, Andrew C. Ritchie, Evan T. Dailey, David G. Zawada, Christine Kranenburg, and Kimberly K. Yates. Accurate bathymetric maps from underwater digital imagery without ground control. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Crutchett:2020:PIS

[628] Thomas Crutchett, Harriet Paterson, Benjamin M. Ford, and Peter Speldewinde. Plastic ingestion in sardines (Sardinops sagax) from Frenchman Bay, Western Australia, highlights a problem in a ubiquitous fish. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Brumley:2020:CTN

[629] Douglas R. Brumley, Francesco Carrara, Andrew M. Hein, George I. Hagstrom, Simon A. Levin, and Roman Stocker. Cutting through the noise: Bacterial chemotaxis in marine microenvironments. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Pardo:2020:ATP

[630] Juan C. F. Pardo, Debra Ramon, Gabriel Stefanelli-Silva, Isa Elegbede, Luciana S. Lima, and Silas C. Principe. Advancing through the pandemic from the perspective of marine graduate researchers: Challenges, solutions, and opportunities. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Tonheim:2020:COM

[631] Susanne Tonheim, Aril Slotte, Leif Andersson, Arild Folkvord, and Florian Berg. Comparison of otolith microstructure of herring larvae and sibling adults reared under identical early life conditions. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Bouchet:2020:SCB

[632] Phil J. Bouchet, Tom B. Letessier, Michael Julian Caley, Scott L. Nichol, Jan M. Hemmi, and Jessica J. Meeuwig. Submerged carbonate banks aggregate pelagic megafauna in offshore tropical Australia. Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Liu:2020:ENE

[633] Cheng-Yue Liu, Fang Zhang, You-Fang Sun, Xiao-Lei Yu, and Hui Huang. Effects of nitrate enrichment on respiration, photosynthesis, and fatty acid composition of reef coral Pocillopora damicornis larvae. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Krause:2020:LLL

[634] Gesche Krause, Christine Wolf, Anne-Kathrin Happe, Jennifer Hauck, Pier Luigi Buttigieg, Nike Fuchs, Jan Scheve, Claudia König, Heidi Wittmer, and Kristina Raab. Lessons learnt from linking global recommendations with localized marine restoration schemes and policy options by using mixed methods. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Srain:2020:FAO

[635] Benjamín M. Srain, Marcus Sobarzo, Giovanni Daneri, Humberto E. González, Giovanni Testa, Laura Farías, Alex Schwarz, Norma Pérez, and Silvio Pantoja-Gutiérrez. Fermentation and anaerobic oxidation of organic carbon in the oxygen minimum zone of the upwelling ecosystem off concepción, in Central Chile. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Liu:2020:PGK

[636] Bingjian Liu, Kun Zhang, Kehua Zhu, Muhammad Shafi, Li Gong, Lihua Jiang, Liqin Liu, Faiz Muhammad, and Zhenming Lü. Population genetics of Konosirus punctatus in Chinese coastal waters inferred from two mtDNA genes (COI and cytb). Frontiers in Marine Science, 7, June 2020. ISSN 2296-7745.

Roelvink:2020:EMC

[637] Dano Roelvink, Bas Huisman, Ahmed Elghandour, Mohamed Ghonim, and Johan Reyns. Efficient modeling of complex sandy coastal evolution at monthly to century time scales. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Castellani:2020:LSV

[638] Giulia Castellani, Fokje L. Schaafsma, Stefanie Arndt, Benjamin A. Lange, Ilka Peeken, Julia Ehrlich, Carmen David, Robert Ricker, Thomas Krumpen, Stefan Hendricks, Sandra Schwegmann, Philippe Massicotte, and Hauke Flores. Large-scale variability of physical and biological seaice properties in polar oceans. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Dankel:2020:MSE

[639] Dorothy J. Dankel, Rachel G. Tiller, Elske Koelma, Vicky W. Y. Lam, and Yajie Liu. The melting snowball effect: a heuristic for sustainable Arctic governance under climate change. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Hoagland:2020:LHF

[640] Porter Hoagland, Barbara Kirkpatrick, Di Jin, Gary Kirkpatrick, Lora E. Fleming, Steven G. Ullmann, Andrew Beet, Gary Hitchcock, Kate K. Harrison, Zongchao C. Li, Bruce Garrison, Roberto E. Diaz, and Vince Lovko. Lessening the hazards of Florida Red Tides: a common sense approach. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kratzer:2020:DOA

[641] Isabella M. F. Kratzer, Ingo Schäfer, Arne Stoltenberg, Jérôme C. Chladek, Lotte Kindt-Larsen, Finn Larsen, and Daniel Stepputtis. Determination of optimal acoustic passive reflectors to reduce bycatch of odontocetes in gillnets. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Shannon:2020:ETV

[642] Lynne J. Shannon, Kelly Ortega-Cisneros, Tarron Lamont, Henning Winker, Robert Crawford, Astrid Jarre, and Marta Coll. Exploring temporal variability in the Southern Benguela ecosystem over the past four decades using a time-dynamic ecosystem model. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ben-Romdhane:2020:CRA

[643] Haïfa Ben-Romdhane, Rima W. Jabado, Edwin Mark Grandcourt, Richard John Obrien Perry, Ayesha Yousef Al Blooshi, Prashanth Reddy Marpu, Taha B. M. J. Ouarda, and Hosni Ghedira. Coral reefs of Abu Dhabi, United Arab Emirates: Analysis of management approaches in light of international best practices and a changing climate. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Bamunawala:2020:HMA

[644] Janaka Bamunawala, Ali Dastgheib, Roshanka Ranasinghe, Ad van der Spek, Shreedhar Maskey, A. Brad Murray, Trang Minh Duong, Patrick L.

Barnard, and T. A. J. G. Sirisena. A holistic modeling approach to project the evolution of inlet-interrupted coastlines over the 21st Century. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Mau:2020:MSI

[645] Susan Mau, Tzu-Hsuan Tu, Marius Becker, Christian dos Santos Ferreira, Jhen-Nien Chen, Li-Hung Lin, Pei-Ling Wang, Saulwood Lin, and Gerhard Bohrmann. Methane seeps and independent methane plumes in the South China Sea offshore Taiwan. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Eger:2020:PPU

[646] Aaron M. Eger, Ezequiel Marzinelli, Paul Gribben, Craig R. Johnson, Cayne Layton, Peter D. Steinberg, Georgina Wood, Brian R. Silliman, and Adriana Vergés. Playing to the positives: Using synergies to enhance kelp forest restoration. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kiko:2020:EZN

[647] Rainer Kiko, Daniele Bianchi, Christian Grenz, Helena Hauss, Morten Iversen, Sanjeev Kumar, Amy Maas, and Carol Robinson. Editorial: Zooplankton and nekton: Gatekeepers of the biological pump. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Wandres:2020:DSS

[648] Moritz Wandres, Jérôme Aucan, Antonio Espejo, Naomi Jackson, Antoine De Ramon N’Yeurt, and Hervé Damlamian. Distant-source swells cause coastal inundation on Fiji’s Coral Coast. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Grose:2020:CCW

[649] Susan O. Grose, Linwood Pendleton, Amanda Leathers, Andrew Cornish, and Sheridan Waitai. Climate change will re-draw the map for marine megafauna and the people who depend on them. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Hall:2020:AUS

[650] Emily R. Hall, Leslie Wickes, Louis E. Burnett, Geoffrey I. Scott, Debra Hernandez, Kimberly K. Yates, Leticia Barbero, Janet J. Reimer, Mohammed Baalousha, Jennifer Mintz, Wei-Jun Cai, J. Kevin Craig, M. Richard DeVoe, William S. Fisher, Terri K. Hathaway, Elizabeth B. Jewett, Zackary Johnson, Paula Keener, Rua S. Mordecai, Scott Noakes, Charlie Phillips, Paul A. Sandifer, Astrid Schnetzer, and Jay Styron.

Acidification in the U.S. Southeast: Causes, potential consequences and the role of the Southeast Ocean and Coastal Acidification Network. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Anugerahanti:2020:PBP

[651] Prima Anugerahanti, Shovonlal Roy, and Keith Haines. Perturbed biology and physics signatures in a 1-D ocean biogeochemical model ensemble. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Christie:2020:DBM

[652] Hartvig Christie, Patrik Kraufvelin, Lucinda Kraufvelin, Niklas Niemi, and Eli Rinde. Disappearing blue mussels — can mesopredators be blamed? Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Fischer:2020:OCU

[653] Philipp Fischer, Holger Brix, Burkard Baschek, Alexandra Kraberg, Markus Brand, Boris Cisewski, Rolf Riethmüller, Gisbert Breitbach, Klas Ove Möller, Jean-Pierre Gattuso, Samir Alliouane, Willem H. van de Poll, and Rob Witbaard. Operating cabled underwater observatories in rough shelf-sea environments: a technological challenge. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Easson:2020:CEA

[654] Cole G. Easson, Kevin M. Boswell, Nicholas Tucker, Joseph D. Warren, and Jose V. Lopez. Combined eDNA and acoustic analysis reflects diel vertical migration of mixed consortia in the Gulf of Mexico. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Wang:2020:SVT

[655] Hui-Yu Wang, Ling Ming Tsang, Fernando P. Lima, Rui Seabra, Monthon Ganmanee, Gray A. Williams, and Benny K. K. Chan. Spatial variation in thermal stress experienced by barnacles on rocky shores: The interplay between geographic variation, tidal cycles and microhabitat temperatures. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Asaduzzaman:2020:MGD

[656] Md Asaduzzaman, Md Abdul Wahab, Md Moshiur Rahman, Mariom, Md Nahiduzzaman, Md Jalilur Rahman, Barun Kanti Roy, Michael J. Phillips, and Li Lian Wong. Morpho-genetic divergence and adaptation of anadromous hilsa shad (Tenualosa ilisha) along their heterogenic migratory habitats. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Huggins:2020:EMF

[657] Jessica L. Huggins, Michael M. Garner, Stephen A. Raverty, Dyanna M. Lambourn, Stephanie A. Norman, Linda D. Rhodes, Joseph K. Gaydos, Jennifer K. Olson, Martin Haulena, and M. Bradley Hanson. The emergence of mucormycosis in free-ranging marine mammals of the Pacific northwest. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Livne:2020:PIO

[658] Leigh Livne, Michal Grossowicz, Dan Tchernov, and Ofira Ayalon. Predicting impacts of offshore monoculture farm expansion in ultraoligotrophic waters of the Levantine Basin. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Dabrowska:2020:PPE

[659] Anna Maria Dabrowska, Józef Mikołaj Wiktor, Małgorzata Merchel, and Józef Maria Wiktor. Planktonic protists of the Eastern Nordic Seas and the Fram Strait: Spatial changes related to hydrography during early summer. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Peel:2020:RMR

[660] Lauren R. Peel, Guy M. W. Stevens, Ryan Daly, Clare A. Keating Daly, Shaun P. Collin, Josep Nogués, and Mark G. Meekan. Regional movements of reef manta rays (Mobula alfredi) in Seychelles waters. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Gredzens:2020:STC

[661] Christian Gredzens and Donna J. Shaver. Satellite tracking can inform population-level dispersal to foraging grounds of post-nesting Kemp’s Ridley sea turtles. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ren:2020:CRP

[662] Shengjie Ren, Peter B. Mather, Binguo Tang, and David A. Hurwood. Comparison of reproductive performance of domesticated Litopenaeus vannamei females reared in recirculating tanks and earthen ponds: an evaluation of reproductive quality of spawns in relation to female body size and spawning order. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Brooke:2020:MRP

[663] Charles G. Brooke, Breanna M. Roque, Claire Shaw, Negeen Najafi, Maria Gonzalez, Abigail Pfefferlen, Vannesa De Anda, David W. Ginsburg, Maddelyn C. Harden, Sergey V. Nuzhdin, Joan King Salwen, Ermias Kebreab, and Matthias Hess. Methane reduction potential of two Pacific coast macroalgae during in vitro ruminant fermentation. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kromkamp:2020:MVB

[664] Jacco C. Kromkamp, Ed Morris, and Rodney M. Forster. Microscale variability in biomass and photosynthetic activity of microphytobenthos during a spring-neap tidal cycle. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Scheinin:2020:UPC

[665] Matias Scheinin and Eero Asmala. Ubiquitous patchiness in chlorophylla concentration in coastal archipelago of Baltic Sea. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Giering:2020:IPS

[666] Sarah L. C. Giering, Brett Hosking, Nathan Briggs, and Morten H. Iversen. The interpretation of particle size, shape, and carbon flux of marine particle images is strongly affected by the choice of particle detection algorithm. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Svedang:2020:CFE

[667] Henrik Svedäng, Viktor Thunell, Ale Pålsson, Sofia A. Wikström, and Martin J. Whitehouse. Compensatory feeding in Eastern Baltic cod (Gadus morhua): Recent shifts in otolith growth and nitrogen content suggest unprecedented metabolic changes. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Auster:2020:SBD

[668] Peter J. Auster, Brooke C. Hodge, Michael P. McKee, and Scott D. Kraus. A scientific basis for designation of the Northeast Canyons and Seamounts Marine National Monument. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Coppin:2020:IWT

[669] Ross Coppin, Christo Rautenbach, Timothy John Ponton, and A. J. Smit. Investigating waves and temperature as drivers of kelp morphology. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Gonzalez-Santana:2020:PDI

[670] David González-Santana, Hélène Planquette, Marie Cheize, Hannah Whitby, Arthur Gourain, Thomas Holmes, Vivien Guyader, Cécile Cathalot, Ewan Pelleter, Yves Fouquet, and Géraldine Sarthou. Processes driving iron and manganese dispersal from the TAG hydrothermal plume (Mid-Atlantic Ridge): Results from a GEOTRACES Process Study. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Tarazona-Janampa:2020:ABT

[671] Ulrike I. Tarazona-Janampa, Allan D. Cembella, María C. Pelayo-Zárate, Silvia Pajares, Laura M. Márquez-Valdelamar, Yuri B. Okolodkov, Jan Tebben, Bernd Krock, and Lorena M. Durán-Riveroll. Associated bacteria and their effects on growth and toxigenicity of the dinoflagellate Prorocentrum lima species complex from epibenthic substrates along Mexican coasts. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Birkmanis:2020:FDS

[672] Charlotte A. Birkmanis, Jennifer J. Freer, Leigh W. Simmons, Julian C. Partridge, and Ana M. M. Sequeira. Future distribution of suitable habitat for pelagic sharks in Australia under climate change models. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Gupta:2020:MEB

[673] Trisha Gupta, Hollie Booth, William Arlidge, Chetan Rao, Muralidharan Manoharakrishnan, Naveen Namboothri, Kartik Shanker, and E. J. Milner-Gulland. Mitigation of elasmobranch bycatch in trawlers: a case study in Indian fisheries. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Grego:2020:DCI

[674] Mateja Grego, Alenka Malej, and Marleen De Troch. The depleted carbon isotopic signature of nematodes and harpacticoids and their place in carbon processing in fish farm sediments. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ibrahim:2020:MSE

[675] Hamed D. Ibrahim, Pengfei Xue, and Elfatih A. B. Eltahir. Multiple salinity equilibria and resilience of Persian/Arabian Gulf Basin salinity to brine discharge. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Abdolahpour:2020:MRT

[676] Maryam Abdolahpour, Marco Ghisalberti, Kathryn McMahon, and Paul Lavery. Material residence time in marine canopies under wave-driven flows. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Montes:2020:DSS

[677] Enrique Montes, Anni Djurhuus, Frank E. Muller-Karger, Daniel Otis, Christopher R. Kelble, and Maria T. Kavanaugh. Dynamic satellite seascapes as a biogeographic framework for understanding phytoplankton assemblages in the Florida Keys National Marine Sanctuary, United States. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Wisz:2020:OMI

[678] Mary S. Wisz, Erin V. Satterthwaite, Maree Fudge, Mibu Fischer, Andrei Polejack, Michael St. John, Stephen Fletcher, and Murray A. Rudd. 100 opportunities for more inclusive ocean research: Cross-disciplinary research questions for sustainable ocean governance and management. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Raykov:2020:EEA

[679] Violin Stoyanov Raykov, Petya Pavlova Ivanova, Cemal Turan, and Ertug Duzgunes. Editorial: Ecosystem approach to fisheries in the Mediterranean and Black Seas - advances in research and technologies. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Rios:2020:CCH

[680] Pilar Ríos, Elena Prado, Francisca C. Carvalho, Francisco Sánchez, Augusto Rodríguez-Basalo, Joana R. Xavier, Teodoro P. Ibarrola, and Javier Cristobo. Community composition and habitat characterization of a rock sponge aggregation (Porifera, Corallistidae) in the Cantabrian Sea. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Meinert:2020:TRD

[681] Corinne R. Meinert, Kimberly Clausen-Sparks, Maëlle Cornic, Tracey T. Sutton, and Jay R. Rooker. Taxonomic richness and diversity of larval fish assemblages in the oceanic Gulf of Mexico: Links to oceanographic conditions. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Bednarsek:2020:CED

[682] Nina Bednaršek, Greg Pelletier, Anise Ahmed, and Richard A. Feely. Chemical exposure due to anthropogenic ocean acidification increases risks for estuarine calcifiers in the Salish Sea: Biogeochemical model scenarios. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Henley:2020:CBS

[683] Sian F. Henley, Emma L. Cavan, Sarah E. Fawcett, Rodrigo Kerr, Thiago Monteiro, Robert M. Sherrell, Andrew R. Bowie, Philip W. Boyd, David K. A. Barnes, Irene R. Schloss, Tanya Marshall, Raquel Flynn, and Shantelle Smith. Changing biogeochemistry of the Southern Ocean and its ecosystem implications. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Miller:2020:PDE

[684] Melissa A. Miller, Megan E. Moriarty, Laird Henkel, Martin Tim Tinker, Tristan L. Burgess, Francesca I. Batac, Erin Dodd, Colleen Young, Michael D. Harris, David A. Jessup, Jack Ames, Patricia A. Conrad, Andrea E. Packham, and Christine K. Johnson. Predators, disease, and environmental change in the nearshore ecosystem: Mortality in southern sea otters (Enhydra lutris nereis) from 1998–2012. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Cvitanovic:2020:EME

[685] Christopher Cvitanovic, E. Ingrid van Putten, Rachel Kelly, Hannah R. Feldman, Tatiana J. van Steveninck, Mary Mackay, Nicholas Badullovich, and Tegan Gourlay. Engaging more effectively with visitors to coastal regions for improved management outcomes: Insights from the Ningaloo Coast, Australia. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Szeligowska:2020:SPP

[686] Marlena Szeligowska, Emilia Trudnowska, Rafał Boehnke, Anna Maria Dabrowska, Józef Maria Wiktor, Sławomir Sagan, and Katarzyna Błachowiak-Samołyk. Spatial patterns of particles and plankton in the warming Arctic fjord (Isfjorden, West Spitsbergen) in seven consecutive mid-summers (2013–2019). Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Tsuchiya:2020:AME

[687] Masashi Tsuchiya, Seiji Miyawaki, Kazumasa Oguri, Takashi Toyofuku, Akihiro Tame, Katsuyuki Uematsu, Koji Takeda, Yuya Sakai, Hiroshi Miyake, and Tadashi Maruyama. Acquisition, maintenance, and ecological roles of kleptoplasts in Planoglabratella opercularis (Foraminifera, Rhizaria). Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Okafor-Yarwood:2020:BEC

[688] Ifesinachi Okafor-Yarwood, Nelly I. Kadagi, Nelson A. F. Miranda, Jacqueline Uku, Isa O. Elegbede, and Ibukun J. Adewumi. The blue economy–cultural livelihood–ecosystem conservation triangle: The African experience. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Montefalcone:2020:ILP

[689] Monica Montefalcone, Carla Morri, and Carlo Nike Bianchi. Influence of local pressures on Maldivian coral reef resilience following repeated bleaching events, and recovery perspectives. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Hunt:2020:QMS

[690] Corallie Hunt, Urška Demšar, Dayton Dove, Craig Smeaton, Rhys Cooper, and William E. N. Austin. Quantifying marine sedimentary carbon: a new spatial analysis approach using seafloor acoustics, imagery, and ground-truthing data in Scotland. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Muhling:2020:PSD

[691] Barbara A. Muhling, Stephanie Brodie, James A. Smith, Desiree Tommasi, Carlos F. Gaitan, Elliott L. Hazen, Michael G. Jacox, Toby D. Auth, and Richard D. Brodeur. Predictability of species distributions deteriorates under novel environmental conditions in the California Current system. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Steinberg:2020:KTA

[692] Rosemary Kate Steinberg, Katherine A. Dafforn, Tracy Ainsworth, and Emma L. Johnston. Know thy anemone: a review of threats to octocorals and anemones and opportunities for their restoration. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Liao:2020:SFD

[693] Jian-Xiang Liao, Chih-Lin Wei, and Moriaki Yasuhara. Species and functional diversity of deep-sea nematodes in a high energy submarine canyon. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Scrosati:2020:LSC

[694] Ricardo A. Scrosati and Julius A. Ellrich. Latitudinal and seasonal changes in intertidal sea surface temperature along the Atlantic Coast of Nova Scotia, Canada. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Takagi:2020:AAG

[695] Satomi Takagi, Yuko Murata, Tomohiko Koshiishi, and Yukio Agatsuma. The amino acids glutamic acid and alanine in feed increase the alanine content in gonads of the sea urchin Mesocentrotus nudus. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Li:2020:RCD

[696] Pingyang Li and Toste Tanhua. Recent changes in deep ventilation of the Mediterranean Sea; evidence from long-term transient tracer observations. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Godard:2020:DBM

[697] Morgan Godard, Claude Manté, Christophe Guinet, Baptiste Picard, and David Nerini. Diving behavior of mirounga leonina: a functional data analysis approach. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Pearson:2020:UPT

[698] Ryan M. Pearson, Jason P. van de Merwe, Michael K. Gagan, and Rod M. Connolly. Unique post-telemetry recapture enables development of multielement isoscapes from barnacle shell for retracing host movement. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Anton:2020:SMT

[699] Andrea Anton, Hanan Almahasheer, Antonio Delgado, Neus Garcias-Bonet, Paloma Carrillo de Albornoz, Núria Marbà, Iris Eline Hendriks, Dorte Krause-Jensen, Vincent Saderne, Kimberlee Baldry, and Carlos M. Duarte. Stunted mangrove trees in the oligotrophic Central Red Sea relate to nitrogen limitation. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Johnson:2020:CET

[700] Kevin M. Johnson, K. A. Sirovy, Sandra M. Casas, Jerome F. La Peyre, and Morgan W. Kelly. Characterizing the epigenetic and transcriptomic responses to Perkinsus marinus infection in the eastern oyster Crassostrea virginica. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Stock:2020:AES

[701] Andy Stock and Ajit Subramaniam. Accuracy of empirical satellite algorithms for mapping phytoplankton diagnostic pigments in the open ocean: a supervised learning perspective. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Almahasheer:2020:PME

[702] Hanan Almahasheer and Carlos M. Duarte. Perceptions of marine environmental issues by Saudi citizens. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Vallet:2020:NHB

[703] Marine Vallet, Yee-Meng Chong, Anne Tourneroche, Grégory Genta-Jouve, Cédric Hubas, Raphael Lami, Claire M. M. Gachon, Tatyana Klochkova, Kok-Gan Chan, and Soizic Prado. Novel α-hydroxy γbutenolides of kelp endophytes disrupt bacterial cell-to-cell signaling. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Briseno-Avena:2020:CVD

[704] Christian Briseño-Avena, Jennifer C. Prairie, Peter J. S. Franks, and Jules S. Jaffe. Comparing vertical distributions of Chl-a fluorescence, marine snow, and taxon-specific zooplankton in relation to density using high-resolution optical measurements. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

True:2020:PHC

[705] Aaron C. True, Donald R. Webster, Marc J. Weissburg, and Jeannette Yen. Persistent hydrodynamic cues elicit orientation-specific behavioral sensitivities and kinematic responses in dispersed crab larvae. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Liu:2020:ICC

[706] Shuhao Liu, Yang Liu, Irene D. Alabia, Yongjun Tian, Zhenjiang Ye, Haiqing Yu, Jianchao Li, and Jiahua Cheng. Impact of climate change on wintering ground of Japanese anchovy (Engraulis japonicus) using marine geospatial statistics. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Tsang:2020:PRO

[707] Hin Hung Tsang, Megan J. Welch, Philip L. Munday, Timothy Ravasi, and Celia Schunter. Proteomic responses to ocean acidification in the brain of juvenile coral reef fish. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Brown:2020:URS

[708] Kristina A. Brown, Johnna M. Holding, and Eddy C. Carmack. Understanding regional and seasonal variability is key to gaining a pan-Arctic perspective on Arctic Ocean freshening. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Pech-Puch:2020:AAP

[709] Dawrin Pech-Puch, Judith Berastegui-Cabrera, Mar Pérez-Povedano, Harold Villegas-Hernández, Sergio Guillén-Hernández, Bastien Cautain, Fernando Reyes, Jerónimo Pachón, Patricia Gómez, Jaime Rodríguez, Carlos Jiménez, and Javier Sánchez-Céspedes. Antiviral and antiproliferative potential of marine organisms from the Yucatan Peninsula, Mexico. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Gibbs:2020:NAC

[710] Max Gibbs, Daniel Leduc, Scott D. Nodder, Andrew Kingston, Andrew Swales, Ashley A. Rowden, Joshu Mountjoy, Greg Olsen, Ron Ovenden, Julie Brown, Sarah Bury, and Brittany Graham. Novel application of a compound-specific stable isotope (CSSI) tracking technique demonstrates connectivity between terrestrial and deep-sea ecosystems via submarine canyons. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Selmoni:2020:UMC

[711] Oliver Selmoni, Gaël Lecellier, Lara Ainley, Antoine Collin, Raimana Doucet, Vaimiti Dubousquet, Hudson Feremaito, Edouard Ito Waia, Stuart Kininmonth, Hélène Magalon, Siola’a Malimali, Ateliana Maugateau, Anders Meibom, Stephen Mosese, Malika René-Trouillefou, Noriyuki Satoh, Madeleine J. H. van Oppen, André Xozamé, Maxime Yékawene, Stéphane Joost, and Véronique Berteaux-Lecellier. Using modern conservation tools for innovative management of coral reefs: The MANACO consortium. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Sokolowska:2020:NAV

[712] Ewa Sokołowska, Magdalena Gozdowska, and Ewa Kulczykowska. Nonapeptides arginine vasotocin and isotocin in fishes: Advantage of bioactive molecules measurement. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Chemel:2020:GNB

[713] Mathilde Chemel, Fanny Noisette, Denis Chabot, Ella Guscelli, Luc Leclerc, and Piero Calosi. Good news — bad news: Combined ocean change drivers decrease survival but have no negative impact on nutritional value and organoleptic quality of the northern shrimp. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kim:2020:MOP

[714] Jee-Hoon Kim, Kyoung-Ho Cho, Hyoung Sul La, Eun Jung Choy, Kohei Matsuno, Sung-Ho Kang, Won Kim, and Eun Jin Yang. Mass occurrence of Pacific copepods in the Southern Chukchi Sea during summer: Implications of the high-temperature Bering summer water. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Yan:2020:RSI

[715] Dong Yan, Kazuhiro Yoshida, Jun Nishioka, Masato Ito, Takenobu Toyota, and Koji Suzuki. Response to sea ice melt indicates high seeding potential of the ice diatom Thalassiosira to spring phytoplankton blooms: a laboratory study on an ice algal community from the Sea of Okhotsk. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Carlson:2020:LMF

[716] Andrew K. Carlson, Daniel I. Rubenstein, and Simon A. Levin. Linking multiscalar fisheries using metacoupling models. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kramer:2020:CMP

[717] Netanel Kramer, Raz Tamir, Gal Eyal, and Yossi Loya. Coral morphology portrays the spatial distribution and population size-structure along a 5–100 m depth gradient. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Wang:2020:AFS

[718] Yibang Wang, Yuanchao Wang, Cui Liang, Hui Zhang, and Weiwei Xian. Assessment of 12 fish species in the Northwest Pacific using the CMSY and BSM methods. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Tan:2020:SRP

[719] Yi Mei Tan, Oliver Dalby, Gary A. Kendrick, John Statton, Elizabeth A. Sinclair, Matthew W. Fraser, Peter I. Macreadie, Chris L. Gillies, Rhys A. Coleman, Michelle Waycott, Kor jent van Dijk, Adriana Vergés, Jeff D. Ross, Marnie L. Campbell, Fleur E. Matheson, Emma L. Jackson, Andrew D. Irving, Laura L. Govers, Rod M. Connolly, Ian M. McLeod, Michael A. Rasheed, Hugh Kirkman, Mogens R. Flindt, Troels Lange, Adam D. Miller, and Craig D. H. Sherman. Seagrass restoration is possible: Insights and lessons from Australia and New Zealand. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ju:2020:ESS

[720] Peilong Ju, Yongjun Tian, Mingru Chen, Shengyun Yang, Yang Liu, Qinwang Xing, and Peng Sun. Evaluating stock status of 16 commercial fish species in the coastal and offshore waters of Taiwan using the CMSY and BSM methods. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Post:2020:MBT

[721] Valerie Post and Dale Squires. Managing bigeye tuna in the Western and Central Pacific Ocean. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Fourrier:2020:RNN

[722] Marine Fourrier, Laurent Coppola, Hervé Claustre, Fabrizio D’Ortenzio, Raphaëlle Sauzède, and Jean-Pierre Gattuso. A regional neural network approach to estimate water-column nutrient concentrations and carbonate system variables in the Mediterranean Sea: CANYON-MED. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745. See corrigendum [1929].

Chance:2020:SII

[723] Rosie Chance, Liselotte Tinel, Amit Sarkar, Alok K. Sinha, Anoop S. Mahajan, Racheal Chacko, P. Sabu, Rajdeep Roy, Tim D. Jickells, David P. Stevens, Martin Wadley, and Lucy J. Carpenter. Surface inorganic iodine speciation in the Indian and Southern Oceans from 12^{o}N to 70^{o}S. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Pennino:2020:CFI

[724] Maria Grazia Pennino, Marta Coll, Marta Albo-Puigserver, Elena Fernández-Corredor, Jeroen Steenbeek, Ana Giráldez, María González, Antonio Esteban, and José M. Bellido. Current and future influence of environmental factors on small pelagic fish distributions in the Northwestern Mediterranean Sea. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Liang:2020:AEFb

[725] Cui Liang, Weiwei Xian, and Daniel Pauly. Assessments of 15 exploited fish stocks in Chinese, South Korean and Japanese waters using the CMSY and BSM methods. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Lea:2020:TTS

[726] James S. E. Lea, Nicolas E. Humphries, Jenny Bortoluzzi, Ryan Daly, Rainer G. von Brandis, Ela Patel, Evan Patel, Christopher R. Clarke, and David W. Sims. At the turn of the tide: Space use and habitat partitioning in two sympatric shark species is driven by tidal phase. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Rybicki:2020:FFE

[727] Sandra Rybicki, Katell G. Hamon, Sarah Simons, and Axel Temming. To fish or not to fish — economic perspectives of the pelagic Northeast Atlantic mackerel and herring fishery. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Guilpin:2020:RVI

[728] Marie Guilpin, Véronique Lesage, Ian McQuinn, Pablo Brosset, Thomas Doniol-Valcroze, Tiphaine Jeanniard du Dot, and Gesche Winkler. Repeated vessel interactions and climate- or fishery-driven changes in prey density limit energy acquisition by foraging blue whales. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ruckdeschel:2020:BDZ

[729] Gennavieve S. Ruckdeschel, Kimberley T. A. Davies, and Tetjana Ross. Biophysical drivers of zooplankton variability on the Scotian Shelf observed using profiling electric gliders. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Guerra-Vargas:2020:STD

[730] Luis Alberto Guerra-Vargas, Lucy Gwen Gillis, and José Ernesto Mancera-Pineda. Stronger together: Do coral reefs enhance seagrass meadows “Blue carbon” potential? Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Muhl:2020:EGT

[731] Ella-Kari Muhl, Ana Carolina Esteves Dias, and Derek Armitage. Experiences with governance in three marine conservation zoning initiatives: Parameters for assessment and pathways forward. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Page-Karjian:2020:ACH

[732] Annie Page-Karjian, Catherine F. Lo, Branson Ritchie, Craig A. Harms, David S. Rotstein, Sushan Han, Sayed M. Hassan, Andreas F. Lehner, John P. Buchweitz, Victoria G. Thayer, Jill M. Sullivan, Emily F. Christiansen, and Justin R. Perrault. Anthropogenic contaminants and histopathological findings in stranded cetaceans in the Southeastern United States, 2012–2018. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

vanderKaaden:2020:SSO

[733] Anna-Selma van der Kaaden, Dick van Oevelen, Max Rietkerk, Karline Soetaert, and Johan van de Koppel. Spatial self-organization as a new perspective on cold-water coral mound development. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ju:2020:SSE

[734] Peilong Ju, Mingru Chen, Yongjun Tian, Yu Zhao, Shengyun Yang, and Jiamei Xiao. Stock status estimating of 5 shark species in the waters around Taiwan using a length-based Bayesian biomass estimation (LBB) method. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Lee:2020:RGI

[735] Benjamin B. Lee, Eric J. Schott, Donald C. Behringer, Jamie Bojko, Andrew Kough, and Louis V. Plough. Rapid genetic identification of the blue crab Callinectes sapidus and other Callinectes spp. using restriction enzyme digestion and High Resolution Melt (HRM) assays. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Listmann:2020:EEI

[736] Luisa Listmann, Giannina S. I. Hattich, Birte Matthiessen, and Thorsten B. H. Reusch. Eco-evolutionary interaction in competing phytoplankton: Nutrient driven genotype sorting likely explains dominance shift and species responses to CO_{2}. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Wrightson:2020:QIC

[737] Lewis Wrightson and Alessandro Tagliabue. Quantifying the impact of climate change on marine diazotrophy: Insights from earth system models. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Kytinou:2020:GSR

[738] Eleni Kytinou, Maria Sini, Yiannis Issaris, and Stelios Katsanevakis. Global systematic review of methodological approaches to analyze coastal shelf food webs. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Erbe:2020:EIS

[739] Christine Erbe, Joshua Nathan Smith, Jessica V. Redfern, and David Peel. Editorial: Impacts of shipping on marine fauna. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Planas:2020:MES

[740] Miquel Planas, Ike Olivotto, María Jesús González, Rosaria Laurà, and Matteo Zarantoniello. A multidisciplinary experimental study on the effects of breeders diet on newborn seahorses (Hippocampus guttulatus). Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Boyd:2020:PCC

[741] Robin Boyd, Robert Thorpe, Kieran Hyder, Shovonlal Roy, Nicola Walker, and Richard Sibly. Potential consequences of climate and management scenarios for the Northeast Atlantic mackerel fishery. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Wang:2020:SSA

[742] Yuanchao Wang, Cui Liang, Yibang Wang, Weiwei Xian, and Maria Lourdes Palomares. Stock status assessments for 12 exploited fishery species in the Tsushima Warm Current Region, Southwest Japan and East China, using the CMSY and BSM methods. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Subramaniam:2020:TDF

[743] Roshni C. Subramaniam, Jessica Melbourne-Thomas, Stuart P. Corney, Karen Alexander, Clara Péron, Philippe Ziegler, and Kerrie M. Swadling. Time-dynamic food web modeling to explore environmental drivers of ecosystem change on the Kerguelen Plateau. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Gong:2020:TND

[744] Yi Gong, Yunkai Li, Xinjun Chen, and Wei Yu. Trophic niche and diversity of a pelagic squid (Dosidicus gigas): a comparative study using stable isotope, fatty acid, and feeding apparatuses morphology. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Neely:2020:DST

[745] Karen L. Neely, Cindy L. Lewis, and Kevin A. Macaulay. Disparities in spawning times between in situ and ex situ pillar corals. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Goelz:2020:DAT

[746] Taylor Goelz, Troy Hartley, Andrew Scheld, and Inga Carboni. The development of attitudes toward scientific models during a participatory modeling process — the impact of participation and social network structure. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Munoz:2020:ATT

[747] Lidia Muñoz, Eneko Aspillaga, Miquel Palmer, João L. Saraiva, and Pablo Arechavala-Lopez. Acoustic telemetry: a tool to monitor fish swimming behavior in sea-cage aquaculture. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Ardila:2020:MYD

[748] Néstor E. Ardila, Hernando Hernández, Astrid Muñoz-Ortiz, Óscar J. Ramos, Erick Castro, Nacor Bolaños, Anthony Rojas, and Juan A. Sánchez. Multi-year density variation of queen conch (Aliger gigas) on Serrana Bank, Seaflower Biosphere Reserve, Colombia: Implications for fisheries management. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Metz:2020:MJG

[749] Tasha L. Metz, Mandi Gordon, Marc Mokrech, and George Guillen. Movements of juvenile green turtles (Chelonia mydas) in the nearshore waters of the Northwestern Gulf of Mexico. Frontiers in Marine Science, 7, July 2020. ISSN 2296-7745.

Naik:2020:MIV

[750] Subrat Naik, Rajani Kanta Mishra, K. C. Sahu, Aneesh A. Lotliker, U. S. Panda, and Pravakar Mishra. Monsoonal influence and variability of water quality, phytoplankton biomass in the tropical coastal waters — a multivariate statistical approach. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Sweetman:2020:IMC

[751] Andrew K. Sweetman, Barbro T. Haugland, Astri J. S. Kvassnes, and Stefan G. Bolam. Impeded macrofaunal colonization and recovery following marine deposition of inert and organically modified mine-tailings. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ulman:2020:LFT

[752] Aylin Ulman, Mustafa Zengin, Nazli Demirel, and Daniel Pauly. The lost fish of Turkey: a recent history of disappeared species and commercial fishery extinctions for the Turkish Marmara and Black Seas. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745. See corrigendum [1229].

Scrosati:2020:CDC

[753] Ricardo A. Scrosati. Cyclone-driven coastal upwelling and cooling depend on location relative to the cyclone’s path: Evidence from Dorian’s arrival to Atlantic Canada. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Kaplan:2020:EBH

[754] Isaac C. Kaplan, Cecilie Hansen, Hem Nalini Morzaria-Luna, Raphael Girardin, and Kristin N. Marshall. Ecosystem-based harvest control rules for Norwegian and US ecosystems. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Dyomin:2020:MPS

[755] Victor Dyomin, Alexandra Davydova, Sergey Morgalev, Nikolay Kirillov, Alexey Olshukov, Igor Polovtsev, and Sergey Davydov. Monitoring of plankton spatial and temporal characteristics with the use of a submersible digital holographic camera. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Paparella:2020:SMG

[756] Francesco Paparella and Marcello Vichi. Stirring, mixing, growing: Microscale processes change larger scale phytoplankton dynamics. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Cook:2020:IBA

[757] Melissa Cook, Victoria S. Dunch, and Andrew T. Coleman. An interviewbased approach to assess angler practices and sea turtle captures on Mississippi fishing piers. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Reeves:2020:FBO

[758] Simon E. Reeves, Julianna J. Renzi, Emily K. Fobert, Brian R. Silliman, Boze Hancock, and Chris L. Gillies. Facilitating better outcomes: How positive species interactions can improve oyster reef restoration. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Reygondeau:2020:CCI

[759] Gabriel Reygondeau, William W. L. Cheung, Colette C. C. Wabnitz, Vicky W. Y. Lam, Thomas Frölicher, and Olivier Maury. Climate changeinduced emergence of novel biogeochemical provinces. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Hamabata:2020:NOI

[760] Tomoko Hamabata, Ayumi Matsuo, Mitsuhiko P. Sato, Satomi Kondo, Kazunari Kameda, Isao Kawazu, Takuya Fukuoka, Katsufumi Sato, Yoshihisa Suyama, and Masakado Kawata. Natal origin identification of green turtles in the North Pacific by genome-wide population analysis with limited DNA samples. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Zou:2020:EFP

[761] Songbao Zou, Qianqian Zhang, Xiaoli Zhang, Christine Dupuy, and Jun Gong. Environmental factors and pollution stresses select bacterial populations in association with protists. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Wiener:2020:CSR

[762] Carlie Wiener, Lars Bejder, David Johnston, Leesa Fawcett, and Paul Wilkinson. Cashing in on spinners: Revenue estimates of wild dolphinswim tourism in the Hawaiian Islands. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Angeletti:2020:CCC

[763] Lorenzo Angeletti, Giorgio Castellan, Paolo Montagna, Alessandro Remia, and Marco Taviani. The “Corsica Channel Cold-Water Coral Province” (Mediterranean Sea). Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Xu:2020:LNM

[764] Xu Xu, Carsten Lemmen, and Kai W. Wirtz. Less nutrients but more phytoplankton: Long-term ecosystem dynamics of the Southern North Sea. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Afonso:2020:DLK

[765] André S. Afonso, Pedro Roque, Leonardo Fidelis, Leonardo Veras, Anxo Conde, Paulo Maranhão, Sérgio Leandro, and Fábio H. V. Hazin. Does lack of knowledge lead to misperceptions? Disentangling the factors modulating public knowledge about and perceptions toward sharks. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Petersen:2020:IMR

[766] Dennis S. Petersen, Stanislav N. Gorb, and Lars Heepe. The influence of material and roughness on the settlement and the adhesive strength of the barnacle balanus improvisus in the Baltic Sea. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Cepeda:2020:DDS

[767] Diego Cepeda, Fernando Pardos, Daniela Zeppilli, and Nuria Sánchez. Dragons of the deep sea: Kinorhyncha communities in a pockmark field at Mozambique Channel, with the description of three new species. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Willson:2020:ELT

[768] Andrew Willson, Blair Witherington, Robert Baldwin, Manjula Tiwari, Thuraya Al Sariri, Ali Al Kiyumi, Suaad Al Harthi, Maïa Sarrouf Willson, Asma Al Bulushi, Ghasi Al Farsi, Juma Al Humaidy, Juma Al Araimi, Lamees A. Daar, Barbara Schroeder, James Perran Ross, and Earl Possardt. Evaluating the long-term trend and management of a globally important loggerhead population nesting on Masirah Island, Sultanate of Oman. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Hernandez-Hernandez:2020:DPD

[769] Nauzet Hernández-Hernández, Javier Arístegui, María F. Montero, Esther Velasco-Senovilla, Federico Baltar, Ángeles Marrero-Díaz, Antonio Martínez-Marrero, and Ángel Rodríguez-Santana. Drivers of plankton distribution across mesoscale eddies at submesoscale range. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Hennige:2020:CRC

[770] Sebastian J. Hennige, Uwe Wolfram, Leslie Wickes, Fiona Murray, J. Murray Roberts, Nicholas A. Kamenos, Sebastian Schofield, Alexander Groetsch, Ewa M. Spiesz, Marie-Eve Aubin-Tam, and Peter J. Etnoyer. Crumbling reefs and cold-water coral habitat loss in a future ocean: Evidence of “Coralporosis” as an indicator of habitat integrity. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Buscher:2020:ALC

[771] Elena Buscher, Darcy L. Mathews, Cheryl Bryce, Kathleen Bryce, Darlene Joseph, and Natalie C. Ban. Applying a low cost, mini remotely operated vehicle (ROV) to assess an ecological baseline of an indigenous seascape in Canada. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Malone:2020:GCE

[772] Thomas C. Malone and Alice Newton. The globalization of cultural eutrophication in the coastal ocean: Causes and consequences. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Baldry:2020:SCM

[773] Kimberlee Baldry, Peter G. Strutton, Nicole A. Hill, and Philip W. Boyd. Subsurface chlorophyll-a maxima in the Southern Ocean. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Schmidt-Roach:2020:BRR

[774] Sebastian Schmidt-Roach, Carlos M. Duarte, Charlotte A. E. Hauser, and Manuel Aranda. Beyond reef restoration: Next-generation techniques for coral gardening, landscaping, and outreach. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Shaver:2020:GTC

[775] Donna J. Shaver, Hilary R. Frandsen, Jeffrey A. George, and Christian Gredzens. Green turtle (Chelonia mydas) nesting underscores the importance of protected areas in the Northwestern Gulf of Mexico. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Busch:2020:CDD

[776] Kathrin Busch, Erik Wurz, Hans Tore Rapp, Kristina Bayer, Andre Franke, and Ute Hentschel. Chloroflexi dominate the deep-sea golf ball sponges Craniella zetlandica and Craniella infrequens throughout different life stages. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Yang:2020:IAP

[777] Zixiang Yang, Han Zhang, Peihong Kang, Yangyang Zhao, and Tiantian Tang. Influence of algal production and decomposition on the carbon isotope signature of labile particulate organic matter on a productive continental shelf under the stress of coastal hypoxia. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Dell:2020:FHS

[778] Claire L. A. Dell, Guilherme O. Longo, Deron E. Burkepile, and Carrie Manfrino. Few herbivore species consume dominant macroalgae on a Caribbean coral reef. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Armengol:2020:PCC

[779] Laia Armengol, Gara Franchy, Alicia Ojeda, and Santiago Hernández-León. Plankton community changes from warm to cold winters in the oligotrophic subtropical ocean. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Sarkar:2020:MNL

[780] Amit Sarkar, Syed Wajih Ahmad Naqvi, Gaute Lavik, Anil Pratihary, Hema Naik, Gayatri Shirodkar, and Marcel M. M. Kuypers. Massive nitrogen loss over the Western Indian Continental Shelf during seasonal anoxia: Evidence from isotope pairing technique. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Urbini:2020:TSV

[781] Lidia Urbini, Gianmarco Ingrosso, Tamara Djakovac, Salvatore Piacentino, and Michele Giani. Temporal and spatial variability of the CO_{2} system in a riverine influenced area of the Mediterranean Sea, the Northern Adriatic. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Gomez:2020:MWC

[782] Gloria Silvana Durán Gómez, Takeyoshi Nagai, and Kotaro Yokawa. Mesoscale warm-core eddies drive interannual modulations of swordfish catch in the Kuroshio Extension System. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

laPuente:2020:GPR

[783] Santiago De la Puente, Rocío López de la Lama, Selene Benavente, Juan Carlos Sueiro, and Daniel Pauly. Growing into poverty: Reconstructing Peruvian small-scale fishing effort between 1950 and 2018. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Laroche:2020:SSS

[784] Olivier Laroche, Oliver Kersten, Craig R. Smith, and Erica Goetze. From sea surface to seafloor: a benthic allochthonous eDNA survey for the abyssal ocean. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Ternon:2020:ERM

[785] Eva Ternon, Benoît Paix, Olivier P. Thomas, Jean-François Briand, and Gérald Culioli. Exploring the role of macroalgal surface metabolites on the settlement of the benthic dinoflagellate ostreopsis cf. ovata. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Alvarado-Rybak:2020:PFC

[786] Mario Alvarado-Rybak, Frederick Toro, Paulette Abarca, Enrique Paredes, Sonia Español-Jiménez, and Mauricio Seguel. Pathological findings in cetaceans sporadically stranded along the Chilean Coast. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Booty:2020:SAE

[787] James M. Booty, Graham J. C. Underwood, Amie Parris, Richard G. Davies, and Trevor J. Tolhurst. Shorebirds affect ecosystem functioning on an intertidal mudflat. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745. See corrigendum [1337].

Xiong:2020:CSS

[788] Tian qi Xiong, Qin sheng Wei, Wei dong Zhai, Cheng long Li, Song yin Wang, Yi xing Zhang, Shuo jiang Liu, and Si qing Yu. Comparing subsurface seasonal deoxygenation and acidification in the Yellow Sea and Northern East China Sea along the north-to-south latitude gradient. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Avigliano:2020:ESB

[789] Esteban Avigliano, Alejandra Vanina Volpedo, and Benjamin D. Walther. Editorial: Studying the biology of aquatic animals through calcified structures. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Sawakuchi:2020:CCD

[790] Henrique O. Sawakuchi, Vania Neu, Nicholas D. Ward, Maria de Lourdes C. Barros, Aline M. Valerio, William Gagne-Maynard, Alan C. Cunha, Diani F. S. Less, Joel E. M. Diniz, Daimio C. Brito, Alex V. Krusche, and Jeffrey E. Richey. Corrigendum: Carbon dioxide emissions along the Lower Amazon River. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See [1].

Mearns-Spragg:2020:BEJ

[791] Andrew Mearns-Spragg, Jessica Tilman, Daniel Tams, and Ashley Barnes. The biological evaluation of jellyfish collagen as a new research tool for the growth and culture of iPSC derived microglia. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Doherty:2020:SES

[792] Philip D. Doherty, Annette Cameron Broderick, Brendan John Godley, K. A. Hart, Q. Phillips, A. Sanghera, Thomas B. Stringell, J. T. Walker, and Peter Bradley Richardson. Spatial ecology of sub-adult green turtles in coastal waters of the Turks and Caicos Islands: Implications for conservation management. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Canning-Clode:2020:WCC

[793] João Canning-Clode, Pedro Sepúlveda, Sílvia Almeida, and João Monteiro. Will COVID-19 containment and treatment measures drive shifts in marine litter pollution? Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Brasier:2020:OOF

[794] Madeleine J. Brasier, Stacey McCormack, Narissa Bax, Jilda A. Caccavo, Emma Cavan, Jessica A. Ericson, Blanca Figuerola, Alyce Hancock, Svenja Halfter, Nicole Hellessey, Juan Höfer, Peter S. Puskic, César Soares de Oliveira, Roshni C. Subramaniam, Jake Wallis, and Christine K. Weldrick. Overcoming the obstacles faced by early career researchers in marine science: Lessons from the marine ecosystem assessment for the Southern Ocean. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Eastman:2020:PIP

[795] Catherine B. Eastman, Jessica A. Farrell, Liam Whitmore, Devon R. Rollinson Ramia, Rachel S. Thomas, Jenifer Prine, Scott F. Eastman, Todd Z. Osborne, Mark Q. Martindale, and David J. Duffy. Plastic ingestion in post-hatchling sea turtles: Assessing a major threat in Florida near shore waters. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Mao:2020:PGS

[796] Xiaowei Mao, Simona Augyte, Mao Huang, Matthew P. Hare, David Bailey, Schery Umanzor, Michael Marty-Rivera, Kelly R. Robbins, Charles Yarish, Scott Lindell, and Jean-Luc Jannink. Population genetics of sugar kelp throughout the Northeastern United States using genome-wide markers. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Hope:2020:ENE

[797] Julie A. Hope, Judi Hewitt, Conrad A. Pilditch, Candida Savage, and Simon F. Thrush. Effect of nutrient enrichment and turbidity on interactions between microphytobenthos and a key bivalve: Implications for higher trophic levels. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Juma:2020:VSC

[798] Gabriel A. Juma, Adiel M. Magana, Githaiga N. Michael, and James G. Kairo. Variation in seagrass carbon stocks between tropical estuarine and marine mangrove-fringed creeks. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Whitt:2020:FVA

[799] Christopher Whitt, Jay Pearlman, Brian Polagye, Frank Caimi, Frank Muller-Karger, Andrea Copping, Heather Spence, Shyam Madhusudhana, William Kirkwood, Ludovic Grosjean, Bilal Muhammad Fiaz, Satinder Singh, Sikandra Singh, Dana Manalang, Ananya Sen Gupta, Alain Maguer, Justin J. H. Buck, Andreas Marouchos, Malayath Aravindakshan Atmanand, Ramasamy Venkatesan, Vedachalam Narayanaswamy, Pierre Testor, Elizabeth Douglas, Sebastien de Halleux, and Siri Jodha Khalsa. Future vision for autonomous ocean observations. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Bahl:2020:SGW

[800] Alexis Bahl, Anand Gnanadesikan, and Marie-Aude S. Pradal. Scaling global warming impacts on ocean ecosystems: Lessons from a suite of earth system models. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Berninsone:2020:SGL

[801] Leonardo G. Berninsone, Pablo Bordino, Martín Gnecco, Mariana Foutel, Alice I. Mackay, and Timothy B. Werner. Switching gillnets to longlines: an alternative to mitigate the bycatch of Franciscana dolphins (Pontoporia blainvillei) in Argentina. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Wong:2020:ADT

[802] Annie P. S. Wong, Susan E. Wijffels, Stephen C. Riser, Sylvie Pouliquen, Shigeki Hosoda, Dean Roemmich, John Gilson, Gregory C. Johnson, Kim Martini, David J. Murphy, Megan Scanderbeg, T. V. S. Udaya Bhaskar, Justin J. H. Buck, Frederic Merceur, Thierry Carval, Guillaume Maze, Cécile Cabanes, Xavier André, Noé Poffa, Igor Yashayaev, Paul M. Barker, Stéphanie Guinehut, Mathieu Belbéoch, Mark Ignaszewski, Molly O’Neil Baringer, Claudia Schmid, John M. Lyman, Kristene E. McTaggart, Sarah G. Purkey, Nathalie Zilberman, Matthew B. Alkire, Dana Swift, W. Brechner Owens, Steven R. Jayne, Cora Hersh, Pelle Robbins, Deb West-Mack, Frank Bahr, Sachiko Yoshida, Philip J. H. Sutton, Romain Cancouët, Christine Coatanoan, Delphine Dobbler, Andrea Garcia Juan, Jerôme Gourrion, Nicolas Kolodziejczyk, Vincent Bernard, Bernard Bourlès, Hervé Claustre, Fabrizio D’Ortenzio, Serge Le Reste, Pierre-Yve Le Traon, Jean-Philippe Rannou, Carole Saout-Grit, Sabrina Speich, Virginie Thierry, Nathalie Verbrugge, Ingrid M. Angel-Benavides, Birgit Klein, Giulio Notarstefano, Pierre-Marie Poulain, Pedro Vélez-Belchí, Toshio Suga, Kentaro Ando, Naoto Iwasaska, Taiyo Kobayashi, Shuhei Masuda, Eitarou Oka, Kanako Sato, Tomoaki Nakamura, Katsunari Sato, Yasushi Takatsuki, Takashi Yoshida, Rebecca Cowley, Jenny L. Lovell, Peter R. Oke, Esmee M. van Wijk, Fiona Carse, Matthew Donnelly, W. John Gould, Katie Gowers, Brian A. King, Stephen G. Loch, Mary Mowat, Jon Turton, E. Pattabhi Rama Rao, M. Ravichandran, Howard J. Freeland, Isabelle Gaboury, Denis Gilbert, Blair J. W. Greenan, Mathieu Ouellet, Tetjana Ross, Anh Tran, Mingmei Dong, Zenghong Liu, Jianping Xu, KiRyong Kang, HyeongJun Jo, Sung-Dae Kim, and Hyuk-Min Park. Argo data 1999–2019: Two million temperature-salinity profiles and subsurface velocity observations from a global array of profiling floats. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Li:2020:MMP

[803] Yang Li, Zifeng Zhan, and Kuidong Xu. Morphology and molecular phylogenetic analysis of deep-sea purple gorgonians (Octocorallia: Victorgorgiidae) from seamounts in the tropical Western Pacific, with description of three new species. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Schoepf:2020:CTV

[804] Verena Schoepf, Maria U. Jung, Malcolm T. McCulloch, Nicole E. White, Michael Stat, and Luke Thomas. Corrigendum: Thermally variable, macrotidal reef habitats promote rapid recovery from mass coral bleaching. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See [347].

Barbeaux:2020:MHS

[805] Steven J. Barbeaux, Kirstin Holsman, and Stephani Zador. Marine heatwave stress test of ecosystem-based fisheries management in the Gulf of Alaska Pacific cod fishery. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Xu:2020:LSL

[806] Jiekai Xu, Jiazhen Sun, John Beardall, and Kunshan Gao. Lower salinity leads to improved physiological performance in the coccolithophorid Emiliania huxleyi, which partly ameliorates the effects of ocean acidification. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Buer:2020:NPC

[807] Anna-Lucia Buer, Daniel Taylor, Per Bergström, Lukas Ritzenhofen, and Annemarie Klemmstein. Nitrogen and phosphorous content in blue mussels (Mytilus spp.) across the Baltic Sea. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Longcore:2020:CBN

[808] Travis Longcore, Dan Duriscoe, Martin Aubé, Andreas Jechow, Christopher C. M. Kyba, and Kellie L. Pendoley. Commentary: Brightness of the night sky affects loggerhead (Caretta caretta) sea turtle hatchling misorientation but not nest site selection. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See [323].

Alms:2020:IDC

[809] Viola Alms and Matthias Wolff. Identification of drivers of change of the Gulf of Nicoya ecosystem (Costa Rica). Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Kwong:2020:CMM

[810] Lian E. Kwong, Natasha Henschke, Evgeny A. Pakhomov, Jason D. Everett, and Iain M. Suthers. Corrigendum: Mesozooplankton and micronekton active carbon transport in contrasting eddies. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See [90].

Fridman:2020:OOC

[811] Sophie Fridman. Ontogeny of the osmoregulatory capacity of teleosts and the role of ionocytes. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Gopal:2020:DCP

[812] Aiswarya Gopal, K. U. Abdul Jaleel, Usha V. Parameswaran, V. N. Sanjeevan, A. V. Saramma, Anilkumar Vijayan, N. Saravanane, G. V. M. Gupta, and M. Sudhakar. Distinctive community patterns with exceptional diversity of polychaetes around a tectonically active archipelago in the tropical Indian Ocean. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Clark:2020:LST

[813] Jennifer S. Clark, Alistair G. B. Poore, Melinda A. Coleman, and Martina A. Doblin. Local scale thermal environment and limited gene flow indicates vulnerability of warm edge populations in a habitat forming macroalga. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Gac:2020:CBO

[814] Jean-Philippe Gac, Pierre Marrec, Thierry Cariou, Christophe Guillerm, Éric Macé, Marc Vernet, and Yann Bozec. Cardinal buoys: an opportunity for the study of air–sea CO_{2} fluxes in coastal ecosystems. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Chen:2020:DPP

[815] Tzong-Yueh Chen, Chao-Chen Lai, Fuh-Kwo Shiah, and Gwo-Ching Gong. Dissolved and particulate primary production and subsequent bacterial C consumption in the Southern East China Sea. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Liu:2020:MIV

[816] Feng Liu, James T. Melton III, Juan M. Lopez-Bautista, and Nansheng Chen. Multiple intraspecific variations of mitochondrial genomes in the green-tide forming alga, Ulva compressa Linnaeus (Ulvophyceae, Chlorophyta). Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Pouso:2020:IAV

[817] Sarai Pouso, Ángel Borja, and María C. Uyarra. An interdisciplinary approach for valuing changes after ecological restoration in marine cultural ecosystem services. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Davies:2020:MME

[818] Harriet N. Davies, Jackie Gould, Renae K. Hovey, Ben Radford, and Gary A. Kendrick. Mapping the marine environment through a crosscultural collaboration. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Berto:2020:CHA

[819] Daniela Berto, Malgorzata Formalewicz, Giordano Giorgi, Federico Rampazzo, Claudia Gion, Benedetta Trabucco, Michele Giani, Marina Lipizer, Slavica Matijevic, Helen Kaberi, Christina Zeri, Oliver Bajt, Nevenka Mikac, Danijela Joksimovic, Andriana F. Aravantinou, Mateja Poje, Magdalena Cara, and Loredana Manfra. Challenges in harmonized assessment of heavy metals in the Adriatic and Ionian Seas. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Steen:2020:ACA

[820] Andrew D. Steen, Stephanie Kusch, Hussain A. Abdulla, Nevenka Cakić, Sarah Coffinet, Thorsten Dittmar, James M. Fulton, Valier Galy, Kai-Uwe Hinrichs, Anitra E. Ingalls, Boris P. Koch, Elizabeth Kujawinski, Zhanfei Liu, Helena Osterholz, Darci Rush, Michael Seidel, Julio Sepúlveda, and Stuart G. Wakeham. Analytical and computational advances, opportunities, and challenges in marine organic biogeochemistry in an era of “Omics”. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Babarro:2020:NIP

[821] Jose M. F. Babarro, Ramón Filgueira, Xosé A. Padín, and M. Angeles Longa Portabales. A novel index of the performance of Mytilus galloprovincialis to improve commercial exploitation in aquaculture. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Sawall:2020:DPC

[822] Yvonne Sawall, Moronke Harris, Mario Lebrato, Marlene Wall, and Ellias Yuming Feng. Discrete pulses of cooler deep water can decelerate coral bleaching during thermal stress: Implications for artificial upwelling during heat stress events. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Fregosi:2020:DWV

[823] Selene Fregosi, Danielle V. Harris, Haruyoshi Matsumoto, David K. Mellinger, Jay Barlow, Simone Baumann-Pickering, and Holger Klinck. Detections of whale vocalizations by simultaneously deployed bottommoored and deep-water mobile autonomous hydrophones. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Carranza:2020:GOR

[824] Alvar Carranza and Philine S. E. zu Ermgassen. A global overview of restorative shellfish mariculture. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Lee:2020:CRS

[825] Sang Yoon Lee and Hwa Jin Lee. Comprehensive RNA-seq analysis to evaluate the pigmentation-related genes involved in albinism of cichlid fish, Aulonocara baenschi. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Vanderklift:2020:UPR

[826] Mathew A. Vanderklift, Christopher Doropoulos, Daniel Gorman, Inês Leal, Antoine J. P. Minne, John Statton, Andrew D. L. Steven, and Thomas Wernberg. Using propagules to restore coastal marine ecosystems. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Armstrong:2020:STP

[827] Amelia J. Armstrong, Asia O. Armstrong, Frazer McGregor, Anthony J. Richardson, Michael B. Bennett, Kathy A. Townsend, Graeme C. Hays, Mike van Keulen, Jessica Smith, and Christine L. Dudgeon. Satellite tagging and photographic identification reveal connectivity between two UNESCO world heritage areas for reef manta rays. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Dvoretsky:2020:EEF

[828] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Effects of environmental factors on the abundance, biomass, and individual weight of juvenile red king crabs in the Barents Sea. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Jevne:2020:BSP

[829] Lone S. Jevne, Margrete S. Øvrelid, Andreas Hagemann, Nina Bloecher, Kristine B. Steinhovden, Anna S. Båtnes, Yngvar Olsen, and Kjell I. Reitan. Biofouling on salmon pen nets and cleaner fish shelters does not harbor planktonic stages of sea lice. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Lesser:2020:NBU

[830] Michael P. Lesser, Jessica K. Jarett, Cara L. Fiore, Megan M. Thompson, M. Sabrina Pankey, and Keir J. Macartney. A new “Business as usual” climate scenario and the stress response of the Caribbean coral Montastraea cavernosa. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Zhai:2020:AEF

[831] Lu Zhai, Cui Liang, and Daniel Pauly. Assessments of 16 exploited fish stocks in Chinese waters using the CMSY and BSM methods. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Vranken:2020:FGD

[832] Cooper Van Vranken, Berthe M. J. Vastenhoud, James P. Manning, Kristian S. Plet-Hansen, Julie Jakoboski, Patrick Gorringe, and Michela Martinelli. Fishing gear as a data collection platform: Opportunities to fill spatial and temporal gaps in operational sub-surface observation networks. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Gonzalez-Mirelis:2021:MDH

[833] Genoveva Gonzalez-Mirelis, Rebecca E. Ross, Jon Albretsen, and Pål Buhl-Mortensen. Modeling the distribution of habitat-forming, deep-sea sponges in the Barents Sea: The value of data. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Murphy:2020:STV

[834] Sinéad Murphy, Marie A. C. Petitguyot, Paul D. Jepson, Rob Deaville, Christina Lockyer, James Barnett, Matthew Perkins, Rod Penrose, Nicholas J. Davison, and Cóilín Minto. Spatio-temporal variability of harbor porpoise life history parameters in the North-East Atlantic. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Madricardo:2020:HDS

[835] Fantina Madricardo, Michol Ghezzo, Nicoletta Nesto, William Joseph Mc Kiver, Gian Claudio Faussone, Riccardo Fiorin, Federico Riccato, Peter Charles Mackelworth, Jelena Basta, Francesca De Pascalis, Aleksandra Kruss, Antonio Petrizzo, and Vanessa Moschino. How to deal with seafloor marine litter: an overview of the state-of-the-art and future perspectives. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Richards:2020:TSS

[836] Travis M. Richards, Tracey T. Sutton, and R. J. David Wells. Trophic structure and sources of variation influencing the stable isotope signatures of meso- and bathypelagic micronekton fishes. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Haynert:2020:FSH

[837] Kristin Haynert, Franziska Gluderer, Melanie M. Pollierer, Stefan Scheu, and Achim Wehrmann. Food spectrum and habitat-specific diets of benthic foraminifera from the Wadden Sea — a fatty acid biomarker approach. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Rosenberger:2020:SWD

[838] Kurt J. Rosenberger, Curt D. Storlazzi, Olivia M. Cheriton, Andrew W. M. Pomeroy, Jeff E. Hansen, Ryan J. Lowe, and Mark L. Buckley. Spectral wave-driven bedload transport across a coral reef flat/lagoon complex. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Henderson:2020:IVS

[839] Angus F. Henderson, Clive R. McMahon, Rob Harcourt, Christophe Guinet, Baptiste Picard, Simon Wotherspoon, and Mark A. Hindell. Inferring variation in southern elephant seal at-sea mortality by modelling tag failure. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Wang:2021:ULT

[840] Xuehui Wang, Yinglin He, Feiyan Du, Mengna Liu, Weilie Bei, Yancong Cai, and Yongsong Qiu. Using LBB tools to assess miter squid stock in the Northeastern South China Sea. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Reed:2020:EEE

[841] Joshua Reed, Robert Harcourt, Leslie New, and Kerstin Bilgmann. Extreme effects of extreme disturbances: a simulation approach to assess population specific responses. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Bell:2020:USA

[842] Tom W. Bell, Nick J. Nidzieko, David A. Siegel, Robert J. Miller, Kyle C. Cavanaugh, Norman B. Nelson, Daniel C. Reed, Dmitry Fedorov, Christopher Moran, Jordan N. Snyder, Katherine C. Cavanaugh, Christie E. Yorke, and Maia Griffith. The utility of satellites and autonomous remote sensing platforms for monitoring offshore aquaculture farms: a case study for canopy forming kelps. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Wo:2020:MDM

[843] Jia Wo, Chongliang Zhang, Xindong Pan, Binduo Xu, Ying Xue, and Yiping Ren. Modeling the dynamics of multispecies fisheries: a case study in the coastal water of North Yellow Sea, China. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Zhang:2020:AEF

[844] Shanshan Zhang, Yibang Wang, Yuanchao Wang, Cui Liang, and Weiwei Xian. Assessment of 11 exploited fish and invertebrate populations in the Japan Sea using the CMSY and BSM methods. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Shester:2021:WWD

[845] Geoffrey G. Shester, Ben Enticknap, Brianne Mecum, Ashley Blacow-Draeger, Tara Brock, and Susan Murray. A win–win for deep-sea corals and fishing: Increasing seafloor protections while restoring fishing opportunities off the United States West Coast. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Haunost:2020:CCS

[846] Mathias Haunost, Ulf Riebesell, and Lennart T. Bach. The calcium carbonate shell of Emiliania huxleyi provides limited protection against viral infection. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Weymer:2020:MLH

[847] Bradley A. Weymer, Phillipe A. Wernette, Mark E. Everett, Potpreecha Pondthai, Marion Jegen, and Aaron Micallef. Multi-layered high permeability conduits connecting onshore and offshore coastal aquifers. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Su:2021:CMH

[848] Zimeng Su, Gabriela S. Pilo, Stuart Corney, Neil J. Holbrook, Mao Mori, and Philippe Ziegler. Characterizing marine heatwaves in the Kerguelen Plateau region. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Sandonnini:2021:MAS

[849] Jessica Sandonnini, Yoana Del Pilar Ruso, Emilio Cortés Melendreras, and Francisca Giménez Casalduero. Massive aggregations of serpulidae associated with eutrophication of the Mar Menor, Southeast Iberian Peninsula. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Kreb:2020:LTP

[850] Danielle Kreb, Stanislav Lhota, Lindsay Porter, Alexandra Redman, Imelda Susanti, and Milan Lazecky. Long-term population and distribution dynamics of an endangered Irrawaddy dolphin population in Balikpapan Bay, Indonesia in response to coastal development. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Fazhan:2020:GDH

[851] Hanafiah Fazhan, Khor Waiho, Henrik Glenner, Julia Hwei Zhong Moh, Marina Hassan, and Mhd Ikhwanuddin. Gonadal degeneration and hepatopancreas alteration in orange mud crab Scylla olivacea infected with Sacculina beauforti (Crustacea; Rhizocephala; Sacculinidae). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Campbell:2020:USC

[852] Elizabeth Campbell, Andrea Pasara-Polack, Jeffrey C. Mangel, and Joanna Alfaro-Shigueto. Use of small cetaceans as bait in small-scale fisheries in Peru. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Eger:2020:FIS

[853] Aaron M. Eger, Adriana Vergés, Chang Geun Choi, Hartvig Christie, Melinda A. Coleman, Camilla W. Fagerli, Daisuke Fujita, Masatoshi Hasegawa, Jeong Ha Kim, Mariana Mayer-Pinto, Dan C. Reed, Peter D.

Steinberg, and Ezequiel M. Marzinelli. Financial and institutional support are important for large-scale kelp forest restoration. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Alves:2020:HRS

[854] Daniel Pavani Vicente Alves, Eduardo Bomfin Caldato, Denise Silva de Moura, Roberto P. Zanon dos Santos, and Luigi Jovane. Highresolution sub-bottom and magnetometer data from Southeastern Brazilian Coast. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Felix:2020:IRE

[855] Fernando Félix and Santiago F. Burneo. Imminent risk of extirpation for two bottlenose dolphin communities in the Gulf of Guayaquil, Ecuador. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

David:2020:HDM

[856] C. Gabriel David and Torsten Schlurmann. Hydrodynamic drivers and morphological responses on small coral islands — the Thoondu Spit on Fuvahmulah, The Maldives. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Blythe:2020:FOF

[857] Jessica Blythe and Christopher Cvitanovic. Five organizational features that enable successful interdisciplinary marine research. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Schuhbauer:2020:GFS

[858] Anna Schuhbauer, Daniel J. Skerritt, Naazia Ebrahim, Frédéric Le Manach, and U. Rashid Sumaila. The global fisheries subsidies divide between small- and large-scale fisheries. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Dong:2020:ILK

[859] Wu-Shan Dong, Peng-Fei Liu, Yu Liu, Yu Du, Yan-Hui Bi, and Zhi-Gang Zhou. Immunocytochemical localization of the kinetochore protein Nuf2p on the gametophyte chromosomes of a cultivar of Saccharina (Phaeophyta). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Rattray:2020:SBI

[860] Jayne E. Rattray and Rienk H. Smittenberg. Separation of branched and isoprenoid glycerol dialkyl glycerol tetraether (GDGT) isomers in peat soils and marine sediments using reverse phase chromatography. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Hoeke:2021:SFA

[861] Ron K. Hoeke, Herve Damlamian, Jérome Aucan, and Moritz Wandres. Severe flooding in the atoll nations of Tuvalu and Kiribati triggered by a distant Tropical Cyclone Pam. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Franzitta:2020:TGN

[862] Marco Franzitta, Eduardo Feijão, Maria Teresa Cabrita, Carla Gameiro, Ana Rita Matos, João Carlos Marques, Johannes W. Goessling, Patrick Reis-Santos, Vanessa F. Fonseca, Carlo Pretti, Isabel Caçador, and Bernardo Duarte. Toxicity going nano: Ionic versus engineered Cu nanoparticles impacts on the physiological fitness of the model diatom Phaeodactylum tricornutum. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Falco:2020:COC

[863] Chiara De Falco, Annalisa Bracco, and Claudia Pasquero. Climatic and oceanographic controls on coral bleaching conditions in the Maldivian region. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Becker:2020:EFM

[864] Sarah L. Becker, Teri E. Nicholson, Karl A. Mayer, Michael J. Murray, and Kyle S. Van Houtan. Environmental factors may drive the postrelease movements of surrogate-reared sea otters. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Scalpone:2020:SEW

[865] Cara R. Scalpone, Jessie C. Jarvis, James M. Vasslides, Jeremy M. Testa, and Neil K. Ganju. Simulated estuary-wide response of seagrass (Zostera marina) to future scenarios of temperature and sea level. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Close:2020:OOM

[866] Hilary G. Close and Lillian C. Henderson. Open-ocean minima in δ^{13}C values of particulate organic carbon in the lower euphotic zone. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Cantarero:2020:SFC

[867] Sebastian I. Cantarero, Carlos Henríquez-Castillo, Nadia Dildar, Cristian A. Vargas, Peter von Dassow, Marcela Cornejo-D’Ottone, and Julio Sepúlveda. Size-fractionated contribution of microbial biomass to suspended organic matter in the eastern tropical South Pacific oxygen minimum zone. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Hobbs:2020:ESA

[868] Laura Hobbs, Neil S. Banas, Finlo R. Cottier, Jørgen Berge, and Malin Daase. Eat or sleep: Availability of winter prey explains mid-winter and spring activity in an Arctic Calanus population. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Tinker-Kulberg:2020:EDB

[869] Rachel Tinker-Kulberg, Anthony Dellinger, Terry E. Brady, Lee Robertson, Melinda K. M. Goddard, John Bowzer, Sarah K. Abood, Christopher Kepley, and Kristen Dellinger. Effects of diet on the biochemical properties of limulus amebocyte lysate from horseshoe crabs in an aquaculture setting. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Stewart-Sinclair:2020:BRB

[870] Phoebe J. Stewart-Sinclair, Jemma Purandare, Elisa Bayraktarov, Nathan Waltham, Simon Reeves, John Statton, Elizabeth A. Sinclair, Benjamin M. Brown, Zoë I. Shribman, and Catherine E. Lovelock. Blue restoration — building confidence and overcoming barriers. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Tiralongo:2020:SMM

[871] Francesco Tiralongo, Giuseppina Messina, Bianca Maria Lombardo, Lucia Longhitano, Giovanni Li Volti, and Daniele Tibullo. Skin mucus of marine fish as a source for the development of antimicrobial agents. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Bracco:2020:TFI

[872] Annalisa Bracco, Claire B. Paris, Andrew J. Esbaugh, Kaitlin Frasier, Samantha B. Joye, Guangpeng Liu, Kurt L. Polzin, and Ana Carolina Vaz. Transport, fate and impacts of the deep plume of petroleum hydrocarbons formed during the Macondo Blowout. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Cisewski:2021:VMP

[873] Boris Cisewski, Hjálmar Hátún, Inga Kristiansen, Bogi Hansen, Karin Margretha H. Larsen, Sólvá Káradóttir Eliasen, and Jan Arge Jacobsen. Vertical migration of pelagic and mesopelagic scatterers from ADCP backscatter data in the Southern Norwegian Sea. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

McGovern:2020:TID

[874] Maeve McGovern, Alexey K. Pavlov, Anne Deininger, Mats A. Granskog, Eva Leu, Janne E. Søreide, and Amanda E. Poste. Terrestrial inputs drive seasonality in organic matter and nutrient biogeochemistry in a High Arctic fjord system (Isfjorden, Svalbard). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Manna:2020:CBD

[875] Gabriella La Manna, Fabio Ronchetti, Gianluca Sarà, Alberto Ruiu, and Giulia Ceccherelli. Common bottlenose dolphin protection and sustainable boating: Species distribution modeling for effective coastal planning. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Plon:2020:LTD

[876] Stephanie Plön, Christine Erbe, and Sabine Wintner. Long-term demographic and spatio-temporal trends of Indo–Pacific bottlenose dolphin (Tursiops aduncus) bycatch in bather protection nets off KwaZulu-Natal, South Africa. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Greer:2020:HRS

[877] Adam T. Greer, John C. Lehrter, Benjamin M. Binder, Aditya R. Nayak, Ranjoy Barua, Ana E. Rice, Jonathan H. Cohen, Malcolm N. McFarland, Alexis Hagemeyer, Nicole D. Stockley, Kevin M. Boswell, Igor Shulman, Sergio deRada, and Bradley Penta. High-resolution sampling of a broad marine life size spectrum reveals differing size- and composition-based associations with physical oceanographic structure. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

vanderJagt:2020:AFC

[878] Helga van der Jagt, Ingrid Wiedmann, Nicole Hildebrandt, Barbara Niehoff, and Morten H. Iversen. Aggregate feeding by the copepods Calanus and Pseudocalanus controls carbon flux attenuation in the Arctic Shelf Sea during the productive period. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Janout:2020:VSF

[879] Markus A. Janout, Jens Hölemann, Georgi Laukert, Alexander Smirnov, Thomas Krumpen, Dorothea Bauch, and Leonid Timokhov. On the variability of stratification in the freshwater-influenced Laptev Sea region. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See corrigendum [1340].

Sajna:2020:ARS

[880] Kuttuvan Valappil Sajna, Siya Kamat, and Chelliah Jayabaskaran. Antiproliferative role of secondary metabolites from Aspergillus unguis AG 1.1 (G) isolated from marine macroalgae Enteromorph sp. by inducing intracellular ROS production and mitochondrial membrane potential loss leading to apoptosis. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Ebmer:2020:APC

[881] David Ebmer, María José Navarrete, Pamela Muñoz, Luis Miguel Flores, Ulrich Gärtner, Jan Brabec, Sven Poppert, Anja Taubert, and Carlos Hermosilla. Anthropozoonotic parasites circulating in synanthropic and Pacific colonies of South American sea lions (Otaria flavescens): Noninvasive techniques data and a review of the literature. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Jin:2020:IOA

[882] Peng Jin, David A. Hutchins, and Kunshan Gao. The impacts of ocean acidification on marine food quality and its potential food chain consequences. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Kot:2020:PNC

[883] Brian C. W. Kot, Henry C. L. Tsui, Tabris Y. T. Chung, and Amy P. Y. Lau. Postmortem neuroimaging of cetacean brains using computed tomography and magnetic resonance imaging. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Wang:2020:RMP

[884] Haiqing Wang, Andreas Hagemann, Arne M. Malzahn, Aleksander Handå, Marianne Uhre, Elin Kjørsvik, and Kjell Inge Reitan. The role of manipulating photoperiod and temperature in oocyte development of the polychaete Hediste diversicolor (O. F. Müller, 1976). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Abelson:2020:CRC

[885] Avigdor Abelson, Daniel C. Reed, Graham J. Edgar, Carter S. Smith, Gary A. Kendrick, Robert J. Orth, Laura Airoldi, Brian Silliman, Michael W. Beck, Gesche Krause, Nadav Shashar, Noga Stambler, and Peter Nelson. Challenges for restoration of coastal marine ecosystems in the Anthropocene. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Hickey:2020:BRH

[886] Sharyn M. Hickey, Ben Radford, Chris M. Roelfsema, Karen E. Joyce, Shaun K. Wilson, Daniel Marrable, Kathryn Barker, Mathew Wyatt, Harriet N. Davies, Javier X. Leon, John Duncan, Thomas H. Holmes, Alan J. Kendrick, J. Nikolaus Callow, and Kathy Murray. Between a reef and a hard place: Capacity to map the next coral reef catastrophe. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Bluhm:2020:PAC

[887] Bodil A. Bluhm, Markus A. Janout, Seth L. Danielson, Ingrid Ellingsen, Maria Gavrilo, Jacqueline M. Grebmeier, Russell R. Hopcroft, Katrin B. Iken, Randi B. Ingvaldsen, Lis L. Jørgensen, Ksenia N. Kosobokova, Ron Kwok, Igor V. Polyakov, Paul E. Renaud, and Eddy C. Carmack. The pan-Arctic continental slope: Sharp gradients of physical processes affect pelagic and benthic ecosystems. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Bellanger:2020:AMC

[888] Manuel Bellanger, Cameron Speir, Fabian Blanchard, Kate Brooks, James R. A. Butler, Scott Crosson, Robert Fonner, Sophie Gourguet, Daniel S. Holland, Sakari Kuikka, Bertrand Le Gallic, Rebecca Lent, Gary D. Libecap, Douglas W. Lipton, Prateep Kumar Nayak, David Reid, Pierre Scemama, Robert Stephenson, Olivier Thébaud, and Juliette C. Young. Addressing marine and coastal governance conflicts at the interface of multiple sectors and jurisdictions. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Xu:2020:NEI

[889] Ningning Xu, Kai Xu, Wenlei Wang, Yan Xu, Dehua Ji, Changsheng Chen, and Chaotian Xie. Nutrient enrichment improves growth and food quality of two strains of the economic seaweed Pyropia haitanensis. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

David:2020:ORR

[890] Andrew A. David. Oyster reef restoration and biological invasions: an overlooked or a non-issue? Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Hu:2020:LEG

[891] Zi-Min Hu, Kai-Le Zhong, Florian Weinberger, De-Lin Duan, Stefano G. A. Draisma, and Ester A. Serrão. Linking ecology to genetics to better understand adaptation and evolution: a review in marine macrophytes. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Souravlias:2020:DFM

[892] Dimitris Souravlias, Ioannis Dafnomilis, Jens Ley, Gerrit Assbrock, Mark B. Duinkerken, Rudy R. Negenborn, and Dingena L. Schott. Design framework for a modular floating container terminal. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

vanEmmerik:2020:MRM

[893] Tim van Emmerik, Justin van Klaveren, Lourens J. J. Meijer, Joost W. Krooshof, Dixie Ann A. Palmos, and Maria Antonia Tanchuling. Manila River mouths act as temporary sinks for macroplastic pollution. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Pelletier:2020:AEC

[894] Dominique Pelletier. Assessing the effectiveness of coastal Marine Protected Area management: Four learned lessons for science uptake and upscaling. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Doi:2020:SAS

[895] Takeshi Doi, Masami Nonaka, and Swadhin Behera. Skill assessment of seasonal-to-interannual prediction of sea level anomaly in the North Pacific based on the SINTEX-F climate model. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Ma:2020:CVP

[896] Shuyang Ma, Yongjun Tian, Jianchao Li, Haiqing Yu, Jiahua Cheng, Peng Sun, Caihong Fu, Yang Liu, and Yoshiro Watanabe. Climate variability patterns and their ecological effects on ecosystems in the Northwestern North Pacific. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Morley:2020:GDS

[897] Simon A. Morley, Doris Abele, David K. A. Barnes, César A. Cárdenas, Cedric Cotté, Julian Gutt, Sian F. Henley, Juan Höfer, Kevin A. Hughes, Stephanie M. Martin, Carlos Moffat, Marilyn Raphael, Sharon E. Stammerjohn, Coleen C. Suckling, Vivitskaia J. D. Tulloch, Cath L. Waller, and Andrew J. Constable. Global drivers on Southern Ocean ecosystems: Changing physical environments and anthropogenic pressures in an earth system. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Young:2020:HAI

[898] Craig S. Young and Christopher J. Gobler. Hypoxia and acidification, individually and in combination, disrupt herbivory and reduce survivorship of the gastropod, lacuna vincta. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Nonaka:2020:ADI

[899] Masami Nonaka, Hideharu Sasaki, Bunmei Taguchi, and Niklas Schneider. Atmospheric-driven and intrinsic interannual-to-decadal variability in the Kuroshio Extension jet and eddy activities. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745. See corrigendum [1321].

Mustaffa:2020:COD

[900] Nur Ili Hamizah Mustaffa, Liisa Kallajoki, Johanna Biederbick, Franziska Isabell Binder, Alexandra Schlenker, and Maren Striebel. Coastal ocean darkening effects via terrigenous DOM addition on plankton: an indoor mesocosm experiment. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Neves:2020:CWS

[901] Bárbara de Moura Neves, Vonda Wareham Hayes, Erin Herder, Kevin Hedges, Cindy Grant, and Philippe Archambault. Cold-water soft corals (Cnidaria: Nephtheidae) as habitat for juvenile basket stars (Echinodermata: Gorgonocephalidae). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Schmidt:2020:SDS

[902] Sabine Schmidt, Gilbert Maudire, Cécile Nys, Joël Sudre, Valérie Harscoat, Gérald Dibarboure, and Frédéric Huynh. Streamlining data and service centers for easier access to data and analytical services: The strategy of ODATIS as the gateway to French marine data. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Cook:2020:MAI

[903] April B. Cook, Andrea M. Bernard, Kevin M. Boswell, Heather Bracken-Grissom, Marta D’Elia, Sergio deRada, Cole G. Easson, David English, Ron I. Eytan, Tamara Frank, Chuanmin Hu, Matthew W. Johnston, Heather Judkins, Chad Lembke, Jose V. Lopez, Rosanna J. Milligan, Jon A. Moore, Bradley Penta, Nina M. Pruzinsky, John A. Quinlan, Travis M. Richards, Isabel C. Romero, Mahmood S. Shivji, Michael Vecchione, Max D. Weber, R. J. David Wells, and Tracey T. Sutton. A multidisciplinary approach to investigate deep-pelagic ecosystem dynamics in the Gulf of Mexico following Deepwater Horizon. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

ODaly:2020:ECF

[904] Stephanie H. O’Daly, Seth L. Danielson, Sarah M. Hardy, Russell R. Hopcroft, Catherine Lalande, Dean A. Stockwell, and Andrew M. P. McDonnell. Extraordinary carbon fluxes on the shallow Pacific Arctic Shelf during a remarkably warm and low sea ice period. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Zhai:2020:OPC

[905] Shaowei Zhai, Yue Wang, Yingxia He, and Xuehao Chen. Oligomeric proanthocyanidins counteracts the negative effects of high level of dietary histamine on American eel (Anguilla rostrata). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Rulent:2020:CWL

[906] Julia Rulent, Francisco M. Calafat, Christopher J. Banks, Lucy May Bricheno, Christine Gommenginger, J. A. Mattias Green, Ivan D. Haigh, Huw Lewis, and Adrien C. H. Martin. Comparing water level estimation in coastal and shelf seas from satellite altimetry and numerical models. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Archer:2020:FSA

[907] Stephanie K. Archer, Amanda S. Kahn, Mary Thiess, Lauren Law, Sally P. Leys, Sophia C. Johannessen, Craig A. Layman, Lily Burke, and Anya Dunham. Foundation species abundance influences food web topology on glass sponge reefs. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Pakiding:2020:CEI

[908] Fitryanti Pakiding, Kartika Zohar, Alberto Y. T. Allo, Sinus Keroman, Deasy Lontoh, Peter H. Dutton, and Manjula Tiwari. Community engagement: an integral component of a multifaceted conservation approach for the transboundary Western Pacific leatherback. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Marshall:2020:EDH

[909] Christopher D. Marshall, Joshua A. Cullen, Mehsin Al-Ansi, Shafeeq Hamza, and Mohamed A. R. Abdel-Moati. Environmental drivers of habitat use by hawksbill turtles (Eretmochelys imbricata) in the Arabian Gulf (Qatar). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Mentges:2020:MPG

[910] Andrea Mentges, Curtis Deutsch, Christoph Feenders, Sinikka T. Lennartz, Bernd Blasius, and Thorsten Dittmar. Microbial physiology governs the oceanic distribution of dissolved organic carbon in a scenario of equal degradability. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Stone:2020:LCC

[911] Hally B. Stone, Neil S. Banas, Parker MacCready, Raphael M. Kudela, and Bridget Ovall. Linking chlorophyll concentration and wind patterns using satellite data in the Central and Northern California Current system. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Mack:2020:SMM

[912] Leoni Mack, Jenni Attila, Eva Aylagas, Arne Beermann, Angel Borja, Daniel Hering, Maria Kahlert, Florian Leese, Robin Lenz, Maiju Lehtiniemi, Antonia Liess, Urmas Lips, Olli-Pekka Mattila, Kristian Meissner, Timo Pyhälahti, Outi Setälä, Jennifer S. Strehse, Laura Uusitalo, Anna Willstrand Wranne, and Sebastian Birk. A synthesis of marine monitoring methods with the potential to enhance the status assessment of the Baltic Sea. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Parsa:2020:RMB

[913] Mahdi Parsa, Timothy J. Emery, Ashley J. Williams, and Simon Nicol. A robust métier-based approach to classifying fishing practices within commercial fisheries. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Kaile:2020:PRD

[914] Namrata Kaile, Mathilde Lindivat, Javier Elio, Gunnar Thuestad, Quentin G. Crowley, and Ingunn Alne Hoell. Preliminary results from detection of microplastics in liquid samples using flow cytometry. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Pan:2020:SST

[915] Dawei Pan, Xiaoyan Ding, Haitao Han, Shenghui Zhang, and Chenchen Wang. Species, spatial-temporal distribution, and contamination assessment of trace metals in typical mariculture area of North China. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Oakes:2021:USI

[916] Rosie L. Oakes, Catherine V. Davis, and Jocelyn A. Sessa. Using the stable isotopic composition of Heliconoides inflatus pteropod shells to determine calcification depth in the Cariaco Basin. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ramirez-Gonzalez:2020:OMT

[917] Jorge Ramírez-González, Nicolas Moity, Solange Andrade-Vera, and Harry Reyes. Overexploitation and more than a decade of failed management leads to no recovery of the Galápagos sea cucumber fishery. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Heymans:2020:ODT

[918] Johanna J. Heymans, Alida Bundy, Villy Christensen, Marta Coll, Kim de Mutsert, Elizabeth A. Fulton, Chiara Piroddi, Yunne-Jai Shin, Jeroen Steenbeek, and Morgane Travers-Trolet. The ocean decade: a true ecosystem modeling challenge. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Franco:2020:ADP

[919] Davide Di Franco, Katrin Linse, Huw J. Griffiths, Christian Haas, Hanieh Saeedi, and Angelika Brandt. Abundance and distributional patterns of benthic peracarid crustaceans from the Atlantic sector of the Southern Ocean and Weddell Sea. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Messyasz:2020:CBP

[920] Adriana Messyasz, Stephanie M. Rosales, Ryan S. Mueller, Teresa Sawyer, Adrienne M. S. Correa, Andrew R. Thurber, and Rebecca Vega Thurber. Coral bleaching phenotypes associated with differential abundances of nucleocytoplasmic large DNA viruses. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Martin:2020:FAE

[921] Morgan Jennifer Martin, Tess Gridley, Jean-Paul Roux, and Simon Harvey Elwen. First abundance estimates of Heaviside’s (Cephalorhynchus heavisidii) and dusky (Lagenorhynchus obscurus) dolphins off Namibia using a novel visual and acoustic line transect survey. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Feng:2020:MCB

[922] Ellias Y. Feng, Yvonne Sawall, Marlene Wall, Mario Lebrato, and Yao Fu. Modeling coral bleaching mitigation potential of water vertical translocation — an analogue to geoengineered artificial upwelling. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Horstmann:2021:PDM

[923] Cora Hörstmann, Pier Luigi Buttigieg, Pauline Simpson, Jay Pearlman, and Anya M. Waite. Perspectives on documenting methods to create ocean best practices. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Karakas:2020:SSB

[924] Ferhat Karakas, Jordan Wingate, Leocadio Blanco-Bercial, Amy E. Maas, and David W. Murphy. Swimming and sinking behavior of warm water pelagic snails. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Koner:2020:EIC

[925] Prabhat K. Koner. Enhancing information content in the satellitederived daytime infrared sea surface temperature dataset using a transformative approach. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Archana:2020:MUC

[926] Anand Archana and David Michael Baker. Multifunctionality of an urbanized coastal marine ecosystem. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Schulze:2020:SFM

[927] Melissa J. Schulze, Sophie von der Heyden, David Japp, Larvika Singh, Deon Durholtz, Veronica K. Kapula, Hilkka O. N. Ndjaula, and Romina Henriques. Supporting fisheries management with genomic tools: a case study of kingklip (Genypterus capensis) off Southern Africa. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Garteizgogeascoa:2020:CSE

[928] María Garteizgogeascoa, Lotta C. Kluger, Isabel E. Gonzales, Gerardo Damonte, and Michael Flitner. Contextualizing scenarios to explore social-ecological futures: a three step participatory case study for the Humboldt Current Upwelling System. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Jhugroo:2020:SSL

[929] Khushboo Jhugroo, Joanne O’Callaghan, Craig L. Stevens, Helen S. Macdonald, Fiona Elliott, and Mark Gregory Hadfield. Spatial structure of low salinity submesoscale features and their interactions with a coastal current. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Jevne:2020:ENA

[930] Lone Sunniva Jevne, Silje Forbord, and Yngvar Olsen. The effect of nutrient availability and light conditions on the growth and intracellular nitrogen components of land-based cultivated Saccharina latissima (Phaeophyta). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Gaebel:2020:RSC

[931] Christine Gaebel, Corinne Baulcomb, David E. Johnson, and J. Murray Roberts. Recognising stakeholder conflict and encouraging consensus of ‘science-based management’ approaches for marine biodiversity beyond national jurisdiction (BBNJ). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Kuriyama:2020:ISS

[932] Peter T. Kuriyama, George Sugihara, Andrew R. Thompson, and Brice X. Semmens. Identification of shared spatial dynamics in temperature, salinity, and ichthyoplankton community diversity in the California Current system with empirical dynamic modeling. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

McKenzie:2020:TDI

[933] Trevor McKenzie, Michael Twardowski, Nathan Briggs, Aditya R. Nayak, Kevin M. Boswell, and Fraser Dalgleish. Three-dimensional imaging lidar for characterizing particle fields and organisms in the mesopelagic zone. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

McQuaid:2020:UHC

[934] Kirsty A. McQuaid, Martin J. Attrill, Malcolm R. Clark, Amber Cobley, Adrian G. Glover, Craig R. Smith, and Kerry L. Howell. Using habitat classification to assess representativity of a protected area network in a large, data-poor area targeted for deep-sea mining. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

daCunha:2020:MIS

[935] Divino B. da Cunha, Luis Fernando S. Rodrigues-Filho, João Braúllio de Luna Sales, Pericles Rêgo, Cleonilde Queiroz, Iracilda Sampaio, and Marcelo Vallinoto. Molecular inferences on Scomberomorus brasiliensis, from the Western South Atlantic, based on two mitochondrial genes. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Chasqui:2020:BEU

[936] Luis Chasqui, Katherine Mejía-Quintero, and Juan David González. Biodiversity and ecological units of the mesophotic coral ecosystems in San Andrés Island, SeaFlower Biosphere Reserve. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Schwing:2020:SDB

[937] Patrick T. Schwing, Paul A. Montagna, Samantha B. Joye, Claire B. Paris, Erik E. Cordes, Craig R. McClain, Joshua P. Kilborn, and Steve A. Murawski. A synthesis of deep benthic faunal impacts and resilience following the Deepwater Horizon oil spill. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Rodriguez-Casariego:2020:GWD

[938] Javier A. Rodríguez-Casariego, Alex E. Mercado-Molina, Daniel Garcia-Souto, Ivanna M. Ortiz-Rivera, Christian Lopes, Iliana B. Baums, Alberto M. Sabat, and Jose M. Eirin-Lopez. Genome-wide DNA methylation analysis reveals a conserved epigenetic response to seasonal environmental variation in the staghorn coral Acropora cervicornis. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Both:2020:DSD

[939] Adrianus Both, Carrie J. Byron, Barry Costa-Pierce, Christopher C. Parrish, and Damian C. Brady. Detrital subsidies in the diet of Mytilus edulis; macroalgal detritus likely supplements essential fatty acids. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Villanueva:2020:EPL

[940] Roger Villanueva, Vladimir V. Laptikhovsky, Stuart B. Piertney, Fernando Ángel Fernández-Álvarez, Martin A. Collins, Jonathan D. Ablett, and Alejandro Escánez. Extended pelagic life in a bathybenthic octopus. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Mir:2020:EPL

[941] Ishfaq Nazir Mir, P. P. Srivastava, Irfan Ahmad Bhat, Jaffer Yousuf Dar, and Deepak Agarwal. Expression profiling of lipolytic and longchain fatty acid biosynthesis genes in catfish larvae fed with graded lipid levels. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Gaskins:2020:MSM

[942] Leo C. Gaskins, Avery B. Paxton, and Brian R. Silliman. Megafauna in salt marshes. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Martinez:2020:EDD

[943] Stephane Martinez, Maya Lalzar, Eli Shemesh, Shai Einbinder, Beverly Goodman Tchernov, and Dan Tchernov. Effect of different derivatization protocols on the calculation of trophic position using amino acids compound-specific stable isotopes. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Molina:2020:SDK

[944] Verónica Molina, Lucy Belmar, Héctor A. Levipan, Salvador Ramírez-Flandes, Cristóbal Anguita, Alexander Galán, Ivonne Montes, and Osvaldo Ulloa. Spatiotemporal distribution of key pelagic microbes in a seasonal oxygen-deficient coastal upwelling system of the Eastern South Pacific Ocean. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Cuevas:2020:FSD

[945] Eduardo Cuevas, Nathan F. Putman, Abigail Uribe-Martínez, Melania C. López-Castro, Vicente Guzmán-Hernández, Sandra A. Gallegos-Fernández, María de los Ángeles Liceaga-Correa, Jorge A. Trujillo-Córdova, Raúl de Jesús González-Díaz-Mirón, Ana Negrete-Phillipe, Héctor H. Acosta-Sánchez, Rosa C. Martínez-Portugal, Martha López-Hernández, Patricia Huerta-Rodríguez, and Jim Silver. First spatial distribution analysis of male sea turtles in the Southern Gulf of Mexico. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Beam:2020:MMM

[946] Jacob P. Beam, Sarabeth George, Nicholas R. Record, Peter D. Countway, David T. Johnston, Peter R. Girguis, and David Emerson. Mud, microbes, and macrofauna: Seasonal dynamics of the iron biogeochemical cycle in an intertidal mudflat. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Boschetti:2020:SFC

[947] Fabio Boschetti, Catherine M. Bulman, Alistair J. Hobday, Elizabeth A. Fulton, Stephanie Contardo, Hector Lozano-Montes, Lucy M. Robinson, Anthony D. M. Smith, Joanna Strzelecki, and E. Ingrid van Putten. Sectoral futures are conditional on choices of global and national scenarios — Australian marine examples. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Latsos:2020:ENS

[948] Christos Latsos, Jasper van Houcke, and Klaas R. Timmermans. The effect of nitrogen starvation on biomass yield and biochemical constituents of Rhodomonas sp. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Franco:2020:SOW

[949] Bárbara C. Franco, Vincent Combes, and Victoria González Carman. Subsurface ocean warming hotspots and potential impacts on marine species: The Southwest South Atlantic Ocean case study. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Pages:2020:PEC

[950] Rémi Pagès, Melika Baklouti, Nicolas Barrier, Mohamed Ayache, Florence Sevault, Samuel Somot, and Thierry Moutin. Projected effects of climate-induced changes in hydrodynamics on the biogeochemistry of the Mediterranean Sea under the RCP 8.5 regional climate scenario. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Dougan:2020:NPP

[951] Katherine E. Dougan, Mark C. Ladd, Corinne Fuchs, Rebecca Vega Thurber, Deron E. Burkepile, and Mauricio Rodriguez-Lanetty. Nutrient pollution and predation differentially affect innate immune pathways in the coral Porites porites. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Li:2020:FDB

[952] Yi-Feng Li, Xin Zhu, Zhi-Yang Cheng, Xiao Liang, You-Ting Zhu, Dan-Dan Feng, Sergey Dobretsov, and Jin-Long Yang. 2(5H)-Furanone disrupts bacterial biofilm formation and indirectly reduces the settlement of plantigrades of the mussel Mytilus coruscus. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Rohmana:2020:ITR

[953] Qurratu A’Yunin Rohmana, Andrew M. Fischer, John Cumming, Boyd D. Blackwell, and John Gemmill. Increased transparency and resource prioritization for the management of pollutants from wastewater treatment plants: a national perspective from Australia. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Pazzaglia:2020:DWE

[954] Jessica Pazzaglia, Alex Santillán-Sarmiento, Stephanie B. Helber, Miriam Ruocco, Antonio Terlizzi, Lázaro Marín-Guirao, and Gabriele Procaccini. Does warming enhance the effects of eutrophication in the seagrass Posidonia oceanica? Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Tang:2020:IAA

[955] Jia Tang, Xingzhen Ni, Jianqing Wen, Lingui Wang, Jian Luo, and Zhi Zhou. Increased ammonium assimilation activity in the scleractinian coral Pocillopora damicornis but not its symbiont after acute heat stress. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Mitchell:2020:MPC

[956] Emily G. Mitchell and Simon Harris. Mortality, population and community dynamics of the glass sponge dominated community “The forest of the weird” from the Ridge Seamount, Johnston Atoll, Pacific Ocean. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Heasman:2020:ENZ

[957] Kevin G. Heasman, Nicholas Scott, Jessica A. Ericson, David I. Taylor, and Bela H. Buck. Extending New Zealand’s marine shellfish aquaculture into exposed environments — adapting to modern anthropogenic challenges. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Fu:2020:CEF

[958] Caihong Fu, Yi Xu, Chuanbo Guo, Norm Olsen, Arnaud Grüss, Huizhu Liu, Nicolas Barrier, Philippe Verley, and Yunne-Jai Shin. The cumulative effects of fishing, plankton productivity, and marine mammal consumption in a marine ecosystem. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Lui:2020:IKH

[959] Hon-Kit Lui, Chen-Tung Arthur Chen, Wei-Ping Hou, Jian-Ming Liau, Wen-Chen Chou, You-Lin Wang, Chau-Ron Wu, Jay Lee, Yi-Chia Hsin, and Yan-Yu Choi. Intrusion of Kuroshio helps to diminish coastal hypoxia in the coast of Northern South China Sea. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Katsanevakis:2020:TRA

[960] Stelios Katsanevakis, Marta Coll, Simonetta Fraschetti, Sylvaine Giakoumi, David Goldsborough, Vesna Mačić, Peter Mackelworth, Gil Rilov, Vanessa Stelzenmüller, Paolo G. Albano, Amanda E. Bates, Stanislao Bevilacqua, Elena Gissi, Virgilio Hermoso, Antonios D. Mazaris, Cristina Pita, Valentina Rossi, Yael Teff-Seker, and Katherine Yates. Twelve recommendations for advancing marine conservation in European and contiguous seas. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Cardin:2020:SAR

[961] Vanessa Cardin, Achim Wirth, Maziar Khosravi, and Miroslav Gačić. South Adriatic recipes: Estimating the vertical mixing in the Deep Pit. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Guimaraes-Costa:2020:DBA

[962] Aurycéia Guimarães-Costa, Fabíola S. Machado, José Amorim Reis-Filho, Marcelo Andrade, Rafael Guilherme Araújo, Eglé Miranda Ramos Corrêa, Iracilda Sampaio, and Tommaso Giarrizzo. DNA barcoding for the assessment of the taxonomy and conservation status of the fish bycatch of the Northern Brazilian shrimp trawl fishery. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Alava:2020:MBB

[963] Juan José Alava. Modeling the bioaccumulation and biomagnification potential of microplastics in a cetacean foodweb of the Northeastern Pacific: a prospective tool to assess the risk exposure to plastic particles. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Troselj:2021:DDC

[964] Joško Trošelj, Junichi Ninomiya, Satoshi Takewaka, and Nobuhito Mori. Dynamical downscaling of coastal dynamics for two extreme storm surge events in Japan. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Mazzoccoli:2020:MPM

[965] Michela Mazzoccoli, Marco Altosole, Veronica Vigna, Barbara Bosio, and Elisabetta Arato. Marine pollution mitigation by waste oils recycling onboard ships: Technical feasibility and need for new policy and regulations. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Logan:2020:SMH

[966] Ryan K. Logan, Jeremy J. Vaudo, Lara L. Sousa, Mark Sampson, Bradley M. Wetherbee, and Mahmood S. Shivji. Seasonal movements and habitat use of juvenile smooth hammerhead sharks in the Western North Atlantic Ocean and significance for management. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Downey-Wall:2020:OAI

[967] Alan M. Downey-Wall, Louise P. Cameron, Brett M. Ford, Elise M. McNally, Yaamini R. Venkataraman, Steven B. Roberts, Justin B. Ries, and Katie E. Lotterhos. Ocean acidification induces subtle shifts in gene expression and DNA methylation in mantle tissue of the eastern oyster (Crassostrea virginica). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Martinez:2020:ESD

[968] Stephane Martinez, Yuval Kolodny, Eli Shemesh, Federica Scucchia, Reinat Nevo, Smadar Levin-Zaidman, Yossi Paltiel, Nir Keren, Dan Tchernov, and Tali Mass. Energy sources of the depth-generalist mixotrophic coral Stylophora pistillata. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Lloret-Lloret:2020:SDH

[969] Elena Lloret-Lloret, Joan Navarro, Joan Giménez, Nieves López, Marta Albo-Puigserver, Maria Grazia Pennino, and Marta Coll. The seasonal distribution of a highly commercial fish is related to ontogenetic changes in its feeding strategy. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Williams:2020:FDS

[970] Alan Williams, Franziska Althaus, Kylie Maguire, Mark Green, Candice Untiedt, Phil Alderslade, Malcolm R. Clark, Nicholas Bax, and Thomas A. Schlacher. The fate of deep-sea coral reefs on seamounts in a fishery-seascape: What are the impacts, what remains, and what is protected? Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Allen:2020:COC

[971] James G. Allen, David A. Siegel, Norman B. Nelson, and Stuart Halewood. Controls on ocean color spectra observed during the North Atlantic aerosols and marine ecosystems study (NAAMES). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Oliveira:2020:MEP

[972] Juliana Oliveira, Afonso Belchior, Verônica D. da Silva, Ana Rotter, Željko Petrovski, Pedro L. Almeida, Nídia D. Lourenço, and Susana P. Gaudêncio. Marine environmental plastic pollution: Mitigation by microorganism degradation and recycling valorization. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Ruhl:2020:IAP

[973] Saskia Rühl, Charlie E. L. Thompson, Ana M. Queirós, and Stephen Widdicombe. Intra-annual patterns in the benthic-pelagic fluxes of dissolved and particulate matter. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Knights:2020:OAM

[974] Antony M. Knights, Matthew J. Norton, Anaëlle J. Lemasson, and Natasha Stephen. Ocean acidification mitigates the negative effects of increased sea temperatures on the biomineralization and crystalline ultrastructure of Mytilus. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Wagner:2020:CRH

[975] Daniel Wagner, Alan M. Friedlander, Richard L. Pyle, Cassandra M. Brooks, Kristina M. Gjerde, and T. ‘Aulani Wilhelm. Coral reefs of the High Seas: Hidden biodiversity hotspots in need of protection. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Friedland:2020:CPC

[976] Kevin D. Friedland, Ryan E. Morse, Nancy Shackell, Jamie C. Tam, Janelle L. Morano, John R. Moisan, and Damian C. Brady. Changing physical conditions and lower and upper trophic level responses on the US Northeast Shelf. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

McMahan:2020:GVL

[977] Marissa D. McMahan, Graham D. Sherwood, and Jonathan H. Grabowski. Geographic variation in life-history traits of Black Sea bass (Centropristis striata) during a rapid range expansion. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Veron:2020:DTV

[978] Matthieu Véron, Erwan Duhamel, Michel Bertignac, Lionel Pawlowski, Martin Huret, and Loïc Baulier. Determinism of temporal variability in size at maturation of sardine Sardina pilchardus in the Bay of Biscay. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Coll:2020:AGE

[979] Marta Coll, Jeroen Steenbeek, Maria Grazia Pennino, Joe Buszowski, Kristin Kaschner, Heike K. Lotze, Yannick Rousseau, Derek P. Tittensor, Carl Walters, Reg A. Watson, and Villy Christensen. Advancing global ecological modeling capabilities to simulate future trajectories of change in marine ecosystems. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Somoza:2020:MSC

[980] Luis Somoza, Teresa Medialdea, Francisco J. González, António Calado, Andreia Afonso, Mónica Albuquerque, María Asensio-Ramos, Renato Bettencourt, Iker Blasco, Jose A. Candón, Marina Carreiro-Silva, Constantino Cid, Cristina De Ignacio, Enrique López-Pamo, Sara Machancoses, Bruno Ramos, Luisa Pinto Ribeiro, Blanca Rincón-Tomás, Esther Santofimia, Miguel Souto, Inês Tojeira, Cláudia Viegas, and Pedro Madureira. Multidisciplinary scientific cruise to the Northern Mid-Atlantic Ridge and Azores Archipelago. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Kreiss:2020:FSP

[981] Cornelia M. Kreiss, Eleni Papathanasopoulou, Katell G. Hamon, John K. Pinnegar, Sandra Rybicki, Giulia Micallef, Andrzej Tabeau, Alhambra M. Cubillo, and Myron A. Peck. Future socio-political scenarios for aquatic resources in Europe: an operationalized framework for aquaculture projections. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Pinnegar:2021:FSP

[982] John K. Pinnegar, Katell G. Hamon, Cornelia M. Kreiss, Andrzej Tabeau, Sandra Rybicki, Eleni Papathanasopoulou, Georg H. Engelhard, Tyler D. Eddy, and Myron A. Peck. Future socio-political scenarios for aquatic resources in Europe: a common framework based on shared-socioeconomic-pathways (SSPs). Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Travers-Trolet:2020:RDF

[983] Morgane Travers-Trolet, Pierre Bourdaud, Mathieu Genu, Laure Velez, and Youen Vermard. The risky decrease of fishing reference points under climate change. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Vargas-Yanez:2020:CHP

[984] Manuel Vargas-Yáñez, Melanie Juza, Rosa Balbín, Pedro Velez-Belchí, M. Carmen García-Martínez, Francina Moya, and Alonso Hernández-Guerra. Climatological hydrographic properties and water mass transports in the Balearic Channels from repeated observations over 1996– 2019. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Tinker-Kulberg:2020:EIO

[985] Rachel Tinker-Kulberg, Anthony L. Dellinger, Lisa C. Gentit, Bryan A. Fluech, Cole A. Wilder, Ivy L. Spratling, David J. Stasek, Christopher L.

Kepley, Lee Robertson, Melinda K. Goddard, Terry E. Brady, Lance Tōland, and Kristen Dellinger. Evaluation of indoor and outdoor aquaculture systems as alternatives to harvesting hemolymph from random wild capture of horseshoe crabs. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Dunlop:2020:EHA

[986] Katherine Dunlop, Astrid Harendza, Liv Plassen, and Nigel Keeley. Epifaunal habitat associations on mixed and hard bottom substrates in coastal waters of Northern Norway. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Benavides:2020:BSG

[987] Mar Benavides and Julie Robidart. Bridging the spatiotemporal gap in diazotroph activity and diversity with high-resolution measurements. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Quemmerais-Amice:2020:MTM

[988] Frédéric Quemmerais-Amice, Julien Barrere, Marie La Rivière, Gabriel Contin, and Denis Bailly. A methodology and tool for mapping the risk of cumulative effects on benthic habitats. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Stenson:2020:HSM

[989] Garry B. Stenson, Tore Haug, and Mike O. Hammill. Harp seals: Monitors of change in differing ecosystems. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Chen:2020:NPB

[990] Wan-Lin Chen, Kuo-Ping Chiang, and Sheng-Fang Tsai. Neglect of presence of bacteria leads to inaccurate growth parameters of the oligotrich ciliate Strombidium sp. during grazing experiments on nanoflagellates. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Nef:2020:CSM

[991] Charlotte Nef, Céline Henry, Élodie Nicolau, Jean-Baptiste Bérard, Fabienne Hervé, Amandine M. N. Caruana, Raymond Kaas, Francis Mairet, and Matthieu Garnier. Cobalamin scarcity modifies carbon allocation and impairs DMSP production through methionine metabolism in the haptophyte microalgae Tisochrysis lutea. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Valdivia:2021:HDS

[992] Nelson Valdivia, Moisés A. Aguilera, and Bernardo R. Broitman. High dimensionality of the stability of a marine benthic ecosystem. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Zhang:2020:SPR

[993] Shuwen Zhang, Cheng Li, Shunyan Cheung, Mingming Sun, Shuqun Song, Wang Guo, Cui Guo, Gongguo Wu, and Hongbin Liu. Snapshot of peptidomics of the Red Tide forming species Noctiluca scintillans. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Mayers:2020:PCF

[994] Kyle M. J. Mayers, Alex J. Poulton, Kay Bidle, Kimberlee Thamatrakoln, Brittany Schieler, Sarah L. C. Giering, Seona R. Wells, Glen A. Tarran, Dan Mayor, Matthew Johnson, Ulf Riebesell, Aud Larsen, Assaf Vardi, and Elizabeth L. Harvey. The possession of coccoliths fails to deter microzooplankton grazers. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Sahri:2020:UCE

[995] Achmad Sahri, Putu Liza Kusuma Mustika, Purwanto Purwanto, Albertinka J. Murk, and Meike Scheidat. Using cost-effective surveys from platforms of opportunity to assess cetacean occurrence patterns for marine park management in the heart of the Coral Triangle. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Lim:2020:SGC

[996] Kah Kheng Lim, Susann Rossbach, Nathan R. Geraldi, Sebastian Schmidt-Roach, Ester A. Serrão, and Carlos M. Duarte. The small giant clam, Tridacna maxima exhibits minimal population genetic structure in the Red Sea and genetic differentiation from the Gulf of Aden. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Kamenos:2020:CCR

[997] Nicholas A. Kamenos and Sebastian J. Hennige. Commentary: Commentary on Reconstructing Four Centuries of Temperature-Induced Bleaching on the Great Barrier Reef by Hoegh-Guldberg et al. 2019 and DeCarlo 2020. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745. See [134].

Anderson:2020:NPG

[998] Giulia Anderson, Monal Lal, Brian Stockwell, John Hampton, Neville Smith, Simon Nicol, and Ciro Rico. No population genetic structure of skipjack tuna (Katsuwonus pelamis) in the tropical Western and Central Pacific assessed using single nucleotide polymorphisms. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Li:2020:CDR

[999] Jing-Jing Li, Sheng-Hui Huang, Zheng-Yi Liu, and Yuan-Xin Bi. Climate-driven range shifts of brown seaweed Sargassum horneri in the Northwest Pacific. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Ford:2020:LHI

[1000] Amanda K. Ford, Jean-Baptiste Jouffray, Albert V. Norström, Bradley R. Moore, Maggy M. Nugues, Gareth J. Williams, Sonia Bejarano, Franck Magron, Christian Wild, and Sebastian C. A. Ferse. Local human impacts disrupt relationships between benthic reef assemblages and environmental predictors. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Fulham:2020:DDE

[1001] Mariel Fulham, Michelle Power, and Rachael Gray. Diversity and distribution of Escherichia coli in three species of free-ranging Australian pinniped pups. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Morgan:2020:LLD

[1002] Kyle M. Morgan, Molly A. Moynihan, Nivedita Sanwlani, and Adam D. Switzer. Light limitation and depth-variable sedimentation drives vertical reef compression on turbid coral reefs. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Seelmann:2020:EOC

[1003] Katharina Seelmann, Tobias Steinhoff, Steffen Aßmann, and Arne Körtzinger. Enhance ocean carbon observations: Successful implementation of a novel autonomous total alkalinity analyzer on a ship of opportunity. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Jack:2020:HMA

[1004] Maria Eugenia Molina Jack, Rigers Bakiu, Ana Castelli, Branko Čermelj, Maja Fafan– del, Christina Georgopoulou, Giordano Giorgi, Athanasia Iona, Damir Ivankovic, Martina Kralj, Elena Partescano, Alice Rotini, Melita Velikonja, and Marina Lipizer. Heavy metals in the Adriatic– Ionian Seas: a case study to illustrate the challenges in data management when dealing with regional datasets. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Uffman-Kirsch:2020:NPS

[1005] Lisa B. Uffman-Kirsch, Benjamin J. Richardson, and Elizabeth Ingrid van Putten. A new paradigm for social license as a path to marine sustainability. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Dellisanti:2020:DPR

[1006] Walter Dellisanti, Ryan H. L. Tsang, Put Ang, Jiajun Wu, Mark L. Wells, and Leo L. Chan. A diver-portable respirometry system for in-situ shortterm measurements of coral metabolic health and rates of calcification. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Hauck:2020:CCO

[1007] Judith Hauck, Moritz Zeising, Corinne Le Quéré, Nicolas Gruber, Dorothee C. E. Bakker, Laurent Bopp, Thi Tuyet Trang Chau, Özgür Gürses, Tatiana Ilyina, Peter Landschützer, Andrew Lenton, Laure Resplandy, Christian Rödenbeck, Jörg Schwinger, and Roland Séférian. Consistency and challenges in the ocean carbon sink estimate for the global carbon budget. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Kim:2020:USS

[1008] Sang Wha Kim, Won Hee Hong, Se Jin Han, Jun Kwon, Heejun Ko, Sung Bin Lee, Sib Sankar Giri, Sang Guen Kim, Byung Yeop Kim, Goo Jang, Byeong Chun Lee, Dong Wan Kim, and Se Chang Park. Use of synthetic salmon GnRH and domperidone (ovaprim(R)) in sharks: Preparation for ex situ conservation. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See corrigendum [1354].

Appah:2020:NRH

[1009] J. K. M. Appah, A. Lim, K. Harris, R. O’Riordan, L. O’Reilly, and A. J. Wheeler. Are non-reef habitats as important to benthic diversity and composition as coral reef and rubble habitats in submarine canyons? Analysis of controls on benthic megafauna distribution in the Porcupine Bank Canyon, NE Atlantic. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

deMitcheson:2020:DMI

[1010] Yvonne Sadovy de Mitcheson, Patrick L. Colin, Steven J. Lindfield, and Asap Bukurrou. A decade of monitoring an Indo–Pacific grouper spawning aggregation: Benefits of protection and importance of survey design. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Sarker:2020:ATL

[1011] Indrani Sarker, Lisa R. Moore, Ian T. Paulsen, and Sasha G. Tetu. Assessing the toxicity of leachates from weathered plastics on photosynthetic marine bacteria prochlorococcus. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Tengfei:2021:MFN

[1012] Du Tengfei, Xiao Yongshuang, Zhao Haixia, Zhou Li, Liu Qinghua, Wang Xueying, Li Jun, Xu Shihong, Wang Yanfeng, Yu Jiachen, Wu Lele, Wang Yunong, and Gao Guang. Multiple fetal nutritional patterns before parturition in viviparous fish Sebastes schlegelii (Hilgendorf, 1880). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Scheffold:2020:QCO

[1013] Maike Iris Esther Scheffold and Inga Hense. Quantifying contemporary organic carbon stocks of the Baltic Sea ecosystem. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Vikstrom:2020:SIB

[1014] Kevin Vikström, Ines Bartl, Jan Karlsson, and Johan Wikner. Strong influence of baseline respiration in an oligotrophic coastal ecosystem. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Nayak:2021:RHA

[1015] Aditya R. Nayak, Ed Malkiel, Malcolm N. McFarland, Michael S. Twardowski, and James M. Sullivan. A review of holography in the aquatic sciences: In situ characterization of particles, plankton, and small scale biophysical interactions. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Perez:2020:DSH

[1016] Jose Angel A. Perez, Lucas Gavazzoni, Luis Henrique P. de Souza, Paulo Y. Gomes Sumida, and Hiroshi Kitazato. Deep-sea habitats and megafauna on the slopes of the São Paulo Ridge, SW Atlantic. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Hall:2020:RRS

[1017] Megan R. Hall, James W. Moffett, and Andrew Y. Gracey. RNAseq reveals sensitive, concentration-dependent transcriptional markers of copper in Mytilus californianus larvae and adults. Frontiers in Marine Science, 7, August 2020. ISSN 2296-7745.

Obst:2020:MBO

[1018] Matthias Obst, Katrina Exter, A. Louise Allcock, Christos Arvanitidis, Alizz Axberg, Maria Bustamante, Ibon Cancio, Diego Carreira-Flores, Eva Chatzinikolaou, Giorgos Chatzigeorgiou, Nathan Chrismas, Melody S. Clark, Thierry Comtet, Thanos Dailianis, Neil Davies, Klaas Deneudt, Oihane Diaz de Cerio, Ana Fortič, Vasilis Gerovasileiou, Pascal I. Hablützel, Kleoniki Keklikoglou, Georgios Kotoulas, Rafal Lasota, Barbara R. Leite, Stéphane Loisel, Laurent Lévêque, Liraz Levy, Magdalena Malachowicz, Borut Mavrič, Christopher Meyer, Jonas Mortelmans, Joanna Norkko, Nicolas Pade, Anne Marie Power, Andreja Ramšak, Henning Reiss, Jostein Solbakken, Peter A. Staehr, Per Sundberg, Jakob Thyrring, Jesus S. Troncoso, Frédérique Viard, Roman Wenne, Eleni Ioanna Yperifanou, Malgorzata Zbawicka, and Christina Pavloudi. A marine biodiversity observation network for genetic monitoring of hard-bottom communities (ARMS-MBON). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Stigebrandt:2020:EBS

[1019] Anders Stigebrandt and Ambjörn Andersson. The eutrophication of the Baltic Sea has been boosted and perpetuated by a major internal phosphorus source. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Maas:2020:MZE

[1020] Amy E. Maas, Shuting Liu, Luis M. Bolaños, Brittany Widner, Rachel Parsons, Elizabeth B. Kujawinski, Leocadio Blanco-Bercial, and Craig A. Carlson. Migratory zooplankton excreta and its influence on prokaryotic communities. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Golubkov:2020:EEG

[1021] Mikhail S. Golubkov, Larisa F. Litvinchuk, and Sergey M. Golubkov. Effects of environmental gradients on the biomass of alien Cercopagis pengoi in the Neva Estuary. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Degenhardt:2020:SDM

[1022] Julius Degenhardt, Leon Dlugosch, Janis Ahrens, Melanie Beck, Hannelore Waska, and Bert Engelen. Seasonal dynamics of microbial diversity at a sandy high energy beach reveal a resilient core community. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Randle:2020:SCT

[1023] Janna L. Randle, Anny Cárdenas, Hagen M. Gegner, Maren Ziegler, and Christian R. Voolstra. Salinity-conveyed thermotolerance in the coral model aiptasia is accompanied by distinct changes of the bacterial microbiome. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Clavel-Henry:2020:SDA

[1024] Morane Clavel-Henry, Chiara Piroddi, Federico Quattrocchi, Diego Macias, and Villy Christensen. Spatial distribution and abundance of mesopelagic fish biomass in the Mediterranean Sea. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Jefferson:2020:ATI

[1025] Thomas A. Jefferson and Jeffrey E. Moore. Abundance and trends of Indo–Pacific finless porpoises (Neophocaena phocaenoides) in Hong Kong waters, 1996–2019. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Luna:2021:GCW

[1026] Fábia de Oliveira Luna, Caitlin E. Beaver, Coralie Nourisson, Robert K. Bonde, Fernanda L. N. Attademo, Adriana Vieira Miranda, Juan Pablo Torres-Florez, Glaucia Pereira de Sousa, José Zanon Passavante, and Margaret E. Hunter. Genetic connectivity of the West Indian manatee in the southern range and limited evidence of hybridization with Amazonian manatees. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Bakir:2020:MCI

[1027] Adil Bakir, Carl D. van der Lingen, Fiona Preston-Whyte, Ashok Bali, Yonela Geja, Jon Barry, Yandiswa Mdazuka, Gcobani Mooi, Denise Doran, Freya Tooley, Rogan Harmer, and Thomas Maes. Microplastics in commercially important small pelagic fish species from South Africa. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Yue:2020:RSS

[1028] Shidong Yue, Xiaomei Zhang, Shaochun Xu, Yu Zhang, Peng Zhao, Xiaodong Wang, and Yi Zhou. Reproductive strategies of the seagrass Zostera japonica under different geographic conditions in Northern China. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Kuroda:2020:USS

[1029] Hiroshi Kuroda, Toshihiko Saito, Toshiki Kaga, Akinori Takasuka, Yasuhiro Kamimura, Sho Furuichi, and Takuya Nakanowatari. Unconventional sea surface temperature regime around Japan in the 2000s–2010s:

Potential influences on major fisheries resources. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Bolinesi:2020:SRC

[1030] Francesco Bolinesi, Maria Saggiomo, Francisco Ardini, Pasquale Castagno, Angelina Cordone, Giannetta Fusco, Paola Rivaro, Vincenzo Saggiomo, and Olga Mangoni. Spatial-related community structure and dynamics in phytoplankton of the Ross Sea, Antarctica. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Jakobsson-Thor:2020:CIT

[1031] Stina Jakobsson-Thor, Janina Brakel, Gunilla B. Toth, and Henrik Pavia. Complex interactions of temperature, light and tissue damage on seagrass wasting disease in Zostera marina. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Bajjouk:2020:HLC

[1032] Touria Bajjouk, Cecile Jauzein, Lucas Drumetz, Mauro Dalla Mura, Audrey Duval, and Stanislas F. Dubois. Hyperspectral and lidar: Complementary tools to identify benthic features and assess the ecological status of Sabellaria alveolata reefs. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Ghaderiardakani:2020:MDA

[1033] Fatemeh Ghaderiardakani, Maria Liliana Quartino, and Thomas Wichard. Microbiome-dependent adaptation of seaweeds under environmental stresses: a perspective. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Warren:2021:PAM

[1034] Victoria E. Warren, Ana Širović, Craig McPherson, Kimberly T. Goetz, Craig A. Radford, and Rochelle Constantine. Passive acoustic monitoring reveals spatio-temporal distributions of Antarctic and pygmy blue whales around Central New Zealand. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ryan:2020:TBB

[1035] Peter G. Ryan, Eleanor A. Weideman, Vonica Perold, and Coleen L. Moloney. Toward balancing the budget: Surface macro-plastics dominate the mass of particulate pollution stranded on beaches. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Hajializadeh:2020:SCF

[1036] Parima Hajializadeh, Mohsen Safaie, Reza Naderloo, Mehdi Ghodrati Shojaei, Johanna Gammal, Anna Villnäs, and Alf Norkko. Species composition and functional traits of macrofauna in different mangrove habitats in the Persian Gulf. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Brigolin:2021:EDS

[1037] Daniele Brigolin, Christophe Rabouille, Clément Demasy, Bruno Bombled, Gaël Monvoisin, and Roberto Pastres. Early diagenesis in sediments of the Venice Lagoon (Italy) and its relationship to hypoxia. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Meekan:2020:AGW

[1038] Mark G. Meekan, Brett M. Taylor, Emily Lester, Luciana C. Ferreira, Ana M. M. Sequeira, Alistair D. M. Dove, Matthew J. Birt, Alex Aspinall, Kim Brooks, and Michele Thums. Asymptotic growth of whale sharks suggests sex-specific life-history strategies. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Endo:2020:HSP

[1039] Hikaru Endo, Eri Inomata, Xu Gao, Junji Kinoshita, Yoichi Sato, and Yukio Agatsuma. Heat stress promotes nitrogen accumulation in meristems via apical blade erosion in a brown macroalga with intercalary growth. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Singh:2020:NDM

[1040] Rabinder Singh, Chetan Paliwal, Asha A. Nesamma, Alka Narula, and Pannaga P. Jutur. Nutrient deprivation mobilizes the production of unique tocopherols as a stress-promoting response in a new indigenous isolate Monoraphidium sp. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Simon-Nutbrown:2020:SDM

[1041] Cornelia Simon-Nutbrown, Peter M. Hollingsworth, Teresa F. Fernandes, Lisa Kamphausen, John M. Baxter, and Heidi L. Burdett. Species distribution modeling predicts significant declines in coralline algae populations under projected climate change with implications for conservation policy. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Ajani:2020:GWI

[1042] Penelope A. Ajani, Claire H. Davies, Ruth S. Eriksen, and Anthony J. Richardson. Global warming impacts micro-phytoplankton at a longterm Pacific Ocean coastal station. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Landsberg:2020:SCT

[1043] Jan H. Landsberg, Yasunari Kiryu, Esther C. Peters, Patrick W. Wilson, Noretta Perry, Yvonne Waters, Kerry E. Maxwell, Lindsay K. Huebner, and Thierry M. Work. Stony Coral Tissue Loss Disease in Florida is associated with disruption of host–zooxanthellae physiology. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Gao:2020:MMP

[1044] Tianxiang Gao, Yiping Ying, Qiaoli Yang, Na Song, and Yongshuang Xiao. The mitochondrial markers provide new insights into the population demographic history of Coilia nasus with two ecotypes (anadromous and freshwater). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Barrett:2020:MPD

[1045] Justine Barrett, Zanna Chase, Jing Zhang, Mark M. Banaszak Holl, Kathryn Willis, Alan Williams, Britta D. Hardesty, and Chris Wilcox. Microplastic pollution in deep-sea sediments from the Great Australian Bight. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Miyama:2021:MHS

[1046] Toru Miyama, Shoshiro Minobe, and Hanako Goto. Marine heatwave of sea surface temperature of the Oyashio region in summer in 2010–2016. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ferreira:2020:CPC

[1047] Afonso Ferreira, Raul R. Costa, Tiago S. Dotto, Rodrigo Kerr, Virginia M. Tavano, Ana C. Brito, Vanda Brotas, Eduardo R. Secchi, and Carlos R. B. Mendes. Changes in phytoplankton communities along the Northern Antarctic Peninsula: Causes, impacts and research priorities. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Perry:2020:SHI

[1048] Cameron T. Perry, Elizabeth Clingham, D. Harry Webb, Rafael de la Parra, Simon J. Pierce, Annalea Beard, LeeAnn Henry, Beth Taylor, Kenickie Andrews, Rhys Hobbs, Gonzalo Araujo, and Alistair D. M.

Dove. St. Helena: an important reproductive habitat for whale sharks (Rhincodon typus) in the Central South Atlantic. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Watson:2020:MSW

[1049] Jordan T. Watson, Diana L. Stram, and Jim Harmon. Mitigating seafood waste through a bycatch donation program. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Martini:2020:BUS

[1050] Séverine Martini, Darrin T. Schultz, Lonny Lundsten, and Steven H. D. Haddock. Bioluminescence in an undescribed species of carnivorous sponge (Cladorhizidae) from the deep sea. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Pastor:2020:STM

[1051] Ane Pastor, Janus Larsen, Christian Mohn, Camille Saurel, Jens Kjerulf Petersen, and Marie Maar. Sediment transport model quantifies plume length and light conditions from mussel dredging. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Bishop:2020:IMC

[1052] Amanda M. Bishop, Casey L. Brown, Renae Sattler, and Markus Horning. An integrative method for characterizing marine habitat features associated with predation: a case study on juvenile Steller sea lions (Eumetopias jubatus). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Aoki:2021:SRF

[1053] Lillian R. Aoki, Karen J. McGlathery, Patricia L. Wiberg, Matthew P. J. Oreska, Amelie C. Berger, Peter Berg, and Robert J. Orth. Seagrass recovery following marine heat wave influences sediment carbon stocks. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Gerringa:2020:DTM

[1054] Loes J. A. Gerringa, Anne-Carlijn Alderkamp, Gert van Dijken, Patrick Laan, Rob Middag, and Kevin R. Arrigo. Dissolved trace metals in the Ross Sea. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

LeTraon:2020:PNP

[1055] Pierre-Yves Le Traon, Fabrizio D’Ortenzio, Marcel Babin, Edouard Leymarie, Claudie Marec, Sylvie Pouliquen, Virginie Thierry, Cecile Cabanes, Hervé Claustre, Damien Desbruyères, Leo Lacour, Jose-Luis Lagunas, Guillaume Maze, Herle Mercier, Christophe Penkerc’h, Noe Poffa, Antoine Poteau, Louis Prieur, Virginie Racapé, Achim Randelhoff, Eric Rehm, Catherine Marie Schmechtig, Vincent Taillandier, Thibaut Wagener, and Xiaogang Xing. Preparing the new phase of Argo: Scientific achievements of the NAOS Project. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Prieto:2020:VCA

[1056] Daniel Prieto, Kristina Arranz, Iñaki Urrutxurtu, Enrique Navarro, Miren Bego Urrutia, and Irrintzi Ibarrola. Variable capacity for acute and chronic thermal compensation of physiological rates contributes to interindividual differences in growth rate in mussels (Mytilus galloprovincialis). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Andre:2020:PNP

[1057] Xavier André, Pierre-Yves Le Traon, Serge Le Reste, Vincent Dutreuil, Edouard Leymarie, Damien Malardé, Claudie Marec, Jérôme Sagot, Martin Amice, Marcel Babin, Hervé Claustre, Arnaud David, Fabrizio D’Ortenzio, Nicolas Kolodziejczyk, José Luis Lagunas, Marc Le Menn, Bertrand Moreau, David Nogré, Christophe Penkerc’h, Antoine Poteau, Corentin Renaut, Christophe Schaeffer, Vincent Taillandier, and Virginie Thierry. Preparing the new phase of Argo: Technological developments on profiling floats in the NAOS Project. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Houlihan:2020:DBL

[1058] Erin P. Houlihan, Nadjejda Espinel-Velasco, Christopher E. Cornwall, Conrad A. Pilditch, and Miles D. Lamare. Diffusive boundary layers and ocean acidification: Implications for sea urchin settlement and growth. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Pethybridge:2020:CFA

[1059] Heidi R. Pethybridge, Elizabeth A. Fulton, Alistair J. Hobday, Julia Blanchard, Catherine M. Bulman, Ian R. Butler, William W. L. Cheung, Leo X. C. Dutra, Rebecca Gorton, Trevor Hutton, Richard Matear, Hector Lozano-Montes, Eva E. Plagányi, Cecilia Villanueva, and Xuebin Zhang. Contrasting futures for Australia’s fisheries stocks under IPCC RCP8.5 emissions — a multi-ecosystem model approach. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Obregon:2020:WYS

[1060] Clara Obregón, Ryan Admiraal, Ingrid van Putten, Michael Hughes, James R. Tweedley, and Neil R. Loneragan. Who you speak to matters: Information sharing and the management of a small-scale fishery. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Figueiras:2020:SVM

[1061] Francisco G. Figueiras, Isabel G. Teixeira, María Froján, Diana Zúñiga, Belén Arbones, and Carmen G. Castro. Seasonal variability in the microbial plankton community in a semienclosed bay affected by upwelling: The role of a nutrient trap. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Yati:2020:MEV

[1062] Emi Yati, Shoshiro Minobe, Nathan Mantua, Shin ichi Ito, and Emanuele Di Lorenzo. Marine ecosystem variations over the North Pacific and their linkage to large-scale climate variability and change. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Mattsson:2021:NEP

[1063] Karin Mattsson, Frida Björkroth, Therese Karlsson, and Martin Hassellöv. Nanofragmentation of expanded polystyrene under simulated environmental weathering (thermooxidative degradation and hydrodynamic turbulence). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Malick:2020:EDS

[1064] Michael J. Malick, Samantha A. Siedlecki, Emily L. Norton, Isaac C. Kaplan, Melissa A. Haltuch, Mary E. Hunsicker, Sandra L. Parker-Stetter, Kristin N. Marshall, Aaron M. Berger, Albert J. Hermann, Nicholas A. Bond, and Stéphane Gauthier. Environmentally driven seasonal forecasts of Pacific hake distribution. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Bravo:2020:ISO

[1065] Gonzalo Bravo, Juan Pablo Livore, and Gregorio Bigatti. The importance of surface orientation in biodiversity monitoring protocols: The case of Patagonian rocky reefs. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Hernvann:2020:CST

[1066] Pierre-Yves Hernvann, Didier Gascuel, Arnaud Grüss, Jean-Noël Druon, Dorothée Kopp, Ilan Perez, Chiara Piroddi, and Marianne Robert. The Celtic Sea through time and space: Ecosystem modeling to unravel fishing and climate change impacts on food-web structure and dynamics. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Rodde:2020:PTF

[1067] Charles Rodde, Marc Vandeputte, François Allal, Mathieu Besson, Frédéric Clota, Alain Vergnet, John A. H. Benzie, and Hugues de Verdal.

Population, temperature and feeding rate effects on individual feed efficiency in European sea bass (Dicentrarchus labrax). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Ghaderpour:2020:EST

[1068] Sirwe Ghaderpour and Alicia Estevez. Effect of short-term rotifer enrichment with marine phospholipids on growth, survival, and composition of meager (Argyrosomus regius) larvae. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Steneck:2020:RCM

[1069] Robert S. Steneck, Suzanne N. Arnold, Robert Boenish, Ramón de León, Peter J. Mumby, Douglas B. Rasher, and Margaret W. Wilson. Response: Commentary: Managing recovery resilience in coral reefs against climateinduced bleaching and hurricanes: a 15 year case study from Bonaire, Dutch Caribbean. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See [448].

Crear:2020:ESP

[1070] Daniel P. Crear, Brian E. Watkins, Marjorie A. M. Friedrichs, Pierre St-Laurent, and Kevin C. Weng. Estimating shifts in phenology and habitat use of cobia in Chesapeake Bay under climate change. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Bamunawala:2020:PAI

[1071] Janaka Bamunawala, Ali Dastgheib, Roshanka Ranasinghe, Ad van der Spek, Shreedhar Maskey, A. Brad Murray, Patrick L. Barnard, Trang Minh Duong, and T. A. J. G. Sirisena. Probabilistic application of an integrated catchment-estuary-coastal system model to assess the evolution of inlet-interrupted coasts over the 21st Century. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Mackay:2020:MMB

[1072] Mary Mackay, Elizabeth Ingrid van Putten, Satoshi Yamazaki, Sarah Jennings, and Hugh Sibly. Me and my behavior: an experiment on individual characteristics and compliance behavior in recreational fishing. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Li:2020:TMP

[1073] Yundong Li, Falin Zhou, Jianhua Huang, Lishi Yang, Song Jiang, Qibin Yang, and Shigui Jiang. Transcriptome and miRNA profiles of black tiger shrimp, Penaeus monodon, under different salinity conditions. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Watson:2020:WWD

[1074] Sally J. Watson, Helen Neil, Marta Ribó, Geoffroy Lamarche, Lorna J. Strachan, Kevin MacKay, Steve Wilcox, Tim Kane, Alan Orpin, Scott Nodder, Arne Pallentin, and Tilmann Steinmetz. What we do in the shallows: Natural and anthropogenic seafloor geomorphologies in a drowned river valley, New Zealand. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Shin:2020:IAC

[1075] JiYun Shin, Suyun Noh, and SungHyun Nam. Intraseasonal abyssal current variability of bottom-trapped topographic Rossby waves in the Southwestern East Sea (Japan Sea). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Qu:2020:HPG

[1076] Zhe Qu, Thomas C. N. Leung, Wenyan Nong, Ho Yin Yip, Ivy H. T. Lee, Siu Gin Cheung, Ngai Sai Ming, Wai Lok So, William G. Bendena, Stephen S. Tobe, and Jerome H. L. Hui. Hemolymph proteomics and gut microbiota of horseshoe crabs Tachypleus tridentatus and Carcinoscorpius rotundicauda. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Saulter:2020:ANR

[1077] Andrew N. Saulter, Christopher Bunney, Robert R. King, and Jennifer Waters. An application of NEMOVAR for regional wave model data assimilation. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Lampe:2021:STV

[1078] Vanessa Lampe, Eva-Maria Nöthig, and Markus Schartau. Spatiotemporal variations in community size structure of Arctic protist plankton in the Fram Strait. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Sullender:2021:SMM

[1079] Benjamin K. Sullender, Kelly Kapsar, Aaron Poe, and Martin Robards. Spatial management measures alter vessel behavior in the Aleutian Archipelago. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Romano:2020:WSS

[1080] Chiara Romano, Amandine Nunes-Jorge, Nadine Le Bris, Greg W. Rouse, Daniel Martin, and Christian Borowski. Wooden stepping stones:

Diversity and biogeography of deep-sea wood boring Xylophagaidae (Mollusca: Bivalvia) in the North-East Atlantic Ocean, with the description of a new genus. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See corrigendum [1356].

Bibi:2020:DAP

[1081] Riaz Bibi, Hee Yoon Kang, Dongyoung Kim, Jaebin Jang, Goutam Kumar Kundu, Young Kyun Kim, and Chang-Keun Kang. Dominance of autochthonous phytoplankton-derived particulate organic matter in a low-turbidity temperate estuarine embayment, Gwangyang Bay, Korea. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Zeller:2020:CAO

[1082] Mary A. Zeller, Bryce R. Van Dam, Christian Lopes, and John S. Kominoski. Carbonate-associated organic matter is a detectable dissolved organic matter source in a subtropical seagrass meadow. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Ferra:2020:UAA

[1083] Carmen Ferrà, Anna Nora Tassetti, Enrico Nicola Armelloni, Alessandro Galdelli, Giuseppe Scarcella, and Gianna Fabi. Using AIS to attempt a quantitative evaluation of unobserved trawling activity in the Mediterranean Sea. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Drown:2020:ISH

[1084] Melissa K. Drown, Amanda N. DeLiberto, Douglas L. Crawford, and Marjorie F. Oleksiak. An innovative setup for high-throughput respirometry of small aquatic animals. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Rainville:2020:ISA

[1085] Luc Rainville, Jeremy Wilkinson, Mary Ellen J. Durley, Scott Harper, Julia DiLeo, Martin J. Doble, Andrew Fleming, David Forcucci, Hans Graber, John T. Hargrove, John Haverlack, Nick Hughes, Brett Hembrough, Martin O. Jeffries, Craig M. Lee, Brendon Mendenhall, David McCormmick, Sofia Montalvo, Adam Stenseth, Geoffrey B. Shilling, Harper L. Simmons, James E. Toomey, and John Woods. Improving situational awareness in the Arctic Ocean. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Kurihara:2020:PAA

[1086] Haruko Kurihara, Yuri Suhara, Izumi Mimura, and Yimnang Golbuu. Potential acclimatization and adaptive responses of adult and transgeneration coral larvae from a naturally acidified habitat. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Giampaoletti:2020:SDB

[1087] Jacopo Giampaoletti, Frine Cardone, Giuseppe Corriero, Maria F. Gravina, and Luisa Nicoletti. Sharing and distinction in biodiversity and ecological role of bryozoans in Mediterranean mesophotic bioconstructions. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Li:2020:DSM

[1088] Jiangtao Li, Jingyu Yang, Mingxue Sun, Lei Su, Hu Wang, Jianqi Gao, and Shijie Bai. Distribution and succession of microbial communities along the dispersal pathway of hydrothermal plumes on the Southwest Indian Ridge. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Saulia:2020:SSD

[1089] Emmrick Saulia, Mar Benavides, Britt Henke, Kendra Turk-Kubo, Haley Cooperguard, Olivier Grosso, Anne Desnues, Martine Rodier, Cécile Dupouy, Lasse Riemann, and Sophie Bonnet. Seasonal shifts in diazotrophs players: Patterns observed over a two-year time series in the New Caledonian Lagoon (western tropical South Pacific Ocean). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Marchesi:2020:VMP

[1090] Maria C. Marchesi, Matías S. Mora, Silvana L. Dans, Mariano A. Coscarella, and Rolando González-José. Vertebral morphology in partially sympatric dolphins: a 3D approach. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Sweet:2020:RFA

[1091] William V. Sweet, Ayesha S. Genz, Jayantha Obeysekera, and John J. Marra. A regional frequency analysis of tide gauges to assess Pacific coast flood risk. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Passerotti:2020:ILH

[1092] Michelle S. Passerotti, Allen H. Andrews, and Lisa J. Natanson. Inferring life history characteristics of the oceanic whitetip shark Carcharhinus longimanus from vertebral bomb radiocarbon. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Becker:2020:GSR

[1093] Susan Becker, Michio Aoyama, E. Malcolm S. Woodward, Karel Bakker, Stephen Coverly, Claire Mahaffey, and Toste Tanhua. GO-SHIP repeat hydrography nutrient manual: The precise and accurate determination of dissolved inorganic nutrients in seawater, using continuous flow analysis methods. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Dang:2020:SUP

[1094] Viet D. H. Dang, Pak-Yin Cheung, Chia-Ling Fong, Aziz J. Mulla, Jia-Ho Shiu, Che-Hung Lin, and Yoko Nozawa. Sea urchins play an increasingly important role for coral resilience across reefs in Taiwan. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Angiolillo:2020:IML

[1095] Michela Angiolillo and Tomaso Fortibuoni. Impacts of marine litter on Mediterranean reef systems: From shallow to deep waters. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Gorman:2020:AHC

[1096] Richard Gorman. Atlantic horseshoe crabs and endotoxin testing: Perspectives on alternatives, sustainable methods, and the 3Rs (replacement, reduction, and refinement). Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Garcia-Corral:2021:PCM

[1097] Lara S. Garcia-Corral, Carlos M. Duarte, and Susana Agusti. Plankton community metabolism in Western Australia: Estuarine, coastal and oceanic surface waters. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Weber:2020:ERM

[1098] Laura Weber, Maickel Armenteros, Melissa Kido Soule, Krista Longnecker, Elizabeth B. Kujawinski, and Amy Apprill. Extracellular reef metabolites across the protected Jardines de la Reina, Cuba Reef System. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Holt:2021:EVD

[1099] Marla M. Holt, Jennifer B. Tennessen, Eric J. Ward, M. Bradley Hanson, Candice K. Emmons, Deborah A. Giles, and Jeffrey T. Hogan. Effects of vessel distance and sex on the behavior of endangered killer whales. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Gan:2020:ITR

[1100] Zhibin Gan, Peng Xu, Xinzheng Li, and Chunsheng Wang. Integrative taxonomy reveals two new species of stalked barnacle (Cirripedia, Thoracica) from seamounts of the Western Pacific with a review of barnacles distributed in seamounts worldwide. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Lin:2020:FHB

[1101] Chien-Hsiang Lin, Jing-Siang Lin, Kuo-Shu Chen, Meng-Hsien Chen, Chiee-Young Chen, and Chih-Wei Chang. Feeding habits of bigeye tuna (Thunnus obesus) in the Western Indian Ocean reveal a size-related shift in its fine-scale piscivorous diet. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Bargione:2020:AGS

[1102] Giada Bargione, Claudio Vasapollo, Fortunata Donato, Massimo Virgili, Andrea Petetta, and Alessandro Lucchetti. Age and growth of striped venus clam Chamelea gallina (Linnaeus, 1758) in the Mid-Western Adriatic Sea: a comparison of three laboratory techniques. Frontiers in Marine Science, 7, September 2020. ISSN 2296-7745.

Langlet:2020:FER

[1103] Dewi Langlet, Vincent M. P. Bouchet, Riccardo Riso, Yohei Matsui, Hisami Suga, Yoshihiro Fujiwara, and Hidetaka Nomaki. Foraminiferal ecology and role in nitrogen benthic cycle in the hypoxic Southeastern Bering Sea. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Barbour:2020:EBF

[1104] Nicole Barbour, George L. Shillinger, Aimee L. Hoover, Sean A. Williamson, Victoria J. Coles, Dong Liang, William F. Fagan, and Helen Bailey. Environmental and biological factors influencing dispersal of neonate leatherback turtles (Dermochelys coriacea) from an endangered Costa Rican nesting population. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Martinez-Vicente:2020:SSA

[1105] Victor Martinez-Vicente, Andrey Kurekin, Carolina Sá, Vanda Brotas, Ana Amorim, Vera Veloso, Junfang Lin, and Peter I. Miller. Sensitivity of a satellite algorithm for harmful algal bloom discrimination to the use of laboratory bio-optical data for training. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Favoretto:2020:BIP

[1106] Fabio Favoretto, Ismael Mascareñas-Osorio, Lorena León-Deniz, Carlos González-Salas, Horacio Pérez-España, Mariana Rivera-Higueras, Miguel-Ángel Ruiz-Zárate, Alejandro Vega-Zepeda, Harold Villegas-Hernández, and Octavio Aburto-Oropeza. Being isolated and protected is better than just being isolated: a case study from the Alacranes Reef, Mexico. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Pantazi:2020:TAK

[1107] Vyronia Pantazi, Alessandro Mannini, Paraskevas Vasilakopoulos, Kostas Kapiris, Persefoni Megalofonou, and Stefanos Kalogirou. That’s all I know: Inferring the status of extremely data-limited stocks. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Kennedy:2021:HTS

[1108] Brian R. C. Kennedy, Lara Hakam, Jan Witting, Regen Milani, Sue Taei, Tim Smith, Erin Taylor, Tooreka Teemari, and Randi D. Rotjan. Historical trends of sperm whale (Physeter macrocephalus) distribution in the Phoenix Archipelago. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Giorgetti:2020:AVD

[1109] Alessandra Giorgetti, Marina Lipizer, Maria Eugenia Molina Jack, Neil Holdsworth, Hans Mose Jensen, Luminita Buga, George Sarbu, Athanasia Iona, Julie Gatti, Martin Larsen, Lotta Fyrberg, Ann Kristin Østrem, and Reiner Schlitzer. Aggregated and validated datasets for the European Seas: The contribution of EMODnet chemistry. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Boo:2020:CMG

[1110] Ga Hun Boo, Mayalen Zubia, Jeffery R. Hughey, Alison R. Sherwood, Mutue T. Fujii, Sung Min Boo, and Kathy Ann Miller. Complete mitochondrial genomes reveal population-level patterns in the widespread red alga Gelidiella fanii (Gelidiales, Rhodophyta). Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Kvale:2021:CPD

[1111] Karin Kvale, Wolfgang Koeve, and Nadine Mengis. Calcifying phytoplankton demonstrate an enhanced role in greenhouse atmospheric CO_{2} regulation. Frontiers in Marine Science, 7, January 2021. ISSN 2296-

Hauser-Davis:2020:SCL

[1112] Rachel Ann Hauser-Davis, Lívia Figueiredo, Leila Lemos, Jailson Fulgêncio de Moura, Rafael C. C. Rocha, Tatiana Saint’Pierre, Roberta L. Ziolli, and Salvatore Siciliano. Subcellular cadmium, lead and mercury compartmentalization in Guiana dolphins (Sotalia guianensis) from Southeastern Brazil. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Munoz-Colmenero:2021:EDE

[1113] Marta Muñoz-Colmenero, Ana Sánchez, Begoña Correa, Francisco G. Figueiras, Jose L. Garrido, and Carmen G. Sotelo. Evaluation of DNA extraction methods and bioinformatic pipelines for marine nano- and pico-eukaryotic plankton analysis. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Papale:2020:SRV

[1114] Elena Papale, Carmelo Fanizza, Giuseppa Buscaino, Maria Ceraulo, Giulia Cipriano, Roberto Crugliano, Rosario Grammauta, Martina Gregorietti, Vito Renò, Pasquale Ricci, Francesca C. Santacesaria, Rosalia Maglietta, and Roberto Carlucci. The social role of vocal complexity in striped dolphins. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Dai:2020:APC

[1115] Jihong Dai, Weihao Ou, Guijuan Yu, Qinghui Ai, Wenbing Zhang, Kangsen Mai, and Yanjiao Zhang. The antimicrobial peptide cecropin AD supplement alleviated soybean meal-induced intestinal inflammation, barrier damage, and microbial dysbiosis in juvenile turbot, Scophthalmus maximus. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Kang:2020:ADK

[1116] Myounghee Kang, Rina Fajaryanti, Wuju Son, Jeong-Hoon Kim, and Hyoung Sul La. Acoustic detection of krill scattering layer in the Terra Nova Bay Polynya, Antarctica. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Klugel:2020:HAC

[1117] Andreas Klügel, Heinrich Villinger, Miriam Römer, Norbert Kaul, Sebastian Krastel, Kai-Frederik Lenz, and Paul Wintersteller. Hydrothermal activity at a Cretaceous seamount, Canary Archipelago, caused by rejuvenated volcanism. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Falsone:2021:HFC

[1118] Fabio Falsone, Danilo Scannella, Michele Luca Geraci, Vita Gancitano, Sergio Vitale, and Fabio Fiorentino. How fishery collapses: The case of Lepidopus caudatus (Pisces: Trichiuridae) in the Strait of Sicily (Central Mediterranean). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Lin:2020:IHS

[1119] Peng Lin, Chen Xu, Luni Sun, Wei Xing, and Peter H. Santschi. Incorporation of hydroxamate siderophore and associated Fe into marine particles in natural seawater. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Howell:2020:BIG

[1120] Kerry L. Howell, Ana Hilário, A. Louise Allcock, David M. Bailey, Maria Baker, Malcolm R. Clark, Ana Colaço, Jon Copley, Erik E. Cordes, Roberto Danovaro, Awantha Dissanayake, Elva Escobar, Patricia Esquete, Austin J. Gallagher, Andrew R. Gates, Sylvie M. Gaudron, Christopher R. German, Kristina M. Gjerde, Nicholas D. Higgs, Nadine Le Bris, Lisa A. Levin, Elisabetta Manea, Craig McClain, Lenaick Menot, Nelia C. Mestre, Anna Metaxas, Rosanna J. Milligan, Agnes W. N. Muthumbi, Bhavani E. Narayanaswamy, Sofia P. Ramalho, Eva Ramirez-Llodra, Laura M. Robson, Alex D. Rogers, Javier Sellanes, Julia D. Sigwart, Kerry Sink, Paul V. R. Snelgrove, Paris V. Stefanoudis, Paulo Y. Sumida, Michelle L. Taylor, Andrew R. Thurber, Rui P. Vieira, Hiromi K. Watanabe, Lucy C. Woodall, and Joana R. Xavier. A blueprint for an inclusive, global deep-sea ocean decade field program. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Wang:2020:NSI

[1121] Yuqi Wang, Fuyou Xu, and Zhanbiao Zhang. Numerical simulation of inline forces on a bottom-mounted circular cylinder under the action of a specific freak wave. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Mannini:2020:ERE

[1122] Alessandro Mannini, Cecilia Pinto, Christoph Konrad, Paraskevas Vasilakopoulos, and Henning Winker. “The elephant in the room”: Exploring natural mortality uncertainty in statistical catch at age models. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Aniceto:2020:AMD

[1123] Ana Sofia Aniceto, Geir Pedersen, Raul Primicerio, Martin Biuw, Ulf Lindstrøm, and Lionel Camus. Arctic marine data collection using oceanic gliders: Providing ecological context to cetacean vocalizations. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Yi:2020:DMC

[1124] Jun Yi, Linus Shing Him Lo, and Jinping Cheng. Dynamics of microbial community structure and ecological functions in estuarine intertidal sediments. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Glaus:2020:III

[1125] Kerstin B. J. Glaus, Sharon A. Appleyard, Brian Stockwell, Juerg M. Brunnschweiler, Mahmood Shivji, Eric Clua, Amandine D. Marie, and Ciro Rico. Insights into insular isolation of the bull shark, Carcharhinus leucas (Müller and Henle, 1839), in Fijian waters. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Kindong:2020:AGS

[1126] Richard Kindong, Haozhan Wang, Feng Wu, Xiaojie Dai, and Siquan Tian. Age, growth, and sexual maturity of the crocodile shark, Pseudocarcharias kamoharai, from the Eastern Atlantic Ocean. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Carbonara:2020:GAV

[1127] Pierluigi Carbonara, Andrea Bellodi, Michele Palmisano, Antonello Mulas, Cristina Porcu, Walter Zupa, Marilena Donnaloia, Roberto Carlucci, Letizia Sion, and Maria Cristina Follesa. Growth and age validation of the thornback ray (Raja clavata Linnaeus, 1758) in the South Adriatic Sea (Central Mediterranean). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Schultzhaus:2020:DSC

[1128] Janna N. Schultzhaus, Chenyue Wang, Shrey Patel, Madeline Smerchansky, Daniel Phillips, Chris R. Taitt, Dagmar H. Leary, Judson Hervey, Gary H. Dickinson, Christopher R. So, Jenifer M. Scancella, Kathryn J. Wahl, and Christopher M. Spillmann. Distribution of select cement proteins in the acorn barnacle Amphibalanus amphitrite. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Schneider:2020:ETS

[1129] Lisa K. Schneider, Kevin J. Flynn, Peter M. J. Herman, Tineke A. Troost, and Willem Stolte. Exploring the trophic spectrum: Placing mixoplankton into marine protist communities of the Southern North Sea. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Helble:2020:FWS

[1130] Tyler A. Helble, Regina A. Guazzo, Gabriela C. Alongi, Cameron R. Martin, Stephen W. Martin, and E. Elizabeth Henderson. Fin whale song patterns shift over time in the Central North Pacific. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Pati:2020:EPP

[1131] Siddhartha Pati, Salwa Shahimi, Hisham Atan Edinur, Bryan Raveen Nelson, Diptikanta Acharya, and Bisnu Prasad Dash. Extraction of people’s perception toward horseshoe crab existence in Northeast Coast of India. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Levy:2020:ISF

[1132] Tom Levy, Sherry L. Tamone, Rivka Manor, Eliahu D. Aflalo, Menachem Y. Sklarz, Vered Chalifa-Caspi, and Amir Sagi. The IAG-switch and further transcriptomic insights into sexual differentiation of a protandric shrimp. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Schumacher:2020:BSP

[1133] Johanna Schumacher, Lisa Bergqvist, Floris M. van Beest, Jacob Carstensen, Bo Gustafsson, Berit Hasler, Vivi Fleming, Henrik Nygård, Kristîne Pakalniete, Alexander Sokolov, Marianne Zandersen, and Gerald Schernewski. Bridging the science-policy gap — toward better integration of decision support tools in coastal and marine policy implementation. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Zhang:2020:SVP

[1134] Zhaoru Zhang, Meng Zhou, Yisen Zhong, Guosen Zhang, Shan Jiang, Yonghui Gao, Ruifeng Zhang, and Walker O. Smith. Spatial variations of phytoplankton biomass controlled by river plume dynamics over the Lower Changjiang Estuary and adjacent shelf based on high-resolution observations. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Tourneroche:2020:BFM

[1135] Anne Tourneroche, Raphaël Lami, Gaëtan Burgaud, Isabelle Domart-Coulon, Wei Li, Claire Gachon, Marc Gèze, Dominique Boeuf, and Soizic Prado. The bacterial and fungal microbiota of Saccharina latissima (Laminariales, Phaeophyceae). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Xiao:2020:PGS

[1136] Yao Xiao, Ting Xu, Jin Sun, Yan Wang, Wai Chuen Wong, Yick Hang Kwan, Chong Chen, Jian-Wen Qiu, and Pei-Yuan Qian. Population genetic structure and gene expression plasticity of the deep-sea vent and seep squat lobster Shinkaia crosnieri. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Bucci:2020:IVT

[1137] Andre F. Bucci, Andrew C. Thomas, and Ivona Cetinić. Interannual variability in the thermal habitat of Alexandrium catenella in the Bay of Fundy and the implications of climate change. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Stefanni:2021:MSL

[1138] Sergio Stefanni, Diana Catarino, Pedro A. Ribeiro, Mafalda Freitas, Gui M. Menezes, Francis Neat, and David Stanković. Molecular systematics of the long-snouted deep water dogfish (Centrophoridae, Deania) with implications for identification, taxonomy, and conservation. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Mao:2020:API

[1139] Chongyuan Mao, Robert R. King, Rebecca Reid, Matthew J. Martin, and Simon A. Good. Assessing the potential impact of changes to the Argo and moored buoy arrays in an operational ocean analysis system. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Powers:2020:OPP

[1140] Leanne C. Powers, Rossana Del Vecchio, Neil V. Blough, Natasha Mc-Donald, Philippe Schmitt-Kopplin, and Michael Gonsior. Optical properties and photochemical transformation of the dissolved organic matter released by sargassum. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Li:2020:EWD

[1141] Ming Li, Renjian Li, Wei-Jun Cai, Jeremy M. Testa, and Chunqi Shen. Effects of wind-driven lateral upwelling on estuarine carbonate chemistry. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Baechler:2020:MER

[1142] Britta R. Baechler, Elise F. Granek, Scott J. Mazzone, Max Nielsen-Pincus, and Susanne M. Brander. Microplastic exposure by razor clam recreational harvester-consumers along a sparsely populated coastline. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Petrik:2020:LPF

[1143] Colleen M. Petrik, Charles A. Stock, Ken H. Andersen, P. Daniël van Denderen, and James R. Watson. Large pelagic fish are most sensitive to climate change despite pelagification of ocean food webs. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Tomaselli:2021:DMT

[1144] Pietro D. Tomaselli, Martin Dixen, Rodolfo Bolaños Sanchez, and Jacob Tornfeldt Sørensen. A decision-making tool for planning o&M activities of offshore wind farms using simulated actual decision drivers. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Luque:2020:PSS

[1145] Patricia Lastra Luque, Saburo Sakai, Hilario Murua, and Haritz Arrizabalaga. Protocol for sampling sequential fin spine growth intervals for isotope analysis in the Atlantic bluefin tuna. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Caballero:2020:IHR

[1146] Ainhoa Caballero, Sandrine Mulet, Nadia Ayoub, Ivan Manso-Narvarte, Xabier Davila, Christine Boone, Florence Toublanc, and Anna Rubio. Integration of HF radar observations for an enhanced coastal mean dynamic topography. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Lucey:2020:MSE

[1147] Noelle Lucey, Eileen Haskett, and Rachel Collin. Multi-stressor extremes found on a tropical coral reef impair performance. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Stimmelmayr:2020:GWE

[1148] Raphaela Stimmelmayr and Frances M. D. Gulland. Gray whale (Eschrichtius robustus) health and disease: Review and future directions. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Harianto:2021:IAW

[1149] Januar Harianto, Joshua Aldridge, Sergio A. Torres Gabarda, Richard J. Grainger, and Maria Byrne. Impacts of acclimation in warm-low pH conditions on the physiology of the sea urchin Heliocidaris erythrogramma and carryover effects for juvenile offspring. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Mackay:2020:IBI

[1150] Mary Mackay, Britta Denise Hardesty, and Chris Wilcox. The intersection between illegal fishing, crimes at sea, and social well-being. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Omori:2020:EAD

[1151] Yuko Omori, Akira Saeki, Shigeki Wada, Yuji Inagaki, and Takeo Hama. Experimental analysis of diurnal variations in humic-like fluorescent dissolved organic matter in surface seawater. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Li:2020:MFD

[1152] Ao Li, Chaogang Wang, Wei Wang, Ziyan Zhang, Mingkun Liu, Zhicai She, Zhen Jia, Guofan Zhang, and Li Li. Molecular and fitness data reveal local adaptation of southern and northern estuarine oysters (Crassostrea ariakensis). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Syrjala:2020:UAE

[1153] Joonas Syrjälä, Risto Kalliola, and Jukka Pajala. Underwater acoustic environment of coastal sea with heavy shipping traffic: NE Baltic Sea during wintertime. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Srinivas:2020:DFP

[1154] Tatiparthi Srinivas, Soniya Sukumaran, S. Neetu, and K. Ramesh Babu. Diversity and functional patterns of benthic amphipods in the coralline intertidal zones of a marine national park, India. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Gugliandolo:2020:PMT

[1155] Enrico Gugliandolo, Patrizia Licata, Rosalia Crupi, Ambrogina Albergamo, Amel Jebara, Vincenzo Lo Turco, Angela Giorgia Potortì, Hedi Ben Mansour, Salvatore Cuzzocrea, and Giuseppa Di Bella. Plasticizers as microplastics tracers in Tunisian marine environment. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Patanasatienkul:2020:CET

[1156] Thitiwan Patanasatienkul, Marina K. V. C. Delphino, and Krishna K. Thakur. Comparing the effectiveness of traditional and alternative baits in Prince Edward Island, Canada lobster fishery. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Li:2020:SSV

[1157] Haibo Li, Xue Chen, Michel Denis, Yuan Zhao, Lingfeng Huang, Zengjie Jiang, Wuchang Zhang, and Tian Xiao. Seasonal and spatial variation of pelagic microbial food web structure in a semi-enclosed temperate bay. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Guerra-Castro:2020:SBM

[1158] Edlin Guerra-Castro, Gema Hidalgo, Raúl E. Castillo-Cupul, María Muciño-Reyes, Elsa Noreña-Barroso, Jaime Quiroz-Deaquino, Maite Mascaro, and Nuno Simoes. Sandy beach macrofauna of Yucatán State (Mexico) and oil industry development in the Gulf of Mexico: First approach for detecting environmental impacts. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Martins:2020:MRM

[1159] José Carlos Martins, Dany Domínguez-Pérez, Catarina Azevedo, Ana Catarina Braga, Pedro Reis Costa, Hugo Osório, Vitor Vasconcelos, and Alexandre Campos. Molecular responses of mussel Mytilus galloprovincialis associated to accumulation and depuration of marine biotoxins okadaic acid and dinophysistoxin-1 revealed by shotgun proteomics. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Fairchild:2021:HRC

[1160] William Fairchild and Burke Hales. High-resolution carbonate system dynamics of Netarts Bay, OR from 2014 to 2019. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Brauko:2020:MHS

[1161] Kalina M. Brauko, Alex Cabral, Natasha V. Costa, Juliana Hayden, Carlos E. P. Dias, Edilene S. Leite, Renan D. Westphal, Carolina M. Mueller, Jason M. Hall-Spencer, Regina R. Rodrigues, Leonardo R. Rörig, Paulo R. Pagliosa, Alessandra L. Fonseca, Orestes E. Alarcon, and Paulo A. Horta. Marine heatwaves, sewage and eutrophication combine to trigger deoxygenation and biodiversity loss: a SW Atlantic case study. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Stephens:2020:OMC

[1162] Brandon M. Stephens, Keri Opalk, Daniel Petras, Shuting Liu, Jacqueline Comstock, Lihini I. Aluwihare, Dennis A. Hansell, and Craig A. Carlson. Organic matter composition at Ocean Station Papa affects its bioavailability, bacterioplankton growth efficiency and the responding taxa. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Hou:2020:MMI

[1163] Gang Hou, Jinrun Wang, Zuozhi Chen, Jinlong Zhou, Wangsu Huang, and Hui Zhang. Molecular and morphological identification and seasonal distribution of eggs of four Decapterus fish species in the Northern South China Sea: a key to conservation of spawning ground. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Kerpen:2020:WID

[1164] Nils B. Kerpen, Torsten Schlurmann, Alexander Schendel, Jannek Gundlach, Daniel Marquard, and Markus Hüpgen. Wave-induced distribution of microplastic in the surf zone. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Reid:2021:TCT

[1165] Ross M. Reid, Andrea L. Reid, David A. Lovejoy, and Peggy R. Biga. Teneurin C-terminal associated peptide (TCAP)-3 increases metabolic activity in zebrafish. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Fogarty:2020:AHA

[1166] Hannah E. Fogarty, Christopher Cvitanovic, Alistair J. Hobday, and Gretta T. Pecl. An assessment of how Australian fisheries management plans account for climate change impacts. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745. See erratum [2739].

Dobbelaere:2020:CEH

[1167] Thomas Dobbelaere, Erinn M. Muller, Lewis J. Gramer, Daniel M. Holstein, and Emmanuel Hanert. Coupled epidemio-hydrodynamic modeling to understand the spread of a deadly coral disease in Florida. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Li:2020:TRD

[1168] Huiru Li, Lydia Scheschonk, Sandra Heinrich, Klaus Valentin, Lars Harms, Gernot Glöckner, Erwan Corre, and Kai Bischof. Transcriptomic responses to darkness and the survival strategy of the kelp Saccharina latissima in the early polar night. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Kindong:2020:SSA

[1169] Richard Kindong, Chunxia Gao, Njomoue Achille Pandong, Qiuyun Ma, Siquan Tian, Feng Wu, and Ousmane Sarr. Stock status assessments of five small pelagic species in the Atlantic and Pacific Oceans using the length-based Bayesian estimation (LBB) method. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Shettigar:2020:ESS

[1170] Nithin Achutha Shettigar, Biswa Bhattacharya, Lörinc Mészáros, Anna Spinosa, and Ghada El Serafy. 3D ensemble simulation of seawater temperature — an application for aquaculture operations. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Otero:2020:PDE

[1171] Jaime Otero, Xosé Antón Álvarez-Salgado, and Antonio Bode. Phytoplankton diversity effect on ecosystem functioning in a coastal upwelling system. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Stroeve:2021:MSM

[1172] Julienne Stroeve, Martin Vancoppenolle, Gaelle Veyssiere, Marion Lebrun, Giulia Castellani, Marcel Babin, Michael Karcher, Jack Landy, Glen E. Liston, and Jeremy Wilkinson. A multi-sensor and modeling approach for mapping light under sea ice during the ice-growth season. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Wilson:2021:LTS

[1173] Megan N. Wilson, Adrian E. Laufer, Erin M. Howard, and Jennifer A. T. K. Wong-Ala. Lessons from the trenches: Students’ perspectives of their own marine transdisciplinary education. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Themudo:2020:LGE

[1174] Gonçalo Espregueira Themudo, Luís Q. Alves, André M. Machado, Mónica Lopes-Marques, Rute R. da Fonseca, Miguel Fonseca, Raquel Ruivo, and L. Filipe C. Castro. Losing genes: The evolutionary remodeling of cetacea skin. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Mariz:2020:UME

[1175] João Pedro V. Mariz and Mariana F. Nery. Unraveling the molecular evolution of blood coagulation genes in fishes and cetaceans. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Boute:2021:EEP

[1176] Pim G. Boute, Maarten Soetaert, José A. Reid Navarro, and Martin J. Lankheet. Effects of electrical pulse stimulation on behaviour and survival of marine benthic invertebrates. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Scrosati:2020:LBD

[1177] Ricardo A. Scrosati, Heather L. MacDonald, César A. Córdova, and Graciela N. Casas. Length and biomass data for Atlantic and Pacific seaweeds from both hemispheres. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Rossbach:2021:DAT

[1178] Susann Rossbach, Andrea Anton, and Carlos M. Duarte. Drivers of the abundance of Tridacna spp. giant clams in the Red Sea. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Liu:2021:DIS

[1179] Yifan Liu, Qinghua Liu, Shihong Xu, Yanfeng Wang, Chengcheng Feng, Chunyan Zhao, Zongcheng Song, and Jun Li. A deep insight of spermatogenesis and hormone levels of aqua-cultured turbot (Scophthalmus maximus). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Teff-Seker:2020:DAD

[1180] Yael Teff-Seker, Peter C. Mackelworth, Tomás Vega Fernández, John McManus, Jungho Nam, Arthur O. Tuda, and Drasko Holcer. Do Alternative Dispute Resolution (ADR) and Track Two processes support transboundary marine conservation? Lessons from six case studies of maritime disputes. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Fruet:2021:DVS

[1181] Pedro F. Fruet, Luciana M. Möller, and Eduardo R. Secchi. Dynamics and viability of a small, estuarine-resident population of Lahille’s bottlenose dolphins from Southern Brazil. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Levine:2020:ODC

[1182] Robert M. Levine, Kristen E. Fogaren, Johna E. Rudzin, Christopher J. Russoniello, Dax C. Soule, and Justine M. Whitaker. Open data, collaborative working platforms, and interdisciplinary collaboration: Building an early career scientist community of practice to leverage Ocean Observatories Initiative data to address critical questions in marine science. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Fontela:2020:NAW

[1183] Marcos Fontela, Fiz F. Pérez, Herlé Mercier, and Pascale Lherminier. North Atlantic western boundary currents are intense dissolved organic carbon streams. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Henderson:2020:DSC

[1184] Mark J. Henderson, David D. Huff, and Mary M. Yoklavich. Deep-sea coral and sponge taxa increase demersal fish diversity and the probability of fish presence. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Barrios-Garrido:2021:GDS

[1185] Hector Barrios-Garrido, Kareen De Turris-Morales, and Ninive Edilia Espinoza-Rodriguez. Guiana dolphin (Sotalia guianensis) in the Maracaibo Lake System, Venezuela: Conservation, threats, and population overview. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Bousquet:2020:STO

[1186] Olivier Bousquet, Mayeul Dalleau, Marion Bocquet, Philippe Gaspar, Soline Bielli, Stéphane Ciccione, Elisabeth Remy, and Arthur Vidard. Sea turtles for ocean research and monitoring: Overview and initial results of the STORM project in the Southwest Indian Ocean. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Li:2020:SAC

[1187] Zhibing Li, Xiaohua Wang, Jianyu Hu, Fernando Pinheiro Andutta, and Zhiqiang Liu. A study on an anticyclonic–cyclonic eddy pair off Fraser Island, Australia. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Kersting:2020:RDR

[1188] Diego K. Kersting, Maite Vázquez-Luis, Baptiste Mourre, Fatima Z. Belkhamssa, Elvira Álvarez, Tatjana Bakran-Petricioli, Carmen Barberá, Agustín Barrajón, Emilio Cortés, Salud Deudero, José R. García-March, Salvatore Giacobbe, Francisca Giménez-Casalduero, Luis González, Santiago Jiménez-Gutiérrez, Silvija Kipson, Javier Llorente, Diego Moreno, Patricia Prado, Juan A. Pujol, Jordi Sánchez, Andrea Spinelli, José M. Valencia, Nardo Vicente, and Iris E. Hendriks. Recruitment disruption and the role of unaffected populations for potential recovery after the Pinna nobilis mass mortality event. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Caputi:2020:SFW

[1189] Simona Sporta Caputi, Giulio Careddu, Edoardo Calizza, Federico Fiorentino, Deborah Maccapan, Loreto Rossi, and Maria Letizia Costantini. Seasonal food web dynamics in the Antarctic benthos of Tethys Bay (Ross Sea): Implications for biodiversity persistence under different seasonal sea-ice coverage. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745. See corrigendum [2815].

Fragkopoulou:2021:BTT

[1190] Eliza Fragkopoulou, Ester A. Serrão, Paulo A. Horta, Gabrielle Koerich, and Jorge Assis. Bottom trawling threatens future climate refugia of rhodoliths globally. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Raoult:2020:PGR

[1191] Vincent Raoult, Clive N. Trueman, Kelsey M. Kingsbury, Bronwyn M. Gillanders, Matt K. Broadhurst, Jane E. Williamson, Ivan Nagelkerken, David J. Booth, Victor Peddemors, Lydie I. E. Couturier, and Troy F. Gaston. Predicting geographic ranges of marine animal populations using stable isotopes: a case study of great hammerhead sharks in Eastern Australia. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Guo:2020:IAV

[1192] Xianghui Guo, Xue Song, Ying Gao, Yaohua Luo, Yi Xu, Tao Huang, and Lifang Wang. Inter-annual variability of the carbonate system in the hypoxic Upper Pearl River Estuary in winter. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Murawski:2021:SDH

[1193] Steven A. Murawski, Joshua P. Kilborn, Adriana C. Bejarano, David Chagaris, David Donaldson, Frank J. Hernandez, Timothy C. MacDonald, Craig Newton, Ernst Peebles, and Kelly L. Robinson. A synthesis of Deepwater Horizon impacts on coastal and nearshore living marine resources. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Dalton:2020:MRL

[1194] Kathryn Dalton, Marlena Skrobe, Henry Bell, Benjamin Kantner, Dave Berndtson, Leopoldo C. Gerhardinger, and Patrick Christie. Marinerelated learning networks: Shifting the paradigm toward collaborative ocean governance. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Jesus:2021:TKA

[1195] Manuella Dultra Jesus, João Bráullio de Luna Sales, Rodrigo Silvestre Martins, Jonathan Stuart Ready, Tarcisio Augusto S. Costa, Jonathan D. Ablett, and Alexandre Schiavetti. Traditional knowledge aids description when resolving the taxonomic status of unsettled species using classical and molecular taxonomy: The case of the shallow-water octopus Callistoctopus furvus (Gould, 1852) from the Western Atlantic Ocean. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Koutsouveli:2020:RBG

[1196] Vasiliki Koutsouveli, Paco Cárdenas, Maria Conejero, Hans Tore Rapp, and Ana Riesgo. Reproductive biology of Geodia species (Porifera, Tetractinellida) from Boreo-Arctic North-Atlantic deep-sea sponge grounds. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Vollbrecht:2021:LTP

[1197] Carleigh Vollbrecht, Paula Moehlenkamp, Jamison M. Gove, Anna B. Neuheimer, and Margaret A. McManus. Long-term presence of the island mass effect at Rangiroa Atoll, French Polynesia. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Drennan:2021:AFP

[1198] Regan Drennan, Thomas G. Dahlgren, Katrin Linse, and Adrian G. Glover. Annelid fauna of the Prince Gustav Channel, a previously icecovered seaway on the Northeastern Antarctic Peninsula. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Wang:2020:DLR

[1199] Wenqiang Wang, Yujue Xu, Shuyan Chi, Peng Yang, Kangsen Mai, and Fei Song. Dietary lysine regulates body growth performance via the nutrient-sensing signaling pathways in largemouth bass (Micropterus salmoides). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

LaCasella:2021:MDP

[1200] Erin L. LaCasella, Michael P. Jensen, Christine A. Madden Hof, Ian P. Bell, Amy Frey, and Peter H. Dutton. Mitochondrial DNA profiling to combat the illegal trade in tortoiseshell products. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Tedetti:2020:IFD

[1201] Marc Tedetti, Lionel Bigot, Jean Turquet, Catherine Guigue, Nicolas Ferretto, Madeleine Goutx, and Pascale Cuet. Influence of freshwater discharges on biogeochemistry and benthic communities of a coral reef ecosystem (La Réunion Island, Indian Ocean). Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Hendrickson:2021:RSS

[1202] Brianna N. Hendrickson, Sarah D. Brooks, Daniel C. O. Thornton, Richard H. Moore, Ewan Crosbie, Luke D. Ziemba, Craig A. Carlson, Nicholas Baetge, Jessica A. Mirrielees, and Alyssa N. Alsante. Role of sea surface microlayer properties in cloud formation. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

deKluijver:2021:IMC

[1203] Anna de Kluijver, Martijn C. Bart, Dick van Oevelen, Jasper M. de Goeij, Sally P. Leys, Sandra R. Maier, Manuel Maldonado, Karline Soetaert, Sander Verbiest, and Jack J. Middelburg. An integrative model of carbon and nitrogen metabolism in a common deep-sea sponge (Geodia barretti). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Freire-Penaherrera:2020:SBA

[1204] Andrea Martha Freire-Peñaherrera, Ana Tirapé, Jerry Landívar-Zambrano, and Juan Manuel Cevallos-Cevallos. A strain of Bacillus amyloliquefaciens can prevent Vibrio vulnificus colonization in Crassostrea gigas oysters. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Zhang:2020:DTC

[1205] Qian Zhang, Min Su, Peng Yao, Yongping Chen, Marcel J. F. Stive, and Zheng B. Wang. Dynamics of a tidal current system in a marginal sea: a case study of the Yellow Sea, China. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Tervo:2021:HSH

[1206] Outi M. Tervo, Susanne Ditlevsen, Manh C. Ngô, Nynne H. Nielsen, Susanna B. Blackwell, Terrie M. Williams, and Mads Peter Heide-Jørgensen. Hunting by the stroke: How foraging drives diving behavior and locomotion of East-Greenland narwhals (Monodon monoceros). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Yan:2021:GHG

[1207] Weijie Yan, Yingjie Qiao, Jiangbo Qu, Xiumei Liu, Quanqi Zhang, and Xubo Wang. The hsp40 gene family in Japanese flounder: Identification, phylogenetic relationships, molecular evolution analysis, and expression patterns. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Curnick:2020:BES

[1208] David J. Curnick, Samantha Andrzejaczek, David M. P. Jacoby, Daniel M. Coffey, Aaron B. Carlisle, Taylor K. Chapple, Francesco Ferretti, Robert J. Schallert, Timothy White, Barbara A. Block, Heather J. Koldewey, and Ben Collen. Behavior and ecology of silky sharks around the Chagos Archipelago and evidence of Indian Ocean wide movement. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Bell:2021:PSD

[1209] Thomas G. Bell, Jack G. Porter, Wei-Lei Wang, Michael J. Lawler, Emmanuel Boss, Michael J. Behrenfeld, and Eric S. Saltzman. Predictability of seawater DMS during the North Atlantic aerosol and marine ecosystem study (NAAMES). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Wiese:2020:IMV

[1210] Anne Wiese, Joanna Staneva, Ha Thi Minh Ho-Hagemann, Sebastian Grayek, Wolfgang Koch, and Corinna Schrum. Internal model variability of ensemble simulations with a regional coupled wave-atmosphere model GCOAST. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Tepsich:2020:WWP

[1211] Paola Tepsich, Anna Borroni, Margherita Zorgno, Massimiliano Rosso, and Aurélie Moulins. Whale watching in the pelagos sanctuary: Status and quality assessment. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Watson:2021:AGS

[1212] David G. Watson, Patrick P. Pomeroy, and Malcolm W. Kennedy. Atlantic grey seal milk shows continuous changes in key metabolites and indicators of metabolic transition in pups from birth to weaning. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ducker:2020:HPO

[1213] James Ducker and Laura J. Falkenberg. How the Pacific oyster responds to ocean acidification: Development and application of a meta-analysis based adverse outcome pathway. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Smallegange:2020:UPL

[1214] Isabel M. Smallegange, Marta Flotats Avilés, and Kim Eustache. Unusually paced life history strategies of marine megafauna drive atypical sensitivities to environmental variability. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Choi:2021:EPA

[1215] Seong Seok Choi, Yong Bae Seo, Soo-Wan Nam, and Gun-Do Kim. Enhanced production of astaxanthin by co-culture of Paracoccus haeundaensis and lactic acid bacteria. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Meiling:2020:PRD

[1216] Sonora Meiling, Erinn M. Muller, Tyler B. Smith, and Marilyn E. Brandt. 3D photogrammetry reveals dynamics of Stony Coral Tissue Loss Disease (SCTLD) lesion progression across a thermal stress event. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Hanz:2021:SVN

[1217] Ulrike Hanz, Lindsay Beazley, Ellen Kenchington, Gerard Duineveld, Hans Tore Rapp, and Furu Mienis. Seasonal variability in near-bed environmental conditions in the Vazella pourtalesii glass sponge grounds of the Scotian Shelf. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See corrigendum [2475].

Villalobos:2020:OAO

[1218] Cristina Villalobos, Brooke A. Love, and M. Brady Olson. Ocean acidification and ocean warming effects on Pacific herring (Clupea pallasi) early life stages. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

McHuron:2020:PAB

[1219] Elizabeth A. McHuron, Katie Luxa, Noel A. Pelland, Kirstin Holsman, Rolf Ream, Tonya Zeppelin, and Jeremy T. Sterling. Practical application of a bioenergetic model to inform management of a declining fur seal population and their commercially important prey. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Medina-Calderon:2021:HSI

[1220] Jairo Humberto Medina-Calderón, José Ernesto Mancera-Pineda, Edward Castañeda-Moya, and Víctor H. Rivera-Monroy. Hydroperiod and salinity interactions control mangrove root dynamics in a karstic oceanic island in the Caribbean Sea (San Andres, Colombia). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Hoffmayer:2021:SOH

[1221] Eric R. Hoffmayer, Jennifer A. McKinney, James S. Franks, Jill M. Hendon, William B. Driggers, Brett J. Falterman, Benjamin Galuardi, and Michael E. Byrne. Seasonal occurrence, horizontal movements, and habitat use patterns of whale sharks (Rhincodon typus) in the Gulf of Mexico. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Penca:2020:MSC

[1222] Jerneja Penca. Mainstreaming sustainable consumption of seafood through enhanced mandatory food labeling. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Sisma-Ventura:2021:PEM

[1223] Guy Sisma-Ventura, Nurit Kress, Jacob Silverman, Yaron Gertner, Tal Ozer, Eli Biton, Ayah Lazar, Isaac Gertman, Eyal Rahav, and Barak Herut. Post-eastern Mediterranean transient oxygen decline in the deep waters of the southeast Mediterranean Sea supports weakening of ventilation rates. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Cesar-Ribeiro:2020:OEF

[1224] Caio Cesar-Ribeiro, Fernanda R. Piedras, Leticia C. da Cunha, Domênica T. de Lima, Luana Q. Pinho, and Gleyci A. O. Moser. Is oligotrophy an equalizing factor driving microplankton species functional diversity within agulhas rings? Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Dungan:2021:STE

[1225] Ashley M. Dungan, Madeleine J. H. van Oppen, and Linda L. Blackall. Short-term exposure to sterile seawater reduces bacterial community diversity in the sea anemone, Exaiptasia diaphana. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Hesp:2020:CGC

[1226] Kylie Hesp, John L. Flores Alvarez, Ana-Maria Alexandru, Jip van der Linden, Dirk E. Martens, Rene H. Wijffels, and Shirley A. Pomponi. CRISPR/Cas12a-mediated gene editing in Geodia barretti sponge cell culture. Frontiers in Marine Science, 7, December 2020. ISSN 2296-

Untiedt:2021:PRW

[1227] Candice Bobby Untiedt, Andrea M. Quattrini, Catherine S. McFadden, Phil A. Alderslade, Eric Pante, and Christopher P. Burridge. Phylogenetic relationships within Chrysogorgia (Alcyonacea: Octocorallia), a morphologically diverse genus of octocoral, revealed using a target enrichment approach. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Cadier:2020:ICW

[1228] Charles Cadier, Elisa Bayraktarov, Renee Piccolo, and Maria Fernanda Adame. Indicators of coastal wetlands restoration success: a systematic review. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Ulman:2020:CLF

[1229] Aylin Ulman, Mustafa Zengin, Nazli Demirel, and Daniel Pauly. Corrigendum: The lost fish of Turkey: a recent history of disappeared species and commercial fishery extinctions for the Turkish Marmara and Black Seas. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745. See [752].

Saldivar-Lucio:2021:FTC

[1230] Romeo Saldívar-Lucio, Armando Trasviña-Castro, Narriman Jiddawi, Ratana Chuenpagdee, Lars Lindström, Svein Jentoft, Julia Fraga, and Maricela de la Torre-Castro. Fine-tuning climate resilience in marine socio-ecological systems: The need for accurate space-time representativeness to identify relevant consequences and responses. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Villagra:2021:EEW

[1231] Damian Villagra, Ana García-Cegarra, Diego I. Gallardo, and Aldo S. Pacheco. Energetic effects of whale-watching boats on humpback whales on a breeding ground. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Fingas:2020:DAC

[1232] Merv F. Fingas, Kaan Yetilmezsoy, and Majid Bahramian. Development of an algorithm for chemically dispersed oil spills. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Agostini:2021:GME

[1233] Sylvain Agostini, Fanny Houlbrèque, Tom Biscéré, Ben P. Harvey, Joshua M. Heitzman, Risa Takimoto, Wataru Yamazaki, Marco Milazzo, and Riccardo Rodolfo-Metalpa. Greater mitochondrial energy production provides resistance to ocean acidification in “Winning” hermatypic corals. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Jiang:2021:NAW

[1234] Shan Jiang, Jie Jin, Shuo Jiang, Ying Wu, Jianing Wang, Ju Chen, Zhenqiu Zhang, Sumei Liu, Yan Chang, Lingyan Wang, and Jing Zhang. Nitrogen in atmospheric wet depositions over the East Indian Ocean and West Pacific Ocean: Spatial variability, source identification, and potential influences. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Amrein:2020:IWW

[1235] Arielle M. Amrein, Hector M. Guzman, Katie C. Surrey, Beth Polidoro, and Leah R. Gerber. Impacts of whale watching on the behavior of humpback whales (Megaptera novaeangliae) in the coast of Panama. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Giddens:2021:NGS

[1236] Jonatha Giddens, Alan Turchik, Whitney Goodell, Michelle Rodriguez, and Denley Delaney. The National Geographic Society deep-sea camera system: a low-cost remote video survey instrument to advance biodiversity observation in the deep ocean. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Greene:2020:CDT

[1237] Austin Greene, Megan J. Donahue, Jamie M. Caldwell, Scott F. Heron, Erick Geiger, and Laurie J. Raymundo. Coral disease time series highlight size-dependent risk and other drivers of white syndrome in a multi-species model. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Morato:2020:EAM

[1238] Telmo Morato, Pedro Afonso, Gui M. Menezes, Ricardo S. Santos, and Mónica A. Silva. Editorial: The Azores marine ecosystem: an open window into North Atlantic open ocean and deep-sea environments. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745.

Valdes-Castro:2021:ETP

[1239] Valentina Valdés-Castro and Camila Fernandez. Effect of three pesticides used in salmon farming on ammonium uptake in Central-Southern and Northern Patagonia, Chile. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Najdek:2020:EIC

[1240] Mirjana Najdek, Marino Korlević, Paolo Paliaga, Marsej Markovski, Ingrid Ivančić, Ljiljana Iveša, Igor Felja, and Gerhard J. Herndl. Effects of the invasion of Caulerpa cylindracea in a Cymodocea nodosa meadow in the Northern Adriatic Sea. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Szymkowiak:2020:ABM

[1241] Marysia Szymkowiak and Melissa Rhodes-Reese. Adaptive behaviors to marine ecosystem shifts: Examining fishermen’s strategies in response to abundant juvenile sablefish (Anoplopoma fimbria) in Alaska. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Song:2020:OPN

[1242] Wenya Song, Yu Qiao, Jingyi Dong, William A. Bourland, Tengteng Zhang, and Xiaotian Luo. Ontogeny and phylogeny of a new hypotrichous ciliate (Protista, Ciliophora), Metaurostylopsis alrasheidi n. sp., with establishment of a new genus Monourostylopsi n. gen. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Chen:2020:SCG

[1243] Xiao Chen, Chundi Wang, Bo Pan, Borong Lu, Chao Li, Zhuo Shen, Alan Warren, and Lifang Li. Single-cell genomic sequencing of three peritrichs (Protista, Ciliophora) reveals less biased stop codon usage and more prevalent programmed ribosomal frameshifting than in other ciliates. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Gulland:2020:PCH

[1244] Frances M. D. Gulland, Ailsa J. Hall, Gina M. Ylitalo, Kathleen M. Colegrove, Tenaya Norris, Pádraig J. Duignan, Barbie Halaska, Karina Acevedo Whitehouse, Linda J. Lowenstine, Alissa C. Deming, and Teresa K. Rowles. Persistent contaminants and herpesvirus OtHV1 are positively associated with cancer in wild California sea lions (Zalophus californianus). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Ison:2021:TPS

[1245] Sierra Ison, Theo Ison, Patricia Marti-Puig, Katherine Needham, Michael K. Tanner, and J. Murray Roberts. Tourist preferences for seamount conservation in the Galápagos Marine Reserve. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Handley:2021:MIB

[1246] Jonathan Handley, Marie-Morgane Rouyer, Elizabeth J. Pearmain, Victoria Warwick-Evans, Katharina Teschke, Jefferson T. Hinke, Heather Lynch, Louise Emmerson, Colin Southwell, Gary Griffith, César A. Cárdenas, Aldina M. A. Franco, Phil Trathan, and Maria P. Dias. Marine important bird and biodiversity areas for penguins in Antarctica, targets for conservation action. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Yang:2020:EOA

[1247] Sung-Yin Yang, Chih-Ying Lu, Sen-Lin Tang, Rocktim Ramen Das, Kazuhiko Sakai, Hideyuki Yamashiro, and Shan-Hua Yang. Effects of ocean acidification on coral endolithic bacterial communities in Isopora palifera and Porites lobata. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Isinibilir:2020:MCE

[1248] Melek Isinibilir, Leonid Svetlichny, Taras Mykitchak, Ezgi E. Türkeri, Kamil Mert Eryalçın, Onur Doğan, Gülşah Can, Esin Yüksel, and Ahmet E. Kideys. Microplastic consumption and its effect on respiration rate and motility of Calanus helgolandicus from the Marmara Sea. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Dias:2020:HCB

[1249] Vítor Dias, Frederico Oliveira, Joana Boavida, Ester A. Serrão, Jorge M. S. Gonçalves, and Márcio A. G. Coelho. High coral bycatch in bottomset gillnet coastal fisheries reveals rich coral habitats in Southern Portugal. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Catala:2020:AAP

[1250] Teresa S. Catalá, Pamela E. Rossel, Félix Álvarez-Gómez, Jan Tebben, Félix L. Figueroa, and Thorsten Dittmar. Antioxidant activity and phenolic content of marine dissolved organic matter and their relation to molecular composition. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Idan:2020:SCC

[1251] Tal Idan, Liron Goren, Sigal Shefer, and Micha Ilan. Sponges in a changing climate: Survival of Agelas oroides in a warming Mediterranean Sea. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Dahdouh-Guebas:2020:PPM

[1252] Farid Dahdouh-Guebas, Gordon N. Ajonina, A. Aldrie Amir, Dominic A. Andradi-Brown, Irfan Aziz, Thorsten Balke, Edward B. Barbier, Stefano Cannicci, Simon M. Cragg, Marília Cunha-Lignon, David J. Curnick, Carlos M. Duarte, Norman C. Duke, Charlie Endsor, Sara Fratini, Ilka C. Feller, François Fromard, Jean Hugé, Mark Huxham, James G. Kairo, Tadashi Kajita, Kandasamy Kathiresan, Nico Koedam, Shing Yip Lee, Hsing-Juh Lin, Jock R. Mackenzie, Mwita M. Mangora, Cyril Marchand, Tarik Meziane, Todd E. Minchinton, Nathalie Pettorelli, Jaime Polanía, Gianluca Polgar, Meenakshi Poti, Jurgenne Primavera, Alfredo Quarto, Stefanie M. Rog, Behara Satyanarayana, Yara Schaeffer-Novelli, Mark Spalding, Tom Van der Stocken, Dominic Wodehouse, Jean W. H. Yong, Martin Zimmer, and Daniel A. Friess. Public perceptions of mangrove forests matter for their conservation. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Cziesielski:2021:IBN

[1253] Maha J. Cziesielski, Carlos M. Duarte, Nojood Aalismail, Yousef Al-Hafedh, Andrea Anton, Faiyah Baalkhuyur, Andrew C. Baker, Thorsten Balke, Iliana B. Baums, Michael Berumen, Vasiliki I. Chalastani, Brendan Cornwell, Daniele Daffonchio, Karen Diele, Ehtesaam Farooq, Jean-Pierre Gattuso, Song He, Catherine E. Lovelock, Elizabeth Mcleod, Peter I. Macreadie, Nuria Marba, Cecilia Martin, Marcelle Muniz-Barreto, Kirshnakumar P. Kadinijappali, Perdana Prihartato, Lotfi Rabaoui, Vincent Saderne, Sebastian Schmidt-Roach, David J. Suggett, Michael Sweet, John Statton, Sam Teicher, Stacey M. Trevathan-Tackett, Thadickal V. Joydas, Razan Yahya, and Manuel Aranda. Investing in blue natural capital to secure a future for the Red Sea ecosystems. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Fuhr:2021:TWM

[1254] Michael Fuhr, Georgi Laukert, Yang Yu, Dirk Nürnberg, and Martin Frank. Tracing water mass mixing from the equatorial to the North Pacific Ocean with dissolved neodymium isotopes and concentrations. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Chow:2020:ASR

[1255] Jennifer Chow, Andrew J. Thompson, Fahad Iqbal, Wali Zaidi, and Naweed I. Syed. The antidepressant sertraline reduces synaptic transmission efficacy and synaptogenesis between identified Lymnaea neurons. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Chen:2021:CCN

[1256] Hao Chen, Lusheng Xin, Lin Wang, Huan Zhang, Rui Liu, Hao Wang, Xue Qiao, Lingling Wang, and Linsheng Song. Ca^{2+} /Calmodulin-NOS/ NO-TNFs pathway hallmarks the inflammation response of oyster during aerial exposure. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Hamylton:2020:WCA

[1257] Sarah M. Hamylton, Zhexuan Zhou, and Lei Wang. What can artificial intelligence offer coral reef managers? Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Steffen:2020:MBM

[1258] Jennifer B. M. Steffen, Halina I. Falfushynska, Helen Piontkivska, and Inna M. Sokolova. Molecular biomarkers of the mitochondrial quality control are differently affected by hypoxia-reoxygenation stress in marine bivalves Crassostrea gigas and Mytilus edulis. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Dong:2020:UFA

[1259] Jingyi Dong, Xinpeng Fan, Tengyue Zhang, Saleh A. Al-Farraj, Thorsten Stoeck, Honggang Ma, and Lifang Li. Ultrastructural features of an abundant and ubiquitous marine ciliate, Uronychia binucleata (Protista, Ciliophora, Euplotida). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Zdun:2021:SSC

[1260] Agnieszka Zdun, Joanna Stoń-Egiert, Dariusz Ficek, and Mirosława Ostrowska. Seasonal and spatial changes of primary production in the Baltic Sea (Europe) based on in situ measurements in the period of 1993–2018. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Liu:2020:TMP

[1261] Mingjian Liu, Chundi Wang, Xiaozhong Hu, Zhishuai Qu, Limin Jiang, Saleh A. Al-Farraj, Hamed A. El-Serehy, Alan Warren, and Weibo Song. Taxonomy and molecular phylogeny of three species of scuticociliates from China: Citrithrix smalli gen. nov., sp. nov., Homalogastra binucleata sp. nov. and Uronema orientalis Pan et al., 2015 (Protozoa, Ciliophora, Oligohymenophorea), with the proposal of a new family, Citrithrixidae fam. nov. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Beemelmanns:2021:DMD

[1262] Anne Beemelmanns, Laia Ribas, Dafni Anastasiadi, Javier Moraleda-Prados, Fábio S. Zanuzzo, Matthew L. Rise, and A. Kurt Gamperl. DNA methylation dynamics in Atlantic salmon (Salmo salar) challenged with high temperature and moderate hypoxia. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Tanioka:2020:TDS

[1263] Tatsuro Tanioka, Cédric G. Fichot, and Katsumi Matsumoto. Toward determining the spatio-temporal variability of upper-ocean ecosystem stoichiometry from satellite remote sensing. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Katneni:2020:NIS

[1264] Vinaya Kumar Katneni, Mudagandur S. Shekhar, Ashok Kumar Jangam, Sudheesh K. Prabhudas, Karthic Krishnan, Nimisha Kaikkolante, Balasubramanian C. Paran, Dushyant Singh Baghel, Vijayan K. Koyadan, Joykrushna Jena, and Trilochan Mohapatra. Novel isoform sequencing based full-length transcriptome resource for Indian white shrimp, Penaeus indicus. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Turon:2020:MTE

[1265] Marta Turon, Carlos Angulo-Preckler, Adrià Antich, Kim Præbel, and Owen S. Wangensteen. More than expected from old sponge samples: a natural sampler DNA metabarcoding assessment of marine fish diversity in Nha Trang Bay (Vietnam). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Morrison:2020:FCD

[1266] Katelin M. Morrison, Heidi Kristina Meyer, Emyr Martyn Roberts, Hans Tore Rapp, Ana Colaço, and Christopher Kim Pham. The first cut is the deepest: Trawl effects on a deep-sea sponge ground are pronounced four years on. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Garcia-Sanchez:2021:VHR

[1267] Guillermo García-Sánchez, Ana M. Mancho, Antonio G. Ramos, Josep Coca, Begoña Pérez-Gómez, Enrique Álvarez-Fanjul, Marcos G. Sotillo, Manuel García-León, Víctor J. García-Garrido, and Stephen Wiggins. Very high resolution tools for the monitoring and assessment of environmental hazards in coastal areas. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Albuquerque:2021:TSN

[1268] Rui Albuquerque, Antonio Bode, Juan Ignacio González-Gordillo, Carlos M. Duarte, and Henrique Queiroga. Trophic structure of neuston across tropical and subtropical oceanic provinces assessed with stable isotopes. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Suffridge:2020:EVB

[1269] Christopher P. Suffridge, Luis M. Bolaños, Kristin Bergauer, Alexandra Z. Worden, Jeff Morré, Michael J. Behrenfeld, and Stephen J. Giovannoni. Exploring vitamin B1 cycling and its connections to the microbial community in the North Atlantic Ocean. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Chagaris:2020:ERP

[1270] David Chagaris, Katie Drew, Amy Schueller, Matt Cieri, Joana Brito, and Andre Buchheister. Ecological reference points for Atlantic menhaden established using an ecosystem model of intermediate complexity. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Olsen:2020:EFO

[1271] Erik Olsen, Isaac C. Kaplan, Cecilie Hansen, Elizabeth Fulton, Michael J. Fogarty, Jamie C. Tam, and Saskia A. Otto. Editorial: Future oceans under multiple stressors: From global change to anthropogenic impact. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Hsu:2020:STO

[1272] Tsai-Hsuan Tony Hsu, Lilian Carlu, Yunli Eric Hsieh, Tzu-Yu Angel Lai, Ching-Wei Wang, Ching-Yun Huang, Shan-Hua Yang, Pei-Ling Wang, Nicolas Sturaro, and Vianney Denis. Stranger things: Organismal traits of two octocorals associated with singular Symbiodiniaceae in a highlatitude coral community from Northern Taiwan. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Nachtsheim:2021:SCH

[1273] Dominik A. Nachtsheim, Sacha Viquerat, Nadya C. Ramírez-Martínez, Bianca Unger, Ursula Siebert, and Anita Gilles. Small cetacean in a human high-use area: Trends in harbor porpoise abundance in the North Sea over two decades. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Manjon-Cabeza:2021:AOA

[1274] Maria Eugenia Manjón-Cabeza, Pilar Ríos, Laura María García-Guillén, Aurora Macías-Ramírez, Francisco Sánchez, Augusto Rodríguez-Básalo, Teodoro P. Ibarrola, and Javier Cristobo. Asteroids and ophiuroids associated with sponge aggregations as a key to marine habitats. A comparative analysis between Avilés Canyons System and El Cachucho, Marine Protected Area. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Sejr:2021:SSF

[1275] Mikael K. Sejr, Kim N. Mouritsen, Dorte Krause-Jensen, Birgit Olesen, Martin Emil Blicher, and Jakob Thyrring. Small scale factors modify impacts of temperature, ice scour and waves and drive rocky intertidal community structure in a Greenland fjord. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Alory:2021:CUL

[1276] Gaël Alory, Casimir Yélognissè Da-Allada, Sandrine Djakouré, Isabelle Dadou, Julien Jouanno, and Dorelle Prudence Loemba. Coastal upwelling limitation by onshore geostrophic flow in the Gulf of Guinea around the Niger River plume. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

vanSebille:2021:DSD

[1277] Erik van Sebille, Erik Zettler, Nicolas Wienders, Linda Amaral-Zettler, Shane Elipot, and Rick Lumpkin. Dispersion of surface drifters in the tropical Atlantic. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Smith:2020:LTE

[1278] David R. Smith, Joshua J. Newhard, Conor P. McGowan, and C. Alyssa Butler. The long-term effect of bleeding for limulus amebocyte lysate on annual survival and recapture of tagged horseshoe crabs. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Howell:2021:CES

[1279] Daniel Howell, Amy M. Schueller, Jacob W. Bentley, Andre Buchheister, David Chagaris, Matthew Cieri, Katie Drew, Mathieu G. Lundy, Debbi Pedreschi, David G. Reid, and Howard Townsend. Combining ecosystem and single-species modeling to provide ecosystem-based fisheries management advice within current management systems. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See commentary [2600].

Nemoy:2021:NAS

[1280] Philip Nemoy, Ehud Spanier, and Dror L. Angel. Nitrification activity of the sponge Chondrosia reniformis under elevated concentrations of ammonium. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Boivin-Rioux:2021:PEC

[1281] Aude Boivin-Rioux, Michel Starr, Joël Chassé, Michael Scarratt, William Perrie, and Zhenxia Long. Predicting the effects of climate change on the occurrence of the toxic dinoflagellate Alexandrium catenella along Canada’s East Coast. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ardyna:2020:UIP

[1282] Mathieu Ardyna, C. J. Mundy, Nicolas Mayot, Lisa C. Matthes, Laurent Oziel, Christopher Horvat, Eva Leu, Philipp Assmy, Victoria Hill, Patricia A. Matrai, Matthew Gale, Igor A. Melnikov, and Kevin R. Arrigo. Under-ice phytoplankton blooms: Shedding light on the “Invisible” part of Arctic primary production. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Sorenson:2020:RPB

[1283] Eva Sörenson, Hanna Farnelid, Elin Lindehoff, and Catherine Legrand. Resource partitioning between phytoplankton and bacteria in the coastal Baltic Sea. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

deOliveira:2020:LGD

[1284] Vanessa K. M. de Oliveira, Drienne M. Faria, Haydée A. Cunha, Teresa E. C. dos Santos, Adriana C. Colosio, Lupércio A. Barbosa, Mylla Carla C. Freire, and Ana Paula C. Farro. Low genetic diversity of the endangered Franciscana (Pontoporia blainvillei) in its northernmost, isolated population (FMAIa, Espírito Santo, Brazil). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Guragain:2020:PHR

[1285] Prashanna Guragain, Bjørnar Sporsheim, Astrid Skjesol, Anna Solvang Båtnes, Yngvar Olsen, Atle M. Bones, and Per Winge. Phenylalanine hydroxylase RNAi knockdown negatively affects larval development, molting and swimming performance of salmon lice. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Liang:2020:DPB

[1286] Qihang Liang, Xianpeng Su, Fang Wang, Baishan Zhu, and Mingdi He. The developmental plasticity of boldness and aggressiveness in juvenile and adult swimming crab (Portunus trituberculatus). Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Yebra:2020:ZBD

[1287] Lidia Yebra, Estefanía Espejo, Sébastien Putzeys, Ana Giráldez, Francisco Gómez-Jakobsen, Pablo León, Soluna Salles, Pedro Torres, and Jesús M. Mercado. Zooplankton biomass depletion event reveals the importance of small pelagic fish top-down control in the Western Mediterranean coastal waters. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Serodio:2020:EAC

[1288] João Serôdio, David M. Paterson, Vona Méléder, and Wim Vyverman. Editorial: Advances and challenges in microphytobenthos research: From cell biology to coastal ecosystem function. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Lamont:2021:VSC

[1289] Margaret M. Lamont and Darren Johnson. Variation in species composition, size and fitness of two multi-species sea turtle assemblages using different neritic habitats. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Gallagher:2021:SCD

[1290] Austin J. Gallagher, Oliver N. Shipley, Maurits P. M. van Zinnicq Bergmann, Jacob W. Brownscombe, Craig P. Dahlgren, Michael G. Frisk, Lucas P. Griffin, Neil Hammerschlag, Sami Kattan, Yannis P. Papastamatiou, Brendan D. Shea, Steven T. Kessel, and Carlos M. Duarte. Spatial connectivity and drivers of shark habitat use within a large Marine Protected Area in the Caribbean, The Bahamas Shark Sanctuary. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ioannou:2020:TDE

[1291] Artemis Ioannou, Alexandre Stegner, Franck Dumas, and Briac Le Vu. Three-dimensional evolution of mesoscale anticyclones in the lee of Crete. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Zhu:2021:PCU

[1292] Xia Zhu. The plastic cycle — an unknown branch of the carbon cycle. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Ardiningsih:2021:ISF

[1293] Indah Ardiningsih, Kechen Zhu, Pablo Lodeiro, Martha Gledhill, Gert-Jan Reichart, Eric P. Achterberg, Rob Middag, and Loes J. A. Gerringa. Iron speciation in Fram Strait and over the Northeast Greenland Shelf: an inter-comparison study of voltammetric methods. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Jusadi:2021:PUM

[1294] Dedi Jusadi, Julie Ekasari, Muhammad Agus Suprayudi, Mia Setiawati, and Ichsan Achmad Fauzi. Potential of underutilized marine organisms for aquaculture feeds. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Sandulli:2021:EEB

[1295] Roberto Sandulli, Jeroen Ingels, Daniela Zeppilli, Andrew Kvassnes Sweetman, Sarah Hardy Mincks, Furu Mienis, and Wei Chin-Lin. Editorial: Extreme benthic communities in the age of global change. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Kim:2021:TSV

[1296] Ho-Jung Kim, Hyung Jeek Kim, Eun-Jin Yang, Kyoung-Ho Cho, Jinyoung Jung, Sung-Ho Kang, Kyung-Eun Lee, Sosul Cho, and Dongseon Kim. Temporal and spatial variations in particle fluxes on the Chukchi Sea and East Siberian Sea slopes from 2017 to 2018. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Li:2021:EGP

[1297] Shuangfei Li, Boyu Wang, Liangxu Liu, Yingjie Song, Chuhan Lv, Xingyu Zhu, Yanqing Luo, Christopher H. K. Cheng, Huapu Chen, Xuewei Yang, and Tao Li. Enhanced growth performance physiological and biochemical indexes of Trachinotus ovatus fed with marine microalgae Aurantiochytrium sp. rich in n-3 polyunsaturated fatty acids. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Zhou:2021:CFT

[1298] Shijie Zhou, Serge M. Garcia, Jeppe Kolding, and Michael J. Plank. Commentary: Fishing without a trace? Assessing the balanced harvest approach using EcoTroph. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See [612].

Kohlbach:2021:MTM

[1299] Doreen Kohlbach, Haakon Hop, Anette Wold, Katrin Schmidt, Lukas Smik, Simon T. Belt, Amalia Keck Al-Habahbeh, Matthias Woll, Martin Graeve, Anna Maria Dabrowska, Agnieszka Tatarek, Angus Atkinson, and Philipp Assmy. Multiple trophic markers trace dietary carbon sources in Barents Sea zooplankton during late summer. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Bang:2021:QCR

[1300] Ashley H. Y. Bang, Chao-Yang Kuo, Colin Kuo-Chang Wen, Kah-Leng Cherh, Ming-Jay Ho, Nien-Yun Cheng, Yen-Chia Chen, and Chaolun Allen Chen. Quantifying coral reef resilience to climate change and human development: an evaluation of multiple empirical frameworks. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See corrigendum [1954].

Hammann:2021:OPD

[1301] Liv Hammann, Brian Silliman, and Bernd Blasius. Optimal planting distance in a simple model of habitat restoration with an Allee effect. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Idan:2020:DDM

[1302] Tal Idan, Liron Goren, Sigal Shefer, Itzchak Brickner, and Micha Ilan. Does depth matter? Reproduction pattern plasticity in two common sponge species found in both mesophotic and shallow waters. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Zhang:2021:ACR

[1303] Kui Zhang, Jun Zhang, Dengfu Shi, and Zuozhi Chen. Assessment of coral reef fish stocks from the Nansha Islands, South China Sea, using length-based Bayesian biomass estimation. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Zhang:2020:SMG

[1304] Tengyue Zhang, Yurui Wang, Ting Cheng, Jiyang Ma, Peter Vd’ačný, Weibo Song, and Chen Shao. Systematics and multi-gene phylogeny of the subfamily Nothoholostichinae (Ciliophora, Hypotrichia), with integrative description of a new marine species Nothoholosticha luporinii n. sp. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Jones:2021:SSR

[1305] Hollis R. Jones, Rebecca A. Briggs, Alison Krepp, and Elizabeth Rohring. Strategies for successful research to application projects: a case study of the National Sea Grant College Program. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Spisla:2021:ELO

[1306] Carsten Spisla, Jan Taucher, Lennart T. Bach, Mathias Haunost, Tim Boxhammer, Andrew L. King, Bettany D. Jenkins, Joselynn R. Wallace, Andrea Ludwig, Jana Meyer, Paul Stange, Fabrizio Minutolo, Kai T.

Lohbeck, Alice Nauendorf, Verena Kalter, Silke Lischka, Michael Sswat, Isabel Dörner, Stefanie M. H. Ismar-Rebitz, Nicole Aberle, Jaw C. Yong, Jean-Marie Bouquet, Anna K. Lechtenbörger, Peter Kohnert, Michael Krudewig, and Ulf Riebesell. Extreme levels of ocean acidification restructure the plankton community and biogeochemistry of a temperate coastal ecosystem: a mesocosm study. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Rahav:2020:NES

[1307] Eyal Rahav, Ofrat Raveh, Keren Yanuka-Golub, Natalia Belkin, Peleg Astrahan, Meor Maayani, Noam Tsumi, Yael Kiro, Barak Herut, Jacob Silverman, and Dror L. Angel. Nitrate-enrichment structures phytoplankton communities in the shallow Eastern Mediterranean coastal waters. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Davie-Martin:2020:SSV

[1308] Cleo L. Davie-Martin, Stephen J. Giovannoni, Michael J. Behrenfeld, William B. Penta, and Kimberly H. Halsey. Seasonal and spatial variability in the biogenic production and consumption of volatile organic compounds (VOCs) by marine plankton in the North Atlantic Ocean. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Goren:2021:MIM

[1309] Liron Goren, Tal Idan, Sigal Shefer, and Micha Ilan. Macrofauna inhabiting massive demosponges from shallow and mesophotic habitats along the Israeli Mediterranean Coast. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Blasco:2020:SGC

[1310] Gordon D. Blasco, Danielle M. Ferraro, Richard S. Cottrell, Benjamin S. Halpern, and Halley E. Froehlich. Substantial gaps in the current fisheries data landscape. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Bigorre:2021:ACA

[1311] Sebastien P. Bigorre and Albert J. Plueddemann. The annual cycle of air–sea fluxes in the northwest tropical Atlantic. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Reverter:2020:MMB

[1312] Miriam Reverter, Sven Rohde, Christelle Parchemin, Nathalie Tapissier-Bontemps, and Peter J. Schupp. Metabolomics and marine biotechnology: Coupling metabolite profiling and organism biology for the discovery of new compounds. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Cardenas:2020:SMD

[1313] Paco Cárdenas. Surface microornamentation of demosponge sterraster spicules, phylogenetic and paleontological implications. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Martins:2021:MMC

[1314] Eva Martins, Hans Tore Rapp, Joana R. Xavier, Gabriela S. Diogo, Rui L. Reis, and Tiago H. Silva. Macro and microstructural characteristics of North Atlantic deep-sea sponges as bioinspired models for tissue engineering scaffolding. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Leitao:2020:EVF

[1315] Francisco Leitão, Ruben Hernan Roa-Ureta, and Fernando Cánovas. Editorial: Vulnerability of fisheries to climate change. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745.

Estevez:2020:MCC

[1316] Daniel Estévez, Eugene Kozminsky, Antonio Carvajal-Rodríguez, Armando Caballero, Rui Faria, Juan Galindo, and Emilio Rolán-Alvarez. Mate choice contributes to the maintenance of shell color polymorphism in a marine snail via frequency-dependent sexual selection. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Ogundare:2021:VSA

[1317] Margaret Ojone Ogundare, Agneta Fransson, Melissa Chierici, Warren R. Joubert, and Alakendra N. Roychoudhury. Variability of sea-air carbon dioxide flux in autumn across the Weddell Gyre and offshore Dronning Maud Land in the Southern Ocean. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Wysocki:2020:VSM

[1318] Ina Tessnow von Wysocki and Alice B. M. Vadrot. The voice of science on marine biodiversity negotiations: a systematic literature review. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Dayanidhi:2021:EDM

[1319] Divya L. Dayanidhi, Beatrice C. Thomas, Joshua S. Osterberg, Mallissa Vuong, Giselle Vargas, Sarah K. Kwartler, Emma Schmaltz, Meagan M.

Dunphy-Daly, Thomas F. Schultz, Daniel Rittschof, William C. Eward, Christopher Roy, and Jason A. Somarelli. Exploring the diversity of the marine environment for new anti-cancer compounds. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Bouchet:2021:RAC

[1320] Phil J. Bouchet, Deborah Thiele, Sarah A. Marley, Kelly Waples, and Frank Weisenberger. Regional assessment of the conservation status of snubfin dolphins (Orcaella heinsohni) in the Kimberley Region, Western Australia. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Nonaka:2020:CAD

[1321] Masami Nonaka, Hideharu Sasaki, Bunmei Taguchi, and Niklas Schneider. Corrigendum: Atmospheric-driven and intrinsic interannual-to-decadal variability in the Kuroshio Extension jet and eddy activities. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See [899].

Cavanagh:2021:FRS

[1322] Rachel D. Cavanagh, Jess Melbourne-Thomas, Susie M. Grant, David K. A. Barnes, Kevin A. Hughes, Svenja Halfter, Michael P. Meredith, Eugene J. Murphy, Rowan Trebilco, and Simeon L. Hill. Future risk for Southern Ocean ecosystem services under climate change. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See corrigendum [2601].

Asghar:2021:MCD

[1323] Usman Asghar, Yong Chi, Yunyi Gao, Borong Lu, Yaohan Jiang, Ruitao Gong, Honggang Ma, Khaled A. S. Al-Rasheid, and Feng Gao. Morphogenesis of the ciliature during sexual process of conjugation in the ciliated protist euplotes raikovi. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Liu:2020:NCT

[1324] Weiwei Liu, Jiamei Jiang, Yehui Tan, and Xiaofeng Lin. Novel contributions to the taxonomy of the ciliates genus Euplotes (Ciliophora, Euplotida): Redescription of two poorly known species, with a brief note on the distributions of this genus in coastal waters of Southern China. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

vonAmmon:2020:PCS

[1325] Ulla von Ammon, Andrew Jeffs, Anastasija Zaiko, Aimee van der Reis, Deb Goodwin, Lynnath E. Beckley, Emmanuel Malpot, and Xavier Pochon. A portable cruising speed net: Expanding global collection of sea surface plankton data. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Kleinhaus:2020:CWO

[1326] Karine Kleinhaus, Christian R. Voolstra, Anders Meibom, Yael Amitai, Hezi Gildor, and Maoz Fine. A closing window of opportunity to save a unique marine ecosystem. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Ma:2021:MMP

[1327] Mingzhen Ma, Yuqing Li, Honggang Ma, Khaled A. S. Al-Rasheid, Alan Warren, Yurui Wang, and Ying Yan. Morphology and molecular phylogeny of four trachelocercid ciliates (Protozoa, Ciliophora, Karyorelictea) found in marine coastal habitats of Northern China, with description of a new genus, two new species and a new combination. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See corrigendum [3072].

Amitai:2021:ESD

[1328] Yael Amitai, Yosef Ashkenazy, and Hezi Gildor. The effect of the source of deep water in the Eastern Mediterranean on Western Mediterranean intermediate and deep water. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Carrasco:2021:DSA

[1329] Sergio A. Carrasco, Macarena Bravo, Christian M. Ibáñez, and Germán Zapata-Hernández. Discrete spawning aggregations of the loliginid squid Doryteuthis gahi reveal life-history interactions of a dwarf morphotype at the center of its distribution range. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Langenkamper:2020:CPB

[1330] Daniel Langenkämper, Martin Zurowietz, Timm Schoening, and Tim W. Nattkemper. Corrigendum: BIIGLE 2.0 —- browsing and annotating large marine image collections. Frontiers in Marine Science, 7, October 2020. ISSN 2296-7745. See [2].

Krause-Jensen:2020:ICC

[1331] Dorte Krause-Jensen, Philippe Archambault, Jorge Assis, Inka Bartsch, Kai Bischof, Karen Filbee-Dexter, Kenneth H. Dunton, Olga Maximova, Sunna Björk Ragnarsdóttir, Mikael K. Sejr, Uliana Simakova, Vassily Spiridonov, Susse Wegeberg, Mie H. S. Winding, and Carlos M. Duarte.

Imprint of climate change on pan-Arctic marine vegetation. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745. See corrigendum [2531].

Chabanne:2021:IBS

[1332] Delphine B. H. Chabanne, Simon J. Allen, William Bruce Sherwin, Hugh Finn, and Michael Krützen. Inconsistency between socio-spatial and genetic structure in a coastal dolphin population. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Solano:2021:UWR

[1333] Neyra Solano, Ines Lopez-Ercilla, Francisco J. Fernandez-Rivera Melo, and Jorge Torre. Unveiling women’s roles and inclusion in Mexican Small-Scale Fisheries (SSF). Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See corrigendum [2188].

Martinez:2020:CRG

[1334] Elodie Martinez, Thomas Gorgues, Matthieu Lengaigne, Clement Fontana, Raphaëlle Sauzède, Christophe Menkes, Julia Uitz, Emanuele Di Lorenzo, and Ronan Fablet. Corrigendum: Reconstructing global chlorophyll-a variations using a non-linear statistical approach. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See [566].

Looi:2021:ECH

[1335] Ley Juen Looi, Ahmad Zaharin Aris, Noorain Mohd Isa, Fatimah Md. Yusoff, and Hazzeman Haris. Elemental composition and health risk assessment of giant mudskipper (Periophthalmodon schlosseri) from the intertidal zone of the West Coast of Peninsular Malaysia. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Shan:2021:DBF

[1336] Binbin Shan, Yan Liu, Changping Yang, Yu Zhao, Gongjun Zhang, Qiaer Wu, and Dianrong Sun. DNA barcoding of fish in Mischief Reef — fish diversity of a reef fish community from Nansha Islands. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Booty:2020:CSA

[1337] James M. Booty, Graham J. C. Underwood, Amie Parris, Richard G. Davies, and Trevor J. Tolhurst. Corrigendum: Shorebirds affect ecosystem functioning on an intertidal mudflat. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745. See [787].

Yan:2021:GDI

[1338] Wei Yan, Xuejin Feng, Wei Zhang, Muhammad Zohaib Nawaz, Tingwei Luo, Rui Zhang, and Nianzhi Jiao. Genomes of diverse isolates of Prochlorococcus high-light-adapted clade II in the Western Pacific Ocean. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Silberberger:2021:SIM

[1339] Marc J. Silberberger, Katarzyna Koziorowska-Makuch, Karol Kuliński, and Monika Kedra. Stable isotope mixing models are biased by the choice of sample preservation and pre-treatment: Implications for studies of aquatic food webs. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Janout:2020:CVS

[1340] Markus A. Janout, Jens Hölemann, Georgi Laukert, Alexander Smirnov, Thomas Krumpen, Dorothea Bauch, and Leonid Timokhov. Corrigendum: On the variability of stratification in the freshwater-influenced Laptev Sea region. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See [879].

Luo:2021:TTO

[1341] Xiaotian Luo, Jie Huang, William A. Bourland, Hamed A. El-Serehy, Saleh A. Al-Farraj, Xumiao Chen, and Weibo Song. Taxonomy of three oxytrichids (Protozoa, Ciliophora, Hypotrichia), with establishment of the new species Rubrioxytricha guangzhouensis spec. nov. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Zhou:2021:CWC

[1342] Jun Zhou, Zhuo-Yi Zhu, Huan-Ting Hu, Gui-Ling Zhang, and Qian-Qian Wang. Clarifying water column respiration and sedimentary oxygen respiration under oxygen depletion off the Changjiang Estuary and adjacent East China Sea. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

McHugh:2021:SAL

[1343] Katherine A. McHugh, Aaron A. Barleycorn, Jason B. Allen, Kim Bassos-Hull, Gretchen Lovewell, Denise Boyd, Anna Panike, Carolyn Cush, Deborah Fauquier, Blair Mase, Robert C. Lacy, Michelle R. Greenfield, Daniel I. Rubenstein, Ann Weaver, Abby Stone, Lisa Oliver, Kent Morse, and Randall S. Wells. Staying alive: Long-term success of bottlenose dolphin interventions in Southwest Florida. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Exeter:2021:SLD

[1344] Owen M. Exeter, Thaung Htut, Christopher R. Kerry, Maung Maung Kyi, Me’ira Mizrahi, Rachel A. Turner, Matthew J. Witt, and Anthony W. J. Bicknell. Shining light on data-poor coastal fisheries. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Fu:2021:ERE

[1345] Zhengyi Fu, Rui Yang, Shengjie Zhou, Zhenhua Ma, and Tao Zhang. Effects of rotifers enriched with different enhancement products on larval performance and jaw deformity of golden pompano larvae Trachinotus ovatus (Linnaeus, 1758). Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745.

Hjerne:2020:CCD

[1346] Olle Hjerne, Susanna Hajdu, Ulf Larsson, Andrea S. Downing, and Monika Winder. Corrigendum: Climate driven changes in timing, composition and magnitude of the Baltic Sea phytoplankton spring bloom. Frontiers in Marine Science, 7, November 2020. ISSN 2296-7745. See [7].

Araujo:2021:CSA

[1347] Rita Araújo, Fatima Vázquez Calderón, Javier Sánchez López, Isabel Costa Azevedo, Annette Bruhn, Silvia Fluch, Manuel Garcia Tasende, Fatemeh Ghaderiardakani, Tanel Ilmjärv, Martial Laurans, Micheal Mac Monagail, Silvio Mangini, César Peteiro, Céline Rebours, Tryggvi Stefansson, and Jörg Ullmann. Current status of the algae production industry in Europe: an emerging sector of the blue bioeconomy. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Wakeham:2020:ECT

[1348] Stuart G. Wakeham, Cindy Lee, Carol Arnosti, Thorsten Dittmar, Kai-Uwe Hinrichs, and Ann Pearson. Editorial: Current topics in marine organic biogeochemical research. Frontiers in Marine Science, 7, December 2020. ISSN 2296-7745.

Chen:2021:SLC

[1349] Peng Chen, Jiahao Chen, and Zhiqiang Hu. Software-in-the-loop combined reinforcement learning method for dynamic response analysis of FOWTs. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Bastardie:2021:RCE

[1350] Francois Bastardie, Elliot J. Brown, Eider Andonegi, Robert Arthur, Esther Beukhof, Jochen Depestele, Ralf Döring, Ole Ritzau Eigaard, Isabel García-Barón, Marcos Llope, Hugo Mendes, GerJan Piet, and David Reid. A review characterizing 25 ecosystem challenges to be addressed by an ecosystem approach to fisheries management in Europe. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Fernandez-Alvarez:2021:MMA

[1351] Fernando Á. Fernández-Álvarez, Pilar Sánchez, and Roger Villanueva. Morphological and molecular assessments of bobtail squids (Cephalopoda: Sepiolidae) reveal a hidden history of biodiversity. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Skirving:2021:EFC

[1352] William Skirving, Andrew Pomeroy, Robert McCall, John Marra, and Curt Storlazzi. Editorial: Flooding on coral reef-lined coasts: Current state of knowledge and future challenges. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Gernez:2021:ERS

[1353] Pierre Gernez, Stephanie C. J. Palmer, Yoann Thomas, and Rodney Forster. Editorial: Remote sensing for aquaculture. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Kim:2021:CUS

[1354] Sang Wha Kim, Won Hee Hong, Se Jin Han, Jun Kwon, Heejun Ko, Sung Bin Lee, Sib Sankar Giri, Sang Guen Kim, Byung Yeop Kim, Goo Jang, Byeong Chun Lee, Dong Wan Kim, and Se Chang Park. Corrigendum: Use of synthetic salmon GnRH and domperidone (ovaprim(R)) in sharks: Preparation for ex situ conservation. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745. See [1008].

Saha:2021:ECM

[1355] Mahasweta Saha, Elisa Berdalet, Phillipe Potin, and Bernardo A. P. Da Gama. Editorial: Chemically mediated interactions between marine macrophytes and microbes. Frontiers in Marine Science, 7, January 2021. ISSN 2296-7745.

Romano:2021:CWS

[1356] Chiara Romano, Amandine Nunes-Jorge, Nadine Le Bris, Greg W. Rouse, Daniel Martin, and Christian Borowski. Corrigendum: Wooden stepping stones: Diversity and biogeography of deep-sea wood boring Xylophagaidae (Mollusca: Bivalvia) in the North-East Atlantic Ocean, with the description of a new genus. Frontiers in Marine Science, 7, February 2021. ISSN 2296-7745. See [1080].

Taylor:2021:DPG

[1357] James Taylor, Colin Devey, Morgane Le Saout, Sven Petersen, Tom Kwasnitschka, Inmaculada Frutos, Katrin Linse, Anne-Nina Lörz, Dominik Pałgan, Anne H. Tandberg, Jörundur Svavarsson, Daniel Thorhallsson, Adrianna Tomkowicz, Hrönn Egilsdóttir, Stefán Á. Ragnarsson, Jasmin Renz, Elena L. Markhaseva, Sabine Gollner, Eva Paulus, Jon Kongsrud, Jan Beermann, Kevin M. Kocot, Karin Meißner, Alexander Bartholomä, Leon Hoffman, Pauline Vannier, Viggó T. Marteinsson, Hans T. Rapp, Guillermo Díaz-Agras, Ramiro Tato, and Saskia Brix. The discovery and preliminary geological and faunal descriptions of three new Steinahóll Vent Sites, Reykjanes Ridge, Iceland. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ryan-Keogh:2021:PPU

[1358] Thomas J. Ryan-Keogh and Charlotte M. Robinson. Phytoplankton photophysiology utilities: a Python toolbox for the standardization of processing active chlorophyll-a fluorescence data. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gorman:2021:ASI

[1359] Kristen B. Gorman, Kate E. Ruck, Tony D. Williams, and William R. Fraser. Advancing the sea ice hypothesis: Trophic interactions among breeding Pygoscelis penguins with divergent population trends throughout the Western Antarctic Peninsula. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Bisack:2021:SMR

[1360] Kathryn D. Bisack and Gisele M. Magnusson. Spatial management to reduce entanglement risk to North Atlantic right whales in fishing gear: a case study of U.S. Northeast lobster fishery 2002–2009. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sanchez-Jimenez:2021:IVP

[1361] Astrid Sánchez-Jiménez, Douglas MacMillan, Matthias Wolff, Achim Schlüter, and Marie Fujitani. The importance of values in predicting and encouraging environmental behavior: Reflections from a Costa Rican small-scale fishery. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Knutsen:2021:DTM

[1362] Helene Rønquist Knutsen, Sune Riis Sørensen, Peter Munk, Tora Bardal, and Elin Kjørsvik. Digestive tract and the muscular pharynx/esophagus in wild leptocephalus larvae of European eel (Anguilla anguilla). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Carus:2021:UAS

[1363] Jana Carus, Carmen Arndt, Boris Schröder, Moritz Thom, Raúl Villanueva, and Maike Paul. Using artificial seagrass for promoting positive feedback mechanisms in seagrass restoration. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Santamaria:2021:SCG

[1364] Jorge Santamaría, Raül Golo, Emma Cebrian, María García, and Alba Vergés. Stressful conditions give rise to a novel and cryptic filamentous form of Caulerpa cylindracea. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Zhou:2021:SFI

[1365] Mingxi Zhou, Ralf Bachmayer, and Brad DeYoung. Surveying a floating iceberg with the USV SEADRAGON. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Macheriotou:2021:DSN

[1366] Lara Macheriotou, Annelien Rigaux, Karine Olu, Daniela Zeppilli, Sofie Derycke, and Ann Vanreusel. Deep-sea nematodes of the Mozambique Channel: Evidence of random community assembly dynamics in seep sediments. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See corrigendum [2355].

Armelloni:2021:DPA

[1367] Enrico Nicola Armelloni, Martina Scanu, Francesco Masnadi, Gianpaolo Coro, Silvia Angelini, and Giuseppe Scarcella. Data poor approach for the assessment of the main target species of Rapido trawl fishery in Adriatic Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Price:2021:FSH

[1368] David M. Price, Aaron Lim, Alexander Callaway, Markus P. Eichhorn, Andrew J. Wheeler, Claudio Lo Iacono, and Veerle A. I. Huvenne. Finescale heterogeneity of a cold-water coral reef and its influence on the distribution of associated taxa. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Weitzman:2021:CRI

[1369] Benjamin Weitzman, Brenda Konar, Katrin Iken, Heather Coletti, Daniel Monson, Robert Suryan, Thomas Dean, Dominic Hondolero, and Mandy Lindeberg. Changes in rocky intertidal community structure during a marine heatwave in the Northern Gulf of Alaska. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Nabi:2021:DHC

[1370] Ghulam Nabi, Shahid Ahmad, Richard William McLaughlin, Yujiang Hao, Suliman Khan, Naveed Ahmad, Saeed Ahmad, Muhammad Shoaib Kiani, Yuefeng Wu, and Dongming Li. Deteriorating habitats and conservation strategies to repopulate the endangered Indus River dolphin (Platanista gangetica minor); a lesson learned from the conservation practices of the Yangtze finless porpoise (Neophocaena asiaeorientalis). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pezner:2021:LVT

[1371] Ariel K. Pezner, Travis A. Courtney, Heather N. Page, Sarah N. Giddings, Cory M. Beatty, Michael D. DeGrandpre, and Andreas J. Andersson. Lateral, vertical, and temporal variability of seawater carbonate chemistry at Hog Reef, Bermuda. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Campbell:2021:CDD

[1372] Theo I. Campbell, James R. Tweedley, Danielle J. Johnston, and Neil R. Loneragan. Crab diets differ between adjacent estuaries and habitats within a sheltered marine embayment. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Hernandez-Elizarraga:2021:CMA

[1373] Víctor Hugo Hernández-Elizárraga, Norma Olguín-López, Rosalina Hernández-Matehuala, Juan Caballero-Pérez, César Ibarra-Alvarado, and Alejandra Rojas-Molina. Comprehensive metatranscriptome analysis of the reef-building Holobiont Millepora complanata. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Liu:2021:ASG

[1374] Yiwen Liu, Chongliang Zhang, Binduo Xu, Ying Xue, Yiping Ren, and Yong Chen. Accounting for seasonal growth in per-recruit analyses: a case study of four commercial fish in coastal China Seas. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Horn:2021:TIM

[1375] Sabine Horn, Cédric L. Meunier, Vera Fofonova, Karen H. Wiltshire, Subrata Sarker, Bernadette Pogoda, and Harald Asmus. Toward improved model capacities for assessment of climate impacts on coastal bentho-pelagic food webs and ecosystem services. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Alonso:2021:BSD

[1376] David Alonso, Martha Vides-Casado, Francisco Arias-Isaza, Hernando Zambrano, Emilio Rodriguez, Venus Rocha-Gutierrez, Pilar Herron, and Anny Castillo. Behind the scenes for the designation of the Corales de Profundidad National Natural Park of Colombia. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Dayan:2021:HES

[1377] Hugo Dayan, Goneri Le Cozannet, Sabrina Speich, and Rémi Thiéblemont. High-end scenarios of sea-level rise for coastal risk-averse stakeholders. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Albo-Puigserver:2021:CLH

[1378] Marta Albo-Puigserver, Maria Grazia Pennino, Jose María Bellido, Ana Isabel Colmenero, Ana Giráldez, Manuel Hidalgo, John Gabriel Ramírez, Jeroen Steenbeek, Pedro Torres, Marta Cousido-Rocha, and Marta Coll. Changes in life history traits of small pelagic fish in the Western Mediterranean Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cozar:2021:MLW

[1379] Andrés Cózar, Stefano Aliani, Oihane C. Basurko, Manuel Arias, Atsuhiko Isobe, Konstantinos Topouzelis, Anna Rubio, and Carmen Morales-Caselles. Marine litter windrows: a strategic target to understand and manage the ocean plastic pollution. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Piwetz:2021:ECC

[1380] Sarah Piwetz, Thomas A. Jefferson, and Bernd Würsig. Effects of coastal construction on Indo–Pacific humpback dolphin (Sousa chinensis) behavior and habitat-use off Hong Kong. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Tiesi:2021:IWM

[1381] Alessandro Tiesi, Arturo Pucillo, Davide Bonaldo, Antonio Ricchi, Sandro Carniel, and Mario Marcello Miglietta. Initialization of WRF model simulations with Sentinel-1 wind speed for severe weather events. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Yamaguchi:2021:UGB

[1382] Kazuaki Yamaguchi and Shigehiro Kuraku. Unbiasing genome-based analyses of selection: an example using iconic shark species. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See commentary [2473].

Moon:2021:FQU

[1383] Katie Moon, Christopher Cvitanovic, Deborah A. Blackman, Ivan R. Scales, and Nicola K. Browne. Five questions to understand epistemology and its influence on integrative marine research. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Sweet:2021:SSV

[1384] Michael Sweet, Mark Bulling, Dorsa Varshavi, Gavin R. Lloyd, Andris Jankevics, Lukáš Najdekr, Ralf J. M. Weber, Mark R. Viant, and Jamie Craggs. Species-specific variations in the metabolomic profiles of Acropora hyacinthus and Acropora millepora mask acute temperature stress effects in adult coral colonies. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Cosentino:2021:TEC

[1385] Mel Cosentino and Lucrecia Souviron-Priego. Think of the early career researchers! Saving the oceans through collaborations. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Mecho:2021:EDM

[1386] Ariadna Mecho, Boris Dewitte, Javier Sellanes, Simon van Gennip, Erin E. Easton, and Joao B. Gusmao. Environmental drivers of mesophotic echinoderm assemblages of the Southeastern Pacific Ocean. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Soto-Cortes:2021:WWM

[1387] Laura V. Soto-Cortés, Andrea Luna-Acosta, and Diana Lucía Maya. Whale-watching management: Assessment of sustainable governance in Uramba Bahía Málaga National Natural Park, Valle del Cauca. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Diaz:2021:EDS

[1388] María José Díaz, Christian Buschbaum, Paul E. Renaud, and Markus Molis. Effects of detached seaweeds on structure and function of Arctic intertidal soft-bottom communities. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Reed:2021:IGR

[1389] Adam J. Reed, Jasmin A. Godbold, Martin Solan, and Laura J. Grange. Invariant gametogenic response of dominant infaunal bivalves from the Arctic under ambient and near-future climate change conditions. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Vaz:2021:CLE

[1390] Ana C. Vaz, Robin Faillettaz, and Claire B. Paris. A coupled Lagrangian– Earth system model for predicting oil photooxidation. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Lockhart:2021:CAA

[1391] Susanne J. Lockhart and John Hocevar. Combined abundance of all vulnerable marine ecosystem indicator taxa inadequate as sole determiner of vulnerability, Antarctic Peninsula. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Lin:2021:ABE

[1392] Chi hung Lin, Hsing-Juh Lin, Jian-Ping Suen, and Lien-Siang Chou. Association between estuary characteristics and activities of the critically endangered Indo–Pacific humpback dolphin (Sousa chinensis). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Hamon:2021:FSP

[1393] Katell G. Hamon, Cornelia M. Kreiss, John K. Pinnegar, Heleen Bartelings, Jurgen Batsleer, Ignacio A. Catalán, Dimitrios Damalas, Jan-Jaap Poos, Sandra Rybicki, Sevrine F. Sailley, Vasiliki Sgardeli, and Myron A. Peck. Future socio-political scenarios for aquatic resources in Europe: an operationalized framework for marine fisheries projections. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Gundersen:2021:VPS

[1394] Hege Gundersen, Eli Rinde, Trine Bekkby, Kasper Hancke, Janne K. Gitmark, and Hartvig Christie. Variation in population structure and standing stocks of kelp along multiple environmental gradients and implications for ecosystem services. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bastesen:2021:RLC

[1395] Eivind Bastesen, Marte Haave, Gidske L. Andersen, Gaute Velle, Gunhild Bødtker, and Charlotte Gannefors Krafft. Rapid landscape changes in plastic bays along the Norwegian coastline. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Koen-Alonso:2021:CMC

[1396] Mariano Koen-Alonso, Ulf Lindstrøm, and Andrew Cuff. Comparative modeling of cod-capelin dynamics in the Newfoundland–Labrador Shelves and Barents Sea ecosystems. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Clementson:2021:EOY

[1397] Lesley A. Clementson, Anthony J. Richardson, Wayne A. Rochester, Kadija Oubelkheir, Bingqing Liu, Eurico J. D’Sa, Luiz Felipe Mendes Gusmão, Penelope Ajani, Thomas Schroeder, Phillip W. Ford, Michele A. Burford, Emily Saeck, and Andrew D. L. Steven. Effect of a once in 100year flood on a subtropical coastal phytoplankton community. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Mekkes:2021:EOA

[1398] Lisette Mekkes, Guadalupe Sepúlveda-Rodríguez, Gintarė Bielkinaitė, Deborah Wall-Palmer, Geert-Jan A. Brummer, Linda K. Dämmer, Jef Huisman, Emiel van Loon, Willem Renema, and Katja T. C. A. Peijnenburg. Effects of ocean acidification on calcification of the sub-Antarctic pteropod Limacina retroversa. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Rees:2021:MRS

[1399] Christine Rees, Julie Janssens, Kendall Sherrin, Peter Hughes, Stephen Tibben, Merinda McMahon, Jack McDonald, Alicia Camac, Cassie Schwanger, and Andreas Marouchos. Method for reproducible shipboard segmented flow analysis ammonium measurement using an in-house reference material for quality control. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Liu:2021:LTV

[1400] Dan Liu, Yongjun Tian, Shuyang Ma, Jianchao Li, Peng Sun, Zhenjiang Ye, Caihong Fu, Kuowei Lan, and Shijie Zhou. Long-term variability of piscivorous fish in China Seas under climate change with implication for fisheries management. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Dabin:2021:VCE

[1401] Willy Dabin, Bastien Rochowski, Michel Daudon, and Vincent Ridoux. Vaginal calculi in the Eastern North Atlantic common dolphins Delphinus delphis, induction mechanisms and possible effects on fecundity. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Moser:2021:SPI

[1402] Annabell Moser, Iain Pheasant, William N. MacPherson, Bhavani E. Narayanaswamy, and Andrew K. Sweetman. Sediment profile imaging: Laboratory study into the sediment smearing effect of a penetrating plate. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Morganti:2021:SPR

[1403] Teresa Maria Morganti, Marta Ribes, Raz Moskovich, Jeremy Brian Weisz, Gitai Yahel, and Rafel Coma. In situ pumping rate of 20 marine demosponges is a function of osculum area. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See corrigendum [2648].

Xu:2021:DCR

[1404] Jingping Xu, Jianhua Zhao, Fei Wang, Yanlong Chen, and Zhongping Lee. Detection of coral reef bleaching based on Sentinel-2 multi-temporal imagery: Simulation and case study. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Figuerola:2021:RMA

[1405] Blanca Figuerola, Alyce M. Hancock, Narissa Bax, Vonda J. Cummings, Rachel Downey, Huw J. Griffiths, Jodie Smith, and Jonathan S. Stark. A review and meta-analysis of potential impacts of ocean acidification on marine calcifiers from the Southern Ocean. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Meschini:2021:EAG

[1406] Marta Meschini, Francesca Prati, Ginevra A. Simoncini, Valentina Airi, Erik Caroselli, Fiorella Prada, Chiara Marchini, Mariana Machado Toffolo, Simone Branchini, Viviana Brambilla, Claudia Covi, and Stefano Goffredo. Environmental awareness gained during a citizen science project in touristic resorts is maintained after 3 years since participation. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Geraci:2021:ACF

[1407] Michele L. Geraci, Fabio Falsone, Vita Gancitano, Danilo Scannella, Fabio Fiorentino, and Sergio Vitale. Assessing cephalopods fisheries in the Strait of Sicily by using poor data modeling. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Quattrocchi:2021:ONS

[1408] Giovanni Quattrocchi, Simone Simeone, Andrea Pes, Roberto Sorgente, Alberto Ribotti, and Andrea Cucco. An operational numerical system for oil stranding risk assessment in a high-density vessel traffic area. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Miller:2021:CSP

[1409] Melissa A. Miller, Megan E. Moriarty, Pádraig J. Duignan, Tanja S. Zabka, Erin Dodd, Francesca I. Batac, Colleen Young, Angelina Reed, Michael D. Harris, Katherine Greenwald, Raphael M. Kudela, Michael J. Murray, Frances M. D. Gulland, Peter E. Miller, Kendra Hayashi, Catherine T. Gunther-Harrington, Martin T. Tinker, and Sharon Toy-Choutka. Clinical signs and pathology associated with domoic acid toxicosis in southern sea otters (Enhydra lutris nereis). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Barcelo:2021:TVE

[1410] Caren Barceló, Richard D. Brodeur, Lorenzo Ciannelli, Elizabeth A. Daly, Craig M. Risien, Gonzalo S. Saldías, and Jameal F. Samhouri. Time-varying epipelagic community seascapes: Assessing and predicting species composition in the Northeastern Pacific Ocean. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Melaniuk:2021:EFV

[1411] Katarzyna Melaniuk. Effectiveness of fluorescent viability assays in studies of Arctic cold seep foraminifera. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Polte:2021:RRS

[1412] Patrick Polte, Tomas Gröhsler, Paul Kotterba, Lena von Nordheim, Dorothee Moll, Juan Santos, Paco Rodriguez-Tress, Yury Zablotski, and Christopher Zimmermann. Reduced reproductive success of Western Baltic herring (Clupea harengus) as a response to warming winters. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Livore:2021:ASL

[1413] Juan Pablo Livore, María M. Mendez, Eduardo Klein, Lorena Arribas, and Gregorio Bigatti. Application of a simple, low-cost, low-tech method to monitor intertidal rocky shore assemblages on a broad geographic scale. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Barboza:2021:LDG

[1414] Carlos A. M. Barboza, Gustavo Mattos, Abílio Soares-Gomes, Ilana Rosental Zalmon, and Leonardo Lopes Costa. Low densities of the ghost crab Ocypode quadrata related to large scale human modification of sandy shores. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Fujii:2021:CMF

[1415] Masahiko Fujii, Shintaro Takao, Takuto Yamaka, Tomoo Akamatsu, Yamato Fujita, Masahide Wakita, Akitomo Yamamoto, and Tsuneo Ono. Continuous monitoring and future projection of ocean warming, acidification, and deoxygenation on the subarctic coast of Hokkaido, Japan. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Kourantidou:2021:EAN

[1416] Melina Kourantidou, Porter Hoagland, Aaron Dale, and Megan Bailey. Equitable allocations in northern fisheries: Bridging the divide for Labrador Inuit. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Rossi:2021:NGO

[1417] Paolo Rossi, Massimo Ponti, Sara Righi, Cristina Castagnetti, Roberto Simonini, Francesco Mancini, Panagiotis Agrafiotis, Leonardo Bassani, Fabio Bruno, Carlo Cerrano, Paolo Cignoni, Massimiliano Corsini, Pierre Drap, Marco Dubbini, Joaquim Garrabou, Andrea Gori, Nuno Gracias, Jean-Baptiste Ledoux, Cristina Linares, Torcuato Pulido Mantas, Fabio Menna, Erica Nocerino, Marco Palma, Gaia Pavoni, Alessandro Ridolfi, Sergio Rossi, Dimitrios Skarlatos, Tali Treibitz, Eva Turicchia, Matan Yuval, and Alessandro Capra. Needs and gaps in optical underwater technologies and methods for the investigation of marine animal forest 3d-structural complexity. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Mullowney:2021:HMD

[1418] Darrell R. J. Mullowney, Krista D. Baker, and Julia R. Pantin. Hard to manage? Dynamics of soft-shell crab in the Newfoundland and Labrador snow crab fishery. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Liu:2021:MMA

[1419] Kwang-Ming Liu, Chiao-Bin Wu, Shoou-Jeng Joung, Wen-Pei Tsai, and Kuan-Yu Su. Multi-model approach on growth estimation and association with life history trait for elasmobranchs. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bonsell:2021:SCD

[1420] Christina Bonsell and Kenneth H. Dunton. Slow community development enhances abiotic limitation of benthic community structure in a High Arctic kelp bed. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Testa:2021:PBR

[1421] Giovanni Testa, Andrea Piñones, and Leonardo R. Castro. Physical and biogeochemical regionalization of the Southern Ocean and the CCAMLR zone 48.1. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Yasuda:2021:GSG

[1422] Nina Yasuda, Yuko F. Kitano, Hiroki Taninaka, Satoshi Nagai, Takuma Mezaki, and Hiroshi Yamashita. Genetic structure of the Goniopora lobata and G. djiboutiensis species complex is better explained by oceanography than by morphological characteristics. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hernandez-Hernandez:2021:STS

[1423] Nauzet Hernández-Hernández, Yeray Santana-Falcón, Sheila Estrada-Allis, and Javier Arístegui. Short-term spatiotemporal variability in picoplankton induced by a submesoscale front south of Gran Canaria (Canary Islands). Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Feng:2021:HVD

[1424] Li Feng, Qianqian Song, Qun Jiang, and Zhiyong Li. The horizontal and vertical distribution of deep-sea sediments fungal community in the South China Sea. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Bowman:2021:WCM

[1425] Jeff S. Bowman, Benjamin A. S. Van Mooy, Daniel P. Lowenstein, Helen F. Fredricks, Colleen M. Hansel, Rebecca Gast, James R. Collins, Nicole Couto, and Hugh W. Ducklow. Whole community metatranscriptomes and lipidomes reveal diverse responses among Antarctic phytoplankton to changing ice conditions. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Kraft:2021:FAI

[1426] Kaisa Kraft, Jukka Seppälä, Heidi Hällfors, Sanna Suikkanen, Pasi Ylöstalo, Sílvia Anglès, Sami Kielosto, Harri Kuosa, Lauri Laakso, Martti Honkanen, Sirpa Lehtinen, Johanna Oja, and Timo Tamminen. First application of IFCB high-frequency imaging-in-flow cytometry to investigate bloom-forming filamentous cyanobacteria in the Baltic Sea. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Schernewski:2021:UME

[1427] Gerald Schernewski, Hagen Radtke, Rahel Hauk, Christian Baresel, Mikael Olshammar, and Sonja Oberbeckmann. Urban microplastics emissions: Effectiveness of retention measures and consequences for the Baltic Sea. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Voerman:2021:PNS

[1428] Sofie E. Voerman, Paul E. Gribben, and Tim M. Glasby. Positive and negative species interactions shape recruitment patterns of a range expanding native alga. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Park:2021:GPS

[1429] Sang Rul Park, Kyeonglim Moon, Seung Hyeon Kim, and Kun-Seop Lee. Growth and photoacclimation strategies of three Zostera species along a vertical gradient: Implications for seagrass zonation patterns. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Friedland:2021:ENM

[1430] René Friedland, Diego Macias, Gianpiero Cossarini, Ute Daewel, Claude Estournel, Elisa Garcia-Gorriz, Bruna Grizzetti, Marilaure Grégoire, Bo Gustafson, Sofia Kalaroni, Onur Kerimoglu, Paolo Lazzari, Hermann Lenhart, Gennadi Lessin, Ilja Maljutenko, Svetla Miladinova, Bärbel Müller-Karulis, Thomas Neumann, Ove Parn, Johannes Pätsch, Chiara Piroddi, Urmas Raudsepp, Corinna Schrum, Christoph Stegert, Adolf Stips, Kostas Tsiaras, Caroline Ulses, and Luc Vandenbulcke. Effects of nutrient management scenarios on marine eutrophication indicators: a pan-European, multi-model assessment in support of the Marine Strategy Framework Directive. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Tai:2021:OAA

[1431] Travis C. Tai, U. Rashid Sumaila, and William W. L. Cheung. Ocean acidification amplifies multi-stressor impacts on global marine invertebrate fisheries. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Piroddi:2021:ENM

[1432] Chiara Piroddi, Ekin Akoglu, Eider Andonegi, Jacob W. Bentley, Igor Celić, Marta Coll, Donna Dimarchopoulou, René Friedland, Kim de Mutsert, Raphael Girardin, Elisa Garcia-Gorriz, Bruna Grizzetti, P.-Y. Hernvann, Johanna J. Heymans, Bärbel Müller-Karulis, Simone Libralato, Christopher P. Lynam, Diego Macias, Svetla Miladinova, Fabien Moullec, Andreas Palialexis, Ove Parn, Natalia Serpetti, Cosimo Solidoro, Jeroen Steenbeek, Adolf Stips, Maciej T. Tomczak, Morgane Travers-Trolet, and Athanassios C. Tsikliras. Effects of nutrient management scenarios on marine food webs: a pan-European assessment in support of the marine strategy framework directive. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Barry:2021:VSA

[1433] Savanna C. Barry, A. Challen Hyman, Charles A. Jacoby, Laura K. Reynolds, Michal Kowalewski, and Thomas K. Frazer. Variation in seagrass-associated macroinvertebrate communities along the Gulf Coast of Peninsular Florida: an exploration of patterns and ecological consequences. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Macusi:2021:PVC

[1434] Edison D. Macusi, Kezia L. Camaso, Anna Barboza, and Erna S. Macusi. Perceived vulnerability and climate change impacts on small-scale fisheries in Davao Gulf, Philippines. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Krumhansl:2021:MMT

[1435] Kira A. Krumhansl, Michael Dowd, and Melisa C. Wong. Multiple metrics of temperature, light, and water motion drive gradients in eelgrass productivity and resilience. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

daSilva:2021:SCV

[1436] Alex Costa da Silva, Alexis Chaigneau, Alina N. Dossa, Gerard Eldin, Moacyr Araujo, and Arnaud Bertrand. Surface circulation and vertical structure of upper ocean variability around Fernando de Noronha Archipelago and Rocas Atoll during spring 2015 and fall 2017. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Fahlman:2021:HDM

[1437] Andreas Fahlman, Michael J. Moore, and Randall S. Wells. How do marine mammals manage and usually avoid gas emboli formation and gas embolic pathology? Critical clues from studies of wild dolphins. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Kawai:2021:LLA

[1438] Yoshimi Kawai. Low-level atmospheric responses to the sea surface temperature fronts in the Chukchi and Bering Seas. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhang:2021:RSD

[1439] Qian Zhang, Zheng Gong, Changkuan Zhang, Jessica Lacy, Bruce Jaffe, Beibei Xu, and Xindi Chen. The role of surges during periods of very shallow water on sediment transport over tidal flats. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Xavier:2021:DDP

[1440] Joana R. Xavier, David J. Rees, Raquel Pereira, Ana Colaço, Christopher K. Pham, and Francisca C. Carvalho. Diversity, distribution and phylogenetic relationships of deep-sea lithistids (Porifera, Heteroscleromorpha) of the Azores Archipelago. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Gammage:2021:SBA

[1441] Louise Carin Gammage and Astrid Jarre. Scenario-based approaches to change management in fisheries can address challenges with scale and support the implementation of an ecosystem approach to fisheries management. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Niemi:2021:BIO

[1442] Andrea Niemi, Nina Bednaršek, Christine Michel, Richard A. Feely, William Williams, Kumiko Azetsu-Scott, Wojciech Walkusz, and James D. Reist. Biological impact of ocean acidification in the Canadian Arctic: Widespread severe pteropod shell dissolution in Amundsen Gulf. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Samuels:2021:SHT

[1443] Toby Samuels, Tatiana A. Rynearson, and Sinéad Collins. Surviving heatwaves: Thermal experience predicts life and death in a Southern Ocean diatom. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Currie:2021:IVH

[1444] Jens J. Currie, Jessica A. McCordic, Grace L. Olson, Abigail F. Machernis, and Stephanie H. Stack. The impact of vessels on humpback whale behavior: The benefit of added whale watching guidelines. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bentley:2021:RFA

[1445] Jacob W. Bentley, Mathieu G. Lundy, Daniel Howell, Steven E. Beggs, Alida Bundy, Francisco de Castro, Clive J. Fox, Johanna J. Heymans, Christopher P. Lynam, Debbi Pedreschi, Pia Schuchert, Natalia Serpetti, Johnny Woodlock, and David G. Reid. Refining fisheries advice with stock-specific ecosystem information. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bednarsek:2021:STOa

[1446] Nina Bednaršek, Piero Calosi, Richard A. Feely, Richard Ambrose, Maria Byrne, Kit Yu Karen Chan, Sam Dupont, Jacqueline L. Padilla-Gamiño, John I. Spicer, Faycal Kessouri, Miranda Roethler, Martha Sutula, and Stephen B. Weisberg. Synthesis of thresholds of ocean acidification impacts on echinoderms. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Leurs:2021:IFN

[1447] Guido Leurs, Karin J. van der Reijden, Sidi Yahya Cheikhna Lemrabott, Iça Barry, Diosnes Manuel Nonque, Han Olff, Samuel Ledo Pontes, Aissa Regalla, and Laura L. Govers. Industrial fishing near West African Marine Protected Areas and its potential effects on mobile marine predators. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Holdsworth:2021:FCO

[1448] Amber M. Holdsworth, Li Zhai, Youyu Lu, and James R. Christian. Future changes in oceanography and biogeochemistry along the Canadian Pacific continental margin. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Jay:2021:VLF

[1449] Chadwick V. Jay, Sara J. Iverson, and Anthony S. Fischbach. Variability of lipids and fatty acids in Pacific walrus blubber. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Weaver:2021:EAD

[1450] Ann Weaver. An ethology of adaptation: Dolphins stop feeding but continue socializing in construction-degraded habitat. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Medina-Feliciano:2021:DGS

[1451] Joshua G. Medina-Feliciano, Stacy Pirro, Jose E. García-Arrarás, Vladimir Mashanov, and Joseph F. Ryan. Draft genome of the sea cucumber Holothuria glaberrima, a model for the study of regeneration. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Reiss:2021:GBE

[1452] Christian S. Reiss, Anthony M. Cossio, Jennifer Walsh, George R. Cutter, and George M. Watters. Glider-based estimates of meso-zooplankton biomass density: a fisheries case study on Antarctic krill (Euphausia superba) around the Northern Antarctic Peninsula. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Pinheiro:2021:IMS

[1453] Hudson T. Pinheiro, Jose A. C. C. Nunes, E. O. C. Coni, E. C. G. Almeida, C. L. S. Sampaio, Carlos E. L. Ferreira, Pedro M. Meirelles, M. Hostim-Silva, Pedro H. C. Pereira, Vinicius J. Giglio, J. L. Gasparini, L. A. Rocha, and Camilo M. Ferreira. An inverted management strategy for the fishery of endangered marine species. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Rabouille:2021:EDH

[1454] Christophe Rabouille, Bruno Lansard, Shannon M. Owings, Nancy N. Rabalais, Bruno Bombled, Edouard Metzger, Julien Richirt, Eryn M. Eitel, Anthony D. Boever, Jordon S. Beckler, and Martial Taillefert. Early diagenesis in the hypoxic and acidified zone of the Northern Gulf of Mexico: Is organic matter recycling in sediments disconnected from the water column? Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Sanchez-Arcilla:2021:CBC

[1455] Agustin Sanchez-Arcilla, Joanna Staneva, Luigi Cavaleri, Merete Badger, Jean Bidlot, Jacob T. Sorensen, Lars B. Hansen, Adrien Martin, Andy Saulter, Manuel Espino, Mario M. Miglietta, Marc Mestres, Davide Bonaldo, Paolo Pezzutto, Johannes Schulz-Stellenfleth, Anne Wiese, Xiaoli Larsen, Sandro Carniel, Rodolfo Bolaños, Saleh Abdalla, and Alessandro Tiesi. CMEMS-based coastal analyses: Conditioning, coupling and limits for applications. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Bart:2021:DSS

[1456] Martijn C. Bart, Meggie Hudspith, Hans Tore Rapp, Piet F. M. Verdonschot, and Jasper M. de Goeij. A deep-sea sponge loop? Sponges transfer dissolved and particulate organic carbon and nitrogen to associated fauna. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Tippenhauer:2021:SSS

[1457] Sandra Tippenhauer, Markus Janout, Manita Chouksey, Sinhue Torres-Valdes, Allison Fong, and Thorben Wulff. Substantial sub-surface chlorophyll patch sustained by vertical nutrient fluxes in Fram Strait observed with an autonomous underwater vehicle. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Gillett:2021:CCS

[1458] David J. Gillett, Lisa Gilbane, and Kenneth C. Schiff. Characterizing community structure of benthic infauna from the Continental Slope of the Southern California Bight. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Slooten:2021:DPS

[1459] Elisabeth Slooten and Stephen M. Dawson. Delays in protecting a small endangered cetacean: Lessons learned for science and management. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Li:2021:AYC

[1460] Wen-Ta Li, Lien-Siang Chou, Hue-Ying Chiou, I-Hua Chen, and Wei-Cheng Yang. Analyzing 13 years of cetacean strandings: Multiple stressors to cetaceans in Taiwanese waters and their implications for conservation and future research. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Yang:2021:ASE

[1461] Wei-Cheng Yang, Chi-Fang Chen, Yee-Chien Chuah, Chun-Ru Zhuang, I-Hua Chen, T. Aran Mooney, Jeff Stott, Myra Blanchard, I-Fan Jen, and Lien-Siang Chou. Anthropogenic sound exposure-induced stress in captive dolphins and implications for cetacean health. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Berger:2021:MGC

[1462] Alexis Berger, Patricia Blackwelder, Tamara Frank, Tracey T. Sutton, Nina M. Pruzinsky, Natalie Slayden, and Jose V. Lopez. Microscopic and genetic characterization of bacterial symbionts with bioluminescent potential in pyrosoma atlanticum. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Ashe:2021:DNA

[1463] Erin Ashe, Rob Williams, Alexandra Morton, and Philip S. Hammond. Disentangling natural and anthropogenic forms of mortality and serious injury in a poorly studied pelagic dolphin. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bonneville:2021:SSH

[1464] Claire Daisy Bonneville, Solène Derville, Jolanda A. Luksenburg, Marc Oremus, and Claire Garrigue. Social structure, habitat use and injuries of Indo–Pacific bottlenose dolphins (Tursiops aduncus) reveal isolated, coastal, and threatened communities in the South Pacific. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Mustika:2021:RAM

[1465] Putu Liza Kusuma Mustika, Rob Williams, Hanggar Prasetio Kadarisman, Andri Oktapianus Purba, I Putu Ranu Fajar Maharta, Deny Rahmadani, Elok Faiqoh, and I Made Iwan Dewantama. A rapid assessment of the marine megafauna biodiversity around South Bali, Indonesia. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Tiongson:2021:PRE

[1466] Angelico Jose C. Tiongson, Jean Asuncion Utzurrum, and Manuel Eduardo L. de la Paz. Patterns of research effort and extinction risk of marine mammals in the Philippines. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Waska:2021:MTD

[1467] Hannelore Waska, Heike Simon, Soeren Ahmerkamp, Janek Greskowiak, Janis Ahrens, Stephan L. Seibert, Kai Schwalfenberg, Oliver Zielinski, and Thorsten Dittmar. Molecular traits of dissolved organic matter in the subterranean estuary of a high-energy beach: Indications of sources and sinks. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

vanderMolen:2021:PMP

[1468] Johan van der Molen, Sonja M. van Leeuwen, Laura L. Govers, Tjisse van der Heide, and Han Olff. Potential micro-plastics dispersal and accumulation in the North Sea, with application to the MSC zoe incident. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bos:2021:SSP

[1469] Ryan P. Bos, Tracey T. Sutton, and Tamara M. Frank. State of satiation partially regulates the dynamics of vertical migration. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Allen:2021:CNN

[1470] Ann N. Allen, Matt Harvey, Lauren Harrell, Aren Jansen, Karlina P. Merkens, Carrie C. Wall, Julie Cattiau, and Erin M. Oleson. A Convolutional Neural Network for Automated Detection of Humpback Whale Song in a Diverse, Long-Term Passive Acoustic Dataset. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See erratum [2605].

Indivero:2021:SCD

[1471] Julia Indivero, Allison N. Myers-Pigg, and Nicholas D. Ward. Seasonal changes in the drivers of water physico-chemistry variability of a small freshwater tidal river. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Anstead:2021:PEA

[1472] Kristen A. Anstead, Katie Drew, David Chagaris, Amy M. Schueller, Jason E. McNamee, Andre Buchheister, Geneviève Nesslage, Jim H. Uphoff, Jr., Michael J. Wilberg, Alexei Sharov, Micah J. Dean, Jeffrey Brust, Michael Celestino, Shanna Madsen, Sarah Murray, Max Appelman, Joseph C. Ballenger, Joana Brito, Ellen Cosby, Caitlin Craig, Corrin Flora, Kurt Gottschall, Robert J. Latour, Eddie Leonard, Ray Mroch, Josh Newhard, Derek Orner, Chris Swanson, Jeff Tinsman, Edward D. Houde, Thomas J. Miller, and Howard Townsend. The path to an ecosystem approach for forage fish management: a case study of Atlantic menhaden. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Shore:2021:RFF

[1473] Amanda Shore, Jordan A. Sims, Michael Grimes, Lauren I. Howe-Kerr, Carsten G. B. Grupstra, Shawn M. Doyle, Lauren Stadler, Jason B. Sylvan, Kathryn E. F. Shamberger, Sarah W. Davies, Lory Z. Santiago-Vázquez, and Adrienne M. S. Correa. On a reef far, far away: Anthropogenic impacts following extreme storms affect sponge health and bacterial communities. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745. See corrigendum [2441].

Drew:2021:BMC

[1474] Katie Drew, Matthew Cieri, Amy M. Schueller, Andre Buchheister, David Chagaris, Geneviève Nesslage, Jason E. McNamee, and James H. Uphoff. Balancing model complexity, data requirements, and management objectives in developing ecological reference points for Atlantic menhaden. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Xu:2021:TSH

[1475] Peng Xu, Haiwei Bai, Xiaoyong Xie, Chun-Chieh Wang, Xing Huang, Xueping Wang, Mingming Zhang, Zhenyu Ye, Junhua Zhu, Wenquan Zhen, Siu Gin Cheung, Paul K. S. Shin, and Kit Yue Kwan. Tri-spine horseshoe crab aquaculture, ranching and stock enhancement: Perspectives and challenges. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bruno:2021:MMC

[1476] Daphne de Albuquerque Bruno, Isabel Q. Willmer, Lucia Helena S. de S. Pereira, Rafael C. C. Rocha, Tatiana D. Saint’Pierre, Paula Baldassin, Ana Carolina S. Scarelli, Amanda Dias Tadeu, Fábio V. Correia, Enrico M. Saggioro, Leila S. Lemos, Salvatore Siciliano, and Rachel Ann Hauser-Davis. Metal and metalloid contamination in green sea turtles (Chelonia mydas) found stranded in Southeastern Brazil. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Saunders:2021:QPF

[1477] Benjamin J. Saunders, Ronen Galaiduk, Karina Inostroza, Elisabeth M. V. Myers, Jordan S. Goetze, Mark Westera, Luke Twomey, Denise McCorry, and Euan S. Harvey. Quantifying patterns in fish assemblages and habitat use along a deep submarine canyon-valley feature using a remotely operated vehicle. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Rock:2021:IUS

[1478] Brianna M. Rock and Barnabas H. Daru. Impediments to understanding seagrasses’ response to global change. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Silva:2021:ODT

[1479] Marcus Silva, Moacyr Araujo, Fábio Geber, Carmen Medeiros, Julia Araujo, Carlos Noriega, and Alex Costa da Silva. Ocean dynamics and topographic upwelling around the Aracati Seamount–North Brazilian Chain from in situ observations and modeling results. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Costello:2021:IMB

[1480] Katie E. Costello, Sharon A. Lynch, Ruth M. O’Riordan, Rob McAllen, and Sarah C. Culloty. The importance of marine bivalves in invasive host–parasite introductions. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Chuang:2021:MGR

[1481] Po-Shun Chuang, Kota Ishikawa, and Satoshi Mitarai. Morphological and genetic recovery of coral polyps after bail-out. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Zorica:2021:DCI

[1482] Barbara Zorica, Daria Ezgeta-Balić, Olja Vidjak, Vedran Vuletin, Marija Šestanović, Igor Isajlović, Vanja Čikeš Keč, Nedo Vrgoč, and Chris Harrod. Diet composition and isotopic analysis of nine important fisheries resources in the Eastern Adriatic Sea (Mediterranean). Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Hop:2021:WSD

[1483] Haakon Hop, Anette Wold, Amelie Meyer, Allison Bailey, Maja Hatlebakk, Slawomir Kwasniewski, Peter Leopold, Piotr Kuklinski, and Janne E. Søreide. Winter-spring development of the zooplankton community below sea ice in the Arctic Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Zhulay:2021:FPB

[1484] Irina Zhulay, Bodil A. Bluhm, Paul E. Renaud, Renate Degen, and Katrin Iken. Functional pattern of benthic epifauna in the Chukchi Borderland, Arctic Deep Sea. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Masuda:2021:IIN

[1485] Takako Masuda, Ondřej Prášil, Virginia E. Villafañe, Macarena S. Valiñas, Keisuke Inomura, and E. Walter Helbling. Impact of increased nutrients and lowered pH on photosynthesis and growth of three marine phytoplankton communities from the coastal South West Atlantic (Patagonia, Argentina). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bell:2021:LHE

[1486] James B. Bell, Vladimir Laptikhovsky, David K. A. Barnes, Ramon Benedet, Amelia E. Bridges, James Glass, Warren Glass, Rodney Green, Simon A. Morley, Simon Robertson, Georgia R. Robson, Harry Stanton, Freja Azzopardi, Clive N. Trueman, Oliver Yates, and Martin A. Collins. Life history and ecology of bluenose warehou (Hyperoglyphe antarctica, Centrolophidae) in the Southern Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Milisenda:2021:IPH

[1487] Giacomo Milisenda, Germana Garofalo, Fabio Fiorentino, Francesco Colloca, Francesc Maynou, Alessandro Ligas, Claudia Musumeci, Luís Bentes, Jorge M. S. Gonçalves, Karim Erzini, Tommaso Russo, Lorenzo D’Andrea, and Sergio Vitale. Identifying persistent hot spot areas of undersized fish and crustaceans in Southern European waters: Implication for fishery management under the discard ban regulation. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Dalby:2021:CSD

[1488] Oliver Dalby, Isadora Sinha, Richard K. F. Unsworth, Len J. McKenzie, Benjamin L. Jones, and Leanne C. Cullen-Unsworth. Citizen science driven big data collection requires improved and inclusive societal engagement. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Juza:2021:MSR

[1489] Mélanie Juza and Joaquín Tintoré. Multivariate sub-regional ocean indicators in the Mediterranean Sea: From event detection to climate change estimations. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Ludwig:2021:SRC

[1490] Kim E. Ludwig, Mags Daly, Stephanie Levesque, and Simon D. Berrow. Survival rates and capture heterogeneity of bottlenose dolphins (Tursiops truncatus) in the Shannon Estuary, Ireland. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Fire:2021:URT

[1491] Spencer E. Fire, Glenn A. Miller, Edna R. Sabater, and Randall S. Wells. Utility of Red Tide (Karenia brevis) monitoring data as a predictive tool to estimate brevetoxin accumulation in live, free-ranging marine mammals. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Combes:2021:SCP

[1492] Magali Combes, Sandrine Vaz, Anthony Grehan, Telmo Morato, Sophie Arnaud-Haond, Carlos Dominguez-Carrió, Alan Fox, José Manuel González-Irusta, David Johnson, Oisín Callery, Andrew Davies, Laurence Fauconnet, Ellen Kenchington, Covadonga Orejas, J. Murray Roberts, Gerald Taranto, and Lénaick Menot. Systematic conservation planning at an ocean basin scale: Identifying a viable network of deep-sea protected areas in the North Atlantic and the Mediterranean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wurz:2021:HDW

[1493] Erik Wurz, Lindsay Beazley, Barry MacDonald, Ellen Kenchington, Hans Tore Rapp, and Ronald Osinga. The hexactinellid deep-water sponge Vazella pourtalesii (Schmidt, 1870) (Rossellidae) copes with temporarily elevated concentrations of suspended natural sediment. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Mieruch:2021:SDL

[1494] Sebastian Mieruch, Serdar Demirel, Simona Simoncelli, Reiner Schlitzer, and Steffen Seitz. SalaciaML: a deep learning approach for supporting ocean data quality control. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Wallace:2021:EMM

[1495] Ryan B. Wallace, Bradley J. Peterson, and Christopher J. Gobler. Ecosystem metabolism modulates the dynamics of hypoxia and acidification across temperate coastal habitat types. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Bourbonnais:2021:PAT

[1496] Annie Bourbonnais, Claudia Frey, Xin Sun, Laura A. Bristow, Amal Jayakumar, Nathaniel E. Ostrom, Karen L. Casciotti, and Bess B. Ward. Protocols for assessing transformation rates of nitrous oxide in the water column. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Tamayo-Leiva:2021:IEW

[1497] Javier Tamayo-Leiva, Jerónimo Cifuentes-Anticevic, Pilar Aparicio-Rizzo, José Ignacio Arroyo, Italo Masotti, and Beatriz Díez. Influence of estuarine water on the microbial community structure of Patagonian fjords. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wardiatno:2021:NMB

[1498] Yusli Wardiatno, Bambang Riyanto, Nopa Aris Iskandar, Sonja Kleinertz, Peter Funch, and Fery Kurniawan. A new marine biomaterial: The shell of mangrove horseshoe crabs, Carcinoscorpius rotundicauda (Latreille, 1802) emphasizing its physico-chemical characteristics. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Addamo:2021:GCM

[1499] Anna Maria Addamo, Serena Zaccara, Günter Försterra, Juan Höfer, Ricardo García-Jiménez, Giuseppe Crosa, and Annie Machordom. Genetic conservation management of marine resources and ecosystems of Patagonian fjords. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Troch:2021:SPB

[1500] Matthias Troch, Sebastien Bertrand, Benjamin Amann, Dawei Liu, Juan A. Placencia, and Carina B. Lange. Sediment provenance in the Baker-Martínez Fjord System (Chile, 48^{o}S) indicated by magnetic susceptibility and inorganic geochemistry. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Tomita:2021:AEG

[1501] Hiroyuki Tomita, Kunio Kutsuwada, Masahisa Kubota, and Tsutomu Hihara. Advances in the estimation of global surface net heat flux based on satellite observation: J-OFURO3 V1.1. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Marshall:2021:SCD

[1502] Matthew G. Marshall, Anne M. Kellerman, Jemma L. Wadham, Jon R. Hawkings, Giovanni Daneri, Rodrigo Torres, Helena V. Pryer, Alexander Beaton, Hong Chin Ng, Alejandra Urra, Laura F. Robinson, and Robert G. M. Spencer. Seasonal changes in dissolved organic matter composition in a Patagonian fjord affected by glacier melt inputs. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Montero:2021:FAA

[1503] Paulina Montero, Martina Coppari, Federico Betti, Giorgio Bavestrello, and Giovanni Daneri. Feeding of Aulacomya atra under different organic matter sources (autochthonous and allochthonous) in a Chilean Patagonia fjord ecosystem. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Prado:2021:SGR

[1504] Elena Prado, Javier Cristobo, Augusto Rodríguez-Basalo, Pilar Ríos, Cristina Rodríguez-Cabello, and Francisco Sánchez. In situ growth rate assessment of the hexactinellid sponge Asconema setubalense using 3D photogrammetric reconstruction. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Crawford:2021:BAC

[1505] David W. Crawford, Paulina Montero, and Giovanni Daneri. Blooms of Alexandrium catenella in coastal waters of Chilean Patagonia: Is subantarctic surface water involved? Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Rodriguez-Basalo:2021:HRS

[1506] Augusto Rodríguez-Basalo, Elena Prado, Francisco Sánchez, Pilar Ríos, María Gómez-Ballesteros, and Javier Cristobo. High resolution spatial distribution for the hexactinellid sponges Asconema setubalense and Pheronema carpenteri in the Central Cantabrian Sea. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Eveleth:2021:ASH

[1507] Rachel Eveleth, David M. Glover, Matthew C. Long, Ivan D. Lima, Alison P. Chase, and Scott C. Doney. Assessing the skill of a high-resolution marine biophysical model using geostatistical analysis of mesoscale ocean chlorophyll variability from field observations and remote sensing. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Rios:2021:IKB

[1508] Pilar Ríos, Javier Cristobo, Emily Baker, Lindsay Beazley, Timothy Culwick, and Ellen Kenchington. Increasing knowledge of biodiversity on the orphan seamount: a new species of Tedania (Tedaniopsis) Dendy, 1924. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Shackell:2021:CGI

[1509] Nancy L. Shackell, David M. Keith, and Heike K. Lotze. Challenges of gauging the impact of area-based fishery closures and OECMs: a case study using long-standing Canadian groundfish closures. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Buil:2021:DDE

[1510] Mercedes Pozo Buil, Michael G. Jacox, Jerome Fiechter, Michael A. Alexander, Steven J. Bograd, Enrique N. Curchitser, Christopher A. Edwards, Ryan R. Rykaczewski, and Charles A. Stock. A dynamically downscaled ensemble of future projections for the California Current system. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Solvang:2021:ACI

[1511] Torfinn Solvang, Emil Scott Bale, Ole Jacob Broch, Aleksander Handå, and Morten Omholt Alver. Automation concepts for industrial-scale production of seaweed. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Galan:2021:AFR

[1512] Alexander Galán, Gonzalo S. Saldías, Andrea Corredor-Acosta, Richard Muñoz, Carlos Lara, and José Luis Iriarte. Argo Float reveals biogeochemical characteristics along the freshwater gradient off Western Patagonia. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Squires:2021:CSB

[1513] Dale Squires, Rebecca Lent, Peter H. Dutton, Laurent Dagorn, and Lisa T. Ballance. Credit systems for bycatch and biodiversity conservation. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Squires:2021:MBN

[1514] Dale Squires, Lisa T. Ballance, Laurent Dagorn, Peter H. Dutton, and Rebecca Lent. Mitigating bycatch: Novel insights to multidisciplinary approaches. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Safiin:2021:POB

[1515] Nik Siti Zaimah Safiin, Saleem Mustafa, Fui Fui Ching, and Rossita Shapawi. Palm oil-based enriched diets for the rotifer, Brachionus plicatilis, improved the growth of Asian seabass (Lates calcarifer) larvae. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Hajbane:2021:CGP

[1516] Sara Hajbane, Bruna Calmanovici, Julia Reisser, Adam Jolly, Vyvyan Summers, Francesco Ferrari, Anas Ghadouani, and Charitha Pattiaratchi. Coastal garbage patches: Fronts accumulate plastic films at Ashmore Reef Marine Park (Pulau Pasir), Australia. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Garcia-Lafuente:2021:HWM

[1517] Jesús García-Lafuente, Simone Sammartino, I. Emma Huertas, Susana Flecha, Ricardo F. Sánchez-Leal, Cristina Naranjo, Irene Nadal, and María Jesús Bellanco. Hotter and weaker Mediterranean outflow as a response to basin-wide alterations. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Dvoretsky:2021:CRW

[1518] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Cucumaria in Russian waters of the Barents Sea: Biological aspects and aquaculture potential. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kyvelou:2021:FSE

[1519] Stella Sofia I. Kyvelou and Dimitrios G. Ierapetritis. Fostering spatial efficiency in the marine space, in a socially sustainable way: Lessons learnt from a soft multi-use assessment in the Mediterranean. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Keil:2021:CSZ

[1520] Katherine E. Keil, Terrie Klinger, Julie E. Keister, and Anna K. McLaskey. Comparative sensitivities of zooplankton to ocean acidification conditions in experimental and natural settings. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Marcos:2021:RES

[1521] Concepción Marcos, David Díaz, Katharina Fietz, Aitor Forcada, Amanda Ford, José Antonio García-Charton, Raquel Goñi, Philippe Lenfant, Sandra Mallol, David Mouillot, María Pérez-Marcos, Oscar Puebla, Stephanie Manel, and Angel Pérez-Ruzafa. Reviewing the ecosystem services, societal goods, and benefits of Marine Protected Areas. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Li:2021:IRM

[1522] Delei Li, Joanna Staneva, Jean-Raymond Bidlot, Sebastian Grayek, Yuchao Zhu, and Baoshu Yin. Improving regional model skills during typhoon events: a case study for Super Typhoon Lingling over the Northwest Pacific Ocean. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Guo:2021:FST

[1523] Chenying Guo, Shin ichi Ito, Michio Yoneda, Hajime Kitano, Hitoshi Kaneko, Megumi Enomoto, Tomoya Aono, Masahiro Nakamura, Takashi Kitagawa, Nicholas C. Wegner, and Emmanis Dorval. Fish specialize their metabolic performance to maximize bioenergetic efficiency in their local environment: Conspecific comparison between two stocks of Pacific chub mackerel (Scomber japonicus). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Tao:2021:IPF

[1524] Lily S. R. Tao, Yanny K. Y. Mak, Valerie C. M. Ho, Ronia C. t. Sham, Tommy T. Y. Hui, Danny C. P. Lau, and Kenneth M. Y. Leung. Improvements of population fitness and trophic status of a benthic predatory fish following a trawling ban. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kim:2021:MAI

[1525] Ah Ran Kim, Tae-Ho Yoon, Chung Il Lee, Chang-Keun Kang, and Hyun-Woo Kim. Metabarcoding analysis of ichthyoplankton in the East/Japan Sea using the novel fish-specific universal primer set. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Almond:2021:SIA

[1526] Peter M. Almond, Katrin Linse, Simon Dreutter, Susie M. Grant, Huw J. Griffiths, Rowan J. Whittle, Melanie Mackenzie, and William D. K. Reid. In-situ image analysis of habitat heterogeneity and benthic biodiversity in the Prince Gustav Channel, Eastern Antarctic Peninsula. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Banas:2021:RBB

[1527] Neil S. Banas, Eva Friis Møller, Kristin L. Laidre, Malene Simon, Ingrid H. Ellingsen, and Torkel Gissel Nielsen. Reconciling behavioural, bioenergetic, and oceanographic views of bowhead whale predation on overwintering copepods at an Arctic hotspot (Disko Bay, Greenland). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lan:2021:ECC

[1528] Kuo-Wei Lan, Yan-Lun Wu, Lu-Chi Chen, Muhamad Naimullah, and Tzu-Hsiang Lin. Effects of climate change in marine ecosystems based on the spatiotemporal age structure of top predators: a case study of bigeye tuna in the Pacific Ocean. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Izquierdo:2021:STA

[1529] Francisco Izquierdo, Iosu Paradinas, Santiago Cerviño, David Conesa, Alexandre Alonso-Fernández, Francisco Velasco, Izaskun Preciado, Antonio Punzón, Fran Saborido-Rey, and Maria Grazia Pennino. Spatiotemporal assessment of the European hake (Merluccius merluccius) recruits in the Northern Iberian Peninsula. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Liu:2021:SMA

[1530] Guangpeng Liu, Annalisa Bracco, and Alexandra Sitar. Submesoscale mixing across the mixed layer in the Gulf of Mexico. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Drummen:2021:CDM

[1531] Ingo Drummen and Gerrit Olbert. Conceptual design of a modular floating multi-purpose island. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Casadevall:2021:EMA

[1532] Margarida Casadevall, Conxi Rodríguez-Prieto, Jordi Torres, Catarina Eira, Michel Marengo, Pierre Lejeune, Roberto Merciai, and Jonathan Richir. Editorial: Marine aquaculture impacts on marine biota. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Jevne:2021:PPS

[1533] Lone S. Jevne, Maria Guttu, Anna S. Båtnes, Yngvar Olsen, and Kjell I. Reitan. Planktonic and parasitic sea lice abundance on three commercial salmon farms in Norway throughout a production cycle. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Adamczak:2021:BCC

[1534] Stephanie K. Adamczak, Rachel R. Holser, Daniel P. Costa, Elizabeth J. Berens McCabe, and Randall S. Wells. Body composition of common bottlenose dolphins in Sarasota Bay, Florida. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Lenz:2021:PDB

[1535] Elizabeth Ann Lenz, Lucy A. Bartlett, Anastasios Stathakopoulos, and Ilsa B. Kuffner. Physiological differences in bleaching response of the coral Porites astreoides along the Florida Keys Reef Tract during hightemperature stress. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:ACE

[1536] Zhiqiang Liu, Jianping Gan, Hui Wu, Jianyu Hu, Zhongya Cai, and Yongfei Deng. Advances on coastal and estuarine circulations around the Changjiang Estuary in the recent decades (2000–2020). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Kolbasov:2021:NSW

[1537] Gregory A. Kolbasov, Alexandra S. Savchenko, William A. Newman, and Benny K. K. Chan. A new species of Waginella (Crustacea: Thecostraca: Ascothoracida) parasitic on a stalked crinoid from Tasman Sea, with notes on morphology of related genera. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Silva:2021:TCS

[1538] Rodolfo Silva, Hocine Oumeraci, M. Luisa Martínez, Valeria Chávez, Debora Lithgow, Brigitta I. van Tussenbroek, Helena F. M. W. van Rijswick, and Tjeerd J. Bouma. Ten commandments for sustainable, safe, and W/Healthy sandy coasts facing global change. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bella:2021:RAC

[1539] Calogero Di Bella, Antonino Calagna, Gaetano Cammilleri, Pietro Schembri, Daniela Lo Monaco, Valentina Ciprì, Luisa Battaglia, Giuseppe Barbera, Vincenzo Ferrantelli, Saloua Sadok, Ines Tliba, Gianluigi Maria Lo Dico, and Rosalinda Allegro. Risk assessment of cadmium, lead, and mercury on human health in relation to the consumption of farmed sea bass in Italy: a meta-analytical approach. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Pillans:2021:LTA

[1540] Richard D. Pillans, Wayne Rochester, Russell C. Babcock, Damian P. Thomson, Michael D. E. Haywood, and Mathew A. Vanderklift. Longterm acoustic monitoring reveals site fidelity, reproductive migrations, and sex specific differences in habitat use and migratory timing in a large coastal shark (Negaprion acutidens). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Barcelo:2021:MSP

[1541] Andrea Barceló, Jonathan Sandoval-Castillo, Karen A. Stockin, Kerstin Bilgmann, Catherine R. M. Attard, Nikki Zanardo, Guido J. Parra, Krista Hupman, Isabella M. Reeves, Emma L. Betty, Gabriela Tezanos-Pinto, Luciano B. Beheregaray, and Luciana M. Möller. A matter of scale: Population genomic structure and connectivity of fisheries at-risk common dolphins (Delphinus delphis) from Australasia. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Zhang:2021:TSE

[1542] Yunbang Zhang, Jian Gao, Yunhai Zhang, Yuanchao Zou, and Xiaojuan Cao. Transcriptome sequencing of the endangered species elongate loach (Leptobotia elongata) from the Yangtze River: De novo transcriptome assembly, annotation, identification and validation of EST-SSR markers. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Simoes-Sousa:2021:BES

[1543] Iury T. Simoes-Sousa, Ilson Carlos A. Silveira, Amit Tandon, Glenn R. Flierl, Cesar H. A. Ribeiro, and Renato P. Martins. The Barreirinhas Eddies: Stable energetic anticyclones in the near-equatorial South Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Szymkowiak:2021:CFI

[1544] Marysia Szymkowiak. A conceptual framework for incorporating human dimensions into integrated ecosystem assessments. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Wang:2021:SIF

[1545] Dayang Wang and E. Eric Adams. Secondary intrusion formation of multiphase plumes. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Parra:2021:APB

[1546] Guido J. Parra, Kerstin Bilgmann, Katharina J. Peters, and Luciana M. Möller. Abundance and potential biological removal of common dolphins subject to fishery impacts in South Australian waters. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Han:2021:WBE

[1547] Lulu Han, Yinghui Wang, Yunping Xu, Yasong Wang, Yiling Zheng, and Jianqiang Wu. Water- and base-extractable organic matter in sediments from Lower Yangtze River–Estuary–East China Sea continuum: Insight into accumulation of organic carbon in the river-dominated margin. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Peltier:2021:WPW

[1548] Helene Peltier, Matthieu Authier, Florence Caurant, Willy Dabin, Pierre Daniel, Cecile Dars, Fabien Demaret, Eleonore Meheust, Olivier Van Canneyt, Jerome Spitz, and Vincent Ridoux. In the wrong place at the wrong time: Identifying spatiotemporal co-occurrence of bycaught common dolphins and fisheries in the Bay of Biscay (NE Atlantic) from 2010 to 2019. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745. See corrigendum [2560].

Carlen:2021:SMH

[1549] Ida Carlén, Laetitia Nunny, and Mark P. Simmonds. Out of sight, out of mind: How conservation is failing European porpoises. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Haughey:2021:DHP

[1550] Rebecca Haughey, Timothy N. Hunt, Daniella Hanf, Cecilia Passadore, Ryan Baring, and Guido J. Parra. Distribution and habitat preferences of Indo–Pacific bottlenose dolphins (Tursiops aduncus) inhabiting coastal waters with mixed levels of protection. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Mann:2021:ECM

[1551] Janet Mann, Vivienne Foroughirad, Molly H. F. McEntee, Madison L. Miketa, Taylor C. Evans, Caitlin Karniski, Ewa Krzyszczyk, Eric M. Patterson, John C. Strohman, and Megan M. Wallen. Elevated calf mortality and long-term responses of wild bottlenose dolphins to extreme climate events: Impacts of foraging specialization and provisioning. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Marlowe:2021:DCS

[1552] Celia Marlowe, Kieran Hyder, Martin D. J. Sayer, and Jan Kaiser. Divers as citizen scientists: Response time, accuracy and precision of water temperature measurement using dive computers. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Zhang:2021:MCS

[1553] Dongsheng Zhang, Yadong Zhou, Jichao Yang, Thomas Linley, Ruiyan Zhang, Bo Lu, Peng Xu, Chengcheng Shen, Shiquan Lin, Yueyun Wang, Dong Sun, Xiaogu Wang, and Chunsheng Wang. Megafaunal community structure from the abyssal to hadal zone in the Yap Trench. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Triest:2021:CLC

[1554] Ludwig Triest and Tom Van der Stocken. Coastal landform constrains dispersal in mangroves. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Graells:2021:RCU

[1555] Giorgia Graells, Nao Nakamura, Juan L. Celis-Diez, Nelson A. Lagos, Pablo A. Marquet, Patricio Pliscoff, and Stefan Gelcich. A review on coastal urban ecology: Research gaps, challenges, and needs. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Verutes:2021:MSD

[1556] Gregory M. Verutes, Sarah E. Tubbs, Nick Selmes, Darren R. Clark, Peter Walker, and Oliver Clements. Modeling seasonal distribution of Irrawaddy dolphins (Orcaella brevirostris) in a transnational important marine mammal area. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Klugel:2021:BSD

[1557] Andreas Klügel, Miriam Römer, Paul Wintersteller, Kai-Frederik Lenz, Sebastian Krastel, Heinrich Villinger, and Norbert Kaul. Bathymetric and seismic data, heat flow data, and age constraints of Le Gouic Seamount, Northeastern Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Secchi:2021:SEH

[1558] Eduardo R. Secchi, M. J. Cremer, D. Danilewicz, and J. Lailson-Brito. A synthesis of the ecology, human-related threats and conservation perspectives for the endangered Franciscana dolphin. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Fader:2021:CMD

[1559] Joseph E. Fader, Brianna W. Elliott, and Andrew J. Read. The challenges of managing depredation and bycatch of toothed whales in pelagic longline fisheries: Two U.S. case studies. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Blanken:2021:FBF

[1560] Hauke Blanken, Caterina Valeo, Charles Hannah, Usman T. Khan, and Tamás Juhász. A fuzzy-based framework for assessing uncertainty in drift prediction using observed currents and winds. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Elliser:2021:RSS

[1561] Cindy R. Elliser and Anna Hall. Return of the Salish Sea harbor porpoise, Phocoena phocoena: Knowledge gaps, current research, and what we need to do to protect their future. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hein:2021:PUC

[1562] Margaux Y. Hein, Tali Vardi, Elizabeth C. Shaver, Sylvain Pioch, Lisa Boström-Einarsson, Mohamed Ahmed, Gabriel Grimsditch, and Ian M. McLeod. Perspectives on the use of coral reef restoration as a strategy to support and improve reef ecosystem services. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Perez-Ortega:2021:DWB

[1563] Betzi Perez-Ortega, Rebecca Daw, Brennan Paradee, Emma Gimbrere, and Laura J. May-Collado. Dolphin-watching boats affect whistle frequency modulation in bottlenose dolphins. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

French-McCay:2021:VOT

[1564] Deborah P. French-McCay, Malcolm L. Spaulding, Deborah Crowley, Daniel Mendelsohn, Jeremy Fontenault, and Matthew Horn. Validation of oil trajectory and fate modeling of the Deepwater Horizon oil spill. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Lai:2021:PRB

[1565] Keng Po Lai, Nathan Tam, Simon Yuan Wang, William Ka Fai Tse, Xiao Lin, Ting Fung Chan, Yin Tong, Jianwen Zhang, Doris Wai Ting Au, Rudolf Shiu Sun Wu, and Richard Yuen Chong Kong. Proteomic response of the brain to hypoxic stress in marine medaka fish (Oryzias melastigma). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Byun:2021:OAO

[1566] Do-Seong Byun, Jin-Yong Jeong, Duk jin Kim, Sungmin Hong, Kyu-Tae Lee, and Kitack Lee. Ocean and atmospheric observations at the remote Ieodo Ocean Research Station in the Northern East China Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pomianowski:2021:NTA

[1567] Konrad Pomianowski, Artur Burzyński, and Ewa Kulczykowska. A de novo transcriptome assembly of the European flounder (Platichthys flesus): The preselection of transcripts encoding active forms of enzymes. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Graham:2021:MAS

[1568] Caroline Graham, Evgeny A. Pakhomov, and Brian P. V. Hunt. Metaanalysis of salmon trophic ecology reveals spatial and interspecies dynamics across the North Pacific Ocean. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Gordon:2021:STV

[1569] Sophie E. Gordon, Max Wingfield, Patrick F. Smallhorn-West, Siola’a Malimali, Tu’ikologahau Halafihi, David S. Schoeman, and Paul C. Southgate. Spatio-temporal variation in growth performance and condition of the winged pearl oyster Pteria penguin. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Liu:2021:SAP

[1570] Ming Liu, Youyuan Chen, Yadong Wu, Jiameng Guo, Ping Sun, and Zhiming Zhang. Synergistic action of plants and microorganism in integrated floating bed on eutrophic brackish water purification in coastal estuary areas. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bowlby:2021:BPR

[1571] Heather D. Bowlby, Hugues P. Benoît, Warren Joyce, James Sulikowski, Rui Coelho, Andrés Domingo, Enric Cortés, Fabio Hazin, David Macias, Gérard Biais, Catarina Santos, and Brooke Anderson. Beyond postrelease mortality: Inferences on recovery periods and natural mortality from electronic tagging data for discarded lamnid sharks. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

UrbanR:2021:MBG

[1572] Jorge Urbán R., Esther Jiménez-López, Héctor M. Guzmán, and Lorena Viloria-Gómora. Migratory behavior of an Eastern North Pacific gray whale from Baja California Sur to Chirikov Basin, Alaska. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Arias:2021:CAS

[1573] Adrian Arias, Brock J. Bergseth, Damian Weekers, and Chris Wilcox. Countering avoidance strategies used by fishers to avoid detection during illegal fishing. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Flikkema:2021:LIA

[1574] Maarten M. B. Flikkema, Fen-Yu (Vicky) Lin, Pernille P. J. van der Plank, Jos Koning, and Olaf Waals. Legal issues for artificial floating islands. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Gregson:2021:MOS

[1575] Benjamin H. Gregson, Boyd A. McKew, Robert D. Holland, Timothy J. Nedwed, Roger C. Prince, and Terry J. McGenity. Marine oil snow, a microbial perspective. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Steenbeek:2021:UGT

[1576] Jeroen Steenbeek, Dalai Felinto, Mike Pan, Joe Buszowski, and Villy Christensen. Using gaming technology to explore and visualize management impacts on marine ecosystems. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Zielinski:2021:AED

[1577] Tymon Zielinski, Ezio Bolzacchini, Karen Evans, Luca Ferrero, Klaudia Gregorczyk, Tomasz Kijewski, Izabela Kotynska-Zielinska, Patrycja Mrowiec, Barbara Oleszczuk, Paulina Pakszys, Ewa Piechowska, Joanna Piwowarczyk, Jan Sobieszczanski, and Marcin Wichorowski. Abundance of environmental data vs. low public interest in climate and ocean issues. Where is the missing link? Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Pearlman:2021:ESO

[1578] Jay Pearlman, Pier Luigi Buttigieg, Mark Bushnell, Claudia Delgado, Juliet Hermes, Emma Heslop, Cora Hörstmann, Kirsten Isensee, Johannes Karstensen, Arno Lambert, Ana Lara-Lopez, Frank Muller-Karger, Cristian Munoz Mas, Francoise Pearlman, Peter Pissierssens, Rachel Przeslawski, Pauline Simpson, Jordan van Stavel, and Ramasamy Venkatesan. Evolving and sustaining ocean best practices to enable interoperability in the UN Decade of Ocean Science for Sustainable Development. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Johri:2021:RDD

[1579] Shaili Johri, Isabella Livingston, Anjani Tiwari, Jitesh Solanki, Anissa Busch, Isabel Moreno, Sam R. Fellows, Michael P. Doane, and Elizabeth A. Dinsdale. Reducing data deficiencies: Preliminary elasmobranch fisheries surveys in India, identify range extensions and large proportions of female and juvenile landings. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Diviacco:2021:MCS

[1580] Paolo Diviacco, Antonio Nadali, Massimiliano Iurcev, Rodrigo Carbajales, Alessandro Busato, Alessandro Pavan, Mihai Burca, Lorenzo Grio, Massimiliano Nolich, Andrea Molinaro, and Francesca Malfatti. MaD-Crow, a citizen science infrastructure to monitor water quality in the Gulf of Trieste (North Adriatic Sea). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Turicchia:2021:RCMa

[1581] Eva Turicchia, Massimo Ponti, Gianfranco Rossi, Martina Milanese, Cristina Gioia Di Camillo, and Carlo Cerrano. The Reef Check Mediterranean Underwater Coastal Environment Monitoring Protocol. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kim:2021:LSL

[1582] Yong-Yub Kim, Bong-Gwan Kim, Kwang Young Jeong, Eunil Lee, Do-Seong Byun, and Yang-Ki Cho. Local sea-level rise caused by climate change in the Northwest Pacific marginal seas using dynamical downscaling. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Horton:2021:RSO

[1583] Tammy Horton, Leigh Marsh, Brian J. Bett, Andrew R. Gates, Daniel O. B. Jones, Noëlie M. A. Benoist, Simone Pfeifer, Erik Simon-Lledó, Jennifer M. Durden, Leen Vandepitte, and Ward Appeltans. Recommendations for the standardisation of open taxonomic nomenclature for image-based identifications. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Feyrer:2021:OPM

[1584] Laura Joan Feyrer, Madison Stewart, Jas Yeung, Colette Soulier, and Hal Whitehead. Origin and persistence of markings in a long-term photoidentification dataset reveal the threat of entanglement for endangered northern bottlenose whales (Hyperoodon ampullatus). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Mendez:2021:MMF

[1585] María M. Mendez, Juan P. Livore, Federico Márquez, and Gregorio Bigatti. Mass mortality of foundation species on rocky shores: Testing a methodology for a continental monitoring program. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Liu:2021:PSC

[1586] Zhiqiang Liu, Jianping Gan, Jianyu Hu, Hui Wu, Zhongya Cai, and Yongfei Deng. Progress of studies on circulation dynamics in the East China Sea: The Kuroshio exchanges with the shelf currents. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Shanks:2021:OEO

[1587] Alan L. Shanks. Observational evidence and open questions on the role of internal tidal waves on the concentration and transport of floating plastic debris. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Montseny:2021:AER

[1588] Maria Montseny, Cristina Linares, Marina Carreiro-Silva, Lea-Anne Henry, David Billett, Erik E. Cordes, Christopher J. Smith, Nadia Papadopoulou, Meri Bilan, Fanny Girard, Heidi L. Burdett, Ann Larsson, Susanna Strömberg, Núria Viladrich, James P. Barry, Patricia Baena, Antonio Godinho, Jordi Grinyó, Andreu Santín, Telmo Morato, Andrew K. Sweetman, Josep-Maria Gili, and Andrea Gori. Active ecological restoration of cold-water corals: Techniques, challenges, costs and future directions. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Huang:2021:SMB

[1589] Wei-Jen Huang, Wei-Jun Cai, and Xinping Hu. Seasonal mixing and biological controls of the carbonate system in a river-dominated continental shelf subject to eutrophication and hypoxia in the Northern Gulf of Mexico. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Zhu:2021:NRO

[1590] Rong Zhu, Zhishuai Qu, Qi Zhang, Sabine Filker, Thorsten Stoeck, Fengchao Li, and Xiaozhong Hu. A new record of Oxytricha granulifera granulifera Foissner and Adam, 1983 (Protozoa, Ciliophora, Oxytrichidae) from a hot spring in Iceland, with notes on its abnormal form during cultivation. Frontiers in Marine Science, 8, February 2021. ISSN 2296-

Garcia-Soto:2021:MCS

[1591] Carlos Garcia-Soto, Jan J. C. Seys, Oliver Zielinski, J. A. Busch, S. I. Luna, Jose Carlos Baez, C. Domegan, K. Dubsky, I. Kotynska-Zielinska, P. Loubat, Francesca Malfatti, G. Mannaerts, Patricia McHugh, P. Monestiez, Gro I. van der Meeren, and G. Gorsky. Marine citizen science: Current state in Europe and new technological developments. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Giesecke:2021:GHB

[1592] Ricardo Giesecke, Jacobo Martín, Andrea Piñones, Juan Höfer, Jose Garcés-Vargas, Ximena Flores-Melo, Emilio Alarcón, Xavier Durrieu de Madron, François Bourrin, and Humberto E. González. General hydrography of the Beagle Channel, a subantarctic interoceanic passage at the southern tip of South America. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Rehren:2021:SSM

[1593] Jennifer Rehren, Maria Grazia Pennino, Marta Coll, Narriman Jiddawi, and Christopher Muhando. Supporting spatial management of datapoor, small-scale fisheries with a Bayesian approach. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kearney:2021:UGS

[1594] Kelly A. Kearney, Steven J. Bograd, Elizabeth Drenkard, Fabian A. Gomez, Melissa Haltuch, Albert J. Hermann, Michael G. Jacox, Isaac C. Kaplan, Stefan Koenigstein, Jessica Y. Luo, Michelle Masi, Barbara Muhling, Mercedes Pozo Buil, and Phoebe A. Woodworth-Jefcoats. Using global-scale earth system models for regional fisheries applications. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Champion:2021:CDR

[1595] Curtis Champion, Stephanie Brodie, and Melinda A. Coleman. Climatedriven range shifts are rapid yet variable among recreationally important coastal-pelagic fishes. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Feng:2021:CEI

[1596] Yuanyuan Feng, Fei Chai, Mark L. Wells, Yan Liao, Pengfei Li, Ting Cai, Ting Zhao, Feixue Fu, and David A. Hutchins. The combined effects of increased pCO_{2} and warming on a coastal phytoplankton assemblage: From species composition to sinking rate. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Watling:2021:VME

[1597] Les Watling and Peter J. Auster. Vulnerable marine ecosystems, communities, and indicator species: Confusing concepts for conservation of seamounts. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Brasier:2021:RSO

[1598] Madeleine J. Brasier, David Barnes, Narissa Bax, Angelika Brandt, Anne B. Christianson, Andrew J. Constable, Rachel Downey, Blanca Figuerola, Huw Griffiths, Julian Gutt, Susanne Lockhart, Simon A. Morley, Alexandra L. Post, Anton Van de Putte, Hanieh Saeedi, Jonathan S. Stark, Michael Sumner, and Catherine L. Waller. Responses of Southern Ocean seafloor habitats and communities to global and local drivers of change. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Eyal:2021:MCE

[1599] Gal Eyal, Jack H. Laverick, Pim Bongaerts, Oren Levy, and John M. Pandolfi. Mesophotic coral ecosystems of the Great Barrier Reef are understudied and underexplored. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Escalas:2021:ESW

[1600] Arthur Escalas, Jean-Christophe Auguet, Amandine Avouac, Raphaël Seguin, Antoine Gradel, Lucie Borrossi, and Sébastien Villéger. Ecological specialization within a carnivorous fish family is supported by a herbivorous microbiome shaped by a combination of gut traits and specific diet. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Gibson:2021:NIS

[1601] Kesley J. Gibson, Matthew K. Streich, Tara S. Topping, and Gregory W. Stunz. New insights into the seasonal movement patterns of shortfin mako sharks in the Gulf of Mexico. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Koch:2021:STM

[1602] Hanna R. Koch, Bailey Wallace, Allyson DeMerlis, Abigail S. Clark, and Robert J. Nowicki. 3D scanning as a tool to measure growth rates of live coral microfragments used for coral reef restoration. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Pein:2021:SSB

[1603] Johannes Pein, Annika Eisele, Tina Sanders, Ute Daewel, Emil V. Stanev, Justus E. E. van Beusekom, Joanna Staneva, and Corinna Schrum. Seasonal stratification and biogeochemical turnover in the freshwater reach of a partially mixed dredged estuary. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Rey-Baquero:2021:UEW

[1604] Maria Paula Rey-Baquero, Laura Valentina Huertas-Amaya, Kerri D. Seger, Natalia Botero-Acosta, Andrea Luna-Acosta, Christina E. Perazio, John K. Boyle, Sarah Rosenthal, and Ann Carole Vallejo. Understanding effects of whale-watching vessel noise on humpback whale song in the North Pacific Coast of Colombia with propagation models of masking and acoustic data observations. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Solomonsz:2021:SED

[1605] Josh Solomonsz, Jess Melbourne-Thomas, Andrew Constable, Rowan Trebilco, Ingrid van Putten, and Lyn Goldsworthy. Stakeholder engagement in decision making and pathways of influence for Southern Ocean ecosystem services. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Kauko:2021:PEC

[1606] Hanna M. Kauko, Tore Hattermann, Thomas Ryan-Keogh, Asmita Singh, Laura de Steur, Agneta Fransson, Melissa Chierici, Tone Falkenhaug, Elvar H. Hallfredsson, Gunnar Bratbak, Tatiana Tsagaraki, Terje Berge, Qin Zhou, and Sebastien Moreau. Phenology and environmental control of phytoplankton blooms in the Kong Håkon VII Hav in the Southern Ocean. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Santos-Carvallo:2021:IWW

[1607] Macarena Santos-Carvallo, Fernanda Barilari, María José Pérez-Alvarez, Laura Gutiérrez, Guido Pavez, Héctor Araya, Cristobal Anguita, Claudia Cerda, and Maritza Sepúlveda. Impacts of whale-watching on the short-term behavior of fin whales (Balaenoptera physalus) in a Marine Protected Area in the Southeastern Pacific. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Papke:2021:DES

[1608] Erin Papke, Bailey Wallace, Sarah Hamlyn, and Robert Nowicki. Differential effects of substrate type and genet on growth of microfragments of Acropora palmata. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Thorstenson:2021:EFV

[1609] Candice A. Thorstenson and Matthias S. Ullrich. Ecological fitness of Vibrio cholerae, Vibrio parahaemolyticus, and Vibrio vulnificus in a small-scale population dynamics study. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Skomal:2021:HVM

[1610] Gregory Skomal, Heather Marshall, Benjamin Galuardi, Lisa Natanson, Camrin D. Braun, and Diego Bernal. Horizontal and vertical movement patterns and habitat use of juvenile porbeagles (Lamna nasus) in the Western North Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Tommasi:2021:CSC

[1611] Desiree Tommasi, Yvonne deReynier, Howard Townsend, Chris J. Harvey, William H. Satterthwaite, Kristin N. Marshall, Isaac C. Kaplan, Stephanie Brodie, John C. Field, Elliott L. Hazen, Stefan Koenigstein, Joshua Lindsay, Kathleen Moore, Barbara Muhling, Lisa Pfeiffer, James A. Smith, Jonathan Sweeney, Brian Wells, and Michael G. Jacox. A case study in connecting fisheries management challenges with models and analysis to support ecosystem-based management in the California Current ecosystem. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bolanos:2021:SMC

[1612] Luis M. Bolaños, Chang Jae Choi, Alexandra Z. Worden, Nicholas Baetge, Craig A. Carlson, and Stephen Giovannoni. Seasonality of the microbial community composition in the North Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Whitehouse:2021:BIF

[1613] George A. Whitehouse, Kerim Y. Aydin, Anne B. Hollowed, Kirstin K. Holsman, Wei Cheng, Amanda Faig, Alan C. Haynie, Albert J. Hermann, Kelly A. Kearney, André E. Punt, and Timothy E. Essington. Bottom– up impacts of forecasted climate change on the Eastern Bering Sea food web. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Naqvi:2021:DMS

[1614] S. Wajih A. Naqvi. Deoxygenation in marginal seas of the Indian Ocean. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Kaplan:2021:MSE

[1615] Isaac C. Kaplan, Sarah K. Gaichas, Christine C. Stawitz, Patrick D. Lynch, Kristin N. Marshall, Jonathan J. Deroba, Michelle Masi, Jon K. T. Brodziak, Kerim Y. Aydin, Kirstin Holsman, Howard Townsend, Desiree Tommasi, James A. Smith, Stefan Koenigstein, Mariska Weijerman, and Jason Link. Management strategy evaluation: Allowing the light on the hill to illuminate more than one species. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Alves:2021:AOR

[1616] José M. R. Alves, Ricardo Tomé, Rui M. A. Caldeira, and Pedro M. A. Miranda. Asymmetric ocean response to atmospheric forcing in an island wake: a 35-year high-resolution study. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Fumagalli:2021:LBM

[1617] Maddalena Fumagalli, Marta Guerra, Tom Brough, William Carome, Rochelle Constantine, James Higham, Will Rayment, Elisabeth Slooten, Karen Stockin, and Steve Dawson. Looking back to move forward: Lessons from three decades of research and management of cetacean tourism in New Zealand. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

deMutsert:2021:EEM

[1618] Kim de Mutsert, Kristy A. Lewis, Eric D. White, and Joe Buszowski. End-to-end modeling reveals species-specific effects of large-scale coastal restoration on living resources facing climate change. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Urban:2021:CWW

[1619] R. Jorge Urbán and Lorena Viloria-Gómora. Challenges of whale watching and swim with dolphins in Mexico. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Sripan:2021:MBC

[1620] Donlaporn Sripan, Alisa Wilantho, Khunnalack Khitmoh, Doonyapong Wongsawaeng, Jamal Ouazzani, Suchana Chavanich, Sissades Tongsima, and Naraporn Somboonna. Marine bacteria community in a 150-m depth Tachai Island, the Southeast Andaman Sea of Thailand. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Toro:2021:STE

[1621] Frederick Toro, Jaime Alarcón, Bárbara Toro-Barros, Gabriela Mallea, Juan Capella, Clara Umaran-Young, Paulette Abarca, Nelly Lakestani, Claudia Peña, Mario Alvarado-Rybak, Franco Cruz, Yerko Vilina, and Jorge Gibbons. Spatial and temporal effects of whale watching on a tourism-naive resident population of bottlenose dolphins (Tursiops truncatus) in the Humboldt Penguin National Reserve, Chile. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Dunlop:2021:APL

[1622] Rebecca A. Dunlop, Janelle Braithwaite, Lars O. Mortensen, and Catriona M. Harris. Assessing population-level effects of anthropogenic disturbance on a marine mammal population. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Varghese:2021:GLC

[1623] Hilary Kates Varghese, Kim Lowell, and Jennifer Miksis-Olds. Globallocal-comparison approach: Understanding marine mammal spatial behavior by applying spatial statistics and hypothesis testing to passive acoustic data. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ferguson:2021:RTD

[1624] Caroline E. Ferguson. A rising tide does not lift all boats: Intersectional analysis reveals inequitable impacts of the seafood trade in fishing communities. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Kaiser:2021:ESO

[1625] Patricia Kaiser, Wilhelm Hagen, Wilken-Jon von Appen, Barbara Niehoff, Nicole Hildebrandt, and Holger Auel. Effects of a submesoscale oceanographic filament on zooplankton dynamics in the Arctic marginal ice zone. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Mehanna:2021:LWR

[1626] Sahar F. Mehanna and Alam Eldeen Farouk. Length–weight relationship of 60 fish species from the Eastern Mediterranean Sea, Egypt (GFCM-GSA 26). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rodriguez-Salazar:2021:BSG

[1627] Julieta Rodríguez-Salazar, Antonio Loza, Katya Ornelas-Ocampo, Rosa Maria Gutierrez-Rios, and Liliana Pardo-López. Bacteria from the Southern Gulf of Mexico: Baseline, diversity, hydrocarbon-degrading potential and future applications. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Pinchuk:2021:DTT

[1628] Alexei I. Pinchuk, Sonia D. Batten, and Wesley W. Strasburger. Doliolid (Tunicata, Thaliacea) blooms in the Southeastern Gulf of Alaska as a result of the recent marine heat wave of 2014–2016. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Qu:2021:GDT

[1629] Zhishuai Qu, Hongbo Pan, Congcong Wang, Honggang Ma, Thorsten Stoeck, and Xiaozhong Hu. Gastronautidae Deroux, 1994 and Trithigmostoma Jankowski, 1967: Evolutionary links among cyrtophorian ciliates (Protista, Ciliophora, and Phyllopharyngea). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Arnes-Urgelles:2021:ECV

[1630] Camila Arnés-Urgellés, Pelayo Salinas de León, Etienne Rastoin-Laplane, Leandro Vaca-Pita, Jenifer Suárez-Moncada, and Diego Páez-Rosas. The effects of climatic variability on the feeding ecology of the scalloped hammerhead shark (Sphyrna lewini) in the tropical Eastern Pacific. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Lemonnier:2021:DPA

[1631] Hugues Lemonnier, Florence Royer, Florian Caradec, Etienne Lopez, Clarisse Hubert, Émilie Rabiller, Térence Desclaux, Jean-Michel Fernandez, and Françoise Andrieux-Loyer. Diagenetic processes in aquaculture ponds showing metal accumulation on shrimp gills. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Lewis:2021:UME

[1632] Kristy A. Lewis, Kenneth A. Rose, Kim de Mutsert, Shaye Sable, Cameron Ainsworth, Damian C. Brady, and Howard Townsend. Using multiple ecological models to inform environmental decision-making. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Briceno:2021:AWC

[1633] Yurasi Briceño, Leonardo Sánchez, Fernando Trujillo, Lorenzo von Fersen, and Sonsireé Ramírez. Aquatic wildmeat consumption of Guiana dolphins (Sotalia guianensis) in Lake Maracaibo System, Venezuela. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Medellin-Mora:2021:UCD

[1634] Johanna Medellín-Mora, Rubén Escribano, Andrea Corredor-Acosta, Pamela Hidalgo, and Wolfgang Schneider. Uncovering the composition and diversity of pelagic copepods in the oligotrophic blue water of the South Pacific Subtropical Gyre. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lawson:2021:BMS

[1635] Christopher L. Lawson, Matthew D. Taylor, James A. Smith, Nicholas L. Payne, Jayson M. Semmens, Iain M. Suthers, and Stephanie Brodie. Bioenergetic model sensitivity to diet diversity across space, time and ontogeny. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Egger:2021:RAF

[1636] Matthias Egger, Lauren Quiros, Giulia Leone, Francesco Ferrari, Christiana M. Boerger, and Michelle Tishler. Relative abundance of floating plastic debris and neuston in the Eastern North Pacific Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Schlenker:2021:RPM

[1637] Lela S. Schlenker, Robin Faillettaz, John D. Stieglitz, Chi Hin Lam, Ronald H. Hoenig, Georgina K. Cox, Rachael M. Heuer, Christina Pasparakis, Daniel D. Benetti, Claire B. Paris, and Martin Grosell. Remote predictions of mahi-mahi (Coryphaena hippurus) spawning in the open ocean using summarized accelerometry data. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Dumala:2021:TMN

[1638] Hanna Dumała, Michał Łuszczuk, Joanna Piwowarczyk, and Tymon Zieliński. Transnational municipal networks as a mechanism for marine governance toward climate change adaptation and mitigation: Between potential and practice. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Deng:2021:DFL

[1639] Zichao Deng, Shouchang Chen, Ping Zhang, Xu Zhang, Jonathan M. Adams, Qiaoqi Luo, and Xin Lin. Dynamics of free-living and attached bacterial assemblages in Skeletonema sp. diatom cultures at elevated temperatures. Frontiers in Marine Science, 8, July 2021. ISSN 2296-

Abrahams:2021:VCU

[1640] Amieroh Abrahams, Robert W. Schlegel, and Albertus J. Smit. Variation and change of upwelling dynamics detected in the world’s eastern boundary upwelling systems. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Washburn:2021:PMB

[1641] Travis W. Washburn, Lenaick Menot, Paulo Bonifácio, Ellen Pape, Magdalena Błażewicz, Guadalupe Bribiesca-Contreras, Thomas G. Dahlgren, Tomohiko Fukushima, Adrian G. Glover, Se Jong Ju, Stefanie Kaiser, Ok Hwan Yu, and Craig R. Smith. Patterns of macrofaunal biodiversity across the clarion-clipperton zone: an area targeted for seabed mining. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bue:2021:IMR

[1642] Nadia Lo Bue, Mairi M. R. Best, Davide Embriaco, Dilumie Abeysirigunawardena, Laura Beranzoli, Richard K. Dewey, Paolo Favali, Ming Feng, Martin Heesemann, Ulpu Leijala, Diarmuid Ó’Conchubhair, Martin Scherwath, Enrico Scoccimarro, and Thomas Wernberg. The importance of marine research infrastructures in capturing processes and impacts of extreme events. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Calvo-Martin:2021:RST

[1643] Elisa Calvo-Martin, Xosé Antón Álvarez-Salgado, Carlos Rocha, and J. Severino Pino Ibánhez. Reactive solute transport through two contrasting subterranean estuary exit sites in the Ría de Vigo (NW Iberian Peninsula). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Orfila:2021:ICC

[1644] Alejandro Orfila, Claudia P. Urbano-Latorre, Juan M. Sayol, Stephanie Gonzalez-Montes, Alejandro Caceres-Euse, Ismael Hernández-Carrasco, and Ángel G. Muñoz. On the impact of the Caribbean Counter Current in the Guajira Upwelling System. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Fraschetti:2021:WMI

[1645] Simonetta Fraschetti, Chris McOwen, Loredana Papa, Nadia Papadopoulou, Meri Bilan, Christoffer Boström, Pol Capdevila, Marina Carreiro-Silva, Laura Carugati, Emma Cebrian, Marta Coll, Thanos Dailianis, Roberto Danovaro, Francesco De Leo, Dario Fiorentino, Karine Gagnon, Cristina Gambi, Joaquim Garrabou, Vasilis Gerovasileiou, Bernat Hereu, Silvija Kipson, Jonne Kotta, Jean-Baptiste Ledoux, Cristina Linares, Juliette Martin, Alba Medrano, I. Montero-Serra, Telmo Morato, Antonio Pusceddu, Katerina Sevastou, Christopher J. Smith, Jana Verdura, and Giuseppe Guarnieri. Where is more important than how in coastal and marine ecosystems restoration. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Delroisse:2021:PDE

[1646] Jérôme Delroisse, Laurent Duchatelet, Patrick Flammang, and Jérôme Mallefet. Photophore distribution and enzymatic diversity within the photogenic integument of the cookie-cutter shark Isistius brasiliensis (Chondrichthyes: Dalatiidae). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Cardenas:2021:TKP

[1647] Susana Cárdenas, María Virginia Gabela-Flores, Arielle Amrein, Katie Surrey, Leah R. Gerber, and Hector M. Guzmán. Tourist knowledge, pro-conservation intentions, and tourist concern for the impacts of whalewatching in Las Perlas Archipelago, Panama. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Qiao:2021:SOP

[1648] Dongsheng Qiao, Guangning Zhi, Haizhi Liang, Dezhi Ning, Jun Yan, and Binbin Li. Scaling orchestration in physical model test of oscillating buoy wave energy converter. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Breckwoldt:2021:LWA

[1649] Annette Breckwoldt, Priscila F. M. Lopes, and Samiya A. Selim. Look who’s asking — reflections on participatory and transdisciplinary marine research approaches. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Palladino:2021:SCM

[1650] Giorgia Palladino, Elena Biagi, Simone Rampelli, Margherita Musella, Federica D’Amico, Silvia Turroni, Patrizia Brigidi, Gian Marco Luna, and Marco Candela. Seasonal changes in microbial communities associated with the jewel anemone corynactis viridis. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Li:2021:VCD

[1651] Hengde Li, Yangfan Wang, Qiang Xing, Qifan Zeng, Liang Zhao, Yaqun Zhang, Xiaoli Hu, and Zhenmin Bao. Variance component decomposition for growth traits of the bay scallop (Argopecten irradians irradians). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Keeley:2021:SIB

[1652] Nigel Keeley, Olivier Laroche, Murray Birch, and Xavier Pochon. A substrate-independent benthic sampler (SIBS) for hard and mixedbottom marine habitats: a proof-of-concept study. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ip:2021:LDP

[1653] Yuen K. Ip and Shit F. Chew. Light-dependent phenomena and related molecular mechanisms in giant clam-dinoflagellate associations: a review. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Little:2021:TDM

[1654] Mark Little, Emma E. George, Milou G. I. Arts, Jade Shivak, Sean Benler, Joel Huckeba, Zachary A. Quinlan, Vittorio Boscaro, Benjamin Mueller, Ana Georgina Cobián Güemes, Maria Isabel Rojas, Brandie White, Daniel Petras, Cynthia B. Silveira, Andreas F. Haas, Linda Wegley Kelly, Mark J. A. Vermeij, Robert A. Quinn, Patrick J. Keeling, Pieter C. Dorrestein, Forest Rohwer, and Ty N. F. Roach. Threedimensional molecular cartography of the Caribbean reef-building coral Orbicella faveolata. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Huang:2021:MEA

[1655] Ning Huang, Shuai Chen, Ming He, Qi Song, Lina Hou, Yan Zhao, Shuo Zhao, and Miao Miao. Molecular evolutionary analyses of euplotes species living in freshwater and marine habitats: a mitogenomic perspective. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Schlegel:2021:DMH

[1656] Robert W. Schlegel, Eric C. J. Oliver, and Ke Chen. Drivers of marine heatwaves in the Northwest Atlantic: The role of air–sea interaction during onset and decline. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Zhang:2021:EES

[1657] Yurong Zhang, Nianjun Xu, and Yahe Li. Effect of the extracts of Sargassum fusiforme on Red Tide microalgae in East China Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lim:2021:PAS

[1658] Kah Kheng Lim, Susann Rossbach, Nathan R. Geraldi, Ester A. Serrão, and Carlos M. Duarte. Phylogeographic analysis suggests a recent population bottleneck in the rare Red Sea Tridacna squamosina. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Hildebrand:2021:CBA

[1659] Tabea Hildebrand, Christian Fischer, Veronica Relano, Richard Gorman, and Craig Stevens. Cost-benefit analysis of a hybrid biophysical approach to wave energy extraction: Bio-oscillator. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Varela-Alvarez:2021:GVS

[1660] Elena Varela-Álvarez, João Loureiro, Patrick G. Meirmans, Mariana Castro, and Ester A. Serrão. Genomes vary in size and spatial patterns within chimeric blades of Porphyra spp. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Yang:2021:WIL

[1661] Nina Yang, Carlin A. Merkel, Yu-An Lin, Naomi M. Levine, Nicholas J. Hawco, Hai-Bo Jiang, Ping-Ping Qu, Michelle A. DeMers, Eric A. Webb, Fei-Xue Fu, and David A. Hutchins. Warming iron-limited oceans enhance nitrogen fixation and drive biogeographic specialization of the globally important cyanobacterium crocosphaera. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Leon-Munoz:2021:CLC

[1662] Jorge León-Muñoz, Rodrigo Aguayo, Rafael Marcé, Núria Catalán, Stefan Woelfl, Jorge Nimptsch, Ivan Arismendi, Camila Contreras, Doris Soto, and Alejandro Miranda. Climate and land cover trends affecting freshwater inputs to a fjord in Northwestern Patagonia. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Heim:2021:EFP

[1663] Vital Heim, Félicie Dhellemmes, Matthew J. Smukall, Samuel H. Gruber, and Tristan L. Guttridge. Effects of food provisioning on the daily ration and dive site use of great hammerhead sharks, Sphyrna mokarran. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ye:2021:RBS

[1664] Leiping Ye, Andrew J. Manning, James Holyoke, Jorge A. Penaloza-Giraldo, and Tian-Jian Hsu. The role of biophysical stickiness on oilmineral flocculation and settling in seawater. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Wu:2021:ICI

[1665] Yang Wu, Zhaomin Wang, and Chengyan Liu. Impacts of changed ice– ocean stress on the North Atlantic Ocean: Role of ocean surface currents. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Kolbusz:2021:CPC

[1666] Jessica Kolbusz, Simon de Lestang, Tim Langlois, and Charitha Pattiaratchi. Changes in Panulirus cygnus settlement along Western Australia using a long time series. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Mari:2021:IEH

[1667] Lorenzo Mari, Paco Melià, Marino Gatto, and Renato Casagrandi. Identification of ecological hotspots for the seagrass Posidonia oceanica via metapopulation modeling. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Szalaj:2021:DRS

[1668] Dorota Szalaj, Alexandra Silva, Pedro Ré, and Henrique Cabral. Detecting regime shifts in the Portuguese Continental Shelf ecosystem within the last three decades. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Friedland:2021:ROP

[1669] Kevin D. Friedland, Elizabeth T. Methratta, Andrew B. Gill, Sarah K. Gaichas, Tobey H. Curtis, Evan M. Adams, Janelle L. Morano, Daniel P. Crear, M. Conor McManus, and Damian C. Brady. Resource occurrence and productivity in existing and proposed wind energy lease areas on the Northeast US shelf. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Dukhovskoy:2021:DCC

[1670] Dmitry S. Dukhovskoy, Steven L. Morey, Eric P. Chassignet, Xu Chen, Victoria J. Coles, Linlin Cui, Courtney K. Harris, Robert Hetland, Tian-Jian Hsu, Andrew J. Manning, Michael Stukel, Kristen Thyng, and Jiaze Wang. Development of the CSOMIO coupled ocean-oil-sediment-biology model. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Lauton:2021:OBI

[1671] Gustavo Lauton, Charitha Bandula Pattiaratchi, and Carlos A. D. Lentini. Observations of breaking internal tides on the Australian North West Shelf Edge. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Xu:2021:NSE

[1672] Hongzhou Xu, Jian Shen, Dongxiao Wang, Lin Luo, and Bo Hong. Nonlinearity of subtidal estuarine circulation in the Pearl River Estuary, China. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Huang:2021:MAB

[1673] Ting-Hsuan Huang, Wei-Jun Cai, Penny Vlahos, Douglas W. R. Wallace, Ernie R. Lewis, and Chen-Tung Arthur Chen. The Mid-Atlantic Bight dissolved inorganic carbon system observed in the March 1996 DOE Ocean Margins Program (OMP) — a baseline study. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Su:2021:SLP

[1674] Jian Su, Elin Andrée, Jacob W. Nielsen, Steffen M. Olsen, and Kristine S. Madsen. Sea level projections from IPCC special report on the ocean and cryosphere call for a new climate adaptation strategy in the Skagerrak– Kattegat Seas. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Cui:2021:FOP

[1675] Linlin Cui, Courtney K. Harris, and Danielle R. N. Tarpley. Formation of oil-particle-aggregates: Numerical model formulation and calibration. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ditria:2021:AVF

[1676] Ellen M. Ditria, Rod M. Connolly, Eric L. Jinks, and Sebastian Lopez-Marcano. Annotated video footage for automated identification and counting of fish in unconstrained seagrass habitats. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Rotter:2021:EMB

[1677] Ana Rotter, Michéle Barbier, Francesco Bertoni, Atle M. Bones, M. Leonor Cancela, Jens Carlsson, Maria F. Carvalho, Marta Cegłowska, Jerónimo Chirivella-Martorell, Meltem Conk Dalay, Mercedes Cueto, Thanos Dailianis, Irem Deniz, Ana R. Díaz-Marrero, Dragana Drakulovic, Arita Dubnika, Christine Edwards, Hjörleifur Einarsson, Ayşegül Erdo an, Orhan Tufan Eroldo an, David Ezra, Stefano Fazi, Richard J. FitzGerald, Laura M. Gargan, Susana P. Gaudêncio, Marija Gligora Udovič, Nadica Ivošević DeNardis, Rósa Jónsdóttir, Marija Kataržytė, Katja Klun, Jonne Kotta, Leila Ktari, Zrinka Ljubešić, Lada Lukić Bilela, Manolis Mandalakis, Alexia Massa-Gallucci, Inga Matijošytė, Hanna Mazur-Marzec, Mohamed Mehiri, Søren Laurentius Nielsen, Lucie Novoveská, Donata Overlingė, Giuseppe Perale, Praveen Ramasamy, Céline Rebours, Thorsten Reinsch, Fernando Reyes, Baruch Rinkevich, Johan Robbens, Eric Röttinger, Vita Rudovica, Jerica Sabotič, Ivo Safarik, Siret Talve, Deniz Tasdemir, Xenia Theodotou Schneider, Olivier P. Thomas, Anna Toruńska-Sitarz, Giovanna Cristina Varese, and Marlen I. Vasquez. The essentials of marine biotechnology. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Butler:2021:DSV

[1678] Jack Butler, Camille M. L. S. Pagniello, Jules S. Jaffe, P. Ed Parnell, and Ana Širović. Diel and seasonal variability in kelp forest soundscapes off the Southern California coast. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Panto:2021:CTT

[1679] Gabriella Pantó, Francesca Pasotti, Lara Macheriotou, and Ann Vanreusel. Combining traditional taxonomy and metabarcoding: Assemblage structure of nematodes in the shelf sediments of the Eastern Antarctic Peninsula. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Cummings:2021:RSB

[1680] Vonda J. Cummings, David A. Bowden, Matthew H. Pinkerton, N. Jane Halliday, and Judi E. Hewitt. Ross Sea benthic ecosystems: Macroand mega-faunal community patterns from a multi-environment survey. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Diercks:2021:RRD

[1681] Arne R. Diercks, Isabel C. Romero, Rebekka A. Larson, Patrick Schwing, Austin Harris, Samantha Bosman, Jeffrey P. Chanton, and Gregg Brooks. Resuspension, redistribution, and deposition of oil-residues to offshore depocenters after the Deepwater Horizon oil spill. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Perez-Alvarez:2021:CPP

[1682] MJosé Pérez-Alvarez, Sebastián Kraft, Nicolás I. Segovia, Carlos Olavarría, Sergio Nigenda-Morales, Jorge Urbán R., Lorena Viloria-Gómora, Frederick Archer, Rodrigo Moraga, Maritza Sepúlveda, Macarena Santos-Carvallo, Guido Pavez, and Elie Poulin. Contrasting phylogeographic patterns among Northern and Southern Hemisphere fin whale populations with new data from the Southern Pacific. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Revelard:2021:SSS

[1683] Adèle Révelard, Emma Reyes, Baptiste Mourre, Ismael Hernández-Carrasco, Anna Rubio, Pablo Lorente, Christian De Lera Fernández, Julien Mader, Enrique Álvarez-Fanjul, and Joaquín Tintoré. Sensitivity of skill score metric to validate Lagrangian simulations in coastal areas: Recommendations for search and rescue applications. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Li:2021:EST

[1684] Jingxuan Li, Rene M. Boiteau, Lydia Babcock-Adams, Marianne Acker, Zhongchang Song, Matthew R. McIlvin, and Daniel J. Repeta. Elementselective targeting of nutrient metabolites in environmental samples by inductively coupled plasma mass spectrometry and electrospray ionization mass spectrometry. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Wright:2021:LSL

[1685] Peter J. Wright, Alice Doyle, John B. Taggart, and Andrew Davie. Linking scales of life-history variation with population structure in Atlantic cod. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Pereira:2021:RCI

[1686] Renato Crespo Pereira and Ricardo da Gama Bahia. Rhodoliths: Can its importance on a large scale be promoted by a microscale and invisible phenomenon? Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Stephenson:2021:QSO

[1687] Robert L. Stephenson, Alistair J. Hobday, Edward H. Allison, Derek Armitage, Kate Brooks, Alida Bundy, Christopher Cvitanovic, Mark Dickey-Collas, Natalia de Miranda Grilli, Catalina Gomez, Astrid Jarre, Laura Kaikkonen, Rachel Kelly, Romain López, Ella-Kari Muhl, Maria Grazia Pennino, Jamie C. Tam, and Ingrid van Putten. The quilt of sustainable ocean governance: Patterns for practitioners. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Robeck:2021:CSP

[1688] Todd R. Robeck, Karen J. Steinman, Celeste B. Parry, Forrest M. Gomez, and Eric D. Jensen. Comparisons of serum progesterone and progestagen concentrations in normal and abnormal bottlenose dolphin (Tursiops truncatus) pregnancies. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Huntington:2021:CMN

[1689] Kathleen Ann Burek Huntington, Verena Ann Gill, Amanda M. Berrian, Tracey Goldstein, Pam Tuomi, Barbara A. Byrne, Kristin Worman, and Jonna Mazet. Causes of mortality of northern sea otters (Enhydra lutris kenyoni) in Alaska from 2002 to 2012. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Briscoe:2021:DTC

[1690] Dana K. Briscoe, Calandra N. Turner Tomaszewicz, Jeffrey A. Seminoff, Denise M. Parker, George H. Balazs, Jeffrey J. Polovina, Masanori Kurita, Hitoshi Okamoto, Tomomi Saito, Marc R. Rice, and Larry B. Crowder. Dynamic thermal corridor may connect endangered loggerhead sea turtles across the Pacific Ocean. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Smith:2021:CDS

[1691] James A. Smith, Desiree Tommasi, Heather Welch, Elliott L. Hazen, Jonathan Sweeney, Stephanie Brodie, Barbara Muhling, Stephen M. Stohs, and Michael G. Jacox. Comparing dynamic and static time-area closures for bycatch mitigation: a management strategy evaluation of a swordfish fishery. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Drazen:2021:RVC

[1692] Jeffrey C. Drazen, Astrid B. Leitner, Daniel O. B. Jones, and Erik Simon-Lledó. Regional variation in communities of demersal fishes and scavengers across the CCZ and Pacific Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kelly:2021:MCC

[1693] Linda Wegley Kelly, Craig E. Nelson, Lihini I. Aluwihare, Milou G. I. Arts, Pieter C. Dorrestein, Irina Koester, Shayle B. Matsuda, Daniel Petras, Zachary A. Quinlan, and Andreas F. Haas. Molecular commerce on coral reefs: Using metabolomics to reveal biochemical exchanges underlying Holobiont biology and the ecology of coastal ecosystems. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pisera:2021:LDD

[1694] Andrzej Pisera and Vasilis Gerovasileiou. Lithistid demosponges of deepwater origin in marine caves of the North-Eastern Mediterranean Sea. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Griffin:2021:NFP

[1695] Lucas P. Griffin, Grace A. Casselberry, Kristen M. Hart, Adrian Jordaan, Sarah L. Becker, Ashleigh J. Novak, Bryan M. DeAngelis, Clayton G. Pollock, Ian Lundgren, Zandy Hillis-Starr, Andy J. Danylchuk, and Gregory B. Skomal. A novel framework to predict relative habitat selection in aquatic systems: Applying machine learning and resource selection functions to acoustic telemetry data from multiple shark species. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

deMendonca:2021:CPR

[1696] Sarah N. de Mendonça and Anna Metaxas. Comparing the performance of a remotely operated vehicle, a drop camera, and a trawl in capturing deep-sea epifaunal abundance and diversity. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Kasperski:2021:ASC

[1697] Stephen Kasperski, Geret S. DePiper, Alan C. Haynie, Suzana Blake, Lisa L. Colburn, Amy Freitag, Michael Jepson, Mandy Karnauskas, Kirsten M. Leong, Douglas Lipton, Michelle Masi, Cameron Speir, Howard Townsend, and Mariska Weijerman. Assessing the state of coupled social-ecological modeling in support of ecosystem based fisheries management in the United States. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Thomson:2021:ATA

[1698] Paul G. Thomson, Richard Pillans, Fabrice R. A. Jaine, Robert G. Harcourt, Michael D. Taylor, Charitha B. Pattiaratchi, and Dianne L. McLean. Acoustic telemetry around Western Australia’s oil and gas infrastructure helps detect the presence of an elusive and endangered migratory giant. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lowman:2021:NSS

[1699] Brooke A. Lowman, Andrew W. Jones, Jeffrey P. Pessutti, Anna M. Mercer, John P. Manderson, and Benjamin Galuardi. Northern shortfin squid (Illex illecebrosus) fishery footprint on the Northeast US continental shelf. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Davidson:2021:HOE

[1700] Keith Davidson, Callum Whyte, Dmitry Aleynik, Andrew Dale, Steven Gontarek, Andrey A. Kurekin, Sharon McNeill, Peter I. Miller, Marie Porter, Rachel Saxon, and Sarah Swan. HABreports: Online early warning of harmful algal and biotoxin risk for the Scottish shellfish and finfish aquaculture industries. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Mitchell:2021:IAM

[1701] Molly Mitchell, Robert E. Isdell, Julie Herman, and Christine Tombleson. Impact assessment and management challenges of key rural human health infrastructure under sea level rise. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Williams:2021:FST

[1702] Sara D. Williams, Cory S. Walter, and Erinn M. Muller. Fine scale temporal and spatial dynamics of the Stony Coral Tissue Loss Disease outbreak within the Lower Florida Keys. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Finotto:2021:PAN

[1703] Licia Finotto, Terence Ivan Walker, and Richard David Reina. Prolonged alteration of neonate traits following maternal exposure to fishingcapture stress during late pregnancy in a chondrichthyan species. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Ivanov:2021:OWF

[1704] Evgeny Ivanov, Arthur Capet, Emil De Borger, Steven Degraer, Eric J. M. Delhez, Karline Soetaert, Jan Vanaverbeke, and Marilaure Grégoire. Offshore wind farm footprint on organic and mineral particle flux to the bottom. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Reum:2021:DJM

[1705] Jonathan C. P. Reum, Howard Townsend, Sarah Gaichas, Skyler Sagarese, Isaac C. Kaplan, and Arnaud Grüss. It’s not the destination, it’s the journey: Multispecies model ensembles for ecosystem approaches to fisheries management. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Asner:2021:AHD

[1706] Gregory P. Asner, Nicholas R. Vaughn, Shawna A. Foo, Ethan Shafron, Joseph Heckler, and Roberta E. Martin. Abiotic and human drivers of reef habitat complexity throughout the main Hawaiian Islands. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See corrigendum

Sandifer:2021:CCC

[1707] Paul A. Sandifer and Geoffrey I. Scott. Coastlines, coastal cities, and climate change: a perspective on urgent research needs in the United States. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Caruso:2021:SHT

[1708] Carlo Caruso, Kira Hughes, and Crawford Drury. Selecting heat-tolerant corals for proactive reef restoration. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Yoshida:2021:LFA

[1709] Kazuhiro Yoshida, Andreas Seger, Matthew Corkill, Petra Heil, Kristen Karsh, Andrew McMinn, and Koji Suzuki. Low Fe availability for photosynthesis of sea-ice algae: Ex situ incubation of the ice diatom Fragilariopsis cylindrus in low-Fe sea ice using an ice tank. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Borger:2021:OWF

[1710] Emil De Borger, Evgeny Ivanov, Arthur Capet, Ulrike Braeckman, Jan Vanaverbeke, Marilaure Grégoire, and Karline Soetaert. Offshore windfarm footprint of sediment organic matter mineralization processes. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Haward:2021:NSC

[1711] Marcus Haward and Bianca Haas. The need for social considerations in SDG 14. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See corrigendum [2704].

Bowden:2021:AHS

[1712] David A. Bowden, Owen F. Anderson, Ashley A. Rowden, Fabrice Stephenson, and Malcolm R. Clark. Assessing habitat suitability models for the deep sea: Is our ability to predict the distributions of seafloor fauna improving? Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Brekhman:2021:PAP

[1713] Vera Brekhman, Maya Ofek-Lalzar, Stephen D. Atkinson, Gema Alama-Bermejo, Keren Maor-Landaw, Assaf Malik, Jerri L. Bartholomew, and Tamar Lotan. Proteomic analysis of the parasitic cnidarian Ceratonova shasta (Cnidaria: Myxozoa) reveals diverse roles of actin in motility and spore formation. Frontiers in Marine Science, 8, April 2021. ISSN 2296-

Loviknes:2021:FHD

[1714] Sunniva Løviknes, Knut H. Jensen, Bjørn A. Krafft, Valantine Anthonypillai, and Leif Nøttestad. Feeding hotspots and distribution of fin and humpback whales in the Norwegian Sea from 2013 to 2018. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Fieler:2021:EDC

[1715] Reinhold Fieler, Michael Greenacre, Sanna Matsson, Luiza Neves, Silje Forbord, and Kasper Hancke. Erosion dynamics of cultivated kelp, Saccharina latissima, and implications for environmental management and carbon sequestration. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cernadas-Martin:2021:CIV

[1716] Sara Cernadas-Martín, Konstantine J. Rountos, Janet A. Nye, Michael G. Frisk, and Ellen K. Pikitch. Composition and intraspecific variability in summer flounder (Paralichthys dentatus) diets in a eutrophic estuary. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Perelman:2021:MSL

[1717] Jessica N. Perelman, Eric Firing, Jesse M. A. van der Grient, Benjamin A. Jones, and Jeffrey C. Drazen. Mesopelagic scattering layer behaviors across the clarion-clipperton zone: Implications for deep-sea mining. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Estrada-Saldivar:2021:ESC

[1718] Nuria Estrada-Saldívar, Blanca A. Quiroga-García, Esmeralda Pérez-Cervantes, Omar O. Rivera-Garibay, and Lorenzo Alvarez-Filip. Effects of the Stony Coral Tissue Loss Disease outbreak on coral communities and the benthic composition of Cozumel reefs. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Ferrer:2021:ENE

[1719] Erica M. Ferrer, Leticia M. Cavole, Simona Clausnitzer, Daniela F. Dias, Tashiana C. Osborne, Rishi Sugla, and Emma Harrison. Entering negotiations: Early-career perspectives on the UN conference of parties and the unfolding climate crisis. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Matsui:2021:DAP

[1720] Hideaki Matsui, Taku Sasaki, Toru Kobari, Viliame Waqalevu, Kazuma Kikuchi, Manabu Ishikawa, and Tomonari Kotani. DHA accumulation in the polar lipids of the euryhaline copepod Pseudodiaptomus inopinus and its transfer to Red Sea bream Pagrus major larvae. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Gazulla:2021:DGA

[1721] Carlota R. Gazulla, Paula López-Sendino, Agostinho Antunes, Didier Aurelle, Ignasi Montero-Serra, Jean-Marie Dominici, Cristina Linares, Joaquim Garrabou, and Jean-Baptiste Ledoux. Demo-genetic approach for the conservation and restoration of a habitat-forming octocoral: The case of red coral, Corallium rubrum, in the Réserve Naturelle de Scandola. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Miralles:2021:IYK

[1722] Laura Miralles, Aitor Ibabe, Mónica González, Eva García-Vázquez, and Yaisel J. Borrell. “If you know the enemy and know yourself”: Addressing the problem of biological invasions in ports through a new NIS invasion threat score, routine monitoring, and preventive action plans. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Karki:2021:MSD

[1723] Sita Karki, Ricardo Bermejo, Robert Wilkes, Michéal Mac Monagail, Eve Daly, Mark Healy, Jenny Hanafin, Alastair McKinstry, Per-Erik Mellander, Owen Fenton, and Liam Morrison. Mapping spatial distribution and biomass of intertidal Ulva blooms using machine learning and earth observation. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Pennino:2021:MLI

[1724] Maria Grazia Pennino, Stephanie Brodie, André Frainer, Priscila F. M. Lopes, Jon Lopez, Kelly Ortega-Cisneros, Samiya Selim, and Natasa Vaidianu. The missing layers: Integrating sociocultural values into marine spatial planning. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Valenzuela:2021:DTO

[1725] Jacob J. Valenzuela, Justin Ashworth, Allison Cusick, Raffaela M. Abbriano, E. Virginia Armbrust, Mark Hildebrand, Mónica V. Orellana, and Nitin S. Baliga. Diel transcriptional oscillations of a plastid antiporter reflect increased resilience of Thalassiosira pseudonana in elevated CO_{2}. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Rodemann:2021:IED

[1726] Jonathan R. Rodemann, W. Ryan James, Rolando O. Santos, Bradley T. Furman, Zachary W. Fratto, Valentina Bautista, Jan Lara Hernandez, Natasha M. Viadero, Joshua O. Linenfelser, Lulu A. Lacy, Margaret O. Hall, Christopher R. Kelble, Christopher Kavanagh, and Jennifer S. Rehage. Impact of extreme disturbances on suspended sediment in Western Florida Bay: Implications for seagrass resilience. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Mara:2021:COA

[1727] Paraskevi Mara, Virginia P. Edgcomb, Taylor R. Sehein, David Beaudoin, Chuck Martinsen, Christina Lovely, Bridget Belcher, Rebecca Cox, Meghan Curran, Claire Farnan, Peter Giannini, Sarah Lott, Kyle Paquette, Anna Pinckney, Natalie Schafer, Tonna-Marie Surgeon-Rogers, and Daniel R. Rogers. Comparison of oyster aquaculture methods and their potential to enhance microbial nitrogen removal from coastal ecosystems. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Huang:2021:MCM

[1728] Yufei Huang, Sheng Yang, Xuebing Bai, Yonghong Shi, and Qiusheng Chen. Molecular and cellular mechanisms of lipid droplet breakdown in the liver of Chinese soft-shelled turtle (Pelodiscus sinensis). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

McGoran:2021:SSS

[1729] Alexandra R. McGoran, James S. Maclaine, Paul F. Clark, and David Morritt. Synthetic and semi-synthetic microplastic ingestion by mesopelagic fishes from Tristan da Cunha and St Helena, South Atlantic. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Jayaram:2021:ART

[1730] Chiranjivi Jayaram, Rajdeep Roy, Neethu Chacko, Debadatta Swain, Ramunaidu Punnana, S. Bandyopadhyay, S. B. Choudhury, and Dibyendu Dutta. Anomalous reduction of the total suspended matter during the COVID-19 lockdown in the hooghly estuarine system. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Mallefet:2021:BLL

[1731] Jérôme Mallefet, Darren W. Stevens, and Laurent Duchatelet. Bioluminescence of the largest luminous vertebrate, the kitefin shark, Dalatias licha: First insights and comparative aspects. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Wang:2021:LRP

[1732] Zhengui Wang, Fei Chai, Huijie Xue, Xiao Hua Wang, Yinglong J. Zhang, Richard Dugdale, and Frances Wilkerson. Light regulation of phytoplankton growth in San Francisco Bay studied using a 3D sediment transport model. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Long:2021:ESP

[1733] Chao Long, Yuehuan Zhang, Yunqing Li, Jun Li, Zihua Zhou, Yanping Qin, Xingyou Li, Haitao Ma, Jinkuan Wei, Yinyin Zhou, Zohaib Noor, Lijuan Long, and Ziniu Yu. Effects of Symbiodiniaceae phylotypes in clades A–E on progeny performance of two giant clams (Tridacna squamosa and T. crocea) during early history life stages in the South China Sea. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Shimada:2021:STR

[1734] Takahiro Shimada, Carlos M. Duarte, Abdulaziz M. Al-Suwailem, Lyndsey K. Tanabe, and Mark G. Meekan. Satellite tracking reveals nesting patterns, site fidelity, and potential impacts of warming on major green turtle rookeries in the Red Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pombo-Ayora:2021:FMC

[1735] Lucía Pombo-Ayora and Jose Tavera. Are feeding modes concealing morphofunctional diversity? The case of the new world parrotfishes. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Wang:2021:CHD

[1736] Kui Wang, Wei-Jun Cai, Jianfang Chen, David Kirchman, Bin Wang, Wei Fan, and Daji Huang. Climate and human-driven variability of summer hypoxia on a large river-dominated shelf as revealed by a hypoxia index. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Bao:2021:IEE

[1737] Nanou Bao and Kunshan Gao. Interactive effects of elevated CO_{2} concentration and light on the picophytoplankton synechococcus. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Townhill:2021:CCI

[1738] Bryony L. Townhill, Elena Couce, James Bell, Stuart Reeves, and Oliver Yates. Climate change impacts on Atlantic Oceanic island tuna fisheries. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Andrews:2021:CFL

[1739] Evan J. Andrews, Sarah Wolfe, Prateep K. Nayak, and Derek Armitage. Coastal fishers livelihood behaviors and their psychosocial explanations:

Implications for fisheries governance in a changing world. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Womble:2021:HSS

[1740] Jamie N. Womble, Perry J. Williams, Robert W. McNabb, Anupma Prakash, Rudiger Gens, Benjamin S. Sedinger, and Cheyenne R. Acevedo. Harbor seals as sentinels of ice dynamics in tidewater glacier fjords. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ceccarelli:2021:HMN

[1741] Daniela M. Ceccarelli, Kate Davey, Geoffrey P. Jones, Peter T. Harris, Sione V. Matoto, Jason Raubani, and Leanne Fernandes. How to meet new global targets in the offshore realms: Biophysical guidelines for offshore networks of no-take Marine Protected Areas. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hou:2021:FFF

[1742] Gang Hou, Yanying Chen, Sijin Wang, Jinrun Wang, Weitao Chen, and Hui Zhang. Formalin-fixed fish larvae could be effectively identified by DNA barcodes: a case study on thousands of specimens in South China Sea. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Thompson:2021:SVP

[1743] Christopher D. H. Thompson, Jessica J. Meeuwig, Judith Brown, Andrew J. Richardson, Alan M. Friedlander, Peter I. Miller, and Sam B. Weber. Spatial variation in pelagic wildlife assemblages in the Ascension Island Marine Protected Area: Implications for monitoring and management. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Pan:2021:MAE

[1744] Wanni Pan, Chuanxin Qin, Tao Zuo, Gang Yu, Wentao Zhu, Hongmei Ma, and Shigai Xi. Is metagenomic analysis an effective way to analyze fish feeding habits? A case of the yellowfin sea bream Acanthopagrus latus (Houttuyn) in Daya Bay. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Blythe:2021:POG

[1745] Jessica L. Blythe, Derek Armitage, Nathan J. Bennett, Jennifer J. Silver, and Andrew M. Song. The politics of ocean governance transformations. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gooday:2021:BDA

[1746] Andrew J. Gooday, Franck Lejzerowicz, Aurélie Goineau, Maria Holzmann, Olga Kamenskaya, Hiroshi Kitazato, Swee-Cheng Lim, Jan Pawlowski, Teresa Radziejewska, Zofia Stachowska, and Brygida Wawrzyniak-Wydrowska. The biodiversity and distribution of abyssal benthic foraminifera and their possible ecological roles: a synthesis across the clarion-clipperton zone. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wear:2021:BAC

[1747] Emma K. Wear, Matthew J. Church, Beth N. Orcutt, Christine N. Shulse, Markus V. Lindh, and Craig R. Smith. Bacterial and archaeal communities in polymetallic nodules, sediments, and bottom waters of the abyssal clarion-clipperton zone: Emerging patterns and future monitoring considerations. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Sharma:2021:MMS

[1748] Shivika Sharma, Vikas Sharma, and Subhankar Chatterjee. Microplastics in the Mediterranean Sea: Sources, pollution intensity, sea health, and regulatory policies. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Deininger:2021:CRS

[1749] Anne Deininger, Angela H. Martin, Juan C. F. Pardo, Paul R. Berg, Jyotirmoy Bhardwaj, Diana Catarino, Albert Fernández-Chacón, Karen Martinez-Swatson, Kotaro Ono, Rebekah A. Oomen, Marte Sodeland, Tonje Knutsen Sørdalen, Ann-Elin Wårøy Synnes, Susanna Huneide Thorbjørnsen, and Jonas Thormar. Coastal research seen through an early career lens — a perspective on barriers to interdisciplinarity in Norway. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Castro:2021:SFB

[1750] Leonardo R. Castro, Humberto E. González, José Garcés-Vargas, and Pamela Barrientos. Separate feeding between the pelagic stage of the squat lobster Munida gregaria and the larger sized zooplankton crustacean groups in the Beagle Channel as revealed by stable isotopes. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wang:2021:CIP

[1751] Sufang Wang, Xiaopei Yu, Shunqin Zhang, Hongyu Jin, Zhongfa Chen, Zhihua Lin, and Yongbo Bao. Cu^{2+} inhibits the peroxidase and antibacterial activity of homodimer hemoglobin from blood clam Tegillarca granosa by destroying its heme pocket structure. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Selim:2021:IAP

[1752] Samiya Ahmed Selim, Marion Glaser, Francheska Ilse Tacke, Mubashshira Rahman, and Nesar Ahmed. Innovative aquaculture for the poor to adjust to environmental change in coastal Bangladesh? Barriers and options for progress. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Betti:2021:HBM

[1753] Federico Betti, Francesco Enrichetti, Giorgio Bavestrello, Andrea Costa, Alessandra Moreni, Marzia Bo, Paula Ortiz Saini, and Giovanni Daneri. Hard-bottom megabenthic communities of a Chilean fjord system: Sentinels for climate change? Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Almunia:2021:UAI

[1754] Javier Almunia, Patricia Delponti, and Fernando Rosa. Using automatic identification system (AIS) data to estimate whale watching effort. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Green:2021:SON

[1755] Hannah L. Green, Helen S. Findlay, Jamie D. Shutler, Peter E. Land, and Richard G. J. Bellerby. Satellite observations are needed to understand ocean acidification and multi-stressor impacts on fish stocks in a changing Arctic Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Villalobos:2021:DDD

[1756] Vicente I. Villalobos, Nelson Valdivia, Günter Försterra, Stacy Ballyram, Juan Pablo Espinoza, Jemma L. Wadham, Katherine Burgos-Andrade, and Vreni Häussermann. Depth-dependent diversity patterns of rocky subtidal macrobenthic communities along a temperate fjord in Northern Chilean Patagonia. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Dosell:2021:UEV

[1757] Abbie Dosell, Dan Edwards, Amanda Gregory, Alwyn Ponteen, Joseph O’Garro, Laura Cornick, and Jane M. Hawkridge. Using evidence from voluntary fisheries data collection programmes to support marine spatial planning and resolve multiple-use conflicts. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kelley:2021:ARP

[1758] Dan E. Kelley, Jaimie Harbin, and Clark Richards. argoFloats: an R package for analyzing Argo data. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Endredi:2021:WWM

[1759] Anett Endrédi, Katalin Patonai, János Podani, Simone Libralato, and Ferenc Jordán. Who is where in marine food webs? A trait-based analysis of network positions. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Estevez:2021:APJ

[1760] Rodrigo A. Estévez, Gabriel Jerez, and Stefan Gelcich. Assessing procedural justice in the administration of small-scale benthic fisheries in Chile. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bates:2021:EEL

[1761] Eileen H. Bates, Lindsay Alma, Tamas Ugrai, Alexander Gagnon, Michael Maher, Paul McElhany, and Jacqueline L. Padilla-Gamiño. Evaluation of the effect of local water chemistry on trace metal accumulation in Puget Sound shellfish shows that concentration varies with species, size, and location. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Quigley:2021:VFT

[1762] Kate M. Quigley, Magena Marzonie, Blake Ramsby, David Abrego, Grant Milton, Madeleine J. H. van Oppen, and Line K. Bay. Variability in fitness trade-offs amongst coral juveniles with mixed genetic backgrounds held in the wild. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Garrido:2021:SFP

[1763] Ignacio Garrido, Luis Miguel Pardo, Ladd E. Johnson, and Dirk Schories. Selective feeding by a predatory sea star across a depth gradient in Northern Patagonia, Chile. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Tonelli:2021:CPS

[1764] Marcos Tonelli, Camila Negrão Signori, Amanda Bendia, Juliana Neiva, Bruno Ferrero, Vivian Pellizari, and Ilana Wainer. Climate projections for the Southern Ocean reveal impacts in the marine microbial communities following increases in sea surface temperature. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Leitner:2021:TSR

[1765] Astrid B. Leitner, Jeffrey C. Drazen, and Craig R. Smith. Testing the seamount refuge hypothesis for predators and scavengers in the Western Clarion–Clipperton Zone. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Gimenez:2021:SMB

[1766] Joan Giménez, Gavin E. Arneill, Ashley Bennison, Enrico Pirotta, Hans D. Gerritsen, Thomas W. Bodey, Stuart Bearhop, Keith C. Hamer, Stephen Votier, and Mark Jessopp. Sexual mismatch between vesselassociated foraging and discard consumption in a marine top predator. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Sharma:2021:DBV

[1767] Nikunj Sharma, Daris Pazhukkunnel Simon, Aracely Maribel Diaz-Garza, Elisa Fantino, Anis Messaabi, Fatma Meddeb-Mouelhi, Hugo Germain, and Isabel Desgagné-Penix. Diatoms biotechnology: Various industrial applications for a greener tomorrow. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Hrabar:2021:PPA

[1768] Jerko Hrabar, Hrvoje Smodlaka, Somayeh Rasouli-Dogaheh, Mirela Petrić, Željka Trumbić, Lauren Palmer, Kristen Sakamaki, Tina Pavelin, and Ivona Mladineo. Phylogeny and pathology of anisakids parasitizing stranded California sea lions (Zalophus californianus) in Southern California. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Deutsch:2021:EAP

[1769] Ynon Deutsch, Lior Gur, Ilana Berman Frank, and David Ezra. Endophytes from algae, a potential source for new biologically active metabolites for disease management in aquaculture. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Kang:2021:TRF

[1770] Yoonja Kang, Matthew J. Harke, Dianna L. Berry, Jackie L. Collier, Steven W. Wilhelm, Sonya T. Dyhrman, and Christopher J. Gobler. Transcriptomic responses of four pelagophytes to nutrient (N, P) and light stress. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Chen:2021:QAC

[1771] Shuangling Chen, Adrienne J. Sutton, Chuanmin Hu, and Fei Chai. Quantifying the atmospheric CO_{2} forcing effect on surface ocean pCO_{2} in the North Pacific Subtropical Gyre in the past two decades. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Urbina-Barreto:2021:WMW

[1772] Isabel Urbina-Barreto, Rémi Garnier, Simon Elise, Romain Pinel, Pascal Dumas, Vincent Mahamadaly, Mathilde Facon, Sophie Bureau, Christophe Peignon, Jean-Pascal Quod, Eric Dutrieux, Lucie Penin, and Mehdi Adjeroud. Which method for which purpose? A comparison of line intercept transect and underwater photogrammetry methods for coral reef surveys. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Biagi:2021:IPD

[1773] Elena Biagi, Margherita Musella, Giorgia Palladino, Valeria Angelini, Sauro Pari, Chiara Roncari, Daniel Scicchitano, Simone Rampelli, Silvia Franzellitti, and Marco Candela. Impact of plastic debris on the gut microbiota of Caretta caretta from Northwestern Adriatic Sea. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

VandePutte:2021:DME

[1774] Anton P. Van de Putte, Huw J. Griffiths, Cassandra Brooks, Pip Bricher, Maxime Sweetlove, Svenja Halfter, and Ben Raymond. From data to marine ecosystem assessments of the Southern Ocean: Achievements, challenges, and lessons for the future. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Morato:2021:NAB

[1775] Telmo Morato, Christopher K. Pham, Laurence Fauconnet, Gerald H. Taranto, Giovanni Chimienti, Erik Cordes, Carlos Dominguez-Carrió, Pablo Durán Muñoz, Hronn Egilsdottir, José-Manuel González-Irusta, Anthony Grehan, Dierk Hebbeln, Lea-Anne Henry, Georgios Kazanidis, Ellen Kenchington, Lenaick Menot, Tina N. Molodtsova, Covadonga Orejas, Berta Ramiro-Sánchez, Manuela Ramos, J. Murray Roberts, Luís Rodrigues, Steve W. Ross, José L. Rueda, Mar Sacau, David Stirling, and Marina Carreiro-Silva. North Atlantic basin-scale multi-criteria assessment database to inform effective management and protection of vulnerable marine ecosystems. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Torres-Freyermuth:2021:HIS

[1776] Alec Torres-Freyermuth, Gabriela Medellín, and Paulo Salles. Human impact on the spatiotemporal evolution of beach resilience on the Northwestern Yucatan Coast. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Stegert:2021:EUR

[1777] Christoph Stegert, Hermann-Josef Lenhart, Anouk Blauw, René Friedland, Wera Leujak, and Onur Kerimoglu. Evaluating uncertainties in reconstructing the pre-eutrophic state of the North Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Gao:2021:STV

[1778] Hui Gao, Hui Zhao, Guoqi Han, and Changming Dong. Spatio-temporal variations of winter phytoplankton blooms northwest of the Luzon Island in the South China Sea. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745.

Dreujou:2021:DES

[1779] Elliot Dreujou, Nicolas Desroy, Julie Carrière, Lisa Tréau de Coeli, Christopher W. McKindsey, and Philippe Archambault. Determining the ecological status of benthic coastal communities: a case in an anthropized sub-Arctic area. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Sudo:2021:DTC

[1780] Kenji Sudo, T. E. Angela L. Quiros, Anchana Prathep, Cao Van Luong, Hsing-Juh Lin, Japar Sidik Bujang, Jillian Lean Sim Ooi, Miguel D. Fortes, Muta Harah Zakaria, Siti Maryam Yaakub, Yi Mei Tan, Xiaoping Huang, and Masahiro Nakaoka. Distribution, temporal change, and conservation status of tropical seagrass beds in Southeast Asia: 2000– 2020. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Contractor:2021:EFL

[1781] Steefan Contractor and Moninya Roughan. Efficacy of feedforward and LSTM neural networks at predicting and gap filling coastal ocean timeseries: Oxygen, nutrients, and temperature. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Dudik:2021:BBO

[1782] Olesia Dudik, Sara Amorim, Joana R. Xavier, Hans Tore Rapp, Tiago H. Silva, Ricardo A. Pires, and Rui L. Reis. Bioactivity of biosilica obtained from North Atlantic deep-sea sponges. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Derycke:2021:DMS

[1783] Sofie Derycke, Sara Maes, Laure Van den Bulcke, Joran Vanhollebeke, Jan Wittoeck, Hans Hillewaert, Bart Ampe, Annelies Haegeman, Kris Hostens, and Annelies De Backer. Detection of macrobenthos species with metabarcoding is consistent in bulk DNA but dependent on body size and sclerotization in eDNA from the ethanol preservative. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Tilot:2021:TDS

[1784] Virginie Tilot, Klaas Willaert, Bleuenn Guilloux, Wenting Chen, Clement Y. Mulalap, François Gaulme, Tamatoa Bambridge, Kimberley Peters, and Arthur Dahl. Traditional dimensions of seabed resource management in the context of deep sea mining in the Pacific: Learning from the socio-ecological interconnectivity between island communities and the ocean realm. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Romero:2021:MMB

[1785] Isabel C. Romero, Jeffrey P. Chanton, Gregg R. Brooks, Samantha Bosman, Rebekka A. Larson, Austin Harris, Patrick Schwing, and Arne Diercks. Molecular markers of biogenic and oil-derived hydrocarbons in deep-sea sediments following the Deepwater Horizon spill. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Niner:2021:BWL

[1786] Holly J. Niner and Sophia N. Wassermann. Better for whom? Leveling the injustices of international conferences by moving online. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Borja:2021:BIS

[1787] Angel Borja, Iratxe Menchaca, Joxe Mikel Garmendia, Javier Franco, Joana Larreta, Yolanda Sagarminaga, Yasmin Schembri, Rocío González, Rocío Antón, Tamara Micallef, Sarah Camilleri, Oihana Solaun, Ainhize Uriarte, and María C. Uyarra. Big insights from a small country: The added value of integrated assessment in the marine environmental status evaluation of Malta. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Maldonado:2021:MNE

[1788] Manuel Maldonado, María López-Acosta, Kathrin Busch, Beate M. Slaby, Kristina Bayer, Lindsay Beazley, Ute Hentschel, Ellen Kenchington, and Hans Tore Rapp. A microbial nitrogen engine modulated by bacteriosyncytia in hexactinellid sponges: Ecological implications for deep-sea communities. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Dias:2021:VCD

[1789] Albertina Dias, Siby Kurian, Suresh Thayapurath, and Anil K. Pratihary. Variations of colored dissolved organic matter in the Mandovi Estuary, Goa, during spring inter-monsoon: a comparison with COVID-19 outbreak imposed lockdown period. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Camacho-Munoz:2021:HVP

[1790] Dolores Camacho-Muñoz, Radisti A. Praptiwi, Linda A. Lawton, and Christine Edwards. High value phycotoxins from the dinoflagellate prorocentrum. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wu:2021:EEM

[1791] Yang Wu, Zhaomin Wang, Chengyan Liu, and Liangjun Yan. Energetics of eddy-mean flow interactions in the amery ice shelf cavity. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Jordan:2021:URB

[1792] Catherine Jordan, Caroline Cusack, Michelle C. Tomlinson, Andrew Meredith, Ryan McGeady, Rafael Salas, Clynton Gregory, and Peter L. Croot. Using the red band difference algorithm to detect and monitor a Karenia spp. bloom off the South Coast of Ireland, June 2019. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Flynn:2021:SDR

[1793] Kevin J. Flynn, Douglas C. Speirs, Michael R. Heath, and Aditee Mitra. Subtle differences in the representation of consumer dynamics have large effects in marine food web models. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Verna:2021:TEP

[1794] Danielle E. Verna, Mark S. Minton, and Gregory M. Ruiz. Trade exports predict regional ballast water discharge by ships in San Francisco Bay. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Bongaerts:2021:RGL

[1795] Pim Bongaerts, Caroline E. Dubé, Katharine E. Prata, Johanna C. Gijsbers, Michelle Achlatis, and Alejandra Hernandez-Agreda. Reefscape genomics: Leveraging advances in 3D imaging to assess fine-scale patterns of genomic variation on coral reefs. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Monfort:2021:TDS

[1796] Tiffany Monfort, Adrien Cheminée, Olivier Bianchimani, Pierre Drap, Arthur Puzenat, and Thierry Thibaut. The three-dimensional structure of Mediterranean shallow rocky reefs: Use of photogrammetry-based descriptors to assess its influence on associated teleost assemblages. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Richaume:2021:PMH

[1797] Justine Richaume, Adrien Cheminée, Pierre Drap, Patrick Bonhomme, Frederic Cadene, Bruno Ferrari, Virginie Hartmann, Noëmie Michez, and Olivier Bianchimani. 3D photogrammetry modeling highlights efficient reserve effect apparition after 5 years and stillness after 40 for red coral (Corallium rubrum) conservation in French MPAs. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Mayorga-Martinez:2021:CGM

[1798] Melissa Mayorga-Martínez, Javier Bello-Pineda, Héctor Perales-Valdivia, Horacio Pérez-España, and William D. Heyman. Characterizing geomorphology of mesophotic coral reef ecosystems in the Southwestern Gulf of Mexico: Implications for conservation and management. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Palstra:2021:ASS

[1799] Arjan P. Palstra, Pablo Arechavala-Lopez, Yuanxu Xue, and Ana Roque. Accelerometry of seabream in a sea-cage: Is acceleration a good proxy for activity? Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Song:2021:PCS

[1800] Weiwei Song, Ronghua Li, Yun Zhao, Herve Migaud, Chunlin Wang, and Michaël Bekaert. Pharaoh cuttlefish, Sepia pharaonis, genome reveals unique reflectin camouflage gene set. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Gomez-Garcia:2021:QED

[1801] Miguel de Jesús Gómez-García, María del Carmen Blázquez-Moreno, Joshua David Stewart, Vianey Leos-Barajas, Iliana Araceli Fonseca-Ponce, Aldo Alfonso Zavala-Jiménez, Karen Fuentes, and James T. Ketchum. Quantifying the effects of diver interactions on manta ray behavior at their aggregation sites. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Kohlbach:2021:WCD

[1802] Doreen Kohlbach, Katrin Schmidt, Haakon Hop, Anette Wold, Amalia Keck Al-Habahbeh, Simon T. Belt, Matthias Woll, Martin Graeve, Lukas Smik, Angus Atkinson, and Philipp Assmy. Winter carnivory and diapause counteract the reliance on ice algae by Barents Sea zooplankton. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Jacob:2021:UIB

[1803] Benjamin Jacob and Emil V. Stanev. Understanding the impact of bathymetric changes in the German Bight on coastal hydrodynamics: One step toward realistic morphodynamic modeling. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cleguer:2021:NMU

[1804] Christophe Cleguer, Natalie Kelly, Julian Tyne, Martin Wieser, David Peel, and Amanda Hodgson. A novel method for using small unoccupied aerial vehicles to survey wildlife species and model their density distribution. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Xu:2021:EGC

[1805] Kexin Xu, Yinghui Wang, Shanshan Lian, Naina Hu, Xiaomei Chen, Xiaoting Dai, Lingling Zhang, Shi Wang, Jingjie Hu, Xiaoli Hu, and Zhenmin Bao. Expansion of C1Q genes in Zhikong scallop and their expression profiling after exposure to the toxic dinoflagellates. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Heyworth:2021:GCS

[1806] Stephanie M. Heyworth, James B. Bell, Christopher M. Wade, Geórgenes Cavalcante, Nicholas Robinson, Emma Young, James Glass, and David A. Feary. Genetic connectivity of seamount populations of bluenose warehou (Hyperoglyphe antarctica). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Archer-Rand:2021:RAS

[1807] Simeon Archer-Rand, Paul Whomersley, Joey O’Connor, and Abbie Dosell. Rapid assessment of seabed habitats around Pitcairn Island in aid of activity management during the COVID-19 global pandemic. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Ames:2021:FED

[1808] Cheryl Lewis Ames, Aki H. Ohdera, Sophie M. Colston, Allen G. Collins, William K. Fitt, André C. Morandini, Jeffrey S. Erickson, and Gary J.

Vora. Fieldable environmental DNA sequencing to assess jellyfish biodiversity in nearshore waters of the Florida Keys, United States. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Jones:2021:SST

[1809] Benjamin L. Jones, Lina M. Nordlund, Richard K. F. Unsworth, Narriman S. Jiddawi, and Johan S. Eklöf. Seagrass structural traits drive fish assemblages in small-scale fisheries. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Vargas-Yanez:2021:LTC

[1810] Manuel Vargas-Yáñez, Mélanie Juza, M. Carmen García-Martínez, Francina Moya, Rosa Balbín, Enrique Ballesteros, María Muñoz, Elena Tel, Josep Pascual, Pedro Vélez-Belchí, and Jordi Salat. Long-term changes in the water mass properties in the Balearic Channels over the period 1996–2019. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Campana:2021:PNS

[1811] Sara Campana, Meggie Hudspith, David Lankes, Anna de Kluijver, Celine Demey, Jorien Schoorl, Samira Absalah, Marcel T. J. van der Meer, Benjamin Mueller, and Jasper M. de Goeij. Processing of naturally sourced macroalgal- and coral-dissolved organic matter (DOM) by high and low microbial abundance encrusting sponges. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hadj-Hammou:2021:RET

[1812] Jeneen Hadj-Hammou, David Mouillot, and Nicholas A. J. Graham. Response and effect traits of coral reef fish. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Chiodi:2021:EPA

[1813] Andrew M. Chiodi, Chidong Zhang, Edward D. Cokelet, Qiong Yang, Calvin W. Mordy, Chelle L. Gentemann, Jessica N. Cross, Noah Lawrence-Slavas, Christian Meinig, Michael Steele, Don E. Harrison, Phyllis J. Stabeno, Heather M. Tabisola, Dongxiao Zhang, Eugene F. Burger, Kevin M. O’Brien, and Muyin Wang. Exploring the Pacific Arctic seasonal ice zone with saildrone USVs. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

deOliveira:2021:CCW

[1814] Larissa Macedo Cruz de Oliveira, Aaron Lim, Luis A. Conti, and Andrew J. Wheeler. 3D classification of cold-water coral reefs: a comparison of classification techniques for 3D reconstructions of cold-water coral reefs and seabed. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Bagsit:2021:ESF

[1815] Farisal U. Bagsit, Eugene Frimpong, Rebecca G. Asch, and Harold M. Monteclaro. Effect of a seasonal fishery closure on sardine and mackerel catch in the Visayan Sea, Philippines. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Turpie:2021:ESA

[1816] Kevin R. Turpie, Steven G. Ackleson, Kristin B. Byrd, and Tiffany A. H. Moisan. Editorial: Science and applications of coastal remote sensing. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Cui:2021:LTT

[1817] Kaixuan Cui, Yi Dong, Xiaoxia Sun, Li Zhao, Haijian Du, Jia Liu, Chaofeng Wang, Chen Liang, Yicong Zhao, Si Chen, Jun Xuan, Suheng Li, Yuan Zhao, and Tian Xiao. Long-term temporal and spatial distribution of coliform bacteria in Jiaozhou Bay associated with human activities and environmental governance. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Iwasa-Arai:2021:UAT

[1818] Tammy Iwasa-Arai, Silvana G. L. Siqueira, Juliana L. Segadilha, and Fosca P. P. Leite. The unique Amphipoda and Tanaidacea (Crustacea: Peracarida) associated with the brown algae Dictyota sp. from the oceanic Trindade Island, Southwestern Atlantic, with biogeographic and phylogenetic insights. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ishikawa:2021:IOC

[1819] Naoto F. Ishikawa, Nanako O. Ogawa, Yoshito Chikaraishi, Moto omi Yamaguchi, Katsunori Fujikura, Yosuke Miyairi, Yusuke Yokoyama, Toshi Nagata, and Naohiko Ohkouchi. Influences of ocean currents on the diets of demersal fish communities in the Western North Pacific revealed by their muscle carbon and nitrogen isotopic compositions. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ramos:2021:MPL

[1820] Jhostin Ramos, Marco Boto, Juan Felipe Blanco-Libreros, and José M. Riascos. The mangrove periwinkle Littoraria angulifera (Mollusca: Littorinidae) in the Urabá Gulf (Colombian Caribbean): Finding ways in an urbanizing coast? Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Romano:2021:TVP

[1821] Filomena Romano, Katerina Symiakaki, and Paraskevi Pitta. Temporal variability of planktonic ciliates in a coastal oligotrophic environment: Mixotrophy, size classes and vertical distribution. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Pini-Fitzsimmons:2021:HRS

[1822] Joni Pini-Fitzsimmons, Nathan A. Knott, and Culum Brown. Heterarchy reveals social organization of a smooth stingray (Bathytoshia brevicaudata) population in a provisioned food context. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Colombo:2021:FLG

[1823] Stefanie M. Colombo, Mohamed Emam, Brian C. Peterson, Jennifer R. Hall, Gary Burr, Zeyu Zhang, and Matthew L. Rise. Freshwater, landlocked Grand Lake strain of Atlantic salmon (Salmo salar L.) as a potential genetic source of long chain polyunsaturated fatty acids synthesis. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Griffiths:2021:BAR

[1824] Huw J. Griffiths, Paul Anker, Katrin Linse, Jamie Maxwell, Alexandra L. Post, Craig Stevens, Slawek Tulaczyk, and James A. Smith. Breaking all the rules: The first recorded hard substrate sessile benthic community far beneath an Antarctic ice shelf. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Yamamoto-Kawai:2021:OAS

[1825] Michiyo Yamamoto-Kawai, Soichiro Ito, Haruko Kurihara, and Jota Kanda. Ocean acidification state in the highly eutrophic Tokyo Bay, Japan: Controls on seasonal and interannual variability. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Jiang:2021:RNP

[1826] Shan Jiang, Jie Jin, Ying Wu, Yixue Zhang, Yongjun Wei, Carlos Rocha, Juan Severino Pino Ibánhez, and Jing Zhang. Response of nitrate processing to bio-labile dissolved organic matter supply under variable oxygen conditions in a sandy beach seepage face. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Quigg:2021:ADD

[1827] Antonietta Quigg, Peter H. Santschi, Chen Xu, Kai Ziervogel, Manoj Kamalanathan, Wei-Chun Chin, Adrian B. Burd, Andrew Wozniak, and Patrick G. Hatcher. Aggregation and degradation of dispersants and oil by microbial exopolymers (ADDOMEx): Toward a synthesis of processes and pathways of marine oil snow formation in determining the fate of hydrocarbons. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Huang:2021:EPI

[1828] Ruiping Huang, Jiazhen Sun, Yunlan Yang, Xiaowen Jiang, Zhen Wang, Xue Song, Tifeng Wang, Di Zhang, He Li, Xiangqi Yi, Shouchang Chen, Nanou Bao, Liming Qu, Rui Zhang, Nianzhi Jiao, Yahui Gao, Bangqin Huang, Xin Lin, Guang Gao, and Kunshan Gao. Elevated pCO_{2} impedes succession of phytoplankton community from diatoms to dinoflagellates along with increased abundance of viruses and bacteria. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Plon:2021:SAW

[1829] Stephanie Plön, Shanan Atkins, Vic Cockcroft, Danielle Conry, Sasha Dines, Simon Elwen, Enrico Gennari, Keshni Gopal, Tess Gridley, Sandra Hörbst, Bridget S. James, Gwenith Penry, Meredith Thornton, O. Alejandra Vargas-Fonseca, and Els Vermeulen. Science alone won’t do it! South Africa’s endangered humpback dolphins Sousa plumbea face complex conservation challenges. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Garcia-Soto:2021:OOC

[1830] Carlos Garcia-Soto, Lijing Cheng, Levke Caesar, S. Schmidtko, Elizabeth B. Jewett, Alicia Cheripka, Ignatius Rigor, Ainhoa Caballero, Sanae Chiba, Jose Carlos Báez, Tymon Zielinski, and John Patrick Abraham. An overview of ocean climate change indicators: Sea surface temperature, ocean heat content, ocean pH, dissolved oxygen concentration, Arctic sea ice extent, thickness and volume, sea level and strength of the AMOC (Atlantic Meridional Overturning Circulation). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Gutierrez-Munoz:2021:PTC

[1831] Paula Gutiérrez-Muñoz, Alice E. M. Walters, Sarah J. Dolman, and Graham J. Pierce. Patterns and trends in cetacean occurrence revealed by Shorewatch, a land-based citizen science program in Scotland (United Kingdom). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Mwangombe:2021:CRC

[1832] Michael Gilbert Mwango’mbe, Jane Spilsbury, Steve Trott, Judith Nyunja, Nina Wambiji, Tim Collins, Inês Gomes, and Sergi Pérez-Jorge.

Cetacean research and citizen science in Kenya. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Schlappy:2021:TTC

[1833] Marie-Lise Schläppy, Lucy M. Robinson, Victoria Camilieri-Asch, and Karen Miller. Trash or treasure? Considerations for future ecological research to inform oil and gas decommissioning. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Baltar:2021:WRO

[1834] Federico Baltar, Xosé A. Alvarez-Salgado, Javier Arístegui, Ronald Benner, Dennis A. Hansell, Gerhard J. Herndl, and Christian Lønborg. What is refractory organic matter in the ocean? Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

vanderReijden:2021:CIS

[1835] Karin J. van der Reijden, Leo Koop, Sebastiaan Mestdagh, Mirjam Snellen, Peter M. J. Herman, Han Olff, and Laura L. Govers. Conservation implications of Sabellaria spinulosa reef patches in a dynamic sandy-bottom environment. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Mieszkowska:2021:IPC

[1836] Nova Mieszkowska, Michael T. Burrows, Stephen J. Hawkins, and Heather Sugden. Impacts of pervasive climate change and extreme events on rocky intertidal communities: Evidence from long-term data. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Wu:2021:SDP

[1837] Lei Wu, Jiqiu Li, Alan Warren, and Xiaofeng Lin. Species diversity of the pleurostomatid ciliate genus Amphileptus (Ciliophora, Haptoria), with notes on the taxonomy and molecular phylogeny of three species. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Eyre:2021:OSF

[1838] J. E. Jack Reeves Eyre, Xubin Zeng, and Kai Zhang. Ocean surface flux algorithm effects on earth system model energy and water cycles. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Trotta:2021:ROM

[1839] Francesco Trotta, Ivan Federico, Nadia Pinardi, Giovanni Coppini, Salvatore Causio, Eric Jansen, Doroteaciro Iovino, and Simona Masina. A relocatable ocean modeling platform for downscaling to shelf-coastal areas to support disaster risk reduction. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Nehmens:2021:EST

[1840] Melissa C. Nehmens, Rebecca M. Varney, Alexis M. Janosik, and David A. Ebert. An exploratory study of telomere length in the deepsea shark, Etmopterus granulosus. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Dauvin:2021:TES

[1841] Jean-Claude Dauvin, Leandro Sampaio, Ana Maria Rodrigues, and Victor Quintino. Taxonomy and ecology of sympatric Ampelisca species (Crustacea, Amphipoda) from the Strait of Gibraltar to the Strait of Dover, North-Eastern Atlantic. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Moraes:2021:BPS

[1842] Gabriela V. Moraes, Mary Colleen Hannon, Douglas M. M. Soares, Cassius V. Stevani, Anja Schulze, and Anderson G. Oliveira. Bioluminescence in polynoid scale worms (Annelida: Polynoidae). Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See corrigendum [2343].

Geiger:2021:OSS

[1843] Erick F. Geiger, Scott F. Heron, William J. Hernández, Jamie M. Caldwell, Kim Falinski, Tova Callender, Austin L. Greene, Gang Liu, Jacqueline L. De La Cour, Roy A. Armstrong, Megan J. Donahue, and C. Mark Eakin. Optimal spatiotemporal scales to aggregate satellite ocean color data for nearshore reefs and tropical coastal waters: Two case studies. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Torres-Irineo:2021:STD

[1844] Edgar Torres-Irineo, Silvia Salas, Jorge Iván Euán-Ávila, Leopoldo E. Palomo, Daniel R. Quijano Quiñones, Eva Coronado, and Rocío Joo. Spatio-temporal determination of small-scale vessels’ fishing grounds using a vessel monitoring system in the Southeastern Gulf of Mexico. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bernhard:2021:IMS

[1845] Joan M. Bernhard, Johannes C. Wit, Victoria R. Starczak, David J. Beaudoin, William G. Phalen, and Daniel C. McCorkle. Impacts of multiple stressors on a benthic foraminiferal community: a long-term experiment assessing response to ocean acidification, hypoxia and warming. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bojorquez-Tapia:2021:AEM

[1846] Luis A. Bojórquez-Tapia, Germán Ponce-Díaz, Daniela Pedroza-Páez, Antonio J. Díaz de León, and Francisco Arreguín-Sánchez. Application of exploratory modeling in support of transdisciplinary inquiry: Regulation of fishing bycatch of loggerhead sea turtles in Gulf of Ulloa, Mexico. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Minuti:2021:ITS

[1847] Jay J. Minuti, Charlee A. Corra, Brian S. Helmuth, and Bayden D. Russell. Increased thermal sensitivity of a tropical marine gastropod under combined CO_{2} and temperature stress. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Roelfsema:2021:WGE

[1848] Chris M. Roelfsema, Mitchell Lyons, Nicholas Murray, Eva M. Kovacs, Emma Kennedy, Kathryn Markey, Rodney Borrego-Acevedo, Alexandra Ordoñez Alvarez, Chantel Say, Paul Tudman, Meredith Roe, Jeremy Wolff, Dimosthenis Traganos, Gregory P. Asner, Brianna Bambic, Brian Free, Helen E. Fox, Zoe Lieb, and Stuart R. Phinn. Workflow for the generation of expert-derived training and validation data: a view to global scale habitat mapping. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Correia:2021:PCD

[1849] Ana M. Correia, Diana Sousa-Guedes, Ágatha Gil, Raul Valente, Massimiliano Rosso, Isabel Sousa-Pinto, Neftalí Sillero, and Graham J. Pierce. Predicting cetacean distributions in the Eastern North Atlantic to support marine management. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Qiu:2021:FFI

[1850] Guoqiang Qiu, Xiaogang Xing, Fei Chai, Xiao-Hai Yan, Zhiyu Liu, and Haili Wang. Far-field impacts of a super typhoon on upper ocean phytoplankton dynamics. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Herman:2021:MNS

[1851] Peter M. J. Herman, J. J. Simeon Moons, Jeroen W. M. Wijsman, Arjen P. Luijendijk, and Tom Ysebaert. A mega-nourishment (sand motor) affects landscape diversity of subtidal benthic fauna. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hsu:2021:CRE

[1852] Astrid J. Hsu, Sara Pruckner, Erin V. Satterthwaite, Lauren V. Weatherdon, Keiran Hadley, and Elizabeth Thanh Tam Nguyen. Challenges and recommendations for equitable use of aerial tools for mangrove research. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ganser:2021:GDM

[1853] Maximilian H. Ganser, Dominik Forster, Weiwei Liu, Xiaofeng Lin, Thorsten Stoeck, and Sabine Agatha. Genetic diversity in marine planktonic ciliates (Alveolata, Ciliophora) suggests distinct geographical patterns — data from Chinese and European coastal waters. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Shaw:2021:TDM

[1854] Rachel L. Shaw, Tobey H. Curtis, Gregory Metzger, Michael P. McCallister, Alisa Newton, G. Christopher Fischer, and Matthew J. Ajemian. Three-dimensional movements and habitat selection of young white sharks (Carcharodon carcharias) across a temperate continental shelf ecosystem. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Kusdianto:2021:MHB

[1855] Heru Kusdianto, Chitrasak Kullapanich, Matanee Palasuk, Suppakarn Jandang, Kobchai Pattaragulwanit, Jamal Ouazzani, Suchana Chavanich, Voranop Viyakarn, and Naraporn Somboonna. Microbiomes of healthy and bleached corals during a 2016 thermal bleaching event in the Upper Gulf of Thailand. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Triest:2021:AGD

[1856] Ludwig Triest, Alieza Del Socorro, Vincent Jay Gado, Analyn M. Mazo, and Tim Sierens. Avicennia genetic diversity and fine-scaled structure influenced by coastal proximity of mangrove fragments. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sanjurjo-Rivera:2021:EPP

[1857] Enrique Sanjurjo-Rivera, Sarah L. Mesnick, Sara Ávila-Forcada, Oriana Poindexter, Rebecca Lent, Vanda Felbab-Brown, Andrés M. Cisneros-Montemayor, Dale Squires, U. Rashid Sumaila, Gordon Munro, Rafael Ortiz-Rodriguez, Ramses Rodriguez, and Jade F. Sainz. An economic perspective on policies to save the vaquita: Conservation actions, wildlife trafficking, and the structure of incentives. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Pan:2021:VFS

[1858] Jun Pan, Fangping Cheng, Fei Yu, Yongqiang Shi, Fan Sun, Guangcheng Si, Chuanjie Wei, Xinyuan Diao, and Yongfang Zhao. Vertical fine-scale distribution of Calanus sinicus in the Yellow Sea cold water mass during the over-summering process. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Owusu:2021:CAH

[1859] Kwabena A. Owusu, Micaela M. Kulesz, and Agostino Merico. Communication among harvesters leads to sustainable fishing behaviour in a continuous time common pool resource experiment. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Zamelczyk:2021:DAP

[1860] Katarzyna Zamelczyk, Agneta Fransson, Melissa Chierici, Elizabeth Jones, Julie Meilland, Griselda Anglada-Ortiz, and Helene Hodal Lødemel. Distribution and abundances of planktic foraminifera and shelled pteropods during the polar night in the sea-ice covered Northern Barents Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sotillo:2021:EOC

[1861] Marcos G. Sotillo, Baptiste Mourre, Marc Mestres, Pablo Lorente, Roland Aznar, Manuel García-León, Maria Liste, Alex Santana, Manuel Espino, and Enrique Álvarez. Evaluation of the operational CMEMS and coastal downstream ocean forecasting services during the Storm Gloria (January 2020). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Hatje:2021:PSA

[1862] Vanessa Hatje, Raiza L. B. Andrade, Carina Costa de Oliveira, Andrei Polejack, and Thandiwe Gxaba. Pollutants in the South Atlantic Ocean: Sources, knowledge gaps and perspectives for the decade of ocean science. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Furtado:2021:MSS

[1863] Luciana C. Furtado, Anelize Bauermeister, Rafael de Felicio, Raquel Ortega, Francisco das Chagas L. Pinto, João Agostinho Machado-Neto, Daniela B. B. Trivella, Otilia D. L. Pessoa, Diego V. Wilke, Norberto P. Lopes, Paula C. Jimenez, and Leticia V. Costa-Lotufo. Marine Streptomyces sp. isolated from the Brazilian endemic tunicate Euherdmania sp. produces dihydroeponemycin and analogs with potent antiglioma activity. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Zheng:2021:EOA

[1864] Xinqing Zheng, Chenying Wang, Huaxia Sheng, Gaofeng Niu, Xu Dong, Lingling Yuan, and Tuo Shi. Effects of ocean acidification on carbon and nitrogen fixation in the hermatypic coral Galaxea fascicularis. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Anderson:2021:INH

[1865] James M. Anderson, Echelle S. Burns, Emily N. Meese, Thomas J. Farrugia, Brian S. Stirling, Connor F. White, Ryan K. Logan, John O’Sullivan, Chuck Winkler, and Christopher G. Lowe. Interannual nearshore habitat use of young of the year white sharks off Southern California. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Mynott:2021:DER

[1866] Frances Mynott, Jemma-Anne Lonsdale, and Tammy Stamford. Developing an ecological risk assessment to effectively manage marine resources in data-limited locations: a case study for St Helena sand extraction. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Borowicz:2021:SSW

[1867] Alex Borowicz, Heather J. Lynch, Tyler Estro, Catherine Foley, Bento Gonçalves, Katelyn B. Herman, Stephanie K. Adamczak, Ian Stirling, and Lesley Thorne. Social sensors for wildlife: Ecological opportunities in the era of camera ubiquity. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Redfern:2021:OMM

[1868] Jessica V. Redfern, Kelly A. Kryc, Lena Weiss, Brooke C. Hodge, Orfhlaith O’Brien, Scott D. Kraus, Ester Quintana-Rizzo, and Peter J. Auster. Opening a marine monument to commercial fishing compromises species protections. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Molony:2021:WCW

[1869] Brett W. Molony, Damian P. Thomson, and Ming Feng. What can we learn from the 2010/11 Western Australian marine heatwave to better understand risks from the one forecast in 2020/21? Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

vandenBergh:2021:NNG

[1870] Jarrett van den Bergh, Ved Chirayath, Alan Li, Juan L. Torres-Pérez, and Michal Segal-Rozenhaimer. NeMO-Net — gamifying 3D labeling of multi-modal reference datasets to support automated marine habitat mapping. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Mok:2021:PDA

[1871] Jin-Sook Mok, Ayeon Choi, Bomina Kim, Sung-Uk An, Won-Chan Lee, Hyung Chul Kim, Jonguk Kim, Cheolho Yoon, and Jung-Ho Hyun. Phosphorus dynamics associated with organic carbon mineralization by reduction of sulfate and iron in sediment exposed to fish farming. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Han:2021:SEI

[1872] Guoqing Han, Changming Dong, Jingsong Yang, Joel Sommeria, Alexandre Stegner, and Rui M. A. Caldeira. Strain evolution and instability of an anticyclonic eddy from a laboratory experiment. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Adewumi:2021:ENU

[1873] Ibukun Jacob Adewumi. Exploring the nexus and utilities between regional and global ocean governance architecture. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Weiand:2021:AOG

[1874] Laura Weiand, Sebastian Unger, Julien Rochette, Alexander Müller, and Barbara Neumann. Advancing ocean governance in marine regions through stakeholder dialogue processes. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Gizzi:2021:DOK

[1875] Francesca Gizzi, João Gama Monteiro, Rodrigo Silva, Susanne Schäfer, Nuno Castro, Silvia Almeida, Sahar Chebaane, Alejandro Bernal-Ibáñez, Filipe Henriques, Ignacio Gestoso, and João Canning-Clode. Disease outbreak in a keystone grazer population brings hope to the recovery of macroalgal forests in a barren dominated island. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Chatzinikolaou:2021:MPG

[1876] Eva Chatzinikolaou, Kleoniki Keklikoglou, and Panos Grigoriou. Morphological properties of gastropod shells in a warmer and more acidic future ocean using 3D micro-computed tomography. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Degger:2021:NCR

[1877] Natalie Degger, Andrew Hudson, Vladimir Mamaev, Mish Hamid, and Ivica Trumbic. Navigating the complexity of regional ocean governance through the large marine ecosystems approach. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Su:2021:MII

[1878] Danielle Su, Sarath Wijeratne, and Charitha Bandula Pattiaratchi. Monsoon influence on the island mass effect around The Maldives and Sri Lanka. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhang:2021:III

[1879] Y. Stacy Zhang, Rachel K. Gittman, Sarah E. Donaher, Stacy N. Trackenberg, T. van der Heide, and Brian R. Silliman. Inclusion of intra- and interspecific facilitation expands the theoretical framework for seagrass restoration. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Villanueva:2021:BBN

[1880] Roger Villanueva, Montserrat Coll-Lladó, Laure Bonnaud-Ponticelli, Sergio A. Carrasco, Oscar Escolar, Fernando Á. Fernández-Álvarez, Ian G. Gleadall, Jaruwat Nabhitabhata, Nicolás Ortiz, Carlos Rosas, Pilar Sánchez, Janet R. Voight, and Jim Swoger. Born with bristles: New insights on the Kölliker’s organs of octopus skin. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Potdukhe:2021:VPV

[1881] Trupti V. Potdukhe, Jane M. Caffrey, Mackenzie J. Rothfus, Carrie E. Daniel, Michael E. Swords, Barbara B. Albrecht, Wade H. Jeffrey, and Lisa A. Waidner. Viable putative Vibrio vulnificus and parahaemolyticus in the Pensacola and Perdido Bays: Water column, sediments, and invertebrate biofilms. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Lorente:2021:PHF

[1882] Pablo Lorente, Jue Lin-Ye, Manuel García-León, Emma Reyes, Maria Fernandes, Marcos Garcia Sotillo, Manuel Espino, Maria Isabel Ruiz, Vicente Gracia, Susana Perez, Roland Aznar, Andres Alonso-Martirena, and Enrique Álvarez-Fanjul. On the performance of high frequency radar in the Western Mediterranean during the record-breaking Storm Gloria. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Power:2021:RPC

[1883] Hannah E. Power, Andrew W. M. Pomeroy, Michael A. Kinsela, and Thomas P. Murray. Research priorities for coastal geoscience and engineering: a collaborative exercise in priority setting from Australia. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Zilioli:2021:ERL

[1884] Martina Zilioli, Caterina Bergami, Paola Carrara, Cristiano Fugazza, Alessandro Oggioni, Alessandra Pugnetti, and Paolo Tagliolato Acquaviva d’Aragona. Enabling the reuse of long-term marine biological observations in essential variables frameworks through a practical approach. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Alvarez-Aleman:2021:CME

[1885] Anmari Alvarez-Alemán, Eddy García Alfonso, James A. Powell, Charles A. Jacoby, James D. Austin, and Thomas K. Frazer. Causes of mortality for endangered Antillean manatees in Cuba. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Xia:2021:HCI

[1886] Meng Xia, Tom Carruthers, Richard Kindong, Libin Dai, Zhe Geng, Xiaojie Dai, and Feng Wu. How can information contribute to management? Value of Information (VOI) analysis on Indian Ocean striped marlin (Kajikia audax). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Krestoff:2021:MDE

[1887] Elizabeth S. Krestoff, James P. Creecy, Wayne D. Lord, Michelle L. Haynie, James A. Coyer, and Kate Sampson. Mitochondrial DNA evaluation and species identification of Kemp’s Ridley sea turtle (Lepidochelys kempii) bones after a 3-year exposure to submerged marine and terrestrial environments. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Million:2021:CLP

[1888] Wyatt C. Million, Sibelle O’Donnell, Erich Bartels, and Carly D. Kenkel. Colony-level 3D photogrammetry reveals that total linear extension and initial growth do not scale with complex morphological growth in the branching coral, Acropora cervicornis. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Li:2021:ASF

[1889] Yuru Li, Shuyang Ma, Caihong Fu, Yongjun Tian, Jianchao Li, Peng Sun, Zhenjiang Ye, and Yang Liu. Appraising the status of fish community structure in the Yellow Sea based on an indicator-testing framework. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

deAlfonso:2021:SGS

[1890] Marta de Alfonso, Jue Lin-Ye, José M. García-Valdecasas, Susana Pérez-Rubio, M. Yolanda Luna, Daniel Santos-Muñoz, M. Isabel Ruiz, Begoña Pérez-Gómez, and Enrique Álvarez-Fanjul. Storm Gloria: Sea state evolution based on in situ measurements and modeled data and its impact on extreme values. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Chen:2021:GOE

[1891] Xiaoyan Chen, Ge Chen, Linyao Ge, Baoxiang Huang, and Chuanchuan Cao. Global oceanic eddy identification: a deep learning method from Argo profiles and altimetry data. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Aljahdali:2021:EFC

[1892] Mohammed O. Aljahdali, Abdullahi B. Alhassan, and Zhaohui Zhang. Environmental factors causing stress in Avicennia marina mangrove in rabigh lagoon along the Red Sea: Based on a multi-approach study. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lardies:2021:GVP

[1893] Marco Antonio Lardies, Paz Caballero, Cristián Duarte, and María Josefina Poupin. Geographical variation in phenotypic plasticity of intertidal sister limpet’s species under ocean acidification scenarios. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ktari:2021:FPC

[1894] Leila Ktari, Cheima Mdallel, Brahim Aoun, Leila Chebil Ajjabi, and Saloua Sadok. Fucoxanthin and phenolic contents of six dictyotales from the Tunisian Coasts with an emphasis for a green extraction using a supercritical CO_{2} method. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Goncalves:2021:CHM

[1895] Cátia Gonçalves and Pedro M. Costa. Cephalotoxins: a hotspot for marine bioprospecting? Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Mariani:2021:CAI

[1896] Patrizio Mariani, Ralf Bachmayer, Sokol Kosta, Ermanno Pietrosemoli, Murat V. Ardelan, Douglas P. Connelly, Eric Delory, Jay S. Pearlman, George Petihakis, Fletcher Thompson, and Alessandro Crise. Collaborative automation and IoT technologies for coastal ocean observing systems. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Murano:2021:IBN

[1897] Carola Murano, Elisa Bergami, Giulia Liberatori, Anna Palumbo, and Ilaria Corsi. Interplay between nanoplastics and the immune system of the Mediterranean Sea urchin Paracentrotus lividus. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Perez-Gomez:2021:USL

[1898] Begoña Pérez-Gómez, Manuel García-León, Javier García-Valdecasas, Emanuela Clementi, César Mösso Aranda, Susana Pérez-Rubio, Simona Masina, Giovanni Coppini, Rafael Molina-Sánchez, Angel Muñoz-Cubillo, Ana García Fletcher, José Francisco Sánchez González, Agustín Sánchez-Arcilla, and Enrique Álvarez Fanjul. Understanding sea level processes during Western Mediterranean Storm Gloria. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Minhat:2021:ODB

[1899] Fatin Izzati Minhat, Suresh M. Ghandhi, Nurul Syahirah Mohd Ahzan, Norizmaira Abdul Haq, Omar Abdul Rahman Abdul Manaf, Shinazamreena Mhd Sabohi, Lee Hin Lee, Mohd Fadzil Akhir, and Maizah Mohd Abdullah. The occurrence and distribution of benthic foraminifera in tropical waters along the Strait of Malacca. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Passaro:2021:ABS

[1900] Marcello Passaro, Felix L. Müller, Julius Oelsmann, Laura Rautiainen, Denise Dettmering, Michael G. Hart-Davis, Adili Abulaitijiang, Ole B. Andersen, Jacob L. Høyer, Kristine S. Madsen, Ida Margrethe Ringgaard, Jani Särkkä, Rory Scarrott, Christian Schwatke, Florian Seitz, Laura Tuomi, Marco Restano, and Jérôme Benveniste. Absolute Baltic Sea level trends in the satellite altimetry era: a revisit. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Duffy:2021:ESC

[1901] Henry J. Duffy, Tom B. Letessier, Heather J. Koldewey, Terence P. Dawson, and Robert A. Irving. Ensuring the sustainability of coastal smallscale fisheries at Pitcairn Island (South Pacific) within a large scale notake MPA. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Maicu:2021:DUC

[1902] Francesco Maicu, Jacopo Alessandri, Nadia Pinardi, Giorgia Verri, Georg Umgiesser, Stefano Lovo, Saverio Turolla, Tiziana Paccagnella, and Andrea Valentini. Downscaling with an unstructured coastal-ocean model to the Goro Lagoon and the Po River Delta branches. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Couch:2021:CCC

[1903] Courtney S. Couch, Thomas A. Oliver, Rhonda Suka, Mia Lamirand, Mollie Asbury, Corinne Amir, Bernardo Vargas-Ángel, Morgan Winston, Brittany Huntington, Frances Lichowski, Ariel Halperin, Andrew Gray, Joao Garriques, and Jennifer Samson. Comparing coral colony surveys from in-water observations and structure-from-motion imagery shows low methodological bias. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hu:2021:CDR

[1904] Xinping Hu, Hongming Yao, Cory J. Staryk, Melissa R. McCutcheon, Michael S. Wetz, and Lily Walker. Corrigendum: Disparate responses of carbonate system in two adjacent subtropical estuaries to the influence of Hurricane Harvey — a case study. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745. See [130].

Refulio-Coronado:2021:CMS

[1905] Sonia Refulio-Coronado, Katherine Lacasse, Tracey Dalton, Austin Humphries, Suchandra Basu, Hirotsugu Uchida, and Emi Uchida. Coastal and marine socio-ecological systems: a systematic review of the literature. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Lotliker:2021:DCD

[1906] Aneesh Anandrao Lotliker, Sanjiba Kumar Baliarsingh, R. Venkat Shesu, Alakes Samanta, R. Chandrasekhar Naik, and T. M. Balakrishnan Nair. Did the coronavirus disease 2019 lockdown phase influence coastal water quality parameters off major Indian cities and river basins? Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Yeom:2021:IDC

[1907] Jisu Yeom, Nayeon Park, Raehyuk Jeong, and Wonchoel Lee. Integrative description of cryptic Tigriopus species from Korea using MALDI-TOF MS and DNA barcoding. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Schulz-Stellenfleth:2021:OPG

[1908] Johannes Schulz-Stellenfleth, Silvia Foerderreuther, Jochen Horstmann, and Joanna Staneva. Optimisation of parameters in a German Bight circulation model by 4DVAR assimilation of current and water level observations. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kiszka:2021:FEE

[1909] Jeremy J. Kiszka, Michelle Caputo, Paula Méndez-Fernandez, and Russell Fielding. Feeding ecology of elusive Caribbean killer whales inferred from Bayesian stable isotope mixing models and whalers’ ecological knowledge. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Paredes-Coral:2021:MGR

[1910] Evelyn Paredes-Coral, Melita Mokos, Ann Vanreusel, and Tim Deprez. Mapping global research on ocean literacy: Implications for science, policy, and the blue economy. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Callejas:2021:ECA

[1911] Ileana A. Callejas, Christine M. Lee, Deepak R. Mishra, Stacey L. Felgate, Claire Evans, Abel Carrias, Andria Rosado, Robert Griffin, Emil A. Cherrington, Mariam Ayad, Megha Rudresh, Benjamin P. Page, and Jennifer A. Jay. Effect of COVID-19 anthropause on water clarity in the Belize Coastal Lagoon. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See corrigendum [2633].

Sarin:2021:AIC

[1912] T. S. Sarin, V. Vinoj, D. Swain, K. Landu, and E. Suhas. Aerosol induced changes in sea surface temperature over the Bay of Bengal due to COVID-19 lockdown. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Krause:2021:DMA

[1913] Douglas J. Krause, Jefferson T. Hinke, Michael E. Goebel, and Wayne L. Perryman. Drones minimize Antarctic predator responses relative to ground survey methods: an appeal for context in policy advice. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Harvey:2021:IEU

[1914] Julio B. J. Harvey, John P. Ryan, and Yanwu Zhang. Influences of extreme upwelling on a coastal retention zone. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Gulland:2021:HWN

[1915] Frances M. D. Gulland. How to work on a non-endangered species and contribute to cetacean conservation: an example by the sarasota dolphin research program. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745.

Meeuwig:2021:PBB

[1916] Jessica Jane Meeuwig, Christopher David Harburn Thompson, John Andrew Forrest, Hanna Jabour Christ, Tom Bech Letessier, and Dirk Jason Meeuwig. Pulling back the blue curtain: a pelagic monitoring program for the blue belt. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wimart-Rousseau:2021:SIV

[1917] Cathy Wimart-Rousseau, Thibaut Wagener, Marta Álvarez, Thierry Moutin, Marine Fourrier, Laurent Coppola, Laure Niclas-Chirurgien, Patrick Raimbault, Fabrizio D’Ortenzio, Xavier Durrieu de Madron, Vincent Taillandier, Franck Dumas, Pascal Conan, Mireille Pujo-Pay, and Dominique Lefèvre. Seasonal and interannual variability of the CO_{2} system in the Eastern Mediterranean Sea: a case study in the North Western Levantine Basin. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Thorsen:2021:INL

[1918] Sandra W. Thorsen, Marianne Holmer, Cintia O. Quintana, Thomas Valdemarsen, and Erik Kristensen. Internal nutrient loading controls macroalgal and cyanobacterial succession in a coastal lagoon restored by managed realignment of agricultural land. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Li:2021:MLI

[1919] Ming Li, Ren Zhang, and Kefeng Liu. Machine learning incorporated with causal analysis for short-term prediction of sea ice. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Fu:2021:INP

[1920] Zhengyi Fu, Shengjie Zhou, Gang Yu, Rui Yang, Mingyang Han, and Zhenhua Ma. Influence of non-protein diets on hepatic metabolism and endocrine in barramundi (Lates calcarifer). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bernal-Ibanez:2021:RSU

[1921] Alejandro Bernal-Ibáñez, Eva Cacabelos, Ricardo Melo, and Ignacio Gestoso. The role of sea-urchins in marine forests from Azores, Webbnesia, and Cabo Verde: Human pressures, climate-change effects and restoration opportunities. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Rockwood:2021:MWD

[1922] R. Cotton Rockwood, Jeffrey D. Adams, Sean Hastings, Jessica Morten, and Jaime Jahncke. Modeling whale deaths from vessel strikes to reduce the risk of fatality to endangered whales. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

DeLeo:2021:TRD

[1923] Danielle M. DeLeo, Amanda Glazier, Santiago Herrera, Alexandria Barkman, and Erik E. Cordes. Transcriptomic responses of deep-sea corals experimentally exposed to crude oil and dispersant. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Brauge:2021:MTM

[1924] Thomas Brauge, Sylvain Trigueros, Arnaud Briet, Sabine Debuiche, Guylaine Leleu, Benoit Gassilloud, Amandine Wilhelm, Jean-Sébastien Py, and Graziella Midelet. MALDI-TOF mass spectrometry fingerprinting performance versus 16S rDNA sequencing to identify bacterial microflora from seafood products and sea water samples. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Thomas:2021:EPA

[1925] Gareth E. Thomas, Stefan G. Bolam, Jan L. Brant, Rodney Brash, Freya Goodsir, Clare Hynes, Terry J. McGenity, Paul S. O. McIlwaine, and Boyd A. McKew. Evaluation of polycyclic aromatic hydrocarbon pollution from the HMS Royal Oak shipwreck and effects on sediment microbial community structure. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Hogg:2021:MAS

[1926] Oliver T. Hogg, Anna-Leena Downie, Rui P. Vieira, and Chris Darby. Macrobenthic assessment of the South Sandwich Islands reveals a biogeographically distinct polar archipelago. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Collins:2021:UPN

[1927] Claire Collins, Ana Nuno, Annette Broderick, David J. Curnick, Asha de Vos, Thomas Franklin, David M. P. Jacoby, Chris Mees, James Moir-Clark, John Pearce, and Tom B. Letessier. Understanding persistent noncompliance in a remote, large-scale Marine Protected Area. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Denamiel:2021:CDE

[1928] Cléa Denamiel, Xun Huan, and Ivica Vilibić. Conceptual design of extreme sea-level early warning systems based on uncertainty quantification and engineering optimization methods. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Fourrier:2021:CRN

[1929] Marine Fourrier, Laurent Coppola, Hervé Claustre, Fabrizio D’Ortenzio, Raphaëlle Sauzède, and Jean-Pierre Gattuso. Corrigendum: A regional neural network approach to estimate water-column nutrient concentrations and carbonate system variables in the Mediterranean Sea: CANYON-MED. Frontiers in Marine Science, 8, January 2021. ISSN 2296-7745. See [722].

Clarke:2021:CSS

[1930] Penny J. Clarke, Hannah C. Cubaynes, Karen A. Stockin, Carlos Olavarría, Asha de Vos, Peter T. Fretwell, and Jennifer A. Jackson. Cetacean strandings from space: Challenges and opportunities of very high resolution satellites for the remote monitoring of cetacean mass strandings. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:SIR

[1931] Yingjie Wang, Na Liu, and Zhanhai Zhang. Sea ice reduction during winter of 2017 due to oceanic heat supplied by Pacific water in the Chukchi Sea, West Arctic Ocean. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bednarsek:2021:STOb

[1932] Nina Bednaršek, Richard Ambrose, Piero Calosi, Richard K. Childers, Richard A. Feely, Steven Y. Litvin, W. Christopher Long, John I. Spicer, Jasna Štrus, Jennifer Taylor, Faycal Kessouri, Miranda Roethler, Martha Sutula, and Stephen B. Weisberg. Synthesis of thresholds of ocean acidification impacts on decapods. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Edwards:2021:PSA

[1933] Mickie R. Edwards, Susana Cárdenas-Alayza, Michael J. Adkesson, Mya Daniels-Abdulahad, and Amy C. Hirons. Peruvian fur seals as archivists of El Niño Southern Oscillation effects. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Viloria-Gomora:2021:GVB

[1934] Lorena Viloria-Gómora, Jorge Urbán R., Braulio Leon-Lopez, and Eduardo Romero-Vivas. Geographic variation in Bryde’s whale Be4 calls in the Gulf of California: an insight to population dynamics. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Xiao:2021:CAD

[1935] Canbo Xiao, Wei Fan, Ying Chen, Yao Zhang, Kai Tang, and Nianzhi Jiao. Could artificial downwelling/upwelling mitigate oceanic deoxygenation in Western Subarctic North Pacific? Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chang:2021:CAP

[1936] Feng-Hsun Chang, Jinny Wu Yang, Ariana Chih-Hsien Liu, Hsiao-Pei Lu, Gwo-Ching Gong, Fuh-Kwo Shiah, and Chih hao Hsieh. Community assembly processes as a mechanistic explanation of the predator–prey diversity relationship in marine microbes. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ben-Zvi:2021:PBO

[1937] Or Ben-Zvi, Daniel Wangpraseurt, Omri Bronstein, Gal Eyal, and Yossi Loya. Photosynthesis and bio-optical properties of fluorescent mesophotic corals. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Coello-Camba:2021:PNP

[1938] Alexandra Coello-Camba and Susana Agustí. Picophytoplankton niche partitioning in the warmest oligotrophic sea. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Franklin:2021:SBH

[1939] Trish Franklin, Wally Franklin, Lyndon Brooks, Peter Harrison, Adam A. Pack, and Phillip J. Clapham. Social behaviour of humpback whales (Megaptera novaeangliae) in Hervey Bay, Eastern Australia, a preferential female stopover during the southern migration. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Abdulhussain:2021:ITP

[1940] Ali H. Abdulhussain, Kathryn B. Cook, Andrew D. Turner, Adam M. Lewis, Thomas S. Bibby, and Daniel J. Mayor. The influence of the toxin-producing dinoflagellate, Alexandrium catenella (1119/27), on the survival and reproduction of the marine copepod, Acartia tonsa, during prolonged exposure. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Zhu:2021:SAS

[1941] Lixin Zhu, Changzi Ge, Zhaoyang Jiang, Chunli Liu, Gang Hou, and Zhenlin Liang. Stock assessment of small yellow croaker (Larimichthys polyactis) off the coast of China using per-recruit analysis based on Bayesian inference. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Quintana:2021:PSE

[1942] Anastasia C. E. Quintana, Alfredo Giron-Nava, Samuel Urmy, Alli N. Cramer, Santiago Domínguez-Sánchez, Salvador Rodríguez-Van Dyck, Octavio Aburto-Oropeza, Xavier Basurto, and Amy Hudson Weaver. Positive social-ecological feedbacks in community-based conservation. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hossen:2021:ECG

[1943] Shaharior Hossen, Zahid Parvez Sukhan, Yusin Cho, and Kang Hee Kho. Effects of cryopreservation on gene expression and post thaw sperm quality of Pacific abalone, Haliotis discus hannai. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ross:2021:DMR

[1944] Rebecca E. Ross, Genoveva Gonzalez-Mirelis, Pablo Lozano, and Pål Buhl-Mortensen. Discerning the management-relevant ecology and distribution of sea pens (Cnidaria: Pennatulacea) in Norway and beyond. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Elizondo:2021:HPM

[1945] Elani B. Elizondo, Joanna C. Carey, Alia N. Al-Haj, Ariel E. Lugo, and Robinson W. Fulweiler. High productivity makes mangroves potentially important players in the tropical silicon cycle. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Redford:2021:BFI

[1946] Michael Redford, Sally Rouse, Peter Hayes, and Thomas A. Wilding. Benthic and fish interactions with pipeline protective structures in the North Sea. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Reigel:2021:MRB

[1947] Alicia M. Reigel, David A. Paz-García, and Michael E. Hellberg. Microbiome of a reef-building coral displays signs of acclimation to a stressful shallow hydrothermal vent habitat. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Eger:2021:RIC

[1948] Sondra L. Eger, Robert L. Stephenson, Derek Armitage, Wesley Flannery, and Simon C. Courtenay. Revisiting integrated coastal and marine management in Canada: Opportunities in the Bay of Fundy. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ortega-Cisneros:2021:ASA

[1949] Kelly Ortega-Cisneros, Kevern L. Cochrane, Nina Rivers, and Warwick H. H. Sauer. Assessing South Africa’s potential to address climate change impacts and adaptation in the fisheries sector. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Margheritini:2021:IMC

[1950] Lucia Margheritini, Per Møldrup, Rasmus Lund Jensen, Kirstine Meyer Frandsen, Yovko Ivanov Antonov, Ken Kawamoto, Lis Wollesen de Jonge, Raffaele Vaccarella, Trine Larsen Bjørgård, and Morten Enggrob Simonsen. Innovative material can mimic coral and boulder reefs properties. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Li:2021:MSF

[1951] Yongren Li, Baolu Zhang, Shuang Liang, and Yongjun Guo. Meiotic spindle formation following inhibition of first polar body formation in the Zhikong scallop (Chlamys farreri Jones et Preston). Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Sutton:2021:EDP

[1952] Tracey T. Sutton, Kevin M. Boswell, Heather D. Bracken-Grissom, Jose V. Lopez, Michael Vecchione, and Marsh Youngbluth. Editorial: Deep pelagic ecosystem dynamics in a highly impacted water column: The Gulf of Mexico after Deepwater Horizon. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Kang:2021:FWT

[1953] Hee Yoon Kang, Young-Jae Lee, Changseong Kim, Dongyoung Kim, Doo-Ho Kim, Jun-Ho Kim, Dong-Lim Choi, and Chang-Keun Kang. Food web trophic structure at marine ranch sites off the East Coast of Korea. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bang:2021:CQC

[1954] Ashley H. Y. Bang, Chao-Yang Kuo, Colin Kuo-Chang Wen, Kah-Leng Cherh, Ming-Jay Ho, Nien-Yun Cheng, Yen-Chia Chen, and Chaolun Allen Chen. Corrigendum: Quantifying coral reef resilience to climate change and human development: an evaluation of multiple empirical frameworks. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See [1300].

Cheng:2021:GWC

[1955] Yunying Cheng, Jinlin Chen, Irum Mukhtar, and Jianming Chen. Genome-wide characterization of the nuclear receptor gene family in Macrostomum lignano imply its evolutionary diversification. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cardoso:2021:MEM

[1956] Cláudio Cardoso and Rui M. A. Caldeira. Modeling the exposure of the Macaronesia Islands (NE Atlantic) to marine plastic pollution. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Madigan:2021:ITS

[1957] Daniel J. Madigan, Oliver N. Shipley, Aaron B. Carlisle, Heidi Dewar, Owyn E. Snodgrass, and Nigel E. Hussey. Isotopic tracers suggest limited trans-oceanic movements and regional residency in North Pacific blue sharks (Prionace glauca). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Roelvink:2021:CRR

[1958] Floortje E. Roelvink, Curt D. Storlazzi, Ap R. van Dongeren, and Stuart G. Pearson. Coral reef restorations can be optimized to reduce coastal flooding hazards. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Muller:2021:MCT

[1959] Alexandre Muller, Camille Poitrimol, Flávia L. D. Nunes, Aurélien Boyé, Amelia Curd, Nicolas Desroy, Louise B. Firth, Laura Bush, Andrew J. Davies, Fernando P. Lima, Martin P. Marzloff, Claudia Meneghesso, Rui Seabra, and Stanislas F. Dubois. Musical chairs on temperate reefs: Species turnover and replacement within functional groups explain regional diversity variation in assemblages associated with honeycomb worms. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Varghese:2021:SAB

[1960] Hilary Kates Varghese, Kim Lowell, Jennifer Miksis-Olds, Nancy Di-Marzio, David Moretti, and Larry Mayer. Spatial analysis of beaked whale foraging during two 12 kHz multibeam echosounder surveys. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Dickens:2021:DUM

[1961] John Dickens, Philip R. Hollyman, Tom Hart, Gemma V. Clucas, Eugene J. Murphy, Sally Poncet, Philip N. Trathan, and Martin A. Collins. Developing UAV monitoring of South Georgia and the South Sandwich Islands’ iconic land-based marine predators. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Slattery:2021:GFS

[1962] Marc Slattery and Michael P. Lesser. Gorgonians are foundation species on sponge-dominated mesophotic coral reefs in the Caribbean. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Heuer:2021:ETA

[1963] Rachael M. Heuer, John D. Stieglitz, Christina Pasparakis, Ian C. Enochs, Daniel D. Benetti, and Martin Grosell. The effects of temperature acclimation on swimming performance in the pelagic mahi-mahi (Coryphaena hippurus). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Azra:2021:CSA

[1964] Mohamad Nor Azra, Victor Tosin Okomoda, Meisam Tabatabaei, Marina Hassan, and Mhd Ikhwanuddin. The contributions of shellfish aquaculture to global food security: Assessing its characteristics from a future food perspective. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Kanki:2021:NMI

[1965] Takayuki Kanki, Kenta Nakamoto, Jun Hayakawa, Takashi Kitagawa, and Tomohiko Kawamura. A new method for investigating relationships between distribution of sessile organisms and multiple terrain variables by photogrammetry of subtidal bedrocks. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Painting:2021:SHM

[1966] Suzanne Jane Painting, Eleanor K. Haigh, Jennifer A. Graham, Simon A. Morley, Leeann Henry, Elizabeth Clingham, Rhys Hobbs, Frances Mynott, Philippe Bersuder, David I. Walker, and Tammy Stamford. St Helena marine water quality: Background conditions and development of assessment levels for coastal pollutants. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Nicolaus:2021:SDA

[1967] Marcel Nicolaus, Mario Hoppmann, Stefanie Arndt, Stefan Hendricks, Christian Katlein, Anja Nicolaus, Leonard Rossmann, Martin Schiller, and Sandra Schwegmann. Snow depth and air temperature seasonality on sea ice derived from snow buoy measurements. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Lobeto:2021:PDS

[1968] Hector Lobeto, Melisa Menendez, and Iñigo J. Losada. Projections of directional spectra help to unravel the future behavior of wind waves. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Liu:2021:GSI

[1969] Mingming Liu, Mingli Lin, Xiaoming Tang, Lijun Dong, Peijun Zhang, David Lusseau, and Songhai Li. Group size of Indo–Pacific humpback dolphins (Sousa chinensis): an examination of methodological and biogeographical variances. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zheng:2021:MDP

[1970] Zhe Zheng, Yongshan Liao, Jianming Ye, Chuangye Yang, Linda Adzigbli, Qingheng Wang, Xiaodong Du, and Yuewen Deng. Microbiota diversity in pearl oyster Pinctada fucata martensii intestine and its aquaculture environment. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Ip:2021:SDP

[1971] Yuen K. Ip, Mel V. Boo, Jeslyn S. T. Poo, Wai P. Wong, and Shit F. Chew. Sodium-dependent phosphate transporter protein 1 is involved in the active uptake of inorganic phosphate in nephrocytes of the kidney and the translocation of Pi into the tubular epithelial cells in the outer mantle of the giant clam, Tridacna squamosa. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Feng:2021:AMC

[1972] Xi Feng, Ya Gong, Meng-Qi Ye, and Zong-Jun Du. Antibiotic modulation of capsular exopolysaccharide in Pelagicoccus enzymogenes sp. nov. isolated from marine sediment. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Lanpher:2021:VMP

[1973] Kaycie B. Lanpher and Kimberly J. Popendorf. Variability of microbial particulate ATP concentrations in subeuphotic microbes due to underlying metabolic strategies in the South Pacific Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:PTP

[1974] Owen R. Liu and Renato Molina. The persistent transboundary problem in marine natural resource management. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Blaisdell:2021:DSS

[1975] Josiah Blaisdell, Hillary L. Thalmann, Willem Klajbor, Yue Zhang, Jessica A. Miller, Benjamin J. Laurel, and Maria T. Kavanaugh. A dynamic stress-scape framework to evaluate potential effects of multiple environmental stressors on Gulf of Alaska juvenile Pacific cod. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lee:2021:GMS

[1976] Dong-Hun Lee, Jung-Hyun Kim, Yung Mi Lee, Ji-Hoon Kim, Young Keun Jin, Charles Paull, Jong-Sik Ryu, and Kyung-Hoon Shin. Geochemical and microbial signatures of siboglinid tubeworm habitats at an active mud volcano in the Canadian Beaufort Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cabrerizo:2021:TFA

[1977] Marco J. Cabrerizo, Emilio Marañón, Cristina Fernández-González, Adrián Alonso-Núñez, Henrik Larsson, and María Aranguren-Gassis. Temperature fluctuation attenuates the effects of warming in estuarine microbial plankton communities. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Breton:2021:SVB

[1978] Elsa Breton, Urania Christaki, Benoit Sautour, Oscar Demonio, Dimitra-Ioli Skouroliakou, Gregory Beaugrand, Laurent Seuront, Loïck Kléparski, Adrien Poquet, Antoine Nowaczyk, Muriel Crouvoisier, Sophie Ferreira, David Pecqueur, Christophe Salmeron, Jean-Michel Brylinski, Arnaud Lheureux, and Eric Goberville. Seasonal variations in the biodiversity, ecological strategy, and specialization of diatoms and copepods in a coastal system with Phaeocystis blooms: The key role of trait trade-offs. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cloete:2021:WBC

[1979] Ryan Cloete, Jean C. Loock, Natasha R. van Horsten, Susanne Fietz, Thato N. Mtshali, Hélène Planquette, and Alakendra N. Roychoudhury. Winter biogeochemical cycling of dissolved and particulate cadmium in the Indian sector of the Southern Ocean (GEOTRACES GIpr07 transect). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Thorburn:2021:SOV

[1980] James Thorburn, Peter J. Wright, Edward Lavender, Jane Dodd, Francis Neat, Julien G. A. Martin, Caroline Lynam, and Mark James. Seasonal and ontogenetic variation in depth use by a critically endangered benthic elasmobranch and its implications for spatial management. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See corrigendum [3018].

Fuirst:2021:CFM

[1981] Matthew Fuirst, Christopher S. Ward, Caroline Schwaner, Zoie Diana, Thomas F. Schultz, and Daniel Rittschof. Compositional and functional microbiome variation between tubes of an intertidal polychaete and surrounding marine sediment. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Neely:2021:RPD

[1982] Karen L. Neely, Cynthia L. Lewis, Kathleen S. Lunz, and Lystina Kabay. Rapid population decline of the pillar coral Dendrogyra cylindrus along the Florida Reef Tract. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ryan:2021:RLF

[1983] John P. Ryan, John E. Joseph, Tetyana Margolina, Leila T. Hatch, Alyson Azzara, Alexis Reyes, Brandon L. Southall, Andrew DeVogelaere, Lindsey E. Peavey Reeves, Yanwu Zhang, Danelle E. Cline, Brent Jones, Paul McGill, Simone Baumann-Pickering, and Alison K. Stimpert. Reduction of low-frequency vessel noise in Monterey Bay National Marine Sanctuary during the COVID-19 pandemic. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Sampatakaki:2021:IIV

[1984] Angeliki Sampatakaki, Vassilis Zervakis, Ioannis Mamoutos, Elina Tragou, Alexandra Gogou, Maria Triantaphyllou, and Nikolaos Skliris. Investigation of the inherent variability of the Mediterranean Sea under contrasting extreme climatic conditions. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Suokhrie:2021:MEP

[1985] Thejasino Suokhrie, Rajeev Saraswat, and Rajiv Nigam. Multiple ecological parameters affect living benthic foraminifera in the river-influenced West-Central Bay of Bengal. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Balazy:2021:SWS

[1986] Piotr Balazy, Marti J. Anderson, Maciej Chelchowski, Maria Włodarska-Kowalczuk, Piotr Kuklinski, and Jørgen Berge. Shallow-water scavengers of polar night and day — an Arctic time-lapse photography study. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bonifacio:2021:DDS

[1987] Paulo Bonifácio, Lenka Neal, and Lénaïck Menot. Diversity of deep-sea scale-worms (Annelida, Polynoidae) in the Clarion–Clipperton Fracture Zone. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lorenzo-Felipe:2021:EDG

[1988] Álvaro Lorenzo-Felipe, Hyun Suk Shin, Sergio León-Bernabeu, Cathaysa Pérez-García, María Jesús Zamorano, Jaume Pérez-Sánchez, and Juan Manuel Afonso-López. The effect of the deformity genetic background of the breeders on the spawning quality of gilthead seabream (Sparus aurata L.). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Xu:2021:SIT

[1989] Wenzhe Xu, Yongxin Dang, Siu Gin Cheung, Zhinan Zhang, Jun Sun, Ankang Teng, and Paul K. S. Shin. Stable isotope tracer addition reveals the trophic status of benthic infauna at an intertidal area adjacent to a seagrass bed. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Mizuta:2021:SEC

[1990] Genta Mizuta, Yasushi Fukamachi, Daisuke Simizu, Yoshimasa Matsumura, Yujiro Kitade, Daisuke Hirano, Masakazu Fujii, Yoshifumi Nogi, and Kay I. Ohshima. Seasonal evolution of Cape Darnley bottom water revealed by mooring measurements. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bode:2021:AAC

[1991] Antonio Bode, Victoria Besada, Begoña Pérez-Fernández, and Lucía Viñas. Amino acid δ^{15}N can detect diet effects on pollution risks for yellow-legged gulls overlooked by trophic position. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Cheng:2021:RTD

[1992] Qichao Cheng, Wanqing Hu, and Zhiyi Bai. Research trends of development on pearl bivalve mollusks based on a bibliometric network analysis in the past 25 years. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Russnak:2021:PTT

[1993] Vanessa Russnak, Mauricio Rodriguez-Lanetty, and Ulf Karsten. Photophysiological tolerance and thermal plasticity of genetically different symbiodiniaceae endosymbiont species of cnidaria. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Boschen-Rose:2021:NMA

[1994] Rachel E. Boschen-Rose and Ana Colaço. Northern Mid-Atlantic Ridge hydrothermal habitats: a systematic review of knowledge status for environmental management. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Haye:2021:TSR

[1995] Pilar A. Haye, Xavier Turon, and Nicolás I. Segovia. Time or space? Relative importance of geographic distribution and interannual variation in three lineages of the Ascidian pyura chilensis in the Southeast Pacific Coast. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Neely:2021:SLU

[1996] Karen L. Neely, Cynthia L. Lewis, Keri O’Neil, Cheryl M. Woodley, Jennifer Moore, Zach Ransom, Amelia Moura, Ken Nedimyer, and David Vaughan. Saving the last unicorns: The genetic rescue of Florida’s pillar corals. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lundevall-Zara:2021:SAE

[1997] Maysoon Lundevall-Zara, Erik Lundevall-Zara, and Volker Brüchert. Sea-air exchange of methane in shallow inshore areas of the Baltic Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Doong:2021:ALP

[1998] Dong-Jiing Doong, Yang-Ming Fan, Jin Yih Chen, and Chia Chuen Kao. Analysis of long-period hazardous waves in the Taiwan Marine Environment Monitoring Service. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Paradinas:2021:NCT

[1999] Lola M. Paradinas, Neil A. James, Brian Quinn, Andrew Dale, and Bhavani E. Narayanaswamy. A new collection tool-kit to sample microplastics from the marine environment (sediment, seawater, and biota) using citizen science. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Murawski:2021:OCM

[2000] Jens Murawski, Jun She, Christian Mohn, Vilnis Frishfelds, and Jacob Woge Nielsen. Ocean circulation model applications for the estuarycoastal-open sea continuum. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Frishfelds:2021:MTF

[2001] Vilnis Frishfelds, Juris Sennikovs, Uldis Bethers, Jens Murawski, and Andrejs Timuhins. Modeling transit flow through port gates and connecting channel in Baltic Sea — Liepaja Port — Liepaja Lake System. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Wang:2021:GRM

[2002] Junjie Wang, Zihua Zhou, Haitao Ma, Jun Li, Yanping Qin, Jinkuan Wei, Xingyou Li, Qingliang Liao, Yunqing Li, Gongpengyang Shi, Yinyin Zhou, Yuehuan Zhang, and Ziniu Yu. Genetic recombination of the mantle color pattern of two boring giant clam (Tridacna crocea) strains. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Cui:2021:ITS

[2003] Wenxiao Cui, Qin Yang, Yin Zhang, Ardavan Farhadi, Huan Fang, Huaiping Zheng, Shengkang Li, Yueling Zhang, Mhd Ikhwanuddin, and Hongyu Ma. Integrative transcriptome sequencing reveals the molecular difference of maturation process of ovary and testis in mud crab Scylla paramamosain. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Meng:2021:PLO

[2004] Yao Meng, Hailong Liu, Ruiqiang Ding, Pengfei Lin, Mengrong Ding, and Pengfei Wang. The predictability limit of ocean mesoscale eddy tracks in the Kuroshio Extension region. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Connolly:2021:IAA

[2005] Rod M. Connolly, David V. Fairclough, Eric L. Jinks, Ellen M. Ditria, Gary Jackson, Sebastian Lopez-Marcano, Andrew D. Olds, and Kristin I. Jinks. Improved accuracy for automated counting of a fish in baited underwater videos for stock assessment. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Heide-Jorgensen:2021:BRS

[2006] Mads Peter Heide-Jørgensen, Susanna B. Blackwell, Outi M. Tervo, Adeline L. Samson, Eva Garde, Rikke G. Hansen, Manh Cu’ò’ng Ngô, Alexander S. Conrad, Per Trinhammer, Hans C. Schmidt, Mikkel-Holger S. Sinding, Terrie M. Williams, and Susanne Ditlevsen. Behavioral response study on seismic airgun and vessel exposures in narwhals. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wang:2021:DCA

[2007] Feipeng Wang, Bangqin Huang, Yuyuan Xie, Shujie Cai, Xiuxiu Wang, and Jingli Mu. Diversity, composition, and activities of nano- and picoeukaryotes in the Northern South China Sea with influences of Kuroshio intrusion. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Xu:2021:DDE

[2008] Hai’e Xu, Wanshuang Cao, Hongliang Sun, Shougang Zhang, Pan Li, Surong Jiang, and Caiyun Zhong. Dose-dependent effects of di-(2-ethylhexyl) phthalate (DEHP) in mussel Mytilus galloprovincialis. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Howell:2021:FHG

[2009] Lyndsey N. Howell and Donna J. Shaver. Foraging habits of green sea turtles (Chelonia mydas) in the Northwestern Gulf of Mexico. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bairstow:2021:IAK

[2010] Fiona Bairstow, Sven Gastauer, Luke Finley, Tom Edwards, C. Tom A. Brown, So Kawaguchi, and Martin J. Cox. Improving the accuracy of krill target strength using a shape catalog. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Orfanidis:2021:ENA

[2011] Sotiris Orfanidis, Fabio Rindi, Emma Cebrian, Simonetta Fraschetti, Ina Nasto, Ergun Taskin, Silvia Bianchelli, Vasileios Papathanasiou, Maria Kosmidou, Annalisa Caragnano, Soultana Tsioli, Stefano Ratti, Erika Fabbrizzi, Jana Verdura, Laura Tamburello, Sajmir Beqiraj, Lefter Kashta, Denada Sota, Apostolos Papadimitriou, Ezzeddine Mahmoudi, Hajdar Kiçaj, Konstantinos Georgiadis, Amel Hannachi, and Roberto Danovaro. Effects of natural and anthropogenic stressors on fucalean brown seaweeds across different spatial scales in the Mediterranean Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Concepcion:2021:ALF

[2012] Ronnie Concepcion, Elmer Dadios, Argel Bandala, Isabel Caçador, Vanessa F. Fonseca, and Bernardo Duarte. Applying limnological feature-based machine learning techniques to chemical state classification in marine transitional systems. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Jueterbock:2021:PMM

[2013] Alexander Jueterbock, Antoine J. P. Minne, J. Mark Cock, Melinda A. Coleman, Thomas Wernberg, Lydia Scheschonk, Ralf Rautenberger, Jie Zhang, and Zi-Min Hu. Priming of marine macrophytes for enhanced restoration success and food security in future oceans. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

He:2021:DFR

[2014] Xuejia He, Ziyu Wang, Ziqing Bai, Liuyu Han, and Mianrun Chen. Diel feeding rhythm and grazing selectivity of small-sized copepods in a subtropical embayment, the Northern South China Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Zhang:2021:DLE

[2015] Jiandong Zhang, Anyi Hu, Yingting Sun, Qingsong Yang, Junde Dong, Lijuan Long, and Sijun Huang. Dispersal limitation expands the diversity of coral microbiome metacommunity in the South China Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lu:2021:PMI

[2016] Cheng-Peng Lu, Chao-Guang Wei, Jun-Quan Zhu, Dao-Jun Tang, Chun-Lin Wang, and Cong-Cong Hou. PtP38 may increase the immune ability of Portunus trituberculatus stimulated by LPS imitating a gram-negative bacterial infection. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Somaweera:2021:PDE

[2017] Ruchira Somaweera, Vinay Udyawer, Michael L. Guinea, Daniela M. Ceccarelli, Rohan H. Clarke, Michelle Glover, Mathew Hourston, John Keesing, Arne Redsted Rasmussen, Kate Sanders, Richard Shine, Damian P. Thomson, and Bruce L. Webber. Pinpointing drivers of extirpation in sea snakes: a synthesis of evidence from Ashmore Reef. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Hernandez-Orts:2021:VTW

[2018] Jesús S. Hernández-Orts, David I. Hernández-Mena, Camila Pantoja, Roman Kuchta, Néstor A. García, Enrique A. Crespo, and Rocío Loizaga. A visitor of tropical waters: First record of a Clymene dolphin (Stenella clymene) off the Patagonian Coast of Argentina, with comments on diet and metazoan parasites. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Fox:2021:CPG

[2019] James Fox, Michael J. Behrenfeld, Nils Haëntjens, Alison Chase, Sasha J. Kramer, Emmanuel Boss, Lee Karp-Boss, Nerissa L. Fisher, W. Bryce Penta, Toby K. Westberry, and Kimberly H. Halsey. Corrigendum: Phytoplankton growth and productivity in the Western North Atlantic: Observations of regional variability from the NAAMES field campaigns. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See [128].

Lopez:2021:EDD

[2020] Chelsea Nicole Lopez and Dennis Arthur Hansell. Evidence of deep DOC enrichment via particle export beneath subarctic and northern subtropical fronts in the North Pacific. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Freire:2021:RMC

[2021] Kátia Meirelles Felizola Freire, Zafira da Silva de Almeida, José Raimundo Emanoel Trindade Amador, José Augusto Aragão, Ana Rosa da Rocha Araújo, Antônio Olinto Ávila da Silva, Bianca Bentes, Marcus Henrique Carneiro, Julien Chiquieri, Cezar Augusto Freire Fernandes, Marina Bezerra Figueiredo, Maurício Hostim-Silva, Érica Antunes Jimenez, Karina Annes Keunecke, Priscila Fabiana Macedo Lopes, Jocemar Tomasino Mendonça, Joelson Musiello-Fernandes, George Olavo, Camila Primitivo, Matheus Marcos Rotundo, Raynara Filho Santana, Rodrigo Sant’Ana, Guilherme Scheidt, Luis Maurício Abdon da Silva, Isaac Trindade-Santos, Gonzalo Velasco, and Marcelo Vianna. Reconstruction of marine commercial landings for the Brazilian industrial and artisanal fisheries from 1950 to 2015. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rabbani:2021:SSF

[2022] Golam Rabbani, Bertrand Chengxiang Yan, Nicole Li Ying Lee, Jillian Lean Sim Ooi, Jen Nie Lee, Danwei Huang, and Benjamin J. Wainwright. Spatial and structural factors shape seagrass-associated bacterial communities in Singapore and Peninsular Malaysia. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Wege:2021:DHS

[2023] Mia Wege, Horst Bornemann, Arnoldus Schytte Blix, Erling Sverre Nordøy, Louise Biddle, and Marthán Nieuwoudt Bester. Distribution and habitat suitability of Ross Seals in a warming ocean. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Moroni:2021:ENP

[2024] Federico Moroni, Fernando Naya-Català, M. Carla Piazzon, Simona Rimoldi, Josep Calduch-Giner, Alberto Giardini, Inés Martínez, Fabio Brambilla, Jaume Pérez-Sánchez, and Genciana Terova. The effects of nisin-producing Lactococcus lactis strain used as probiotic on gilthead sea bream (Sparus aurata) growth, gut microbiota, and transcriptional response. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Minelli:2021:OML

[2025] Annalisa Minelli, Alessandro Sarretta, Alessandro Oggioni, Caterina Bergami, Mauro Bastianini, Fabrizio Bernardi Aubry, Elisa Camatti, and Alessandra Pugnetti. Opening marine long-term ecological science: Lesson learned from the LTER-Italy Site Northern Adriatic Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cook:2021:UDS

[2026] Melissa Cook, Jaymie L. Reneker, Redwood W. Nero, Brian A. Stacy, David S. Hanisko, and Zhankun Wang. Use of drift studies to understand seasonal variability in sea turtle stranding patterns in Mississippi. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Westerbom:2021:ERS

[2027] Mats Westerbom, Patrik Kraufvelin, Olli Mustonen, and Eliecer Díaz. Explaining recruitment stochasticity at a species’ range margin. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Andrades:2021:NRW

[2028] Ryan Andrades, Jean-Christophe Joyeux, Raphael M. Macieira, Bruno Spacek Godoy, José Amorim Reis-Filho, Andrew L. Jackson, and Tommaso Giarrizzo. Niche-relationships within and among intertidal reef fish species. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wenley:2021:SPC

[2029] Jess Wenley, Kim Currie, Scott Lockwood, Blair Thomson, Federico Baltar, and Sergio E. Morales. Seasonal prokaryotic community linkages between surface and deep ocean water. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Prakash:2021:AWQ

[2030] K. Vijay Prakash, Ch. S. Geetha Vimala, T. Preethi Latha, Chiranjivi Jayaram, P. V. Nagamani, and Ch. N. V. Laxmi. Assessment of water quality along the southeast coast of India during COVID-19 lockdown. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Madarcos:2021:ULP

[2031] John Roderick V. Madarcos, Lota A. Creencia, Bethany R. Roberts, Mathew P. White, Johana Nayoan, Karyn Morrissey, and Lora E. Fleming. Understanding local perceptions of the drivers/pressures on the coastal marine environment in Palawan, Philippines. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Liu:2021:ISG

[2032] Yang Liu, Xinmei Xia, Yongjun Tian, Irene D. Alabia, Shuyang Ma, Peng Sun, and Sei-Ichi Saitoh. Influence of spawning ground dynamics on the long-term abundance of Japanese flying squid (Todarodes pacificus) winter cohort. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Huo:2021:HEP

[2033] Qianyu Huo, Jing Liu, You Cheng, Bin Cao, Ming Lei, and Yunde Liu. Hemoglobin extracted from Perinereis aibuhitensis acts as a cell culture medium supplement to reduce apoptosis. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Durbach:2021:CMC

[2034] Ian N. Durbach, Catriona M. Harris, Cameron Martin, Tyler A. Helble, E. Elizabeth Henderson, Glenn Ierley, Len Thomas, and Stephen W. Martin. Changes in the movement and calling behavior of minke whales (Balaenoptera acutorostrata) in response to Navy training. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Georgiades:2021:RVB

[2035] Eugene Georgiades, Chris Scianni, Ian Davidson, Mario N. Tamburri, Matthew R. First, Gregory Ruiz, Kevin Ellard, Marty Deveney, and Daniel Kluza. The role of vessel biofouling in the translocation of marine pathogens: Management considerations and challenges. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Martin:2021:CSD

[2036] Stephanie M. Martin, Marta Soeffker, Andy Schofield, Rhys Hobbs, Trevor Glass, and Simon A. Morley. Cetaceans sightings during research cruises in three remote Atlantic British Overseas Territories. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Mayfield:2021:SPA

[2037] Anderson B. Mayfield, Catalina Aguilar, Graham Kolodziej, Ian C. Enochs, and Derek P. Manzello. Shotgun proteomic analysis of thermally challenged reef corals. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Page:2021:AHE

[2038] Tessa M. Page, Ellie Bergstrom, and Guillermo Diaz-Pulido. Acclimation history of elevated temperature reduces the tolerance of coralline algae to additional acute thermal stress. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bridges:2021:BAC

[2039] Amelia E. H. Bridges, David K. A. Barnes, James B. Bell, Rebecca E. Ross, and Kerry L. Howell. Benthic assemblage composition of South Atlantic seamounts. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sasso:2021:LTE

[2040] Christopher R. Sasso, Paul M. Richards, Scott R. Benson, Michael Judge, Nathan F. Putman, Derke Snodgrass, and Brian A. Stacy. Leatherback turtles in the Eastern Gulf of Mexico: Foraging and migration behavior during the autumn and winter. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Conley:2021:NDM

[2041] Daniel D. Conley and Erin N. R. Hollander. A non-destructive method to create a time series of surface area for coral using 3D photogrammetry. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Ludt:2021:MMR

[2042] William B. Ludt. Missing in the middle: a review of equatorially disjunct marine taxa. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Chen:2021:ASC

[2043] Yuqing Chen, Xuena Huang, Yiyong Chen, and Aibin Zhan. Ammonia stress coping strategy in a highly invasive ascidian. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Anglada-Ortiz:2021:PFP

[2044] Griselda Anglada-Ortiz, Katarzyna Zamelczyk, Julie Meilland, Patrizia Ziveri, Melissa Chierici, Agneta Fransson, and Tine L. Rasmussen. Planktic foraminiferal and pteropod contributions to carbon dynamics in the Arctic Ocean (North Svalbard Margin). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ibanhez:2021:DNE

[2045] J. Severino P. Ibánhez, Xosé Antón Álvarez-Salgado, and Carlos Rocha. Does nitrate enrichment accelerate organic matter turnover in subterranean estuaries? Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Cruz:2021:HRM

[2046] Elisabet R. Cruz, Rita Nolasco, Xosé A. Padin, Miguel Gilcoto, Jose M. F. Babarro, Jesús Dubert, and Fiz F. Pérez. A high-resolution modeling study of the circulation patterns at a coastal embayment: Ría de Pontevedra (NW Spain) under upwelling and downwelling conditions. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Humphries:2021:TAM

[2047] Fran Humphries, Muriel Rabone, and Marcel Jaspars. Traceability approaches for marine genetic resources under the proposed ocean (BBNJ) treaty. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Cocci:2021:FMT

[2048] Paolo Cocci, Emanuele Troli, Mauro Angeletti, and Francesco Alessandro Palermo. Field monitoring of Tritia mutabilis (Linnaeus, 1758) egg capsule deposition and intracapsular embryonic patterns using artificial substrates and machine learning-based approaches. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

McTigue:2021:REG

[2049] Nathan D. McTigue, Quentin A. Walker, and Carolyn A. Currin. Refining estimates of greenhouse gas emissions from salt marsh “Blue carbon” erosion and decomposition. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Washburn:2021:EHT

[2050] Travis W. Washburn, Daniel O. B. Jones, Chih-Lin Wei, and Craig R. Smith. Environmental heterogeneity throughout the clarion-clipperton zone and the potential representativity of the APEI network. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Chatzievangelou:2021:IDV

[2051] Damianos Chatzievangelou, Nixon Bahamon, Séverine Martini, Joaquin del Rio, Giorgio Riccobene, Michael Tangherlini, Roberto Danovaro, Fabio C. De Leo, Benoit Pirenne, and Jacopo Aguzzi. Integrating diel vertical migrations of bioluminescent deep scattering layers into monitoring programs. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lott:2021:HLB

[2052] Christian Lott, Andreas Eich, Dorothée Makarow, Boris Unger, Miriam van Eekert, Els Schuman, Marco Segre Reinach, Markus T. Lasut, and Miriam Weber. Half-life of biodegradable plastics in the marine environment depends on material, habitat, and climate zone. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Studivan:2021:SET

[2053] Michael S. Studivan, Alycia Shatters, Danielle L. Dodge, Jeff L. Beal, and Joshua D. Voss. Synergistic effects of thermal stress and estuarine discharge on transcriptomic variation of Montastraea cavernosa corals in Southeast Florida. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Randall:2021:ECS

[2054] Carly J. Randall, Christine Giuliano, Andrew J. Heyward, and Andrew P. Negri. Enhancing coral survival on deployment devices with microrefugia. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

vanPutten:2021:DIS

[2055] Ingrid van Putten, Rachel Kelly, Rachel D. Cavanagh, Eugene J. Murphy, Annette Breckwoldt, Stephanie Brodie, Christopher Cvitanovic, Mark Dickey-Collas, Lisa Maddison, Jess Melbourne-Thomas, Haritz Arrizabalaga, Kumiko Azetsu-Scott, Lynnath E. Beckley, Richard Bellerby, Andrew J. Constable, Greg Cowie, Karen Evans, Marion Glaser, Julie Hall, Alistair J. Hobday, Nadine M. Johnston, Joel K. Llopiz, Franz Mueter, Frank E. Muller-Karger, Kevin C. Weng, Dieter Wolf-Gladrow, and José C. Xavier. A decade of incorporating social sciences in the Integrated Marine Biosphere Research Project (IMBeR): Much done, much to do? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:IMO

[2056] Sheng Liu, Qinggang Xue, Hongqiang Xu, and Zhihua Lin. Identification of main oyster species and comparison of their genetic diversity in Zhejiang Coast, South of Yangtze River Estuary. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Santin:2021:MCP

[2057] Andreu Santín, Jordi Grinyó, Maria Jesús Uriz, Claudio Lo Iacono, Josep Maria Gili, and Pere Puig. Mediterranean coral provinces as a sponge diversity reservoir: Is there a Mediterranean cold-water coral sponge fauna? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Sanchez-Marono:2021:ASI

[2058] Sonia Sánchez-Maroño, Andrés Uriarte, Leire Ibaibarriaga, and Leire Citores. Adapting simple index-based catch rules for data-limited stocks to short-lived fish stocks’ characteristics. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

OBrien:2021:ECS

[2059] Charlotte L. O’Brien, Peter T. Spooner, Jack H. Wharton, Eirini Papachristopoulou, Nicolas Dutton, David Fairman, Rebecca Garratt, Tianying Li, Francesco Pallottino, Fiona Stringer, and David J. R. Thornalley. Exceptional 20th Century shifts in deep-sea ecosystems are spatially heterogeneous and associated with local surface ocean variability. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Campanella:2021:FIM

[2060] Fabio Campanella, Martin A. Collins, Emma F. Young, Vladimir Laptikhovsky, Paul Whomersley, and Jeroen van der Kooij. First insight of meso- and bentho-pelagic fish dynamics around remote seamounts in the South Atlantic Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bartzke:2021:IPH

[2061] Gerhard Bartzke, Lennart Siemann, Robert Büssing, Paride Nardone, Katinka Koll, Dierk Hebbeln, and Katrin Huhn. Investigating the prevailing hydrodynamics around a cold-water coral colony using a physical and a numerical approach. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Omori:2021:MIS

[2062] Kristen L. Omori, Cindy A. Tribuzio, Elizabeth A. Babcock, and John M. Hoenig. Methods for identifying species complexes using a novel suite of multivariate approaches and multiple data sources: a case study with Gulf of Alaska rockfish. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Hollingsworth:2021:SVA

[2063] Anita L. Hollingsworth, Daniel O. B. Jones, and C. Robert Young. Spatial variability of abyssal nitrifying microbes in the North-Eastern Clarion–Clipperton Zone. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Reineman:2021:COA

[2064] Dan R. Reineman, Kellee Koenig, Nik Strong-Cvetich, and John N. Kittinger. Conservation opportunities arise from the co-occurrence of surfing and key biodiversity areas. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Rudd:2021:CLM

[2065] Merrill B. Rudd, Jason M. Cope, Chantel R. Wetzel, and James Hastie. Catch and length models in the stock synthesis framework: Expanded application to data-moderate stocks. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Hao:2021:TNO

[2066] You-You Hao, Zhuo-Yi Zhu, Fu-Tao Fang, Tihana Novak, Milan Čanković, Enis Hrustić, Zrinka Ljubešić, Ming Li, Jin-Zhou Du, Rui-Feng Zhang, and Blaženka Gašparović. Tracing nutrients and organic matter changes in eutrophic Wenchang (China) and oligotrophic Krka (Croatia) estuaries: a comparative study. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Stadtlander:2021:BRI

[2067] Christian T. K.-H. Stadtländer. Book review: International Marine Mammal Law. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Laming:2021:CRB

[2068] Sven R. Laming, Magdalini Christodoulou, Pedro Martinez Arbizu, and Ana Hilário. Comparative reproductive biology of deep-sea ophiuroids inhabiting polymetallic-nodule fields in the clarion-clipperton fracture zone. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Nehir:2021:ICD

[2069] Münevver Nehir, Mario Esposito, Christian Begler, Carsten Frank, Oliver Zielinski, and Eric P. Achterberg. Improved calibration and data processing procedures of OPUS optical sensor for high-resolution in situ monitoring of nitrate in seawater. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Blumenthal:2021:CIS

[2070] Janice M. Blumenthal, Jane L. Hardwick, Timothy J. Austin, Annette C. Broderick, Paul Chin, Lucy Collyer, Gina Ebanks-Petrie, Leah Grant, Lorri D. Lamb, Jeremy Olynik, Lucy C. M. Omeyer, Alejandro Prat-Varela, and Brendan J. Godley. Cayman Islands sea turtle nesting population increases over 22 years of monitoring. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Anderson:2021:RFI

[2071] Emily R. Anderson, Jack Butler, and Mark J. Butler. Response of fish and invertebrate larvae to backreef sounds at varying distances: Implications for habitat restoration. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Harvey:2021:FAA

[2072] Euan S. Harvey, Stephanie L. Watts, Benjamin J. Saunders, Damon Driessen, Laura A. F. Fullwood, Michael Bunce, Se Songploy, Jes Kettratad, Paweena Sitaworawet, Sarin Chaiyakul, Travis S. Elsdon, and Michael J. Marnane. Fish assemblages associated with oil and gas platforms in the Gulf of Thailand. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Tamis:2021:DEI

[2073] Jacqueline E. Tamis, Ruud H. Jongbloed, Gerjan J. Piet, and Robbert G. Jak. Developing an environmental impact assessment for floating island applications. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Bergman:2021:MOR

[2074] Jessica L. Bergman, William Leggat, and Tracy D. Ainsworth. The metaorganism response of the environmental generalist Pocillopora damicornis exposed to differential accumulation of heat stress. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Polejack:2021:IOS

[2075] Andrei Polejack. The importance of ocean science diplomacy for ocean affairs, global sustainability, and the UN decade of ocean science. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Ma:2021:HDP

[2076] Xiao Ma, Charles A. Jacoby, and Kevin B. Johnson. High densities of a prochlorophyte (unresolved species) inhibit grazing by the herbivorous copepod Parvocalanus crassirostris. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ma:2021:GCP

[2077] Xiao Ma, Charles A. Jacoby, and Kevin B. Johnson. Grazing by the copepod Parvocalanus crassirostris on Picochlorum sp. at harmful bloom densities and the role of particle size. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Landolfi:2021:EMN

[2078] Angela Landolfi, Sophie Rabouille, and Beatriz Mouriño-Carballido. Editorial: Marine N2 fixation: Recent discoveries and future challenges. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Fernandez-Betelu:2021:FFE

[2079] Oihane Fernandez-Betelu, Isla M. Graham, Kate L. Brookes, Barbara J. Cheney, Tim R. Barton, and Paul M. Thompson. Far-field effects of impulsive noise on coastal bottlenose dolphins. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Haunost:2021:ICC

[2080] Mathias Haunost, Ulf Riebesell, Francesco D’Amore, Ole Kelting, and Lennart T. Bach. Influence of the calcium carbonate shell of coccolithophores on ingestion and growth of a dinoflagellate predator. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

McCullough:2021:ASB

[2081] Jennifer L. K. McCullough, Johanna L. K. Wren, Erin M. Oleson, Ann N. Allen, Zachary A. Siders, and Erik S. Norris. An acoustic survey of beaked whales and Kogia spp. in the Mariana Archipelago using drifting recorders. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Kelly:2021:IBU

[2082] Kyla J. Kelly, Fei-Xue Fu, Xiaowen Jiang, He Li, Dong Xu, Nina Yang, Michelle A. DeMers, Joshua D. Kling, Kunshan Gao, Naihao Ye, and David A. Hutchins. Interactions between ultraviolet B radiation, warming, and changing nitrogen source may reduce the accumulation of toxic Pseudo-nitzschia multiseries biomass in future coastal oceans. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

TinHan:2021:SOP

[2083] Thomas C. TinHan and R. J. David Wells. Spatial and ontogenetic patterns in the trophic ecology of juvenile bull sharks (Carcharhinus leucas) from the Northwest Gulf of Mexico. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sukhan:2021:TSA

[2084] Zahid Parvez Sukhan, Md. Rajib Sharker, Yusin Cho, Shaharior Hossen, Kap Seong Choi, and Kang Hee Kho. Thermal stress affects gonadal maturation by regulating GnRH, GnRH receptor, APGWamide, and serotonin receptor gene expression in male Pacific abalone, Haliotis discus hannai during breeding season. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ferron:2021:CTS

[2085] Bruno Ferron, Pascale Bouruet-Aubertot, Katrin Schroeder, Harry L. Bryden, Yannis Cuypers, and Mireno Borghini. Contribution of thermohaline staircases to deep water mass modifications in the Western Mediterranean Sea from microstructure observations. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Wright:2021:YTB

[2086] Serena R. Wright, David Righton, Joachim Naulaerts, Robert J. Schallert, Chris A. Griffiths, Taylor Chapple, Daniel Madigan, Vladimir Laptikhovsky, Victoria Bendall, Rhys Hobbs, Doug Beare, Elizabeth Clingham, Barbara Block, and Martin A. Collins. Yellowfin tuna behavioural ecology and catchability in the South Atlantic: The right place at the right time (and depth). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Puccinelli:2021:OPU

[2087] Eleonora Puccinelli, Fany Sardenne, Laure Pecquerie, Sarah E. Fawcett, Eric Machu, and Philippe Soudant. Omega-3 pathways in upwelling systems: The link to nitrogen supply. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:MIS

[2088] Kwang-Ming Liu, Ya-Wen Huang, and Hua-Hsun Hsu. Management implications for skates and rays based on analysis of life history parameters. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Burnham:2021:EMM

[2089] Rianna E. Burnham, Svein Vagle, Caitlin O’Neill, and Krista Trounce. The efficacy of management measures to reduce vessel noise in critical habitat of southern resident killer whales in the Salish Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pelusio:2021:IBD

[2090] Nicole Francesca Pelusio, Daniel Scicchitano, Luca Parma, Francesco Dondi, Elisa Brini, Federica D’Amico, Marco Candela, Manuel Yúfera, Neda Gilannejad, Francisco Javier Moyano, Pier Paolo Gatta, and Alessio Bonaldo. Interaction between dietary lipid level and seasonal temperature changes in gilthead sea bream Sparus aurata: Effects on growth, fat deposition, plasma biochemistry, digestive enzyme activity, and gut bacterial community. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Benhemma-LeGall:2021:BSR

[2091] Aude Benhemma-Le Gall, Isla M. Graham, Nathan D. Merchant, and Paul M. Thompson. Broad-scale responses of harbor porpoises to piledriving and vessel activities during offshore windfarm construction. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Grare:2021:IWG

[2092] Laurent Grare, Nicholas M. Statom, Nick Pizzo, and Luc Lenain. Instrumented wave gliders for air–sea interaction and upper ocean research. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Pati:2021:ICC

[2093] Siddhartha Pati, Tanmay Sarkar, Hassan I. Sheikh, Kaushik Kumar Bharadwaj, Pranab Kishor Mohapatra, Anil Chatterji, Bisnu Prasad Dash, Hisham Atan Edinur, and Bryan Raveen Nelson. γ-irradiated chitosan from Carcinoscorpius rotundicauda (Latreille, 1802) improves the shelf life of refrigerated aquatic products. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Cusano:2021:SCB

[2094] Dana A. Cusano, David Paton, Michael J. Noad, and Rebecca A. Dunlop. Socially complex breeding interactions in humpback whales are mediated using a complex acoustic repertoire. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Finucci:2021:DSD

[2095] Brittany Finucci, Clinton A. J. Duffy, Tom Brough, Malcolm P. Francis, Marco Milardi, Matthew H. Pinkerton, Grady Petersen, and Fabrice Stephenson. Drivers of spatial distributions of basking shark (Cetorhinus maximus) in the Southwest Pacific. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Astarloa:2021:RCO

[2096] Amaia Astarloa, Maite Louzao, Joana Andrade, Lucy Babey, Simon Berrow, Oliver Boisseau, Tom Brereton, Ghislain Dorémus, Peter G. H. Evans, Nicola K. Hodgins, Mark Lewis, Jose Martinez-Cedeira, Malin L. Pinsky, Vincent Ridoux, Camilo Saavedra, M. Begoña Santos, James T. Thorson, James J. Waggitt, Dave Wall, and Guillem Chust. The role of climate, oceanography, and prey in driving decadal spatio-temporal patterns of a highly mobile top predator. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Moeller:2021:CCC

[2097] Mareen Moeller, Sascha Pawlowski, Mechtild Petersen-Thiery, Ingo B. Miller, Samuel Nietzer, Yannik Heisel-Sure, Matthias Y. Kellermann, and Peter J. Schupp. Challenges in current coral reef protection — possible impacts of UV filters used in sunscreens, a critical review. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Nikouli:2021:DLE

[2098] Eleni Nikouli, Konstantinos Ar. Kormas, Yang Jin, Yngvar Olsen, Ingrid Bakke, and Olav Vadstein. Dietary lipid effects on gut microbiota of first feeding Atlantic salmon (Salmo salar L.). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Carlson:2021:MAC

[2099] Daniel F. Carlson, Gareth Carr, J. L. Crosbie, Peter Lundgren, Nicolas Peissel, Pippa Pett, Will Turner, and Søren Rysgaard. The 2017 mission Arctic citizen science sailing expedition conductivity, temperature, and depth profiles in Western Greenland and Baffin Bay. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Amores:2021:CFM

[2100] Angel Amores, Marta Marcos, Rodrigo Pedreros, Gonéri Le Cozannet, Sophie Lecacheux, Jérémy Rohmer, Jochen Hinkel, Geronimo Gussmann, Thomas van der Pol, Ali Shareef, and Zammath Khaleel. Coastal flooding in the Maldives induced by mean sea-level rise and wind-waves: From global to local coastal modelling. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Nian:2021:IPA

[2101] Rui Nian, Qiang Yuan, Hui He, Xue Geng, Chi-Wei Su, Bo He, and Amaury Lendasse. The identification and prediction in abundance variation of Atlantic cod via long short-term memory with periodicity, time– frequency co-movement, and lead-lag effect across sea surface temperature, sea surface salinity, catches, and prey biomass from 1919 to 2016. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Jacoby:2021:SNA

[2102] David M. P. Jacoby, Bethany S. Fairbairn, Bryan S. Frazier, Austin J. Gallagher, Michael R. Heithaus, Steven J. Cooke, and Neil Hammerschlag. Social network analysis reveals the subtle impacts of tourist provisioning on the social behavior of a generalist marine apex predator. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Rueger:2021:NFU

[2103] Theresa Rueger, Rebecca Branconi, Catheline Y. M. Froehlich, Siobhan J. Heatwole, Marian Y. L. Wong, and Peter M. Buston. The next frontier in understanding the evolution of coral reef fish societies. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Olinger:2021:SMS

[2104] Lauren K. Olinger, Wendy K. Strangman, Steven E. McMurray, and Joseph R. Pawlik. Sponges with microbial symbionts transform dissolved organic matter and take up organohalides. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

MosqueraGimenez:2021:COC

[2105] Ángela Mosquera Giménez, Pedro Vélez-Belchí, Jesús Rivera, Safo Piñeiro, Noelia Fajar, Verónica Caínzos, Rosa Balbín, Juan Antonio Jiménez Aparicio, Carlos Dominguez-Carrió, Jordi Blasco-Ferre, Marina Carreiro-Silva, Telmo Morato, Patricia Puerta, and Covadonga Orejas. Corrigendum: Ocean circulation over North Atlantic underwater features in the path of the Mediterranean outflow water: The Ormonde and Formigas Seamounts, and the Gazul Mud Volcano. Frontiers in Marine Science, 8, February 2021. ISSN 2296-7745. See [9].

Mansour:2021:EEP

[2106] Joost Samir Mansour and Konstantinos Anestis. Eco-evolutionary perspectives on mixoplankton. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Walker:2021:OSC

[2107] Brian K. Walker, Nicholas R. Turner, Hunter K. G. Noren, Samantha F. Buckley, and Kelly A. Pitts. Optimizing Stony Coral Tissue Loss Disease (SCTLD) intervention treatments on Montastraea cavernosa in an endemic zone. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Choi:2021:GDP

[2108] Hee kyu Choi, Ji Eun Jang, Seo Yeon Byeon, Yu Rim Kim, Dale Maschette, Sangdeok Chung, Seok-Gwan Choi, Hyun-Woo Kim, and Hyuk Je Lee. Genetic diversity and population structure of the Antarctic toothfish, Dissostichus mawsoni, using mitochondrial and microsatellite DNA markers. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Jones-Williams:2021:RIM

[2109] Kirstie Jones-Williams, Tamara S. Galloway, Victoria L. Peck, and Clara Manno. Remote, but not isolated — microplastics in the sub-surface waters of the Canadian Arctic Archipelago. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Graiff:2021:SPP

[2110] Angelika Graiff, Inka Bartsch, Karin Glaser, and Ulf Karsten. Seasonal photophysiological performance of adult Western Baltic Fucus vesiculosus (Phaeophyceae) under ocean warming and acidification. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Fernandes-Salvador:2021:CSF

[2111] Jose A. Fernandes-Salvador, Keith Davidson, Marc Sourisseau, Marta Revilla, Wiebke Schmidt, Dave Clarke, Peter I. Miller, Paola Arce, Raúl Fernández, Luz Maman, Alexandra Silva, Callum Whyte, Maria Mateo, Patricia Neira, Marcos Mateus, Manuel Ruiz-Villarreal, Luis Ferrer, and Joe Silke. Current status of forecasting toxic harmful algae for the North-East Atlantic shellfish aquaculture industry. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Callies:2021:ULD

[2112] Ulrich Callies, Markus Kreus, Wilhelm Petersen, and Yoana G. Voynova. On using Lagrangian drift simulations to aid interpretation of in situ monitoring data. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Brana:2021:TES

[2113] Carlos Brais Carballeira Braña, Kristine Cerbule, Paula Senff, and Insa Kristina Stolz. Towards environmental sustainability in marine finfish aquaculture. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Liu:2021:RRR

[2114] Fang Liu, Wenyuan Shi, Haihui Ye, An Liu, and Zhihuang Zhu. RNAi reveals role of insulin-like androgenic gland hormone 2 (IAG2) in sexual differentiation and growth in hermaphrodite shrimp. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Mote:2021:DSA

[2115] Sambhaji Mote, Vishal Gupta, Kalyan De, Afreen Hussain, Kuldeep More, Mandar Nanajkar, and Baban Ingole. Differential symbiodiniaceae association with coral and coral-eroding sponge in a bleaching impacted marginal coral reef environment. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Doering:2021:SIS

[2116] Kristin Doering, Claudia Ehlert, Katharina Pahnke, Martin Frank, Ralph Schneider, and Patricia Grasse. Silicon isotope signatures of radiolaria reveal taxon-specific differences in isotope fractionation. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Polyakov:2021:CBA

[2117] Igor V. Polyakov, Matthew B. Alkire, Bodil A. Bluhm, Kristina A. Brown, Eddy C. Carmack, Melissa Chierici, Seth L. Danielson, Ingrid Ellingsen, Elizaveta A. Ershova, Katarina Gårdfeldt, Randi B. Ingvaldsen, Andrey V. Pnyushkov, Dag Slagstad, and Paul Wassmann. Corrigendum: Borealization of the Arctic Ocean in response to anomalous advection from sub-Arctic seas. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See [593].

Paulus:2021:SLD

[2118] Eva Paulus. Shedding light on deep-sea biodiversity — a highly vulnerable habitat in the face of anthropogenic change. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Kunze:2021:MDI

[2119] Charlotte Kunze, Mirco Wölfelschneider, and Lena Rölfer. Multiple driver impacts on rocky intertidal systems: The need for an integrated approach. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lishchenko:2021:ASA

[2120] Fedor Lishchenko and Jessica Briony Jones. Application of shape analyses to recording structures of marine organisms for stock discrimination and taxonomic purposes. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Flynn:2021:MET

[2121] Kevin J. Flynn, Susan A. Kimmance, Darren R. Clark, Aditee Mitra, Luca Polimene, and William H. Wilson. Modelling the effects of traits and abiotic factors on viral lysis in phytoplankton. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Fanning:2021:CIR

[2122] Lucia Fanning, Robin Mahon, Sanya Compton, Chris Corbin, Patrick Debels, Milton Haughton, Sherry Heileman, Nicole Leotaud, Patrick Mc-Conney, Manuel Perez Moreno, Terrence Phillips, and Cesar Toro. Challenges to implementing regional ocean governance in the wider Caribbean region. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Rogers:2021:MGR

[2123] Alex D. Rogers, Amy Baco, Elva Escobar-Briones, Duncan Currie, Kristina Gjerde, Judith Gobin, Marcel Jaspars, Lisa Levin, Katrin Linse, Muriel Rabone, Eva Ramirez-Llodra, Javier Sellanes, Timothy M. Shank, Kerry Sink, Paul V. R. Snelgrove, Michelle L. Taylor, Daniel Wagner, and Harriet Harden-Davies. Marine genetic resources in areas beyond national jurisdiction: Promoting marine scientific research and enabling equitable benefit sharing. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See corrigendum [2702].

Yan:2021:ATI

[2124] Xiaobo Yan, Yu Chen, Xiaohui Dong, Beiping Tan, Hao Liu, Shuang Zhang, Shuyan Chi, Qihui Yang, Hongyu Liu, and Yuanzhi Yang. Ammonia toxicity induces oxidative stress, inflammatory response and apoptosis in hybrid grouper ([Venus sign] Epinephelus fuscoguttatus × [male sign] E. lanceolatu). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Takata:2021:GWS

[2125] Kenji Takata, Fumihito Iwase, Akira Iguchi, Hideaki Yuasa, Hiroki Taninaka, Nozomu Iwasaki, Kouji Uda, Tomohiko Suzuki, Masanori Nonaka, Taisei Kikuchi, and Nina Yasuda. Genome-wide SNP data revealed notable spatial genetic structure in the deep-sea precious coral Corallium japonicum. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Chassignet:2021:TML

[2126] Eric P. Chassignet, Xiaobiao Xu, and Olmo Zavala-Romero. Tracking marine litter with a global ocean model: Where does it go? Where does it come from? Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Liang:2021:IAM

[2127] Shuang Liang, Weiwei You, Xuan Luo, Jinwei Ke, Miaoqin Huang, Yongjun Guo, and Caihuan Ke. Integrated analysis of mRNA and miRNA changes in two Haliotis diversicolor genotypes and their hybrid. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Macedo:2021:MOR

[2128] Maria Wanna Figueiredo Sena Macedo, Nicolau Brito da Cunha, Juliana Araújo Carneiro, Rosiane Andrade da Costa, Sergio Amorim de Alencar, Marlon Henrique Cardoso, Octávio Luiz Franco, and Simoni Campos Dias. Marine organisms as a rich source of biologically active peptides. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

IJsseldijk:2021:PHP

[2129] Lonneke L. IJsseldijk, Kees C. J. Camphuysen, Guido O. Keijl, Gerard Troost, and Geert Aarts. Predicting harbor porpoise strandings based on near-shore sightings indicates elevated temporal mortality rates. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pipithkul:2021:HCG

[2130] Supisara Pipithkul, Sota Ishizu, Akifumi Shimura, Hiroyuki Yokochi, Satoshi Nagai, Hironobu Fukami, and Nina Yasuda. High clonality and geographically separated cryptic lineages in the threatened temperate coral, Acropora pruinosa. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Dedman:2021:IIC

[2131] Craig J. Dedman, Marwa M. I. Rizk, Joseph A. Christie-Oleza, and Gemma-Louise Davies. Investigating the impact of cerium oxide nanoparticles upon the ecologically significant marine cyanobacterium prochlorococcus. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Freestone:2021:SSC

[2132] David Freestone. The Sargasso Sea Commission: an evolving new paradigm for High Seas ecosystem governance? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Calle-Trivino:2021:AFI

[2133] Johanna Calle-Triviño, Aarón Israel Muñiz-Castillo, Camilo Cortés-Useche, Megan Morikawa, Rita Sellares-Blasco, and Jesús Ernesto Arias-González. Approach to the functional importance of Acropora cervicornis in outplanting sites in the Dominican Republic. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Han:2021:TSR

[2134] Mingyang Han, Zhengyi Fu, Rui Yang, Jian G. Qin, and Zhenhua Ma. Temperature significantly regulates gene expressions of vertebrae chondrocytes and osteoblasts in juvenile golden pompano (Trachinotus ovatus). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:CGG

[2135] Feng Liu and James T. Melton. Chloroplast genomes of the green-tide forming alga Ulva compressa: Comparative Chloroplast genomics in the genus Ulva (Ulvophyceae, Chlorophyta). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ramenzoni:2021:CGT

[2136] Victoria C. Ramenzoni. Co-governance, transregional maritime conventions, and indigenous customary practices among subsistence fishermen in Ende, Indonesia. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Fielding:2021:AAS

[2137] Russell Fielding and Jeremy J. Kiszka. Artisanal and aboriginal subsistence whaling in Saint Vincent and the Grenadines (Eastern Caribbean): History, catch characteristics, and needs for research and management. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Ziervogel:2021:SRA

[2138] Kai Ziervogel, Julia Sweet, Andrew R. Juhl, and Uta Passow. Sediment resuspension and associated extracellular enzyme activities measured ex situ: a mechanism for benthic-pelagic coupling in the deep Gulf of Mexico. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Williams:2021:PSC

[2139] Stacey M. Williams, Jorge García-Sais, and Jorge Sabater-Clavell. Prevalence of Stony Coral Tissue Loss Disease at El Seco, a mesophotic reef system off Vieques Island, Puerto Rico. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Jin:2021:IBS

[2140] Jae-Youll Jin, Jong Dae Do, Jin-Soon Park, Jun Seok Park, Byunggil Lee, Sung-Doo Hong, Sung-Ju Moon, Keun Choon Hwang, and Yeon S. Chang. Intelligent buoy system (INBUS): Automatic lifting observation system for macrotidal coastal waters. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Park:2021:COU

[2141] Young-Gyu Park, Seongbong Seo, Dong Guk Kim, Joocheul Noh, and Hyuk Min Park. Coastal observation using a vertical profiling system at the Southern Coast of Korea. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Natnan:2021:IOT

[2142] Maya E. Natnan, Chen-Fei Low, Chou-Min Chong, Hamidun Bunawan, and Syarul N. Baharum. Integration of Omics tools for understanding the fish immune response due to microbial challenge. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Schadewell:2021:FME

[2143] Yvonne Schadewell and Clare I. M. Adams. Forensics meets ecology — environmental DNA offers new capabilities for marine ecosystem and fisheries research. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Schere:2021:JCW

[2144] Constance M. Schéré, Kate Schreckenberg, Terence P. Dawson, and Nikoleta Jones. It’s just conservation: To what extent are Marine Protected Areas in the Irish Sea equitably governed and managed? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Papaioannou:2021:AWW

[2145] Eva A. Papaioannou, Rebecca L. Selden, Julia Olson, Bonnie J. McCay, Malin L. Pinsky, and Kevin St. Martin. Not all those who wander are lost — responses of fishers’ communities to shifts in the distribution and abundance of fish. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Quinones:2021:SDM

[2146] Daniel R. Quijano Quiñones, Jorge A. López-Rocha, Isis Hernández-Herrera, and Edgar Torres-Irineo. Spatial dynamics modeling of smallscale fishing fleets with a random walk approach. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Sykora-Bodie:2021:UFM

[2147] Seth T. Sykora-Bodie, Jorge G. Álvarez-Romero, Javier A. Arata, Alistair Dunn, Jefferson T. Hinke, Grant Humphries, Christopher Jones, Pål Skogrand, Katharina Teschke, Philip N. Trathan, Dirk Welsford, Natalie C. Ban, Grant Murray, and David A. Gill. Using forecasting methods to incorporate social, economic, and political considerations into Marine Protected Area planning. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Guzman-Urieta:2021:SPL

[2148] Edgar O. Guzmán-Urieta and Eric Jordán-Dahlgren. Spatial patterns of a lethal white syndrome outbreak in Pseudodiploria strigosa. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Morales-Bojorquez:2021:IBM

[2149] Enrique Morales-Bojórquez, Manuel O. Nevárez-Martínez, Gabriela García-Alberto, Héctor Villalobos, Hugo Aguirre-Villaseñor, Estefani Larios-Castro, Sergio S. González-Peláez, Dana I. Arizmendi-Rodríguez, and María de los Ángeles Martínez-Zavala. Interaction between marine fauna and the small pelagic fishery in the coastal environment of the Gulf of California, Mexico. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Benaltabet:2021:HMR

[2150] Tal Benaltabet, Eldad Gutner-Hoch, and Adi Torfstein. Heavy metal, rare earth element and Pb isotope dynamics in mussels during a depuration experiment in the Gulf of Aqaba, Northern Red Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Audenhaege:2021:HRV

[2151] Loïc Van Audenhaege, Emmeline Broad, Katharine R. Hendry, and Veerle A. I. Huvenne. High-resolution vertical habitat mapping of a deep-sea cliff offshore Greenland. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ormaza-Gonzailez:2021:CIB

[2152] Franklin I. Ormaza-Gonzaìlez, Divar Castro-Rodas, and Peter J. Statham. COVID-19 impacts on beaches and coastal water pollution at selected sites in Ecuador, and management proposals post-pandemic. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Mateus:2021:MSC

[2153] Marcos Mateus, Miguel Remondes, Lígia Pinto, and Alexandra Silva. Monitoring SARS-CoV-2 as a microbiological risk in shellfish aquaculture. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

McCordic:2021:BSS

[2154] Jessica A. McCordic, Annamaria I. DeAngelis, Logan R. Kline, Candace McBride, Giverny G. Rodgers, Timothy J. Rowell, Jeremy Smith, Jenni A. Stanley, Allison Stokoe, and Sofie M. Van Parijs. Biological sound sources drive soundscape characteristics of two Australian marine parks. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chang:2021:CPC

[2155] Grace Chang, Genevra Harker-Klimeš, Kaustubha Raghukumar, Brian Polagye, Joseph Haxel, James Joslin, Frank Spada, and Garrett Staines. Clearing a path to commercialization of marine renewable energy technologies through public–private collaboration. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Sunanda:2021:LTT

[2156] N. Sunanda, J. Kuttippurath, R. Peter, Kunal Chakraborty, and A. Chakraborty. Long-term trends and impact of SARS-CoV-2 COVID-19 lockdown on the primary productivity of the North Indian Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Soper:2021:ERS

[2157] Deanna M. Soper. Education and research: a symbiosis to better understand a novel coral disease. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Passos:2021:BAT

[2158] Flávio Dias Passos, Paulo Vinicius Ferraz Corrêa, and Marcel Sabino Miranda. Better alone than in bad company: New species of Caudofoveate Limifossorid (Mollusca, Aplacophora) unravels patterns of distribution hidden in the deep Atlantic. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Sellares-Blasco:2021:ACR

[2159] Rita I. Sellares-Blasco, Maria F. Villalpando, Sergio D. Guendulain-García, and Aldo Croquer. Assisted coral reproduction in the Dominican Republic: a successful story to replicate in the Caribbean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sylvester:2021:DCS

[2160] Zephyr T. Sylvester, Matthew C. Long, and Cassandra M. Brooks. Detecting climate signals in Southern Ocean krill growth habitat. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Haver:2021:LVA

[2161] Samara M. Haver, Jeffrey D. Adams, Leila T. Hatch, Sofie M. Van Parijs, Robert P. Dziak, Joseph Haxel, Scott A. Heppell, Megan F. McKenna, David K. Mellinger, and Jason Gedamke. Large vessel activity and lowfrequency underwater sound benchmarks in United States waters. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Nayak:2021:ESS

[2162] Aditya R. Nayak, Houshuo Jiang, Margaret L. Byron, James M. Sullivan, Malcolm N. McFarland, and David W. Murphy. Editorial: Small scale spatial and temporal patterns in particles, plankton, and other organisms. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745.

Martin:2021:HWM

[2163] Saskia C. Martin, Ana S. Aniceto, Heidi Ahonen, Geir Pedersen, and Ulf Lindstrøm. Humpback whale (Megaptera novaeangliae) song on a subarctic feeding ground. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lafage:2021:SCM

[2164] Denis Lafage, Alexandre Carpentier, Sylvain Duhamel, Christine Dupuy, Eric Feunteun, Sandric Lesourd, and Julien Pétillon. Site characteristics more than vegetation type influence food web structure of intertidal salt marshes. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Herrera:2021:DCC

[2165] Inma Herrera, Manuel Carrillo, Marcial Cosme de Esteban, and Ricardo Haroun. Distribution of cetaceans in the geonameCanary Islands (Northeast Atlantic Ocean): Implications for the Natura 2000 network and future conservation measures. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Han:2021:PNA

[2166] Yuru Han, Dongsheng Zhang, Chunsheng Wang, and Yadong Zhou. Out of the Pacific: a new alvinellid worm (Annelida: Terebellida) from the Northern Indian Ocean hydrothermal vents. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Meszaros:2021:CCI

[2167] Lőrinc Mészáros, Frank van der Meulen, Geurt Jongbloed, and Ghada El Serafy. Climate change induced trends and uncertainties in phytoplankton spring bloom dynamics. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

VanWynen:2021:SHU

[2168] Cassie M. VanWynen, Morgan V. Hightshoe, Nicole D. Fogarty, Craig P. Dahlgren, and David S. Gilliam. Should hybrids be used in coral nurseries? A case study comparing Caribbean Acropora spp. and their hybrid in The Bahamas. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

ONeil:2021:RES

[2169] Keri L. O’Neil, Rachel M. Serafin, Joshua T. Patterson, and Jamie R. K. Craggs. Repeated ex situ spawning in two highly disease susceptible corals in the family Meandrinidae. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Humanes:2021:EFS

[2170] Adriana Humanes, Elizabeth A. Beauchamp, John C. Bythell, Mitch K. Carl, Jamie R. Craggs, Alasdair J. Edwards, Yimnang Golbuu, Liam Lachs, Helios M. Martinez, Pawel Palmowski, Faith Paysinger, Janna L. Randle, Eveline van der Steeg, Michael Sweet, Achim Treumann, and James R. Guest. An experimental framework for selectively breeding corals for assisted evolution. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Melville-Rea:2021:RPR

[2171] Hannah Melville-Rea, Clare Eayrs, Nasser Anwahi, John A. Burt, Denise Holland, Fatin Samara, Francesco Paparella, Ahmed Hassan Al Murshidi, Maryam Rashed Al-Shehhi, and David M. Holland. A roadmap for policy-relevant sea-level rise research in the United Arab Emirates. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lin:2021:ESC

[2172] Wei-Jen Lin, Ming-Chih Chiu, Chiao-Wen Lin, and Hsing-Juh Lin. Effects of sediment characteristics on carbon dioxide fluxes based on interacting factors in unvegetated tidal flats. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Branoff:2021:RUV

[2173] Benjamin L. Branoff and Marconi Campos-Cerqueira. The role of urbanness, vegetation structure, and scale in shaping Puerto Rico’s acoustically active mangrove fauna communities. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Blanco-Libreros:2021:TMA

[2174] Juan F. Blanco-Libreros and Karla Ramírez-Ruiz. Threatened mangroves in the Anthropocene: Habitat fragmentation in urban coastalscapes of Pelliciera spp. (Tetrameristaceae) in Northern South America. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ulman:2021:LPL

[2175] Aylin Ulman, Holden E. Harris, Nikos Doumpas, Hasan Deniz Akbora, Sara A. A Al Mabruk, Ernesto Azzurro, Michel Bariche, Burak Ali Çiçek, Alan Deidun, Nazli Demirel, Alexander Q. Fogg, Stelios Katsavenakis, Demetris Kletou, Periklis Kleitou, Athina Papadopoulou, Jamila Ben Souissi, Jason M. Hall-Spencer, Francesco Tiralongo, and Taner Yildiz. Low pufferfish and lionfish predation in their native and invaded ranges suggests human control mechanisms may be necessary to control their Mediterranean abundances. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Maneval:2021:GND

[2176] Paul Maneval, Charles A. Jacoby, Holden E. Harris, and Thomas K. Frazer. Genotype, nursery design, and depth influence the growth of Acropora cervicornis fragments. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wang:2021:CPA

[2177] Zhongqiu Wang, Hao Tang, Bent Herrmann, and Liuxiong Xu. Catch pattern for Antarctic krill (Euphausia superba) of different commercial trawls in similar times and overlapping fishing grounds. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Grzelak:2021:HMD

[2178] Katarzyna Grzelak, Daniela Zeppilli, Mauricio Shimabukuro, and Martin V. Sørensen. Hadal mud dragons: First insight into the diversity of kinorhyncha from the Atacama Trench. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ashlock:2021:TPO

[2179] Lauren Ashlock, Marisol García-Reyes, Chelle Gentemann, Sonia Batten, and William Sydeman. Temperature and patterns of occurrence and abundance of key copepod taxa in the Northeast Pacific. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Meiling:2021:VSR

[2180] Sonora S. Meiling, Erinn M. Muller, Danielle Lasseigne, Ashley Rossin, Alex J. Veglia, Nicholas MacKnight, Bradford Dimos, Naomi Huntley, Adrienne M. S. Correa, Tyler Burton Smith, Daniel M. Holstein, Laura D. Mydlarz, Amy Apprill, and Marilyn E. Brandt. Variable species responses to experimental Stony Coral Tissue Loss Disease (SCTLD) exposure. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Xu:2021:NIC

[2181] Tingting Xu, Yingxia Li, Xiujuan Shan, Jingwei Hao, Qiang Wu, Kathy F. J. Tang, Qingli Zhang, and Cuiluan Yao. Natural infection of covert mortality nodavirus in small yellow croaker in coastal water. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cavada-Blanco:2021:FEB

[2182] Francoise Cavada-Blanco, Aldo Cróquer, Edgard Yerena, and Jon P. Rodríguez. Flow of economic benefits from coral reefs in a multi-use Caribbean Marine Protected Area using network theory. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Lejzerowicz:2021:EBS

[2183] Franck Lejzerowicz, Andrew John Gooday, Inés Barrenechea Angeles, Tristan Cordier, Raphaël Morard, Laure Apothéloz-Perret-Gentil, Lidia Lins, Lenaick Menot, Angelika Brandt, Lisa Ann Levin, Pedro Martinez Arbizu, Craig Randall Smith, and Jan Pawlowski. Eukaryotic biodiversity and spatial patterns in the clarion-clipperton zone and other abyssal regions: Insights from sediment DNA and RNA metabarcoding. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Utama:2021:EFE

[2184] Febrianto W. Utama, Xavier Hoenner, Britta Denise Hardesty, David Peel, Jessica H. Ford, Vanessa Adams, and Chris Wilcox. Estimating fishing effort and spatio-temporal distribution of longline vessels in the Indian Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Asner:2021:CAH

[2185] Gregory P. Asner, Nicholas R. Vaughn, Shawna A. Foo, Ethan Shafron, Joseph Heckler, and Roberta E. Martin. Corrigendum: Abiotic and human drivers of reef habitat complexity throughout the main Hawaiian Islands. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See [1706].

Hanna:2021:CSS

[2186] Megan E. Hanna, Erin M. Chandler, Brice X. Semmens, Tomoharu Eguchi, Garrett E. Lemons, and Jeffrey A. Seminoff. Citizen-sourced sightings and underwater photography reveal novel insights about green sea turtle distribution and ecology in Southern California. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Durden:2021:MEW

[2187] Jennifer M. Durden, Meagan Putts, Sarah Bingo, Astrid B. Leitner, Jeffrey C. Drazen, Andrew J. Gooday, Daniel O. B. Jones, Andrew K. Sweetman, Travis W. Washburn, and Craig R. Smith. Megafaunal ecology of the Western Clarion Clipperton Zone. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Solano:2021:CUW

[2188] Neyra Solano, Ines Lopez-Ercilla, Francisco J. Fernandez-Rivera Melo, and Jorge Torre. Corrigendum: Unveiling women’s roles and inclusion in Mexican Small-Scale Fisheries (SSF). Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See [1333].

vonOppeln-Bronikowski:2021:ONO

[2189] Nicolai von Oppeln-Bronikowski, Mingxi Zhou, Taimaz Bahadory, and Brad de Young. Overview of a new ocean glider navigation system: OceanGNS. Frontiers in Marine Science, 8, November 2021. ISSN 2296-

Bingham:2021:KPF

[2190] Julia A. Bingham, Saul Milne, Grant Murray, and Terry Dorward. Knowledge pluralism in First Nations’ salmon management. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Truong:2021:SOS

[2191] Gary Truong and Tracey L. Rogers. Seasonal occurrence of sympatric blue whale subspecies: the Chilean and Southeast Indian Ocean pygmy blue whales with the Antarctic blue whale. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sukumaran:2021:MMN

[2192] Soniya Sukumaran, Tejal Vijapure, Jyoti Mulik, and Hurmine Ridha. Marine macrobenthos of NorthWest India — reviewing the known and unknown. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Painting:2021:MWQ

[2193] Suzanne J. Painting, Paul Nelson, Andy J. Smith, Carolyn A. Graves, Andy Powell, Philippe Bersuder, Samuel Bullen, and Simeon Archer-Rand. Marine water quality at Diego Garcia: a preliminary study of pollution levels in coastal and lagoon waters. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Stoever:2021:DOP

[2194] Jana Stoever, Katrin Rehdanz, and Ikechukwu Charles Okoli. A drop in the ocean: Patchy evidence on the societal benefits of marine litter reduction. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Dahihande:2021:DSC

[2195] Azraj S. Dahihande and Narsinh L. Thakur. Differences in the structural components influence the pumping capacity of marine sponges. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ghosh:2021:DES

[2196] Moumita Ghosh and Sumit Mandal. Disentangling the effect of seasonal dynamics on meiobenthic community structure from River Matla of Sundarbans Estuarine System, India. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Saprykina:2021:WCV

[2197] Yana V. Saprykina, S. V. Samiksha, and Sergey Yu. Kuznetsov. Wave climate variability and occurrence of mudbanks along the Southwest Coast of India. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Vilhena:2021:BST

[2198] Jéssica C. E. Vilhena, Ana Amorim, Lourenço Ribeiro, Bernardo Duarte, and Maíra Pombo. Baseline study of trace element concentrations in sediments of the intertidal zone of Amazonian oceanic beaches. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Sheehan:2021:RPA

[2199] E. V. Sheehan, L. A. Holmes, B. F. R. Davies, A. Cartwright, A. Rees, and M. J. Attrill. Rewilding of protected areas enhances resilience of marine ecosystems to extreme climatic events. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Sautya:2021:DPB

[2200] Sabyasachi Sautya, Santosh Gaikwad, Sanofar Khokher, Umesh Kumar Pradhan, Soumya Chatterjee, Amita Choudhury, Balaram Sahu, and Suman Attri. Distribution pattern of the benthic meiofaunal community along the depth gradient of the Western Indian Continental Margin, including the OMZ and abyssal plain. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Menna:2021:CTP

[2201] Milena Menna, Riccardo Gerin, Giulio Notarstefano, Elena Mauri, Antonio Bussani, Massimo Pacciaroni, and Pierre-Marie Poulain. On the circulation and thermohaline properties of the Eastern Mediterranean Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Binder:2021:FSA

[2202] Benjamin M. Binder, J. Christopher Taylor, Kurtis Gregg, and Kevin M. Boswell. Fish spawning aggregations in the Southeast Florida coral reef ecosystem conservation area: a case study synthesis of user reports, literature, and field validation efforts. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Bednarsek:2021:IAO

[2203] Nina Bednaršek, Kerry-Ann Naish, Richard A. Feely, Claudine Hauri, Katsunori Kimoto, Albert J. Hermann, Christine Michel, Andrea Niemi, and Darren Pilcher. Integrated assessment of ocean acidification risks to pteropods in the northern high latitudes: Regional comparison of exposure, sensitivity and adaptive capacity. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Jeronimo:2021:OTP

[2204] Daniel Jerónimo, Ana Isabel Lillebø, Felisa Rey, Henrique Koga II, M. Rosário M. Domingues, and Ricardo Calado. Optimizing the timeframe to produce polychaetes (Hediste diversicolor) enriched with essential fatty acids under different combinations of temperature and salinity. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Liu:2021:IPV

[2205] Mingming Liu, Mingli Lin, David Lusseau, and Songhai Li. Intrapopulation variability in group size of Indo–Pacific humpback dolphins (Sousa chinensis). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Quiros:2021:BCE

[2206] T. E. Angela L. Quiros, Kenji Sudo, Reynante V. Ramilo, Helbert G. Garay, Muammar Princess G. Soniega, Alvin Baloloy, Ariel Blanco, Ayin Tamondong, Kazuo Nadaoka, and Masahiro Nakaoka. Blue carbon ecosystem services through a vulnerability lens: Opportunities to reduce social vulnerability in fishing communities. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Yu:2021:GSM

[2207] Fangjie Yu, Zhiyuan Zhuang, Jie Yang, and Ge Chen. A glider simulation model based on optimized support vector regression for efficient coordinated observation. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Barus:2021:HMA

[2208] Beta Susanto Barus, Kai Chen, Minggang Cai, Rongmao Li, Huorong Chen, Cong Li, Jun Wang, and Sha-Yen Cheng. Heavy metal adsorption and release on polystyrene particles at various salinities. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Blasiak:2021:EPS

[2209] Robert Blasiak, Alice Dauriach, Jean-Baptiste Jouffray, Carl Folke, Henrik Österblom, Jan Bebbington, Frida Bengtsson, Amar Causevic, Bas Geerts, Wenche Grønbrekk, Patrik J. G. Henriksson, Sofia Käll, Duncan Leadbitter, Darian McBain, Guillermo Ortuño Crespo, Helen Packer, Isao Sakaguchi, Lisen Schultz, Elizabeth R. Selig, Max Troell, José Villalón, Colette C. C. Wabnitz, Emmy Wassénius, Reg A. Watson, Nobuyuki Yagi, and Beatrice Crona. Evolving perspectives of stewardship in the seafood industry. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Pizarro:2021:EET

[2210] Oscar Pizarro and Leonard Pace. Editorial: Emerging technologies with high impact for ocean sciences, ecosystem management, and environmental conservation. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Zhang:2021:FCM

[2211] Qianchu Zhang and Boris Katsnelson. A fish chorus on the margin of New Jersey Atlantic Continental Shelf. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wang:2021:PVT

[2212] Jianing Wang, Fan Wang, Youyu Lu, Qiang Ma, Larry J. Pratt, and Zhixiang Zhang. Pathways, volume transport, and seasonal variability of the lower deep limb of the Pacific meridional overturning circulation at the Yap–Mariana Junction. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Mihanovic:2021:OPR

[2213] Hrvoje Mihanović, Ivica Vilibić, Jadranka Šepić, Frano Matić, Zrinka Ljubešić, Elena Mauri, Riccardo Gerin, Giulio Notarstefano, and Pierre-Marie Poulain. Observation, preconditioning and recurrence of exceptionally high salinities in the Adriatic Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Amenyogbe:2021:PPI

[2214] Eric Amenyogbe, Jian sheng Huang, Gang Chen, and Wei zheng Wang. Probiotic potential of indigenous (Bacillus sp. RCS1, Pantoea agglomerans RCS2, and Bacillus cereus strain RCS3) isolated from cobia fish (Rachycentron canadum) and their antagonistic effects on the growth of pathogenic Vibrio alginolyticus, Vibrio harveyi, Streptococcus iniae, and Streptococcus agalactiae. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Ning:2021:SMT

[2215] Pengfei Ning, Cuicui Zhang, Xuefeng Zhang, and Xiaoyi Jiang. Shortto medium-term sea surface height prediction in the Bohai Sea using an optimized simple recurrent unit deep network. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Xu:2021:ATD

[2216] Guangjun Xu, Wenhong Xie, Changming Dong, and Xiaoqian Gao. Application of three deep learning schemes into oceanic eddy detection. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Wilford:2021:QSA

[2217] Dylan Charles Wilford, Jennifer L. Miksis-Olds, S. Bruce Martin, Daniel R. Howard, Kim Lowell, Anthony P. Lyons, and Michael James Smith. Quantitative soundscape analysis to understand multidimensional features. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Durante:2021:MTI

[2218] Sara Durante, Paolo Oliveri, Rajesh Nair, and Stefania Sparnocchia. Mixing in the Tyrrhenian interior due to thermohaline staircases. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sinha:2021:EOS

[2219] Anirban Sinha and Ryan Abernathey. Estimating ocean surface currents with machine learning. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Banister:2021:RFR

[2220] Raymond B. Banister and Robert van Woesik. Ranking 67 Florida reefs for survival of Acropora cervicornis outplants. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Muller:2021:BIS

[2221] Carolin Müller. Not as bad as it seems? A literature review on the case of microplastic uptake in fish. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Nestorowicz:2021:IHC

[2222] Iga-Maria Nestorowicz, Frederico Oliveira, Pedro Monteiro, Luís Bentes, Nuno Sales Henriques, Ricardo Aguilar, Barbara Horta e Costa, and Jorge M. S. Gonçalves. Identifying habitats of conservation priority in the São Vicente Submarine Canyon in Southwestern Portugal. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Magozzi:2021:CSS

[2223] Sarah Magozzi, Simon R. Thorrold, Leah Houghton, Victoria A. Bendall, Stuart Hetherington, Gonzalo Mucientes, Lisa J. Natanson, Nuno Queiroz, Miguel N. Santos, and Clive N. Trueman. Compound-specific stable isotope analysis of amino acids in pelagic shark vertebrae reveals baseline, trophic, and physiological effects on bulk protein isotope records. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Baker:2021:RMA

[2224] Dana Baker, Grant Murray, Jackson Kaijage, Arielle Levine, David Gill, and Enock Makupa. Relationships matter: Assessing the impacts of a Marine Protected Area on human wellbeing and relational values in Southern Tanzania. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Principe:2021:PSD

[2225] Silas C. Principe, André L. Acosta, João E. Andrade, and Tito M. C. Lotufo. Predicted shifts in the distributions of Atlantic reef-building corals in the face of climate change. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Atwood:2021:CGP

[2226] Trisha B. Atwood, Andrew Witt, Juan Mayorga, Edd Hammill, and Enric Sala. Corrigendum: Global patterns in marine sediment carbon stocks. Frontiers in Marine Science, 8, March 2021. ISSN 2296-7745. See [267].

Rodriguez-Ramos:2021:VNP

[2227] Tamara Rodríguez-Ramos, Mar Nieto-Cid, Adrià Auladell, Elisa Guerrero-Feijóo, and Marta M. Varela. Vertical niche partitioning of archaea and bacteria linked to shifts in dissolved organic matter quality and hydrography in North Atlantic waters. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Hunnam:2021:UTT

[2228] Kimberley Hunnam, Imelda Carlos, Michael P. Hammer, Joctan Dos Reis Lopes, David J. Mills, and Natasha Stacey. Untangling tales of tropical sardines: Local knowledge from fisheries in Timor-Leste. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Dunn:2021:SSS

[2229] Charlotte Dunn, James Theriault, Leigh Hickmott, and Diane Claridge. Slower ship speed in The Bahamas due to COVID-19 produces a dramatic reduction in ocean sound levels. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Delroisse:2021:LDS

[2230] Jérôme Delroisse, Laurent Duchatelet, Patrick Flammang, and Jérôme Mallefet. Leaving the dark side? Insights into the evolution of luciferases. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bosse:2021:MIC

[2231] Karl R. Bosse, Michael J. Sayers, Robert A. Shuchman, John Lekki, and Roger Tokars. Measuring the impact of the COVID-19 shutdown on Great Lakes water quality using remote sensing. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kotta:2021:SSP

[2232] Jonne Kotta, Kaire Torn, Tiina Paalme, Merli Rätsep, Kaire Kaljurand, Martin Teeveer, and Ilmar Kotta. Scale-specific patterns of the production of the charophyte Chara aspera in the brackish Baltic Sea: Linking individual and community production and biomass growth. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Aavaste:2021:ITC

[2233] Age Aavaste, Liis Sipelgas, Rivo Uiboupin, and Kristi Uudeberg. Impact of thermohaline conditions on vertical variability of optical properties in the Gulf of Finland (Baltic Sea): Implications for water quality remote sensing. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Quintana-Rizzo:2021:SDA

[2234] Ester Quintana-Rizzo, Andrea A. Cabrera, Jenniffer Ortiz-Wolford, and Vanessa Dávila. Spatial distribution and abundance of small cetaceans in the Pacific waters of Guatemala. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bouchet:2021:ARS

[2235] Phil J. Bouchet, Catriona M. Harris, and Len Thomas. Assessing the role of sampling uncertainty when predicting behavioral responses of tagged cetaceans exposed to naval sonar. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Methratta:2021:DBS

[2236] Elizabeth T. Methratta. Distance-based sampling methods for assessing the ecological effects of offshore wind farms: Synthesis and application to fisheries resource studies. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Salauddin:2021:EES

[2237] Md Salauddin, John J. O’Sullivan, Soroush Abolfathi, and Jonathan M. Pearson. Eco-engineering of seawalls — an opportunity for enhanced climate resilience from increased topographic complexity. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Fiorentino:2021:HCW

[2238] Fabio Fiorentino and Sergio Vitale. How can we reduce the overexploitation of the Mediterranean resources? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Toledo-Guedes:2021:PCS

[2239] Kilian Toledo-Guedes, Alberto Brito, and Carlos Garcia de Leaniz. Phenotypic convergence in sea bass (Dicentrarchus labrax) escaping from fish farms: The onset of feralization? Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

DeClippele:2021:MSC

[2240] Laurence H. De Clippele and Denise Risch. Measuring sound at a coldwater coral reef to assess the impact of COVID-19 on noise pollution. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Moland:2021:RAD

[2241] Even Moland, Albert Fernández-Chacón, Tonje Knutsen Sørdalen, David Villegas-Ríos, Susanna Huneide Thorbjørnsen, Kim Tallaksen Halvorsen, Mats Huserbråten, Esben Moland Olsen, Portia Joy Nillos Kleiven, Alf Ring Kleiven, Halvor Knutsen, Sigurd Heiberg Espeland, Carla Freitas, and Jan Atle Knutsen. Restoration of abundance and dynamics of coastal fish and lobster within Northern Marine Protected Areas across two decades. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gabriele:2021:USL

[2242] Christine M. Gabriele, Dimitri W. Ponirakis, and Holger Klinck. Underwater sound levels in Glacier Bay during reduced vessel traffic due to the COVID-19 pandemic. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Breeze:2021:CUN

[2243] Heather Breeze, Shihan Li, Emmaline C. Marotte, James A. Theriault, Jessica Wingfield, and Jinshan Xu. Changes in underwater noise and vessel traffic in the approaches to Halifax Harbor, Nova Scotia, Canada. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Enright:2021:ETP

[2244] Sarah Ryan Enright, Ricardo Meneses-Orellana, and Inti Keith. The Eastern Tropical Pacific Marine Corridor (CMAR): The emergence of a voluntary regional cooperation mechanism for the conservation and sustainable use of marine biodiversity within a fragmented regional ocean governance landscape. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ransome:2021:VSL

[2245] Nicola Ransome, Neil R. Loneragan, Luis Medrano-González, Fernando Félix, and Joshua N. Smith. Vessel strikes of large whales in the eastern tropical Pacific: a case study of regional underreporting. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cramer:2021:RCH

[2246] Katie L. Cramer and John N. Kittinger. Reef conservation off the hook: Can market interventions make coral reef fisheries more sustainable? Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Boaventura:2021:POL

[2247] Diana Boaventura, Ana Teresa Neves, Jaime Santos, Paula Colares Pereira, Cristina Luís, António Monteiro, Alexandra Cartaxana, Stephen John Hawkins, Maria Filomena Caldeira, and António Ponces de Carvalho. Promoting ocean literacy in elementary school students through investigation activities and citizen science. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Felix:2021:MDC

[2248] Pedro M. Félix, Ana Pombo, Francisco Azevedo e Silva, Tomás Simões, Tiago A. Marques, Ricardo Melo, Cristina Rocha, João Sousa, Eliana Venâncio, José L. Costa, and Ana C. Brito. Modelling the distribution of a commercial NE-Atlantic sea cucumber, Holothuria mammata: Demographic and abundance spatio-temporal patterns. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chattopadhyay:2021:DPM

[2249] Devapriya Chattopadhyay, Deepjay Sarkar, and Madhura Bhattacherjee. The distribution pattern of marine bivalve death assemblage from the western margin of Bay of Bengal and its oceanographic determinants. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Neely:2021:SLT

[2250] Karen L. Neely, Colin P. Shea, Kevin A. Macaulay, Emily K. Hower, and Michelle A. Dobler. Short- and long-term effectiveness of coral disease treatments. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Anugerahanti:2021:EOB

[2251] Prima Anugerahanti, Onur Kerimoglu, and S. Lan Smith. Enhancing ocean biogeochemical models with phytoplankton variable composition. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Park:2021:NME

[2252] Yeongbin Park, Chanhyung Jeon, Hajin Song, Youngseok Choi, Jeong-Yeob Chae, Eun-Joo Lee, Jin Sung Kim, and Jae-Hun Park. Novel method for the estimation of vertical temperature profiles using a coastal acoustic tomography system. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chang:2021:DSR

[2253] Ssu-Wei Chang, Ruei-Gu Chen, Tsung-Han Liu, Yao-Chang Lee, Chih-Shin Chen, Tai-Sheng Chiu, and Chia-Ying Ko. Dietary shifts and risks of artifact ingestion for Argentine shortfin squid Illex argentinus in the Southwest Atlantic. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Turicchia:2021:RCMb

[2254] Eva Turicchia, Massimo Ponti, Gianfranco Rossi, and Carlo Cerrano. The Reef Check Med Dataset on key Mediterranean marine species 2001– 2020. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Cereja:2021:TPE

[2255] Rui Cereja, Vanda Brotas, Joana P. C. Cruz, Marta Rodrigues, and Ana C. Brito. Tidal and physicochemical effects on phytoplankton community variability at Tagus Estuary (Portugal). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Guastella:2021:RBD

[2256] Roberta Guastella, Agnese Marchini, Antonio Caruso, Julian Evans, Miriam Cobianchi, Claudia Cosentino, Leonardo Langone, Rita Lecci, and Nicoletta Mancin. Reconstructing bioinvasion dynamics through micropaleontologic analysis highlights the role of temperature change as a driver of alien foraminifera invasion. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ribas-Deulofeu:2021:PGS

[2257] Lauriane Ribas-Deulofeu, Pierre-Alexandre Château, Vianney Denis, and Chaolun Allen Chen. Portraying gradients of structural complexity in coral reefs using fine-scale depth profiles. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

deMagalhaes:2021:MPG

[2258] Cláudia Raposo de Magalhães, Denise Schrama, Chatsirin Nakharuthai, Surintorn Boonanuntanasarn, Dominique Revets, Sébastien Planchon, Annette Kuehn, Marco Cerqueira, Raquel Carrilho, Ana Paula Farinha, and Pedro M. Rodrigues. Metabolic plasticity of gilthead seabream under different stressors: Analysis of the stress responsive hepatic proteome and gene expression. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Kriegl:2021:MPA

[2259] Michael Kriegl, Xochitl E. Elías Ilosvay, Christian von Dorrien, and Daniel Oesterwind. Marine Protected Areas: At the crossroads of nature conservation and fisheries management. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Sanborn:2021:ISS

[2260] Thomas Sanborn and Julia Jung. Intersecting social science and conservation. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Luque:2021:CFB

[2261] Pau Luque, Lluís Gómez-Pujol, Marta Marcos, and Alejandro Orfila. Coastal flooding in the Balearic Islands during the Twenty-First Century caused by sea-level rise and extreme events. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Jemai:2021:RAF

[2262] Ahlem Jemai, Jochen Wollschläger, Daniela Voß, and Oliver Zielinski. Radiometry on Argo Floats: From the multispectral state-of-the-art on the step to hyperspectral technology. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Nielsen:2021:NOI

[2263] Ole Nielsen, Thaís C. S. Rodrigues, Vsevolod L. Popov, Kuttichantran Subramaniam, Thomas B. Waltzek, Dyanna M. Lambourn, and Stephen Raverty. A novel orthoreovirus isolated from dead stranded harbor seals from Puget Sound, Washington State, United States. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Minich:2021:IMH

[2264] Jeremiah J. Minich, Barbara Nowak, Abigail Elizur, Rob Knight, Stewart Fielder, and Eric E. Allen. Impacts of the marine hatchery built environment, water and feed on mucosal microbiome colonization across ontogeny in yellowtail kingfish, Seriola lalandi. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Hills:2021:OOU

[2265] Jeremy M. Hills, Claire Lajaunie, Payal N. Maharaj, and Pierre Mazzega. Orchestrating the ocean using a national ocean policy: The case of the Solomon Islands. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ghandour:2021:MLH

[2266] Ibrahim M. Ghandour, Aaid G. Al-Zubieri, Ali S. Basaham, Ammar A. Mannaa, Talha A. Al-Dubai, and Brian G. Jones. Mid-Late Holocene paleoenvironmental and sea level reconstruction on the Al Lith Red Sea Coast, Saudi Arabia. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Perez-Botello:2021:IHC

[2267] Antar Mijail Pérez-Botello, Maite Mascaró, and Nuno Simões. The importance of home cleaning: Sediment transport by alpheid shrimps provides a competitive advantage to their host anemones. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Deng:2021:HPA

[2268] Zhengzhi Deng, Pinjie Wang, and Pengda Cheng. Hydrodynamic performance of an asymmetry OWC device mounted on a box-type breakwater. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Barreto:2021:PIA

[2269] Daniel Andres Lizarbe Barreto, Ricardo Chevarria Saravia, Takeyoshi Nagai, and Takafumi Hirata. Phytoplankton increase along the Kuroshio due to the large meander. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lipcius:2021:MOR

[2270] Romuald N. Lipcius, Yi Zhang, Jingyi Zhou, Leah B. Shaw, and Junping Shi. Modeling oyster reef restoration: Larval supply and reef geometry jointly determine population resilience and performance. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Rotem-Dai:2021:LTE

[2271] Noa Rotem-Dai, Simy Weil, Yariv Greenshpan, Shai Abehsera, Rivka Manor, Rivka Ofir, Roi Gazit, Benyamin Rosental, Eliahu D. Aflalo, and Amir Sagi. Lentiviral-transduced ectopic expression of androgenic hormone in a crustacean hematopoietic primary cell culture. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Descoteaux:2021:MDS

[2272] Raphaëlle Descôteaux, Elizaveta Ershova, Owen S. Wangensteen, Kim Præbel, Paul E. Renaud, Finlo Cottier, and Bodil A. Bluhm. Meroplankton diversity, seasonality and life-history traits across the Barents Sea polar front revealed by high-throughput DNA barcoding. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Bouveroux:2021:CCA

[2273] Thibaut Bouveroux, Nicolas Loiseau, Adam Barnett, Natasha D. Marosi, and Juerg M. Brunnschweiler. Companions and casual acquaintances: The nature of associations among bull sharks at a shark feeding site in Fiji. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Dudley:2021:MCC

[2274] Peter N. Dudley, Tanya L. Rogers, Mark M. Morales, Amanda D. Stoltz, Casey J. Sheridan, Anne K. Beulke, Carrie Pomeroy, and Mark H. Carr. A more comprehensive climate vulnerability assessment framework for fisheries social-ecological systems. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Syme:2021:DCM

[2275] Jonathan Syme, Jeremy J. Kiszka, and Guido J. Parra. Dynamics of cetacean mixed-species groups: a review and conceptual framework for assessing their functional significance. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Singh:2021:KAM

[2276] Charu Lata Singh, Megan J. Huggett, Paul S. Lavery, Christin Säwström, and Glenn A. Hyndes. Kelp-associated microbes facilitate spatial subsidy in a detrital-based food web in a shoreline ecosystem. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Esposito:2021:SVD

[2277] Giovanni Esposito, Maristella Berta, Luca Centurioni, T. M. Shaun Johnston, John Lodise, Tamay Özgökmen, Pierre-Marie Poulain, and Annalisa Griffa. Submesoscale vorticity and divergence in the Alboran Sea: Scale and depth dependence. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Verwega:2021:PMD

[2278] Maria-Theresia Verwega, Carola Trahms, Avan N. Antia, Thorsten Dickhaus, Enno Prigge, Martin H. U. Prinzler, Matthias Renz, Markus Schartau, Thomas Slawig, Christopher J. Somes, and Arne Biastoch. Perspectives on marine data science as a blueprint for emerging data science disciplines. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wassmann:2021:ETU

[2279] Paul Wassmann, Dorte Krause-Jensen, Bodil A. Bluhm, and Markus Janout. Editorial: Towards a unifying pan-Arctic perspective of the contemporary and future Arctic Ocean. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Guarnaccia:2021:DCB

[2280] Andrew M. Guarnaccia, Sara Rose Krivoshik, John S. Sparks, David F. Gruber, and Jean P. Gaffney. Discovery and characterization of a bilirubin inducible green fluorescent protein from the moray eel Gymnothorax zonipectis. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Tuo:2021:TLH

[2281] Sing how Tuo, Margaret R. Mulholland, Yukiko Taniuchi, Houng-Yung Chen, Wann-Neng Jane, Yen-Huei Lin, and Yuh ling Lee Chen. Trichome lengths of the heterocystous N2-Fixing cyanobacteria in the tropical marginal seas of the Western North Pacific. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Relva:2021:TEA

[2282] Julieta Vigliano Relva and Julia Jung. Through the eyes of another: Using a narrative lens to navigate complex social-ecological systems and to embrace multiple ways of knowing. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hsieh:2021:EHT

[2283] Hsueh-Han Hsieh, Ming-Hsiu Chuang, Yung-Yen Shih, W. Sanjaya Weerakkody, Wei-Jen Huang, Chin-Chang Hung, François L. L. Muller, R. R. M. K. P. Ranatunga, and D. S. Wijethunga. Eutrophication and hypoxia in tropical Negombo Lagoon, Sri Lanka. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Xue:2021:IST

[2284] Chengfeng Xue, Yang Yang, Peipei Zhao, Dongyun Wei, Jianhua Gao, Peng Sun, Zhiyang Huang, and Jianjun Jia. Impact of ship traffic on the characteristics of shelf sediments: an Anthropocene prospective. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Goldberg:2021:PCL

[2285] Daphna Shapiro Goldberg, Gil Rilov, Sébastien Villéger, and Jonathan Belmaker. Predation cues lead to reduced foraging of invasive Siganus rivulatus in the Mediterranean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Adyasari:2021:SDF

[2286] Dini Adyasari, Daniel Montiel, Behzad Mortazavi, and Natasha Dimova. Storm-driven fresh submarine groundwater discharge and nutrient fluxes from a barrier island. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wang:2021:ULM

[2287] Yuanchao Wang, Cui Liang, Weiwei Xian, and Yibang Wang. Using the LBB method for the assessments of seven fish stocks from the Yangtze Estuary and its adjacent waters. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ko:2021:VSP

[2288] Young Ho Ko, Geun-Ha Park, Dongseon Kim, and Tae-Wook Kim. Variations in seawater pCO_{2} associated with vertical mixing during tropical cyclone season in the northwestern subtropical Pacific Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zavialov:2021:POS

[2289] Peter O. Zavialov. Are periodic oscillations of sea surface height inherent to river plumes? Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Jeon:2021:ARD

[2290] Chanhyung Jeon, Jae-Hun Park, Maureen Kennelly, Erran Sousa, D. Randolph Watts, Eun-Joo Lee, Taewook Park, and Thomas Peacock. Advanced remote data acquisition using a Pop-Up Data Shuttle (PDS) to report data from Current- and Pressure-Recording Inverted Echo Sounders (CPIES). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Tew:2021:ISB

[2291] Kwee Siong Tew, Jimmy Kuo, Jing-O Cheng, Fung-Chi Ko, Pei-Jie Meng, Anderson B. Mayfield, and Pi-Jen Liu. Impacts of seagrass on benthic microalgae and phytoplankton communities in an experimentally warmed coral reef mesocosm. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Farhadi:2021:RMS

[2292] Ardavan Farhadi, Wenxiao Cui, Huaiping Zheng, Shengkang Li, Yueling Zhang, Mhd Ikhwanuddin, and Hongyu Ma. The regulatory mechanism of sexual development in decapod crustaceans. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745.

Whalan:2021:DPO

[2293] Steve Whalan, Marji Puotinen, Mary Wakeford, Iain Parnum, and Karen Miller. Distribution of the pearl oyster Pinctada maxima off Eighty Mile Beach, Western Australia. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Barcelo-Llull:2021:FSO

[2294] Bàrbara Barceló-Llull, Ananda Pascual, Antonio Sánchez-Román, Eugenio Cutolo, Francesco d’Ovidio, Gina Fifani, Enrico Ser-Giacomi, Simón Ruiz, Evan Mason, Fréderic Cyr, Andrea Doglioli, Baptiste Mourre, John T. Allen, Eva Alou-Font, Benjamín Casas, Lara Díaz-Barroso, Franck Dumas, Laura Gómez-Navarro, and Cristian Muñoz. Fine-scale ocean currents derived from in situ observations in anticipation of the upcoming SWOT altimetric mission. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Tiedemann:2021:DIC

[2295] Maik Tiedemann, Aril Slotte, Richard D. M. Nash, Erling K. Stenevik, and Olav S. Kjesbu. Drift indices confirm that rapid larval displacement is essential for recruitment success in high-latitude oceans. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Rosenau:2021:IHR

[2296] Nicholas A. Rosenau, Holly Galavotti, Kimberly K. Yates, Curtis C. Bohlen, Christopher W. Hunt, Matthew Liebman, Cheryl A. Brown, Stephen R. Pacella, John L. Largier, Karina J. Nielsen, Xinping Hu, Melissa R. McCutcheon, James M. Vasslides, Matthew Poach, Tom Ford, Karina Johnston, and Alex Steele. Integrating high-resolution coastal acidification monitoring data across seven United States estuaries. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Mejri:2021:EFA

[2297] Sahar Chauffour Mejri, Réjean Tremblay, Céline Audet, Paul S. Wills, and Marty Riche. Essential fatty acid requirements in tropical and coldwater marine fish larvae and juveniles. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Mirasbekov:2021:CIF

[2298] Yersultan Mirasbekov, Aigerim Abdimanova, Kuanysh Sarkytbayev, Kanat Samarkhanov, Aidyn Abilkas, Daria Potashnikova, Galina Arbuz, Zhanpeis Issayev, Ivan A. Vorobjev, Dmitry V. Malashenkov, and Natasha S. Barteneva. Combining imaging flow cytometry and molecular biological methods to reveal presence of potentially toxic algae at the Ural River in Kazakhstan. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Behan:2021:ESD

[2299] Jamie Behan, Bai Li, and Yong Chen. Examining scale dependent environmental effects on American lobster (Homarus americanus) spatial distribution in a changing Gulf of Maine. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Michael:2021:RFN

[2300] Sarah A. Michael, David T. S. Hayman, Rachael Gray, and Wendi D. Roe. Risk factors for New Zealand sea lion (Phocarctos hookeri) pup mortality: Ivermectin improves survival for conservation management. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Acabado:2021:VNF

[2301] Cristy S. Acabado, Yu-Hsin Cheng, Ming-Huei Chang, and Chung-Chi Chen. Vertical nitrate flux induced by Kelvin–Helmholtz billows over a seamount in the Kuroshio. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Miccoli:2021:TCE

[2302] Andrea Miccoli, Emanuele Mancini, Marco Boschi, Francesca Provenza, Veronica Lelli, Francesco Tiralongo, Monia Renzi, Antonio Terlizzi, Simone Bonamano, and Marco Marcelli. Trophic, chemo-ecological and sex-specific insights on the relation between Diplodus sargus (Linnaeus, 1758) and the invasive Caulerpa cylindracea (Sonder, 1845). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

McPherson:2021:MLS

[2303] Rebecca A. McPherson, Craig L. Stevens, Joanne M. O’Callaghan, Andrew J. Lucas, and Jonathan D. Nash. Mechanisms of lateral spreading in a near-field buoyant river plume entering a fjord. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Tran:2021:HDP

[2304] Hung Quang Tran, Markéta Prokešová, Mahyar Zare, Tatyana Gebauer, Antonia Concetta Elia, Elena Colombino, Ilario Ferrocino, Christian Caimi, Francesco Gai, Laura Gasco, and Vlastimil Stejskal. How does pikeperch Sander lucioperca respond to dietary insect meal Hermetia illucens? Investigation on gut microbiota, histomorphology, and antioxidant biomarkers. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Poupard:2021:IGO

[2305] Marion Poupard, Helena Symonds, Paul Spong, and Hervé Glotin. Intragroup orca call rate modulation estimation using compact four hydrophones array. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Montalbetti:2021:MBH

[2306] Enrico Montalbetti, Tom Biscéré, Christine Ferrier-Pagès, Fanny Houlbrèque, Ivan Orlandi, Matilde Forcella, Paolo Galli, Marina Vai, and Davide Seveso. Manganese benefits heat-stressed corals at the cellular level. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ok:2021:PBD

[2307] Jin Hee Ok, Hae Jin Jeong, Ji Hyun You, Hee Chang Kang, Sang Ah Park, An Suk Lim, Sung Yeon Lee, and Se Hee Eom. Phytoplankton bloom dynamics in incubated natural seawater: Predicting bloom magnitude and timing. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Park:2021:SCM

[2308] Hyoungchul Park and Jin Hwan Hwang. A standard criterion for measuring turbulence quantities using the four-receiver acoustic Doppler velocimetry. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Guerreiro:2021:BGC

[2309] José Guerreiro. The blue growth challenge to maritime governance. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Armes:2021:CGI

[2310] April C. Armes and Alison Buchan. Cyclic di-GMP is integrated into a hierarchal quorum sensing network regulating antimicrobial production and biofilm formation in Roseobacter clade member Rhodobacterales strain Y4I. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Traylor-Knowles:2021:GER

[2311] Nikki Traylor-Knowles, Michael T. Connelly, Benjamin D. Young, Katherine Eaton, Erinn M. Muller, Valerie J. Paul, Blake Ushijima, Allyson DeMerlis, Melissa K. Drown, Ashley Goncalves, Nicholas Kron, Grace A. Snyder, Cecily Martin, and Kevin Rodriguez. Gene expression response to Stony Coral Tissue Loss Disease transmission in M. cavernosa and O. faveolata from Florida. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Franz:2021:COO

[2312] Guilherme Franz, Carlos A. E. Garcia, Janini Pereira, Luiz Paulo de Freitas Assad, Marcelo Rollnic, Luis Hamilton P. Garbossa, Letícia Cotrim da Cunha, Carlos A. D. Lentini, Paulo Nobre, Alexander Turra, Janice R. Trotte-Duhá, Mauro Cirano, Segen F. Estefen, José Antonio M. Lima, Afonso M. Paiva, Mauricio A. Noernberg, Clemente A. S. Tanajura, José Luiz Moutinho, Francisco Campuzano, Ella S. Pereira, André Cunha Lima, Luís F. F. Mendonça, Helder Nocko, Leandro Machado, João B. R. Alvarenga, Renato P. Martins, Carina Stefoni Böck, Raquel Toste, Luiz Landau, Tiago Miranda, Francisco dos Santos, Júlio Pellegrini, Manuela Juliano, Ramiro Neves, and Andrei Polejack. Coastal ocean observing and modeling systems in Brazil: Initiatives and future perspectives. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Le:2021:IYD

[2313] Hong Minh Le, Karen Bekaert, Ruth Lagring, Bart Ampe, Ann Ruttens, Karien De Cauwer, Kris Hostens, and Bavo De Witte. 4DEMON: Integrating 40 years of data on PCB and metal contamination in marine sediments of the Belgian part of the North Sea. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Pang:2021:GSK

[2314] Ying Pang, Xiaojun Zhang, Jianbo Yuan, Xiaoxi Zhang, Manwen Su, and Fuhua Li. Glycogen synthase kinase 3 gene is important in growth and molting of the Pacific white shrimp Litopenaeus vannamei. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Kasten:2021:PMC

[2315] Paula Kasten, Stuart R. Jenkins, and Ronaldo A. Christofoletti. Participatory monitoring — a citizen science approach for coastal environments. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhang:2021:HFM

[2316] Zhipeng Zhang, Hongzhou Xu, Philip A. Vetter, Qiang Xie, Xiaohui Xie, Wei Song, and Chuan Tian. High-frequency motions in the Southeastern South China Sea during winter–spring 2018/2019. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Dahlgren:2021:STP

[2317] Craig Dahlgren, Valeria Pizarro, Krista Sherman, William Greene, and Joseph Oliver. Spatial and temporal patterns of Stony Coral Tissue Loss Disease outbreaks in The Bahamas. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Jimenez-Quiroz:2021:ECM

[2318] María del Carmen Jiménez-Quiroz, Francisco Javier Barrón-Barraza, Rafael Cervantes-Duarte, and René Funes-Rodríguez. Environmental considerations for the management of the bivalve fisheries of Bahía Magdalena (Mexico), a coastal lagoon at the southern end of the California Current. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kolodziej:2021:ISC

[2319] Graham Kolodziej, Michael S. Studivan, Arthur C. R. Gleason, Chris Langdon, Ian C. Enochs, and Derek P. Manzello. Impacts of Stony Coral Tissue Loss Disease (SCTLD) on coral community structure at an inshore patch reef of the Upper Florida Keys using photomosaics. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Philipp:2021:FER

[2320] Carolin Philipp, Bianca Unger, Sonja M. Ehlers, Jochen H. E. Koop, and Ursula Siebert. First evidence of retrospective findings of microplastics in harbour porpoises (Phocoena phocoena) from German waters. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Franceschini:2021:END

[2321] Simone Franceschini, Alessandro Cau, Lorenzo D’Andrea, Maria Cristina Follesa, and Tommaso Russo. Eating near the dump: Identification of nearby plastic hotspot as a proxy for potential microplastic contamination in the Norwegian lobster (Nephrops norvegicus). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Costa:2021:DDE

[2322] Sophia V. Costa, Stephanie J. Hibberts, Danielle A. Olive, Kayla A. Budd, Alexys E. Long, Sonora S. Meiling, Madyson B. Miller, Kelsey M. Vaughn, Claudia I. Carrión, Maksym B. Cohen, Annie E. Savage, Matthew F. Souza, Lorraine Buckley, Kristin W. Grimes, Renata Platenberg, Tyler B. Smith, Jeremiah Blondeau, and Marilyn E. Brandt. Diversity and disease: The effects of coral diversity on prevalence and impacts of Stony Coral Tissue Loss Disease in Saint Thomas, U.S. Virgin Islands. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Lin:2021:DMC

[2323] Xianzhi Lin, Simin Hu, Yong Liu, Li Zhang, Hui Huang, and Sheng Liu. Disturbance-mediated changes in coral reef habitat provoke a positive feeding response in a major coral reef detritivore, Ctenochaetus striatus. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Arrowsmith:2021:FIH

[2324] Lucy M. Arrowsmith, Charan Kumar Paidi, Farukhkha Husenkha Bloch, Sajan John, Binod Chandra Choudhury, Rahul Kaul, Ana M. M. Sequeira, Charitha B. Pattiaratchi, and Mark G. Meekan. First insights into the horizontal movements of whale sharks (Rhincodon typus) in the Northern Arabian Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Bonnet-Lebrun:2021:SMS

[2325] Anne-Sophie Bonnet-Lebrun, Maria P. Dias, Richard A. Phillips, José P. Granadeiro, M. de L. Brooke, Olivier Chastel, Thomas A. Clay, Annette L. Fayet, Olivier Gilg, Jacob González-Solís, Tim Guilford, Sveinn A. Hanssen, April Hedd, Audrey Jaeger, Johannes Krietsch, Johannes Lang, Matthieu Le Corre, Teresa Militão, Børge Moe, William A. Montevecchi, Hans-Ulrich Peter, Patrick Pinet, Matt J. Rayner, Tim Reid, José Manuel Reyes-González, Peter G. Ryan, Paul M. Sagar, Niels M. Schmidt, David R. Thompson, Rob van Bemmelen, Yutaka Watanuki, Henri Weimerskirch, Takashi Yamamoto, and Paulo Catry. Seabird migration strategies: Flight budgets, diel activity patterns, and lunar influence. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Maurer:2021:DMQ

[2326] Tanya L. Maurer, Joshua N. Plant, and Kenneth S. Johnson. Delayedmode quality control of oxygen, nitrate, and pH data on SOCCOM biogeochemical profiling floats. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Galparsoro:2021:DCE

[2327] Ibon Galparsoro and Ángel Borja. Defining cost-effective solutions in designing Marine Protected Areas, using systematic conservation planning. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Bradshaw:2021:PDB

[2328] Clare Bradshaw, Martin Jakobsson, Volker Brüchert, Stefano Bonaglia, Carl-Magnus Mörth, Julia Muchowski, Christian Stranne, and Mattias Sköld. Physical disturbance by bottom trawling suspends particulate matter and alters biogeochemical processes on and near the seafloor. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cabrera-Brufau:2021:DDE

[2329] Miguel Cabrera-Brufau, Laura Arin, Maria Montserrat Sala, Pedro Cermeño, and Cèlia Marrasé. Diatom dominance enhances resistance of phytoplanktonic POM to mesopelagic microbial decomposition. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bashir:2021:PCC

[2330] Saidu M. Bashir, Sam Kimiko, Chu-Wa Mak, James Kar-Hei Fang, and David Gonçalves. Personal care and cosmetic products as a potential source of environmental contamination by microplastics in a densely populated Asian city. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Reynes:2021:PGL

[2331] Lauric Reynes, Didier Aurelle, Cristele Chevalier, Christel Pinazo, Myriam Valero, Stéphane Mauger, Stéphane Sartoretto, Aurélie Blanfuné, Sandrine Ruitton, Charles-François Boudouresque, Marc Verlaque, and Thierry Thibaut. Population genomics and Lagrangian modeling shed light on dispersal events in the Mediterranean endemic Ericaria zosteroides (= Cystoseira zosteroides) (Fucales). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Toimil:2021:VUC

[2332] Alexandra Toimil, Paula Camus, Iñigo J. Losada, and Moises Alvarez-Cuesta. Visualising the uncertainty cascade in multi-ensemble probabilistic coastal erosion projections. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Hildebrand:2021:DGW

[2333] Lisa Hildebrand, Kim S. Bernard, and Leigh G. Torres. Do gray whales count calories? Comparing energetic values of gray whale prey across two different feeding grounds in the Eastern North Pacific. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Agus:2021:MBA

[2334] Blondine Agus, Laura Carugati, Andrea Bellodi, Rita Cannas, Alessandro Cau, Jacopo Cera, Elisabetta Coluccia, Riccardo Melis, Stefano Ruiu, and Danila Cuccu. Molecular and biological analysis on Ommastrephes caroli findings in the Central Western Mediterranean Sea (Sardinian Waters) including first age investigation using eye lenses and beaks. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Tubio:2021:SAR

[2335] Ana Tubío, Pablo Pita, Carlota Barañano, and Sebastián Villasante. A state-of-the-art review of marine ecosystem services in the Rías Baixas Natura 2000 Network (Galicia, NW Spain). Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Silva:2021:PAF

[2336] Priscila M. Silva, Célia M. Teixeira, Cristina Pita, Henrique N. Cabral, and Susana França. Portuguese artisanal fishers’ knowledge about elasmobranchs — a case study. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pratte:2021:CPE

[2337] Zoe A. Pratte and Christina A. Kellogg. Comparison of preservation and extraction methods on five taxonomically disparate coral microbiomes. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kitolelei:2021:VMN

[2338] Salanieta Kitolelei, Randy Thaman, Joeli Veitayaki, Annette Breckwoldt, and Susanna Piovano. Na Vuku Makawa ni Qoli: Indigenous Fishing Knowledge (IFK) in Fiji and the Pacific. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Li:2021:EDB

[2339] Yi-Yen Li, Feng-Jie Su, Yi-Jui Hsieh, Tai-Chun Huang, and Yung-Song Wang. Embryo development and behavior in sea urchin (Tripneustes gratilla) under different light emitting diodes condition. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Roca-Marti:2021:QDS

[2340] Montserrat Roca-Martí, Viena Puigcorbé, Maxi Castrillejo, Núria Casacuberta, Jordi Garcia-Orellana, J. Kirk Cochran, and Pere Masqué.

Quantifying ^{210}Po/^{210}Pb disequilibrium in seawater: a comparison of two precipitation methods with differing results. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Rekstad:2021:OHC

[2341] Matias Einer Rekstad, Sanna Majaneva, Åshild Løvas Borgersen, and Nicole Aberle. Occurrence and habitat characteristics of Aurelia sp. polyps in a high-latitude fjord. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Grosse:2021:SAA

[2342] Julia Grosse, Eva-Maria Nöthig, Sinhué Torres-Valdés, and Anja Engel. Summertime amino acid and carbohydrate patterns in particulate and dissolved organic carbon across Fram Strait. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Moraes:2021:CBP

[2343] Gabriela V. Moraes, Mary Colleen Hannon, Douglas M. M. Soares, Cassius V. Stevani, Anja Schulze, and Anderson G. Oliveira. Corrigendum: Bioluminescence in polynoid scale worms (Annelida: Polynoidae). Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745. See [1842].

Escudero:2021:CRG

[2344] Mireille Escudero, Borja G. Reguero, Edgar Mendoza, Fernando Secaira, and Rodolfo Silva. Coral reef geometry and hydrodynamics in beach erosion control in North Quintana Roo, Mexico. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pinto:2021:CPD

[2345] Bruno Pinto and Paula Castro. Contesting political decisions involving environmental issues: a case study in Portugal based on the press about offshore oil and gas drilling. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Cruz:2021:KSC

[2346] Pedro Manuel Carrasco De La Cruz. The knowledge status of coastal and marine ecosystem services - challenges, limitations and lessons learned from the application of the ecosystem services approach in management. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Freeling:2021:AMS

[2347] Benjamin S. Freeling and Sean D. Connell. Animal minds, social change, and the future of fisheries science. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Liu:2021:CMM

[2348] Jun Liu and Haibin Zhang. Combining multiple markers in environmental DNA metabarcoding to assess deep-sea benthic biodiversity. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Inaba:2021:PJJ

[2349] Nobuharu Inaba, Takuma Matsumoto, Hiroshi Kawai, Yuji Anaguchi, and Kohei Matsuno. Predation of juvenile Japanese sea cucumber Apostichopus japonicus by kelp crab Pugettia ferox. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Azevedo:2021:GSI

[2350] Leonardo Azevedo, Luís Matias, Francesco Turco, Renan Tromm, and Álvaro Peliz. Geostatistical seismic inversion for temperature and salinity in the Madeira Abyssal Plain. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Eirin-Lopez:2021:EME

[2351] Jose M. Eirin-Lopez and Hollie Putnam. Editorial: Marine environmental epigenetics. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Pajuelo:2021:CES

[2352] Eloísa Pajuelo, Sandra Arjona, Ignacio D. Rodríguez-Llorente, Enrique Mateos-Naranjo, Susana Redondo-Gómez, Francisco Merchán, and Salvadora Navarro-Torre. Coastal ecosystems as sources of biofertilizers in agriculture: From genomics to application in an urban orchard. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Mahon:2021:SSP

[2353] Robin Mahon and Lucia Fanning. Scoping science-policy arenas for regional ocean governance in the wider Caribbean region. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rotter:2021:NTF

[2354] Ana Rotter, Susana P. Gaudêncio, Katja Klun, Jan-Niklas Macher, Olivier P. Thomas, Irem Deniz, Christine Edwards, Ernesta Grigalionyte-Bembič, Zrinka Ljubešić, Johan Robbens, Giovanna Cristina Varese, and Marlen I. Vasquez. A new tool for faster construction of marine biotechnology collaborative networks. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Macheriotou:2021:CDS

[2355] Lara Macheriotou, Annelien Rigaux, Karine Olu, Daniela Zeppilli, Sofie Derycke, and Ann Vanreusel. Corrigendum: Deep-sea nematodes of the Mozambique Channel: Evidence of random community assembly dynamics in seep sediments. Frontiers in Marine Science, 8, April 2021. ISSN 2296-7745. See [1366].

Fiechter:2021:PSC

[2356] Jerome Fiechter, Mercedes Pozo Buil, Michael G. Jacox, Michael A. Alexander, and Kenneth A. Rose. Projected shifts in 21st Century sardine distribution and catch in the California Current. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Vazquez:2021:RUS

[2357] Elsa Vázquez, Sarah A. Woodin, David S. Wethey, Laura G. Peteiro, and Celia Olabarria. Reproduction under stress: Acute effect of low salinities and heat waves on reproductive cycle of four ecologically and commercially important bivalves. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Franca:2021:HDC

[2358] Susana França, Vanessa F. Fonseca, Susanne E. Tanner, Rita P. Vasconcelos, Patrick Reis-Santos, Anabela Maia, Miguel Ruano, Inês Cardoso, Sofia Henriques, Miguel Pessanha Pais, Pedro Brandão, Ana C. Brito, Paola Castellanos, Paula Chainho, Bernardo Quintella, Luís da Costa, Maria José Costa, José Lino Costa, and Henrique N. Cabral. Historical data in the CoastNet Geoportal: Documenting fish assemblages in Portuguese estuaries. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Coelho:2021:MMM

[2359] Pedro Coelho, Frederico Carvalho, Thomas Goulding, Paula Chainho, and José Guerreiro. Management models of the Manila clam (Ruditapes philippinarum) fisheries in invaded European coastal systems. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See corrigendum [3136].

Ridall:2021:SFL

[2360] Aaron Ridall and Jeroen Ingels. Suitability of free-living marine nematodes as bioindicators: Status and future considerations. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hou:2021:GIT

[2361] Cuihong Hou, Shougang Song, Lulu Zhu, Lili Shi, Beiping Tan, and Shuang Zhang. Growth, immunity, and transcriptome response to dietary wnt/β-catenin pathway activator TWS119 in Penaeus vannamei. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Liu:2021:DDE

[2362] Zi-Yan Liu, Hong-Ling Yang, Ling-Hao Hu, Wei Yang, Chun-Xiang Ai, and Yun-Zhang Sun. Dose-dependent effects of histamine on growth, immunity and intestinal health in juvenile grouper (Epinephelus coioides). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Jordan:2021:PRT

[2363] Christian Jordan, Jan Visscher, and Torsten Schlurmann. Projected responses of tidal dynamics in the North Sea to sea-level rise and morphological changes in the Wadden Sea. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Russo:2021:LHC

[2364] Elisabetta Russo, Marco Anelli Monti, Giacomo Toninato, Claudio Silvestri, Alessandra Raffaetà, and Fabio Pranovi. Lockdown: How the COVID-19 pandemic affected the fishing activities in the Adriatic Sea (Central Mediterranean Sea). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Drake:2021:PTV

[2365] Jeana L. Drake, Assaf Malik, Yotam Popovits, Oshra Yosef, Eli Shemesh, Jarosław Stolarski, Dan Tchernov, Daniel Sher, and Tali Mass. Physiological and transcriptomic variability indicative of differences in key functions within a single coral colony. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Ma:2021:DAA

[2366] Li Ma, Shangjin Tan, Hongbin Liu, Shuh-Ji Kao, Minhan Dai, and Jin-Yu Terence Yang. Distribution and activity of ammonia-oxidizers on the size-fractionated particles in the Pearl River Estuary. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Plaganyi:2021:IIC

[2367] Éva Plagányi, Roy Aijun Deng, Mark Tonks, Nicole Murphy, Sean Pascoe, Steven Edgar, Kinam Salee, Trevor Hutton, Laura Blamey, and Leo Dutra. Indirect impacts of COVID-19 on a tropical lobster fishery’s harvest strategy and supply chain. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Santos:2021:MHU

[2368] Catarina C. Santos, Andrés Domingo, John Carlson, Lisa J. Natanson, Paulo Travassos, David Macías, Enric Cortés, Philip Miller, Fábio Hazin, Federico Mas, Josetxu Ortiz de Urbina, Pedro G. Lino, and Rui Coelho. Movements, habitat use, and diving behavior of shortfin mako in the Atlantic Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Avila:2021:MBA

[2369] Conxita Avila and Carlos Angulo-Preckler. A minireview on biodiscovery in Antarctic marine benthic invertebrates. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Travers-Smith:2021:VNP

[2370] Hana Travers-Smith, Fernanda Giannini, Akash R. Sastri, and Maycira Costa. Validation of non-photochemical quenching corrections for chlorophyll-a measurements aboard ships of opportunity. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Caldarescu:2021:IDD

[2371] Diana E. Caldarescu, Thomas Brey, Doris Abele, Lars Beierlein, Gerrit Lohmann, and Monica Ionita. The influence of depth-dependent seasonal temperature variability on growth signal in Arctica islandica. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Zheng:2021:MCE

[2372] Peihua Zheng, Xiuxia Zhang, Dongmei Wang, Juntao Li, Zelong Zhang, Yaopeng Lu, Jianan Xian, Anli Wang, and Lei Wang. Molecular characterization and expression analysis of a novel glutaredoxin 3 gene in Pacific white shrimp (Litopenaeus vannamei). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rabaoui:2021:DDD

[2373] Lotfi Rabaoui, Ruben H. Roa-Ureta, Lamia Yacoubi, Yu-Jia Lin, Rommel Maneja, Thadickal V. Joydas, Premlal Panickan, Jinoy Gopalan, Ronald Loughland, Perdana K. Prihartato, Ali Qassem, Tyas I. Hikmawan, Bruno Diaz Lopez, and Mohammed A. Qurban. Diversity, distribution, and density of marine mammals along the Saudi waters of the Arabian Gulf: Update from a multi-method approach. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Garcia-Davis:2021:BAM

[2374] Sara García-Davis, Carolina P. Reyes, Irene Lagunes, José M. Padrón, Eugenio Fraile-Nuez, José J. Fernández, and Ana R. Díaz-Marrero. Bioprospecting antiproliferative marine microbiota from submarine volcano Tagoro. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Garcia-Rodriguez:2021:SIA

[2375] Emiliano García-Rodríguez, Sharon Z. Herzka, Oscar Sosa-Nishizaki, Christopher G. Lowe, and John B. O’Sullivan. Stable isotope analysis of juvenile white sharks inside a nursery area reveals foraging in demersalinshore habitats and trophic overlap with sympatric sharks. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Spillman:2021:NOS

[2376] Claire M. Spillman and Grant A. Smith. A new operational seasonal thermal stress prediction tool for coral reefs around Australia. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Barnes:2021:BFD

[2377] Natasha Maria Barnes, Samir R. Damare, and Belle Damodara Shenoy. Bacterial and fungal diversity in sediment and water column from the abyssal regions of the Indian Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Gac:2021:DDC

[2378] Jean-Philippe Gac, Pierre Marrec, Thierry Cariou, Emilie Grosstefan, Éric Macé, Peggy Rimmelin-Maury, Marc Vernet, and Yann Bozec. Decadal dynamics of the CO_{2} system and associated ocean acidification in coastal ecosystems of the North East Atlantic Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Aspillaga:2021:HTT

[2379] Eneko Aspillaga, Robert Arlinghaus, Martina Martorell-Barceló, Margarida Barcelo-Serra, and Josep Alós. High-throughput tracking of social networks in marine fish populations. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Planas:2021:PBD

[2380] Miquel Planas, Ike Olivotto, María Jesús González, Rosaria Laurà, Carlo Angeletti, Adolfo Amici, and Matteo Zarantoniello. Pre-breeding diets in the seahorse Hippocampus reidi: How do they affect fatty acid profiles, energetic status and histological features in newborn? Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Vedor:2021:ODV

[2381] Marisa Vedor, Gonzalo Mucientes, Sofia Hernández-Chan, Rui Rosa, Nick Humphries, David W. Sims, and Nuno Queiroz. Oceanic diel vertical movement patterns of blue sharks vary with water temperature and productivity to change vulnerability to fishing. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gallardo:2021:SSS

[2382] Camila Gallardo, Nicolas C. Ory, María de los Ángeles Gallardo, Marcel Ramos, Luis Bravo, and Martin Thiel. Sea-surface slicks and their effect on the concentration of plastics and zooplankton in the coastal waters of Rapa Nui (Easter Island). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fernandez:2021:MFS

[2383] Marc Fernandez, Filipe Alves, Rita Ferreira, Jan-Christopher Fischer, Paula Thake, Nuno Nunes, Rui Caldeira, and Ana Dinis. Modeling finescale cetaceans’ distributions in oceanic islands: Madeira Archipelago as a case study. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kato:2021:ERS

[2384] Seiji Kato, Fred G. Rose, Fu-Lung Chang, David Painemal, and William L. Smith. Evaluation of regional surface energy budget over ocean derived from satellites. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Haussermann:2021:SFH

[2385] Vreni Häussermann, Stacy Anushka Ballyram, Günter Försterra, Claudio Cornejo, Christian M. Ibáñez, Javier Sellanes, Aris Thomasberger, Juan Pablo Espinoza, and Francine Beaujot. Species that fly at a higher game: Patterns of deep–water emergence along the Chilean Coast, including a global review of the phenomenon. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kinnby:2021:CCI

[2386] Alexandra Kinnby, Gunilla B. Toth, and Henrik Pavia. Climate change increases susceptibility to grazers in a foundation seaweed. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Anderson:2021:NRC

[2387] James M. Anderson, Alyssa J. Clevenstine, Brian S. Stirling, Echelle S. Burns, Emily N. Meese, Connor F. White, Ryan K. Logan, John O’Sullivan, Patrick T. Rex, Jack May, Kady Lyons, Chuck Winkler, Emiliano García-Rodríguez, Oscar Sosa-Nishizaki, and Christopher G. Lowe. Non-random co-occurrence of juvenile white sharks (Carcharodon carcharias) at seasonal aggregation sites in Southern California. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:SML

[2388] Wei Zhang, Beiping Tan, Junming Deng, Xiaohui Dong, Qihui Yang, Shuyan Chi, Hongyu Liu, Shuang Zhang, Shiwei Xie, and Haitao Zhang. The single-molecule long-read sequencing of intestine after soy mealinduced enteritis in juvenile pearl gentian grouper, Epinephelus fuscoguttatus[Venus sign] × Epinephelus lanceolatus [male sign]. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wollschlager:2021:ECC

[2389] Jochen Wollschläger, Patrick J. Neale, Rebecca L. North, Maren Striebel, and Oliver Zielinski. Editorial: Climate change and light in aquatic ecosystems: Variability & ecological consequences. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Figueroa:2021:TIC

[2390] Marcela Montserrat Landero Figueroa, Miles J. G. Parsons, Benjamin J. Saunders, Ben Radford, and Iain M. Parnum. Testing the improvement of coral reef associated fish distribution models based on multibeam bathymetry by adding seafloor backscatter data. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Weiss:2021:DCS

[2391] Michael N. Weiss, Samuel Ellis, and Darren P. Croft. Diversity and consequences of social network structure in toothed whales. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Yuyama:2021:ARI

[2392] Ikuko Yuyama, Tomihiko Higuchi, and Michio Hidaka. Application of RNA interference technology to acroporid juvenile corals. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Buasakaew:2021:WBC

[2393] Nutcha Buasakaew, Benny K. K. Chan, and Kringpaka Wangkulangkul. Why are barnacles common on intertidal rocks but rare in rock pools? Effect of water temperature, salinity, and continuous submergence on barnacle survival in Indian Ocean rock pools. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Lemley:2021:EMC

[2394] Daniel A. Lemley, Stephen J. Lamberth, Warren Manuel, Monique Nunes, Gavin M. Rishworth, Lara van Niekerk, and Janine B. Adams. Effective management of closed hypereutrophic estuaries requires catchment-scale interventions. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Garcia-Salinas:2021:DSC

[2395] Pablo García-Salinas, Victor Gallego, and Juan F. Asturiano. Development of sperm cryopreservation protocols for sharks and rays: New tools for elasmobranch conservation. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Melvin:2021:CGS

[2396] Jessica Melvin, Madeline Bury, Justine Ammendolia, Charles Mather, and Max Liboiron. Critical gaps in shoreline plastics pollution research. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Cheng:2021:WIC

[2397] Peng Cheng, Han Wang, Xin Nie, Siyang Zhu, Zhoupeng Chen, Xingyi Wu, Anlu Zhang, and Jing Wang. What are the impacts of a coastal zone protection policy on farmers’ livelihood capital? Empirical analysis from the perspective of farmer participation. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Pelletier:2021:SWB

[2398] Dominique Pelletier, David Roos, Marc Bouchoucha, Thomas Schohn, William Roman, Charles Gonson, Thomas Bockel, Liliane Carpentier, Bastien Preuss, Abigail Powell, Jessica Garcia, Matthias Gaboriau, Florent Cadé, Coline Royaux, Yvan Le Bras, and Yves Reecht. A standardized workflow based on the STAVIRO unbaited underwater video system for monitoring fish and habitat essential biodiversity variables in coastal areas. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

McParland:2021:OTB

[2399] Erin L. McParland, Harriet Alexander, and Winifred M. Johnson. The osmolyte ties that bind: Genomic insights into synthesis and breakdown of organic osmolytes in marine microbes. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Candela:2021:URS

[2400] Alberto Candela, Kevin Edelson, Michelle M. Gierach, David R. Thompson, Gail Woodward, and David Wettergreen. Using remote sensing and in situ measurements for efficient mapping and optimal sampling of coral reefs. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cowley:2021:IQC

[2401] Rebecca Cowley, Rachel E. Killick, Tim Boyer, Viktor Gouretski, Franco Reseghetti, Shoichi Kizu, Matthew D. Palmer, Lijing Cheng, Andrea Storto, Marc Le Menn, Simona Simoncelli, Alison M. Macdonald, and Catia M. Domingues. International quality-controlled ocean database (IQuOD) v0.1: The temperature uncertainty specification. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Basan:2021:SGB

[2402] Fritjof Basan, Jens-Georg Fischer, and Dennis Kühnel. Soundscapes in the German Baltic Sea before and during the Covid-19 pandemic. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kong:2021:DPB

[2403] Weiyuan Kong, Liping Zhou, Georges Aumaître, Didier Bourlès, and Karim Keddadouche. Dissolved and particulate beryllium isotopes in the Pearl River Estuary: Their geochemical behavior in estuarine water and potential contributions from anthropogenic sources. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Sinnickson:2021:EIR

[2404] Dylan Sinnickson, David Chagaris, and Micheal Allen. Exploring impacts of river discharge on forage fish and predators using Ecopath with Ecosim. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Moradi:2021:NMB

[2405] Nasrollah Moradi, Isabell Klawonn, Morten H. Iversen, Frank Wenzhöfer, Hans-Peter Grossart, Helle Ploug, Gerhard Fischer, and Arzhang Khalili. A novel measurement-based model for calculating diffusive fluxes across substrate-water interfaces of marine aggregates, sediments and biofilms. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lovindeer:2021:CLM

[2406] Raisha Lovindeer, Lawrence Abbott, Hannah Medina, and Katherine R. M. Mackey. Costs and limitations of marine synechococcus blue-green chromatic acclimation. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Osiecka:2021:UWM

[2407] Anna N. Osiecka, Susanna Quer, Aleksandra Wróbel, and Katarzyna Osiecka-Brzeska. Unpaid work in marine science: a snapshot of the earlycareer job market. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Yang:2021:CCG

[2408] Mei-Jie Yang, Jie Feng, Hao Song, Zheng-Lin Yu, Pu Shi, Jian Liang, Zhi Hu, Cong Zhou, Xiao-Lin Wang, and Tao Zhang. The characteristic of critical genes in neuroendocrine system and their regulation on food habit transition and metamorphosis of veined rapa whelk Rapana venosa (Valenciennes, 1846). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hoellein:2021:MIF

[2409] Timothy Hoellein, Catherine Rovegno, Amy V. Uhrin, Ed Johnson, and Carlie Herring. Microplastics in invasive freshwater mussels (Dreissena sp.): Spatiotemporal variation and occurrence with chemical contaminants. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Richardson:2021:GCD

[2410] Kelsey Richardson, Britta Denise Hardesty, Joanna Zofia Vince, and Chris Wilcox. Global causes, drivers, and prevention measures for lost fishing gear. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rittschof:2021:ECD

[2411] Daniel Rittschof. Editorial: Cross-disciplinary marine research: New tools and societal interfaces. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Schuback:2021:STV

[2412] Nina Schuback, Philippe D. Tortell, Ilana Berman-Frank, Douglas A. Campbell, Aurea Ciotti, Emilie Courtecuisse, Zachary K. Erickson, Tetsuichi Fujiki, Kimberly Halsey, Anna E. Hickman, Yannick Huot, Maxime Y. Gorbunov, David J. Hughes, Zbigniew S. Kolber, C. Mark Moore, Kevin Oxborough, Ondřej Prášil, Charlotte M. Robinson, Thomas J. Ryan-Keogh, Greg Silsbe, Stefan Simis, David J. Suggett, Sandy Thomalla, and Deepa R. Varkey. Single-turnover variable chlorophyll fluorescence as a tool for assessing phytoplankton photosynthesis and primary productivity: Opportunities, caveats and recommendations. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Chen:2021:OOU

[2413] Emily Yi-Shyuan Chen. Often overlooked: Understanding and meeting the current challenges of marine invertebrate conservation. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lengar:2021:SPM

[2414] Živa Lengar, Katja Klun, Iztok Dogsa, Ana Rotter, and David Stopar. Sequestration of polystyrene microplastics by jellyfish mucus. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Jiang:2021:ESP

[2415] Zong-Pei Jiang, Yuqi Tong, Mengmeng Tong, Jiajun Yuan, Qin Cao, and Yiwen Pan. The effects of suspended particulate matter, nutrient, and salinity on the growth of Amphidinium carterae under estuary environmental conditions. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Waiho:2021:IIV

[2416] Khor Waiho, Mhd Ikhwanuddin, Muyassar H. Abualreesh, Alexander Chong Shu-Chien, Sairatul Dahlianis Ishak, Mohamad Jalilah, Ghazali Azmie, and Hanafiah Fazhan. Intra- and interspecific variation in sexual dimorphism patterns of mud crab genus Scylla along the equatorial region. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Han:2021:ADD

[2417] Yijing Han, Gege Hu, Yuying Chen, Lizhu Chen, Daode Yu, Qianqian Zhang, and Dinglong Yang. Antimicrobial defensin and DNA traps in Manila clam Ruditapes philippinarum: Implications for their roles in immune responses. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Goldstein:2021:RRA

[2418] Esther D. Goldstein, Thomas E. Helser, Johanna J. Vollenweider, Ashwin Sreenivasan, and Fletcher F. Sewall. Rapid and reliable assessment of fish physiological condition for fisheries research and management using Fourier Transform Near-Infrared spectroscopy. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhao:2021:SUU

[2419] Qi Zhao, Ziqiang Zheng, Huimin Zeng, Zhibin Yu, Haiyong Zheng, and Bing Zheng. The synthesis of unpaired underwater images for monocular underwater depth prediction. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wootton:2021:CMF

[2420] Nina Wootton, Marta Ferreira, Patrick Reis-Santos, and Bronwyn M. Gillanders. A comparison of microplastic in fish from Australia and Fiji. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Bravo:2021:RVH

[2421] Gonzalo Bravo, Nicolas Moity, Edgardo Londoño-Cruz, Frank Muller-Karger, Gregorio Bigatti, Eduardo Klein, Francis Choi, Lark Parmalee, Brian Helmuth, and Enrique Montes. Robots versus humans: Automated annotation accurately quantifies essential ocean variables of rocky intertidal functional groups and habitat state. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wei:2021:SFB

[2422] Yuqiu Wei, Zhaoyi Zhang, Zhengguo Cui, and Jun Sun. Size-fractionated biogenic silica standing stocks and carbon biomass in the western tropical North Pacific: Evidence for the ecological importance of pico-sized plankton in oligotrophic gyres. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Wei:2021:LMP

[2423] Yuliang Wei, Benxiang Li, Houguo Xu, and Mengqing Liang. Liver metabolome and proteome response of turbot (Scophthalmus maximus) to lysine and leucine in free and dipeptide forms. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Shimabukuro:2021:CDD

[2424] Mauricio Shimabukuro, Orlemir Carrerette, Joan Manel Alfaro-Lucas, Alexandra Elaine Rizzo, Kenneth M. Halanych, and Paulo Yukio Gomes Sumida. Corrigendum: Diversity, distribution and phylogeny of Hesionidae (Annelida) colonizing whale falls: New species of Sirsoe and connections between ocean basins. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745. See [5].

Mello-Rafter:2021:EGE

[2425] Kristen Mello-Rafter, Derek Sowers, Mashkoor Malik, Les Watling, Larry A. Mayer, and Jennifer A. Dijkstra. Environmental and geomorphological effects on the distribution of deep-sea canyon and seamount communities in the Northwest Atlantic. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Johnstone:2021:PCP

[2426] Carolina Johnstone and José Carlos Báez. Placing the COVID-19 pandemic in a marine ecological context: Potential risks for conservation of marine air-breathing animals and future zoonotic outbreaks. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Casas-Monroy:2021:DBW

[2427] Oscar Casas-Monroy and Sarah A. Bailey. Do ballast water management systems reduce phytoplankton introductions to Canadian waters? Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Benkiran:2021:AIA

[2428] Mounir Benkiran, Giovanni Ruggiero, Eric Greiner, Pierre-Yves Le Traon, Elisabeth Rémy, Jean Michel Lellouche, Romain Bourdallé-Badie, Yann Drillet, and Babette Tchonang. Assessing the impact of the assimilation of SWOT observations in a global high-resolution analysis and forecasting system part 1: Methods. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Pierri:2021:DDM

[2429] Cataldo Pierri, Frine Cardone, Giuseppe Corriero, Tamara Lazic, Federico Quattrocchi, Giorgio Alabiso, and Michele Gristina. Density decline in a Mediterranean Seahorse population: Natural fluctuations or new emerging threats? Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Knobel:2021:SDS

[2430] Loreen Knöbel, Jennifer C. Nascimento-Schulze, Trystan Sanders, Dominique Zeus, Claas Hiebenthal, Francisco R. Barboza, Heiko Stuckas, and Frank Melzner. Salinity driven selection and local adaptation in Baltic Sea mytilid mussels. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Aalto:2021:DAO

[2431] Nerea J. Aalto, Karley Campbell, Hans C. Eilertsen, and Hans C. Bernstein. Drivers of atmosphere-ocean CO_{2} flux in Northern Norwegian fjords. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Meunier:2021:DDN

[2432] Valentine Meunier, Sophie Bonnet, Mar Benavides, Andreas Ravache, Olivier Grosso, Christophe Lambert, and Fanny Houlbrèque. Diazotrophderived nitrogen assimilation strategies differ by scleractinian coral species. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Luo:2021:EWC

[2433] Zhifa Luo, Bensheng Huang, Xiaohong Chen, Chao Tan, Jing Qiu, and Guangling Huang. Effects of wave–current interaction on storm surge in the Pearl River Estuary: a case study of Super Typhoon Mangkhut. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wilms:2021:BGI

[2434] Sabrina Natalie Wilms. A beginner’s guide on integrating *Omics approaches to study marine microbial communities: Details and discussions from sample collection to bioinformatics analysis. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Baskaran:2021:NAD

[2435] Mark Baskaran and Katherine Krupp. Novel application of ^{210}Po–^{210}Pb disequilibria to date snow, melt pond, ice core, and ice-rafted sediments in the Arctic Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Xiao:2021:CCE

[2436] Kang Xiao, Xi Wang, Wen-Bin Liu, Ding-Dong Zhang, Xiang-Fei Li, Chun-Nuan Zhang, Wei-Hua Chen, Kenneth Prudence Abasubong, and Guang-Zhen Jiang. Corticosterone can be an essential stress index in channel catfish (Ictalurus punctatus). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cleland:2021:DMC

[2437] Jason Cleland, Georgios Kazanidis, J. Murray Roberts, and Steve W. Ross. Distribution of megabenthic communities under contrasting settings in deep-sea cold seeps near Northwest Atlantic canyons. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

dePuelles:2021:ASZ

[2438] María Luz Fernández de Puelles, Magdalena Gazá, Miguel Cabanellas-Reboredo, Alba González-Vega, Inma Herrera, Carmen Presas-Navarro, Jesús M. Arrieta, and Eugenio Fraile-Nuez. Abundance and structure of the zooplankton community during a post-eruptive process: The case of the submarine volcano Tagoro (El Hierro; Canary Islands), 2013–2018. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Garzia:2021:VDP

[2439] Matteo Garzia, Paolo Mariottini, Daniele Salvi, and Giulia Furfaro. Variation and diagnostic power of the internal transcribed spacer 2 in Mediterranean and Atlantic eolid nudibranchs (Mollusca, Gastropoda). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Rolfer:2021:IRI

[2440] Lena Rölfer, Arianna Liconti, Natalie Prinz, and Clara Antonia Klöcker. Integrated research for integrated ocean management. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Shore:2021:CRF

[2441] Amanda Shore, Jordan A. Sims, Michael Grimes, Lauren I. Howe-Kerr, Carsten G. B. Grupstra, Shawn M. Doyle, Lauren Stadler, Jason B. Sylvan, Kathryn E. F. Shamberger, Sarah W. Davies, Lory Z. Santiago-Vázquez, and Adrienne M. S. Correa. Corrigendum: On a Reef Far, Far Away: Anthropogenic Impacts Following Extreme Storms Affect Sponge Health and Bacterial Communities. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See [1473].

Wang:2021:MLG

[2442] Heng Wang, Yisheng Peng, Chen Wang, Qingke Wen, Jinyong Xu, Zhan Hu, Xing Jia, Xuemei Zhao, Weihang Lian, Stijn Temmerman, Judith Wolf, and Tjeerd Bouma. Mangrove loss and gain in a densely populated urban estuary: Lessons from the Guangdong–Hong Kong–Macao Greater Bay Area. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kregting:2021:SMG

[2443] Louise Kregting, Damon Britton, Craig N. Mundy, and Catriona L. Hurd. Safe in my garden: Reduction of mainstream flow and turbulence by macroalgal assemblages and implications for refugia of calcifying organisms from ocean acidification. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wang:2021:IRI

[2444] Bo Wang, Limin Han, and Hongzhi Zhang. The impact of regional industrial structure upgrading on the economic growth of marine fisheries in China — the perspective of industrial structure advancement and rationalization. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sun:2021:RBM

[2445] Mingshuai Sun, Xianyong Zhao, Yancong Cai, Kui Zhang, and Zuozhi Chen. Relationship between multi-factors and short-term changes in fishery resources. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Zhang:2021:IAD

[2446] Shuai Zhang, Min Li, Jiangfeng Zhu, Shannan Xu, and Zuozhi Chen. An integrated approach to determine the stock structure of spinyhead croaker Collichthys lucidus (Sciaenidae) in Chinese coastal waters. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Serpetti:2021:MSS

[2447] Natalia Serpetti, Steven Benjamins, Stevie Brain, Maurizio Collu, Bethany J. Harvey, Johanna J. Heymans, Adam D. Hughes, Denise Risch, Sophia Rosinski, James J. Waggitt, and Ben Wilson. Modeling small scale impacts of multi-purpose platforms: an ecosystem approach. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rogers:2021:PIA

[2448] Kerrylee Rogers and Neil Saintilan. Processes influencing autocompaction modulate coastal wetland surface elevation adjustment with sealevel rise. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Yang:2021:USD

[2449] Weifeng Yang, Xiufeng Zhao, Laodong Guo, Bangqin Huang, Min Chen, Ziming Fang, Xiao Zhang, and Yusheng Qiu. Utilization of soot and 210Po–^{210}Pb disequilibria to constrain particulate organic carbon fluxes in the Northeastern South China Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Parsons:2021:DPD

[2450] Aoife E. Parsons, Ole B. Samuelsen, Ingrid Askeland Johnsen, Rita Hannisdal, Tore Tjensvoll, and Vivian Husa. Distribution and persistence of diflubenzuron and teflubenzuron in the marine environment around salmonid aquaculture facilities. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Hatun:2021:NSG

[2451] Hjálmar Hátún, Léon Chafik, and Karin Margretha Húsgard Larsen. The Norwegian Sea Gyre — a regulator of Iceland–Scotland ridge exchanges. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Talijancic:2021:PRG

[2452] Igor Talijančić, Iva Žužul, Viktorija Kiridžija, Jasna Šiljić, Jelka Pleadin, Leon Grubišić, and Tanja Šegvić-Bubić. Plastic responses of gilthead seabream Sparus aurata to wild and aquaculture pressured environments. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gleason:2021:PGC

[2453] Arthur C. R. Gleason, Ross Smith, Sam J. Purkis, Kyle Goodrich, Alexandra Dempsey, and Alejandro Mantero. The prospect of global coral reef bathymetry by combining ice, cloud, and land elevation satellite-2 altimetry with multispectral satellite imagery. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Karlovic:2021:FEM

[2454] Thamíris C. Karlovic, Renata R. Gomes, Paulo C. Paiva, Elizabeth A. Babcock, and June F. Dias. Functionality and effectiveness of Marine Protected Areas in Southeastern Brazilian waters for demersal elasmobranchs. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Goverts:2021:ERT

[2455] Zoë Goverts, Paul Nührenberg, and Alex Jordan. Environmental reconstruction and tracking as methods to explore social interactions in marine environments: a test case with the Mediterranean rainbow wrasse Coris julis. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Gu:2021:HEP

[2456] Ting Gu, Dai Jia, Xing Ma, Liying Peng, Guicheng Zhang, Yuqiu Wei, Tingting Lou, and Jun Sun. Hypoxia-enhanced N_{2}O production under ocean acidification in the Bohai Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Goren:2021:SAP

[2457] Liron Goren, Tal Idan, Sigal Shefer, and Micha Ilan. Sponge-associated polychaetes: Not a random assemblage. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Rubin:2021:MMC

[2458] Ewelina T. Rubin, Ian C. Enochs, Colin Foord, Anderson B. Mayfield, Graham Kolodziej, Isabelle Basden, and Derek P. Manzello. Molecular mechanisms of coral persistence within highly urbanized locations in the Port of Miami, Florida. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Choi:2021:PMS

[2459] Sun Min Choi, Jun Young Seo, Guan hong Lee, Xiaoteng Shen, and Ho Kyung Ha. Practical method to screen contaminated holograms of flocs using light intensity. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Jobson:2021:CHB

[2460] Sara Jobson, Jean-François Hamel, Taylor Hughes, and Annie Mercier. Cellular, hormonal, and behavioral responses of the holothuroid Cucumaria frondosa to environmental stressors. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Choi:2021:DES

[2461] Jun Myoung Choi, Wonkook Kim, Tran Thy My Hong, and Young-Gyu Park. Derivation and evaluation of satellite-based surface current. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Diehl:2021:SHI

[2462] Nora Diehl, Michael Y. Roleda, Inka Bartsch, Ulf Karsten, and Kai Bischof. Summer heatwave impacts on the European kelp Saccharina latissima across its latitudinal distribution gradient. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ha:2021:CMD

[2463] Hun Jun Ha and Ho Kyung Ha. Comparison of methods for determining erosion threshold of cohesive sediments using a microcosm system. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Menden-Deuer:2021:PID

[2464] Susanne Menden-Deuer, Wayne Homer Slade, and Heidi Dierssen. Promoting instrument development for new research avenues in ocean science: Opening the black box of grazing. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Guazzo:2021:SFW

[2465] Regina A. Guazzo, Ian N. Durbach, Tyler A. Helble, Gabriela C. Alongi, Cameron R. Martin, Stephen W. Martin, and E. Elizabeth Henderson. Singing fin whale swimming behavior in the Central North Pacific. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Testor:2021:COC

[2466] Pierre Testor, Brad de Young, Daniel L. Rudnick, Scott Glenn, Daniel Hayes, Craig M. Lee, Charitha Pattiaratchi, Katherine Hill, Emma Heslop, Victor Turpin, Pekka Alenius, Carlos Barrera, John A. Barth, Nicholas Beaird, Guislain Bécu, Anthony Bosse, François Bourrin, J. Alexander Brearley, Yi Chao, Sue Chen, Jacopo Chiggiato, Laurent Coppola, Richard Crout, James Cummings, Beth Curry, Ruth Curry, Richard Davis, Kruti Desai, Steve DiMarco, Catherine Edwards, Sophie Fielding, Ilker Fer, Eleanor Frajka-Williams, Hezi Gildor, Gustavo Goni, Dimitri Gutierrez, Peter Haugan, David Hebert, Joleen Heiderich, Stephanie Henson, Karen Heywood, Patrick Hogan, Loïc Houpert, Sik Huh, Mark E. Inall, Masso Ishii, Shin ichi Ito, Sachihiko Itoh, Sen Jan, Jan Kaiser, Johannes Karstensen, Barbara Kirkpatrick, Jody Klymak, Josh Kohut, Gerd Krahmann, Marjolaine Krug, Sam McClatchie, Frédéric Marin, Elena Mauri, Avichal Mehra, Michael P. Meredith, Thomas Meunier, Travis Miles, Julio M. Morell, Laurent Mortier, Sarah Nicholson, Joanne O’Callaghan, Diarmuid O’Conchubhair, Peter Oke, Enric Pallàs-Sanz, Matthew Palmer, JongJin Park, Leonidas Perivoliotis, Pierre-Marie Poulain, Ruth Perry, Bastien Queste, Luc Rainville, Eric Rehm, Moninya Roughan, Nicholas Rome, Tetjana Ross, Simon Ruiz, Grace Saba, Amandine Schaeffer, Martha Schönau, Katrin Schroeder, Yugo Shimizu, Bernadette M. Sloyan, David Smeed, Derrick Snowden, Yumi Song, Sebastian Swart, Miguel Tenreiro, Andrew Thompson, Joaquin Tintore, Robert E. Todd, Cesar Toro, Hugh Venables, Taku Wagawa, Stephanie Waterman, Roy A. Watlington, and Doug Wilson. Corrigendum: OceanGliders: a component of the integrated GOOS. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See [4].

Floren:2021:RFI

[2467] Adonis S. Floren, Ken ichi Hayashizaki, Sumaitt Putchakarn, Piyalap Tuntiprapas, and Anchana Prathep. A review of factors influencing the seagrass-sea cucumber association in tropical seagrass meadows. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Stack:2021:BIC

[2468] Stephanie H. Stack, Kate R. Sprogis, Grace L. Olson, Florence A. Sullivan, Abigail F. Machernis, and Jens J. Currie. The behavioural impacts of commercial swimming with whale tours on humpback whales (Megaptera novaeangliae) in Hervey Bay, Australia. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Garcia-Fernandez:2021:STV

[2469] Cristina García-Fernández, Justin J. Suca, Joel K. Llopiz, Paula Álvarez, Rosario Domínguez-Petit, and Fran Saborido-Rey. Spatial and temporal variability in the occurrence and abundance of European hake larvae, Merluccius merluccius, on the Galician Shelf (NE Atlantic). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhang:2021:IDB

[2470] Yu Zhang, Jianzhong Huang, Xinke Li, Cuiping Song, Jingquan Xia, Yuxiao Ren, Aimin Wang, and Xiubao Li. Improper diving behavior affects physiological responses of Acropora hyacinthus and Porites cylindrica. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Borges:2021:IEP

[2471] Francisco O. Borges, Catarina P. Santos, José R. Paula, Enrique Mateos-Naranjo, Susana Redondo-Gomez, Janine Barbara Adams, Isabel Caçador, Vanessa F. Fonseca, Patrick Reis-Santos, Bernardo Duarte, and Rui Rosa. Invasion and extirpation potential of native and invasive Spartina species under climate change. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Wong:2021:SCV

[2472] Shu-Kuan Wong, Shotaro Suzuki, Yingshun Cui, Ryo Kaneko, Kazuhiro Kogure, and Koji Hamasaki. Sampling constraints and variability in the analysis of bacterial community structures in the sea surface microlayer. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Marra:2021:CUG

[2473] Nicholas J. Marra, Michael J. Stanhope, Nathaniel K. Jue, Vincent P. Richards, Stephen J. O’Brien, Agostinho Antunes, and Mahmood S. Shivji. Commentary: Unbiasing genome-based analyses of selection: an example using iconic shark species. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See [1382].

Possenti:2021:ASG

[2474] Luca Possenti, Matthew P. Humphreys, Dorothee C. E. Bakker, Marcos Cobas-García, Liam Fernand, Gareth A. Lee, Francesco Pallottino, Socratis Loucaides, Matt Charles Mowlem, and Jan Kaiser. Air–sea gas fluxes and remineralization from a novel combination of pH and O2 sensors on a glider. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Hanz:2021:CSV

[2475] Ulrike Hanz, Lindsay Beazley, Ellen Kenchington, Gerard Duineveld, Hans Tore Rapp, and Furu Mienis. Corrigendum: Seasonal variability in near-bed environmental conditions in the Vazella pourtalesii glass sponge grounds of the Scotian Shelf. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See [1217].

Xu:2021:ENA

[2476] Ningning Xu, Wenlei Wang, Yan Xu, Dehua Ji, Changsheng Chen, Chaotian Xie, and Kai Xu. Effects of nutrient availability on the release of dissolved and particulate organic carbon by Pyropia haitanensis and its implications. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Huveneers:2021:CSN

[2477] Charlie Huveneers, Yuri Niella, Michael Drew, Rory McAuley, Paul Butcher, Victor Peddemors, Daniela Waltrick, Chris Dowling, Silas Mountford, Ian Keay, and Matias Braccini. Continental-scale network reveals cross-jurisdictional movements of sympatric sharks with implications for assessment and management. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Gunn:2021:VMH

[2478] Kathryn L. Gunn, Alex Dickinson, Nicky J. White, and Colm cille P. Caulfield. Vertical mixing and heat fluxes conditioned by a seismically imaged oceanic front. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Al-Thani:2021:MEQ

[2479] Roda F. Al-Thani and Bassam T. Yasseen. Microbial ecology of Qatar, the Arabian Gulf: Possible roles of microorganisms. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Grasse:2021:CSI

[2480] Patricia Grasse, Kristin Haynert, Kristin Doering, Sonja Geilert, Janice L. Jones, Mark A. Brzezinski, and Martin Frank. Controls on the silicon isotope composition of diatoms in the Peruvian Upwelling. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Vega-Herrera:2021:SQM

[2481] Albert Vega-Herrera, Marta Llorca, Katerina Savva, Víctor M. León, Esteban Abad, and Marinella Farré. Screening and quantification of Micro(Nano)Plastics and plastic additives in the seawater of Mar Menor Lagoon. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Bosi:2021:RSD

[2482] Sofia Bosi, Göran Broström, and Fabien Roquet. The role of Stokes Drift in the dispersal of North Atlantic surface marine debris. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bam:2021:TPC

[2483] Wokil Bam, Kanchan Maiti, and Mark Baskaran. ^{210}Po and ^{210}Pb as tracers of particle cycling and export in the Western Arctic Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Dayras:2021:MDI

[2484] Paul Dayras, Capucine Bialais, Irina Sadovskaya, Min-Chul Lee, Jae-Seong Lee, and Sami Souissi. Microalgal diet influences the nutritive quality and reproductive investment of the cyclopoid copepod Paracyclopina nana. Frontiers in Marine Science, 8, August 2021. ISSN 2296-

Mowlem:2021:IPL

[2485] Matt Mowlem, Alexander Beaton, Robin Pascal, Allison Schaap, Socratis Loucaides, Sam Monk, Andrew Morris, Christopher L. Cardwell, Sara E. Fowell, Matthew D. Patey, and Patricia López-García. Industry partnership: Lab on chip chemical sensor technology for ocean observing. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fang:2021:SSP

[2486] Jinghui Fang, Wei He, Shan Meng, Zengjie Jiang, Jianguang Fang, Yuze Mao, and Bo Liang. The sediment selectivity of Perinereis aibuhitensis larvae: Active or passive? Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Lee:2021:IEU

[2487] Serena Blyth Lee, Fan Zhang, Charles James Lemckert, and Rodger Tomlinson. Investigations exploring the use of an unstructured-grid, finitevolume modelling approach to simulate coastal circulation in remote island settings — case study region, Vanuatu/New Caledonia. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Ge:2021:ACF

[2488] Ruping Ge, Hongju Chen, Yunyun Zhuang, and Guangxing Liu. Active carbon flux of mesozooplankton in South China Sea and Western Philippine Sea. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:TAR

[2489] Jingjing Zhang, Zhi Wu, Yujie He, Xinhui Li, and Jie Li. Transcriptome analysis reveals impaired fertility and immunity under salinity exposure in juvenile grass carp. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bertheussen:2021:SCA

[2490] Bernt Arne Bertheussen. Sustained competitive advantage based on industry-specific institutional frameworks. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Schill:2021:SSC

[2491] Steven R. Schill, Gregory P. Asner, Valerie Pietsch McNulty, F. Joseph Pollock, Aldo Croquer, Nicholas R. Vaughn, Ximena Escovar-Fadul, George Raber, and Elizabeth Shaver. Site selection for coral reef restoration using airborne imaging spectroscopy. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Birchill:2021:EOB

[2492] Antony J. Birchill, A. D. Beaton, Tom Hull, Jan Kaiser, Matt Mowlem, R. Pascal, A. Schaap, Yoana G. Voynova, C. Williams, and M. Palmer. Exploring ocean biogeochemistry using a lab-on-chip phosphate analyser on an underwater glider. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zou:2021:DBD

[2493] Shanmei Zou, Yachao Bao, Xuemin Wu, and Changhai Wang. DNA barcoding diatoms from China with multiple genes. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Gilannejad:2021:FPM

[2494] Neda Gilannejad, Francisco J. Moyano, Gonzalo Martínez-Rodríguez, and Manuel Yúfera. Feeding protocol modulates the digestive process in Senegalese sole (Solea senegalensis) juveniles. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Meyer:2021:SSL

[2495] Julia Meyer, Ingrid Kröncke, Alexander Bartholomä, Mathias Heckroth, and Gregor Scheiffarth. Small-scale and long-term variability in population dynamics of the cockle Cerastoderma edule in a Southern North Sea tidal flat system. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Rajasilta:2021:LTD

[2496] Marjut Rajasilta, Katja Mäkinen, Suvi Ruuskanen, Jari Hänninen, and Päivi Laine. Long-term data reveal the associations of the egg quality with abiotic factors and female traits in the Baltic herring under variable environmental conditions. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Naya-Catala:2021:MGS

[2497] Fernando Naya-Català, Giulia A. Wiggers, M. Carla Piazzon, Manuel I. López-Martínez, Itziar Estensoro, Josep A. Calduch-Giner, M. Carmen Martínez-Cuesta, Teresa Requena, Ariadna Sitjà-Bobadilla, Marta Miguel, and Jaume Pérez-Sánchez. Modulation of gilthead sea bream gut microbiota by a bioactive egg white hydrolysate: Interactions between bacteria and host lipid metabolism. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Goulding:2021:SEM

[2498] Thomas Goulding, Pedro M. Sousa, Gilda Silva, João Paulo Medeiros, Frederico Carvalho, Inês Metelo, Catarina Freitas, Nuno Lopes, Paula Chainho, and José Lino Costa. Shifts in estuarine macroinvertebrate communities associated with water quality and climate change. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Almahasheer:2021:IAA

[2499] Hanan Almahasheer. Internodal analysis of Avicennia marina in the Western Arabian Gulf. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Stevenne:2021:CSU

[2500] Chloé Stévenne, Maud Micha, Jean-Christophe Plumier, and Stéphane Roberty. Corals and sponges under the light of the Holobiont concept: How microbiomes underpin our understanding of marine ecosystems. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Zhang:2021:ELI

[2501] Xuekai Zhang, Chao Fan, Xingzhi Zhang, Qiongzhen Li, Yangchun Li, and Zhaoping Wang. Effects of light intensity and wavelength on the phototaxis of the Crassostrea gigas ([male sign]) and Crassostrea sikamea ([Venus sign]) hybrid larvae. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Fanelli:2021:IPP

[2502] Emanuela Fanelli, Silvia Bianchelli, Federica Foglini, Miquel Canals, Giorgio Castellan, Queralt Güell-Bujons, Bella Galil, Menachem Goren, Julian Evans, Marie-Claire Fabri, Sandrine Vaz, Tiziana Ciuffardi, Patrick J. Schembri, Lorenzo Angeletti, Marco Taviani, and Roberto Danovaro. Identifying priorities for the protection of deep Mediterranean Sea ecosystems through an integrated approach. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

vonLeesen:2021:AOO

[2503] Gotje von Leesen, Hlynur Bardarson, Sæmundur Ari Halldórsson, Martin J. Whitehouse, and Steven E. Campana. Accuracy of otolith oxygen isotope records analyzed by SIMS as an index of temperature exposure of wild Icelandic cod (Gadus morhua). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Wang:2021:DSC

[2504] Yong Wang, Wenli Li, Qingmei Li, Yingli Zhou, Zhaoming Gao, and Dong Feng. Deep-sea carbonates are a reservoir of fossil microbes previously inhabiting cold seeps. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Xing:2021:EIT

[2505] Lei Xing, Yong Chen, Bai Li, Kisei R. Tanaka, Robert Boenish, Yiping Ren, and Binduo Xu. Evaluating impacts of trophic interactions on the effectiveness of single-species fisheries management. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Tinker:2021:TWS

[2506] Jonathan Tinker and Leon Hermanson. Towards winter seasonal predictability of the North West European Shelf seas. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Adamopoulou:2021:DPF

[2507] Argyro Adamopoulou, Christina Zeri, Francesca Garaventa, Chiara Gambardella, Christos Ioakeimidis, and Elli Pitta. Distribution patterns of floating microplastics in open and coastal waters of the Eastern Mediterranean Sea (Ionian, Aegean, and Levantine Seas). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pertiwi:2021:CNC

[2508] Avi Putri Pertiwi, Chengfa Benjamin Lee, and Dimosthenis Traganos. Cloud-native coastal turbid zone detection using multi-temporal Sentinel-2 data on Google Earth Engine. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Aeby:2021:CSC

[2509] Greta Aeby, Blake Ushijima, Erich Bartels, Cory Walter, Joseph Kuehl, Scott Jones, and Valerie J. Paul. Changing Stony Coral Tissue Loss Disease dynamics through time in Montastraea cavernosa. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Carman:2021:CSI

[2510] Victoria González Carman, Pablo Denuncio, Martina Vassallo, María Paula Berón, Karina C. Álvarez, and Sergio Rodriguez-Heredia. Charismatic species as indicators of plastic pollution in the Río de la Plata Estuarine Area, SW Atlantic. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Aurelio:2021:TSS

[2511] Luísa Aurélio, Susana França, Vera Sequeira, Diana Boaventura, Maria João Correia, Bruno Pinto, Sandra Amoroso, Maria João Feio, Cristina Brito, Paula Chainho, and Luísa Chaves. Tell a story to save a river: Assessing the impact of using a children’s book in the classroom as a tool to promote environmental awareness. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Konik:2021:DSF

[2512] Marta Konik, Mirosław Darecki, Alexey K. Pavlov, Sławomir Sagan, and Piotr Kowalczuk. Darkening of the Svalbard fjords waters observed with satellite ocean color imagery in 1997–2019. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Garcia-Moreiras:2021:EFA

[2513] Iria García-Moreiras, Anabela Oliveira, Ana I. Santos, Paulo B. Oliveira, and Ana Amorim. Environmental factors affecting spatial dinoflagellate cyst distribution in surface sediments off Aveiro-Figueira da Foz (Atlantic Iberian Margin). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Ge:2021:CFL

[2514] Hailong Ge, Haoyu Zhang, Qingyuan Zhao, Fang Li, Haoran Gu, Siying Liu, Haibo Yang, Yun Li, and Zhijian Wang. Construction of a full-length transcriptome resource for the Chinese sucker (Myxocyprinus asiaticus), a rare protected fish, based on isoform sequencing (Iso-Seq). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Duncan:2021:PPS

[2515] Emily M. Duncan, Annette C. Broderick, Kay Critchell, Tamara S. Galloway, Mark Hamann, Colin J. Limpus, Penelope K. Lindeque, David Santillo, Anton D. Tucker, Scott Whiting, Erina J. Young, and Brendan J. Godley. Plastic pollution and small juvenile marine turtles: a potential evolutionary trap. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

deCarvalho:2021:ESI

[2516] Ricardo Cruz de Carvalho, Eduardo Feijão, Irina Duarte, Vanessa Pinto, Marisa Silva, Ana Rita Matos, Anabela Bernardes da Silva, Isabel Caçador, Patrick Reis-Santos, Vanessa F. Fonseca, and Bernardo Duarte. Ecoengineering solutions for the impairment of spreading and growth of invasive Spartina patens in Mediterranean salt marshes. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Machado:2021:RFW

[2517] Inês Machado, José Lino Costa, and Henrique Cabral. Response of foodwebs indicators to human pressures, in the scope of the marine strategy framework directive. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Quattrocchi:2021:TNA

[2518] Giovanni Quattrocchi, Andrea Cucco, Giulia Cerritelli, Resi Mencacci, Giorgia Comparetto, Daniela Sammartano, Alberto Ribotti, and Paolo Luschi. Testing a novel aggregated methodology to assess hydrodynamic impacts on a high-resolution marine turtle trajectory. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Robledo:2021:COR

[2519] Daniel Robledo, Erika Vázquez-Delfín, Yolanda Freile-Pelegrín, Román Manuel Vásquez-Elizondo, Zujaila Nohemy Qui-Minet, and Adán Salazar-Garibay. Challenges and opportunities in relation to sargassum events along the Caribbean Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Brzezinski:2021:NCP

[2520] Mark A. Brzezinski, Ivia Closset, Janice L. Jones, Gregory F. de Souza, and Colin Maden. New constraints on the physical and biological controls on the silicon isotopic composition of the Arctic Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Pain:2021:BHD

[2521] Andrea J. Pain, Jonathan B. Martin, and Caitlin R. Young. Biogeochemical and hydrological drivers of heterogeneous nutrient exports from subterranean estuaries. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Moura:2021:ICR

[2522] Amelia Moura, Brian Beck, Renee Duffey, Lucas McEachron, Margaret Miller, Jennifer Moore, Alison Moulding, and R. Scott Winters. Integrating coral restoration data with a novel coral sample registry. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhong:2021:DWN

[2523] Qiangqiang Zhong, Tao Yu, Hui Lin, Jing Lin, Jianda Ji, Jialin Ni, Jinzhou Du, and Dekun Huang. ^{210}Po–^{210}Pb disequilibrium in the Western North Pacific Ocean: Particle cycling and POC export. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Spence:2021:ULS

[2524] Michael A. Spence, Paul J. Dolder, Richard Nash, and Robert B. Thorpe. The use of a length-structured multispecies model fitted directly to data in near-real time as a viable tool for advice. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

delaVega:2021:BRS

[2525] Camille de la Vega, Claire Mahaffey, David J. Yurkowski, Louisa Norman, Elysia Simpson, Sophie Smout, Steven H. Ferguson, and Rachel M. Jeffreys. Biomarkers in ringed seals reveal recent onset of borealization in the high- compared to the mid-latitude Canadian Arctic. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Tamburini:2021:MNI

[2526] Marco Tamburini, Erica Keppel, Agnese Marchini, Michele F. Repetto, Gregory M. Ruiz, Jasmine Ferrario, and Anna Occhipinti-Ambrogi. Monitoring non-indigenous species in port habitats: First application of a standardized North American protocol in the Mediterranean Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Hu:2021:DEP

[2527] Huina Hu, Xiao Liu, Chunyan Ren, Renming Jia, Yusheng Qiu, Minfang Zheng, and Min Chen. ^{210}Po/^{210}Pb disequilibria and its estimate of particulate organic carbon export around Prydz Bay, Antarctica. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wang:2021:LFL

[2528] Shuangqiang Wang, Ellen Kenchington, Zeliang Wang, and Andrew J. Davies. Life in the fast lane: Modeling the fate of glass sponge larvae in the Gulf Stream. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Vilela:2021:UIL

[2529] Raul Vilela, Claudia Burger, Ansgar Diederichs, Fabian E. Bachl, Lesley Szostek, Anika Freund, Alexander Braasch, Jochen Bellebaum, Brian Beckers, Werner Piper, and Georg Nehls. Use of an INLA latent Gaussian modeling approach to assess bird population changes due to the development of offshore wind farms. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Seo:2021:CBP

[2530] Hojong Seo, DongJoo Joung, and Guebuem Kim. Contrasting behaviors of ^{210}Pb and ^{210}Po in the productive shelf water versus the oligotrophic water. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Krause-Jensen:2021:CIC

[2531] Dorte Krause-Jensen, Philippe Archambault, Jorge Assis, Inka Bartsch, Kai Bischof, Karen Filbee-Dexter, Kenneth H. Dunton, Olga Maximova, Sunna Björk Ragnarsdóttir, Mikael K. Sejr, Uliana Simakova, Vassily Spiridonov, Susse Wegeberg, Mie H. S. Winding, and Carlos M. Duarte. Corrigendum: Imprint of climate change on pan-Arctic marine vegetation. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See [1331].

Ayouche:2021:NLP

[2532] Adam Ayouche, Guillaume Charria, Xavier Carton, Nadia Ayoub, and Sébastien Theetten. Non-linear processes in the Gironde River plume (North-East Atlantic): Instabilities and mixing. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Voolstra:2021:CGA

[2533] Christian R. Voolstra, Kate M. Quigley, Sarah W. Davies, John Everett Parkinson, Raquel S. Peixoto, Manuel Aranda, Andrew C. Baker, Adam R. Barno, Daniel J. Barshis, Francesca Benzoni, Victor Bonito, David G. Bourne, Carol Buitrago-López, Tom C. L. Bridge, Cheong Xin Chan, David J. Combosch, Jamie Craggs, Jörg C. Frommlet, Santiago Herrera, Andrea M. Quattrini, Till Röthig, James D. Reimer, Esther Rubio-Portillo, David J. Suggett, Helena Villela, Maren Ziegler, and Michael Sweet. Consensus guidelines for advancing coral Holobiont genome and specimen voucher deposition. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Guragain:2021:PMC

[2534] Prashanna Guragain, Max Tkachov, Anna Solvang Båtnes, Yngvar Olsen, Per Winge, and Atle M. Bones. Principles and methods of counteracting harmful salmon–arthropod interactions in salmon farming: Addressing possibilities, limitations, and future options. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Aguirre:2021:HFV

[2535] Catalina Aguirre, René Garreaud, Lucy Belmar, Laura Farías, Laura Ramajo, and Facundo Barrera. High-frequency variability of the surface ocean properties off Central Chile during the upwelling season. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Panga:2021:TBE

[2536] Fleurdeliz M. Panga, Jonathan A. Anticamara, Miledel Christine C. Quibilan, Michael P. Atrigenio, and Porfirio M. Aliño. Through the boundaries: Environmental factors affecting reef benthic cover in Marine Protected Areas in the Philippines. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Huang:2021:DCP

[2537] Jianqiang Huang, Ruzhuo Zhong, Chuangye Yang, Qingheng Wang, Yongshan Liao, and Yuewen Deng. Dietary carbohydrate and protein levels affect the growth performance of juvenile peanut worm (Sipunculus nudus): an LC–MS-based metabolomics study. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sun:2021:ALO

[2538] Jiang Sun, Wu Men, Fenfen Wang, and Junwen Wu. Activity levels of 210Po, ^{210}Pb and other radionuclides (^{134}Cs, ^{137}Cs, ^{90}Sr, ^{110m}Ag, ^{238}U, 226Ra and ^{40}K) in marine organisms from coastal waters adjacent to Fuqing and Ningde Nuclear Power Plants (China) and radiation dose assessment. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Gijsman:2021:NBE

[2539] Rik Gijsman, Erik M. Horstman, Daphne van der Wal, Daniel A. Friess, Andrew Swales, and Kathelijne M. Wijnberg. Nature-based engineering: a review on reducing coastal flood risk with mangroves. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

vanderWalt:2021:MHE

[2540] Kerry-Ann van der Walt, Warren M. Potts, Francesca Porri, Alexander C. Winkler, Murray I. Duncan, Michael R. Skeeles, and Nicola C. James. Marine heatwaves exceed cardiac thermal limits of adult sparid fish (Diplodus capensis, Smith 1884). Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

Santamaria:2021:HIA

[2541] Jorge Santamaría, Fiona Tomas, Enric Ballesteros, and Emma Cebrian. Herbivory on the invasive alga Caulerpa cylindracea: The role of omnivorous fishes. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Karimi:2021:MSL

[2542] Armin Agha Karimi, Mohammad Bagherbandi, and Milan Horemuz. Multidecadal sea level variability in the Baltic Sea and its impact on acceleration estimations. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Smith:2021:MPB

[2543] Lauren Smith and William Richard Turrell. Monitoring plastic beach litter by number or by weight: The implications of fragmentation. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Goodbody-Gringley:2021:PPA

[2544] Gretchen Goodbody-Gringley, Federica Scucchia, Rebecca Ju, Alex Chequer, Shai Einbinder, Stephane Martinez, Hagai Nativ, and Tali Mass. Plasticity of Porites astreoides early life history stages suggests mesophotic coral ecosystems act as refugia in Bermuda. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Mokodongan:2021:SAA

[2545] Daniel Frikli Mokodongan, Hiroki Taninaka, La Sara, Taisei Kikuchi, Hideaki Yuasa, Yoshihisa Suyama, and Nina Yasuda. Spatial autocorrelation analysis using MIG-seq data indirectly estimated the gamete and larval dispersal range of the blue coral, Heliopora coerulea, within reefs. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Liu:2021:CEC

[2546] Feifei Liu, Ute Daewel, Annette Samuelsen, Sebastian Brune, Ulrike Hanz, Holger Pohlmann, Johanna Baehr, and Corinna Schrum. Can environmental conditions at North Atlantic deep-sea habitats be predicted several years ahead? — Taking sponge habitats as an example. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Yun:2021:QSR

[2547] Sukyoung Yun, Won Sang Lee, Robert P. Dziak, Lauren Roche, Haruyoshi Matsumoto, Tai-Kwan Lau, Angela Sremba, David K. Mellinger, Joseph H. Haxel, Seung-Goo Kang, Jong Kuk Hong, and Yongcheol Park. Quantifying soundscapes in the Ross Sea, Antarctica using long-term autonomous hydroacoustic monitoring systems. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Melet:2021:ECS

[2548] Angélique Melet, Carlo Buontempo, Matteo Mattiuzzi, Peter Salamon, Pierre Bahurel, George Breyiannis, Samantha Burgess, Laurence Crosnier, Pierre-Yves Le Traon, Lorenzo Mentaschi, Julien Nicolas, Lorenzo Solari, Freja Vamborg, and Evangelos Voukouvalas. European Copernicus services to inform on sea-level rise adaptation: Current status and perspectives. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Naidoo:2021:PRO

[2549] Ashley Naidoo, Hashali Hamukuaya, Mafaniso Hara, Yamkela Mngxe, and Jesper Raakjær. Polycentric regional ocean governance opportunity in the Benguela Current Convention. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lin:2021:MLC

[2550] Peng Lin, Chen Xu, Wei Xing, and Peter H. Santschi. Molecular level characterization of diatom and coccolithophore-associated biopolymers that are binding ^{210}Pb and ^{210}Po in seawater. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745.

pien:2021:HSR

[2551] Anna St pień, Krzysztof Pabis, Robert Sobczyk, and Bjorn Serigstad. High species richness and extremely low abundance of cumacean communities along the shelf and slope of the Gulf of Guinea (West Africa). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Miksis-Olds:2021:OSA

[2552] Jennifer L. Miksis-Olds, Peter J. Dugan, S. Bruce Martin, Holger Klinck, David K. Mellinger, David A. Mann, Dimitri W. Ponirakis, and Olaf Boebel. Ocean sound analysis software for making ambient noise trends accessible (MANTA). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Wall:2021:NWP

[2553] Carrie C. Wall, Samara M. Haver, Leila T. Hatch, Jennifer Miksis-Olds, Rob Bochenek, Robert P. Dziak, and Jason Gedamke. The next wave of passive acoustic data management: How centralized access can enhance science. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Masucci:2021:CSS

[2554] Giovanni D. Masucci, Piera Biondi, and James D. Reimer. A comparison of size, shape, and fractal diversity between coral rubble sampled from natural and artificial coastlines around Okinawa Island, Japan. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

McElligott:2021:ISD

[2555] Megan M. McElligott and Marc O. Lammers. Investigating spinner dolphin (Stenella longirostris) occurrence and acoustic activity in the Maui Nui region. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Neves:2021:TDG

[2556] João Neves, Eva Giacomello, Gui M. Menezes, Jorge Fontes, and Susanne E. Tanner. Temperature-driven growth variation in a deep-sea fish: The case of Pagellus bogaraveo (Brünnich, 1768) in the Azores Archipelago. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lawrence:2021:CST

[2557] Judy Lawrence, Scott Stephens, Paula Blackett, Robert G. Bell, and Rebecca Priestley. Climate services transformed: Decision-making practice for the coast in a changing climate. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Proulx:2021:ITE

[2558] MaryJane Proulx, Lydia Ross, Christina Macdonald, Shayla Fitzsimmons, and Michael Smit. Indigenous traditional ecological knowledge and ocean observing: a review of successful partnerships. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Schmidt:2021:RDA

[2559] Nicole Schmidt, Timothy Dellapenna, and Peng Lin. ^{7}Be/^{210}Pb_{xs} ratioderived age and residence time of suspended sediments in Galveston Bay. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Peltier:2021:CWP

[2560] Helene Peltier, Matthieu Authier, Florence Caurant, Willy Dabin, Pierre Daniel, Cecile Dars, Fabien Demaret, Eleonore Meheust, Olivier Van Canneyt, Jerome Spitz, and Vincent Ridoux. Corrigendum: In the wrong place at the wrong time: Identifying spatiotemporal co-occurrence of bycaught common dolphins and fisheries in the Bay of Biscay (NE Atlantic) from 2010 to 2019. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See [1548].

Islam:2021:SFC

[2561] Mohammad Mahmudul Islam, Amany Begum, Syed Mohammad Aminur Rahman, and Hadayet Ullah. Seasonal fishery closure in the Northern Bay of Bengal causes immediate but contrasting ecological and socioeconomic impacts. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Livernois:2021:OPE

[2562] Mariah C. Livernois, John A. Mohan, Thomas C. TinHan, Travis M. Richards, Brett J. Falterman, Nathan R. Miller, and R. J. David Wells. Ontogenetic patterns of elemental tracers in the vertebrae cartilage of coastal and oceanic sharks. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lee:2021:SRD

[2563] Jeffery Liang-Neng Lee, Kuo-Ping Chiang, and Sheng-Fang Tsai. Sexual reproduction in dinoflagellates — the case of Noctiluca scintillans and its ecological implications. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Choi:2021:ETS

[2564] Seo Yeol Choi, Eun Hye Lee, Ho Young Soh, and Min-Chul Jang. Effects of temperature and salinity on egg production, hatching, and mortality rates in Acartia ohtsukai (Copepoda, Calanoida). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Tanedo:2021:IIE

[2565] Mikhael Clotilde S. Tañedo, Ronald D. Villanueva, Andrew F. Torres, Rachel Ravago-Gotanco, and Maria Lourdes San Diego-McGlone. Individual and interactive effects of ocean warming and acidification on adult favites colemani. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cvitanovic:2021:TCR

[2566] Christopher Cvitanovic, Carina Wyborn, Evora Glenn, Rachel Kelly, Elena Louder, Elizabeth I. van Putten, and Angela Bednarek. Ten considerations for research funders seeking to enhance knowledge exchange and the impact of marine science on policy and practice. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Dao:2021:IST

[2567] Hung N. Dao, Hang T. Vu, Susan Kay, and Sevrine Sailley. Impact of seawater temperature on coral reefs in the context of climate change. A case study of Cu Lao Cham-Hoi An Biosphere Reserve. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Marhoefer:2021:SAA

[2568] Shelby R. Marhoefer, Kyall R. Zenger, Jan M. Strugnell, Murray Logan, Madeleine J. H. van Oppen, Carly D. Kenkel, and Line K. Bay. Signatures of adaptation and acclimatization to reef flat and slope habitats in the coral Pocillopora damicornis. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Legeais:2021:CSL

[2569] Jean-François Legeais, Benoît Meyssignac, Yannice Faugère, Adrien Guerou, Michaël Ablain, Marie-Isabelle Pujol, Claire Dufau, and Gérald Dibarboure. Copernicus sea level space observations: a basis for assessing mitigation and developing adaptation strategies to sea level rise. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Gjerde:2021:PRO

[2570] Kristina M. Gjerde and Siddharth Shekhar Yadav. Polycentricity and regional ocean governance: Implications for the emerging UN agreement on marine biodiversity beyond national jurisdiction. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Himmelstein:2021:MPE

[2571] Joshua Himmelstein, Orencio Duran Vinent, Stijn Temmerman, and Matthew L. Kirwan. Mechanisms of pond expansion in a rapidly submerging marsh. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Valsecchi:2021:FED

[2572] Elena Valsecchi, Antonella Arcangeli, Roberto Lombardi, Elizabeth Boyse, Ian M. Carr, Paolo Galli, and Simon J. Goodman. Ferries and environmental DNA: Underway sampling from commercial vessels provides new opportunities for systematic genetic surveys of marine biodiversity. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Sun:2021:CSL

[2573] Jing Sun, Mingqiang Chen, Zhengyi Fu, Jingru Yang, Shengjie Zhou, Gang Yu, Wenli Zhou, and Zhenhua Ma. A comparative study on low and high salinity tolerance of two strains of Pinctada fucata. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Sole-Jimenez:2021:RGM

[2574] Paula Solé-Jiménez, Fernando Naya-Català, M. Carla Piazzon, Itziar Estensoro, Josep À. Calduch-Giner, Ariadna Sitjà-Bobadilla, Danny Van Mullem, and Jaume Pérez-Sánchez. Reshaping of gut microbiota in gilthead sea bream fed microbial and processed animal proteins as the main dietary protein source. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Coguiec:2021:SVZ

[2575] Estelle Coguiec, Elizaveta A. Ershova, Malin Daase, Tobias R. Vonnahme, Owen S. Wangensteen, Rolf Gradinger, Kim Præbel, and Jørgen Berge. Seasonal variability in the zooplankton community structure in a sub-Arctic fjord as revealed by morphological and molecular approaches. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Freeman:2021:SMI

[2576] Christopher J. Freeman, Cole G. Easson, Cara L. Fiore, and Robert W. Thacker. Sponge–microbe interactions on coral reefs: Multiple evolutionary solutions to a complex environment. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Benoit:2021:CTE

[2577] Jonathan Benoit and Baylor Fox-Kemper. Contextualizing thermal effluent impacts in Narragansett Bay using Landsat-derived surface temperature. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Bribiesca-Contreras:2021:BCA

[2578] Guadalupe Bribiesca-Contreras, Thomas G. Dahlgren, Tammy Horton, Jeffrey C. Drazen, Regan Drennan, Daniel O. B. Jones, Astrid B. Leitner, Kirsty A. McQuaid, Craig R. Smith, Sergio Taboada, Helena Wiklund, and Adrian G. Glover. Biogeography and connectivity across habitat types and geographical scales in Pacific abyssal scavenging amphipods. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Yin:2021:DAD

[2579] Bin Yin, Hongyu Liu, Beiping Tan, Xiaohui Dong, Shuyan Chi, Qihui Yang, and Shuang Zhang. Dynamic alterations of the distal intestinal microbiota, transcriptome, and metabolome of hybrid grouper by β - conglycinin with reconciliations by sodium butyrate in feed. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Poppeschi:2021:USE

[2580] Coline Poppeschi, Guillaume Charria, Eric Goberville, Peggy Rimmelin-Maury, Nicolas Barrier, Sébastien Petton, Maximilian Unterberger, Emilie Grossteffan, Michel Repecaud, Loïc Quéméner, Sébastien Theetten, Jean-François Le Roux, and Paul Tréguer. Unraveling salinity extreme events in coastal environments: a winter focus on the Bay of Brest. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Martinez-Quintana:2021:ELH

[2581] Ángela Martínez-Quintana and Howard R. Lasker. Early life-history dynamics of Caribbean octocorals: The critical role of larval supply and partial mortality. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pnyushkov:2021:SRV

[2582] A. V. Pnyushkov, I. V. Polyakov, G. V. Alekseev, I. M. Ashik, T. M. Baumann, E. C. Carmack, V. V. Ivanov, and R. Rember. A steady regime of volume and heat transports in the Eastern Arctic Ocean in the early 21st Century. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Arossa:2021:IMS

[2583] Silvia Arossa, Alan Barozzi, Matteo Callegari, Shannon G. Klein, Anieka J. Parry, Shiou-Han Hung, Alexandra Steckbauer, Manuel Aranda, Daniele Daffonchio, and Carlos M. Duarte. The internal microenvironment of the symbiotic jellyfish Cassiopea sp. from the Red Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Simm:2021:ECC

[2584] Jonathan Simm, Ben Gouldby, Darren Lumbroso, and Tom Matthewson. Effective coastal climate services — an end-user perspective for resilient infrastructure. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Miller:2021:CND

[2585] K. A. Miller, K. Brigden, D. Santillo, D. Currie, P. Johnston, and K. F. Thompson. Challenging the need for deep seabed mining from the perspective of metal demand, biodiversity, ecosystems services, and benefit sharing. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Bashar:2021:EDS

[2586] Abul Bashar, Neaz A. Hasan, Mohammad Mahfujul Haque, Md. Fazle Rohani, and Md. Sazzad Hossain. Effects of dietary silica nanoparticle on growth performance, protein digestibility, hematology, digestive morphology, and muscle composition of Nile tilapia, Oreochromis niloticus. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kawamura:2021:VSR

[2587] Kaz Kawamura, Satoko Sekida, Koki Nishitsuji, Eiichi Shoguchi, Kanako Hisata, Shigeki Fujiwara, and Noriyuki Satoh. In vitro symbiosis of reef-building coral cells with photosynthetic dinoflagellates. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See corrigendum [3201].

Wang:2021:VEB

[2588] Sikai Wang, Qiang Sheng, Feng Zhao, Tingting Zhang, and Ping Zhuang. Variable effects on benthic community from diking to eradicate invasive plants in the Yangtze Estuary salt marsh. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chen:2021:EDB

[2589] Yongkang Chen, Shuyan Chi, Shuang Zhang, Xiaohui Dong, Qihui Yang, Hongyu Liu, Beiping Tan, and Shiwei Xie. Evaluation of the dietary black soldier fly larvae meal (Hermetia illucens) on growth performance, intestinal health, and disease resistance to Vibrio parahaemolyticus of the Pacific white shrimp (Litopenaeus vannamei). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Fabrizio:2021:ESS

[2590] Mary C. Fabrizio, Troy D. Tuckey, Aaron J. Bever, and Michael L. MacWilliams. The extent of seasonally suitable habitats may limit forage fish production in a temperate estuary. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pierce:2021:CDM

[2591] Jordan Pierce, Mark J. Butler, Yuri Rzhanov, Kim Lowell, and Jennifer A. Dijkstra. Classifying 3-D models of coral reefs using structurefrom-motion and multi-view semantic segmentation. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Castellanos:2021:CDM

[2592] Paola Castellanos, Ana C. Brito, Paula Chainho, Bernardo R. Quintella, Luis da Costa, Susana França, Henrique Cabral, and José Lino Costa. CoastNet dataset from Mondego, Tejo and Mira Estuaries: Multiparametric measurements during 2020. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Gallagher:2021:FAD

[2593] Austin J. Gallagher, Nourah A. Alsudairy, Brendan D. Shea, Nicholas L. Payne, and Carlos M. Duarte. First application of 360-degree camera technology to marine predator bio-logging. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wang:2021:CPE

[2594] Qiang Wang, Sergey Danilov, Dmitry Sidorenko, and Xuezhu Wang. Circulation pathways and exports of Arctic river runoff influenced by atmospheric circulation regimes. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chen:2021:SCO

[2595] Guofeng Chen, Bin Yin, Hongyu Liu, Beiping Tan, Xiaohui Dong, Qihui Yang, Shuyan Chi, and Shuang Zhang. Supplementing chitosan oligosaccharide positively affects hybrid grouper (Epinephelus fuscoguttatus [Venus sign] × E. lanceolatus [male sign]) fed dietary fish meal replacement with cottonseed protein concentrate: Effects on growth, gut microbiota, antioxidant function and immune response. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Alvera-Azcarate:2021:AYD

[2596] Aida Alvera-Azcárate, Dimitry Van der Zande, Alexander Barth, Charles Troupin, Samuel Martin, and Jean-Marie Beckers. Analysis of 23 years of daily cloud-free chlorophyll and suspended particulate matter in the Greater North Sea. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wiesebron:2021:SBD

[2597] Lauren E. Wiesebron, Natalie Steiner, Claudia Morys, Tom Ysebaert, and Tjeerd J. Bouma. Sediment bulk density effects on benthic macrofauna burrowing and bioturbation behavior. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Lai:2021:PBR

[2598] Chao-Chen Lai, Chau-Ron Wu, Chia-Ying Chuang, Jen-Hua Tai, Kuo-Yuan Lee, Hsiang-Yi Kuo, and Fuh-Kwo Shiah. Phytoplankton and bacterial responses to monsoon-driven water masses mixing in the Kuroshio off the East Coast of Taiwan. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Ibanez:2021:MTO

[2599] Christian M. Ibáñez, Mariana Díaz-Santana-Iturrios, Sergio A. Carrasco, Fernando A. Fernández-Álvarez, David A. López-Córdova, Claudio F. Cornejo, Nicolás Ortiz, Francisco Rocha, Erica A. G. Vidal, and María Cecilia Pardo-Gandarillas. Macroevolutionary trade-offs and trends in life history traits of cephalopods through a comparative phylogenetic approach. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Thorpe:2021:CCE

[2600] Robert B. Thorpe, Michael A. Spence, Paul J. Dolder, and Richard D. M. Nash. Commentary: Combining ecosystem and single-species modeling to provide ecosystem-based fisheries management advice within current management systems. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745. See [1279].

Cavanagh:2021:CFR

[2601] Rachel D. Cavanagh, Jess Melbourne-Thomas, Susie M. Grant, David K. A. Barnes, Kevin A. Hughes, Svenja Halfter, Michael P. Meredith, Eugene J. Murphy, Rowan Trebilco, and Simeon L. Hill. Corrigendum: Future risk for Southern Ocean ecosystem services under climate change. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745. See [1322].

Zeller:2021:FEA

[2602] Dirk Zeller, Gabriel M. S. Vianna, Matthew Ansell, Angie Coulter, Brittany Derrick, Krista Greer, Simon-Luc Noël, Maria L. Deng Palomares, Audrey Zhu, and Daniel Pauly. Fishing effort and associated catch per unit effort for small-scale fisheries in the Mozambique Channel region: 1950–2016. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Xia:2021:ETS

[2603] Yuguo Xia, Xinhui Li, Jiping Yang, Shuli Zhu, Zhi Wu, Jie Li, and Yuefei Li. Elevated temperatures shorten the spawning period of silver carp (Hypophthalmichthys molitrix) in a large subtropical river in China. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Liu:2021:EWC

[2604] Bing Liu, Yining Chen, Tinglu Cai, Yan Li, and Li Sun. Estimating waves and currents at the saltmarsh edge using acoustic Doppler velocimeter data. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Anonymous:2021:ECN

[2605] Anonymous. Erratum: A Convolutional Neural Network for Automated Detection of Humpback Whale Song in a Diverse, Long-Term Passive Acoustic Dataset. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745. See [1470].

Zainescu:2021:RJV

[2606] Florin Zăinescu, Edward Anthony, and Alfred Vespremeanu-Stroe. River jets versus wave-driven longshore currents at river mouths. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Harrison:2021:RGM

[2607] Jesse P. Harrison, Panagiota-Myrsini Chronopoulou, Iines S. Salonen, Tom Jilbert, and Karoliina A. Koho. 16S and 18S rRNA gene metabarcoding provide congruent information on the responses of sediment communities to eutrophication. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Wang:2021:CBM

[2608] Han Wang, Qiwen Zheng, Shuang Li, and Xiaoyu Yang. The closer the better? Modeling spatial spillover effects of ecological externalities in coastal mangroves. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Russo:2021:URB

[2609] Stefania Russo, Marco Torri, Bernardo Patti, Patricia Reglero, Diego Álvarez-Berastegui, Angela Cuttitta, and Gianluca Sarà. Unveiling the relationship between sea surface hydrographic patterns and tuna larval distribution in the Central Mediterranean Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Nativ:2021:SEC

[2610] Hagai Nativ, Federica Scucchia, Stephane Martinez, Shai Einbinder, Alex Chequer, Gretchen Goodbody-Gringley, and Tali Mass. In situ estimation of coral recruitment patterns from shallow to mesophotic reefs using an optimized fluorescence imaging system. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Cebrian:2021:RRM

[2611] Emma Cebrian, Laura Tamburello, Jana Verdura, Giuseppe Guarnieri, Alba Medrano, Cristina Linares, Bernat Hereu, Joaquim Garrabou, Carlo Cerrano, Cristina Galobart, and Simonetta Fraschetti. A roadmap for the restoration of Mediterranean macroalgal forests. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Crosman:2021:CCA

[2612] Katherine M. Crosman, Gerald G. Singh, and Sabine Lang. Confronting complex accountability in conservation with communities. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Weber:2021:SCC

[2613] Edward D. Weber, Toby D. Auth, Simone Baumann-Pickering, Timothy R. Baumgartner, Eric P. Bjorkstedt, Steven J. Bograd, Brian J. Burke, José L. Cadena-Ramírez, Elizabeth A. Daly, Martin de la Cruz, Heidi Dewar, John C. Field, Jennifer L. Fisher, Ashlyn Giddings, Ralf Goericke, Eliana Gomez-Ocampo, Jose Gomez-Valdes, Elliot L. Hazen, John Hildebrand, Cheryl A. Horton, Kym C. Jacobson, Michael G. Jacox, Jaime Jahncke, Mati Kahru, Raphe M. Kudela, Bertha E. Lavaniegos, Andrew Leising, Sharon R. Melin, Luis Erasmo Miranda-Bojorquez, Cheryl A. Morgan, Catherine F. Nickels, Rachael A. Orben, Jessica M. Porquez, Elan J. Portner, Roxanne R. Robertson, Daniel L. Rudnick, Keith M. Sakuma, Jarrod A. Santora, Isaac D. Schroeder, Owyn E. Snodgrass, William J. Sydeman, Andrew R. Thompson, Sarah Ann Thompson, Jennifer S. Trickey, Josue Villegas-Mendoza, Pete Warzybok, William Watson, and Samantha M. Zeman. State of the California Current 2019–2020: Back to the future with marine heatwaves? Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Croquer:2021:SDB

[2614] Aldo Cróquer, Ernesto Weil, and Caroline S. Rogers. Similarities and differences between two deadly Caribbean coral diseases: White plague and Stony Coral Tissue Loss Disease. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Rowlands:2021:ECO

[2615] Emily Rowlands, Tamara Galloway, Matthew Cole, Ceri Lewis, Victoria Peck, Sally Thorpe, and Clara Manno. The effects of combined ocean acidification and nanoplastic exposures on the embryonic development of Antarctic krill. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Rosenau:2021:CCV

[2616] Nicholas A. Rosenau, Sarah Gignoux-Wolfsohn, Richard A. Everett, A. Whitman Miller, Mark S. Minton, and Gregory M. Ruiz. Considering commercial vessels as potential vectors of Stony Coral Tissue Loss Disease. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Li:2021:BCU

[2617] Dewang Li, Xiaobo Ni, Kui Wang, Dingyong Zeng, Bin Wang, Haiyan Jin, Hongliang Li, Feng Zhou, Daji Huang, and Jianfang Chen. Biological CO_{2} uptake and upwelling regulate the air–sea CO_{2} flux in the Changjiang plume under south winds in summer. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Singh:2021:GLS

[2618] Gerald G. Singh, Richard S. Cottrell, Tyler D. Eddy, and Andrés Miguel Cisneros-Montemayor. Governing the land–sea interface to achieve sustainable coastal development. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Ferguson:2021:DPM

[2619] Steven H. Ferguson, Jeff W. Higdon, Patricia A. Hall, Rikke Guldborg Hansen, and Thomas Doniol-Valcroze. Developing a precautionary management approach for the Eastern Canada–West Greenland population of bowhead whales (Balaena mysticetus). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:DBP

[2620] Huangchen Zhang, Linbin Zhou, Kaizhi Li, Zhixin Ke, and Yehui Tan. Decreasing biological production and carbon export due to the barrier layer: a case study in the Bay of Bengal. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Campbell:2021:VRF

[2621] Brooke Campbell, Alistair McIlgorm, and Michelle Voyer. Virtual regional fisheries management meetings in the Western and Central Pacific Ocean region: a preliminary appraisal of performance and fit. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Zhao:2021:AIC

[2622] Kangshun Zhao, Min Zhang, Kang Wang, Konghao Zhu, Congjun Xu, Jiayi Xie, and Jun Xu. Aquaculture impacts on China’s marine wild fisheries over the past 30 years. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

El-Tarabily:2021:MEP

[2623] Khaled A. El-Tarabily, Gaber A. Ramadan, Alaa A. Elbadawi, Amira H. Hassan, Saeed Tariq, Eman W. Ghazal, Mohammad I. Abo Gamar, and Synan F. AbuQamar. The marine endophytic polyamine-producing Streptomyces mutabilis UAE1 isolated from extreme niches in the Arabian Gulf promotes the performance of mangrove (Avicennia marina) seedlings under greenhouse conditions. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Nicholls:2021:CLR

[2624] Robert J. Nicholls, Richard P. Beaven, Anne Stringfellow, Daniel Monfort, Gonéri Le Cozannet, Thomas Wahl, Julia Gebert, Matthew Wadey, Arne Arns, Kate L. Spencer, Debra Reinhart, Timo Heimovaara, Víctor M. Santos, Alejandra R. Enríquez, and Samantha Cope. Coastal landfills and rising sea levels: a challenge for the 21st Century. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wang:2021:ISS

[2625] Rui Wang, Puqing Song, Yuan Li, and Longshan Lin. An integrated, size-structured stock assessment of Antarctic krill, Euphausia superba. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Mankad:2021:GEC

[2626] Aditi Mankad, Elizabeth V. Hobman, and Lucy Carter. Genetically engineering coral for conservation: Psychological correlates of public acceptability. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Mai:2021:ART

[2627] Guangming Mai, Jihua Liu, Xiaomin Xia, Xinyue Pang, Binkai Li, Linghui Yu, Yehui Tan, Xingyu Song, and Gang Li. Acutely rising temperature reduces photosynthetic capacity of phytoplankton assemblages in tropical oceans: a large-scale investigation. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Steinberg:2021:WIC

[2628] Rosemary Kate Steinberg, Emma L. Johnston, Teresa Bednarek, Katherine A. Dafforn, and Tracy D. Ainsworth. Its what’s on the inside that counts: an effective, efficient, and streamlined method for quantification of octocoral symbiodiniaceae and chlorophyll. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Kang:2021:WQI

[2629] Yoonja Kang, Chang-Ho Moon, Hyun-Jung Kim, Yang Ho Yoon, and Chang-Keun Kang. Water quality improvement shifts the dominant phytoplankton group from cryptophytes to diatoms in a coastal ecosystem. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lima:2021:CSB

[2630] Leonardo Lima, Stefania Angela Ciliberti, Ali Aydoğdu, Simona Masina, Romain Escudier, Andrea Cipollone, Diana Azevedo, Salvatore Causio, Elisaveta Peneva, Rita Lecci, Emanuela Clementi, Eric Jansen, Mehmet Ilicak, Sergio Cretì, Laura Stefanizzi, Francesco Palermo, and Giovanni Coppini. Climate signals in the Black Sea from a multidecadal eddyresolving reanalysis. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Magris:2021:ELS

[2631] Rafael Almeida Magris. Effectiveness of large-scale Marine Protected Areas in the Atlantic Ocean for reducing fishing activities. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

ORegan:2021:GAC

[2632] Sacha M. O’Regan, Stephanie K. Archer, Sarah K. Friesen, and Karen L. Hunter. A global assessment of climate change adaptation in Marine Protected Area management plans. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Callejas:2021:CEC

[2633] Ileana A. Callejas, Christine M. Lee, Deepak R. Mishra, Stacey L. Felgate, Claire Evans, Abel Carrias, Andria Rosado, Robert Griffin, Emil A. Cherrington, Mariam Ayad, Megha Rudresh, Benjamin P. Page, and Jennifer A. Jay. Corrigendum: Effect of COVID-19 anthropause on water clarity in the Belize Coastal Lagoon. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745. See [1911].

Muras:2021:EAF

[2634] Andrea Muras, Severine Larroze, Celia Mayer, Tânia Teixeira, Reut Wengier, Yehuda Benayahu, and Ana Otero. Evaluation of the antifouling efficacy of Bacillus licheniformis extracts under environmental and natural conditions. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

McCabe:2021:EMK

[2635] Elizabeth J. Berens McCabe, Randall S. Wells, Christina N. Toms, Aaron A. Barleycorn, Krystan A. Wilkinson, and Valeriy I. Palubok. Effects of multiple Karenia brevis Red Tide blooms on a common bottlenose dolphin (Tursiops truncatus) prey fish assemblage: Patterns of resistance and resilience in Sarasota Bay, Florida. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Melbourne-Thomas:2021:DRN

[2636] Jess Melbourne-Thomas, Keith R. Hayes, Alistair J. Hobday, L. Richard Little, Joanna Strzelecki, Damian P. Thomson, Ingrid van Putten, and Sharon E. Hook. Decommissioning research needs for offshore oil and gas infrastructure in Australia. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Jefferson:2021:PPO

[2637] Rebecca Jefferson, Emma McKinley, Holly Griffin, Alison Nimmo, and Stephen Fletcher. Public perceptions of the ocean: Lessons for marine conservation from a global research review. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Hughes:2021:DNB

[2638] Adam D. Hughes. Defining nature-based solutions within the blue economy: The example of aquaculture. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Comesana:2021:AAC

[2639] Sara Comesaña, Floriana Lai, Ann-Elise Olderbakk Jordal, Tiziano Verri, Marit Espe, José L. Soengas, and Ivar Rønnestad. Amino acid carriers of the solute carrier families 7 (SLC7) and 38 (SLC38) are involved in leucine sensing in the brain of Atlantic salmon (Salmo salar). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bennett:2021:ASE

[2640] Nathan J. Bennett, Laure Katz, Whitney Yadao-Evans, Gabby N. Ahmadia, Scott Atkinson, Natalie C. Ban, Neil M. Dawson, Asha de Vos, Juno Fitzpatrick, David Gill, Mael Imirizaldu, Naia Lewis, Sangeeta Mangubhai, Leah Meth, Ella-Kari Muhl, David Obura, Ana K. Spalding, Angelo Villagomez, Daniel Wagner, Alan White, and Aulani Wilhelm. Advancing social equity in and through marine conservation. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Borchert:2021:GGE

[2641] Florian Borchert, Iraj Emadodin, Christof Kluß, Ana Rotter, and Thorsten Reinsch. Grass growth and N_{2}O emissions from soil after application of jellyfish in coastal areas. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lorgen-Ritchie:2021:TDG

[2642] Marlene Lorgen-Ritchie, Michael Clarkson, Lynn Chalmers, John F. Taylor, Herve Migaud, and Samuel A. M. Martin. A temporally dynamic gut microbiome in Atlantic salmon during freshwater recirculating aquaculture system (RAS) production and post-seawater transfer. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Isada:2021:DNC

[2643] Tomonori Isada, Hiroya Abe, Hiromi Kasai, and Masahiro Nakaoka. Dynamics of nutrients and colored dissolved organic matter absorption in a wetland-influenced subarctic coastal region of Northeastern Japan: Contributions from mariculture and eelgrass meadows. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ying:2021:PRP

[2644] Long Ying, Sutinee Sinutok, Pathompong Pramneechote, Pattara Aiyarak, Peter J. Ralph, and Ponlachart Chotikarn. Physiological responses of Pocillopora acuta and Porites lutea under plastic and fishing net stress. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Comesana:2021:MPG

[2645] Antonio Comesaña, Bieito Fernández-Castro, Paloma Chouciño, Emilio Fernández, Antonio Fuentes-Lema, Miguel Gilcoto, María Pérez-Lorenzo, and Beatriz Mouriño-Carballido. Mixing and phytoplankton growth in an upwelling system. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Li:2021:MLR

[2646] Yifan Li, Siguang Liu, Mengyang Liu, Wei Huang, Kai Chen, Yongcheng Ding, Fangzhu Wu, Hongwei Ke, Linghao Lou, Yan Lin, Mingyu Zhang, Fengjiao Liu, Chunhui Wang, and Minggang Cai. Mid-level riverine outflow matters: a case of microplastic transport in the Jiulong River, China. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Villasante:2021:ECT

[2647] Sebastián Villasante, Ana Tubío, Ignacio Gianelli, Pablo Pita, and Antonio García-Allut. Ever changing times: Sustainability transformations of Galician small-scale fisheries. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Morganti:2021:CSP

[2648] Teresa Maria Morganti, Marta Ribes, Raz Moskovich, Jeremy Brian Weisz, Gitai Yahel, and Rafel Coma. Corrigendum: In situ pumping rate of 20 marine demosponges is a function of osculum area. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See [1403].

Spijkerboer:2021:IDI

[2649] Rozanne C. Spijkerboer. The institutional dimension of integration in marine spatial planning: The case of the Dutch North Sea dialogues and agreement. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Thomas:2021:GRL

[2650] Elin A. Thomas, Aoife Molloy, Nova B. Hanson, Monika Böhm, Mary Seddon, and Julia D. Sigwart. A global red list for hydrothermal vent molluscs. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Chen:2021:CSS

[2651] Dake Chen, Jinhai Zheng, Chi Zhang, Dawei Guan, Yuan Li, and Yigang Wang. Critical shear stress for erosion of sand-mud mixtures and pure mud. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Martindale:2021:EGM

[2652] Rowan C. Martindale, Daniel Holstein, Nancy Knowlton, Joshua D. Voss, Anna M. Weiss, and Adrienne M. S. Correa. Editorial: Gulf of Mexico reefs: Past, present and future. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Robeck:2021:MTM

[2653] Todd R. Robeck, Zhe Fei, Amin Haghani, Joseph A. Zoller, Caesar Z. Li, Karen J. Steinman, Stacy Dirocco, Lydia Staggs, Todd Schmitt, Steve Osborn, Gisele Montano, Magdalena Rodriguez, and Steve Horvath. Multi-tissue methylation clocks for age and sex estimation in the common bottlenose dolphin. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Owens:2021:MTT

[2654] Katharine A. Owens and Katie Conlon. Mopping up or turning off the tap? Environmental injustice and the ethics of plastic pollution. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kho:2021:CGL

[2655] Kang Hee Kho, Zahid Parvez Sukhan, Shaharior Hossen, Yusin Cho, Soo Cheol Kim, Md. Rajib Sharker, Hee-Jeong Jung, and Ill-Sup Nou. Construction of a genetic linkage map based on SNP markers, QTL mapping and detection of candidate genes of growth-related traits in Pacific abalone using genotyping-by-sequencing. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Iwasaki:2021:EWI

[2656] Shinsuke Iwasaki and Junichi Otsuka. Evaluation of wave-ice parameterization models in WAVEWATCH III(R) along the coastal area of the Sea of Okhotsk during winter. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Tao:2021:NPD

[2657] Wei Tao, Lixia Niu, Yanhong Dong, Tao Fu, and Quansheng Lou. Nutrient pollution and its dynamic source–sink pattern in the Pearl River Estuary (South China). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Ting:2021:LRS

[2658] Seng Yeat Ting, Nyok-Sean Lau, Ka-Kei Sam, Evan S. H. Quah, Amirrudin B. Ahmad, Mohd-Noor Mat-Isa, and Alexander Chong Shu-Chien. Long-read sequencing reveals the repertoire of long-chain polyunsaturated fatty acid biosynthetic genes in the purple land crab, Gecarcoidea lalandii (H. Milne Edwards, 1837). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Kirkfeldt:2021:RSC

[2659] Trine Skovgaard Kirkfeldt and Catarina Frazão Santos. A review of sustainability concepts in marine spatial planning and the potential to supporting the UN sustainable development goal 14. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Huang:2021:PPT

[2660] Ruhui Huang, Xiaohui Xie, Jianyu Hu, and Zhenyu Sun. Poleward propagation of typhoon-induced near-inertial waves in the Northern South China Sea. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Meissa:2021:CTD

[2661] Beyah Meissa, Mamadou Dia, Braham C. Baye, Moustapha Bouzouma, Ely Beibou, and Rubén H. Roa-Ureta. A comparison of three data-poor stock assessment methods for the pink spiny lobster fishery in Mauritania. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bass:2021:ADM

[2662] Alissa Bass, Thomas Wernberg, Mads Thomsen, and Dan Smale. Another decade of marine climate change experiments: Trends, progress and knowledge gaps. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Ballesteros:2021:CMF

[2663] Ainara Ballesteros, Carina Östman, Andreu Santín, Macarena Marambio, Mridvika Narda, and Josep-Maria Gili. Cnidome and morphological features of Pelagia noctiluca (Cnidaria: Scyphozoa) throughout the different life cycle stages. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Taninaka:2021:PBC

[2664] Hiroki Taninaka, Davide Maggioni, Davide Seveso, Danwei Huang, Abram Townsend, Zoe T. Richards, Sen-Lin Tang, Naohisa Wada, Taisei Kikuchi, Hideaki Yuasa, Megumi Kanai, Stéphane De Palmas, Niphon Phongsuwan, and Nina Yasuda. Phylogeography of blue corals (genus Heliopora) across the Indo-West Pacific. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Morales-Nin:2021:CMW

[2665] Beatriz Morales-Nin, Robert Arlinghaus, and Josep Alós. Contrasting the motivations and wildlife-related value orientations of recreational fishers with participants of other outdoor and indoor recreational activities. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Deutsch:2021:MHS

[2666] Jessica M. Deutsch, Olakunle A. Jaiyesimi, Kelly A. Pitts, Jay Houk, Blake Ushijima, Brian K. Walker, Valerie J. Paul, and Neha Garg. Metabolomics of healthy and Stony Coral Tissue Loss Disease affected Montastraea cavernosa corals. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Galimany:2021:BFB

[2667] Eve Galimany, Jessica Lunt, Christopher J. Freeman, I. Segura-García, M. Mossop, A. Domingos, J. Houk, and Valerie J. Paul. Bivalve feeding on the brown tide Aureoumbra lagunensis in a shallow coastal environment. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

OHalloran:2021:HRB

[2668] Chris O’Halloran and Mary Silver. Health risks and benefits among surfers after exposure to seawater in Monterey Bay, Santa Cruz County, California, United States. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Terova:2021:UGP

[2669] Genciana Terova, Federico Moroni, Micaela Antonini, Stefano Bertacchi, Chiara Pesciaroli, Paola Branduardi, Massimo Labra, Danilo Porro, Chiara Ceccotti, and Simona Rimoldi. Using glycerol to produce European sea bass feed with oleaginous microbial biomass: Effects on growth performance, filet fatty acid profile, and FADS2 gene expression. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

El-Tarabily:2021:CRC

[2670] Khaled A. El-Tarabily, Arjun Sham, Alaa A. Elbadawi, Amira H. Hassan, Bashaer K. K. Alhosani, Mohamed A. El-Esawi, Abdulmajeed S. AlKhajeh, and Synan F. AbuQamar. A consortium of rhizospherecompetent actinobacteria exhibiting multiple plant growth-promoting traits improves the growth of Avicennia marina in the United Arab Emirates. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Safuan:2021:PRS

[2671] Che Din Mohd Safuan, Muhammad Arif Samshuri, Siti Nurtahirah Jaafar, Chun Hong Tan, and Zainudin Bachok. Physiological response of shallow-water hard coral Acropora digitifera to heat stress via fatty acid composition. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Ronspiess:2021:BVP

[2672] Lisa Rönspieß, Günther Nausch, and Detlef Schulz-Bull. Bioavailability of various phosphorus fractions and their seasonality in a eutrophic estuary in the Southern Baltic Sea — a laboratory approach. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Brandt:2021:EII

[2673] Marilyn E. Brandt, Rosmin S. Ennis, Sonora S. Meiling, Joseph Townsend, Kathryn Cobleigh, Adam Glahn, Jason Quetel, Viktor Brandtneris, Leslie M. Henderson, and Tyler B. Smith. The emergence and initial impact of Stony Coral Tissue Loss Disease (SCTLD) in the United States Virgin Islands. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Walter:2021:MIE

[2674] Juline Marta Walter, Louisi Souza de Oliveira, Diogo Antonio Tschoeke, Pedro Milet Meirelles, Maria Helena Campos Baeta Neves, Daniela Batista, Ana Polycarpa Carvalho, Rafaela Dos Santos Costa, Sergey Dobretsov, Ricardo Coutinho, Jean Swings, Cristiane Carneiro Thompson, and Fabiano L. Thompson. Metagenomic insights into ecosystem function in the microbial mats of a large hypersaline coastal lagoon system. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Xiao:2021:CNL

[2675] Ziyu Xiao, Zhaoqing Yang, Taiping Wang, Ning Sun, Mark Wigmosta, and David Judi. Characterizing the non-linear interactions between tide, storm surge, and river flow in the Delaware Bay Estuary, United States. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Rahimnejad:2021:EDP

[2676] Samad Rahimnejad, Konrad Dabrowski, Marisol Izquierdo, Oleksandr Malinovskyi, Jitka Kolářová, and Tomas Policar. Effects of dietary protein and lipid levels on growth, body composition, blood biochemistry, antioxidant capacity and ammonia excretion of European grayling (Thymallus thymallus). Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Guo:2021:VIG

[2677] Jinqiang Guo, Huamao Yuan, Jinming Song, Baoxiao Qu, Jianwei Xing, Qidong Wang, Xuegang Li, Liqin Duan, Ning Li, and Yingxia Wang. Variation of isoprenoid GDGTs in the stratified marine water column: Implications for GDGT-based TEX86 paleothermometry. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Basilone:2021:RSM

[2678] Gualtiero Basilone, Rosalia Ferreri, Salvatore Aronica, Salvatore Mazzola, Angelo Bonanno, Antonella Gargano, Maurizio Pulizzi, Ignazio Fontana, Giovanni Giacalone, Pietro Calandrino, Simona Genovese, and Marco Barra. Reproduction and sexual maturity of European sardine (Sardina pilchardus) in the Central Mediterranean Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lopez:2021:EFR

[2679] Alexander G. López, Raymond G. Najjar, Marjorie A. M. Friedrichs, Michael A. Hickner, and Denice H. Wardrop. Estuaries as filters for riverine microplastics: Simulations in a large, coastal-plain estuary. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Yang:2021:ESV

[2680] Yang Yang, Tong Wang, Sen Yang, Xi Wu, Wenhua Huang, Bin Fan, Junhong Xia, Zining Meng, and Xiaochun Liu. The effects of storage in vitro on functions, transcriptome, proteome, and oxidation resistance of giant grouper sperm. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See corrigendum [3148].

McRae:2021:CPR

[2681] Crystal J. McRae, Anderson B. Mayfield, Wen-Bin Huang, Isabelle M. Côté, and Tung-Yung Fan. Contrasting proteomic responses of adult and larval coral to high temperatures. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chen:2021:SVC

[2682] Chih-Shin Chen. Spatiotemporal variation in the catch composition and abundance of precious corals around Taiwan: Implications for fisheries management. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fulfer:2021:HDG

[2683] Victoria M. Fulfer and Susanne Menden-Deuer. Heterotrophic dinoflagellate growth and grazing rates reduced by microplastic ingestion. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cembella:2021:PDA

[2684] Allan D. Cembella, Lorena M. Durán-Riveroll, Ulrike I. Tarazona-Janampa, Yuri B. Okolodkov, Ricardo García-Sandoval, Bernd Krock, Cora Hörstmann, and Uwe John. Phylogeography and diversity among populations of the toxigenic benthic dinoflagellate prorocentrum from coastal reef systems in Mexico. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ma:2021:DET

[2685] Qiang Ma, Yusheng Qiu, Run Zhang, E Lv, Yipu Huang, and Min Chen. 210Po/^{210}Pb disequilibria in the eastern tropical North Pacific. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ye:2021:MYE

[2686] Min Ye, Jie Nie, Anan Liu, Zhigang Wang, Lei Huang, Hao Tian, Dehai Song, and Zhiqiang Wei. Multi-year ENSO forecasts using parallel convolutional neural networks with heterogeneous architecture. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Eaton:2021:MSC

[2687] Katherine R. Eaton, Jan H. Landsberg, Yasunari Kiryu, Esther C. Peters, and Erinn M. Muller. Measuring Stony Coral Tissue Loss Disease induction and lesion progression within two intermediately susceptible species, Montastraea cavernosa and Orbicella faveolata. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Ali:2021:SMS

[2688] Nadia Gabr Ali, Asmaa Mahmoud El-Nokrashy, Mofeed Yousef Gouda, and Ibrahim Mohamed Aboyadak. Summer mortality syndrome affecting cultured European seabass at Kafrelsheikh Province, Egypt. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Cardozo-Ferreira:2021:ETI

[2689] Gabriel C. Cardozo-Ferreira, Thayane L. Calazans, Larissa J. Benevides, Osmar J. Luiz, Carlos E. L. Ferreira, and Jean-Christophe Joyeux. Ecological traits influencing anthropogenic debris ingestion by herbivorous reef fishes. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sheaves:2021:ECM

[2690] Marcus Sheaves, Carlo Mattone, Rod M. Connolly, Stephanie Hernandez, Ivan Nagelkerken, Nicholas Murray, Michael Ronan, Nathan John Waltham, and Michael Bradley. Ecological constraint mapping: Understanding outcome-limiting bottlenecks for improved environmental decision-making in marine and coastal environments. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cristiani:2021:BMN

[2691] John Cristiani, Emily Rubidge, Coreen Forbes, Ben Moore-Maley, and Mary I. O’Connor. A biophysical model and network analysis of invertebrate community dispersal reveals regional patterns of seagrass habitat connectivity. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chow:2021:WEB

[2692] Chun Hoe Chow, Yung-Yen Shih, Ya-Tang Chien, Jing Yi Chen, Ning Fan, Wei-Chang Wu, and Chin-Chang Hung. The wind effect on biogeochemistry in eddy cores in the Northern South China Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chou:2021:UDP

[2693] Wen-Chen Chou, Lan-Feng Fan, Chang-Chang Yang, Ying-Hsuan Chen, Chin-Chang Hung, Wei-Jen Huang, Yung-Yen Shih, Keryea Soong, Hsiao-Chun Tseng, Gwo-Ching Gong, Hung-Yu Chen, and Cheng-Kuan Su. A unique diel pattern in carbonate chemistry in the seagrass meadows of Dongsha Island: The enhancement of metabolic carbonate dissolution in a semienclosed lagoon. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Sanchez-Arcilla:2021:CAU

[2694] Agustín Sánchez-Arcilla, Vicente Gracia, César Mösso, Iván Cáceres, Daniel González-Marco, and Jesús Gómez. Coastal adaptation and uncertainties: The need of ethics for a shared coastal future. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cappelletto:2021:GOE

[2695] Margherita Cappelletto, Rita Giuffredi, Erasmia Kastanidi, Vassiliki Vassilopoulou, and Alba L’Astorina. Grounding ocean ethics while sharing knowledge and promoting environmental responsibility: Empowering young ambassadors as agents of change. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:LBA

[2696] Kui Zhang, Jiajun Li, Gang Hou, Zirong Huang, Dengfu Shi, Zuozhi Chen, and Yongsong Qiu. Length-based assessment of fish stocks in a data-poor, jointly exploited (China and Vietnam) fishing ground, Northern South China Sea. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Giuffredi:2021:KCC

[2697] Rita Giuffredi, Laura Criscuolo, Amelia De Lazzari, Giovanni Fanelli, Raffaele Giordano, Antonella Petrocelli, Giuseppe Portacci, Alessandra Pugnetti, and Alba L’Astorina. Knowledge co-construction by citizens and researchers to create a SNAPSHOT of the marine environment during and after the Covid-19 lockdown. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Barthelmess:2021:VSS

[2698] Theresa Barthelmeß, Florian Schütte, and Anja Engel. Variability of the sea surface microlayer across a filament’s edge and potential influences on gas exchange. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cortes-Useche:2021:CCR

[2699] Camilo Cortés-Useche, Williams Reyes-Gamboa, José Luís Cabrera-Pérez, Johanna Calle-Triviño, Ana Cerón-Flores, Rodolfo Raigoza-Figueras, Roshni Yathiraj, and Jesús Ernesto Arias-González. Capture, culture and release of postlarvae fishes: Proof-of-concept as a tool approach to support reef management. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Izett:2021:NTM

[2700] Robert W. Izett, Roberta C. Hamme, Craig McNeil, Cara C. M. Manning, Annie Bourbonnais, and Philippe D. Tortell. Δ O_{2} /N_{2}′ as a new tracer of marine net community production: Application and evaluation in the Subarctic Northeast Pacific and Canadian Arctic Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Venkatesan:2021:SIM

[2701] R. Venkatesan, K. Jossia Joseph, C. Anoopa Prasad, M. Kalyani, M. Arul Muthiah, S. Ramasundaram, P. Murugesh, K. Thirumurugan, R. Sundar, B. Kesavakumar, G. Vengatesan, K. Ramesh, M. V. Martin, K. N. Navaneeth, P. Senthilkumar, Biswajit Haldar, Abhishek Tandon, R. Sridharan, S. Sundar Jesuraj, C. Muthukumar, N. Sundaravadivelu, and M. Saravanan. Sustenance of Indian moored buoy network during COVID-19 pandemic — a saga of perseverance. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Rogers:2021:CMG

[2702] Alex D. Rogers, Amy Baco, Elva Escobar-Briones, Duncan Currie, Kristina Gjerde, Judith Gobin, Marcel Jaspars, Lisa Levin, Katrin Linse, Muriel Rabone, Eva Ramirez-Llodra, Javier Sellanes, Timothy M. Shank, Kerry Sink, Paul V. R. Snelgrove, Michelle L. Taylor, Daniel Wagner, and Harriet Harden-Davies. Corrigendum: Marine genetic resources in areas beyond national jurisdiction: Promoting marine scientific research and enabling equitable benefit sharing. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745. See [2123].

Lunt:2021:SSC

[2703] Jessica Lunt, Christopher J. Freeman, Dean S. Janiak, Katrina Bayliss, Michelle Stephens, Eve Galimany, and Valerie J. Paul. Soft-sediment communities of the Northern Indian River Lagoon, FL, United States. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Haward:2021:CNS

[2704] Marcus Haward and Bianca Haas. Corrigendum: The need for social considerations in SDG 14. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745. See [1711].

McKenna:2021:AIS

[2705] Megan F. McKenna, Simone Baumann-Pickering, Annebelle C. M. Kok, William K. Oestreich, Jeffrey D. Adams, Jack Barkowski, Kurt M. Fristrup, Jeremy A. Goldbogen, John Joseph, Ella B. Kim, Anke Kügler, Marc O. Lammers, Tetyana Margolina, Lindsey E. Peavey Reeves, Timothy J. Rowell, Jenni A. Stanley, Alison K. Stimpert, Eden J. Zang, Brandon L. Southall, Carrie C. Wall, Sofie Van Parijs, and Leila T. Hatch. Advancing the interpretation of shallow water marine soundscapes. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Purkiani:2021:NSD

[2706] Kaveh Purkiani, Benjamin Gillard, André Paul, Matthias Haeckel, Sabine Haalboom, Jens Greinert, Henko de Stigter, Martina Hollstein, Matthias Baeye, Annemiek Vink, Laurenz Thomsen, and Michael Schulz. Numerical simulation of deep-sea sediment transport induced by a dredge experiment in the Northeastern Pacific Ocean. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Naik:2021:MFP

[2707] Azza S. Naik, R. D. Whitaker, S. Albrektsen, Runar G. Solstad, Lars Thoresen, and Maria Hayes. Mesopelagic fish protein hydrolysates and extracts: a source of novel anti-hypertensive and anti-diabetic peptides. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Luis:2021:PDF

[2708] Orlando J. Luis and João M. Gago. Perception of dissolved food-related compounds by the sea urchin Paracentrotus lividus (Echinodermata: Echinoidea). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pata:2021:SPI

[2709] Patrick R. Pata and Aletta T. Yñiguez. Spatial planning insights for Philippine coral reef conservation using larval connectivity networks. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Dai:2021:AEC

[2710] Qi Dai, Zhi-Xuan Wang, Yan-Qing Sheng, Zhi-Wen Wu, Yan Qiu, Pei Su, Cai-Huan Ke, and Dan-Qing Feng. 2-arachidonoylglycerol as an endogenous cue negatively regulates attachment of the mussel Perna viridis. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Authier:2021:ECB

[2711] Matthieu Authier, Etienne Rouby, and Kelly Macleod. Estimating cetacean bycatch from non-representative samples (I): a simulation study with regularized multilevel regression and post-stratification. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Tomasetti:2021:BEN

[2712] Stephen J. Tomasetti, Jeffrey R. Kraemer, and Christopher J. Gobler. Brief episodes of nocturnal hypoxia and acidification reduce survival of economically important blue crab (Callinectes sapidus) larvae. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Lewis:2021:SDS

[2713] Savannah L. Lewis, Georges Saliba, Lynn M. Russell, Patricia K. Quinn, Timothy S. Bates, and Michael J. Behrenfeld. Seasonal differences in submicron marine aerosol particle organic composition in the North Atlantic. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Harcourt:2021:RVW

[2714] Rob Harcourt, Mark A. Hindell, Clive R. McMahon, Kimberly T. Goetz, Jean-Benoit Charrassin, Karine Heerah, Rachel Holser, Ian D. Jonsen, Michelle R. Shero, Xavier Hoenner, Rose Foster, Baukje Lenting, Esther Tarszisz, and Matthew Harry Pinkerton. Regional variation in winter foraging strategies by Weddell seals in Eastern Antarctica and the Ross Sea. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

He:2021:DNM

[2715] Wenxuan He, Sizhen Liu, Zhichen Jiang, Jinshui Zheng, Xuegang Li, and Dechao Zhang. The diversity and nitrogen metabolism of culturable nitrate-utilizing bacteria within the oxygen minimum zone of the Changjiang (Yangtze River) Estuary. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Choi:2021:NPS

[2716] Jung Min Choi, Jae Ho Jung, Ki Hong Kim, D. Wayne Coats, and Young Ok Kim. A novel parasitic, syndinean dinoflagellate Euduboscquella triangula infecting the tintinnid Helicostomella longa. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Jaiteh:2021:BEP

[2717] Vanessa Jaiteh, Tom Peatman, Steve Lindfield, Eric Gilman, and Simon Nicol. Bycatch estimates from a Pacific tuna longline fishery provide a baseline for understanding the long-term benefits of a large, blue water marine sanctuary. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Isabel:2021:DET

[2718] Laurie Isabel, David Beauchesne, Chris McKindsey, and Philippe Archambault. Detection of ecological thresholds and selection of indicator taxa for epibenthic communities exposed to multiple pressures. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sudatti:2021:VSC

[2719] Daniela Bueno Sudatti, Aline Santos Oliveira, Bernardo Antonio Perez da Gama, Mutue Toyota Fujii, Silvana Vianna Rodrigues, and Renato Crespo Pereira. Variability in seaweed chemical defense and growth under common garden conditions. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Torres-Perez:2021:RSR

[2720] Juan L. Torres-Pérez, Carlos E. Ramos-Scharrón, William J. Hernández, Roy A. Armstrong, Maritza Barreto-Orta, Jorge Ortiz-Zayas, Liane S. Guild, and Roberto Viqueira. River streamflow, remotely sensed water quality, and benthic composition of previously undescribed nearshore coral reefs in Northern Puerto Rico. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Meynecke:2021:RED

[2721] Jan-Olaf Meynecke, Jasper de Bie, Jan-Lukas Menzel Barraqueta, Elisa Seyboth, Subhra Prakash Dey, Serena B. Lee, Saumik Samanta, Marcello Vichi, Ken Findlay, Alakendra Roychoudhury, and Brendan Mackey. The role of environmental drivers in humpback whale distribution, movement and behavior: a review. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Paoletti:2021:PMF

[2722] Silvia Paoletti, J. Rasmus Nielsen, Claus R. Sparrevohn, Francois Bastardie, and Berthe M. J. Vastenhoud. Potential for mesopelagic fishery compared to economy and fisheries dynamics in current large scale Danish pelagic fishery. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Tait:2021:LGK

[2723] Leigh W. Tait, François Thoral, Matthew H. Pinkerton, Mads S. Thomsen, and David R. Schiel. Loss of giant kelp, Macrocystis pyrifera, driven by marine heatwaves and exacerbated by poor water clarity in New Zealand. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Fang:2021:RAD

[2724] Xingnan Fang, Wei Yu, Xinjun Chen, and Yang Zhang. Response of abundance and distribution of Humboldt squid (Dosidicus gigas) to short-lived eddies in the Eastern Equatorial Pacific Ocean from April to June 2017. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ma:2021:SGA

[2725] Haitao Ma, Dongmei Yu, Shu Xiao, Yanping Qin, Yang Zhang, Jun Li, Yuehuan Zhang, and Ziniu Yu. A significant genetic admixture in farmed populations of the noble scallop Chlamys nobilis revealed by microsatellite DNA analysis in Southern China. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Wu:2021:HHE

[2726] Jiong Wu, Weiwei Zhang, and Chenghua Li. Heat and hypoxia exposure mediates circadian rhythms response via methylation modification in apostichopus japonicas. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Toyoda:2021:SLC

[2727] Takahiro Toyoda, Kei Sakamoto, Norihisa Usui, Nariaki Hirose, Kiyoshi Tanaka, Takaaki Katsumata, Daisuke Takahashi, Masato Niki, Kunio Kutsuwada, Toru Miyama, Hideyuki Nakano, L. Shogo Urakawa, Kensuke K. Komatsu, Yuma Kawakami, and Goro Yamanaka. Surface-layer circulations in Suruga Bay induced by intrusions of Kuroshio Branch water. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Buhl-Mortensen:2021:ESM

[2728] Pål Buhl-Mortensen, Vincent Lecours, and Craig J. Brown. Editorial: Seafloor mapping of the Atlantic Ocean. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Schmidt:2021:FRA

[2729] Neele Schmidt, Yusuf C. El-Khaled, Felix I. Roßbach, and Christian Wild. Fleshy red algae mats influence their environment in the Mediterranean Sea. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cravo:2021:FAS

[2730] Mariana Cravo, Armando J. Almeida, Hamilton Lima, João Azevedo e Silva, Salomão Bandeira, Vilma Machava-António, and José Paula. Fish assemblages in a small mangrove system on Príncipe Island, Gulf of Guinea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Sultana:2021:MRF

[2731] Rumana Sultana, Haseeb Md. Irfanullah, Samiya A. Selim, Syed Tauheed Raihan, Joy Bhowmik, and Shaikh Giasuddin Ahmed. Multilevel resilience of fishing communities of coastal Bangladesh against Covid-19 pandemic and 65-day fishing ban. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Karr:2021:IPC

[2732] Kendra A. Karr, Valerie Miller, Eva Coronado, Nadia C. Olivares-Bañuelos, Martha Rosales, Javier Naretto, Luciano Hiriart-Bertrand, Camila Vargas-Fernández, Romina Alzugaray, Rafael Puga, Servando Valle, L. P. Osman, Julio Chamorro Solís, Marco Ide Mayorga, Doug Rader, and Rod Fujita. Identifying pathways for climate-resilient multispecies fisheries. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hein:2021:LIB

[2733] Christopher J. Hein, Michael S. Fenster, Keryn B. Gedan, Jeff R. Tabar, Emily A. Hein, and Todd DeMunda. Leveraging the interdependencies between barrier islands and backbarrier saltmarshes to enhance resilience to sea-level rise. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Seelanki:2021:EIP

[2734] Vivek Seelanki and Vimlesh Pant. An evaluation of the impact of pandemic driven lockdown on the phytoplankton biomass over the North Indian Ocean using observations and model. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Marsh:2021:SPH

[2735] Robert Marsh, Kwasi Appeaning Addo, Philip-Neri Jayson-Quashigah, Hazel A. Oxenford, Ava Maxam, Romario Anderson, Nikolaos Skliris, Jadu Dash, and Emma L. Tompkins. Seasonal predictions of holopelagic sargassum across the tropical Atlantic accounting for uncertainty in drivers and processes: The SARTRAC ensemble forecast system. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Barrera-Falcon:2021:CSC

[2736] Erick Barrera-Falcon, Rodolfo Rioja-Nieto, Roberto C. Hernández-Landa, and Edgar Torres-Irineo. Comparison of standard Caribbean coral reef monitoring protocols and underwater digital photogrammetry to characterize hard coral species composition, abundance and cover. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Li:2021:ADG

[2737] Li Li, Duc Huy Dang, Xiaojing Wang, Jihua Liu, Yonghua Wu, Aimei Zhu, and Xuefa Shi. Atypical diagenesis and geochemistry of redoxsensitive elements in hydrothermal sediments of the Southern Okinawa Trough. Frontiers in Marine Science, 8, September 2021. ISSN 2296-

Roman:2021:UUM

[2738] Alejandro Román, Antonio Tovar-Sánchez, Irene Olivé, and Gabriel Navarro. Using a UAV-mounted multispectral camera for the monitoring of marine macrophytes. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Anonymous:2021:EAH

[2739] Anonymous. Erratum: An assessment of how Australian fisheries management plans account for climate change impacts. Frontiers in Marine Science, 8, June 2021. ISSN 2296-7745. See [1166].

Untiedt:2021:IBC

[2740] Candice B. Untiedt, Alan Williams, Franziska Althaus, Phil Alderslade, and Malcolm R. Clark. Identifying black corals and octocorals from deepsea imagery for ecological assessments: Trade-offs between morphology and taxonomy. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Taylor:2021:CID

[2741] Ryland B. Taylor, Martha E. Mather, Joseph M. Smith, and Kayla M. Boles. Can identifying discrete behavioral groups with individual-based acoustic telemetry advance the understanding of fish distribution patterns? Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Paulmier:2021:HSC

[2742] Aurelien Paulmier, Gerard Eldin, José Ochoa, Boris Dewitte, Joël Sudre, Véronique Garçon, Jaques Grelet, Kobi Mosquera-Vásquez, Oscar Vergara, and Helmut Maske. High-sustained concentrations of organisms at very low oxygen concentration indicated by acoustic profiles in the oxygen deficit region off Peru. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Krahmann:2021:CBI

[2743] Gerd Krahmann, Damian L. Arévalo-Martínez, Andrew W. Dale, Marcus Dengler, Anja Engel, Nicolaas Glock, Patricia Grasse, Johannes Hahn, Helena Hauss, Mark J. Hopwood, Rainer Kiko, Alexandra N. Loginova, Carolin R. Löscher, Marie Maßmig, Alexandra-Sophie Roy, Renato Salvatteci, Stefan Sommer, Toste Tanhua, and Hela Mehrtens. Climatebiogeochemistry interactions in the tropical ocean: Data collection and legacy. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sands:2021:PAS

[2744] Chester J. Sands, Timothy D. O’Hara, and Rafael Martín-Ledo. Pragmatic assignment of species groups based on primary species hypotheses: The case of a dominant component of the Southern Ocean benthic fauna. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Rudovica:2021:VMW

[2745] Vita Rudovica, Ana Rotter, Susana P. Gaudêncio, Lucie Novoveská, Füsun Akgül, Linn Kristin Akslen-Hoel, Diogo A. M. Alexandrino, Olga Anne, Lauris Arbidans, Miroslava Atanassova, Magdalena Bełdowska, Jacek Bełdowski, Amit Bhatnagar, Oskars Bikovens, Valdis Bisters, Maria F. Carvalho, Teresa S. Catalá, Arita Dubnika, Ayşegül Erdoğan, Laura Ferrans, Berat Z. Haznedaroglu, Roy Hendroko Setyobudi, Bożena Graca, Inga Grinfelde, William Hogland, Efstathia Ioannou, Yahya Jani, Marija Kataržytė, Stefanos Kikionis, Katja Klun, Jonne Kotta, Mait Kriipsalu, Jalel Labidi, Lada Lukić Bilela, Marta Martínez-Sanz, Juliana Oliveira, Ruta Ozola-Davidane, Jovita Pilecka-Ulcugaceva, Kristyna Pospiskova, Céline Rebours, Vassilios Roussis, Amparo López-Rubio, Ivo Safarik, Frank Schmieder, Karina Stankevica, Toomas Tamm, Deniz Tasdemir, Cristiana Torres, Giovanna Cristina Varese, Zane Vincevica-Gaile, Ivar Zekker, and Juris Burlakovs. Valorization of marine waste: Use of industrial by-products and beach wrack towards the production of high added-value products. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kurczyn:2021:AUW

[2746] Jorge. A. Kurczyn, Rodrigo Duran, Emilio Beier, and Alejandro J. Souza. On the advection of upwelled water on the Western Yucatan Shelf. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lam-Gordillo:2021:TFP

[2747] Orlando Lam-Gordillo, Ryan Baring, and Sabine Dittmann. Taxonomic and functional patterns of benthic communities in southern temperate tidal flats. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Tzeng:2021:FCQ

[2748] Mimi W. Tzeng, Oliver Floerl, and Anastasija Zaiko. A framework for compiling quantifications of marine biosecurity risk factors associated with common vessel types. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Spadafore:2021:MVA

[2749] Rachele Spadafore, Ryan Fura, William F. Precht, and Steven V. Vollmer. Multi-variate analyses of coral mortality from the 2014– 2015 Stony Coral Tissue Loss Disease outbreak off Miami-Dade County, Florida. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Shi:2021:SMM

[2750] Ce Shi, Yangfang Ye, Feng Pei, Changkao Mu, and Chunlin Wang. Survival and metabolic modulation of swimming crab Portunus trituberculatus during live transport. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gregoire:2021:GOO

[2751] Marilaure Grégoire, Véronique Garçon, Hernan Garcia, Denise Breitburg, Kirsten Isensee, Andreas Oschlies, Maciej Telszewski, Alexander Barth, Henry C. Bittig, Jacob Carstensen, Thierry Carval, Fei Chai, Francisco Chavez, Daniel Conley, Laurent Coppola, Sean Crowe, Kim Currie, Minhan Dai, Bruno Deflandre, Boris Dewitte, Robert Diaz, Emilio Garcia-Robledo, Denis Gilbert, Alessandra Giorgetti, Ronnie Glud, Dimitri Gutierrez, Shigeki Hosoda, Masao Ishii, Gil Jacinto, Chris Langdon, Siv K. Lauvset, Lisa A. Levin, Karin E. Limburg, Hela Mehrtens, Ivonne Montes, Wajih Naqvi, Aurélien Paulmier, Benjamin Pfeil, Grant Pitcher, Sylvie Pouliquen, Nancy Rabalais, Christophe Rabouille, Virginie Recape, Michaël Roman, Kenneth Rose, Daniel Rudnick, Jodie Rummer, Catherine Schmechtig, Sunke Schmidtko, Brad Seibel, Caroline Slomp, U. Rashid Sumalia, Toste Tanhua, Virginie Thierry, Hiroshi Uchida, Rik Wanninkhof, and Moriaki Yasuhara. A global ocean oxygen database and atlas for assessing and predicting deoxygenation and ocean health in the open and coastal ocean. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhao:2021:PBT

[2752] Linlin Zhao, Shouqiang Wang, Fangrui Lou, Tianxiang Gao, and Zhiqiang Han. Phylogenomics based on transcriptome data provides evidence for the internal phylogenetic relationships and potential terrestrial evolutionary genes of lungfish. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Esenkulova:2021:HAO

[2753] Svetlana Esenkulova, Karyn D. Suchy, Rich Pawlowicz, Maycira Costa, and Isobel A. Pearsall. Harmful algae and oceanographic conditions in the Strait of Georgia, Canada based on citizen science monitoring. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Rogers-Bennett:2021:SCD

[2754] Laura Rogers-Bennett, Robert Klamt, and Cynthia A. Catton. Survivors of climate driven abalone mass mortality exhibit declines in health and reproduction following kelp forest collapse. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Kim:2021:IFD

[2755] Bo Kyung Kim, Misa Jeon, Hyoung Min Joo, Tae-Wan Kim, Sang-Jong Park, Jisoo Park, and Sun-Yong Ha. Impact of freshwater discharge on the carbon uptake rate of phytoplankton during summer (January– February 2019) in Marian Cove, King George Island, Antarctica. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Kume:2021:DFH

[2756] Gen Kume, Taichi Shigemura, Masahiro Okanishi, Junya Hirai, Kazuhiro Shiozaki, Mutsuo Ichinomiya, Tomohiro Komorita, Akimasa Habano, Fumihiro Makino, and Toru Kobari. Distribution, feeding habits, and growth of chub mackerel, Scomber japonicus, larvae during a high-stock period in the Northern Satsunan Area, Southern Japan. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zhang:2021:SAH

[2757] Anguo Zhang, Hongjun Li, Xiaolong Yang, Lili Wang, Yangchun Gao, Mingshan Song, and Xiutang Yuan. Stock assessment of hatcheryreleased clam Meretrix meretrix in an estuary of China from the perspectives of population ecology and genetic diversity. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Yao:2021:GTA

[2758] Hanhan Yao, Zhihua Lin, Yinghui Dong, Xianghui Kong, Lin He, and Liangyi Xue. Gonad transcriptome analysis of the razor clam (Sinonovacula constricta) revealed potential sex-related genes. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Wang:2021:CAD

[2759] Xue Wang, Weijun Wang, Zan Li, Guohua Sun, Tao Xu, Xiaohui Xu, Yanwei Feng, Qihao Luo, Bin Li, and Jianmin Yang. Comprehensive analysis of differentially expressed mRNA, non-coding RNA, and their competitive endogenous RNA network of Pacific oyster Crassostrea gigas with different glycogen content between different environments. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cloyed:2021:WIM

[2760] Carl S. Cloyed, Elizabeth E. Hieb, Kayla DaCosta, Monica Ross, and Ruth H. Carmichael. West Indian manatees use partial migration to expand their geographic range into the Northern Gulf of Mexico. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Lv:2021:MAE

[2761] Huirong Lv, Yanli Liu, Haidong Li, Xiaolong Yin, Ping Wang, Xiaoyu Qu, Yang Gao, Weiye Li, and Zhangjie Chu. Modulation of antioxidant enzymes, heat shock protein, and intestinal microbiota of large yellow croaker (Larimichthys crocea) under acute cold stress. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Zhu:2021:ITM

[2762] Xiaowen Zhu, Ruijuan Hao, Changxu Tian, Junpeng Zhang, Chunhua Zhu, and Guangli Li. Integrative transcriptomics and metabolomics analysis of body color formation in the leopard coral grouper (Plectropomus leopardus). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Cunha:2021:IPS

[2763] Jacinto Cunha, Fábio Santos Cardona, Ana Bio, and Sandra Ramos. Importance of protection service against erosion and storm events provided by coastal ecosystems under climate change scenarios. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Basirati:2021:EZO

[2764] Mohadese Basirati, Romain Billot, Patrick Meyer, and Erwan Bocher. Exact zoning optimization model for marine spatial planning (MSP). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

deSmit:2021:KBS

[2765] Jaco C. de Smit, Muriel Z. M. Brückner, Katherine I. Mesdag, Maarten G. Kleinhans, and Tjeerd J. Bouma. Key bioturbator species within benthic communities determine sediment resuspension thresholds. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Qiu:2021:ESS

[2766] Chunhua Qiu, Hong Liang, Xiujun Sun, Huabin Mao, Dongxiao Wang, Zhenhui Yi, and Anindya Wirasatriya. Extreme sea-surface cooling induced by eddy heat advection during tropical cyclone in the North Western Pacific Ocean. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Olsvik:2021:OSS

[2767] Pål A. Olsvik, Elin Sørhus, Sonnich Meier, Ingrid Torvanger, Maja Thorbjørnsen, Anders Thorsen, Lisbet Sørensen, Per Gunnar Fjelldal, Ørjan Karlsen, Igo G. Guimaraes, and Kai K. Lie. Ontogeny-specific skeletal deformities in Atlantic haddock caused by larval oil exposure. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hochberg:2021:MRC

[2768] Eric J. Hochberg and Michelle M. Gierach. Missing the reef for the corals: Unexpected trends between coral reef condition and the environment at the ecosystem scale. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Vona:2021:SEB

[2769] Iacopo Vona, Cindy M. Palinkas, and William Nardin. Sediment exchange between the created saltmarshes of living shorelines and adjacent submersed aquatic vegetation in the Chesapeake Bay. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Caramanna:2021:SSW

[2770] Giorgio Caramanna, Stefan M. Sievert, and Solveig I. Bühring. Submarine shallow-water fluid emissions and their geomicrobiological imprint: a global overview. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Jones:2021:ULC

[2771] Ross Jones, Mari-Carmen Pineda, Heidi M. Luter, Rebecca Fisher, David Francis, Wojciech Klonowski, and Matthew Slivkoff. Underwater light characteristics of turbid coral reefs of the Inner Central Great Barrier Reef. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Farhadi:2021:NIG

[2772] Ardavan Farhadi, Xi Shi, Yin Zhang, Yueling Zhang, Shengkang Li, Huaiping Zheng, Mhd Ikhwanuddin, and Hongyu Ma. A novel imprinted gene (Sp-Pol) with sex-specific SNP locus and sex-biased expression pattern provides insights into the gonad development of mud crab (Scylla paramamosain). Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Zhu:2021:EWS

[2773] Zhenchang Zhu, Tjeerd J. Bouma, Qin Zhu, Yanpeng Cai, and Zhifeng Yang. Effects of waves and sediment disturbance on seed bank persistence at tidal flats. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bussmann:2021:DPM

[2774] Ingeborg Bussmann, Holger Brix, Götz Flöser, Uta Ködel, and Philipp Fischer. Detailed patterns of methane distribution in the German Bight. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chen:2021:IEI

[2775] Biao Chen, Kefu Yu, Qiucui Yao, Zhiheng Liao, Zhenjun Qin, Xiaopeng Yu, Qian Wu, and Biao He. Insights into the environmental impact on genetic structure and larval dispersal of crown-of-thorns starfish in the South China Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Matejusova:2021:EDB

[2776] Iveta Matejusova, Jennifer Graham, Fiona Bland, Jean-Pierre Lacaze, Guillaume Herman, Lyndsay Brown, Eric Dalgarno, John D. Bishop, Jenni E. Kakkonen, Kirsty F. Smith, and Alex Douglas. Environmental DNA based surveillance for the highly invasive carpet sea squirt Didemnum vexillum: a targeted single-species approach. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Singh:2021:DFF

[2777] Aprajita Singh, Sajjad Karimi, Aleksandar Vidakovic, Johan Dicksved, Markus Langeland, Jorge A. Ferreira, Mohammad J. Taherzadeh, Anders Kiessling, and Torbjörn Lundh. Dietary filamentous fungi and duration of feeding modulates gut microbial composition in rainbow trout (Oncorhynchus mykiss). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bisaro:2021:GCS

[2778] Alexander Bisaro, Jochen Hinkel, Gonéri Le Cozannet, Thomas van der Pol, and Armin Haas. Global climate services: a typology of global decisions influenced by climate risk. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Lei:2021:ACM

[2779] Xinming Lei, Hao Cheng, Yong Luo, Yuyang Zhang, Lei Jiang, Youfang Sun, Guowei Zhou, and Hui Huang. Abundance and characteristics of microplastics in seawater and corals from reef region of Sanya Bay, China. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Song:2021:TPB

[2780] Pu-Qing Song, Hu-Shun Zhang, Xin-Qing Zheng, Yuan-Yuan Li, Long-Shan Lin, and Yuan Li. Trophic plasticity of Bombay duck (Harpadon nehereus) in the South-Central East China Sea based on stable isotope evidence. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pan:2021:EED

[2781] Yen-Ju Pan, Wei-Lung Wang, Jiang-Shiou Hwang, and Sami Souissi. Effects of epibiotic diatoms on the productivity of the calanoid copepod Acartia tonsa (Dana) in intensive aquaculture systems. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cho:2021:EAM

[2782] Suhyung Cho, Sang-Hyeok Cho, So-Ra Ko, Yujin Jeong, Eunju Lee, Sangrak Jin, Bo-Seong Jeong, Byung-Ha Oh, Hee-Mock Oh, Chi-Yong Ahn, and Byung-Kwan Cho. Elucidation of the algicidal mechanism of the marine bacterium Pseudoruegeria sp. M32A2M against the harmful alga Alexandrium catenella based on time-course transcriptome analysis. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Angelini:2021:UDA

[2783] Silvia Angelini, Enrico N. Armelloni, Ilaria Costantini, Andrea De Felice, Igor Isajlović, Iole Leonori, Chiara Manfredi, Francesco Masnadi, Giuseppe Scarcella, Vjekoslav Tičina, and Alberto Santojanni. Understanding the dynamics of ancillary pelagic species in the Adriatic Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Bautista:2021:WTG

[2784] Naim M. Bautista and Amélie Crespel. Within- and trans-generational environmental adaptation to climate change: Perspectives and new challenges. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Goransson:2021:ARC

[2785] Gunnel Göransson, Anna Apler, Anna-Karin Dahlberg, Hjördis Löfroth, Sarah Josefsson, Karin Wiberg, Paul Frogner-Kockum, Per Nylander, Jim Hedfors, and Ian Snowball. Assessing the risk of contaminant dispersion from fibrous sediments of industrial origin. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chowdhury:2021:SOT

[2786] Riyanka Roy Chowdhury, S. Prasanna Kumar, and Arun Chakraborty. Simultaneous occurrence of tropical cyclones in the Northern Indian Ocean: Differential response and triggering mechanisms. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Baek:2021:DLS

[2787] Sang-Soo Baek, JongCheol Pyo, Yong Sung Kwon, Seong-Jun Chun, Seung Ho Baek, Chi-Yong Ahn, Hee-Mock Oh, Young Ok Kim, and Kyung Hwa Cho. Deep learning for simulating harmful algal blooms using ocean numerical model. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

VanDam:2021:OPS

[2788] Bryce Van Dam, Christian Lopes, Mary A. Zeller, Mariana Ribas-Ribas, Hongjie Wang, and Helmuth Thomas. Overstated potential for seagrass meadows to mitigate coastal ocean acidification. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745. See corrigendum [3211].

Gielings:2021:DMM

[2789] Romy Gielings, Maria Fais, Diego Fontaneto, Simon Creer, Filipe Oliveira Costa, Willem Renema, and Jan-Niklas Macher. DNA metabarcoding methods for the study of marine benthic meiofauna: a review. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Nauen:2021:TSG

[2790] Cornelia E. Nauen and Maria Fernanda Arraes Treffner. Translating SSF guidelines into practice with the Small-Scale Fisheries Academy. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Huang:2021:CWE

[2791] Songqian Huang, Yuki Ichikawa, Kazutoshi Yoshitake, Shigeharu Kinoshita, Md Asaduzzaman, Fumito Omori, Kaoru Maeyama, Kiyohito Nagai, Shugo Watabe, and Shuichi Asakawa. Conserved and widespread expression of piRNA-like molecules and PIWI-like genes reveal dual functions of transposon silencing and gene regulation in Pinctada fucata (Mollusca). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Jiang:2021:DQP

[2792] Hongbo Jiang, Jie Bao, Jinghui Liu, Yuwen Chen, Chengcheng Feng, Xiaodong Li, Shuai Huang, and Qijun Chen. Development of a quantitative PCR method for specific and quantitative detection of Enterocytospora artemiae, a microsporidian parasite of Chinese grass shrimp (Palaemonetes sinensis). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Rixen:2021:ONT

[2793] Tim Rixen, Niko Lahajnar, Tarron Lamont, Rolf Koppelmann, Bettina Martin, Justus E. E. van Beusekom, Claire Siddiqui, Keshnee Pillay, and Luisa Meiritz. Oxygen and nutrient trapping in the Southern Benguela upwelling system. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Anilkumar:2021:RFW

[2794] N Anilkumar, Babula Jena, Jenson V. George, Sabu P, Kshitija S, and M Ravichandran. Recent freshening, warming, and contraction of the Antarctic bottom water in the Indian sector of the Southern Ocean. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Galvan:2021:NSE

[2795] Cristina Galván, Araceli Puente, and José A. Juanes. Nested socioecological maps as a spatial planning instrument for estuary conservation and ecosystem-based management. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Phlips:2021:CPR

[2796] Edward J. Phlips, Susan Badylak, Natalie G. Nelson, Lauren M. Hall, Charles A. Jacoby, Margaret A. Lasi, Jean C. Lockwood, and Janice D. Miller. Cyclical patterns and a regime shift in the character of phytoplankton blooms in a restricted sub-tropical lagoon, Indian River Lagoon, Florida, United States. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Bushra:2021:TSA

[2797] Nazla Bushra, Rubayet Bin Mostafiz, Robert V. Rohli, Carol J. Friedland, and Md Adilur Rahim. Technical and social approaches to study shoreline change of Kuakata, Bangladesh. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Torres:2021:ZDC

[2798] Andrew F. Torres, Darryl Anthony M. Valino, and Rachel Ravago-Gotanco. Zooxanthellae diversity and coral-symbiont associations in the Philippine Archipelago: Specificity and adaptability across thermal gradients. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ahmadia:2021:LPI

[2799] Gabby N. Ahmadia, Samantha H. Cheng, Dominic A. Andradi-Brown, Stacy K. Baez, Megan D. Barnes, Nathan J. Bennett, Stuart J. Campbell, Emily S. Darling, Estradivari, David Gill, Erika Gress, Georgina G. Gurney, Vera Horigue, Raymond Jakub, Emma V. Kennedy, Shauna L. Mahajan, Sangeeta Mangubhai, Shayle B. Matsuda, Nyawira A. Muthiga, Michael O. Navarro, Nadia Santodomingo, Henri Vallès, Laura Veverka, Angelo Villagomez, Amelia S. Wenger, and Adaoma Wosu. Limited progress in improving gender and geographic representation in coral reef science. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Beca-Carretero:2021:SAM

[2800] Pedro Beca-Carretero, Tomás Azcárate-García, Marc Julia-Miralles, Clara S. Stanschewski, Freddy Guihéneuf, and Dagmar B. Stengel. Seasonal acclimation modulates the impacts of simulated warming and light reduction on temperate seagrass productivity and biochemical composition. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:CMS

[2801] Tian-Shu Zhang, Ya-Ning Zhang, De-Chen Lu, Ya Gong, and Zong-Jun Du. Carboxylicivirga marinus sp. nov., isolated from marine sediment, and genome insight of the genus Carboxylicivirga. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Bao:2021:TPB

[2802] Xiaokai Bao, Yan Li, Jianbai Zhang, Xipan Chen, Xiaohui Xu, Yanwei Feng, Guohua Sun, Xiumei Liu, Bin Li, Weijun Wang, Zan Li, and Jianmin Yang. Transcriptome profiling based on different time points after hatching provides a core set of gene resource for understanding larval immune response mechanisms against Vibrio anguillarum infection in Amphioctopus fangsiao. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Marriott:2021:ICB

[2803] Sara E. Marriott, Courtney Cox, Rizaller C. Amolo, Dean Apistar, Roquelito H. Mancao, and Kim de Mutsert. Implications of communitybased management of marine reserves in the Philippines for reef fish communities and biodiversity. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hou:2021:SAP

[2804] Gang Hou, Hui Zhang, Jinrun Wang, Yanying Chen, and Jianbin Lin. Stock assessment of 19 Perciformes in the Beibu Gulf, China, using a length-based Bayesian biomass method. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Chang:2021:EIC

[2805] Yi-Jay Chang, Jhen Hsu, Po-Kai Lai, Kuo-Wei Lan, and Wen-Pei Tsai. Evaluation of the impacts of climate change on albacore distribution in the South Pacific Ocean by using ensemble forecast. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Kong:2021:TRF

[2806] Lingling Kong, Pingping Liu, Moli Li, Huizhen Wang, Jiaoxia Shi, Jingjie Hu, Yueru Li, and Xiaoli Hu. Transcriptional responses of flavincontaining monooxygenase genes in scallops exposed to PST-producing dinoflagellates implying their involvements in detoxification. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Chen:2021:DSV

[2807] Tzong-Yueh Chen, Chao-Chen Lai, Jen-Hua Tai, Chia-Ying Ko, and Fuh-Kwo Shiah. Diel to seasonal variation of picoplankton in the tropical South China Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lavoie:2021:GSL

[2808] Diane Lavoie, Nicolas Lambert, Michel Starr, Joël Chassé, Olivier Riche, Yvonnick Le Clainche, Kumiko Azetsu-Scott, Béchir Béjaoui, James R. Christian, and Denis Gilbert. The Gulf of St. Lawrence Biogeochemical Model: a modelling tool for fisheries and ocean management. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pedersen:2021:ORW

[2809] Torstein Pedersen, Nina Mikkelsen, Ulf Lindstrøm, Paul E. Renaud, Marcela C. Nascimento, Marie-Anne Blanchet, Ingrid H. Ellingsen, Lis L. Jørgensen, and Hugues Blanchet. Overexploitation, recovery, and warming of the Barents Sea ecosystem during 1950–2013. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pootakham:2021:NAB

[2810] Wirulda Pootakham, Chutima Sonthirod, Chaiwat Naktang, Narinratana Kongjandtre, Lalita Putchim, Duangjai Sangsrakru, Thippawan Yoocha, and Sithichoke Tangphatsornruang. De novo assembly of the brain coral Platygyra sinensis genome. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zaferani:2021:MAM

[2811] Sara Zaferani and Harald Biester. Mercury accumulation in marine sediments — a comparison of an upwelling area and two large river mouths. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745. See corrigendum [3176].

Appeldoorn:2021:IAV

[2812] Richard S. Appeldoorn, David L. Ballantine, Milton Carlo, Juan J. Cruz Motta, Michael Nemeth, Hector J. Ruiz, Nikolaos V. Schizas, Clark E. Sherman, Ernesto Weil, and Paul M. Yoshioka. Intra-annual variation in mesophotic benthic assemblages on the insular slope of Southwest Puerto Rico as a function of depth and geomorphology. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hou:2021:ASI

[2813] Gang Hou, Jinrun Wang, Limei Liu, Yanying Chen, Chuanhao Pan, Jianbin Lin, and Hui Zhang. Assemblage structure of the ichthyoplankton and its relationship with environmental factors in spring and autumn off the Pearl River Estuary. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Verma:2021:BFR

[2814] Komal Verma, Harshit Singh, Arun Deo Singh, Pradyumna Singh, Rajeev Kumar Satpathy, and Pothuri Divakar Naidu. Benthic foraminiferal response to the millennial-scale variations in monsoon-driven productivity and deep-water oxygenation in the Western Bay of Bengal during the last 45 ka. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Caputi:2021:CSF

[2815] Simona Sporta Caputi, Giulio Careddu, Edoardo Calizza, Federico Fiorentino, Deborah Maccapan, Loreto Rossi, and Maria Letizia Costantini. Corrigendum: Seasonal food web dynamics in the Antarctic benthos of Tethys Bay (Ross Sea): Implications for biodiversity persistence under different seasonal sea-ice coverage. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745. See [1189].

Whoriskey:2021:MBP

[2816] Frederick Whoriskey, Michele Barbier, Mackenzie Mazur, Tobias Hahn, Jacob Kritzer, and Richard Vallee. Maximizing benefits from punctual ocean infrastructure: an ethical perspective. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pang:2021:ABI

[2817] Wenhong Pang, Xiaoyan Zhou, Zhijun Dai, Shushi Li, Hu Huang, and Yaping Lei. ADV-based investigation on bed level changes over a mesomacro tidal beach. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Song:2021:OIS

[2818] Haibin Song, Yi Gong, Shengxiong Yang, and Yongxian Guan. Observations of internal structure changes in shoaling internal solitary waves based on seismic oceanography method. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Maximiliano-Cordova:2021:AIW

[2819] Carmelo Maximiliano-Cordova, M. Luisa Martínez, Rodolfo Silva, Patrick A. Hesp, Roger Guevara, and Rosario Landgrave. Assessing the impact of a winter storm on the beach and dune systems and erosion mitigation by plants. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Acharyya:2021:MMC

[2820] Akash Acharyya, Joydeep Das, and Kazi Nurul Hasan. Melatonin as a multipotent component of fish feed: Basic information for its potential application in aquaculture. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Feng:2021:ISW

[2821] Yingci Feng, Qunshu Tang, Jian Li, Jie Sun, and Wenhuan Zhan. Internal solitary waves observed on the continental shelf in the Northern South China Sea from acoustic backscatter data. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Xiao:2021:TSI

[2822] Wuxin Xiao, Katy Louise Sheen, Qunshu Tang, Jamie Shutler, Richard Hobbs, and Tobias Ehmen. Temperature and salinity inverted for a Mediterranean eddy captured with seismic data, using a spatially iterative Markov Chain Monte Carlo approach. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zydlewski:2021:WWL

[2823] Joseph Zydlewski, Daniel S. Stich, Samuel Roy, Michael Bailey, Timothy Sheehan, and Kenneth Sprankle. What have we lost? Modeling dam impacts on American shad populations through their native range. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sweat:2021:MCD

[2824] L. Holly Sweat, Hunter Alexander, Edward J. Phlips, and Kevin B. Johnson. Mesozooplankton community dynamics and grazing potential across algal bloom cycles in a subtropical estuary. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Chang:2021:SPM

[2825] Sarah L. Chang, Who-Seung Lee, and Stephan B. Munch. Separating paternal and maternal contributions to thermal transgenerational plasticity. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Yao:2021:MCS

[2826] Hongzhi Yao, Xing Li, Yuhao Chen, Guoling Liang, Gao Gao, Huan Wang, Chunlin Wang, and Changkao Mu. Metabolic changes in Scylla paramamosain during adaptation to an acute decrease in salinity. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lu:2021:ABE

[2827] Xin Lu, Jing Xu, Zhaodong Xu, and Xiaoshou Liu. Assessment of benthic ecological quality status using multi-biotic indices based on macrofaunal assemblages in a semi-enclosed bay. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Osadchiev:2021:SIA

[2828] Alexander Osadchiev, Dmitry Frey, Eduard Spivak, Sergey Shchuka, Natalia Tilinina, and Igor Semiletov. Structure and inter-annual variability of the freshened surface layer in the Laptev and East-Siberian Seas during ice-free periods. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Mejia-Quintero:2021:CST

[2829] Katherine Mejía-Quintero, Giomar H. Borrero-Pérez, and Erika Montoya-Cadavid. Callogorgia spp. and their brittle stars: Recording unknown relationships in the Pacific Ocean and the Caribbean Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Yang:2021:LOI

[2830] Rick J. Yang, James T. Liu, Chih-Chieh Su, Yi Chang, Jimmy J. Xu, and Hon-Kit Lui. Land–ocean interaction affected by the monsoon regime change in Western Taiwan Strait. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Saco:2021:ITP

[2831] Amaro Saco, Antonella Panebianco, Sofía Blanco, Beatriz Novoa, Angel P. Diz, and Antonio Figueras. Integration of transcriptomics and proteomics improves the characterization of the role of mussel gills in a bacterial waterborne infection. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Stashchuk:2021:IWD

[2832] Nataliya Stashchuk and Vasiliy Vlasenko. Internal wave dynamics over isolated seamount and its influence on coral larvae dispersion. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Santos:2021:ERE

[2833] Catarina Pereira Santos, Eduardo Sampaio, Beatriz P. Pereira, Maria Rita Pegado, Francisco O. Borges, Carolyn R. Wheeler, Ian A. Bouyoucos, Jodie L. Rummer, Catarina Frazão Santos, and Rui Rosa. Elasmobranch responses to experimental warming, acidification, and oxygen loss — a meta-analysis. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pita:2021:FAI

[2834] Pablo Pita, Gillian B. Ainsworth, Bernardino Alba, Antônio B. Anderson, Manel Antelo, Josep Alós, Iñaki Artetxe, Jérôme Baudrier, José J. Castro, Belén Chicharro, Karim Erzini, Keno Ferter, Mafalda Freitas, Laura García de-la Fuente, José A. García-Charton, María Giménez-Casalduero, Antoni M. Grau, Hugo Diogo, Ana Gordoa, Filipe Henriques, Kieran Hyder, David Jiménez-Alvarado, Paraskevi K. Karachle, Josep Lloret, Martin Laporta, Adam M. Lejk, Arnau L. Dedeu, Pablo Martín-Sosa, Lllibori Martínez, Antoni M. Mira, Beatriz Morales-Nin, Estanis Mugerza, Hans J. Olesen, Anastasios Papadopoulos, João Pontes, José J. Pascual-Fernández, Ariadna Purroy, Milena Ramires, Mafalda Rangel, José Amorim Reis-Filho, Jose L. Sánchez-Lizaso, Virginia Sandoval, Valerio Sbragaglia, Luis Silva, Christian Skov, Iván Sola, Harry V. Strehlow, María A. Torres, Didzis Ustups, Tessa van der Hammen, Pedro Veiga, Leonardo A. Venerus, Thomas Verleye, Sebastián Villasante, Marc Simon Weltersbach, and Lucía Zarauz. First assessment of the impacts of the COVID-19 pandemic on global marine recreational fisheries. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Song:2021:NNP

[2835] Xiaoying Song, Xinyu Zhai, Shuang Hao, Lixia Shang, Yunyan Deng, Zhaoyang Chai, Junhui Chen, Zhangxi Hu, and Ying Zhong Tang. Novel non-paralytic shellfish toxin and non-spirolide toxicity to finfish, brine shrimp, and rotifer observed in a culture of the dinoflagellate Alexandrium insuetum isolated from the coastal water of China. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Hammond:2021:EAM

[2836] Philip S. Hammond, Tessa B. Francis, Dennis Heinemann, Kristy J. Long, Jeffrey E. Moore, André E. Punt, Randall R. Reeves, Maritza Sepúlveda, Gudjón Már Sigurdsson, Margaret C. Siple, Gísli Víkingsson, Paul R. Wade, Rob Williams, and Alexandre N. Zerbini. Estimating the abundance of marine mammal populations. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Golubkov:2021:RBN

[2837] Mikhail Golubkov and Sergey Golubkov. Relationships between Northern Hemisphere teleconnection patterns and phytoplankton productivity in the Neva Estuary (Northeastern Baltic Sea). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhang:2021:RIL

[2838] Hao-Ran Zhang, Yuntao Wang, Peng Xiu, Yiquan Qi, and Fei Chai. Roles of iron limitation in phytoplankton dynamics in the Western and Eastern Subarctic Pacific. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Khuyen:2021:ICM

[2839] Vo Thi Kim Khuyen, Dinh Vu Le, Le Hung Anh, Axel René Fischer, and Christina Dornack. Investigating the correlation of microplastic pollution between seawater and marine salt using micro-Raman spectroscopy. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Dadras:2021:RMA

[2840] Hadiseh Dadras, Tomas Policar, Dmitry Loginov, Sergii Boryshpolets, Yana Loginova, Sabine Sampels, Anatolii Sotnikov, Vitaliy Kholodnyy, Jan Sterba, Miaomiao Xin, and Borys Dzyuba. Relationship of motility activation to lipid composition, protein profile, and swelling rate of burbot Lota lota spermatozoon following change of temperature and osmolality. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ge:2021:FLT

[2841] Hailong Ge, Haoyu Zhang, Lijun Yang, Haoyu Wang, Limei Tu, Zhuojin Jiang, Jing Zheng, Bolin Chen, Juan Chen, Yun Li, and Zhijian Wang. Full-length transcriptome sequencing from the longest-lived freshwater bony fish of the world: Bigmouth buffalo (Ictiobus cyprinellus). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Lin:2021:EWH

[2842] Junfang Lin, Peter I. Miller, Bror F. Jönsson, and Michael Bedington. Early warning of harmful algal bloom risk using satellite ocean color and Lagrangian particle trajectories. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Khan:2021:EOA

[2843] Fahim Ullah Khan, Yueyong Shang, Xueqing Chang, Hui Kong, Amina Zuberi, James K. H. Fang, Wei Liu, Jinxia Peng, Xingzhi Zhang, Menghong Hu, and Youji Wang. Effects of ocean acidification, hypoxia, and warming on the gut microbiota of the thick shell mussel Mytilus coruscus through 16S rRNA gene sequencing. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Liu:2021:STD

[2844] Quanhong Liu, Ren Zhang, Yangjun Wang, Hengqian Yan, and Mei Hong. Short-term daily prediction of sea ice concentration based on deep learning of gradient loss function. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cao:2021:BMO

[2845] Jie Cao, Xuefang Wang, Matthew D. Damiano, Cheng Zhou, and Jiangfeng Zhu. A Bayesian multilevel ordinal regression model for fish maturity data: Difference in maturity ogives of skipjack tuna (Katsuwonus pelamis) between schools in the Western and Central Pacific Ocean. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Yi:2021:CDP

[2846] Xiangqi Yi, Di Zhang, Jiazhen Sun, John Beardall, and Kunshan Gao. Cyanobacteria-dominated phytoplankton in the oligotrophic South China Sea maintain photosynthetic potential despite diurnal photoinactivation of PSII. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Chambault:2021:ETF

[2847] Philippine Chambault, Philippe Gaspar, and Florence Dell’Amico. Ecological trap or favorable habitat? First evidence that immature sea turtles may survive at their range-limits in the North-East Atlantic. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Garcia-Bonilla:2021:HHF

[2848] Erika García-Bonilla, Diego Chaves-Moreno, Diego Riaño-Pachón, Wilson Terán, Alberto Acosta, and Howard Junca. The hologenome of Haliclona fulva (Porifera, Demospongiae) reveals an abundant and diverse viral community. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Sutton:2021:EFI

[2849] Lauren Sutton, Franz J. Mueter, Bodil A. Bluhm, and Katrin Iken. Environmental filtering influences functional community assembly of epibenthic communities. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Pacheco:2021:EWW

[2850] Aldo S. Pacheco, Maritza Sepúlveda, and Peter Corkeron. Editorial: Whale-watching impacts: Science, human dimensions and management. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Villasante:2021:RAC

[2851] Sebastian Villasante, Ana Tubío, Gillian Ainsworth, Pablo Pita, Manel Antelo, and José María Da-Rocha. Rapid assessment of the COVID-19 impacts on the Galician (NW Spain) seafood sector. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Satterthwaite:2021:EFG

[2852] Erin V. Satterthwaite, Nicholas J. Bax, Patricia Miloslavich, Lavenia Ratnarajah, Gabrielle Canonico, Daniel Dunn, Samantha E. Simmons, Roxanne J. Carini, Karen Evans, Valerie Allain, Ward Appeltans, Sonia Batten, Lisandro Benedetti-Cecchi, Anthony T. F. Bernard, Sky Bristol, Abigail Benson, Pier Luigi Buttigieg, Leopoldo Cavaleri Gerhardinger, Sanae Chiba, Tammy E. Davies, J. Emmett Duffy, Alfredo Giron-Nava, Astrid J. Hsu, Alexandra C. Kraberg, Raphael M. Kudela, Dan Lear, Enrique Montes, Frank E. Muller-Karger, Todd D. O’Brien, David Obura, Pieter Provoost, Sara Pruckner, Lisa-Maria Rebelo, Elizabeth R. Selig, Olav Sigurd Kjesbu, Craig Starger, Rick D. Stuart-Smith, Marjo Vierros, John Waller, Lauren V. Weatherdon, Tristan P. Wellman, and Anna Zivian. Establishing the foundation for the global observing system for marine life. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gaspar:2021:IAS

[2853] Lisa Gaspar, Patrick Flammang, Ricardo José, Ricardo Luis, Patrício Ramalhosa, João Monteiro, Natacha Nogueira, João Canning-Clode, and Romana Santos. Interspecific analysis of sea urchin adhesive composition emphasizes variability of glycans conjugated with putative adhesive proteins. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Steinman:2021:CCA

[2854] K. J. Steinman, G. A. Montano, and T. R. Robeck. Characterization of circulating androgens, cortisol and estrogens during normal, abnormal and false pregnancy in bottlenose dolphins (Tursiops truncatus) under managed care. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Shim:2021:RTE

[2855] JeongHee Shim, Mi-Ju Ye, Jae-Hyun Lim, Jung-No Kwon, and Jeong Bae Kim. Red Tide events and seasonal variations in the partial pressure of CO_{2} and related parameters in shellfish-farming bays, Southeastern Coast of Korea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lopez-Garcia:2021:EES

[2856] Patricia Lopez-Garcia, Juan J. Muñoz-Perez, Antonio Contreras, Juan Vidal, Bismarck Jigena, Juan J. Santos, Jeanette Romero, and Francisco Contreras. Error on the estimation of sand size parameters when using small diameter sieves and a solution. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Park:2021:IBK

[2857] Sang Ah Park, Hae Jin Jeong, Jin Hee Ok, Hee Chang Kang, Ji Hyun You, Se Hee Eom, and Eun Chong Park. Interactions between the kleptoplastidic dinoflagellate Shimiella gracilenta and several common heterotrophic protists. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zeghal:2021:PRM

[2858] Emna Zeghal, Annika Vaksmaa, Hortense Vielfaure, Teun Boekhout, and Helge Niemann. The potential role of marine fungi in plastic degradation — a review. Frontiers in Marine Science, 8, November 2021. ISSN 2296-

Carreras:2021:TSM

[2859] Carlos Carreras, Víctor Ordóñez, Àlex García-Cisneros, Owen S. Wangensteen, Creu Palacín, Marta Pascual, and Xavier Turon. The two sides of the Mediterranean: Population genomics of the Black Sea urchin Arbacia lixula (Linnaeus, 1758) in a warming sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yu:2021:EHA

[2860] Zonghe Yu, Yingqiu Zhang, Lei Zhou, and Hongyan Sun. The effects of hypoxia and anthropogenic disturbance on the distribution, abundance and recruitment of the sea cucumber Holothuria leucospilota at the mouth of a subtropical bay, China. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zou:2021:PTR

[2861] Desheng Zou, Junhao Ning, Xia Lu, Xia Wang, Min Chen, Bo Liu, Jun Fang, and Chunde Wang. Physiological and transcriptional responses to acute and chronic thermal stress in the ark shell Scapharca subcrenata. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Liu:2021:SDM

[2862] Tong Liu, Shijun Yu, Xiaoshan Zhu, Ran Liao, Zepeng Zhuo, Yanping He, and Hui Ma. In-situ detection method for microplastics in water by polarized light scattering. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gostischa:2021:MSF

[2863] Julia Gostischa, Alessandro Massolo, and Rochelle Constantine. Multispecies feeding association dynamics driven by a large generalist predator. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Choudhary:2021:ESP

[2864] Babita Choudhary, O. P. Chauhan, and Avinash Mishra. Edible seaweeds: a potential novel source of bioactive metabolites and nutraceuticals with human health benefits. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Chen:2021:SPB

[2865] Ying Chen, Haiyi Shi, and Hui Zhao. Summer phytoplankton blooms induced by upwelling in the Western South China Sea. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Liu:2021:PRA

[2866] An Liu, Wenyuan Shi, Dongdong Lin, and Haihui Ye. A possible role of allatostatin C in inhibiting ecdysone biosynthesis revealed in the mud crab Scylla paramamosain. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Vasconcelos:2021:TFP

[2867] Joana Vasconcelos, Alba Jurado-Ruzafa, José Luis Otero-Ferrer, Antoni Lombarte, Rodrigo Riera, and Víctor M. Tuset. Thinking of fish population discrimination: Population average phenotype vs. population phenotypes. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gong:2021:PRS

[2868] Sanqiang Gong, Gang Li, Xuejie Jin, Dajun Qiu, Jiayuan Liang, Kefu Yu, Yehui Tan, Xiao Ma, and Xiaomin Xia. Photobleaching and recovery of Symbiodiniaceae Effrenium voratum SCS01 reveals life form transformation under thermal stress. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lebbe:2021:DCA

[2869] Théophile Bongarts Lebbe, Hélène Rey-Valette, Éric Chaumillon, Guigone Camus, Rafael Almar, Anny Cazenave, Joachim Claudet, Nicolas Rocle, Catherine Meur-Férec, Frédérique Viard, Denis Mercier, Christine Dupuy, Frédéric Ménard, Bernardo Aliaga Rossel, Lauren Mullineaux, Marie-Alexandrine Sicre, Anna Zivian, Françoise Gaill, and Agathe Euzen. Designing coastal adaptation strategies to tackle sea level rise. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yin:2021:ELH

[2870] Bin Yin, Hongyu Liu, Beiping Tan, Xiaohui Dong, Shuyan Chi, Qihui Yang, and Shuang Zhang. Effects of low- and high-level gossypol and sodium butyrate supplementation under high-level gossypol condition on growth performance and intestinal health of hybrid grouper (Epinephelus fuscoguttatus[Venus sign] × Epinephelus lanceolatus [male sign]). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Listmann:2021:DCA

[2871] Luisa Listmann, Franziska Kerl, Nele Martens, and C-Elisa Schaum. Differences in carbon acquisition could explain adaptive responses in a Baltic Sea pico-phytoplankton. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cesbron:2021:PES

[2872] Guillaume Cesbron, Angélique Melet, Rafael Almar, Anne Lifermann, Damien Tullot, and Laurence Crosnier. Pan-European satellite-derived coastal bathymetry — review, user needs and future services. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kalra:2021:MMD

[2873] Tarandeep S. Kalra, Neil K. Ganju, Alfredo L. Aretxabaleta, Joel A. Carr, Zafer Defne, and Julia M. Moriarty. Modeling marsh dynamics using a 3-D coupled wave-flow-sediment model. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Liggins:2021:IMP

[2874] Libby Liggins and Cory Noble. The Ira Moana Project: a genetic observatory for Aotearoa’s marine biodiversity. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Waycott:2021:GBP

[2875] Michelle Waycott, Kor jent van Dijk, Ainsley Calladine, Eric Bricker, and Ed Biffin. Genomics-based phylogenetic and population genetic analysis of global samples confirms Halophila johnsonii Eiseman as Halophila ovalis (R.Br.) Hook.f. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Osadchiev:2021:WEB

[2876] Alexander Osadchiev, Olga Konovalova, and Alexandra Gordey. Water exchange between the Gulf of Ob and the Kara Sea during ice-free seasons: The roles of river discharge and wind forcing. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hu:2021:CIC

[2877] Liping Hu, Liming Jiang, Qiang Xing, Zujing Yang, Qiang Zhao, Liyong Wang, Xiaoting Huang, and Zhenmin Bao. Chromosome identification and cytogenetic map construction of Zhikong scallop (Chlamys farreri) based on fluorescence in situ hybridization. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Salvi:2021:YAM

[2878] Daniele Salvi, Emanuele Berrilli, Matteo Garzia, and Paolo Mariottini. Yet another mitochondrial genome of the Pacific cupped oyster: The published mitogenome of Alectryonella plicatula (Ostreinae) is based on a misidentified Magallana gigas (Crassostreinae). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Cossarini:2021:HRR

[2879] Gianpiero Cossarini, Laura Feudale, Anna Teruzzi, Giorgio Bolzon, Gianluca Coidessa, Cosimo Solidoro, Valeria Di Biagio, Carolina Amadio, Paolo Lazzari, Alberto Brosich, and Stefano Salon. High-resolution reanalysis of the Mediterranean Sea biogeochemistry (1999–2019). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lu:2021:DAB

[2880] Zhenzhen Lu, Yunfei Sun, Changlun Xiao, and Yongxu Cheng. Dietary analysis based on 18S rDNA, and stable carbon and nitrogen isotopes in juvenile Eriocheir sinensis crabs reared under three feeding modes. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Yi:2021:SHM

[2881] Jun Yi, Linus Shing Him Lo, Hongbin Liu, Pei-Yuan Qian, and Jinping Cheng. Study of heavy metals and microbial communities in contaminated sediments along an urban estuary. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Baumann:2021:EIM

[2882] Moritz Baumann, Jan Taucher, Allanah J. Paul, Malte Heinemann, Mari Vanharanta, Lennart T. Bach, Kristian Spilling, Joaquin Ortiz, Javier Arístegui, Nauzet Hernández-Hernández, Isabel Baños, and Ulf Riebesell. Effect of intensity and mode of artificial upwelling on particle flux and carbon export. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Atmore:2021:SBT

[2883] Lane M. Atmore, Magie Aiken, and Fabricio Furni. Shifting baselines to thresholds: Reframing exploitation in the marine environment. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Goldsmit:2021:KEC

[2884] Jesica Goldsmit, Robert W. Schlegel, Karen Filbee-Dexter, Kathleen A. MacGregor, Ladd E. Johnson, Christopher J. Mundy, Amanda M. Savoie, Christopher W. McKindsey, Kimberly L. Howland, and Philippe Archambault. Kelp in the Eastern Canadian Arctic: Current and future predictions of habitat suitability and cover. Frontiers in Marine Science, 18, October 2021. ISSN 2296-7745.

Ning:2021:EIS

[2885] Dezhi Ning, Lars Johanning, Zhen Gao, Philipp Rudolf Thies, Wei Shi, and Kathleen Aviso. Editorial: Innovation and system integration for offshore renewable energy structures. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhu:2021:GWA

[2886] Xinghai Zhu, Junhao Zhang, Xiujiang Hou, Pingping Liu, Jia Lv, Qiang Xing, Xiaoting Huang, Jingjie Hu, and Zhenmin Bao. A genome-wide association study identifies candidate genes associated with shell color in bay scallop Argopecten irradians irradians. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Kleinhans:2021:SMT

[2887] Maarten G. Kleinhans, Harke Douma, Elisabeth A. Addink, Laura Coumou, Tessa Deggeller, Robert Jentink, Edwin Paree, and Jelmer Cleveringa. Salt marsh and tidal flat area distributions along three estuaries. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Xing:2021:DER

[2888] Qikun Xing, Miriam Bernard, Sylvie Rousvoal, Erwan Corre, Gabriel V. Markov, Akira F. Peters, and Catherine Leblanc. Different early responses of laminariales to an endophytic infection provide insights about kelp host specificity. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Chen:2021:CUP

[2889] Tiantian Chen, Yun Liu, Zhangxi Hu, Shuqun Song, and Caiwen Li. Chloroplast ultrastructure and photosynthetic response of the dinoflagellate Akashiwo sanguinea throughout infection by Amoebophrya sp. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Liu:2021:FCC

[2890] Yang Liu, Jun Sun, Xingzhou Wang, Xiaofang Liu, Xi Wu, Zhuo Chen, Ting Gu, Weiqiang Wang, Linghui Yu, Yu Guo, Yujian Wen, Guodong Zhang, and Guicheng Zhang. Fluorescence characteristics of chromophoric dissolved organic matter in the Eastern Indian Ocean: a case study of three subregions. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Roner:2021:DDS

[2891] Marcella Roner, Massimiliano Ghinassi, Alvise Finotello, Adele Bertini, Nathalie Combourieu-Nebout, Sandra Donnici, Adrian Gilli, Martina Vannacci, Luigi Vigliotti, Luca G. Bellucci, Mariaelena Fedi, Lucia Liccioli, Laura Tommasini, and Andrea D’Alpaos. Detecting the delayed signatures of changing sediment supply in salt-marsh landscapes: The case of the Venice Lagoon (Italy). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ritter:2021:PCA

[2892] Carmen J. Ritter, Leah M. Harper, Jonathan S. Lefcheck, Valerie J. Paul, Ross Whippo, Scott Jones, Matthew B. Ogburn, and J. Emmett Duffy. Patterns of consumption across a Caribbean seascape: Roles of habitat and consumer species composition through time. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pengpeng:2021:DSI

[2893] Wang Pengpeng, Zhang Fang, Guo Dongjie, and Sun Song. Diets and seasonal ingestion rates of Aurelia coerulea (Cnidaria: Scyphozoa) polyps by in situ feeding experiments in Jiaozhou Bay, China. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Tsiaras:2021:MPA

[2894] Kostas Tsiaras, Yannis Hatzonikolakis, Sofia Kalaroni, Annika Pollani, and George Triantafyllou. Modeling the pathways and accumulation patterns of micro- and macro-plastics in the Mediterranean. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hop:2021:IAA

[2895] Haakon Hop, Mikko Vihtakari, Bodil A. Bluhm, Malin Daase, Rolf Gradinger, and Igor A. Melnikov. Ice-associated amphipods in a pan-Arctic scenario of declining sea ice. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Li:2021:AEG

[2896] Wenhao Li, Faxiang Wang, Shouwen Jiang, Binbin Pan, Jiulin Chan, and Qianghua Xu. The adaptive evolution and gigantism mechanisms of the hadal “Supergiant” amphipod Alicella gigantea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Wang:2021:EZW

[2897] Qiong Wang, Jianjian Lv, Xianyun Ren, Jiajia Wang, Shaoting Jia, Yuying He, and Jian Li. Evidences of Z- and W-linked regions on the genome of Fenneropenaeus chinensis. Frontiers in Marine Science, 8, August 2021. ISSN 2296-7745.

Yuan:2021:RSE

[2898] Jigui Yuan, Hungdu Lin, Lisheng Wu, Xuan Zhuang, Junkai Ma, Bin Kang, and Shaoxiong Ding. Resource status and effect of long-term stock enhancement of large yellow croaker in China. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Gauthier:2021:GPA

[2899] Jeff Gauthier, Hélène Marquis, Valérie E. Paquet, Steve J. Charette, Roger C. Levesque, and Nicolas Derome. Genomic perspectives on Aeromonas salmonicida subsp. salmonicida strain 890054 as a model system for pathogenicity studies and mitigation of fish infections. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Guzman:2021:NIG

[2900] Héctor M. Guzmán, Caitlin E. Beaver, and Edgardo Díaz-Ferguson. Novel insights into the genetic population connectivity of transient whale sharks (Rhincodon typus) in Pacific Panama provide crucial data for conservation efforts. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hewitt:2021:RMB

[2901] Olivia H. Hewitt and Hisham M. Shaikh. The rhythm of many: Biological rhythms in the marine environment, from macro-scale planktonic ecosystems to micro-scale Holobionts. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Franks:2021:STV

[2902] Bryan R. Franks, John P. Tyminski, Nigel E. Hussey, Camrin D. Braun, Alisa L. Newton, Simon R. Thorrold, George C. Fischer, Brett McBride, and Robert E. Hueter. Spatio-temporal variability in white shark (Carcharodon carcharias) movement ecology during residency and migration phases in the Western North Atlantic. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Christensen:2021:CDF

[2903] Asbjørn Christensen, Kostas Tsiaras, Jens Murawski, Yannis Hatzonikolakis, Jun She, Michael St. John, Urmas Lips, and Roy Brouwer. A cross disciplinary framework for cost-benefit optimization of marine litter cleanup at regional scale. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Liu:2021:DCR

[2904] Yanzi Liu, Hanhan Yao, Tingting Zhou, Zhihua Lin, and Yinghui Dong. The discovery of circadian rhythm of feeding time on digestive enzymes activity and their gene expression in Sinonovacula constricta within a light/dark cycle. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Chen:2021:EBV

[2905] Lizhu Chen, Chengjie Lv, Bin Li, Huawei Zhang, Lihua Ren, Qianqian Zhang, Xiaoli Zhang, Jiqing Gao, Chunxiao Sun, and Shunxin Hu. Effects of Bacillus velezensis supplementation on the growth performance, immune responses, and intestine microbiota of Litopenaeus vannamei. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Reynolds:2021:YRM

[2906] S. James Reynolds, Colin P. Wearn, B. John Hughes, Roger C. Dickey, Lucy J. H. Garrett, Sean Walls, Fay T. Hughes, Nicola Weber, Sam B. Weber, Eliza H. K. Leat, Kenickie Andrews, Jaime A. Ramos, and Vitor H. Paiva. Year-round movements of sooty terns (Onychoprion fuscatus) nesting within one of the Atlantic’s largest Marine Protected Areas. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:IAT

[2907] Chaogang Wang, Ao Li, Wei Wang, Rihao Cong, Luping Wang, Guofan Zhang, and Li Li. Integrated application of transcriptomics and metabolomics reveals the energy allocation-mediated mechanisms of growth-defense trade-offs in Crassostrea gigas and crassostrea angulata. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ramirez-Amaro:2021:GME

[2908] Sergio Ramírez-Amaro, Francesc Ordines, Ronald Fricke, Ignacio Ruiz-Jarabo, Ignacio Bolado, and Enric Massutí. Genetic and morphological evidence to split the Coris julis species complex (Teleostei: Labridae) into two sibling species: Resurrection of Coris melanura (Lowe, 1839) redescription of Coris julis (Linnaeus, 1758). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Liu:2021:IST

[2909] Hao Liu, Zexun Wei, Ingo Richter, Xunwei Nie, and Chuanshun Li. Influence of salinity and temperature gradients on the variability of the North Brazil Undercurrent. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Perez:2021:SCH

[2910] Maëva Perez, Jin Sun, Qinzeng Xu, and Pei-Yuan Qian. Structure and connectivity of hydrothermal vent communities along the mid-ocean ridges in the West Indian Ocean: a review. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Pearce:2021:SSR

[2911] Jessica Pearce, Matthew W. Fraser, Ana M. M. Sequeira, and Parwinder Kaur. State of shark and ray genomics in an era of extinction. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Giossi:2021:LIC

[2912] Chiara E. Giossi, Sónia Cruz, Felisa Rey, Rúben Marques, Tânia Melo, Maria do Rosário Domingues, and Paulo Cartaxana. Light induced changes in pigment and lipid profiles of bryopsidales algae. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Surugiu:2021:HRB

[2913] Victor Surugiu, Adrian Teacă, Ilie Şvedu, and Pedro A. Quijón. A hotspot in the Romanian Black Sea: Eelgrass beds drive local biodiversity in surrounding bare sediments. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Proud:2021:UPP

[2914] Roland Proud, Camille Le Guen, Richard B. Sherley, Akiko Kato, Yan Ropert-Coudert, Norman Ratcliffe, Simon Jarman, Adam Wyness, John P. Y. Arnould, Ryan A. Saunders, Paul G. Fernandes, Lars Boehme, and Andrew S. Brierley. Using predicted patterns of 3D prey distribution to map king penguin foraging habitat. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

French:2021:SSM

[2915] Lindsay A. French, Stephen R. Midway, David H. Evans, and George H. Burgess. Shark side of the moon: Are shark attacks related to lunar phase? Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Curbelo-Hernandez:2021:ASC

[2916] David Curbelo-Hernández, J. Magdalena Santana-Casiano, Aridane González González, and Melchor González-Dávila. Air–sea CO_{2} exchange in the Strait of Gibraltar. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Levy:2021:PTU

[2917] Tom Levy, Valerio Zupo, Mirko Mutalipassi, Emanuele Somma, Nadia Ruocco, Maria Costantini, Shai Abehsera, Rivka Manor, Vered Chalifa-Caspi, Amir Sagi, and Eliahu D. Aflalo. Protandric transcriptomes to uncover parts of the crustacean sex-differentiation puzzle. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Charrier:2021:GLC

[2918] Bénédicte Charrier, Samuel Boscq, Bradley J. Nelson, and Nino F. Läubli. Growth and labelling of cell wall components of the brown alga ectocarpus in microfluidic chips. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Blenckner:2021:RND

[2919] Thorsten Blenckner, Yosr Ammar, Bärbel Müller-Karulis, Susa Niiranen, Lars Arneborg, and Qiang Li. The risk for novel and disappearing environmental conditions in the Baltic Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fu:2021:CFA

[2920] Lulu Fu, Heming Shi, Wenfang Dai, Hanhan Yao, Yongbo Bao, Zhihua Lin, and Yinghui Dong. Characterization and function analysis of β, βcarotene-9′, 10′-oxygenase 2 (BCDO2) gene in carotenoid metabolism of the red shell hard clam (Meretrix meretrix). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Ren:2021:ELT

[2921] Xianyun Ren, Qiong Wang, Huixin Shao, Yao Xu, Ping Liu, and Jian Li. Effects of low temperature on shrimp and crab physiology, behavior, and growth: a review. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Silva:2021:TOY

[2922] Edson Silva, François Counillon, Julien Brajard, Anton Korosov, Lasse H. Pettersson, Annette Samuelsen, and Noel Keenlyside. Twentyone years of phytoplankton bloom phenology in the Barents, Norwegian, and North Seas. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Marx:2021:MMC

[2923] Lukas Marx, Susana Flecha, Marlene Wesselmann, Carlos Morell, and Iris Eline Hendriks. Marine macrophytes as carbon sinks: Comparison between seagrasses and the non-native alga Halimeda incrassata in the Western Mediterranean (Mallorca). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Siegfried:2021:IST

[2924] Tabitha Siegfried, Christopher Noren, Jackson Reimer, Matthew Ware, Mariana M. P. B. Fuentes, and Susan E. Piacenza. Insights into sea turtle population composition obtained with stereo-video cameras in situ across nearshore habitats in the Northeastern Gulf of Mexico. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zhong:2021:MGE

[2925] Liqiang Zhong, Minghua Wang, Daming Li, Shengkai Tang, and Xiaohui Chen. Mitochondrial genome of Eleutheronema rhadinum with an additional non-coding region and novel insights into the phylogenetics. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Zhou:2021:PTE

[2926] Liang Zhou, Shu Gao, Jianjun Jia, Yang Yang, Changliang Tong, and Aijun Wang. Paleo-typhoon events as indicated by coral reef boulder deposits on the Southern Coast of Hainan Island, China. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Dong:2021:QAH

[2927] Yan-Zou Dong, Tian Xia, Ji-Bin Lin, Ling Wang, Kai Song, and Chun-Xiao Zhang. Quercetin attenuates high-fat diet-induced excessive fat deposition of spotted seabass (Lateolabrax maculatus) through the regulatory for mitochondria and endoplasmic reticulum. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Huang:2021:WCA

[2928] Chenfu Huang, Longhuan Zhu, Gangfeng Ma, Guy A. Meadows, and Pengfei Xue. Wave climate associated with changing water level and ice cover in Lake Michigan. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Whitney:2021:FCC

[2929] Michael M. Whitney, Yan Jia, Kelly L. Cole, Daniel G. MacDonald, and Kimberly D. Huguenard. Freshwater composition and connectivity of the Connecticut River plume during ambient flood tides. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Athanasiou:2021:CAP

[2930] Panagiotis Athanasiou, Ap van Dongeren, Alessio Giardino, Michalis Vousdoukas, Jose A. A. Antolinez, and Roshanka Ranasinghe. A clustering approach for predicting dune morphodynamic response to storms using typological coastal profiles: a case study at the Dutch coast. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Shih:2021:CPP

[2931] Yung-Yen Shih, Fuh-Kwo Shiah, Chao-Chen Lai, Wen-Chen Chou, Jen-Hua Tai, Yu-Shun Wu, Cheng-Yang Lai, Chia-Ying Ko, and Chin-Chang Hung. Comparison of primary production using in situ and satellitederived values at the SEATS Station in the South China Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fang:2021:HEF

[2932] Xiaoyu Fang, Francesco Cozzoli, Sven Smolders, Antony Knights, Tom Moens, Karline Soetaert, and Carl Van Colen. Hindcasting ecosystem functioning change in an anthropogenized estuary: Implications for an era of global change. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Dong:2021:ACC

[2933] Shipeng Dong, Fang Wang, Dongxu Zhang, Liye Yu, Weijia Pu, Xian Xu, and Yicheng Xie. Assessment of the carrying capacity of integrated pond aquaculture of Portunus trituberculatus at the ecosystem level. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Roch:2021:STA

[2934] Marisa Roch, Peter Brandt, Sunke Schmidtko, Filomena Vaz Velho, and Marek Ostrowski. Southeastern Tropical Atlantic changing from subtropical to tropical conditions. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yamhure:2021:HLC

[2935] Gabriela M. Yamhure, Nathalie Reyns, and Jesús Pineda. High larval concentrations and onshore transport of barnacle cyprids associated with thermal stratification. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ru:2021:PLA

[2936] Xiaoshang Ru, Libin Zhang, and Hongsheng Yang. Plasticity of locomotor activity permits energy homeostasis during reproduction in a female sea cucumber. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Schone:2021:OTM

[2937] Bernd R. Schöne and Qian Huang. Ontogenetic δ^{15}N trends and multidecadal variability in shells of the bivalve mollusk, Arctica islandica. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Weerathunga:2021:IPF

[2938] Veran Weerathunga, Wei-Jen Huang, Sam Dupont, Hsueh-Han Hsieh, Nathangi Piyawardhana, Fei-Ling Yuan, Jhe-Syuan Liao, Chia-Yu Lai, Wei-Ming Chen, and Chin-Chang Hung. Impacts of pH on the fitness and immune system of Pacific white shrimp. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Yang:2021:BMP

[2939] Yanan Yang, Peng Zhang, Zhaoxia Cui, and Chenchang Bao. Bone morphogenetic protein 2 is involved in oocyte maturation through an autocrine/paracrine pathway in Scylla paramamosain. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Petersen:2021:PBP

[2940] Lars-Erik Petersen, Matthias Y. Kellermann, Samuel Nietzer, and Peter J. Schupp. Photosensitivity of the bacterial pigment cycloprodigiosin enables settlement in coral larvae — light as an understudied environmental factor. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hanafi-Portier:2021:WIP

[2941] Mélissa Hanafi-Portier, Sarah Samadi, Laure Corbari, Tin-Yam Chan, Wei-Jen Chen, Jhen-Nien Chen, Mao-Ying Lee, Christopher Mah, Thomas Saucède, Catherine Borremans, and Karine Olu. When imagery and physical sampling work together: Toward an integrative methodology of deep-sea image-based megafauna identification. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

DSa:2021:RSD

[2942] Eurico J. D’Sa, Hyun-Cheol Kim, Sun-Yong Ha, and Ishan Joshi. Ross Sea dissolved organic matter optical properties during an austral summer: Biophysical influences. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Nepper-Davidsen:2021:IGS

[2943] Jacob Nepper-Davidsen, Marie Magnusson, Christopher R. K. Glasson, Philip M. Ross, and Rebecca J. Lawton. Implications of genetic structure for aquaculture and cultivar translocation of the kelp Ecklonia radiata in Northern New Zealand. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Chang:2021:NAE

[2944] Tzu-Yin Chang, Hongey Chen, Shih-Chun Hsiao, Han-Lun Wu, and Wei-Bo Chen. Numerical analysis of the effect of binary typhoons on ocean surface waves in waters surrounding Taiwan. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lin:2021:DRI

[2945] Chia-Hao Lin, Po-Ling Yeh, Yu-Chun Wang, and Tsung-Han Lee. Dynamic regulation of ions and amino acids in adult Asian hard clams Meretrix lusoria upon hyperosmotic salinity. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Xie:2021:HRO

[2946] Zhe Xie, Shuaishuai Wei, Haomiao Dong, Hui Chen, Qianqian Zhang, Wei Liu, Jinxia Peng, Inna M. Sokolova, Menghong Hu, and Youji Wang. Hemocyte responses of the oyster Crassostrea hongkongensis exposed to diel-cycling hypoxia and salinity change. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Pegg:2021:EEA

[2947] Nicole Pegg, Irene T. Roca, Danielle Cholewiak, Genevieve E. Davis, and Sofie M. Van Parijs. Evaluating the efficacy of acoustic metrics for understanding baleen whale presence in the Western North Atlantic Ocean. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bierlich:2021:CUA

[2948] K. C. Bierlich, Joshua Hewitt, Clara N. Bird, Robert S. Schick, Ari Friedlaender, Leigh G. Torres, Julian Dale, Jeremy Goldbogen, Andrew J. Read, John Calambokidis, and David W. Johnston. Comparing uncertainty associated with 1-, 2-, and 3D aerial photogrammetry-based body condition measurements of baleen whales. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Jazdzewska:2021:PBD

[2949] Anna M. Jażdżewska, Tammy Horton, Ed Hendrycks, Tomasz Mamos, Amy C. Driskell, Saskia Brix, and Pedro Martínez Arbizu. Pandora’s Box in the deep sea — intraspecific diversity patterns and distribution of two congeneric scavenging amphipods. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Harrison:2021:ICL

[2950] Peter L. Harrison, Dexter W. dela Cruz, Kerry A. Cameron, and Patrick C. Cabaitan. Increased coral larval supply enhances recruitment for coral and fish habitat restoration. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Xu:2021:RWE

[2951] Chunyang Xu, Chunyan Zhou, Kao Ma, Peng Wang, and Xihe Yue. Response of water environment to land reclamation in Jiaozhou Bay, China over the last 150 years. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Jiao:2021:ILD

[2952] Lefei Jiao, Tianmeng Dai, Xinyue Tao, Jingjing Lu, and Qicun Zhou. Influence of light/dark cycles on body color, hepatopancreas metabolism, and intestinal microbiota homeostasis in Litopenaeus vannamei. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Work:2021:VLP

[2953] Thierry M. Work, Tina M. Weatherby, Jan H. Landsberg, Yasunari Kiryu, Samantha M. Cook, and Esther C. Peters. Viral-like particles are associated with endosymbiont pathology in Florida corals affected by Stony Coral Tissue Loss Disease. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Oliveira:2021:USP

[2954] Tiago C. A. Oliveira, Ying-Tsong Lin, and Michael B. Porter. Underwater sound propagation modeling in a complex shallow water environment. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Cooper:2021:TVT

[2955] Joanna K. Cooper, Andrew R. Gorman, M. Hamish Bowman, and Robert O. Smith. Temporal variability of thermohaline fine-structure associated with the subtropical front off the southeast coast of New Zealand in high-frequency short-streamer multi-channel seismic data. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lyu:2021:CSF

[2956] Mingxin Lyu, Xiaolong Gao, Mo Zhang, Shihui Lin, Xuan Luo, Weiwei You, and Caihuan Ke. Comparative study of the feeding characteristics and digestion physiology of Haliotis discus hannai and haliotis gigantea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kundu:2021:PFF

[2957] Goutam Kumar Kundu, Changseong Kim, Dongyoung Kim, Riaz Bibi, Heeyong Kim, and Chang-Keun Kang. Phytoplankton fuel fish food webs in a low-turbidity temperate coastal embayment: a stable isotope approach. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lin:2021:PRF

[2958] Bai an Lin, Yan Jiang, Robert Boenish, Qing Xu, and Min Liu. Population, reproductive and fishery dynamics of spotted box crab (Calappa philargius), a new claw-only fishery species, in the Southern Taiwan Strait, China. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

McMahon:2021:ABO

[2959] Clive R. McMahon, Fabien Roquet, Sophie Baudel, Mathieu Belbeoch, Sophie Bestley, Clint Blight, Lars Boehme, Fiona Carse, Daniel P. Costa, Michael A. Fedak, Christophe Guinet, Robert Harcourt, Emma Heslop, Mark A. Hindell, Xavier Hoenner, Kim Holland, Mellinda Holland, Fabrice R. A. Jaine, Tiphaine Jeanniard du Dot, Ian Jonsen, Theresa R. Keates, Kit M. Kovacs, Sara Labrousse, Philip Lovell, Christian Lydersen, David March, Matthew Mazloff, Megan K. McKinzie, Mônica M. C. Muelbert, Kevin O’Brien, Lachlan Phillips, Esther Portela, Jonathan Pye, Stephen Rintoul, Katsufumi Sato, Ana M. M. Sequeira, Samantha E. Simmons, Vardis M. Tsontos, Victor Turpin, Esmee van Wijk, Danny Vo, Mia Wege, Frederick Gilbert Whoriskey, Kenady Wilson, and Bill Woodward. Animal borne ocean sensors — AniBOS — an essential component of the Global Ocean Observing System. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ikuta:2021:IDP

[2960] Tetsuro Ikuta, Ryota Nakajima, Masashi Tsuchiya, Sanae Chiba, and Katsunori Fujikura. Interdecadal distribution of persistent organic pollutants in deep-sea chemosynthetic bivalves. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Liu:2021:CTP

[2961] Yuan Liu, Zhimei Zhu, Juan Qi, Zhenghong Sui, Erlei Shang, Shu Zhang, and Zeeshan Niaz. Comparative transcriptome profiling reveals insights into the mechanisms related to explosive growth of Alexandrium pacificum. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Nardi:2021:IIM

[2962] Alessandro Nardi, Maura Benedetti, Stefania Gorbi, and Francesco Regoli. Interactive immunomodulation in the Mediterranean mussel Mytilus galloprovincialis under thermal stress and cadmium exposure. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Merkens:2021:CLT

[2963] Karlina Merkens, Simone Baumann-Pickering, Morgan A. Ziegenhorn, Jennifer S. Trickey, Ann N. Allen, and Erin M. Oleson. Characterizing the long-term, wide-band and deep-water soundscape off Hawai’i. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Moore:2021:EBM

[2964] Jeffrey E. Moore, Dennis Heinemann, Tessa B. Francis, Philip S. Hammond, Kristy J. Long, André E. Punt, Randall R. Reeves, Maritza Sepúlveda, Gudjón Már Sigurdsson, Margaret C. Siple, Gísli A. Víkingsson, Paul R. Wade, Rob Williams, and Alexandre N. Zerbini. Estimating bycatch mortality for marine mammals: Concepts and best practices. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Leung:2021:EFM

[2965] Jonathan Y. S. Leung, Ben P. Harvey, and Bayden D. Russell. Editorial: Fitness of marine calcifiers in the future acidifying ocean. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Sulanke:2021:CDQ

[2966] Erik Sulanke and Sandra Rybicki. Community development quotas and support of small-scale fisheries as two key concepts for blue growth in fisheries. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lee:2021:DMR

[2967] Bo Yee Lee, Kaori Wakabayashi, Simon Yung Wa Sin, Susumu Ohtsuka, and Ling Ming Tsang. DNA metabarcoding revealed interspecific dietary difference and prey selectivity in juvenile horseshoe crabs Carcinoscorpius rotundicauda and Tachypleus tridentatus from Hong Kong. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Al-Otaibi:2021:PDV

[2968] Najwa Al-Otaibi, Francisca C. García, and Xosé Anxelu G. Morán. Picoplankton diel variability and estimated growth rates in epipelagic and mesopelagic waters of the Central Red Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Trefry:2021:ERC

[2969] John H. Trefry and Austin L. Fox. Extreme runoff of chemical species of nitrogen and phosphorus threatens a Florida barrier island lagoon. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lawley:2021:RNP

[2970] Jonathan W. Lawley, Anthony R. Carroll, and Carmel McDougall. Rhizostomins: a novel pigment family from rhizostome jellyfish (Cnidaria, Scyphozoa). Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Heo:2021:CGT

[2971] Young Mok Heo, Seung-Yoon Oh, Kyeongwon Kim, Sang-Il Han, Sun Lul Kwon, Yeonjae Yoo, Dongjun Kim, Jong Seong Khim, Seunghyun Kang, Hanbyul Lee, and Jae-Jin Kim. Comparative genomics and transcriptomics depict marine algicolous Arthrinium species as endosymbionts that help regulate oxidative stress in brown algae. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bundy:2021:EMF

[2972] Alida Bundy, Paul E. Renaud, Marta Coll, Stefan Koenigstein, Susa Niiranen, Maria Grazia Pennino, Jamie C. Tam, and Morgane Travers-Trolet. Editorial: Managing for the future: Challenges and approaches for disentangling the relative roles of environmental change and fishing in marine ecosystems. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Grearson:2021:LPS

[2973] Anik G. Grearson, Alison Dugan, Taylor Sakmar, Dominic M. Sivitilli, David H. Gire, Roy L. Caldwell, Cristopher M. Niell, Gül Dölen, Z. Yan Wang, and Bret Grasse. The lesser Pacific striped octopus, Octopus chierchiae: an emerging laboratory model. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Bolanos-Jimenez:2021:PDR

[2974] Jaime Bolaños-Jiménez, Eduardo Morteo, Christian A. Delfín-Alfonso, Pedro F. Fruet, Eduardo R. Secchi, and Javier Bello-Pineda. Population dynamics reveal a core community of the common bottlenose dolphin (Tursiops truncatus) in open waters of the South-Western Gulf of Mexico. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Cavanaugh:2021:ROC

[2975] Kyle C. Cavanaugh, Tom Bell, Maycira Costa, Norah E. Eddy, Lianna Gendall, Mary G. Gleason, Margot Hessing-Lewis, Rebecca Martone, Meredith McPherson, Ondine Pontier, Luba Reshitnyk, Rodrigo Beas-Luna, Mark Carr, Jennifer E. Caselle, Katherine C. Cavanaugh, Rebecca Flores Miller, Sara Hamilton, Walter N. Heady, Heidi K. Hirsh, Rietta Hohman, Lynn Chi Lee, Julio Lorda, James Ray, Daniel C. Reed, Vienna R. Saccomanno, and Sarah B. Schroeder. A review of the opportunities and challenges for using remote sensing for management of surface-canopy forming kelps. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Xue:2021:PTR

[2976] Dong-Xiu Xue, Hai-Yan Wang, and Tao Zhang. Phylogeography and taxonomic revision of the pen shell Atrina pectinata species complex in the South China Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Nian:2021:IPM

[2977] Rui Nian, Yu Cai, Zhengguang Zhang, Hui He, Jingyu Wu, Qiang Yuan, Xue Geng, Yuqi Qian, Hua Yang, and Bo He. The identification and prediction of mesoscale eddy variation via memory in memory with scheduled sampling for sea level anomaly. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Li:2021:DBW

[2978] Xiaopeng Li, Zehan Dai, Pengfei Di, Junxi Feng, Jun Tao, Duofu Chen, Niu Li, and Yazi Li. Distinct bottom-water bacterial communities at methane seeps with various seepage intensities in Haima, South China Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lee:2021:DPN

[2979] Ha-Hyun Lee, Sunggyu Lee, Jung Suk Lee, and Hyo-Bang Moon. Distribution of polychlorinated naphthalenes in sediment from industrialized coastal waters of Korea with the optimized cleanup and GC-MS/MS methods. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:CEC

[2980] Huan Wang, Meng Wang, Shen Fan, Jie Lu, Yi Lan, Menggong Li, Jiangyan Li, Rulong Liu, Jin Sun, Jiasong Fang, Pei-Yuan Qian, Yu-Zhong Zhang, and Weipeng Zhang. Culture enrichment combined with long-read sequencing facilitates genomic understanding of hadal sediment microbes. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Plonus:2021:ASP

[2981] Rene-Marcel Plonus, Stefanie Vogl, and Jens Floeter. Automatic segregation of pelagic habitats. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Han:2021:EBH

[2982] Mingming Han, Yuxin Liu, Tongqing Zhang, Daming Li, Qichen Jiang, and Zakaria Zuraini. Effect of berberine hydrochloride treatment on transcriptomic response in gill tissue of Charybdis japonica infected with Aeromonas hydrophila. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Geissler:2021:HVN

[2983] Laura Geissler, Valentine Meunier, Nils Rädecker, Gabriela Perna, Riccardo Rodolfo-Metalpa, Fanny Houlbrèque, and Christian R. Voolstra. Highly variable and non-complex diazotroph communities in corals from ambient and high CO_{2} environments. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Mintzer:2021:SWT

[2984] Vanessa J. Mintzer and Kristi L. Fazioli. Salinity and water temperature as predictors of bottlenose dolphin (Tursiops truncatus) encounter rates in Upper Galveston Bay, Texas. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ballance:2021:HTD

[2985] Lisa T. Ballance, Tim Gerrodette, Cleridy E. Lennert-Cody, Robert L. Pitman, and Dale Squires. A history of the tuna-dolphin problem: Successes, failures, and lessons learned. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Duong:2021:CCI

[2986] Trang Minh Duong. Climate change induced coastline change adjacent to small tidal inlets. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Shang:2021:EMB

[2987] Yueyong Shang, Xinghuo Wang, Xueqing Chang, Inna M. Sokolova, Shuaishuai Wei, Wei Liu, James K. H. Fang, Menghong Hu, Wei Huang, and Youji Wang. The effect of microplastics on the bioenergetics of the mussel Mytilus coruscus assessed by cellular energy allocation approach. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Wang:2021:IIS

[2988] Yaxin Wang, Yi Gong, Jian Zhang, Yi Tang, Xiaofei Shi, and Jiangao Shi. Intra- and inter-specific variation in edible jellyfish biomarkers and implications for origin traceability and authentication. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sakaue:2021:NIA

[2989] Jiro Sakaue, Kazuhiko Maeda, Michael J. Miller, Ryuichi Sakai, Koh ichi Tahara, Hideki Abe, Kazuya Maeda, and Hitoshi Ida. New insights about the behavioral ecology of the coelacanth Latimeria chalumnae video recorded in the absence of humans off South Africa. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Bruce:2021:IHE

[2990] Peter Bruce, Clare Bradshaw, Yvonne Ohlsson, Anna Sobek, and Anna Christiernsson. Inconsistencies in how environmental risk is evaluated in Sweden for dumping dredged sediment at sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bamunawala:2021:IPM

[2991] Janaka Bamunawala, Ad van der Spek, Ali Dastgheib, A. Brad Murray, and Roshanka Ranasinghe. An integrated, probabilistic modeling approach to assess the evolution of barrier-island systems over the Twenty-First Century. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Tapia-Guerra:2021:FEC

[2992] Jan M. Tapia-Guerra, Cynthia M. Asorey, Erin E. Easton, Daniel Wagner, Matthias Gorny, and Javier Sellanes. First ecological characterization of whip black coral assemblages (Hexacorallia: Antipatharia) in the Easter Island Ecoregion, Southeastern Pacific. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Hauser-Davis:2021:EEC

[2993] Rachel Ann Hauser-Davis and Salvatore Siciliano. Editorial: Emerging concerns, biology and conservation in key marine environments undergoing highly selective pressures. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Gonzalez:2021:EWE

[2994] Juliana A. Gonzalez, Martin A. Coombes, M. Gabriela Palomo, Federico I. Isla, Sabrina A. Soria, and Jorge L. Gutiérrez. Enhanced weathering and erosion of a cohesive shore platform following the experimental removal of mussels. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Miatta:2021:SOM

[2995] Marta Miatta and Paul V. R. Snelgrove. Sedimentary organic matter shapes macrofaunal communities but not benthic nutrient fluxes in contrasting habitats along the Northwest Atlantic continental margin. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhou:2021:FET

[2996] Fanyu Zhou, Junya Hirai, Koji Hamasaki, Sachiko Horii, and Atsushi Tsuda. Feeding ecology of three euphausiid species in the North Pacific Ocean inferred from 18S V9 metabarcoding and stable isotope analysis. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Tilot:2021:COS

[2997] Virginie C. Tilot, Bleuenn Guilloux, Klaas Willaert, Clement Y. Mulalap, Tamatoa Bambridge, Paul D’Arcy, Alexander Mawyer, François Gaulme, Edwige Kacenelenbogen, Alain Jeudy de Grissac, Juan Moreno Navas, and Arthur Dahl. The concept of oceanian sovereignty in the context of deep sea mining in the Pacific region. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Seo:2021:TAI

[2998] Hojong Seo, Guebuem Kim, Young-Il Kim, and Intae Kim. Tracing the atmospheric input of seawater-dissolvable Pb based on the budget of ^{210}Pb in the East Sea (Japan Sea). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Halasan:2021:FEC

[2999] Lorenzo C. Halasan, Paul John L. Geraldino, and Hsiu-Chin Lin. First evidence of cryptic species diversity and population structuring of Selaroides leptolepis in the tropical Western Pacific. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Hunt:2021:SCP

[3000] Corallie A. Hunt, Urška Demšar, Ben Marchant, Dayton Dove, and William E. N. Austin. Sounding out the carbon: The potential of acoustic backscatter data to yield improved spatial predictions of organic carbon in marine sediments. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Van:2021:BFN

[3001] Cong Mai Van, Anh Ngo, Tri Mai, and H. T. Dao. Bamboo fences as a nature-based measure for coastal wetland protection in Vietnam. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Richards:2021:TRN

[3002] Christina L. Richards, Kristen L. Langanke, Jeannie Mounger, Gordon A. Fox, and David B. Lewis. Trait response to nitrogen and salinity in Rhizophora mangle propagules and variation by maternal family and population of origin. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Dong:2021:FTR

[3003] Jian-Yu Dong, Linlin Zhao, Xiaolong Yang, Xin Sun, and Xiumei Zhang. Functional trait responses of macrobenthos to anthropogenic pressure in three temperate intertidal communities. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhou:2021:MEI

[3004] Jiang Zhou, Guidi Zhou, Hailong Liu, Zhuhua Li, and Xuhua Cheng. Mesoscale eddy-induced ocean dynamic and thermodynamic anomalies in the North Pacific. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wade:2021:BPA

[3005] Paul R. Wade, Kristy J. Long, Tessa B. Francis, André E. Punt, Philip S. Hammond, Dennis Heinemann, Jeffrey E. Moore, Randall R. Reeves, Maritza Sepúlveda, Genoa Sullaway, Gudjón Már Sigurdsson, Margaret C. Siple, Gísli A. Víkingsson, Rob Williams, and Alexandre N. Zerbini. Best practices for assessing and managing bycatch of marine mammals. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Liu:2021:SVP

[3006] Haijiao Liu, Yuyao Song, Xiaodong Zhang, Guicheng Zhang, Chao Wu, Xingzhou Wang, Satheeswaran Thangaraj, Dongxiao Wang, Ju Chen, and Jun Sun. Spatial variation in primary production in the Eastern Indian Ocean. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Ladd:2021:SRP

[3007] Cai John Tomos Ladd, Mollie Frances Duggan-Edwards, Jordi F. Pagès, and Martin Wiggers Skov. Saltmarsh resilience to periodic shifts in tidal channels. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Harford:2021:MIH

[3008] William J. Harford, Ricardo Amoroso, Richard J. Bell, Matias Caillaux, Jason Marc Cope, Dawn Dougherty, Natalie Anne Dowling, Frank Hurd, Serena Lomonico, Josh Nowlis, Dan Ovando, Ana M. Parma, Jeremy D. Prince, and Jono R. Wilson. Multi-indicator harvest strategies for datalimited fisheries: a practitioner guide to learning and design. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Jiang:2021:SCA

[3009] Yan Jiang, Bai an Lin, Hao yang He, Guang mao Ding, Li ting Yan, Ge Zhang, Min Liu, and Lian ming Zheng. Species composition and assemblages of ichthyoplankton in Sansha Bay, Fujian Province, China. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Shinzato:2021:NMD

[3010] Chuya Shinzato, Haruhi Narisoko, Koki Nishitsuji, Tomofumi Nagata, Noriyuki Satoh, and Jun Inoue. Novel mitochondrial DNA markers for scleractinian corals and generic-level environmental DNA metabarcoding. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Al-Mansi:2021:STP

[3011] Ahmed M. Al-Mansi, Anas Z. Sambas, Baleegh A. Abukaboos, Ahmed H. Al Zahrani, Ahmed S. Abdulaziz, Ali A. Almasabi, Raed S. Alkreda, and Jeff Miller. Satellite tracking of post-nesting green sea turtles (Chelonia mydas) from Ras Baridi, Red Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Gao:2021:IMH

[3012] Guang Gao, Xin Zhao, Meijia Jiang, and Lin Gao. Impacts of marine heatwaves on algal structure and carbon sequestration in conjunction with ocean warming and acidification. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Samuel:2021:TGP

[3013] Robyn M. Samuel, Raissa Meyer, Pier Luigi Buttigieg, Neil Davies, Nicholas W. Jeffery, Christopher Meyer, Christina Pavloudi, Kathleen Johnson Pitz, Maxime Sweetlove, Susanna Theroux, Jodie van de Kamp, and Alison Watts. Toward a global public repository of community protocols to encourage best practices in biomolecular ocean observing and research. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Ahammad:2021:EFG

[3014] A. K. Shakur Ahammad, Neaz A. Hasan, Mohammad Mahfujul Haque, Abul Bashar, Md. Borhan Uddin Ahmed, Mohammad Ashraful Alam, Md. Asaduzzaman, Md. Abul Bashar, and Yahia Mahmud. Environmental factors and genetic diversity as drivers of early gonadal maturation: a gonadosomatic index based investigation on Indian shad, Tenualosa ilisha population of Bangladesh. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Kim:2021:VAE

[3015] Tae-Kon Kim, Changbin Lim, and Jung-Lyul Lee. Vulnerability analysis of episodic beach erosion by applying storm wave scenarios to a shoreline response model. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Hopkins:2021:MBS

[3016] Grant Hopkins, Ian Davidson, Eugene Georgiades, Oliver Floerl, Donald Morrisey, and Patrick Cahill. Managing biofouling on submerged static artificial structures in the marine environment — assessment of current and emerging approaches. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fang:2021:ISC

[3017] Di an Fang, Xiang ping Xue, Dong po Xu, Xiao hao Wang, Ning ze Sun, Bai dong Zhang, and Peng Ren. Ichthyoplankton species composition and assemblages from the estuary to the Hukou section of the Changjiang River. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Thorburn:2021:CSO

[3018] James Thorburn, Peter J. Wright, Edward Lavender, Jane Dodd, Francis Neat, Julien G. A. Martin, Caroline Lynam, and Mark James. Corrigendum: Seasonal and ontogenetic variation in depth use by a critically endangered benthic elasmobranch and its implications for spatial management. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745. See [1980].

Wesselmann:2021:WTP

[3019] Marlene Wesselmann, Rosa M. Chefaoui, Núria Marbà, Ester A. Serrao, and Carlos M. Duarte. Warming threatens to propel the expansion of the exotic seagrass Halophila stipulacea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:MMR

[3020] Ruixuan Wang, Hongli Mou, Xiaozhi Lin, Hui Zhu, Bing Li, Jiangyong Wang, Muhammad Junaid, and Jun Wang. Microplastics in mollusks: Research progress, current contamination status, analysis approaches, and future perspectives. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Qin:2021:CTP

[3021] Zhen Qin, Qianqian Ge, Jiajia Wang, Mingdong Li, Ping Liu, Jian Li, and Jitao Li. Comparative transcriptomic and proteomic analysis of Exopalaemon carinicauda in response to alkalinity stress. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zurowietz:2021:CTF

[3022] Martin Zurowietz and Tim W. Nattkemper. Current trends and future directions of large scale image and video annotation: Observations from four years of BIIGLE 2.0. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Merten:2021:IAC

[3023] Véronique Merten, Till Bayer, Thorsten B. H. Reusch, Oscar Puebla, Janina Fuss, Julia Stefanschitz, Alexandra Lischka, Helena Hauss, Philipp Neitzel, Uwe Piatkowski, Stephanie Czudaj, Bernd Christiansen, Anneke Denda, and Henk-Jan T. Hoving. An integrative assessment combining deep-sea net sampling, in situ observations and environmental DNA analysis identifies Cabo Verde as a cephalopod biodiversity hotspot in the Atlantic Ocean. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ge:2021:CWD

[3024] Ruping Ge, Hongju Chen, Ling Li, Guangxing Liu, Bangqin Huang, Lisha Wang, Yunyun Zhuang, and Xin Liu. Controls of wind-driven jet on the distribution of pelagic cnidarians in the Midwestern South China Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Borges:2021:CKS

[3025] Rebecca Borges, Indira A. L. Eyzaguirre, Roberta Sá Leitão Barboza, Karin Boos, Marion Glaser, and Priscila F. M. Lopes. Combining knowledge systems helps understand the spatial dynamics of data-limited small-scale fisheries systems in Brazil: a methods analysis. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Yuan:2021:IPC

[3026] Bin Yuan, Fangzhou Yue, Xueye Wang, and Hanxiao Xu. The impact of pollution on China marine fishery culture: an econometric analysis of heterogeneous growth. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Rahimnejad:2021:EVC

[3027] Samad Rahimnejad, Konrad Dabrowski, Marisol Izquierdo, Nima Hematyar, Aiman Imentai, Christoph Steinbach, and Tomas Policar. Effects of vitamin C and E supplementation on growth, fatty acid composition, innate immunity, and antioxidant capacity of rainbow trout (Oncorhynchus mykiss) fed oxidized fish oil. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Barbariol:2021:WWM

[3028] Francesco Barbariol, Silvio Davison, Francesco Marcello Falcieri, Rossella Ferretti, Antonio Ricchi, Mauro Sclavo, and Alvise Benetazzo. Wind waves in the Mediterranean Sea: an ERA5 reanalysis wind-based climatology. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Falace:2021:SMC

[3029] Annalisa Falace, Giuliana Marletta, Gilda Savonitto, Fabio Candotto Carniel, Marina Srijemsi, Stanislao Bevilacqua, Mauro Tretiach, and Giuseppina Alongi. Is the South-Mediterranean canopy-forming Ericaria giacconei (= Cystoseira hyblaea) a loser from ocean warming? Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Shamskhany:2021:EMS

[3030] Arefeh Shamskhany, Zhuoran Li, Preet Patel, and Shooka Karimpour. Evidence of microplastic size impact on mobility and transport in the marine environment: a review and synthesis of recent research. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Agulles:2021:FSB

[3031] Miguel Agulles, Gabriel Jordà, and Piero Lionello. Flooding of sandy beaches in a changing climate. The case of the Balearic Islands (NW Mediterranean). Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Murphy:2021:GPS

[3032] Eugene J. Murphy, Carol Robinson, Alistair J. Hobday, Alice Newton, Marion Glaser, Karen Evans, Mark Dickey-Collas, Stephanie Brodie, and Marion Gehlen. The global pandemic has shown we need an action plan for the ocean. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

VanParijs:2021:NBM

[3033] Sofie M. Van Parijs, Kyle Baker, Jordan Carduner, Jaclyn Daly, Genevieve E. Davis, Carter Esch, Shane Guan, Amy Scholik-Schlomer, Nicholas B. Sisson, and Erica Staaterman. NOAA and BOEM minimum recommendations for use of passive acoustic listening systems in offshore wind energy development monitoring and mitigation programs. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Naugle:2021:VCT

[3034] Melissa S. Naugle, Thomas A. Oliver, Daniel J. Barshis, Ruth D. Gates, and Cheryl A. Logan. Variation in coral thermotolerance across a pollution gradient erodes as coral symbionts shift to more heat-tolerant genera. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ma:2021:PIF

[3035] Mingyu Ma, Mingyu Zhong, Quansheng Zhang, Wei Zhao, Mengxin Wang, and Chengying Luo. Phylogenetic implications and functional disparity in the Chalcone synthase gene family of common seagrass Zostera marina. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:IEA

[3036] Haozhuang Wang, Jürgen Titschack, Claudia Wienberg, Chelsea Korpanty, and Dierk Hebbeln. The importance of ecological accommodation space and sediment supply for cold-water coral mound formation, a case study from the Western Mediterranean Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Kang:2021:CTA

[3037] Hee Chang Kang, Hae Jin Jeong, Sang Ah Park, Jin Hee Ok, Ji Hyun You, Se Hee Eom, Eun Chong Park, Se Hyeon Jang, and Sung Yeon Lee. Comparative transcriptome analysis of the phototrophic dinoflagellate Biecheleriopsis adriatica grown under optimal temperature and cold and heat stress. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhao:2021:ITR

[3038] Jichen Zhao, Minze Liao, Zexu Lin, Yiyi Huang, Yunqi Zhong, Luyao Liu, Guoliang Chen, Zuotao Ni, and Chengbo Sun. Isobaric tags for relative and absolute quantitation-based proteomics analysis provides a new perspective into unsynchronized growth in kuruma shrimp (Marsupenaeus japonicus). Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Aracri:2021:PSB

[3039] Simona Aracri, Marco Contardi, Ilker S. Bayer, Muhammad Zahid, Francesco Giorgio-Serchi, and Adam A. Stokes. Propaedeutic study of biocomposites obtained with natural fibers for oceanographic observing platforms. Frontiers in Marine Science, 8, December 2021. ISSN 2296-

Zeng:2021:SCT

[3040] Lin Zeng, Chao Zhan, Qing Wang, Xianbin Liu, Longsheng Wang, Xueyan Li, Xin Wang, Xiang Yu, Jinzhi Zhang, and Buli Cui. Sediment coarsening in tidal flats and stable coastline of the abandoned Southern Yellow River sub-delta in response to fluvial sediment flux decrease during the past decades. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Davidson:2021:RBS

[3041] Ian Davidson, Patrick Cahill, Arne Hinz, Daniel Kluza, Chris Scianni, and Eugene Georgiades. A review of biofouling of ships’ internal seawater systems. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Berry:2021:UFD

[3042] David S. Berry. Unity or fragmentation in the deep blue: Choices in institutional design for marine biological diversity in areas beyond national jurisdiction. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zhou:2021:TMH

[3043] Qian Zhou, Chen Tu, Jie Yang, Chuancheng Fu, Yuan Li, and Joanna J. Waniek. Trapping of microplastics in halocline and turbidity layers of the semi-enclosed Baltic Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Li:2021:PRE

[3044] Xi Li, Shiguo Li, Jiawei Cheng, Ruiying Fu, and Aibin Zhan. Proteomic response to environmental stresses in the stolon of a highly invasive fouling ascidian. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gonzalez-Aravena:2021:LTP

[3045] Marcelo González-Aravena, Rodolfo Rondon, Alejandro Font, César A. Cárdenas, Jean-Yves Toullec, Erwan Corre, and Kurt Paschke. Low transcriptomic plasticity of Antarctic giant isopod Glyptonotus antarcticus juveniles exposed to acute thermal stress. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Arabshahi:2021:UVA

[3046] Homayon John Arabshahi, Tomaž Trobec, Valentin Foulon, Claire Hellio, Robert Frangež, Kristina Sepčić, Patrick Cahill, and Johan Svenson. Using virtual AChE homology screening to identify small molecules with the ability to inhibit marine biofouling. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Banchi:2021:SFH

[3047] Elisa Banchi, Paola Del Negro, Mauro Celussi, and Francesca Malfatti. Sediment features and human activities structure the surface microbial communities of the Venice Lagoon. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zamoner:2021:IOD

[3048] Julia Biscaia Zamoner, Anaide Wrublevski Aued, Luis Carlos Pinto Macedo-Soares, Vitor André Passos Picolotto, Carlos Alberto Eiras Garcia, and Bárbara Segal. Integrating oceanographic data and benthic community structure temporal series to assess the dynamics of a marginal reef. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Gao:2021:IMM

[3049] Ning Gao, Zhihui Huang, Jianing Xing, Siyi Zhang, and Jing Hou. Impact and molecular mechanism of microplastics on zebrafish in the presence and absence of copper nanoparticles. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Simard:2021:PPC

[3050] Nittya S. M. Simard, Thane A. Militz, Jeff Kinch, and Paul C. Southgate. From past to present: Construction of a dataset documenting mother-of-pearl exports from a Pacific island nation, Papua New Guinea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Oliveira:2021:FRB

[3051] Anderson G. Oliveira, Danilo T. Amaral, Mary Colleen Hannon, and Anja Schulze. First record of bioluminescence in a sipunculan worm. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Jeong:2021:IEL

[3052] Jaewoon Jeong, Marit Stormoen, Krishna K. Thakur, and Crawford W. Revie. Imperfect estimation of Lepeophtheirus salmonis abundance and its impact on salmon lice treatment on Atlantic salmon farms. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zhang:2021:EFV

[3053] Mo Zhang, Xiaolong Gao, Mingxin Lyu, Shihui Lin, Ying Su, Xuan Luo, Weiwei You, and Caihuan Ke. Effects of flow velocity on the growth and survival of Haliotis discus hannai larvae in the recirculating upflow system from the point of energy metabolism. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Naimullah:2021:ENN

[3054] Muhamad Naimullah, Yan-Lun Wu, Ming-An Lee, and Kuo-Wei Lan. Effect of the El Niño–Southern Oscillation (ENSO) cycle on the catches and habitat patterns of three swimming crabs in the Taiwan Strait. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

San:2021:TAG

[3055] Lize San, Baosuo Liu, Bo Liu, Huayang Guo, Liang Guo, Nan Zhang, Kecheng Zhu, Shigui Jiang, and Dianchang Zhang. Transcriptome analysis of gills provides insights into translation changes under hypoxic stress and reoxygenation in golden pompano, Trachinotus ovatus (Linnaeus 1758). Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Kalananthan:2021:BDC

[3056] Tharmini Kalananthan, Ana S. Gomes, Floriana Lai, Ingvill Tolås, Ann-Elise O. Jordal, Sissel Norland, Koji Murashita, and Ivar Rønnestad. Brain distribution of 10 cart transcripts and their response to 4 days of fasting in Atlantic salmon (Salmo salar L.). Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

He:2021:MMC

[3057] Ziying He, Huaye Sun, Xiaoli Yu, Zhushi Yin, Mengxing Wu, Lili Zhao, Zhan Hu, Yisheng Peng, and Shing Yip Lee. Monoculture or mixed culture? Relevance of fine root dynamics to carbon sequestration oriented mangrove afforestation and restoration. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Bai:2021:MCH

[3058] Shijie Bai, Hengchao Xu, and Xiaotong Peng. Microbial communities of the hydrothermal scaly-foot snails from kairei and longqi vent fields. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Silva:2021:SAA

[3059] Marcos V. B. Silva and Tereza C. M. Araújo. The South American advances in submarine canyons studies and their link to the ocean decade. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Behrenfeld:2021:EUM

[3060] Michael J. Behrenfeld, Sarah D. Brooks, Peter Gaube, and Kristina D. A. Mojica. Editorial: Unraveling mechanisms underlying annual plankton blooms in the North Atlantic and their implications for biogenic aerosol properties and cloud formation. Frontiers in Marine Science, 8, September 2021. ISSN 2296-7745.

Jintasataporn:2021:ESC

[3061] Orapint Jintasataporn, Srinoy Chumkam, Supawit Triwutanon, Allan LeBlanc, and Jarin Sawanboonchun. Effects of a single cell protein (Methylococcus capsulatus, Bath) in Pacific white shrimp (Penaeus vannamei) diet on growth performance, survival rate and resistance to Vibrio parahaemolyticus, the causative agent of acute hepatopancreatic necrosis disease. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Cote:2021:FST

[3062] David Cote, Cassandra A. Konecny, Jennica Seiden, Tristan Hauser, Trond Kristiansen, and Ben J. Laurel. Forecasted shifts in thermal habitat for cod species in the Northwest Atlantic and Eastern Canadian Arctic. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Giebel:2021:MGO

[3063] Helge-Ansgar Giebel, Carol Arnosti, Thomas H. Badewien, Insa Bakenhus, John Paul Balmonte, Sara Billerbeck, Leon Dlugosch, Rohan Henkel, Birgit Kuerzel, Jens Meyerjürgens, Felix Milke, Daniela Voss, Gerrit Wienhausen, Matthias Wietz, Holger Winkler, Mathias Wolterink, and Meinhard Simon. Microbial growth and organic matter cycling in the Pacific Ocean along a latitudinal transect between subarctic and subantarctic waters. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhang:2021:CIC

[3064] Liyan Zhang, Jing Zhang, Shigang Liu, Rui Wang, Jiali Xiang, Xing Miao, Ran Zhang, Puqing Song, and Longshan Lin. Characteristics of ichthyoplankton communities and their relationship with environmental factors above the Ninety East Ridge, Eastern Indian Ocean. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Lu:2021:EDI

[3065] Qisheng Lu, Longwei Xi, Yulong Liu, Yulong Gong, Jingzhi Su, Dong Han, Yunxia Yang, Junyan Jin, Haokun Liu, Xiaoming Zhu, and Shouqi Xie. Effects of dietary inclusion of Clostridium autoethanogenum protein on the growth performance and liver health of largemouth bass (Micropterus salmoides). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Assis:2021:PBC

[3066] Jorge Assis, Pierre Failler, Eliza Fragkopoulou, David Abecasis, Gregoire Touron-Gardic, Aissa Regalla, Ebaye Sidina, Herculano Dinis, and Ester A. Serrao. Potential biodiversity connectivity in the network of Marine Protected Areas in Western Africa. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Morais:2021:HPS

[3067] Pedro Morais, Luís Afonso, and Ester Dias. Harnessing the power of social media to obtain biodiversity data about cetaceans in a poorly monitored area. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

dErrico:2021:AMW

[3068] Giuseppe d’Errico, Alessandro Nardi, Maura Benedetti, Marica Mezzelani, Daniele Fattorini, Marta Di Carlo, Lucia Pittura, Maria Elisa Giuliani, Simona Macchia, Valentina Vitiello, Davide Sartori, Alice Scuderi, Lorenzo Morroni, Gianluca Chiaretti, Stefania Gorbi, David Pellegrini, and Francesco Regoli. Application of a multidisciplinary weight of evidence approach as a tool for monitoring the ecological risk of dredging activities. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Goode:2021:FSM

[3069] Savannah L. Goode, Ashley A. Rowden, David A. Bowden, Malcolm R. Clark, and Fabrice Stephenson. Fine-scale mapping of megaepibenthic communities and their patch characteristics on two New Zealand seamounts. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hardin:2021:SRA

[3070] Emily E. Hardin and Mariana M. P. B. Fuentes. A systematic review of acoustic telemetry as a tool to gain insights into marine turtle ecology and aid their conservation. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Waiho:2021:SDL

[3071] Khor Waiho, Hanafiah Fazhan, Alexander Chong Shu-Chien, Muyassar H. Abualreesh, Hongyu Ma, Mohammad Syahnon, Ghazali Azmie, Nurul Jannah Razman, and Mhd Ikhwanuddin. Size distribution, length– weight relationship, and size at morphometric maturity of the mud spiny lobster Panulirus polyphagus (Herbst, 1793) in the Johor Strait. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ma:2021:CMM

[3072] Mingzhen Ma, Yuqing Li, Honggang Ma, Khaled A. S. Al-Rasheid, Alan Warren, Yurui Wang, and Ying Yan. Corrigendum: Morphology and molecular phylogeny of four trachelocercid ciliates (Protozoa, Ciliophora, Karyorelictea) found in marine coastal habitats of Northern China, with description of a new genus, two new species and a new combination. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745. See [1327].

Zhou:2021:IWD

[3073] Guidi Zhou, Zhuhua Li, and Xuhua Cheng. Intrinsic and wind-driven decadal variability of the Kuroshio Extension in altimeter observations. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Rey:2021:HFH

[3074] Wilmer Rey, Pablo Ruiz-Salcines, Paulo Salles, Claudia P. Urbano-Latorre, Germán Escobar-Olaya, Andrés F. Osorio, Juan Pablo Ramírez, Angélica Cabarcas-Mier, Bismarck Jigena-Antelo, and Christian M. Appendini. Hurricane flood hazard assessment for the Archipelago of San Andres, Providencia and Santa Catalina, Colombia. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Clark:2021:SCD

[3075] Kasey E. Clark, Angela Capper, Wei-Ting Liu, Amanda M. Fenner, Alejandro Almanza, Gina Della Togna, Liuris Herrera, Timothy Johns, Valerie J. Paul, Pieter C. Dorrestein, Todd L. Capson, and Marcy J. Balunas. Sequestration and cyanobacterial diet preferences in the opisthobranch molluscs Dolabrifera nicaraguana and Stylocheilus rickettsi. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:ASL

[3076] Verena H. Wang, Carley R. Zapfe, and Frank J. Hernandez. Assemblage structure of larval fishes in epipelagic and mesopelagic waters of the Northern Gulf of Mexico. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Han:2021:DLS

[3077] Qingpeng Han, Xiujuan Shan, Xianshi Jin, Harry Gorfine, Tao Yang, and Chengcheng Su. Data-limited stock assessment for fish species devoid of catch statistics: Case studies for Pampus argenteus and Setipinna taty in the Bohai and Yellow Seas. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:NSE

[3078] Dong Wang, Rong Wan, Zengguang Li, Junbo Zhang, Xiangyu Long, Pengbo Song, Lu Zhai, and Shuo Zhang. The non-stationary environmental effects on spawning habitat of fish in estuaries: a case study of Coilia mystus in the Yangtze Estuary. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhang:2021:INC

[3079] Han Zhang, Lei Wang, Xiang Yang, Zhiwei Lian, Yinbin Qiu, Zhanying Dong, Xuedong Wu, and Xinghua Pan. Identification of novel conopeptides and distinct gene superfamilies in the marine cone snail Conus quercinus. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Adyel:2021:WLM

[3080] Tanveer M. Adyel and Peter I. Macreadie. World’s largest mangrove forest becoming plastic cesspit. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Grasselli:2021:HWD

[3081] Ferrante Grasselli and Laura Airoldi. How and to what degree does physical structure differ between natural and artificial habitats? A multi-scale assessment in marine intertidal systems. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Cecino:2021:TIS

[3082] Giorgia Cecino, Roozbeh Valavi, and Eric A. Treml. Testing the influence of seascape connectivity on marine-based species distribution models. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Vikebo:2021:NAC

[3083] Frode B. Vikebø, Ole Jacob Broch, Clarissa Akemi Kajiya Endo, Håvard G. Frøysa, JoLynn Carroll, Jonas Juselius, and Øystein Langangen. Northeast Arctic cod and prey match-mismatch in a high-latitude spring-bloom system. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Fusi:2021:DOH

[3084] Marco Fusi, Daniele Daffonchio, Jenny Booth, and Folco Giomi. Dissolved oxygen in heterogeneous environments dictates the metabolic rate and thermal sensitivity of a tropical aquatic crab. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yucharoen:2021:EAV

[3085] Mathinee Yucharoen, Sutinee Sinutok, Ponlachart Chotikarn, and Pimchanok Buapet. Experimental assessment of vulnerability to warming in tropical shallow-water marine organisms. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sivadas:2021:PPA

[3086] Sanitha K. Sivadas, K. Ramu, Pravakar Mishra, and Ramana Murthy MV. Potential plastic accumulation zones in the Indian coastal seas. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Vitale:2021:CPR

[3087] Giulia Vitale, Arnau L. Dedeu, Marta Pujol, and Valerio Sbragaglia. Characterizing the profile of recreational fishers who share their catches on social media. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Tsikopoulou:2021:TVE

[3088] Irini Tsikopoulou, Panagiotis D. Dimitriou, Ioannis Karakassis, Nikolaos Lampadariou, Nadia Papadopoulou, and Christopher J. Smith. Temporal variation in the ecological functioning of benthic communities after 20 years in the Eastern Mediterranean. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Fuller:2021:BAD

[3089] Kate M. Fuller, Austin L. Fox, Charles A. Jacoby, and John H. Trefry. Biological abundance and diversity in organic-rich sediments from a Florida barrier island lagoon. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Miranda:2021:CPP

[3090] José Luis López Miranda, Lourdes B. Celis, Miriam Estévez, Valeria Chávez, Brigitta I. van Tussenbroek, Abigail Uribe-Martínez, Eduardo Cuevas, Izarelly Rosillo Pantoja, Luis Masia, Citlali Cauich-Kantun, and Rodolfo Silva. Commercial potential of pelagic Sargassum spp. in Mexico. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Graham:2021:ESW

[3091] Olivia J. Graham, Lillian R. Aoki, Tiffany Stephens, Joshua Stokes, Sukanya Dayal, Brendan Rappazzo, Carla P. Gomes, and C. Drew Harvell. Effects of seagrass wasting disease on eelgrass growth and belowground sugar in natural meadows. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yamaguchi:2021:RBE

[3092] Atsuko Yamaguchi, Keisuke Furumitsu, and Jennifer Wyffels. Reproductive biology and embryonic diapause as a survival strategy for the East Asian endemic eagle ray Aetobatus narutobiei. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Yamada:2021:GVP

[3093] Yosuke Yamada, Hideki Fukuda, Yu Umezawa, and Toshi Nagata. Geographic variation of particle size distribution in the Kuroshio region: Possible causes in the upper water column. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Patel:2021:BNI

[3094] Nirav Patel, Yosuke Yamada, and Farooq Azam. Bacterial nanotubes as intercellular linkages in marine assemblages. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Han:2021:TTE

[3095] MyeongHee Han, Yeon S. Chang, Hyoun-Woo Kang, Dong-Jin Kang, and Yong Sun Kim. Turnover time of the East Sea (Sea of Japan) meridional overturning circulation. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zhong:2021:IRT

[3096] Zehua Zhong and Zhan Hu. The impact of reclamation on tidal flat morphological equilibrium. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lopez:2021:HRS

[3097] Cary B. Lopez, Charles L. Tilney, Eric Muhlbach, Josée N. Bouchard, Maria Célia Villac, Karen L. Henschen, Laura R. Markley, Stephanie Keller Abbe, Sugandha Shankar, Colin P. Shea, Leanne Flewelling, Matthew Garrett, Susan Badylak, Edward J. Phlips, Lauren M. Hall, Margaret A. Lasi, Ashley A. Parks, Richard Paperno, Douglas H. Adams, Dwayne D. Edwards, Jacob E. Schneider, Kyle B. Wald, Autumn R. Biddle, Shawna L. Landers, and Katherine A. Hubbard. High-resolution spatiotemporal dynamics of harmful algae in the Indian River Lagoon (Florida) — a case study of Aureoumbra lagunensis, Pyrodinium bahamense, and Pseudo-nitzschia. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Martins:2021:ICE

[3098] Thais Martins, Paula Santana, Ítalo Lutz, Raimundo da Silva, Aurycéia Guimarães-Costa, Marcelo Vallinoto, Iracilda Sampaio, and Grazielle Evangelista-Gomes. Intensive commercialization of endangered sharks and rays (Elasmobranchii) along the coastal Amazon as revealed by DNA barcode. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Fu:2021:BPA

[3099] Yiqiu Fu, Zonghang Zhang, Zhen Zhang, Fengyuan Shen, Xiuwen Xu, Zheng Li, Yanhao Zhang, and Xiumei Zhang. Boldness predicts aggressiveness, metabolism, and activity in black rockfish Sebastes schlegelii. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Chen:2021:IAF

[3100] Peng Chen, Xinjun Chen, Wei Yu, and Dongming Lin. Interannual abundance fluctuations of two oceanic squids in the Pacific Ocean can be evaluated through their habitat temperature variabilities. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Gaos:2021:HNH

[3101] Alexander R. Gaos, Lauren Kurpita, Hannah Bernard, Luke Sundquist, Cheryl S. King, Joy H. Browning, Eldridge Naboa, Irene K. Kelly, Kelleigh Downs, Tomo Eguchi, George Balazs, Kyle Van Houtan, Devin Johnson, T. Todd Jones, and Summer L. Martin. Hawksbill nesting in Hawai’i: 30-year dataset reveals recent positive trend for a small, yet vital population. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Han:2021:ICI

[3102] Xiaoxin Han, Rolf D. Vogt, Jiaying Zhou, Boyang Zheng, Xue Yu, Jianfeng Feng, and Xueqiang Lu. Increased Cu(II) adsorption onto UV-aged polyethylene, polypropylene, and polyethylene terephthalate microplastic particles in seawater. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Zhang:2021:CGF

[3103] Xingzhi Zhang, Bingcong Ye, Zhifeng Gu, Meng Li, Shouguo Yang, Aimin Wang, and Chunsheng Liu. Comparison in growth, feeding, and metabolism between a fast-growing selective strain and a cultured population of pearl oyster (Pinctada fucata martensii). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Chen:2021:DCR

[3104] Lu-Chi Chen, Jinn-Shing Weng, Muhamad Naimullah, Po-Yuan Hsiao, Chen-Te Tseng, Kuo-Wei Lan, and Che-Chen Chuang. Distribution and catch rate characteristics of narrow-barred Spanish mackerel (Scomberomorus commerson) in relation to oceanographic factors in the waters around Taiwan. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Chen:2021:DNR

[3105] Shunyang Chen, Bingpeng Xing, Weiwei Yu, Bin Chen, Jianji Liao, Wenshuo An, and Guangcheng Chen. Distribution of a newly recorded gastropod species, Mainwaringia leithii (Gastropoda, Littorinidae), in young, rehabilitated mangroves in China. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Fischer:2021:EMD

[3106] Philipp Fischer, Peter Dietrich, Eric P. Achterberg, Norbert Anselm, Holger Brix, Ingeborg Bussmann, Laura Eickelmann, Götz Flöser, Madlen Friedrich, Hendrik Rust, Claudia Schütze, and Uta Koedel. Effects of measuring devices and sampling strategies on the interpretation of monitoring data for long-term trend analysis. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Luan:2021:MDT

[3107] Jing Luan, Chongliang Zhang, Yupeng Ji, Binduo Xu, Ying Xue, and Yiping Ren. Matching data types to the objectives of species distribution modeling: an evaluation with marine fish species. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Anderson:2021:MMP

[3108] Madeline P. B. C. Anderson, Phillip B. Fenberg, Huw J. Griffiths, and Katrin Linse. Macrobenthic Mollusca of the Prince Gustav Channel, Eastern Antarctic Peninsula: an area undergoing colonisation. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Scholz:2021:RCG

[3109] Vincent V. Scholz, Kasper E. Brodersen, Michael Kühl, and Klaus Koren. Resolving chemical gradients around seagrass roots — a review of available methods. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Chu:2021:SMA

[3110] Yu-Ting Chu and Paul B. Brown. Sustainable marine aquaponics: Effects of shrimp to plant ratios and C/N ratios. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Visudtiphole:2021:GLA

[3111] Virak Visudtiphole, Jutatip Khudet, Panomkorn Chaitongsakul, Siwat Plaisen, Jittima Siriwattano, Seansook Laiphrom, Anupong Klaysuban, Thaninthorn Raweeratanapong, Kanchana Sittikankaew, Nakul Rattanaphan, Looksorn Koichai, Panida Unagul, and Umaporn Uawisetwathana. Growth and lipidomic analyses of Penaeus monodon larvae supplemented with Aurantiochytrium limacinum BCC52274. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Nguyen-Duy:2021:VRR

[3112] Tung Nguyen-Duy, Nadia K. Ayoub, Patrick Marsaleix, Florence Toublanc, Pierre De Mey-Frémaux, Violaine Piton, Marine Herrmann, Thomas Duhaut, Manh Cuong Tran, and Thanh Ngo-Duc. Variability of the Red River plume in the Gulf of Tonkin as revealed by numerical modeling and clustering analysis. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Sanchez:2021:ESB

[3113] Juan Armando Sánchez, Sonia Bejarano, and Santiago Herrera. Editorial: SeaFlower Biosphere Reserve: New findings and trends in the largest Caribbean Marine Protected Area. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Gravina:2021:FRB

[3114] Maria Flavia Gravina, Edoardo Casoli, Luigia Donnarumma, Jacopo Giampaoletti, Federica Antonelli, Carlotta Sacco Perasso, and Sandra Ricci. First report on the benthic invertebrate community associated with a bronze naval ram from the First Punic War: a proxy of marine biodiversity. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Drivdal:2021:CDD

[3115] Magnus Drivdal, Erin H. Kunisch, Bodil A. Bluhm, Rolf Gradinger, Stig Falk-Petersen, and Jørgen Berge. Connections to the deep: Deep vertical migrations, an important part of the life cycle of apherusa glacialis, an Arctic ice-associated amphipod. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Johansen:2021:AVC

[3116] Ellie Johansen, Nicole Aberle, Mari-Ann Østensen, and Sanna Majaneva. Assessing the value of a citizen science approach for ctenophore identification. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Li:2021:MIT

[3117] Bingtian Li, Libin Du, Shiqiu Peng, Yibo Yuan, Xiangqian Meng, and Xianqing Lv. Modulation of internal tides by turbulent mixing in the South China Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Pineiro-Corbeira:2021:STN

[3118] Cristina Piñeiro-Corbeira, Laura Iglesias, Raquel Nogueira, Sara Campos, Arturo Jiménez, Marcos Regueira, Rodolfo Barreiro, and Miquel Planas. Structure and trophic niches in mobile epifauna assemblages associated with seaweeds and habitats of syngnathid fishes in Cíes Archipelago (Atlantic Islands Marine National Park, North West Iberia). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Asch:2021:EEA

[3119] Rebecca G. Asch, Johnna M. Holding, Darren J. Pilcher, Sara Rivero-Calle, and Kenneth A. Rose. Editorial: Ecological applications of earth system models and regional climate models. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Sayed:2021:LDR

[3120] Alaa El-Din H. Sayed, Kento Nagata, Takuya Nakazawa, Hiroshi Mitani, Junya Kobayashi, and Shoji Oda. Low dose-rate irradiation of gamma-rays-induced cytotoxic and genotoxic alterations in peripheral erythrocytes of p53-deficient medaka (Oryzias latipes). Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Sokolov:2021:IMU

[3121] Eugene P. Sokolov, Linda Adzigbli, Stephanie Markert, Amanda Bundgaard, Angela Fago, Dörte Becher, Claudia Hirschfeld, and Inna M. Sokolova. Intrinsic mechanisms underlying hypoxia-tolerant mitochondrial phenotype during hypoxia-reoxygenation stress in a marine facultative anaerobe, the blue mussel Mytilus edulis. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Feugere:2021:BSP

[3122] Lauric Feugere, Lauren Angell, James Fagents, Rebecca Nightingale, Kirsty Rowland, Saffiyah Skinner, Jőrg Hardege, Helga Bartels-Hardege, and Katharina C. Wollenberg Valero. Behavioural stress propagation in benthic invertebrates caused by acute pH drop-induced metabolites. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Caronni:2021:FRG

[3123] Sarah Caronni, Maria Anna Delaria, Rodolfo Gentili, Chiara Montagnani, Augusto Navone, Pieraugusto Panzalis, and Sandra Citterio. First report of gametogenesis and spawning for the invasive alga Caulerpa cylindracea in the Tyrrhenian Sea: The key role of water motion and temperature. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Chao:2021:DTS

[3124] Qiqi Chao, Zhifeng Gu, Aimin Wang, Chunsheng Liu, and Yi Yang. Diversity of three small type’s giant clams and their associated endosymbiotic symbiodiniaceae at Hainan and Xisha Islands, South China Sea. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Cui:2021:AGD

[3125] Xinhai Cui, Xiangfeng Zhao, Jintong Li, Ziwei Li, Xia Ren, Yangang Zhao, Xianjun Fu, and Xiuxue Li. Advances in the global distribution, chemical constituents, and pharmacology of Hippocampus, a traditional marine Chinese medicine. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Jannot:2021:UBM

[3126] Jason E. Jannot, Eric J. Ward, Kayleigh A. Somers, Blake E. Feist, Thomas P. Good, Dan Lawson, and James V. Carretta. Using Bayesian models to estimate humpback whale entanglements in the United States West Coast sablefish pot fishery. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Tang:2021:RRB

[3127] Meizhen Tang, Zhijie Lu, Sarath V. Babu, Guang Yang, Yanan Li, Zizheng Xu, Gan Pan, Zhendong Qin, and Li Lin. The regulatory relationships between the gonad-inhibiting hormone and insulin-like androgenic gland hormone-binding protein genes in the eyestalk-androgenic gland-testis axis of Macrobrachium rosenbergii. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Runde:2021:EED

[3128] Brendan J. Runde, Jeffrey A. Buckel, Paul J. Rudershausen, Warren A. Mitchell, Erik Ebert, Jie Cao, and J. Christopher Taylor. Evaluating the effects of a deep-water Marine Protected Area a decade after closure: a multifaceted approach reveals equivocal benefits to reef fish populations. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Wang:2021:DLM

[3129] Haining Wang, Xiaoxue Fu, Chengqian Zhao, Zhendong Luan, and Chaolun Li. A deep learning model to recognize and quantitatively analyze cold seep substrates and the dominant associated species. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Qin:2021:CGP

[3130] Yawen Qin, Chaoqun He, Wenqiang Wang, Peng Yang, Jun Wang, Qinbo Qin, Kangsen Mai, and Fei Song. Changes in growth performance, nutrient metabolism, antioxidant defense and immune response after fishmeal was replaced by low-gossypol cottonseed meal in golden pompano (Trachinotus ovatus). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Mancuso:2021:EMM

[3131] Monique Mancuso, Julian Blasco, and Nunziacarla Spanò. Editorial: Microplastics in the Mediterranean Sea. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Costanza:2021:PRA

[3132] Anna B. Costanza, Chiara Guidino, Jeffrey C. Mangel, Joanna Alfaro-Shigueto, Gregory Verutes, Marjolaine Caillat, Aritree Samanta, and Ellen Hines. Participatory risk assessment of humpback whale (Megaptera novaeangliae) and leatherback turtle (Dermochelys coriacea) bycatch in Northern Peru. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Xu:2021:FIV

[3133] Wen Xu, Yeqiang Shu, Dongxiao Wang, Ju Chen, Jinghong Wang, Qiang Xie, Qiang Wang, Danian Liu, Tingting Zu, and Yunkai He. Features of intraseasonal variability observed in the upper-layer current in the Northern South China Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Guden:2021:MFA

[3134] Rodgee Mae Guden, Sofie Derycke, and Tom Moens. A multi-faceted approach to understand how resource diversity can mediate the coexistence of cryptic marine nematode species. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Erdle:2021:WMF

[3135] Lisa M. Erdle, Dorsa Nouri Parto, David Sweetnam, and Chelsea M. Rochman. Washing machine filters reduce microfiber emissions: Evidence from a community-scale pilot in Parry Sound, Ontario. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Coelho:2021:CMM

[3136] Pedro Coelho, Frederico Carvalho, Thomas Goulding, Paula Chainho, and José Guerreiro. Corrigendum: Management models of the Manila clam (Ruditapes philippinarum) fisheries in invaded European coastal systems. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745. See [2359].

Ollevier:2021:NSD

[3137] Anouk Ollevier, Jonas Mortelmans, Anaïs Aubert, Klaas Deneudt, and Michiel B. Vandegehuchte. Noctiluca scintillans: Dynamics, size measurements and relationships with small soft-bodied plankton in the Belgian part of the North Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

McClanahan:2021:EVT

[3138] Tim Rice McClanahan and Maxwell Kodia Azali. Environmental variability and threshold model’s predictions for coral reefs. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Part:2021:OSD

[3139] Siim Pärt, Harri Kankaanpää, Jan-Victor Björkqvist, and Rivo Uiboupin. Oil spill detection using fluorometric sensors: Laboratory validation and implementation to a FerryBox and a moored SmartBuoy. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lintner:2021:APO

[3140] Michael Lintner, Bianca Lintner, Wolfgang Wanek, Sarina Schmidt, Nina Keul, and Petra Heinz. Assimilation of particular organic matter and dissolved organic or inorganic compounds by Cribroelphidium selseyense (Foraminifera). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Xu:2021:EMM

[3141] Bingying Xu, Yuechong Liu, Kai Chen, Lei Wang, Gladstone Sagada, Arnaud Fabrice Tegomo, Yifei Yang, Yuxiao Sun, Lu Zheng, Sami Ullah, and Qingjun Shao. Evaluation of methanotroph (Methylococcus capsulatus, Bath) bacteria meal (FeedKind(R)) as an alternative protein source for juvenile Black Sea bream, Acanthopagrus schlegelii. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

laMora:2021:IBS

[3142] Magdalena Précoma de la Mora, Nathan J. Bennett, Stuart Fulton, Adrian Munguia-Vega, Cristina Lasch-Thaler, Mariana Walther-Mendoza, José Alberto Zepeda-Domínguez, Elena Marie Finkbeiner, Alison L. Green, Alvin Suárez, Amy Hudson Weaver, Ana Luisa R. Figueroa Carranza, Armando Vega Velázquez, Calina Zepeda, Celerino Montes, David Antonio Fuentes Montalvo, Fiorenza Micheli, Héctor Reyes-Bonilla, Iliana Chollett, Ines Lopez-Ercilla, Juan Francisco Torres Origel, Leonardo Vázquez-Vera, María del Carmen García-Rivas, María del Mar Mancha-Cisneros, María José Espinosa-Romero, Mirian Martín Ruíz, Nur Arafeh-Dalmau, Ollin T. González-Cuellar, Oswaldo Huchim, and Salvador Rodríguez Van Dyck. Integrating biophysical, socio-economic and governance principles into marine reserve design and management in Mexico: From theory to practice. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Blazewicz:2021:PSC

[3143] Magdalena Błażewicz, Aleksandra Jakiel, and Graham J. Bird. Pseudotanais Sars, 1882 (Crustacea: Tanaidacea) from the SE Australian Slope: a gap in our knowledge. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Basran:2021:UCS

[3144] Charla Jean Basran and Gudjón Már Sigurdsson. Using case studies to investigate cetacean bycatch/interaction under-reporting in countries with reporting legislation. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Lee:2021:TVA

[3145] Dae Hyeok Lee, Jee Woong Choi, Sungwon Shin, and H. C. Song. Temporal variability in acoustic behavior of snapping shrimp in the East China Sea and its correlation with ocean environments. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Long:2021:EOF

[3146] Shuisheng Long, Zhihao Li, Xiaohui Dong, Xiaobo Yan, Hao Liu, Beiping Tan, Shuang Zhang, Simiao Pan, Tao Li, Xiangxiang Suo, and Yuanzhi Yang. The effect of oxidized fish oil on the spleen index, antioxidant activity, histology and transcriptome in juvenile hybrid grouper ([Venus sign] Epinephelus fuscoguttatus × [male sign] Epinephelus lanceolatus). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Claustre:2021:OBC

[3147] Hervé Claustre, Louis Legendre, Philip W. Boyd, and Marina Levy. The oceans’ biological carbon pumps: Framework for a research observational community approach. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Yang:2021:CES

[3148] Yang Yang, Tong Wang, Sen Yang, Xi Wu, Wenhua Huang, Bin Fan, Junhong Xia, Zining Meng, and Xiaochun Liu. Corrigendum: The effects of storage in vitro on functions, transcriptome, proteome, and oxidation resistance of giant grouper sperm. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745. See [2680].

Valdivia:2021:BDA

[3149] Nelson Valdivia, José Garcés-Vargas, Ignacio Garrido, Iván Gómez, Pirjo Huovinen, Nelso P. Navarro, Erasmo C. Macaya, and Luis Miguel Pardo. Beta diversity of Antarctic and sub-Antarctic benthic communities reveals a major role of stochastic assembly processes. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

deVerneil:2021:SOD

[3150] Alain de Verneil, John A. Burt, Matthew Mitchell, and Francesco Paparella. Summer oxygen dynamics on a Southern Arabian Gulf coral reef. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Martineac:2021:ACS

[3151] Rachel P. Martineac, Alexey V. Vorobev, Mary Ann Moran, and Patricia M. Medeiros. Assessing the contribution of seasonality, tides, and microbial processing to dissolved organic matter composition variability in a Southeastern U.S. Estuary. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hossain:2021:ISE

[3152] Farzana Hossain, S. M. Majharul Islam, Mohammad Ashaf-Ud-Doulah, Md Shaha Ali, Md Sadiqul Islam, Christopher Brown, and Md Shahjahan. Influences of salinity on embryonic and larval development of striped catfish Pangasianodon hypophthalmus. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Kraft:2021:SSG

[3153] Sebastián Kraft, MJosé Pérez-Alvarez, Carlos Olavarría, Rodrigo Moraga, C. Scott Baker, Debbie Steel, Paul Tixier, Christophe Guinet, Amelia Viricel, Paul Brickle, Marina Costa, Enrique Crespo, Cristian Durante, Rocio Loizaga, and Elie Poulin. From settlers to subspecies: Genetic differentiation in Commerson’s dolphins between South America and the Kerguelen Islands. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Dineshram:2021:OAT

[3154] Ramadoss Dineshram, Shu Xiao, Ginger Wai Kuen Ko, Jun Li, Kovilingal Smrithi, Vengatesen Thiyagarajan, Yang Zhang, and Ziniu Yu. Ocean acidification triggers cell signaling, suppress immune and calcification in the Pacific oyster larvae. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Walters:2021:ETM

[3155] David C. Walters, Joel A. Carr, Alyssa Hockaday, Joshua A. Jones, Eliza McFarland, Katya E. Kovalenko, Matthew L. Kirwan, Donald R. Cahoon, and Glenn R. Guntenspergen. Experimental tree mortality does not induce marsh transgression in a Chesapeake Bay low-lying coastal forest. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Genitsaris:2021:EMC

[3156] Savvas Genitsaris, Konstantinos Ar. Kormas, and Maria Moustaka-Gouni. Editorial: Microbial communities of coastal eutrophic systems. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745.

Han:2021:MCP

[3157] Mingming Han, Tianheng Gao, Yuxin Liu, Zakaria Zuraini, Chenxi Zhu, Tongqing Zhang, Feng Ji, Qichen Jiang, and Xian Sun. Molecular characterization of PRDX family and response of antioxidant enzymes in berberine hydrochloride-treated Charybdis japonica infected with Aeromonas hydrophila. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhang:2021:PTA

[3158] Jun Zhang, Shixia Liu, Chaoyang Hu, Xiaojiao Chen, Xue Sun, and Nianjun Xu. Physiological and transcriptome analysis of exogenous Larginine in the alleviation of high-temperature stress in Gracilariopsis lemaneiformis. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Pang:2021:UTL

[3159] Caryn Z. Pang, Yuen K. Ip, and Shit F. Chew. Using transcript levels of nitrate transporter 2 as molecular indicators to estimate the potentials of nitrate transport in Symbiodinium, Cladocopium, and Durusdinium of the fluted giant clam, Tridacna squamosa. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:LTA

[3160] Tuo Wang, Dongneng Jiang, Hongjuan Shi, Umar Farouk Mustapha, Siping Deng, Zhilong Liu, Wanxin Li, Huapu Chen, Chunhua Zhu, and Guangli Li. Liver transcriptomic analysis of the effects of dietary fish oil revealed a regulated expression pattern of genes in adult female spotted scat (Scatophagus argus). Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Fernandez-Gonzalez:2021:CPD

[3161] Victoria Fernandez-Gonzalez, Sandra Navarro-Mayoral, and Pablo Sanchez-Jerez. Connectivity patterns for direct developing invertebrates in fragmented marine habitats: Fish farms fouling as source population in the establishment and maintenance of local metapopulations. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Reavis:2021:RST

[3162] Janie L. Reavis, H. Seckin Demir, Blair E. Witherington, Michael J. Bresette, Jennifer Blain Christen, Jesse F. Senko, and Sule Ozev. Revealing sea turtle behavior in relation to fishing gear using color-coded spatiotemporal motion patterns with deep neural networks. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Almar:2021:ECU

[3163] Rafael Almar, Marta Marcos, Gonéri Le Cozannet, and Roshanka Ranasinghe. Editorial: Coasts under changing climate: Observations and modeling. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Richardson:2021:DSP

[3164] Katherine Richardson and Jørgen Bendtsen. Distinct seasonal primary production patterns in the Sub-Polar Gyre and surrounding seas. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

vonDassow:2021:DDL

[3165] Peter von Dassow, Paula Valentina Muñoz Farías, Sarah Pinon, Esther Velasco-Senovilla, and Simon Anguita-Salinas. Do differences in latitudinal distributions of species and organelle haplotypes reflect thermal reaction norms within the Emiliania/Gephyrocapsa complex? Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Yang:2021:CGT

[3166] Yanan Yang, Yuanfeng Xu, Peng Zhang, Zhaoxia Cui, and Chenchang Bao. Comparative genomic and transcriptomic analyses of CHHs and their putative receptors in Scylla paramamosain, Portunus trituberculatus, and Eriocheir sinensis. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Pauly:2021:ESM

[3167] Daniel Pauly, Cui Liang, and Weiwei Xian. Editorial: The status of marine fisheries in East Asia. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Khangaonkar:2021:PNP

[3168] Tarang Khangaonkar, Adi Nugraha, Su Kyong Yun, Lakshitha Premathilake, Julie E. Keister, and Julia Bos. Propagation of the 2014–2016 Northeast Pacific Marine Heatwave through the Salish Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Seger:2021:EBA

[3169] Kerri D. Seger, Renata Sousa-Lima, Juan J. Schmitter-Soto, and Edward R. Urban. Editorial: Before-after control-impact (BACI) studies in the ocean. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lacy:2021:AVS

[3170] Robert C. Lacy, Randall S. Wells, Michael D. Scott, Jason B. Allen, Aaron A. Barleycorn, Kim W. Urian, and Suzanne Hofmann. Assessing the viability of the Sarasota Bay community of bottlenose dolphins. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Droste:2021:SIS

[3171] Elise S. Droste, Alex R. Baker, Chan Yodle, Andrew Smith, and Laurens Ganzeveld. Soluble iodine speciation in marine aerosols across the Indian and Pacific Ocean Basins. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Liu:2021:CAC

[3172] Kuiyan Liu, Yang Chen, Zongmei Cui, Shuya Liu, Qing Xu, and Nansheng Chen. Comparative analysis of Chloroplast genomes of Thalassiosira species. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Splinter:2021:COC

[3173] Kristen D. Splinter and Giovanni Coco. Challenges and opportunities in coastal shoreline prediction. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:EFC

[3174] Lin Wang, Shuhong Wang, Chaoshu Zeng, Yuyue Wang, and Chunxiao Zeng. Effects of food concentration and photoperiod on egg production, female life expectancy and population dynamics of the paracalanid copepod, bestiolina amoyensis. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Li:2021:RSA

[3175] Zhujun Li, Tianxiang Gao, and Zhiqiang Han. RNA-seq and analysis of Argyrosomus japonicus under different salinities. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Zaferani:2021:CMA

[3176] Sara Zaferani and Harald Biester. Corrigendum: Mercury accumulation in marine sediments — a comparison of an upwelling area and two large river mouths. Frontiers in Marine Science, 8, October 2021. ISSN 2296-7745. See [2811].

Kim:2021:ISS

[3177] Dongyoung Kim, Rubao Ji, Hyun Je Park, Zhixuan Feng, Jaebin Jang, Chung l Lee, Yun-Ho Kang, and Chang-Keun Kang. Impact of shifting subpolar front on phytoplankton dynamics in the western margin of East/Japan Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Liu:2021:KSL

[3178] Guangpeng Liu, Annalisa Bracco, Andrea M. Quattrini, and Santiago Herrera. Kilometer-scale larval dispersal processes predict metapopulation connectivity pathways for Paramuricea biscaya in the Northern Gulf of Mexico. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Galaska:2021:SGR

[3179] Matthew P. Galaska, Guangpeng Liu, Destiny West, Katie Erickson, Andrea M. Quattrini, Annalisa Bracco, and Santiago Herrera. Seascape genomics reveals metapopulation connectivity network of Paramuricea biscaya in the Northern Gulf of Mexico. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Anderson:2021:SEO

[3180] Christopher M. Anderson, Amber Himes-Cornell, Cristina Pita, Ashleigh Arton, Margaux Favret, Dan Averill, Stephen Stohs, and Catherine Sarah Longo. Social and economic outcomes of fisheries certification: Characterizing pathways of change in canned fish markets. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Dmitrenko:2021:CPD

[3181] Igor A. Dmitrenko, Vladislav Y. Petrusevich, Ksenia Kosobokova, Alexander S. Komarov, Caroline Bouchard, Maxime Geoffroy, Nikolay V. Koldunov, David G. Babb, Sergei A. Kirillov, and David G. Barber. Coastal polynya disrupts the acoustic backscatter diurnal signal over the Eastern Laptev Sea Shelf. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Soares:2021:EMG

[3182] Douglas M. M. Soares, Samir V. F. Atum, Etelvino J. H. Bechara, João C. Setubal, Cassius V. Stevani, and Renato S. Freire. Exploring the microbiota of the Guarapiranga Water Reservoir with long-read sequencing technology. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Liu:2021:RCF

[3183] Fang Liu, Wenyuan Shi, Lin Huang, Guizhong Wang, Zhihuang Zhu, and Haihui Ye. Roles of crustacean female sex hormone 1a in a protandric simultaneous hermaphrodite shrimp. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:IBQ

[3184] Xin Wang, Dan Wang, Tianyu Zhang, Qianqian Zhang, and Jianmin Zhao. iTRAQ-based quantitative proteomic analysis reveals toxicity mechanisms in Chlamys farreri exposed to okadaic acid. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Linse:2021:EEI

[3185] Katrin Linse, Ilka Peeken, and Anne Helene Solberg Tandberg. Editorial: Effects of ice loss on marine biodiversity. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Ponti:2021:EBR

[3186] Massimo Ponti, Cristina Linares, Carlo Cerrano, Riccardo Rodolfo-Metalpa, and Bert W. Hoeksema. Editorial: Biogenic reefs at risk: Facing globally widespread local threats and their interaction with climate change. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Sun:2021:CRG

[3187] Gaigai Sun, Changsen Sun, Jing He, Hanhan Yao, Wenfang Dai, Zhihua Lin, and Yinghui Dong. Characterizing the role of glutamine synthetase gene on ammonia nitrogen detoxification metabolism of the razor clam Sinonovacula constricta. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hong:2021:EEH

[3188] Hyun-Ki Hong, Chang Wan Kim, Jeong-Hwa Kim, Nobuhisa Kajino, and Kwang-Sik Choi. Effect of extreme heatwaves on the mortality and cellular immune responses of purplish bifurcate mussel Mytilisepta virgata (Wiegmann, 1837) (= Septifer virgatus) in indoor mesocosm experiments. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Huang:2021:TLS

[3189] Songqian Huang, Li Zhang, and Xiaojuan Cao. Transcriptional landscape of small non-coding RNAs in somatic and gonadal tissues of brachyuran crabs (genus Eriocheir and Portunus). Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

vanderLinde:2021:LSH

[3190] Krista van der Linde, Ingrid N. Visser, Rick Bout, Chris Lalas, Lara Shepherd, David Hocking, Brittany Finucci, Jim Fyfe, and Matthew Pinkerton. Leopard seals (Hydrurga leptonyx) in New Zealand waters predating on chondrichthyans. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Berenji:2021:EVB

[3191] Sevil Berenji, Prateep Kumar Nayak, and Ankur Shukla. Exploring values and beliefs in a complex coastal social-ecological system: a case of small-scale fishery and dried fish production in Sagar Island, Indian Sundarbans. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Hu:2021:PLP

[3192] Tingcang Hu, Min Luo, Yunping Xu, Shanggui Gong, and Duofu Chen. Production of labile protein-like dissolved organic carbon associated with anaerobic methane oxidization in the Haima Cold Seeps, South China Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhang:2021:PMR

[3193] Jingyan Zhang, Mengqian Zhang, Nishad Jayasundara, Xianyun Ren, Baoquan Gao, Ping Liu, Jian Li, and Xianliang Meng. Physiological and molecular responses in the gill of the swimming crab Portunus trituberculatus during long-term ammonia stress. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Smith:2021:EBC

[3194] Craig R. Smith, Malcolm R. Clark, Erica Goetze, Adrian G. Glover, and Kerry L. Howell. Editorial: Biodiversity, connectivity and ecosystem function across the clarion-clipperton zone: a regional synthesis for an area targeted for nodule mining. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Mikami:2021:EES

[3195] Koji Mikami, Susumu Takio, Yuji Hiwatashi, and Manoj Kumar. Editorial: Environmental stress-promoting responses in algae. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Lin:2021:SMI

[3196] Fan Lin, Lars Asplin, and Hao Wei. Summertime M2 internal tides in the Northern Yellow Sea. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Cowan:2021:WSW

[3197] Emily Cowan and Rachel Tiller. What shall we do with a sea of plastics? A systematic literature review on how to pave the road toward a global comprehensive plastic governance agreement. Frontiers in Marine Science, 8, November 2021. ISSN 2296-7745.

Scrosati:2021:RPR

[3198] Ricardo A. Scrosati and Janelle K. Holt. Recruitment and postrecruitment dynamics of the barnacle Semibalanus balanoides on a waveexposed headland in Atlantic Canada. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Dahdouh-Guebas:2021:MRU

[3199] Farid Dahdouh-Guebas and Stefano Cannicci. Mangrove restoration under shifted baselines and future uncertainty. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhang:2021:EOC

[3200] Jiajia Zhang, Guoyuan Zou, Xuexia Wang, Wencheng Ding, Li Xu, Baoyin Liu, Yunsen Mu, Xuran Zhu, Lianjie Song, and Yanhua Chen. Exploring the occurrence characteristics of microplastics in typical maize farmland soils with long-term plastic film mulching in Northern China. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Kawamura:2021:CVS

[3201] Kaz Kawamura, Satoko Sekida, Koki Nishitsuji, Eiichi Shoguchi, Kanako Hisata, Shigeki Fujiwara, and Noriyuki Satoh. Corrigendum: In vitro symbiosis of reef-building coral cells with photosynthetic dinoflagellates. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745. See [2587].

Fu:2021:IRD

[3202] Ruiying Fu, Xuena Huang, Yiyong Chen, Zaohuang Chen, and Aibin Zhan. Interactive regulations of dynamic methylation and transcriptional responses to recurring environmental stresses during biological invasions. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Bue:2021:EID

[3203] Nadia Lo Bue, Vincenzo Artale, and Katrin Schroeder. Editorial: Impact of deep oceanic processes on circulation and climate variability: Examples from the Mediterranean Sea and the global ocean. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Zhang:2021:RMC

[3204] Hongxia Zhang, Mingshan Song, Lili Wang, Anguo Zhang, Xiaolong Yang, Yongjian Liu, and Xiutang Yuan. Response of microbial communities on culturing plates of post-settlement sea cucumbers to seawater acidification and warming. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Fang:2021:RAL

[3205] HaoHang Fang, XuanShu He, HuLong Zeng, YongJian Liu, LiXia Tian, and Jin Niu. Replacement of astaxanthin with lutein in diets of juvenile Litopenaeus vannamei: Effects on growth performance, antioxidant capacity, and immune response. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Tamburri:2021:TCD

[3206] Mario N. Tamburri, Eugene T. Georgiades, Christopher Scianni, Matthew R. First, Gregory M. Ruiz, and Carolyn E. Junemann. Technical considerations for development of policy and approvals for in-water cleaning of ship biofouling. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Wang:2021:SIA

[3207] Yibang Wang, Cui Liang, Zhaomin Chen, Shude Liu, Hui Zhang, and Weiwei Xian. Spring ichthyoplankton assemblage structure in the Yangtze Estuary under environmental factors. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Osadchiev:2021:RSR

[3208] Alexander Osadchiev, Roman Sedakov, and Alexandra Barymova. Response of a small river plume on wind forcing. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Blanco-Libreros:2021:EEE

[3209] Juan F. Blanco-Libreros, José M. Riascos, Benjamin Branoff, and Eliza C. Heery. Editorial: Eco-evolutionary dynamics of marine biodiversity in human-dominated coastalscapes. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

Emerenciano:2021:BTB

[3210] Maurício G. C. Emerenciano, Anselmo Miranda-Baeza, Marcel Martínez-Porchas, Moisés A. Poli, and Felipe do Nascimento Vieira. Biofloc technology (BFT) in shrimp farming: Past and present shaping the future. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.

VanDam:2021:COP

[3211] Bryce Van Dam, Christian Lopes, Mary A. Zeller, Mariana Ribas-Ribas, Hongjie Wang, and Helmuth Thomas. Corrigendum: Overstated potential for seagrass meadows to mitigate coastal ocean acidification. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745. See [2788].

Wen:2021:HRP

[3212] Zheng-Yong Wen, Chuan-Jie Qin, Yun-Yun Lv, Yan-Ping Li, Yuan-Chao Zou, Sheng-Tao Guo, and Qiong Shi. Homeostasis regulation by potassium channel subfamily K member 3 (KCNK3) in various fishes. Frontiers in Marine Science, 8, December 2021. ISSN 2296-7745.